A Complete Bibliography of Publications in IRE Transactions on Engineering Writing and Speech, IEEE Transactions on Engineering Writing and Speech, and IEEE Transactions on Professional

Communication

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

24 August 2024 Version 1.07

Title word cross-reference

#5 [HM95]. #NoDAPL [SvI18].

$15.00 [Die72]. $6.95 [Mek72]. $64 [Sau93]. ∘ [WRTBRL23].

-A [Daw10]. -er} [Ano74i]. -re} [Ano74i].

1 [Ano61n, ZFWX21]. 10th [Mac02b]. 14 [Ano71a]. 16 [Ano73a]. 1640 [Ram98]. 17 [Ano74a]. 1700 [Fer06]. 17th [Wha04]. 18 [Ano75a]. 1838 [Pod03]. 1879 [Fil05]. 19 [Ano76a, GLCA22, RWAK24]. 1925 [Cun10]. 1950 [Zim01]. 1960 [Ano60a]. 1963 [Ano64p]. 1965 [Ano64r]. 1968 [Ano67a]. 1975 [Ano75m, Dob75]. 1984 [Wor87]. 1990s [BFA05, Lin93]. 1996 [Ano97a, Ano97b].

2 [Ano64j, Her64]. 2.0 [Ric09, WOD^{+}16]. 2005 [Ano04j]. 2009 [Ano09u]. 2022 [Wal23]. 21 [Ano78a]. 21st [Pet14, Wil03, St.04]. 22 [Ano79a]. 23 [Ano80a]. 24 [Ano81b]. 25 [Ano82c]. 26 [Ano83b]. 27 [Ano84a]. 28 [Ano85a]. 29 [Ano86b]. 2nd [Den83b, Flo02d, Flo06d, Hol03, Hou21b, Hue02, Kae03, Kno80, Kry00a, LN98, Ril01, Wei01, Wha01, You10, Zim01, Mar08].

3 [Ano61n]. 3/E [Ber12]. 30 [Ano87a]. 360 [Fre68, Zie69]. 3rd [Bra05, Fre80a, LeC01, LK08, Nor97, Yu24].

4th [Flo02c, Har80a, Jen99b, Pot01a].

500 [JD11, Jon15]. 5th [Dou01, Flo06a, Mac02a].

6 [Ano64j, Her64]. 6-2 [Ano64j, Her64]. 600-year-old [HN93]. 61 [Ano60k]. 6th [Flo02a, Flo07c, Joh02].

’70 [Mil65]. ’72 [Phi72a, Bas72, Bli72, De 72, Dou72, Gar72, Gre72, Ham72, Lin72, LK72, Mac72, McG72, Men72, Pri72, Rob72, Sar72, Shr72, Sum72, Ter72]. ’73 [Ano73m, Phi72b]. ’74 [Ano73m]. 7th [Flo00b].

84 [Ano83-31, Ano84-32, Ano84-33]. 8th [Flo03d, LM02].

9-11 [Cha06]. 9/11 [Str04]. 9th [Flo05a].

= [AM07, BH16].

A&M [Gat04]. A11y [SMP21]. Aaaalll [SSV^{+}23]. AAPT [Got79]. Aaron [Rok19]. ABACUS [Tel68]. ABACUS-AB [Tel68]. Abbreviations [Loe68]. ability [Ano83l, Joe83a]. Ableism [Ben22]. abound [Ort78]. abroad [Kim79, Ter94, Win83]. Absence [JW22]. Absorptive [RD16]. Abstract [Luf71, Sek73]. Abstracting [Ort72, App01b]. abstraction [TCW06]. Abstracts [Ano58a, Ano61a, Ano09s, Ano10m, Ano10n, Ano10s, Ano11f, Arn61, Bro71, Lac62, Mic62, SSM74, Ano77c]. abuse [MG82]. abused [Ano81h]. Academic [Bat18a, CFMM14, Dou00, Hov08, Ley08, Smi20, And84, KW93, Kus97, Meh94, PT05, Sch93]. academic-practitioner [PT05]. Academics [LD20]. Academy [GB16, Boo17]. Accelerator [CSSV20, PH24]. acceptable [Boy81]. Acceptance [ESM98, JCJ^{+}75]. Access [Ano12-28, Ano14-27, Ano14y, Ano14z, Ano15-31, Ano15-32, BH21, Ben22, Die04, SvI18, Urb77, Ano12q, Fie77, Pal75, Pal77, Wei77, St.09]. Accessibility [Hun21, MR21, SMP21, Wil24, CM01]. Accessibility-Centered [SMP21]. Accessible [Str21, Wil24, Rab92]. Accident [Fil05, Win88]. Accidental [Flo04b, Jen99a]. accommodating [TK04]. Accommodation [Hat69]. according [Ren96]. account [Ano87y]. accountability [Bur84a]. accounting [GRS96]. Accumulation [JE95]. Accuracy [Cor83, CI02, Dun82]. accurate [Ant99]. ACE [Cor83]. Achievement [AFH20]. Achieving [Man98a, Mic68, Zem12, Sch93]. Acing [Bar09a]. Acknowledgment [Ano71b]. Acquiring [Ste04]. Acquisition [BG16, Men72]. Acronyms [Ibr89]. Across [CLP08, Lul06, SSW17, TM08, ZB17, For04, Gil03, JDJ99, Ken88, Ken89, OW91, PT05]. Act [Ama04a, Luf69a, Luf71, MPD10, Ock07, Wor87, Ano80i, Luf68]. Acting [Dou00]. Action [Dre16, Koc03a, WLM^{+}22, Koc01, Koc03b, Sch81d, And97, GGS23, Sta20]. Activating [BN07, QK05]. Active [Lie18, MMCL00, App02]. Activities [FPKS^{+}21, Sch66, SU09, Wil64, Ano63r, Fry62, Moo87a]. Activity [GMT08, Nyc97, Whi06]. Acts [YK13, San87a]. Ad [vMPKG15, LMJK14]. AD-HOC [LMJK14]. Adams [Whi78a]. Adapt [FW23]. Adaptation [KCM15, XS24, AC06, Koc01, Kos95].

Adapted [Wat20]. Adapting [Bar14, McC06, SY09]. Adaptive [KT14]. Add [Plu84, Her89]. added [PM98, RR93]. Adding [Sub92]. address [Buc74, Dar82, Den96]. Addressing [Dat91, Den98b, Hun21, RK91]. Adjectives [Sam06]. administration [Cov75]. Administrative [Fre68]. Adopters [Cog15]. Adoption [AZ17, AZDA10, Mun05, Gro95a]. Adorable [Rok19]. adult [Cra85a]. Advances [Red98]. Advancing [GZ97, Ano14-32, Ano14-33, Ano14-34]. advantage [BB99, HR99]. advantages [BS98]. Advertisement [Ano59a, Ano64b, Ano64c, Ano64d, Ano65a, Ano66a, Ano69a, Ano04r, Ano04k, Ano04l, Ano06j, Ano06k, Ano07l, Ano07m, Ano08k, Ano08t, Ano09m, Ano09n, Ano09t, Ano11d, Ano12o, Ano12p, Ano12q, Ano12x, Ano12y, Ano12z, Ano12-27, Ano12-28, Ano14m]. Advertisements [Ano73b, Ano73c, Ano73d, Ano73e, Ano74b, Ano78b, Ano78c, vMPKG15]. Advertising [OL98, Por87]. advice [Ano78p, FP79, FS92, MRE96]. Advisement [SSW17]. Advisory [SSW17]. Advocacy [Dur22, KW23]. Advocating [Ben22]. Aerospace [Ano63k, EBAS18, BPK94, PBK^{+}93]. Affect [KCLG13, MFL17, vSdJ07, Nor00]. Affecting [GAGW15, Gan04, PFM88]. Affective [HK07]. Affects [CWSL19, Ken82b]. Affiliate [Ano67b]. afford [Ano82f, Ano83m, Ano83n, Ano84m, Har75a]. Affordances [Tha24, Die04]. Africa [ACW19, WD09]. After [CC19, Sta20]. After-Hours [CC19]. Again [Ano67e]. Agboka [Gan22]. Age [Flo05d, Gre17, Lan05, Moo05, Wil89, Cre95, Flo03c, Gri84, Li80, Mat91, St.02c, Ram07, Ris98, Spa14, St.01c]. Agencies [Boi18, Sus61]. Agency [SY09, Lid75]. agenda [St.01a]. Agent [LHAL23].

Agent-Based [LHAL23]. Agents [Cog15]. aggravated [Sum72]. Agile [AG21]. aging [Lip04]. ahead [Bur84b, Phi72a, Phi72b]. AIAA [DBM73]. Aid [Bun71, Pet69, Roy07, DH94, Gup83a, Ivo80, Ken64a, Lip75, Web79]. Aided [Lip71, Mor67b, Mor67a, Mus78b, Kai68]. Aids [Tur79, Aug78, Gou73]. aim [Sav88]. aim-write [Sav88]. air [Cor90, Bou58, ZS04]. airlines [Str04]. al [Fer07c, Fer07d, Flo07a, Flo07e, Flo08, Man18, Xu08]. al. [Ano64f]. Alan [Bra03a, Wha04]. Albers [Fil06]. Alcohol [Her95]. Alexander [Pri99]. Alfred [Ano97a, Ano77a, Ano78d, Ano79b, Ano80c, Ano81c, Ano83c, Ano84b, Ano85d, Ano86c, Ano87b, Ano96]. Algebra [Hou21a]. Algorithmic [Hou22b]. algorithms [Hau79, Hou21a]. Aligning [LY98, Roh19]. Alignment [BN10]. Allen [Bra99a, Eri11]. Allyn [And97, Ber12]. Aloud [Coo10, ELdJ12, May19, BR00, KU04, vW00]. Alouds [MEZ12]. Alphanumeric [Bec68]. Alred [Den83c, Den83b, Whi77c, Har80b]. also [Her95]. alternate [Sta73]. alternative [CR84, CE05, KSW78, Zim88]. Alternatives [Pal75, Aic78]. Amant [Boo17, Hin10, Waa09]. Amateur [WC15, Wil89]. ameliorating [Cro86]. America [Bro65, Pei81, Sau93]. American [Whi75, Zim01, Dar82, Dev75, Goo94, HZX24, MG95, Nye99, Par75, Sap08, SF02, SJYZ18, Yu20, Yu23, ZC24, Ano77c, Sim96a, You11]. Among [DLVE23, Bal75, DRG^{+}02, Hut75, MOA02, Rob72, ST99a]. Anaheim [Dob75]. Analyses [Boi18, Daw10, Hou21b]. Analysis [Ano63c, Ant06, Ant18, BP10, BP15, Car12, Car24, Cho21, Cli63, DL17, ELdJ12, FDJW18, FMB18, HS18, Jin18a, JW17, KDBC20, Kon22, Lam16, LBE23, LdJK20, MCC62, MPD10, Mou05, RWAK24, Smi20, Ste22, SWR21, SKZ23, Vic03, WKSH68, WLM^{+}22, YB17, Yu20, Yu23, ZC24, dGKNG06, vdHdJS06, Bat02, BW90, Bel99, BFA05, CSH84, Cra92, Den96, GBF05, Hac85, Hag90, Her95, HT77, HS72, HCKJ03, Luz05, Mac04b, Mar02, Nel90, Ock05, Str99, Sub92, Teb80, TCW06, Tha04, Tho01, WB92, Wan89, Wat90, Wil97, Smi10]. Analyst [dGKGN11]. analytical [AV98, SF99]. Analyze [PMGW21]. Analyzing [MPD10, Whi89, GBF05]. Anatomy [Gou70]. Andrea [Pet13]. Andrew [Den98a]. Andrews [The97]. Angels [Reh98]. Anges [Ril02a]. Anglicisms [Lav06]. Ann [Hou22b]. Annals [BB75]. Anne [Whi79e]. Announcement [Ano70a, Ano75b, Ano83-31]. Annual [Ano58b, Ano58i, KdJH14, Moo87a, dGKNG06, dGKGN11, Bro89, MG95, Pot83]. Answer [Bar09a, HK11]. answering [Dre88]. Answers [Ben61, St.06b, Boy81]. Antecedents [CC19]. anthology [PM84]. anti [FO06]. anti-union [FO06]. Anticipating [ROW19]. antiseptics [CC88]. Anxiety [FVB16]. Any [Ram05a]. Anytime [Per96]. anyway [Par86]. anywhere [Per96]. Apart [Dou00]. Apollo [Mor67a]. apparatus [Koc05]. Appeals [Fri13]. Appendix [MC75a]. Appliances [LdJK20]. applicability [KL06]. applicants [Car90b]. Application [BG16, Ven62, Ama03b, Ano07a, Ano13a, Ano13b, FC79, Hor89]. Applications [Izq24, Pet14, Flo02d, PFM88, Sha94b, Con99, Sto08]. Apply [Bli61, Ano78k]. Applying [Bis96, Lam18]. Appointment [Ano66b]. Appraisal [Hat69]. Apprehension [Lie18]. Approach [BIdJ^{+}21, Bur69, Coo10, Dre16, DGPR10, Fer04b, Fer04c, Flo04c, HLY22, JW17, KT18, Kor75, Lip71, Lof10, LHAL23, Ric09, Sap08, SSS^{+}13, AC85, Ama05, Cas86, Cha02, Deg80, DEBD83, Hub89, Kre88, Lee64, LL03, May82, Moo85, Mus78a, Per70, Ros83, Sko83, Suo13, Tha04, Tho73, TC95, Vel88, Wal99, ZZ05, Zim88, ZS04, vdM97b, Gar19, Nor97, Gra96a, Flo02c]. Approaches [BI18, Car12, Flo05b, Sou96, Tus05, CW90, DR94, Mar97]. approaching [Zur77]. appropriate [Hor91, Mac82, Mac78b]. appropriateness [Dri87]. appropriation [Bae12]. Arabic [Ibr85]. Archibald [Cha81]. architectural [Kva94]. Architecture [Mur00c, OM12]. architectures [Heb97]. Archive [Sto15]. Archives [JW22, Ano77-29]. ARCOL [Mak68]. ARCOL/System [Mak68]. Area [BN10, Nic59]. aren’t [Gro94]. argue [Del81]. Argumentative [PH24]. Arguments [McE18]. Arias [Mas08]. Arias-Oliva [Mas08]. Arieti [Sch80a]. Arigato [Cam07]. Arisoph [BR16]. Arlington [The97]. armadillo [Dob85]. armor [Dob85]. army [Ber92]. Arnold [Mon79]. Art [Ano63-27, GH09, HK70, Jon89, LN98, Man98c, Mar64, Pas60, Ano71d, Ano80j, BR93, Den98b, Eng89, Flo00a, Gre69, Hof87, Hub85, Mur00i, Tra92]. Arthur [Ano80i]. Article [Ano09s, Ano10m, Ano10n, Ano10s, Ano11f, Car59, GE13, Jin18a, Jin18b, Kan12, KL12a, KCLG13, SC12, WHC^{+}12, Wha04, Ant79, Ant99, Ant01, Bau62, FWJ77, Hag90, Hol76, Sta73, Urb77, You80]. Articles [Ano61a, Jin18b, JLS20, KT18, Rau21, MGB89, Mic62, Tho01]. artifact [Tha04]. artist [Vog83]. Artistic [Fry81]. Arts [Joh97a, Fie83, Nag87c, Spe60]. ASAPP [DEBD83]. Ask [Bar09a, Nag87b]. Asking [Sar72]. Aspects [Lin19, Mic61e, Red98, Reg70, Ano84-32, Ano84-33, DF87, Lie77, MG95, Sko79, Wei75]. aspiring [Eck03]. Assembling [GG22a, McE18]. assembly [Sau82]. Assessing [BAB10, Coo10, Dav10, Orr02, YTY12, Gui89a, Koc03b]. Assessment [Boe10, GE13, KW23, MV21, Orr10, Wil10a, Eri11]. Assessments [DSV10]. assigning [Gre87]. assignment [Hag90]. assignments [GG96]. assist [AL03]. Assistance [SV24]. Assisted [GY68, Kon22, LBE23, Hor84]. assisting [NKU03]. Assoc [Joh97b]. Associate [Web65, Ano66b]. Associated [Car20, Str64]. association [JE95, Dev75]. assume [Sou88]. Astrolabe [HN93]. asylum [Mur00f]. asymmetrical [Chi89]. Asynchronous [SJD20, Ock05]. Atlantic [MMVH10, VMM21]. atomic [Tel68]. Atrisco [HM95]. attempt [GG84]. Attention [Hat69, GvdM02, St.01b]. attitude [FG84, Sid87]. Attitudes [KNMM09, KdJH14, CJ89, FWJ77, Pry95]. Attributes [LC17]. Attribution [Smi10]. Audience [CSH84, Kil09, Par06, Per78, RB19, RW86, Tus05, Car89, Cos84, Cox80, Fry87, Gar80a, Hag90, Kan85, Ker80, Kil82, Lon95, Pot84, Sub92, Rav06]. audience-with [Car89]. Audiences [Bat14, Dre17, McC06, Lip04, Rei02, Sch78b]. Audio [Ano77-30, Ese69]. Audio/visual [Ano77-30]. audiovisual [Rua80]. Audit [Izq24]. Augmented [MMW^{+}22]. Augmenting [San91]. Aunt [Nag86]. aural [EA84]. Austerity [Gre17]. Austin [Ano80f]. Australia [AW17, Den96]. Authenticity [DC87]. Author [Ano60b, Ano60c, Ano61b, Ano99a, Ano01a, Ano02a, Ano03a, Hol77, Pad78, FHR87b, Gre87, Lay73, Mic90]. Author-supplied [Hol77]. Authoring [ESPQ15]. Authority [Dos06, Loe95]. Authors [Ano68r, Ano19g, Ano19h, Ano19i, Ano20k, Ano20l, Ano20m, Ano20n, Ano21k, Ano21l, Ano22l, Ano22m, Ano22b, Ano22k, Ano23b, Ano23c, Ano23d, Ano23e, Ano24a, Ano24b, Ano24c, CFMM14, Mur59, Ano61c, Ano61d, Ano61e, Ano61o, Ano61p, Ano61q, Ano62a, Ano62h, Ano63p, Ano63q, Ano64v, Ano64w, Ano64x, Ano65j, Ano66k, Ano66l, Ano67l, Ano68n, Ano68o, Ano68p, Ano69o, Ano69p, Ano69q, Ano70m, Ano70n, Ano71t, Ano71u, Ano71v, Ano77m, Ano77n, Ano77o, Ano77p, Ano78l, Ano78m, Ano78n, Ano79p, Ano79q, Ano79r, Ano80x, Ano80y, Ano80z, Ano80-27, Ano81n, Ano81o, Ano81p, Ano82l, Ano82m, Ano82n, Ano83w, Ano83x, Ano83y, Ano84y, Ano84z, Ano85t, Ano86u, Ano87u, Ano04b, Ano04c, Ano04d, Ano04e, Ano05b, Ano05c, Ano05d]. authors [Ano05e, Ano06b, Ano06c, Ano06d, Ano06e, Ano07c, Ano07d, Ano07e, Ano07f, Ano08b, Ano08c, Ano08d, Ano08e, Ano09b, Ano09c, Ano09d, Ano09e, Ano10b, Ano10c, Ano10d, Ano10e, Ano11a, Ano12f, Ano12g, Ano21i, Ano21j, Dio82, Gei75, Ter65, Zap80]. authors/typists [Zap80]. Authorship [Ada06, Lam14, D’A65]. Autoethnography [BG16]. Automated [Boy68, Bur69, DSV10, DEBD83, HE77, HC88, Kor77, Stu69]. Automatic [GH70, Lam68, Sal68, Die72]. Automation [EBAS18]. Autonomous [Hou22b]. Autoscript [Stu69]. AV [Wan89]. Ava [Flo01a]. available [Ano85y]. averation [Cal74]. Aviation [DSV10, MMW^{+}22]. avoid [Mic90, Wis91]. avoiding [SO98, Woo90]. Award [Ano96, Ano97a, Ano97b, Ano77a, Ano78d, Ano79b, Ano80c, Ano81c, Ano83c, Ano84b, Ano85d, Ano86c, Ano87b]. Aware [MP20]. Awareness [Ano68t, BMH05, Ley08, RB19, DH92, Lon95, LN03, Par06, Sub92]. ax [Wis91]. Ayako [Rok19].

B [Ano07k, Bar22, Ber12, DR10, Flo10, Kre86, Lof08, Waa09, Wis15, Yea09, MC75a]. B.S. [Cam96]. Babel [LH00]. back [Ano60j, Ano62i, Ano72k, Ano73r, Ano75n, Ano77s, Ano77t, Ano77u, Ano80-28, Ano83g, Ano83h, Ano83i, Ano84g, Ano84h, Ano84i, Ano84j, Ano85j, Ano85k, Ano85l, Ano86h, Ano86i, Ano86j, Ano86k, Ano87k, Ano87g, Ano87h, Ano87i, Ano87j, Vel88, Ano72a, Ano73f, Ano75c, Ano76b, Ano76c, Ano77b, Ano79c, Ano79d, Ano80d, Ano83d, Ano83e, Ano83f, Ano84c, Ano84d, Ano84e, Ano84f, Ano85e, Ano85f, Ano85g, Ano85h, Ano85i, Ano86d, Ano86e, Ano86f, Ano86g, Ano87c, Ano87d, Ano87e, Ano87f]. Background [Kes60, Whi77a, JVT01, Sha94a]. Bacon [And97, Ber12]. Bad [Bal18, KL12b, Con78, Mur00g, Teb80, Wis91]. Balaban [Sch79]. Balancing [Gil87a, Ock07, San87a]. Banks [SWR21, SKZ23]. Barbara [Luf73, Wil03]. Barnett [Pen98g, Sch65]. Barrass [Whi78c]. Barriers [AZ17, Bas72, LK72, RFH87]. bars [Wri79]. Barton [LD20]. base [Coy77, GKD77, PDK88]. Based [AZ17, BR16, BC17, BW19, CD98, Dui96, GMT08, Gra96a, HS18, Hen22, Hou21a, HZX24, KZCR10, KT20, KMR14, KCJC08, Lee13, LHAL23, MAK15, Pet14, Ric09, RRW^{+}17, ROW19, Sam06, SWR21, VG16, Vic03, YB17, ZWE21, dGKGN11, BP82, BK03, CM02, Del92, Evi04, Eva00, Fre03b, Hau79, JCF92, LL03, LN03, MSW97, MMCL00, NKU03, Par06, SWB^{+}05, WBW89, War88, DOHM96, Str03]. bases [Bur75, Pad78]. basic [Bri99, Cra83a, Cur80, Jon80, LC79, Tho75, Mic85b, Cha80a, Har80a, Whi79c]. Basics [Bra05, Hol03, Bos87a, Gro94]. Basis [Per63, SSV^{+}23]. Bates [Sch80b]. Baym [Mog09]. Be [Ano63k, BC70, Bro64, Pod84, Dom95, Dul92, FBE^{+}03, Ful80, Joh65, Lou58, Sch80b, Vog83]. bears [Luf73]. Beat [Hou22a]. beats [DC87]. Become [Cha59]. becomes [Gre72]. Becoming [Hay80, SY09]. bedfellows [BCB03]. Been [McN58a]. Beer [Bra99b]. Before [Flo07d, FC79, Hin73, Pak81]. begin [Woo81]. beginner [Ram89]. Beginning [SP87]. Beginnings [PP97]. Behavior [LE18, vdM07, Hol74, Koc05, UN00, Zho05]. Behavioral [AC06, LG14]. Behaviors [SC12, Kol93, RH92]. Behavioural [Ano63c]. Being [KWLS17]. believe [Jon80]. Bell [Cle03]. Belliston [Mus78a]. Bench [Goo05]. Benchmarking [SB95]. Benefits [HK11, Met75, KL06]. Benjamin [Bro65]. Berdanier [Bar22]. Berger [Lof08]. Bernadette [Rou19]. Bernstein [Mek72, Cha65]. Beshara [Bar09a]. Besse [NB04]. Best [HK11, ZJG17]. Best-Known [ZJG17]. Better [Ano80e, Ben61, BC70, BH16, Cho21, Far64, Gon20, Van59, vdGRRv13, Ano14h, AN85, BB86, Dav81a, Den88, Dul93, Eng72, Hin72, JP86, Kie82, Mac97, May82, PP97, Pot82, Rau81, Sch80d, Smi03, Whi78a]. Betty [Den83a]. Between [BJ21b, AG21, AFH20, Boi18, LWMR07, LdJK20, Loe91, LAWI24, Sap08, Wol06, Caw75, DRG^{+}02, Dia14, KVRR06, Lid75, Lin72, Lof10, Luf76, RP98, SU09]. Beware [SGM^{+}24, Mun02]. Beyond [Bra82c, JSZ07, KSNF15, Pat08, Sul89, Bro88a, Den00, Man98b, Mir99, Tuk62, GG22b, St.03]. Bhatt [LD20]. Biafore [DR10]. Bias [Car24, FR13, BK03, Den98b]. bias-reducing [BK03]. Bible [DR10]. bibliographic [Hil89, Pad78, Tel68]. Bibliographies [McC65, Smi10]. Bibliography [Ede61, HH93, Hil85a, Abs82, Bro88b, FS80, Gup73, Kil83, Pra87, Ram89, Sno64, WH82]. Bibliometric [CFMM14]. Big [CE05, DeT94]. Bill [Pod03]. Binder [Zim13]. Bingham [Str84]. Biographies [Ano58e, Ano59c, Ano59d, Ano59e, Ano60b, Ano60c]. biological [Koc05]. Biology [Hut75]. Birth [Dan77, Wit77, Mau79]. Black [AW22, Rob78, Wil10b, Bro75, JW22]. Blank [Ano15c, Ano16c, Ano17b, Ano18b, Ano18c, SM10, Ano83g, Ano83h, Ano83i, Ano84g, Ano84h, Ano84i, Ano84j, Ano85j, Ano85k, Ano85l, Ano86h, Ano86i, Ano86j, Ano86k, Ano87g, Ano87h, Ano87i, Ano87j, Ano12d, Ano12e, Ano13d, Ano13e, Ano14b, Ano14c, Ano15b, Ano16b, Ano17c]. Blank-Page [SM10]. Bleifeld [Rei79]. blended [Bae12, DH04]. Blind [BNB21]. Block [Bro71]. blockbusting [Whi78a]. blocks [Pot02]. Blogging [DV09]. Blogosphere [SFM09]. Blogs [BAB10, Fer07c, Lee13]. Blood [Ano67e]. Bloomstein [Hou22a]. blowing [HWW04].

Blueprint [Dor63]. Boards [SSW17]. Boiarsky [Rou17b]. boilerplate [Hay83]. Bois [Ano80h]. Bonaparte [St.02b]. Bond [Vol82]. Bonfire [Gat04]. Bonni [Bai24]. Bonnie [Nyc97]. Book [Ale20, All97, Ana21, And97, Ano62b, Ano63d, Ano63-27, Ano64e, Ano64f, Ano84k, Ano97d, Bai24, Bal18, Bar14, Bar09a, Bar09b, Bar09c, Bar22, Bec84, Ber12, BA83, Bra01, Bra03a, Bra03b, CD98, Car00, Cha96, Cha23, Cla12, Cod09, Con99, Cuf19, Cuf20, Cun09, Cun10, Daw10, Deb86, Den83a, Dou00, Dou01, Dre16, DOHM96, Dui96, Dur22, Dyr98, Eck03, Ede87, EC00, EC01, Emi99, Eri11, Erv99a, Fer07b, Fer07c, Fer08, Fer09a, Fer09b, Fer10, Fer12, Fie67, Flo00c, Flo00a, Flo00b, Flo01a, Flo01b, Flo01c, Flo01d, Flo02a, Flo02b, Flo02c, Flo03d, Flo03b, Flo03c, Flo03a, Flo06d, Flo07f, Flo07a, Flo07c, Flo08, Flo10, Flo18, Flo20, Fre13, Gan22, Gar19, Gen10, Gir97, Gon20, Goo97a, Gra96a, Gre17, Gro99, Gro01, Gup83b, GG22a, GG22b, GG22c, GGS23, Hei58, Hen22]. Book [Hin10, Hol03, Hou21c, Hou21a, Hou21b, Hou21d, Hou22a, Hou22b, Hou23, Hue02, Hun98, Hyd18, Jan08, Jen99b, Jen00b, Joh65, Joh97a, Joh20, Joh02, Kae03, Kal97, Kam01, Kat84, Kir20, Kle03, Kre86, Kre87a, Kry00b, Kry00a, LM02, LJ10, Lay97, LeC01, LK08, LD20, Lie20, Lof08, LN98, LN00, Mac02a, Mac02b, Mac03a, Mag03, Mal01, Mal11, Man18, Mas08, McC19, McE18, Mic85a, Mic85b, Moe98, Mog09, Moo02, Mur00a, Mur00b, Mur00c, Mur00f, Mur00d, Mur00i, Mur00h, Mur00e, Mur00k, Mur00j, Mur00g, Nad96, Noa03, Nob10, Nor97, Nyc97, OM12, Pin18, Pod03, Pot00a, Pot00b, Pot01b, Pot01a, Ram03, Ram98, Ram00, Red98, Reh98, Rei02, Rei99, Ril00, Ril01, Ril02b, Ris98, Riv85, Roh19, Rok19, RA97, Ros64, Rou17a, Rou19, Rui20, Sac01]. Book [Sch72a, Sch73b, Sch73c, Sch73d, Sch73e, Sch20, Sei72a, Sel98, Sha68, Sim96a, Sim96b, Smi03, Smi10, Sou96, Spa14, Spi96, St.00a, St.01a, St.01b, St.01c, St.02a, St.02b, St.03, St.09, Sta99, Sta13, Sta20, Sto08, Sto15, Str01, Str03, Suo13, The97, Tra12, Vic03, Waa09, Wei01, Wei02a, Wha01, Wha02a, Wha02b, Whi76a, Whi76b, Whi77b, Wil03, Wil10b, Wil98, Win97, Wis15, Wof09, Xu08, Yea09, Yea10, You10, Zet19, Zha21, Zho08, Zim00, Zim01, vdM97a, Far62, Lee62, RLG97, Van77, Van62, Wel82, Yu24]. Booklet [Ano86m]. Books [Car24, McC65, Sch68, Bro86, Gup83b, Mur00k, Mur00j, TL92, Whi79e]. boom [Ort74]. Boosters [YJY24]. Bootstrap [Ano58c]. borders [St.03]. Boss [Hat66, LL09a]. Boston [And97, Nic59]. Both [Lay73]. Bottlenecks [Roh19]. Bottom [Flo07e, ZZL22]. Bottom-Up [ZZL22]. Bottoms [Ste62]. Boundaries [KN05, BTM12, JDJ99, KD00]. Boundary [PS12]. Bowker [Mir99]. Brad [Hou21a]. Brain [Gon20, Joh17, Yee90, Den83a]. brain/mind [Yee90]. Brainstorming [Tay82]. Brand [KWLS17]. Brands [Hou22a]. Braqq [FP79]. Breaches [HBM21]. Breadboarding [Bro80]. Breaking [KHV10, RFH87]. Brent [vdM97a]. Bresch [Zha21]. Brewer [Hou21b]. bricolage [CE05]. Bricoleurs [KW23]. Bridge [HK22, Hou22a, GG84]. Bridgeford [Boo17, Wis15]. Bridges [LAWI24]. Bridging [KD00, Kro80, MMVH10, SJD20, Wil24]. brief [Joh65]. Bring [Ano87k]. Bringing [GZ97]. brings [Eng89]. Britain [HvMK^{+}12]. British [Gra13, Hov14, Rai86, dGKNG06, dGKGN11]. British-English [dGKGN11]. Broad [You11, DH73]. broadcasting [Bar93]. Broadening [Ham96]. brochure [MSD85]. brochures [Whi77d]. Brockmann [Fil05, Kre87a]. Brooke [Dur22]. Brookhaven [Coc04]. Brother [DeT94]. Brown [Fer07c]. Bruce [Sch83]. Brusaw [Den83c, Den83b, Har80b, Whi77c]. budget [Wha86]. budgets [Gre87]. Buehl [Cha23, McE18]. Build [Gon20, McA85, Sto15, You80, Wal97]. Building [CHR04, DR20, Gir06, GC22, LAWI24, LG14, Pat08, Pin18, Sta20, CD98, Kle03, Xu08]. built [ML78]. built-up [ML78]. Bulletin [Cot59]. Bureau [Her95]. Bureaucratic [Ril00]. burnout [Cro86]. Burton [Bai24]. Bush [Pic80]. Business [AW17, ACW19, BP09, Boy81, Car20, DB24, Dos06, DR10, Far62, Fer05a, Fer07a, Flo05e, Flo06a, Flo07a, Gre70, HF84, Hat84, HvMK^{+}12, Hin10, Izq24, Jen99b, Kam01, Kie83, Kil09, KDBC20, Lam05, Lav06, Lin04, LHMN07, MML15, Mar08, PKLS17, PH24, Räi20, Rei81, Ril01, Wes71, AD88, AAR84, BD84, Car90a, Car90b, Cha80a, Cha80b, Cra98, Deg80, Den83c, Dou01, Flo00c, Flo01a, Fre80a, Gir60, HPSR83, Hul80, Hul81, Jac81, Ken82b, Koc03b, LY98, Nor00, ODF92, SR83a, SPR84, Sum72, ULK01, UC01, Wha83, Whi79c, Wol79, Bro83b, Ril00, St.00a, Bra03b, Mur00e, Gra96a]. business-strategies [Flo01a]. businesses [Fre13, Wha01]. Buurman [Hel07]. buy [Pak81]. bypassing [CC88].

C [Ana21, Bar09b, Bea99, Cla12, Cun10, Fer08, Flo07c, Flo20, Hof80, Mir99, Roh19, Sch75b, Sch78b, Sto15, The97, Yea09, Yea10]. C. [Luf64]. Ca [All97]. CAD [Kro80, Rav89]. CAD/CAM [Kro80]. Cadence [Bos87b]. calibration [Clo64]. Calif [Dob75]. Call [Ano64p, Ano75d, Ano77d, Ano77e, Ano77-30, Ano82d, Ano82e, Ano83l, Ano83j, Ano83k, Ano83-31, Ano84l, Ano85m, Ano86n, Ano86o, Ano87l, Ano87m, Ano04j, Ano08j, Ano09k, Ano09j, Ano10j, Ano10k, Ano10l, Ano11c, Bra82a, Joe80a, Joe81b, Joe81c, Joe81a, Joe82a, Joe82b, Joe83a]. Calls [Ano09l]. Cam [Ano83m, Kro80]. Cambridge [Nyc97]. camera [Koc75, Zap80]. camera-ready [Zap80]. camouflage [Sch81b]. campaign [FO06]. Campus [ASHB16, Dur22]. Can [Ano82f, Ano83n, Ano84m, BC70, BC17, Dom95, FW23, Har75a, Kna84, KPV08, LS23, Mil61, OL98, Sti59, ZM23, Ano73y, Bar89, Bar79, Bar82b, Blo87, BD82, Cos78, Dav81a, Gar73, Gle82, Jon89, Lay73, Mic90, Spe60, SS87]. Canada [Con91, Ros94]. Canned [KI06]. Cannibalized [KWLS17]. cannot [Sch80b]. can’t [Ask73, Bur84b]. Capacity [HDKC08, RD16]. Capital [BAB10, SC12]. capitalized [Mal87]. captioned [Cra85b]. Captions [Smi20, Mal85a]. Captivate [Ano07k]. captures [GvdMdJP99, GvdM02]. capturing [PDG^{+}88]. Card [Ken64b]. Card-Catalog [Ken64b]. cards [Ano64g, Ano64h, Ano65k]. Care [SDB^{+}13, WD09, Bon83, Rui20]. Career [AFH20]. Careers [Sta99]. Careful [Cha65]. carefully [Pil81]. Carel [Wei02b]. Caribbean [DGPR11]. Carl [Sch20]. Carlos [Hou21c]. Carole [Sch78d]. Carolina [Cos23b]. Carolyn [Rou17b]. Carroll [DOHM96, FW90]. Carter [Mon79]. cartoon [Rub87]. Casagrande [Nob10]. Case [BH15, BR16, BJ21a, Bed12, BP09, CR06, Cog15, CLP08, Dre17, Fre03b, Hov14, KZCR10, KS14, LMJK14, Mel09, Moo16, NKU03, O’D07, PMGW21, Raj12, RB19, Rou17b, Tri16, Wen69, Wil24, WS04, WARAL20, WRTBRL23, YFW15, ZM23, vdM97a, BBRT86, Ber92, Bro75, CC88, FBE^{+}03, FO06, Gat04, Goo05, Gri94, Gro01, HH82, HM95, KK90, KVRR06, Kus97, Lin90a, Lin93, Mac97, MG95, Mar97, McA85, NB04, ODF92, PM84, Por87, SF99, Str04, SH01, ZS04]. Case-based [Fre03b]. Case-Method [RB19]. Casebook [RA97]. Cases [Ano08j, Ano09k, Ano09j, Ano09l, Ano10j, Ano10k, Ano10l, Ano11c, Gra96b, FBE^{+}03]. Catalog [Ano64g, Ano64h, Ken64b, Ano65k].

Categorical [SKZ23]. Catherine [Bar22]. Causal [LG14]. causes [Mar86]. CCTV [FW78]. CD [Moo02]. CD-ROM [Moo02]. ceiling [Del92]. Cell [KNMM09, Lon14]. Cell-Phone [KNMM09]. Centennial [Ano84n, Ryd85]. Center [LE18, Mac04c, Pav94, Ano77g, Bam73, Wei75, DC07, Rho77]. Centered [BIdJ^{+}21, Fer04c, HVS^{+}22, JSZ07, SSS^{+}13, Sim22, SMP21, Pot01b, Tra12]. Centering [Wal23]. Centers [WP18]. Central [JW22, Mac04a]. cents [Sch82]. centuries [Nee92]. Century [Fil05, Pet14, Ram00, St.01b, Wha04, Der99, Dos06, MMW17, St.04, Wil03, Dre16, Sim96b]. CEO [HZX24]. Certificate [ZM23]. Chaim [Bra03a]. Chairman [Ano64i, Ano64y]. Challenge [Hov14, Reu65, Fla93, RW86]. Challenger [Dom91, Dom95, Dom07, Win88]. Challenges [AG21, Car12, Car16, Eva00, HK11, Wil03, Goo05, Suo13]. challenging [Pie58]. chance [Ano80f, Pas03]. Change [BG16, Cog15, Emi99, Goo96, Hat69, SY09, Cap93, Cra98, Cro80, FO06, Gal01, Gei77, Goo95, VLA96, St.02a, Wha02b]. Change-Up [BG16]. Changed [Mal11]. Changes [Car16, Nor00]. Changing [Baz99, Bli72, Mar64, Mis05, DH04, Ede77]. Channel [SS10, TM08, Ben85]. channels [McG00]. Chapter [Ano58d, Ano63l]. Character [SP87, Cam95]. Characteristic [Ant01]. Characteristics [DG75, Lam14, SS10, TJQN11, YJY24, RCS02, Sch67, Sch75a, SH01]. Characterizing [BMF21]. Charge [Flo04b, Mar77]. Charles [Den83c, Den83b, Fre80a, Har80b, Whi77c, Erv99b, Mic61b, Mic61a]. chart [Blo83, Mac82, Mac78b, Szo82]. Charter [Dre17]. Charters [FW23]. Charts [LO20, Tho67]. Chase [Ano80f]. chat [Pul86]. Chaucer [HN93]. Check [Ano64i, War80, Gro05]. Checked [Ver80]. checkers [Fur90]. checklists [Hap73]. Cheese [LR24]. Chemical [Jin18a, Jin18b, Lid75]. chemists [Men72]. Chile [SSV^{+}23]. Chilean [CSSV20]. China [Che92, DL24, FR13, GDTZ09, Jin86, Lin19, WARAL16, WOD^{+}16, YFW15]. Chinese [ACW19, Cha84, Che92, FC96, HZX24, LdJK20, SJYZ18, SWR21, WKSH68, WS23, XS24, Yu10, YB17, Yu20, Yua21, ZZL22, ZC24]. chip [Wil97]. Chit [Pul86]. Chit-chat [Pul86]. Choice [LL09a, LFT08, Lee00, RD05, SV06]. Choices [Nob10, ZWX20]. Choose [LG14]. Choosing [DV02, Szo82]. Christiane [Fer03]. Christina [Joh97b]. Christopher [Pri99]. chromophobia [Hor91]. Churchill [Mon78]. CID [PEG19]. CIM [MYV^{+}91, MYV92]. CIM/GEMS [MYV^{+}91, MYV92]. Circa [Mil65]. circulation [JE95, Mar77]. Citation [Smi10, Tuk62, HT77, SB79]. citations [Mou75]. Cited [Smi10]. Citizenship [SC12]. City [Dre17]. Civil [Par75]. claims [Hau80]. Clarification [Rei88a, Wil80, Wil82]. Clarity [Kos07, Aug78, Cor83, Dul92, Mue80, Rat85, WR83]. Clarity* [Kos08]. Clarkson [Hor84]. Class [Bed12, Yu11, ZWE21, Cor92]. Classes [DL24, MMCL00]. Classic [DuB08]. classification [App01a, BR92, Doo79, Tay73]. Classroom [Ant06, JD21, KSNF15, MAK15, Ric09, Wat20, All96, For04, Fre03b, Gol03, Hac84, Hor89, IP02, Jen00a, Sur96, Vel88, Zim93, Wil98]. Classrooms [JD23, Pet14, Gwi84]. Clauses [KCM15]. Clear [CEH87, LJ10, Rei81, Lip83, Nai81]. clearance [Ano77g]. clearer [Dul92]. Clearing [Cor90]. Clearly [Bai24]. Cleary [GG22c]. Client [BR16, KSNF15]. Client-Based [BR16]. clients [Kry99]. Cliffs [Die72, Nor97, vdM97a]. Climate [GYK14]. climates [BS89]. clip [Den98b]. Clips [vHK08]. Clobbered [Nob10]. closed [Cra85b, Mal85a]. closed-captioned [Cra85b]. Cloud [KMR14]. Cloud-Based [KMR14]. cloze [Gui89a]. Clue [McC63]. Clue-Word [McC63]. Cluster [MOA02]. CMS [BH15]. co [Kre00]. co-op [Kre00]. Coalitions [Sta20]. Coauthors [Luf69a]. Coded [Bec68]. Codes [Wil10b, Buc89]. Coding [Kor75, LBB18, Rea22]. Coefficient [LDKC23]. Cognate [Cal74]. Cognition [Kil96, KR95, Gir97]. Cognition/Culture/Power [Kil96]. Cognitive [Che17, DW07, KVRR06, SJD20, ST06a, RD05, ST05a, Red98, Vic03]. Coherent [Ano63e]. coherently [Bro83b]. Cohesion [PDL16]. Cohorts [Sta17]. coin [Ano63o]. Cold [Gra75a]. Colin [CD98]. Collaborating [WP18, MSC96]. Collaboration [BSS13, CCL24, GG22b, KN05, LRLM17, MPD10, MMVH10, Mun05, SJD20, Tho01, WLM^{+}22, BMRH05, FR86, GBF05, Goo05, JDJ99, PT05, QK05, Reh94, RL96, RVV^{+}02, Ana21, Dui96]. Collaborations [AME21]. Collaborative [DS89, Fer07c, Gra22, Hou22b, KMR14, Kni08, Lof10, MSDF22, MAK15, Ric09, SHB08, Str89, Bae12, IP02, LN03, LNB^{+}04, LNCL05, Nel03, RY90, RH94, RLS06, SM98, Wal99]. Collabratec [Ano15z, Ano15-27, Ano16t, Ano16u, Ano17y, Ano17z, Ano17-27, Ano17-28, Ano18v, Ano18w]. colleague [MSC96]. colleagues [Har84]. Collecting [LBB18]. Collection [Mel09, RK94]. Collective [Ric09, Dui96]. Collectivist [HK07]. college [Mag03, Mal87, Ste99]. Colleges [How83, How84]. Collin [Den98a]. Collins [Fer07b]. Colloquium [Ano63t]. Colocated [MAK15]. Color [KCLG13, Ort75, Win91, Bro91a, EZ88, Hor91, Tru91, WLMR82]. Colorado [OW91]. Columbia [Dom07]. Combination [Ano17d]. combined [TCW06]. Combining [Bat18a, ELdJ12, Whi87]. Come [Lam14, Wal97]. Comico [Luf69a]. Comico-Tragico-Tutorial [Luf69a].

Comics [Man98c]. Coming

[Doe69, Abe73, Li80]. commands [BRL91]. Comment [Ano64j, Ano64k, Ano64p, Dud61a, Kel61a]. Commentary [Bed07, Hof07, O’D07, Smi10, Baz99, Her64, Rob72]. Comments [Ano58j, Ano63u, Ano70b, Ano70c, Ano71e, Ano71m, Dav93, Doa19, Mil62, Mac04b]. Commercial [Mil62, Cov75, Har75b, KM94]. Commission [Gat04]. Commitment [BJ21b, HK07, ZZL22]. commitments [Pry95]. Common [Phe20, Yer74, MC75a]. CommuGuide\begin{matrix}c \\ ◯ \end{matrix} [Ano86m]. Commun [The97]. Communicate [WLM^{+}22, Ano82f, Ano83l, Ano83m, Ano83n, Ano84m, Cra98, Dav81a, Dio82, Joe79b, Joe83a, Mic84, Moo86b, Win89, Mon79]. Communicating [Ano79e, Bai24, BB99, Cap93, De 81, Dou72, Dud62, Gol03, Gre72, Gro99, HPSR83, LKHC16, RDJM09, Rob79a, RK91, Sko81, SI06, Sta17, BS89, HH93, Loe91, Mac02a, Wha04, Mus78a, Hen22, Rou17a]. Communication [BJ21b, AW17, ACW19, AG21, Ama04b, Ana21, AME21, Ano77-28, Ano78d, Ano78l, Ano78m, Ano78o, Ano78n, Ano79b, Ano80c, Ano80u, Ano80v, Ano80w, Ano80b, Ano81c, Ano81a, Ano81m, Ano81l, Ano82a, Ano82b, Ano83c, Ano83a, Ano84b, Ano85d, Ano86c, Ano86a, Ano87b, Ano96, Ano04n, Ano04o, Ano04p, Ano04q, Ano05k, Ano05l, Ano05m, Ano05n, Ano06o, Ano06p, Ano06q, Ano06r, Ano07o, Ano07p, Ano07q, Ano07r, Ano07s, Ano08o, Ano08p, Ano08q, Ano08r, Ano09o, Ano09p, Ano09q, Ano09r, Ano09x, Ano10o, Ano10p, Ano10q, Ano10r, Ano11e, Ano12r, Ano12s, Ano12t, Ano12u, Ano12v, Ano12w, Ano13n, Ano13o, Ano13p, Ano13q, Ano13r, Ano13s, Ano13t, Ano13u, Ano14p, Ano14q, Ano14r, Ano14s, Ano14t, Ano14u, Ano14v, Ano14w]. Communication [Ano15p, Ano15q, Ano15r, Ano15s, Ano15t, Ano15u, Ano15v, Ano15w, Ano15h, Ano15i, Ano16h, Ano16n, Ano16i, Ano16j, Ano16o, Ano16k, Ano16l, Ano16p, Ano16q, Ano16m, Ano16r, Ano16s, Ano17m, Ano17q, Ano17r, Ano17n, Ano17s, Ano17t, Ano17o, Ano17u, Ano17v, Ano17p, Ano17w, Ano17x, Ano18j, Ano18n, Ano18o, Ano18k, Ano18p, Ano18q, Ano18l, Ano18r, Ano18s, Ano18m, Ano18t, Ano18u, Ano19a, Ano19d, Ano19b, Ano19e, Ano19c, Ano19f, Ano20c, Ano20h, Ano20d, Ano20i, Ano20e, Ano20j, Ano20f, Ano20g, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano24d]. Communication [Ano24e, Ano24f, Ano24g, Ano24h, Ano24i, Bar77, Bat18a, Bay22, Bed07, Bed12, Bel99, BH21, Ben22, Ber12, Bli87, BS16, BC17, BP10, Boe10, BL13, BI18, Boi18, BSS13, Bra04, BH16, BD84, BWPH12, Car12, Car20, CS06, Cho21, Cho18, Cle16, Cod09, Coo10, Cos23b, CLP08, Cun10, Dav10, Daw10, DB24, DL17, De 72, Doa19, Dom07, DL24, DC07, Dyr98, DGPR11, Etz63, Far62, Fer12, FCS10, Flo04c, Flo05b, Flo06b, Flo06c, Flo07b, FPKS^{+}21, FL23, FVB16, Gar72, Gar19, GMG^{+}23, GLL23, GHG91, Goo96, Goo97b, Gra96a, Gre17, GZ97, GDTZ09, GGS23, HS18, HB01, HM61, HW09, Hat66, Her64, HK11, HCKJ03, Hot06, Hou21b, Hov14, Hun98, Hun21, Jam84, JC21, JES13, Joh63, KT14, Kil96, Kil09, KS14, Kir20, KHV10, Koc03b, KCJC08]. Communication [KHH^{+}08, Kur66, LM02, Lam05, Lam14, Lam16, Lam18, LBB18, LP16, Lav06, LWMR07, Lie18, Lof10, LDKC23, LAWI24, LHMN07, Mac04a, Mac04c, Mal07, Man18, MML15, McG00, McG72, Mel08, Mog14, MKBF09, Moo16, Nag87a, NLK10, Orr06, Orr10, Pad07, PD24, PM08, PEG19, Pas60, PKLS17, PDD^{+}24, Pod84, Rai86, Raj12, Ram05b, Ram00, Red10, Ril01, Ril02a, Rob72, Ros15, ROW19, Rou17b, RVvGS08, Sch68, ST99a, SS10, SGM^{+}24, Shr20, SR83a, SB00, SSW17, SLP^{+}79, Spi16, St.00a, SU09, Sta20, SA10, Str21, Swa18, Tha22, Tha24, VT21, VG16, Ven62, VMM21, Wal64, Wal18, WMPBA07, WX06, WC15, Wil03, WS04, Win88, WOD^{+}16, WRTBRL23, Yea09, YK13, Yor78, You10, Yu10, Yu11, Yua21, Zha21, ZZL22, ZM23, vDVvL19, Ama01, Ama00a, ABMF84, Ano77-30, Ano80e]. communication [Ano84-32, Ano84-33, Ano87n, Ano04s, AAR84, AN85, Bal08, BPK94, Bas72, Bat84, Bea99, Ben85, Bis99, Bos87a, Bos92, BW92, Bra03b, Bra93, Bre73, Bri99, Bro79, Bro84, Bro89, Bur75, Cap92, CS96, Cho97b, Cla87, Coc04, CR84, CC88, Con91, CE05, Cus93, DR94, Dav74, DeT94, DRG^{+}02, Del92, Den00, DF87, Dou01, Dou02b, Dri87, ESM98, ES99, Fea77, Fes62, Fil82, FBE^{+}03, Fla93, Fla05, Flo02d, Flo03c, For04, Gil03, Gal01, Gar79b, GRS96, Gil87a, GD85, Goo94, Goo95, GG96, Gre62, Gri91, GR94, Gui94, Gup83a, Gur95, Hac84, HP88, Ham72, HF84, HSV05, HFS94, HS72, Hor75, How83, How84, HM95, Jen00a, JP86, Jon89, KK90, Kes60, Koc01, Koc03a, Koc05, Kre87a, Lan94, Lin72, LK72, Luz05, Mac90, Mac99a, Mag03]. communication [Man98a, Man99a, Mar97, Mar84, Mas85, Max88, McD94, Min69, MOA02, Nag87c, NH05, Nor00, OW91, ODF92, Pal00, PD05, Par06, Par87, Pat08, PY04, Per96, Pfa86, PBK^{+}93, Plu81, Pot84, Pow82a, Pri72, Pul86, RCK^{+}90, Rat80, RH92, RO96, Rob73, RCS02, Rob96, Sal80, San87c, Sch81a, Sch72b, Sch78e, Sch80c, Sel99, Sen62, ST05a, Ske90, Som89, Son62, SPR84, Spe60, Spr97, St.99b, St.00b, St.01a, Str04, Stu92, SH01, Suo13, TWHN00, Tay73, TSS^{+}01, Ter72, Tha04, Tho01, TK04, Tru88, Tsu91, UC01, VLTP95, VLA96, Vis77, Vol82, Web79, Wei95, Woe79, Yee88, Zim83, Zim93, Zim01, vLUA02, All97, Ano77a, Ano77m, Ano77q, Ano77n, Ano77o, Ano77p, Ano08u, Ano09v, Ano09w, Ano22k, Fil06, Flo06d, The97, You10, Sac01, Ano64f]. Communication [Bar98a, Cha23, Eri11, Fer10, GG22a, GG22c, Hou23, Joh20, Joh02, Kam01, Mic64, Str01, Xu07, EC00, ES98, Gro01, Jen00b, Kae03, Kry00a, Moo02, Wha02a, Whi79c, Yu24, Ano73a, Ano74a, Ano75a, Ano76a, Ano78a, Ano79p, Ano79q, Ano79r, Ano79a, Ano80h, Ano80g, Ano80x, Ano80y, Ano80z, Ano80-27, Ano80a, Ano81n, Ano81o, Ano81p, Ano81b, Ano82l, Ano82m, Ano82n, Ano82c, Ano83w, Ano83x, Ano83y, Ano83b, Ano84y, Ano84z, Ano84a, Ano85t, Ano86u, Ano86l, Ano97c, Ano00, Ano04f, Ano04b, Ano04g, Ano04c, Ano04h, Ano04d, Ano04i, Ano04e, Ano05f, Ano05b, Ano05g, Ano05c, Ano05h, Ano05d, Ano05i, Ano05e, Ano06f, Ano06b, Ano06g, Ano06c, Ano06h, Ano06d, Ano06i, Ano06e, Ano07b, Ano07g, Ano07c, Ano07h, Ano07d]. Communication [Ano07i, Ano07e, Ano07j, Ano07f, Ano08a, Ano08f, Ano08b, Ano08g, Ano08c, Ano08h, Ano08d, Ano08i, Ano08e, Ano09a, Ano09f, Ano09b, Ano09g, Ano09c, Ano09h, Ano09d, Ano09i, Ano09e, Ano10a, Ano10f, Ano10b, Ano10g, Ano10c, Ano10h, Ano10d, Ano10i, Ano10e, Ano11b, Ano11a, Ano12c, Ano12h, Ano12f, Ano12i, Ano12g, Ano12j, Ano12k, Ano13c, Ano13f, Ano13g, Ano13h, Ano13i, Ano14a, Ano14d, Ano14e, Ano14f, Ano14g, Ano15a, Ano15h, Ano15d, Ano15i, Ano15e, Ano15f, Ano15g, Ano16a, Ano17a, Ano18a, Ano20a, Ano22a, Ano22b, Ano23a, Ano23b, Ano23f, Ano23c, Ano23d, Ano23e, Ano24a, Ano24b, Ano24c, BG07, Fry87, Whi80a, Ano85a, Ano86b, Ano87a, Van83]. Communication- [GZ97]. communication-based [Del92]. Communication-focused [Koc03b]. Communications [Ano63t, Ano67a, Dob75, Dud61b, Dud61c, Dwo69, Flo08, Fon61, Gia66, Got78, McK66, Wes71, Bli72, Bro89, Cas84, Dul93, Fri73, Goe78, Hil85a, Hil85b, KC81, Lay88, Mar89a, Per70, Phi69b, Sei72c, Shr72, Ste62, Sum72, Wea85, Whi89, Woe80, Woo90, Buc74, Mic65, Fre80a]. communicative [Sch78a, vLUA02, Fer08]. Communicator [Löc22, Ama02, Fis98, Fis99, Flo01d, Her62, Hol74, Lit90, MM00, Mic81, Ols91, Wil89]. Communicators [AD22, Cog15, Gan22, Ste05, And84, EC01, FS80, JM01, Lip04, Mar93, Mun02, RR93, Ske90, Win88, Dav93, Boo17, Cla12, Kal97, Zho08]. Communities [Boi18, GLCA22, SEN10, GRS96, Hea99, JDW96]. Community [BJ21a, Cle16, Gen10, Gra22, LL09b, RD16, RRW^{+}17, Rou17b, ZLCZ17, Kry99, Bar93, Dyr98]. Community-Based [RRW^{+}17]. Comp [Gra75a]. Companies [Raj12]. companion [Flo00c]. Company [KdJH14, Fra89, MG95, MV83, Rab92, War88, YJ01]. Comparative [KC19, RWAK24, WC15]. compared [GvdMdJP99, LD97]. Comparing [Bus91, NNB^{+}22, SA23, WARAL16, Yua21, Che92]. Comparison [Car20, Cos23a, CDM10, Gre62, GDTZ09, KU04, Lin72, MMCL00, UN00, WBW89]. comparisons [Hub85]. Compensatory [Koc01, KCM15]. Competence [Yu12]. Competencies [Mel08, BTM12]. competing [HR99]. competitive [Bli84, BS89, HSR84, St.00a]. competitor [Ano85z]. Complaint [Sch10]. Complementing [BP09]. Complete [Ano97d, Lee62, Mar08, Mic85a]. Completion [Ano97d]. Complex [Dor63, Fer04a, Fil06, Pla71, Cra92, Gol97, HHE97, Pri97, Smi60]. Complexity [GG13, KT14, KL12a, KR95, TCW06]. Compliance [GYK14]. compliments [Mac06]. Component [And14, AB15a, Bat14, Bat18b]. Components [GG22a, Rat85, vdMG04]. compose [Rei79]. Composing [Loe95, Sus85]. Composition [Bra01, Bur69, KG68, Lam68, Mac75b, MAK15, Par75, Sha68, Var77, Ada06, Dob75, Hol77, Kor77, MM73, Kno80]. compositor [FHR87b]. Comprehend [YK13]. comprehension [FS00, KVRR06, Kuo93, LR85]. comprehensive [And84, Ano87y, Ano06j, Ano06k, Ano07l, Ano08k, Ano09m, Ano09n, Eck03, Gro94, Sno64]. compress [Ano78r]. Compressed [Gre69]. Compression [YTY12]. compromise [Ano04r]. Computational [ESPQ15]. Computer [Ano77f, BSS13, Boy68, CB08, Cho97b, Cli63, Cor69, Dui96, FWB86, FVB16, GY68, GMT08, GMG^{+}23, GHR67, GG13, HT77, Hor84, KAOC81, KG68, LBE23, Lip71, Mar86, Mor67a, Mor67b, MC75b, Ols96, Pie71, Pul86, Rea22, Sed69, Sha68, VNC^{+}13, Vic03, Wil98, AACF81, Ano87-28, Ant01, Arm83, Bas97, BCB84, BS73, Bli87, Bro83a, Bro86, BB86, Car90c, CS96, Cra92, Cre95, Cus93, DR94, DeT94, Fee86, Gri84, GR86, Gri83, Gui89b, HP88, HFS94, Jon78, Kau88, Koc03a, Kru00, KM94, Kva94, Lev86a, LM97, Man98b, MYV^{+}91, MYV92, Mus78b, NKU03, Nyc97, Orr99, Pfa86, RLS06, RW86, Rub86a, SB79, Spi86, Sub76, Sub79, Wea85, Kai68, Sal80]. Computer-Aided [Lip71, Mor67b, Mor67a, Mus78b, Kai68]. Computer-Assisted [GY68, LBE23, Hor84]. Computer-Based [Dui96, GMT08, Vic03, NKU03]. Computer-Controlled [Sed69]. computer-delivered [Car90c]. computer-integrated [MYV^{+}91, MYV92]. Computer-Mediated [BSS13, CB08, FVB16, GMG^{+}23, GG13, DR94, HFS94, Koc03a, Kru00, Kva94, RLS06]. computer-network-supported [CS96]. Computerized [Bar80, Bar81, Boy69, Fea77, Jos70, Kun69, Met75, Pet69, MSW97, Tel68]. Computers [Gia66, Sam70, Sel99, Ama01, Jen99a, KF77, MB84, Mar85, Rob79a, RL84, Sim96b]. Concept [Ano97d, DF70, Doe69, DF75, Loe95, Tru91]. Conception [JSZ07, Sul89]. Conceptions [HdJK18]. Concepts [LBE23, Fry87, Pot84, Ril02a, Ter72]. Conceptual [WFSC^{+}23, Gui89b, KVRR06, ZVS01, Whi78a]. Conceptualization [vDVvL19]. Concern [Ben22]. Concerning [Ano87n]. Concerns [ROW19, Ken89, Ros94]. Concise [Flo06d, Ano87y, Moo02, Ril02b]. conclusions [CW90, Sek73]. Concurrent [Coo10, ELdJ12]. conduct [Irw72]. Conducting [Bec85, Hul81]. Conference [Ano63k, Ano67i, Ano73g, Ano75m, Ano81m, Ano07o, Dob75, KM61, LRM^{+}09, PDK88, Ack89, Ano86a, Buc74, LC79, Moo85, Pot83, Ano67a, Ano80b, Ano81a, Ano82a, Ano82b, Ano83a]. conference/workshop [Buc74]. Conferences [Ano63d, Gar79a, Irw72, Ker86, Spi86]. Conferencing [GDTZ09, Fea77, Fee86, FWB86, Lev86a, Pul86]. Confessions [Ger90]. Confidence [LKS13]. confident [Hor91]. Configuration [Roy07]. conflict [Baa94, HS72]. Conformity [KDBC20]. Confronting [AD22]. confronts [Zur75]. Confucian [Lee00]. Confused [Bha12]. confuses [CR73]. Congressional [Ano85b, Ano85c]. Congruency [HvMK^{+}12]. congruent [RK91]. Connecting [CM05, HLY22, TM15]. Connections [Caw75, Ano04j]. Connectivity [CC19]. ConocoPhillips [YFW15]. Conributors [Ano74c]. Consciousness [Nyc97]. consensual [KR95]. Consent [SGM^{+}24]. Consequences [HF08]. Consideration [Rin59]. Considerations [Boe76, FS85, Gui89b, Ken89]. Consistent [ZJG17]. consonant [Mon81b]. consortium [Bro91b]. Constantinides [Pen98b]. Constitutional [Sta13]. construct [IPC01, PY04, TCW06]. Constructing [PH24]. Construction [Arn62, DW07, Kon22, Mor67b, WS23, ZFWX21, Ano80o, Bra78, Bue77, Duk89, KR95]. Constructive [vdHdJS06, Gro03]. Constructivist [vdM97a]. Constructs [AAA19]. consultant [Ger90, Sid87]. consultant-taught [Sid87]. Consulting [KHV10, Spi96, BD84]. Consuming [Nye99]. Contact [FS09, Ama03b, Tuk62]. contacts [Gar79a]. containment [Bea98]. Content [And14, AB15a, AB15b, Ano13v, Bat14, BA16, Bat18b, BP10, Bot70, CNC20, Car24, Cla16, Cog15, Dre16, GL16, GAGW15, GPHDM16, Hac16, Hou21c, JD21, LdJK20, MP20, MFL17, Smi20, WC15, Ano09u, CJ89, HS72, Tho01]. Content-Management [And14, Bat18b, Cog15, GAGW15, GPHDM16]. Contents [Ano59b, Ano60d, Ano60e, Ano61f, Ano61g, Ano61h, Ano62c, Ano64l, Ano65b, Ano79f, Ano79g, Ano79h, Ano79i, Ano80k, Ano80l, Ano80m, Ano80n, Ano81d, Ano81e, Ano81f, Ano81g, Ano82g, Ano82h, Ano82i, Ano82j, Ano83o, Ano83p, Ano83q, Ano83r, Ano84o, Ano84p, Ano84q, Ano84r, Ano85n, Ano13y, Ano14-30, Ano15-35, Ano15-36, Ano16x, Ano16y, Ano17-29, Ano17-30, Ano17-31, Ano18x, Ano18y, Ano18z, Ano18-27, Ano19k, Ano19l, Ano19m, Ano20o, Ano20p, Ano20q, Ano20r, Ano21m, Ano21n, Ano21o, Ano21p, Ano22n, Ano22o, Ano22p, Ano22q, Ano23n, Ano23o, Ano23p, Ano23q, Ano24j, Ano24k, Ano24l, Ano59p, Ano59q, Ano63x, Ano63y, Ano64-30, Ano64-31, Ano65o, Ano66n, Ano66o, Ano67n, Ano68x, Ano68y, Ano68z, Ano69u, Ano69v, Ano69w, Ano70o, Ano71y, Ano71z, Ano71-27]. contents [Ano04t, Ano04u, Ano04v, Ano04w, Ano05o, Ano05p, Ano05q, Ano05r, Ano06s, Ano06t, Ano06u, Ano06v, Ano07u, Ano07v, Ano07w, Ano07x, Ano08v, Ano08w, Ano08x, Ano08y, Ano09y, Ano09z, Ano09-27, Ano09-28, Ano10t, Ano10u, Ano10v, Ano10w, Ano11g, Ano12-33, Ano12-29, Ano12-30, Ano12-31, Ano12-32, Ano13w, Ano13x, Ano13z, Ano14k, Ano14l, Ano14-28, Ano14-29, Ano14-31, Ano15-33, Ano15-34, Ano16v, Ano16w, Ano17-32]. Contesting [Bra93]. Context [DB24, HW09, Kam01, Shr20, Spi96, Den88, EM94, SI06, Nyc97]. Contexts [Ano09x, BSS13, Dav10, KT20, LO07, Mel08, WBB13, BT95, Gro01, Gwi84, Dou00, Bea99]. Contextual [Mun05]. Contextualization [CWSL19, KT14, SC12]. Continuing [Den78]. continuous [BMRH05]. Continuum [Ley08, Sch89]. Contract [KCM15, Bec84]. contract-winning [Bec84]. contractor [Koc03b]. Contracts [PKLS17, GB84b, Gre85, Wha01, Fre13]. Contradictory [Doa19]. Contrast [Izq24, St.99b]. contrasting [Chi89, Kos95]. contrastive [McD94]. contributing [Win88]. contribution [vLUA02]. Contributor [Ano58e, Ano59c, Ano59d, Ano59e]. Contributors [Ano63f, Ano63g, Ano64m, Ano64n, Ano64o, Ano65c, Ano65d, Ano66c, Ano66d, Ano67c, Ano67d, Ano68a, Ano68b, Ano69b, Ano69c, Ano70d, Ano70e, Ano71f, Ano71g, Ano71h, Ano72b, Ano72c, Ano72d, Ano72e, Ano73h, Ano73i, Ano73j, Ano73k, Ano74d, Ano75e, Ano75f, Ano75g]. Control [Bli61, CC22a, SP87, Cap92, D’A65]. Controlled [Sed69, KSW78]. Controversy [SvI18, Til06]. CONTU [KF77]. Conundrum [Kos07, Kos08, Mac99a]. convention [ZZ05]. Conventional [Ken64b, MMCL00, Tho67]. Conventions [BMF21, Ber12, Kae05, Mis05]. Convergence [PD24, LF95, Pod03]. Conversation [GG22b, Phe21, Fre03a, NC81, Gen10]. Conversations [Mog09]. Conversion [Lie20]. Conveying [Dos06]. Conwell [McE61]. Cook [Flo07a]. Cookbook [Bro83b]. COOL [Wei98]. Coombs [You10]. Cooperation [Bal75, Hut75, Lid75, Heu75, Mou75]. cooperative [CS96, DRG^{+}02, Mir99]. Coordination [JJ10, PDL16, LN03]. copies [Urb77]. Coping [DLVE23, KL12a]. Copy [Ano68w, Dev75, Har75a, Mir91]. Copyediting [Wol88, Zim88]. Copying [Wei77, FWJ77, JF03]. Copyright [Ano77g, Tur74, Vis77, Ano77y, Ano77z, Ano77-32, Ano78k, Ano79o, Ano80t, Ano81k, Ano82k, Ano83v, Ano84x, Ano85s, Ano86t, Ano87t, Ano04m, Ano05j, Ano06l, Ano07n, Ano08n, Den99, Gan77, Koc77, Lie77, Lin77a, Pic80, Tal77, Zur75]. Core [Sub76]. Coronas [Mas08]. Corporate [BAB10, Bha12, Bro89, CM02, Flo04c, Goo96, Goo97b, Her89, HZX24, Joh10, KdJH14, Lee13, Lin19, Lon95, Pry95, Ram05a, SFM09, Woo90, Woo80, YFW15, Yu20, Yua21, ZC24, Zol14, ZB15, BW92, Bra03b, BS89, BK03, Cap92, Foy96, Goo95, Gri91, Gur95, LF95, RO96, Som89, Wan89, Whi89, Whi91, Pot00a]. Corporation [SZLY22]. Corporations [SKZ23, YFW15]. Corpus [Ant06, BW19, HZX24, Izq24, KI06, Kon22, LBE23, LMJK14, Orr06, Sam06, YB17, dGKNG06]. Corpus-Assisted [Kon22]. Corpus-Based [BW19, HZX24, Sam06, YB17]. Corpus-Informed [Izq24]. Correct [KCM15]. Correction [Ano61i, Ano69d, Ano71i, GY68, GHR67]. Correctness [Sch75c, Luf76]. correlation [BW92]. Correspondence [Ano62d, Ano75h, Lam16, Rav06]. Corridor [MOA02]. Cortical [Goo05]. Cortright [Sal61]. Cost [Cra68, Dre75, Ese69, Man98a]. Costs [Mac75b, Ste08, Kun73, MM73, Wha86]. Could [Dul92]. Council [Hut75, Ano08s]. countering [Han80]. Counterintuitive [KCM15]. Course [Bay22, CB08, Cos23a, DL24, GPHDM16, LO07, Ric09, SA23, Bro84, Cod04, Cra83a, JF03, KD84, Kre89, Rob78, Top87, Wha83]. Course-Based [Ric09]. Courses [Car24, Cho18, Doa19, GLL23, KO20, Mog14, Van71, BFH03, Cod04, GG96, Irw72, Sid87, WS04]. Courseware [Bur85]. Covello [Hen22]. Cover [Ano18h, Ano18i, Ano58f, Ano59g, Ano59h, Ano60f, Ano60g, Ano60h, Ano60i, Ano60j, Ano61j, Ano61k, Ano61l, Ano61m, Ano62f, Ano62g, Ano62i, Ano64q, Ano65e, Ano65f, Ano65l, Ano66e, Ano66f, Ano67g, Ano67h, Ano67m, Ano68d, Ano68e, Ano68f, Ano69f, Ano69g, Ano69h, Ano70g, Ano72a, Ano72f, Ano72g, Ano72h, Ano72i, Ano72j, Ano72k, Ano72l, Ano72m, Ano72n, Ano73f, Ano73n, Ano73o, Ano73p, Ano73q, Ano73r, Ano73s, Ano73t, Ano73u, Ano73v, Ano74e, Ano74f, Ano74g, Ano74h, Ano75c, Ano75i, Ano75j, Ano75k, Ano75l, Ano75n, Ano76b, Ano76c, Ano76d, Ano76e, Ano76f, Ano77b, Ano77i, Ano77j, Ano77k, Ano77l, Ano77s, Ano77t, Ano77u, Ano77v, Ano77w, Ano77x, Ano78f, Ano78g, Ano78h, Ano78i, Ano79c, Ano79d, Ano79j, Ano79k]. cover [Ano79l, Ano79m, Ano79s, Ano79t, Ano79u, Ano79v, Ano80d, Ano80p, Ano80q, Ano80r, Ano80s, Ano80-28, Ano80-29, Ano80-30, Ano80-31, Ano80-32, Ano81i, Ano81q, Ano83d, Ano83e, Ano83f, Ano83g, Ano83h, Ano83i, Ano83s, Ano83t, Ano83u, Ano83z, Ano83-27, Ano83-28, Ano84c, Ano84d, Ano84e, Ano84f, Ano84g, Ano84h, Ano84i, Ano84j, Ano84t, Ano84u, Ano84v, Ano84w, Ano84-27, Ano84-28, Ano84-29, Ano84-30, Ano85e, Ano85f, Ano85g, Ano85h, Ano85i, Ano85j, Ano85k, Ano85l, Ano85o, Ano85p, Ano85q, Ano85r, Ano85u, Ano85v, Ano85w, Ano85x, Ano86d, Ano86e, Ano86f, Ano86g, Ano86h, Ano86i, Ano86j, Ano86k, Ano86p, Ano86q, Ano86r, Ano86s, Ano86v, Ano86w, Ano86x, Ano86y, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano87h, Ano87i]. cover [Ano87j, Ano87o, Ano87p, Ano87q, Ano87r, Ano87s, Ano87v, Ano87w, Ano87x, Ano08l, Ano08m, Ano12l, Ano12m, Ano12n, Ano12-33, Ano13j, Ano13k, Ano13l, Ano13m, Ano14k, Ano14l, Ano14i, Ano14j, Ano15j, Ano15k, Ano15l, Ano15m, Ano16d, Ano16e, Ano16f, Ano16g, Ano17i, Ano17j, Ano17k, Ano17l, Ano18f, Ano18g, Ega81, Fox81, Mic60, Mic61c, Mic61d]. coverage [Ter73]. COVID [GLCA22, RWAK24]. COVID-19 [GLCA22, RWAK24]. Crack [Wel22]. Crafting [Bra03b]. Crawford [CDL84]. Create [FW23, Cod04, Gol97, LN00]. Creating [Boe10, Har97, Jen02, KT20, LNB^{+}04, SB84, STM09, Wat90, Win99, Hou21c]. Creation [MP20, RL96, LD20, Ano80i]. Creative [Arm82, Bue76, Car16, MFL17, Mic61e, Rin63, Ano80h, Ano80g, Ano80j, Rau83, Ste62, Tru88, Car16]. Creativity [Sch80c, Ven62, LF95, Lay88, Ock05, Ano80f, Sch80a]. Credibility [FS09, MY14, MYT16, WS23]. Crises [MVV09, TVRR20, Som89]. Crisis [Hen22, Kon22, Str04, You10, Coc04, Wil97]. Cristina [Zha21]. criteria [Gat04, Whi79a]. Critical [AG21, GG22a, JW17, JW22, Kon22, Mir91, WPSB13, Yu20, ZB15, Bis99, Bro88b, Joh90]. criticism [Gro03]. Critique [Bos64, Joh10, Kae04, Smi10, Und67, Wat20, Ano73-27, Han83]. critiquing [Rav88, Zim88]. Cross [BMS^{+}15, Flo01a, HvMK^{+}12, RH92, WX06, Yu23, Zem12, AC06, Gil03, KD00, Loe91, Loe95, MM00, Nor00, RKP00, SB00]. Cross-Cultural [WX06, Yu23, HvMK^{+}12, Zem12, AC06]. Cross-Disciplinary [BMS^{+}15, Gil03]. cross-functional [KD00, Loe91, MM00, Nor00, RKP00, SB00]. Cross-gender [RH92]. Croteau [Erv99c]. Crowd [VG16]. Crowd-Based [VG16]. Crowdfunding [VG16]. Crown [Cle03]. Crucial [MO80, Fla93]. CSR [SJYZ18, YB17]. CSS [Kle03]. Cueing [STP22]. Cues [GMG^{+}23, HTB80]. cuff [Nee80]. Cultural [And14, Cor92, Fer08, GAGW15, Gor90, Kae04, Kam01, Kos95, LL09a, LRLM17, LdJK20, LK72, St.04, WX06, Whi14, Yu23, AC06, BT95, Den88, Gri94, HvMK^{+}12, Lee00, Sub92, Zem12, vLUA02, Den98a]. Culturally [MP20]. Culture [Car16, DGPR11, FR13, HK07, JSZ07, Kil96, SEN10, Sim96b, Ama02, BS89, Cap93, Her89, dALT92, Lon95, MSC96, Ols91, RH01, St.99b, St.02a, Tha04, ULK01]. Cultures [Ama01, Car12, GTTX08, Bus91, DRG^{+}02, Reh94, VVL^{+}01, Luf64]. Cunningham [Hou21d, Sch75b, Whi79d]. Curbing [Pen99a]. Current [AB15a, Ano68t, Rin63, Ano73y, Fra75, Wan89]. Curricula [PM08]. curricular [VLA96]. Curriculum [CLP08, DL24, FC24, KW23, Lul06, NLK10, ZB17, Bra93, CR84, Eng89, Fla93, Ivo80, Ken88, Ken89, OW91, Par87]. curse [Sie78]. Customer [HHE97, JJ10, RD16, SMS96]. Customer-Service [JJ10]. Cut [Gri84]. Cuteness [Rok19]. cuts [MM73]. cutting [St.02c]. cutting-edge [St.02c]. cyberculture [Der99]. cyberspace [Dum96, Gur95]. Cycle [Hov08]. cyclical [GR94]. Cynical [Coc04]. Cynicism [Hou22a].

D [Ano63-27, Ano64e, Bra03a, Bro83a, Cla12, Cun09, DR10, Erv99b, Fer07b, Fer10, Flo07c, Flo08, Gre17, Sch78b, Sch80b, Sta13, You10, Zim99, Mic68]. Dakota [SvI18]. Dan [Mag03]. Danes [Gri94]. D’Angelo [Kno80]. danger [Sau92]. Daniel [Flo03d]. Dank [Cam07]. D’Arcy [Ama00b]. Darwinian [KHH^{+}08]. Data [Ant18, Ath22, BIdJ^{+}21, BI18, Boy68, Coy77, GC22, HBM21, KPV08, Kos07, Kos08, Lam18, Lam68, LBB18, Lav06, Sal68, SGM^{+}24, Str21, Wel22, Wor87, Ano78r, BR93, Blo87, DSC78, FWB86, GKD77, GV82, HHE97, Lid75, Mud04, Mey87, Pad78, PDK88, PM04, RK94, RL84, San83, Sch78e, Cuf19, Cuf20]. Data-Analysis [Ant18]. data-base-produced [GKD77]. data-collection [RK94]. Data-Driven [BI18, Lam18, LBB18, Lav06, Mud04]. data-privacy [PM04]. Database [Ack89, CS06, HC88, ST06a, SI06]. databases [Flo03a, Hil89, SD06, Wei93]. David [Ano97a, CD98, Her95, LD20, Whi79c, Bra99b, Mic85a, Roh19, Rou19]. Davis [Ano64e, NB04]. Dawson [Pen98d]. Day [Mic80]. deadline [Ano78r]. Deadly [Kni08]. deaf [Mal87]. Dean [Ama00c]. Dearin [Bra03a]. Death [Ano71l, Ano71x, Weh77]. Debate [Ano63h, SvI18, BB91, Fra75, Sha89]. debates [Pod03]. Debating [ROW19]. Deborah [The97]. Debunk [BD82]. decade [Rus77]. Decades [JSZ07]. Decalogue [Wen85]. Deceitful [KCM15]. decentralized [Ske90]. Deception [FMT12, GMG^{+}23, GG13, St.02b, Zho05]. Deceptive [LO20, LS23]. Deciphering [Buc89]. Decision [All97, Baa07, Fri13, AV98, Get86, HSR84, Riv85, San91]. Decision-Making [Baa07, Fri13, AV98]. Decisions [LBE23, MO80, Stu92, Wil75]. Decks [SNT^{+}14, SNT^{+}15]. Declarative [Mac05b]. declared [Ros79]. Decorum [Mar85]. dedicated [WBW89]. deductively [FS00, SF02]. Deep [Mar02]. defective [WM87]. Defense [St.05b, Fis84, Koc03b, Zim99]. Deficient [Shr20]. Defining [Fis98, Per63]. Definite [Ano07k]. definition [BB91]. Definitions [Bun71, Loe68]. Degree [XS24]. degrees [ST99a]. Deibert [St.09]. delay [She73]. Delays [HBS08]. deliberative [ZS04]. deliver [Jin86, LD84]. delivered [Car90c]. Delivering [Goo97a, Riv85, Wei01]. Delivery [BG16, Mel09]. Demand [Sau93, Hur77]. demographics [ODF92]. Denied [St.09]. Denmark [MYV^{+}91, MYV92]. Dennis [RA97]. Denton [Den83c]. Department [Her95, BD83, Sam88]. Dependency [CWSL19]. Deploying [GAGW15]. depth [Goo05]. Derivative [Van75]. Describing [RH94, Swa18]. Description [Rau21]. Descriptive [KB17]. desert [Ano86-27]. Design [Ara59, Bal18, BCJR87, CB08, DL24, DOHM96, Fer04a, Fer04c, Fer05a, HSA21, HVS^{+}22, Hel07, Hov14, JD11, Jon15, KW23, Lan04, Lin07, Lon14, MP20, MFL17, MAK15, Pin18, RWAK24, Ris98, Roh19, SSS^{+}13, Sim22, Sto08, Sto15, Str21, Tha22, Tra12, WD09, Wat20, Wei02b, Wil24, Wof09, Zet19, ZJG17, Ama05, Ano81v, Bal08, BRL91, BBRT86, Ber92, Bos92, Bur85, BK03, Cas86, Cra83b, Cra92, DRG^{+}02, Die89, Evi04, Eva00, Fee86, Flo03b, Gan04, GB84a, Gri91, Gui89b, Her95, Hur83, Key87, Kos95, Kro80, Kva94, LH00, Lip04, Lip58, Mur00b, Mur00g, ND83, NH05, PDG^{+}88, Par87, Pot01b, Ram03, Rub86b, SM98, SB91, SWB^{+}05, Str99, Swa05, Whi77d, Win91, Yee90, ZVS01, Zim00, vLUA02, RA97, vdM97a, Joh97a, Spy98, Wei02b, Wei02a, Mur00g]. design [Ril02a, Ana21, Gon20, Zim13]. Designed [LDKC23]. Designers [Ale20, Fri13, GB16, vMPKG15, Whi82a]. Designing [Ach24, All96, Ama03a, Cra83a, DH04, DWP85, Fer05a, GG96, Hof05, JD21, JDW96, Lop84, Mar82, MMW^{+}22, Pet14, RDRS04, Rei78, RCB82, Rok19, Ske90, SMP21, vdGRRv13, Bag83, BCB84, Gil03, JD83, GLCA22, Cla12, Goo97a, Hou23, Ril02b, Sch78b]. Desktop [Bro88a, Bro88b, FHR87b, PDK88]. destiny [Zim00]. Details [RDJM09]. Detection [GMT08, MC75b]. determining [Dri87]. develop [Rab92, Sau93]. developed [PFM88]. developers [CM01, Ale20]. Developing [Ano83l, Ant06, BR16, Bar91, Fer04a, HH82, Izq24, Joe83a, KW23, Per63, ROW19, Wha83, Yu10, And84, Eva00, Rua80, Flo01b]. Development [AG21, BG16, Con70, DL17, DM63, DGPR11, Flo06c, Hac16, LL09b, MKBF09, PEG19, Phe21, SJD20, SY09, SP87, WMPBA07, ZB17, Zem21, AD88, Adl00, BM00, Boc98, BMRH05, Die04, Fis98, Fis99, Hac85, LS94, McC94, NKU03, Pak83, PY04, SSNJ05, ST05a, SB91, DOHM96]. Developmental [Rau21]. developments [SLP^{+}79]. Device [LCX12]. Devices [Loe68, JVT01, Mar75]. Devising [Ric09]. dexterity [GW84]. DHTML [Kle03]. diagnostic [Har97]. Diagram [Bro71]. Diagrams [DR10, PKLS17, BS98, ML78, Rav89, Win90]. dialog [Hul81, Jon78, MCWK77]. Dialogic [Dre16, PT05]. Dialogue [Ano64-29, Lon14, SNT^{+}14, Fre03a, TWHN00]. dialogues [EM94, Haw94]. Diamond [Ano17d]. Diary [Flo07f]. dictating [GB79]. Dictator [Cha59]. Dictionaries [Lac63, Lac87]. Dictionary [LN04, Mor67b, RL84, Mic61b]. Did [Lam14, McE58, Pil81]. Difference [FW23, dALT92]. Differences [HvMK^{+}12, Jin18b, KPV08, LdJK20, SSSC08, BR93, KD00, TCW06, TK04]. Different [Car12, KdJH14, LLY18, Sam06, Gre58, His77, Spi94]. Differential [BWPH12]. Differently [LLY18]. difficult [CH89]. Diffusion [LHAL23, ZB15, Ano04s]. Digest [Buc74, Ken63]. digests [Lip75]. Digital [Ano12o, Ano12p, Ano12x, Ano12y, Ano12z, Ano12-27, Ano15x, Ano15y, Flo05d, GHR67, Heb97, Moo05, Rif10, Ris98, Spa14, Bar98b, DH92, St.01c]. Digital-Survey [Rif10]. Digitally [Bec68]. Dilemma [McK66, RM03, Stu92]. DiMarco [Sto08]. Dimension [LL09b, HB01, Hoh60, Sub92]. Dimensional [KG68, MMW^{+}22]. Dimensions [HvMK^{+}12, MV21, The97, TM08, WBB13, Chi89, Fra75, Sch81c]. Direct [YK13]. Directed [NP23]. Directions [AAA19, Wil03]. directness [RM03]. Dirk [Joh17]. disabilities [CM01]. Disability [BH21, Ben22]. disadvantaged [Mal85b, Wal85]. Disciplinarity [BMF21]. Disciplinary [BMS^{+}15, FMB18, Kil96, YJY24, Gil03].

Discipline [CC22a, Flo01c, Har00]. Discipline-and-Control [CC22a]. Disciplined [Mic68]. Disciplines [Car12, PEG19, Smi20, For04, Hel07]. Disclosure [ZNG22, Cot79, GH85]. Discourse [Baa07, DV09, Fer08, FMB18, Flo05e, Kae05, Kon22, LR24, MPD10, Mis05, Mou05, Ols96, Red98, Rei99, Rob18, Sam06, Sho80, Ama03b, RK91, Tra83, WB92, Ber12]. Discover [BG07]. Discovering [SRVR97]. Discrepancies [Wol06]. Discrete [Loe68]. discriminability [WS92]. Discursive [FDJW18, WS23, Nor00, Fer07a]. Discussing [ROW19]. Discussion [Bos64, Jin18a, Jin18b, Dav74, GR89, KL06, Lou58]. Discussions [Ano63d]. Disease [App01a]. disgruntled [Ken82a]. dishing [Cow80]. disorders [Man86]. display [GV82, Kat84]. Displayed [Kor75]. Displaying [LM97, JD83]. Displays [Kos07, Kos08, Aic78]. Disruptions [BJ21a]. disseminated [Bas97]. disseminating [RLG97]. Dissemination [HK70, Sal68, WRTBRL23, BB75, Bha77, Jos75, WM87]. Dissertation [HK22, Mic85a]. dissolve [HS79]. Distance [SJD20, YTY12, CS96, Cod04]. Distinct [GTTX08]. distraction [Bro91a]. Distress [DLVE23]. Distributed [AG21, BSS13, CHME07, MPD10, PHP16, PHW13, SC12, Bel99, BMH05, JDJ99, LN03, LNB^{+}04, VVL^{+}01]. distribution [Mat91, Sta73]. disturbance [Man86]. DITA [ESPQ15, Hou21c]. Dittrich [Fer03]. Diverse [FW23, Gal01]. Diversity [Bay22, BWPH12, Car24, VT21, Wal21, ZWX20, BW92, RW86]. Divide [Wil24, PT05, Mir99]. Division [ZWX20, Wil77b]. Dixie [Ede87]. Do [Ano68w, Ano14h, BN07, Dob85, Flo04b, Kle81, LLY18, WLMR82, WLM^{+}22, ZM23, And84, Bur90, CH89, Cos78, Duk89, Gla04, Gui86, Joh82, O’C62, Red89, Sch67, Sch75a]. doc [Win99]. doctoral [Mic85a].

Document [Hou21a, Kin12, MYV^{+}91, MYV92, SP87, Spy98, Vel89, Wat20, Wei02b, Ano81v, Die72, Flo03a, Gan04, GN94, Han83, Heb97, Her95, Koe77, LH00, MSW97, Nel90, Rat86, RM03, SB91, St.85, Top87, Vog83, Wei93, Whi82a, Win91, Yee90, dJS00]. Document-driven [MYV^{+}91, MYV92]. Documentary [Sal68]. Documentation [Ano80o, Dia14, Flo04d, GLCA22, LJ10, Ode86, AC85, Arn84, Bar90, BJ86, BCB84, BR92, Bra84, Bro86, BB86, Com86, Del00, FW90, Get86, GR86, Gri81, Gud82, Gui86, Gui89b, HHE97, Joe82b, Kap62a, Man98b, Mar86, MV83, McG00, McG86, Mir91, Ora86, Pak81, PFM88, Rav89, Rei88b, Ril02a, RW86, Rub86a, Sch86, SMS96, SB91, SC81, Sty86, Tha04, Wal87, WB92, Gen10, LN98, Sou96, Pak83, Smi10]. Documenting [Lyn83]. Documents [Ano97d, Boi18, Gil03, GQ12, Gra96b, Rav06, SDB^{+}13, Ste22, Tho70, Gol97, HN93, Har97, Hor89, Hor90, Mar89b, RLG97, SSM74, SHS97, UN00, Win89, Nad96]. Does [Hun21, Vel87b, vdM96, dALT92, dJS00]. Doing [Dos06]. Dollars [Sch82]. Dollars-and-cents [Sch82]. Domain [NNB^{+}22, GN94]. domains [KR95]. Dombrowski [Sac01]. Dominance [Ock07]. Domo [Cam07]. Donald [Hou21d, Sch75b, Whi79d]. done [JP86]. Donna [OM12]. Don’t [Ano67e, Ste22, Kle81, Sou72, Sou75, Vog83]. don’ts [Duk89]. donuts [Whi91]. door [Fil82]. Dot [Mal11]. Double [Con58b, Pei81, Mon81b]. Douglas [Flo03b]. Doumont [Cod09]. Down [ZZL22, Ben85, Bos87a, Fei59, Lub80, RFH87]. DP [Van83]. Dr. [Ano63i]. draft [Bue76, Dre83]. Drafting [OL98]. drafts [Gil03, McC92]. Drama [Luf69a, Luf68]. Draw [Mus78a]. Drawing [Den83a]. Drawings [Dav81a, Ano79n, Mag77]. Dreyfuss [Pic80]. Drive [Lie20, Ram05a]. Driven [BI18, Hov08, Lam18, LBB18, Lav06, vdM07, Mud04, May82, MYV^{+}91, MYV92]. Driver [Pad07]. Drowning [Flo07e]. Dual [Ano59i]. Dudley [Kel61a]. Duffy [GG22b]. Duin [Hou22b]. duplicate [San83]. during [Dul93, Rus77]. Dutch [BPK94, dGKNG06, dGKGN11, vMPKG15]. Dutch-English [dGKNG06, dGKGN11]. Duty [Hag61]. Duyne [Flo03b]. Dyad [LRLM17]. dyadic [ES98]. Dynamic [Fil06, Man99a, Reg70, RLG97]. Dynamics [And14, Emi99, Fer09a, GAGW15, Kni08, LD20, PHP16, PHW13, Goo05, Gur95, SM98, Spy98]. Dyslexia [Phe21].

E-Collaboration

[KN05, Mun05, BMRH05, GBF05, RVV^{+}02]. E-communication [Koc05]. E-cooperative [DRG^{+}02]. e-Environment [KC19]. e-Human [Fro06, Mas08]. e-Learning [McC06, Bae12]. E-Mail [Flo07e, Cho97a, HB94, TDS95]. earcons [BRL91]. Earl [Str84]. Early [CBSM^{+}09, Cog15, KL12b, PH24, SAP20]. Early-Stage [PH24, SAP20]. Ease [McE61]. easier [Atl81]. Easy [Lan04, CH89, Lin90b]. Ecodefense [Ros15]. Economic [Waa09, Kin77, Zur77]. Economics [JLS20, Sam06, Day83]. economies [JE95]. Economy [Car16, Ram05a, SU09]. ecosystem [Stu92]. Ecotage [Ros15]. Ed [Die72, DOHM96, Flo02c, GGS23, Hel07, Jen99b, LN98, Mic65, Nor97, Nyc97, Sch75b, The97, vdM97a, Bra05, Dou01, Flo02a, Flo02d, Flo03d, Flo03c, Fre80a, Har80a, Hue02, Joh02, Kae03, Kno80, Kre87a, LM02, LeC01, LK08, Mac02b, Mar08, Pot01a, Wha01, Zim01]. Ed. [Bro79, Zap80, Hol03]. edge [Bli84, HSR84, St.02c]. Edit [Nad96, ACH^{+}98, Atl81, Sch83]. Editing [ACH^{+}98, BNB21, Bli76, Bur90, CSSV20, CC22b, Hou21d, KG68, Tho75, WH98, AT88, Axe73, Bec85, Boc98, Cha77, Dev75, FP95, Har75a, Har84, Kan85, KAOC81, LY98, Lut84, Mas85, Vel87b, WBW89, Mur00a]. Edition [Cla12, DR10, Flo06d, Flo07c, Hou21b, Pet13, Xu07, You10, Den83b, Flo00c, Flo00b, Fre13, Hof80, Kry00a, Mac02a, Wei01, Whi78b, Whi79f, Yu24, Flo05a, Flo06a, Kre86, Mic85b, Bra01, Ril01]. Editor [Ano71w, Bau62, Hay23, Lay73, Pet69, DH73, Gro94, Pow81, San91, Sch79, Smi85, Ano64p, Ano66b, Ano68q, Ano73w, Ano73x, Ano08u, Ano09v, Ano09w, Ano20b, Hay16, Hay21, Hay22, Hay24, Kai68, Nag86, Tra74]. Editor-author [Lay73]. Editorial [Ano59f, Ano62e, Ano63j, Ano67f, Ano68c, Ano68g, Ano69e, Ano69i, Ano70f, Ano70h, Ano71l, Ano71j, Ano71k, Ano73l, Ano73m, Phi72a, Phi72b, Sal62, Sei72b, Sei72c, SA10, Wel75, Wel76, Bam73, Car90c, Cha60, Gan58, Rho77]. Editors [BNB21, Doe69, KB17, St.09, Wis15, AV98, Boc98, Dav85, FP95, Fey88, GR86, Hut75, WBW89, Car73, Wei02b, Flo01b]. Edmond [Rav06]. EDP [Gri81]. Eds [All97, Bar09b, Cha23, Eri11, Fer03, Fer07d, Flo07b, Gre17, Mas08, Mog09, RA97, Rou19, Waa09, Yea09, You10]. Eds. [Pic80]. educate [Bar79, Bar82b]. Education [AME21, Bar14, CD98, Flo20, GB16, LL09b, MMW^{+}22, PD22, Rea22, RVvGS08, Spa14, St.06b, Yu10, ZM23, BTM12, CS96, CR84, CM05, Den78, Orr01, San87c, Bar14, Lay97, Bro79]. Educational [LC17, STP22, vdM97a, LNB^{+}04]. Edward [Ale20, Hou21d, Kat84, Moe98]. Edwards [Den83a]. Effect [BH15, CL19, FS09, Lie18, LHAL23, Mac05b, PDL16, PFD^{+}21, Wes71, ES98, FS92, HMF02, Lon95, MRE96, SMS96, SF02]. Effective [AAR84, Cod09, DPHF17, De 72, EM78, Flo06b, GD85, Gre70, Hor75, Hul80, Kie83, KS14, Kin12, Kir20, Lam05, Pod84, Pow82a, Pra87, Ama03a, Bas72, Bos87b, Bul72, CEH87, CM02, Dou01, Duf87, Eva60, GG96, Gup83a, Hay80, Lee64, Los85, Mal01, Ril99, Ril02b, Sau82, Tho67, Van83, Wei01, Whi79a, Whi79g, Sal61, Fre80a, Fre80b, Mur00b, Zha21]. Effectively [WLM^{+}22, Ano82f, Ano83m, Ano83n, Ano84m, Ken88, Ken89, Mon79, Rau79]. Effectiveness [CL19, Dav67, GDTZ09, KCJC08, PDL16, PFD^{+}21, PHP16, PHW13, BFH03, Cor83, EE78]. Effects [BW19, BWPH12, FMT12, FVB16, GMT08, GG13, HBS08, Hes75, KdJH14, KCM15, LRLM17, LKS13, Ock07, Sha68, STP22, SEN10, TM08, ZZL22, CJ89, Cre95, GvdMdJP99, KS04, KA03, Mac04b, She73, Tay73, TCW06]. Efficacy [Wil24, ZWE21]. efficiency [Cam92]. efficient [Ken64a, Sch82]. Efficiently [TM15]. Effort [WT09, Smi60]. Efforts [Cus93, LL09b, NP23, LY98]. Ego [AB81, ZLCZ17]. Ego-speak [AB81]. Egonets [LP16]. Ehrlich [Ano63-27]. Eidetecker [Ols91]. Eight [Pot82, Mac03b]. eighteenth [Nee92]. Eighty [Har81]. Eimbinder [Ano64p]. Eisenberg [Whi79e]. Eisenson [Ano64f]. Elaborated [FW23]. Electrical [Coi69, Fou59, LN04, Rau21, VLA96, Ano87-28, DRG^{+}02, KC81, Oli58]. electrical/electronics/computer [Ano87-28]. Electricity [Coo65]. Electro [Lev68]. Electro-Optical-Mechanical [Lev68]. Electronic [CDM10, Dor63, DC07, Goo97b, HdJK18, HF08, Hov08, Ker86, KHH^{+}08, KL12a, Lam05, Lan05, Los85, Pie71, Pla71, Red98, Ros94, SDB^{+}13, SP87, TM08, Wes71, All96, EM94, FBE^{+}03, Gol97, Har97, Hea99, Heb97, Hou83, Key87, Koc01, Lin93, Lip58, Man99a, Mur00h, ODF92, PY04, Pri97, Pul86, QK05, RFH87, Spi94, TWHN00, vLUA02, Pot00b, Win97]. Electronics [Coi69, Coo65, Cro59, DM63, Lac63, Lac87, LN04, Ano87-28, Bur75, Kin58]. electrotechnology [Ano85z]. elegance [Whi87]. elegant [Woo79]. Elements [Car90c, Hue02, Rav06, Yer74, Ano64e].

Elevator [MV21]. Eleven [BD82, Mac03b]. elite [Rob79b]. Elizabeth [Pet13]. eloquence [DC87, MD84]. else [Jen99a]. EMACS [PDG^{+}88]. Email [Cun09, PY04, WHC^{+}12, Mal01]. Emailing [LL09a]. Embodiment [TDG^{+}18]. Embodying [MSDF22]. Emergence [Ram98]. Emergencies [ASHB16]. Emergency [ASHB16, Cos23b, GE13, Mar89a]. Emerging [LMJK14, VG16, Boy87, Bur75, Per96]. Emerson [Flo18]. Emily [Ano97b]. Emoji [DB24]. Emoji/Emoticon [DB24]. Emoticon [DB24]. Empathetic [KW23]. Empirical [ASHB16, AG21, Erv99a, GYK14, KMR14, MKBF09, TJQN11, XS24, vDVvL19, Zho05]. Employed [Jen00a]. Employee [CDM10, DC07, GYK14, ODF92]. Employees [LLY18, Sta17]. employment [Ham96]. Empty [HFS94]. Enable [Hov14]. enablers [SSNJ05]. Enabling [Pas03, RLG97]. Enact [PD22]. Enacting [Wel22]. enchanted [JS81]. Encoding [SP87, Win90]. Encountering [Han80]. Encounters [Whi14]. Encyclopedic [Sch70, Sch75d]. end [BMRH05, Der99]. end-user [BMRH05]. endings [PP97]. ends [Ano63r]. energies [Nye99]. Energy [SKZ23, Tel68]. Enforcement [Ste22]. Engagement [Lie20, MSDF22, Moo16, STP22]. Engaging [Bal08, MP20]. Engine [Kil13, LMJK14]. Engineer [Ano68r, Gol58, Hag61, Joh63, Lac63, Lay97, McK66, Ama04a, And81, Ano80j, AN85, Boy87, Gre72, Her64, HS72, McC79, Moo85, Nai81, Nus79, Sau92, VLA96, Wal74, Woe75, Woe79, Kry00b]. Engineering [AME21, Ano58i, Ano59m, Ano59n, Ano59o, Ano60a, Ano61a, Ano61u, Ano61v, Ano61w, Ano62j, Ano63m, Ano63n, Ano64e, Ano64s, Ano64t, Ano64u, Ano65g, Ano66g, Ano66h, Ano66i, Ano66j, Ano67j, Ano67k, Ano68h, Ano68i, Ano68j, Ano68k, Ano68l, Ano68m, Ano69j, Ano69k, Ano69l, Ano69m, Ano69n, Ano70i, Ano70j, Ano70k, Ano70l, Ano71n, Ano71o, Ano71p, Ano71q, Ano71r, Ano71s, Ano71a, Ano97d, Ano08u, Ano09v, Ano09w, Bar98a, BJ21a, BMF21, Bli61, BC17, Bot70, BJT^{+}19, Cam92, CB08, Con58b, CLP08, Dav10, Doc61, Dud71, Duf87, Fei59, FMB18, Flo20, FC24, Fri19, Fuj58, Gra13, Gre58, HK22, Hov14, Hyd18, Jin18a, Jin18b, JD21, Kan12, KT18, KT20, KS14, KO20, KRM24, Lee62, LC17, LO07, LL09b, LDKC23, LN04, Mac61]. Engineering [Mac04c, Man18, Mar70, Mic59, Mic61e, Moo87b, Mor66, Mor67b, Nel03, NLK10, PM08, Rau21, RDJM09, Rok19, Rou17b, Rou19, Sto15, Wee19, WP18, WFSC^{+}23, Yu11, Yu12, ZWE21, Ano84-32, Ano84-33, Ano87-28, Ant99, Ask73, Bar82a, Bra78, Bri99, Bul72, Bur84a, Bur84b, Car87, Caw75, CR84, D’A65, DH94, Den78, Dou72, Eva60, For04, Fra89, Fre03a, Gar72, GV82, Gou60, Har89, Har00, Hil89, IP02, JDJ99, Kae03, KD84, Kli79, Kol93, Kre86, Kre00, KC81, Kry99, Lyn83, Mud04, Mac04b, Mac05a, Mal87, Mat91, MSC96, Mic62, Mic74, MF83, Mic84, Mil65, Mor67a, Orr99, Orr02, Orr03, Pie58, PP97, Rai86, Rat86, Ren96, Ros83, Sei67, Spe60, Str60, Van63, Wal99, Woe80, Zim01, Bai24, Flo20]. Engineers [Ano63v, Ano63w, Arm83, BP09, Bra04, Coi69, Flo05c, FPKS^{+}21, Fou59, HS19, Hou21a, KHV10, Luf64, Mur59, PEG19, RM61, VMM21, And83, Ano82o, Ano82p, Ano82q, Ano83-29, Ano83-30, Ano84-31, BPK94, Bar79, Bar82b, Bli76, Bli87, Blo87, Bro79, DRG^{+}02, Gar72, Gwi84, Kna84, KW93, Mac72, Mey64, Moo86b, Moo87c, NKU03, PBK^{+}93, Rob72, RC89, Sch67, Sch75a, Smi03, Ter65, VLTP95, WP82, Ano64k, Par75, Bar22]. Engines [Kil09]. England [BT95, Mal07]. Englewood [Die72, Nor97, vdM97a]. English [Cun09, LJ10, Ram98, Rav06, Ano84s, AFH20, Che92, Cra85a, Den88, DSV10, Fey88, FC96, Fur90, Gre58, Izq24, JD21, Jon80, KS14, KSW78, Mud04, McD94, MC75a, NKU03, Ock71, Ock75, Orr01, Orr02, Orr03, Räi20, RL78, Sch63, Shi81, Shi82, SHS97, St.99b, TL92, Tsu91, VV16, YJ01, YB17, dGKNG06, dGKGN11, vMPKG15, Sch78a, Whi80b]. English-Language [DSV10]. Enhance [BC17, MMW^{+}22, TWHN00]. Enhanced [Kle03, BW92]. Enhancement [McC19]. Enhancing [DW07, Flo06c, WPSB13]. enough [Gro94]. Enron [HWW04]. Enterprise [CWSL19, KWLS17, Xu08]. Enterprise-Wide [Xu08]. Enterprises [WS23]. enters [Gri84]. Entertain [SWR21]. Enthusiasm [BW90]. Entrepreneur [SNT^{+}14]. Entrepreneurial [LKHC16, PH24]. Entrepreneurs [LP16, SNT^{+}15, WARAL20]. Entrepreneurship [Ano15h, Ano15i, MV21, Spi16, SAP20, VG16, WARAL16]. entry [San83]. Environment [Con70, Gro99, KC19, MYT16, Bro89, Ede77, Gro95b, LF95, MSC96, ODF92, Sau92, Zur77]. Environmental [Boi18, Cre95, GGS23, Moo16]. Environmentalism [Yu20]. Environments [Bat18b, Flo04d, vdM97a, LNB^{+}04]. Envisioning [DOHM96]. EPC [Whi77a]. EPC-II [Whi77a]. ePortfolios [FC24]. Equation [LG14, Ham72]. Equations [ZJG17]. Equilibrium [Hea69]. Equipment [Dor63, Key87]. Equipping [Gan22]. Equitable [Ach24, Gra22]. Equity [Bay22, Car24, Rea22, Wal21]. equivocality [DR94]. Ergonomics [Mar84, Bro78]. Erin [Flo07f]. Erlbaum [Joh97b]. eroded [NB04]. Error [Man99b]. Errors [Fey88, MC75b, Bai81b, MC75a, Rav88]. Escape [Der99]. essay [BCB03, Kil82, Woo79]. essay-writing [BCB03]. essays [ABMF84, Rei99]. Essential [Bed12, Ril02a, Wat90, Mur00b].

Essentials [Ano84s]. EST [Fre03b]. Establishing [DPHF17, KT18, ZZ05]. esteem [Gui94]. Estes [Hou21b]. estimated [Row78]. Estimating [Wal85, ZJG17, Abr77, BCR00]. Estrin [Sch75b]. Ethic [ROW19, Wel22, Stu92]. Ethical [Dre16, Gar19, Par87, SFM09, Chi89, Dav93, DF87, Jen00a, Kre89, Mar93, Mar02, Ros94, Stu92, Wil97, Yee88]. Ethics [AV98, BW92, Cla87, Dom07, Fra81, Gar06, Her95, Mar97, RDJM09, Ros15, Bal08, Buc89, Cam95, Cha02, Dom95, HM94, HWW04, Mic90, RY90, Rub87, TDS95, Sac01, Kre87a, Hun98, Hyd18]. ethnographic [Whi89, Zim93]. Ethnography [HS18]. Ethnography-Based [HS18]. ethnomethodological [TC95]. Ethos [Cam95, Rif10, Til06]. Etiquette [Flo07a]. Eugen [Pen98a]. Eugene [Ano80g]. Eunjeong [Dur22]. European [Bal75, Bar93]. Evaluate [Pak81, Kru97, LN00]. Evaluating [Dru85, Hof05, Per63, RRW^{+}17, Sch89, Tri16, Wil10a, Ama01, Hub89, RCB82, Tha04]. Evaluation [DL24, Fri15, LR24, NNB^{+}22, vSdJ07, Adl00, GR89, JVT01, LD97, LH00, PDG^{+}88, RLS06, Sha82, dJS00, vW00]. evaluations [Ros89]. Evaluative [Sam06]. Evans [Roh19]. Even [Nob10]. Ever [Ano67e]. Every [Lie20]. Everybody [Luf69b, Luf75]. Everyday [LO20]. everyone [Jen99a]. Everything [GL16]. Everywhere [AD22]. Evia [Hou21c, Yea09]. Evidence [BJ21b, Hen22, Lee13, Lin19, Moe98, Til06, WOD^{+}16, ZZL22]. Evidence-Based [Hen22, Lee13]. Evolution [Bed07, Abe73, GR94, Koc05, Mac72]. Evolutionary [PD24, SSSC08]. Evolving [Bar14, Dom07]. EWS [Ano64p, Ano67i, Phi69b, Ano61n, Ano63k, Ano64j, Ano64r, Ano64-28, Ano65i, Ano67o, Ano71a, Her64, Tro66]. EWS-1 [Ano61n].

EWS-14 [Ano71a]. EWS-3 [Ano61n]. Examination [AZDA10, FMT12, Jon15, Dul92, vdM96]. Examinations [SHB08]. Examining [JW22, MKBF09, SS10, SNT^{+}14, St.07, SU09, St.99b, Lam14]. example [Win88]. Excellence [Ano20b, Hay21, Man98a]. EXE [GQ12]. Executing [vHK08]. Executive [Kle81, NLK10, Bon83]. Exemplars [Daw10, Hou21b]. Exercise [BH16]. Exercises [vHK08, Gla04]. Existing [Boy68, GG84]. Expand [Ano17e, Ano17f, Ano17g, Ano17h, Ano18d, Ano18e]. Expanding [Bou62, Fry87, KN05, St.00b]. Expansion [TM08]. Expatriates [ACW19]. expectation [HCKJ03, Rub86b]. Expectations [JC21, Bar82a, Que86]. Expected [LFT08]. Experience [Ale20, BG16, FL23, GG13, HdJK18, KT20, Mer05, Pet13, RRW^{+}17, RVvGS08, Wil24, WBB13, WJ13, ZB17, Zem21, ZM23, vdGRRv13, Fla85, Jen02, Men72, Ser88, Vel88, Pin18, Ano80h]. Experienced [LFT08]. Experiences [Hou23, Mun05, Kre00]. Experiential [Sou88, Zem12]. Experiment [Ven62, BB75, Wat90, Whi77a]. Experimental [BL13, CDM10, Dav67, FR13, GY68, KCJC08, Mir91, Ock05]. Experiments [Die72, Mus78b, DBM73]. Expert [BCJR87, NNB^{+}22, LD97, McG86, WBW89]. expert-focused [LD97]. expert-novice [WBW89]. expertise [Mac04b]. Experts [BP15, Fla93, Han80]. Explain [KPV08]. Explanation [Ols96, Man01]. Explanations [Moe98]. Explanatory [Ols96, Win89]. Exploding [Pod03]. exploration [YJ01]. Explorations [Ram00]. Explorative [Roy07]. Exploratory [CBSM^{+}09, Fri15, Fri19, GTTX08, VVL^{+}01, WARAL20, SSNJ05]. Explore [Ano06j, Ano06k, Ano07l, Ano08k, Ano09m, Ano09n]. Exploring [Ber12, Boe10, HS18, HS72, HLY22, JLS20, Jon89, Kae05, LWMR07, MEZ12, SU09, Tha24, dALT92, St.01a]. expository [FS99]. exposure [Con87]. Expressing [Dos06]. Expression [Tri16, Mus78b]. Expressions [Ano68u]. Extending [LCX12, ACH^{+}98]. Extensible [App01b]. extrapolation [Mil65]. Eye [ELdJ12, LS23, Men23, Rob96]. Eye-Tracking [ELdJ12]. Eyes [Bea59, Bro87, Car89].

F [Ano63d, Bra99b, Daw10, Hou21b, Joe79a, Sch20, Whi80b]. Face [CHME07, Dom07, Fie68, GDTZ09, SA23, Vel90, VNC^{+}13, WT09, Mar82, RLS06, ULK01]. Face-to-Face [CHME07, Fie68, GDTZ09, SA23, WT09, RLS06]. Facebook [CC22a]. faces [Flo73]. facet [Sch81a]. Facilitate [RB19]. Facilitating [PS12]. Facilitation [Ken64b]. Facilitators [AZ17]. Facing [GG13, Phe21]. fact [Row78]. Factors [ASHB16, Gan04, Hov14, KNMM09, Kal97, PFM88, XS24, ZNG22, ZLCZ17, ZB15, vSdJ07, Cle87, Fes62, Par87, Ren96]. Faculty [LDKC23, PEG19, Phe21, ZB17, Zem21, HFS94, Ren96]. Faculty-Facing [Phe21]. Fahim [Dur22]. fail [BRS81]. failure [HM95]. failures [Win88]. Fair [Ano77h, Sop75]. familiar [EE78]. Familiarity [MML15]. Family [MML15]. Fantasy [WC15]. Faraday [Ano78p]. Farce [Luf71]. Farrar [Mek72]. Farrell [Ano64j]. Fast [Lan04]. Faster [Lin90a]. Fatal [Luf71]. fate [St.02b]. faxmodem [SH93]. Feasibility [GH70, Pla71, Bec84]. Features [Mis05, Ant01, Chi89, ESM98]. February [Phi69b]. federal [AD88, Blo87]. Federalese [Ano78e]. Feedback [CL19, Doa19, Mel09, NP23, BMRH05, Dre75, Evi04]. feeding [Bon83]. Feeling [Lam18]. fellowships [Ano85b, Ano85c]. female [Bro93]. Feminisms [MSDF22]. Feminist [AW22, Bos92]. Fern [Kre87a]. Ferreira [Pin18]. fiber [Li80]. Field [Jin18a, Jin18b, LP16, Zim13, Ano08t, Ano09t, Orr99, Pie58, Son62, TSS^{+}01, RA97]. Fielden [Flo07e]. Fields [Rou19, Sto15]. Fifteen [Geo83, Mal07]. Fifteen-Year [Mal07]. fifth [Mic85b]. Figure [Smi20]. figures [HMF02]. filing [FC79, HE77, Mau79]. Film [vHK08]. films [Mur58]. filter [Dat91]. Filtering [St.09]. Financial [dGKGN11]. Financing [Cru73]. find [Blo87]. Finding [Cos23a, SMP21, Gro94]. Findings [AG21, BESFW22, Yu10, Gat04]. finish [CW90]. Finnish [YJ01]. FireAnt [Ant18]. Firearms [Her95]. Firm [Fri19, Bat84]. Firms [Yua21, ZZL22]. First [Dre16, FR13, Ivo80, Roc59, Sim96b, AACF81, Bro65, Dre83, Ram00, St.01b, MMW17]. first-time [AACF81]. Fisher [Sch78d]. Fit [BJ21b, BG16, Hun21, BMH05, DM81, KVRR06]. Five [Ano63k, Sch68, Wri81]. Fix [Bal18]. Flaherty [Zha21]. Flammia [Suo13]. flashed [Mal87]. Flat [Aic78]. Flat-panel [Aic78]. Flesch [Joh65]. Flexible [FW23]. Flint [Kon22]. flip [Blo83]. flip-chart [Blo83]. Florence [Cos23b]. Florida [GLCA22, ODF92]. Flow [Reg70, Hol74, Jen02, Koc03b]. Floyd [Fer03]. Fluency [Den98a]. Focus [Ano09x]. Focused [KO20, Koc03b, LD97, Sch89]. foe [Ano80-33]. fog [Ver80]. folios [Wei98]. Follett [Whi75]. follow [Ber82]. Following [KDBC20, Mac03b]. Font [SP87]. footnotes [JVT01]. for-profit [Ham96]. Foraging [HBS08]. Force [And14, Kni08, Ste22, Bou58]. Forecasting [Hat69, Pot77]. foreign [Cor92, Kus97, Luf65, McD94, MSC96]. forensic [Boy87]. Foreword [Joe77a, Ano84n]. Forewords [dGKNG06]. Forgotten [Flo18]. Form [Ano68u, Bra05, Bro71, Hol03, Ano79o, Ano80t, Ano81k, Ano82k, Ano83v, Ano84x, Ano85s, Ano86t, Ano87t, Ano04m, Ano05j, Ano06l, Ano06m, Ano06n, Ano07n, Ano07t, Ano08n, Ano12a, Ano12b, HS72, Min69]. Form-2nd [Hol03]. formal [TCW06]. format [GB84a, SJ99, Ber68]. formats [BKSM82, Ter73]. formatting [KA03]. forms [RH94, San83]. Formula [Nor58]. formulaic [Mac06]. Formulas [Cor69, Ano81h, BRS81, McC87, Plu81, Red81]. Fortune [JD11, Jon15]. Forty [Blo83]. Forty-six [Blo83]. Forum [Ano81h, Ano86z, Ano87z, Ano87-27, Olk87, Spi94]. Fostering [Bay22, KW23, PKLS17]. foundation [Fry62, SWB^{+}05]. foundational [Mar97]. Four [Bag83, Gol82a, Jon78, LCX12, Smi20, SSNJ05, vdMG04]. fourth [Flo03c]. Fragmentary [Web65]. Frame [BG16, Zim93]. Framework [Abr77, CC22b, GG22a, HK22, KZCR10, Lam18, Pad07, BK03, GBF05, OC97, ZVS01]. Frameworks [BESFW22]. Framing [SvI18]. Franca [Räi20]. Frank [Kno80, Whi79c]. Franklin [Bro65]. fraught [Cal74]. Free [GL16, Sam70]. Free-Text [Sam70]. Freedman [Arn69]. Freedom [Cha06, CC22a, Flo07f, Sha89]. Freeman [Cha80a]. Freeware [Ant06, Ant18]. Frequency [Gra13]. Frequent [Boy81]. Frey [Wha01]. Friend [Ano68u, Ano80-33]. Friendly [JD21, Flo00a]. Front [Ano12-33, Ano18h, Ano18i, Ano65l, Ano67m, Ano72l, Ano72m, Ano72n, Ano73s, Ano73t, Ano73u, Ano73v, Ano74h, Ano77v, Ano77w, Ano77x, Ano79s, Ano79t, Ano79u, Ano79v, Ano80-29, Ano80-30, Ano80-31, Ano80-32, Ano81q, Ano83z, Ano83-27, Ano83-28, Ano84-27, Ano84-28, Ano84-29, Ano84-30, Ano85u, Ano85v, Ano85w, Ano85x, Ano86v, Ano86w, Ano86x, Ano86y, Ano87v, Ano87w, Ano87x, Ano58f, Ano59g, Ano59h, Ano60g, Ano60h, Ano60i, Ano61k, Ano61l, Ano61m, Ano62f, Ano62g, Ano64q, Ano65e, Ano65f, Ano66e, Ano66f, Ano67g, Ano67h, Ano68d, Ano68e, Ano68f, Ano69f, Ano69g, Ano69h, Ano70g, Ano72f, Ano72g, Ano72h, Ano72i, Ano72j, Ano73n, Ano73o, Ano73p, Ano73q, Ano74e, Ano74f, Ano74g, Ano75i, Ano75j, Ano75k, Ano75l, Ano76d, Ano76e]. Front [Ano76f, Ano77i, Ano77j, Ano77k, Ano77l, Ano78f, Ano78g, Ano78h, Ano78i, Ano79j, Ano79k, Ano79l, Ano79m, Ano80p, Ano80q, Ano80r, Ano80s, Ano81i, Ano83s, Ano83t, Ano83u, Ano84t, Ano84u, Ano84v, Ano84w, Ano85o, Ano85p, Ano85q, Ano85r, Ano86p, Ano86q, Ano86r, Ano86s, Ano87o, Ano87p, Ano87q, Ano87r, Ano87s, Ano08l, Ano08m, Ano12l, Ano12m, Ano12n, Ano13j, Ano13k, Ano13l, Ano13m, Ano14k, Ano14l, Ano14i, Ano14j, Ano15j, Ano15k, Ano15l, Ano15m, Ano16d, Ano16e, Ano16f, Ano16g, Ano17i, Ano17j, Ano17k, Ano17l, Ano18f, Ano18g]. frontline [Man98a]. Fuel [Ano14m, Ano15n, Ano15o]. fulfillment [Flo75]. Fully [DSV10]. Function [Ara59, Gra13, Str64, Cha60]. Functional [KMR14, Gup83a, KD00, Loe91, Loe95, MM00, Nor00, RKP00, SB00]. Functions [LCX12, Mac06, Rob73]. Fundable [Zim99]. Fundamentals [Ale20, Flo80, LK08]. funding [Meh94]. Funnel [FW90]. Funny [Nag87b]. Further [Ano71m]. fusion [Baa94]. Fusions [BH21]. Future [AAA19, AFH20, Jos70, Ano62e, Ano73y, Lit90, Met73, She73, McC19, Flo07b]. Futures [Car16, Hou22b, Tha22].

G [Bar98a, Bar09c, Bar22, Bra03a, Daw10, Fer07a, Flo03d, Hel07, Jan08, Sch20, Sto15, Str84, Wha04, Ano64r, Ano65i, Ano73m, Tro66]. G-EWS [Ano64r, Ano65i, Tro66]. G-PC [Ano73m]. G. [Mir99]. Gainesville [GLCA22]. Gaining [Lof08]. Gains [Sch17]. Galaxy [CFMM14]. Game [BC17, Jen02]. Game-Based [BC17]. Gamers [May19]. games [Cas86]. Gamification [KWLS17]. Gap [Cos23a, MMVH10, GG84, HCKJ03, Kro80, Hou22a]. Gaps [Mic59]. Garbage [Cow80]. Gasser [Mir99]. gatekeeper [San91]. Gateway [KO20, Sto09]. gather [Cra83c]. Gathering [BMRH05, Fie68, SSSC08, BK03, HHE97]. gatherings [Bak82]. Gauging [Mac99b]. GE [DC07]. gee [Bro88a]. Gehani [Cle03]. GEMS [MYV^{+}91, MYV92]. Gender [BW19, Bos92, Bra93, Cor90, IP02, dALT92, Lin99, RH92, TK04]. gender-neutral [Cor90]. Gendered [BT95]. General [MMW^{+}22, dGKNG06, Kil82]. Generalization [Luf72]. Generated [GL16, SDB^{+}13, WC15, SRVR97]. Generating [Dre83]. Generation [AZDA10, Joh10, Ols96, Str64, BBRT86, Bro86, JM01, Lip75, Mun02, SD06, TS81, JC21]. Generational [Shr20, Sta17]. Generative [BH21]. Generator [Bur69, PDK88]. Generic [YB17]. Genetics [Sta13]. Genre [Bha12, FMB18, Kil96, Lul06, Luz05, Orr99, Pal00, Sch10, SF99, SCW20, SWR21, Wol06, Yu23, vMPKG15, Baz99, Den96, SO98, Wal99]. Genres [Fer04b, Kry99]. Geoffrey [Sch78a]. Geographically [AG21, BSS13]. geometry [Tru91]. George [Ano71d, Arn69, Hou21b, Pic80, Rob78, Whi78d]. Georges [Bro83b]. geospatio [KVRR06]. geospatio-temporal [KVRR06]. Gerald [Den83c, Den83b, Har80b, Whi77c]. Gerard [Die72]. Gerre [Whi77d]. gestalt [Tru91]. Get [Ano15-29, Ano15-30, Bar09a, Nob10, Ros64, Ant79, Ber82, Bur84b, Van77, Wol82, Mal01]. Get-A-Member [Ano15-29, Ano15-30]. Getting [And83, Bai81a, Bos87a, Fer07c, JP86, Moo86a, Fla93, Pro94]. giant [Cox93]. Gibby [Hof80]. Gibson [Whi87]. Gilbert [Bea99]. Giles [Fer10]. Giner [Zet19]. Giroux [Mek72]. give [Sch80d]. Giving [Swa80]. Glass [HSA21, Cow80, Del92]. Global [Ano09x, Ano14n, Ano14o, Bat14, Bat18a, BP09, DPHF17, Dyr98, Gro99, LJ10, LRLM17, Mel08, Mel09, Mun05, PKLS17, PDL16, PDD^{+}24, Rav06, RWAK24, SJD20, St.04, St.06b, St.09, BTM12, LY98, PD05, St.02c, Suo13, UC01, YJ01, Ach24, DC07, Gro01]. Globalization [St.05b, Bal08, St.03]. Globalizing [Bar09b]. Globally [Ano19j]. Glossary [Mai70]. Glosses [Smi10]. Go [BN07, SAP20, Hun81, Sch81d, Tra92]. Goal [vdM07, Gla04]. Goal-Driven [vdM07]. Goal-Orientation [vdM07]. goal-orientedness [Gla04]. Goal-Setting [vdM07]. Goals [Fil06, FG84]. Gobbledegook [Fuj58, MKM81]. Godwin [Gan22]. goes [Pot00b]. Going [Den00, Lam14, Mel09, St.02c]. Goldsmith [Ano77a, Ano78d, Ano79b, Ano80c, Ano81c, Ano83c, Ano84b, Ano85d, Ano86c, Ano87b, Ano96, Ano97a]. Gonzales [Hou23]. Gonzalez [Bai24]. Good [Car90a, Car90b, Car59, FC79, Gol58, Reu65, AC85, Ano87k, Cox80, Dec84, Fei79, Mil81, Mur00d, Mur00g, Ulr84]. Goodman [Pet13]. Goswami [Ede87]. Governance [CC22a]. Government [Boi18, Ril00, Rou17b, Sus61, Ano77-34, BD84, CH89, Fre13, GB84b, Gre85, Lid75, Wha01, Fre80b]. grace [WR83]. Grading [LDKC23, Wol88, Zim88]. Graduate [AW22, DL24, KS14, LC17, LO07, PD22, ZM23, BD83, EA84, GHG91]. graduates [Kre00]. Graham [Bai24]. Grammar [KHV10, Nob10, Rav88, Man02, Ort78, Sch78a]. Grammatical [Kli79]. GRANDiose [FWB86]. Grant [Lan05]. Grantsmanship [Sch20]. graph [DV02]. Graphic [Fie83, Loe68, Sha82, Har90, Her95, Mus78b, Spe60]. Graphical [Roy07, St.02b]. Graphics [AM07, BP82, Hof05, Hof07, Lem83, Aic78, Fil82, HP88, Joe81b, Joe81c, LM97, Tru91, Win89]. Graphics-based [BP82]. graphing [Vol82]. Graphs [LO20, LS23, MP79]. Grasping [GH09]. Gray [Lip04, Ort75, Hun98]. Great [FDJW18, Flo04a, Mac68, Bro65, Mac75a, Mir99, Sha89, HvMK^{+}12]. Greater [EE78, GMG^{+}23]. Greeks [Pet59]. Greer [Hau80]. grind [Dob85]. Grissom [Pen98c]. Gross [Bra03a, Jan08, Wha04]. Group [All97, Ano58i, Ano59m, Ano59n, Ano59o, Ano61u, Ano61v, Ano61w, Ano62j, Ano63l, Ano63m, Ano63n, Ano64t, Ano64u, Ano77a, Ano77m, Ano77q, Ano77n, Ano77o, Ano77p, CL19, Fie65, GG13, PHW13, Ano77-28, Goo05, Heu75, LN03, RCS02, RLS06, Ano64s, Ano65g, Ano66g, Ano66h, Ano66i, Ano66j, Ano67j, Ano67k, Ano67o, Ano68h, Ano68i, Ano68j, Ano68k, Ano68l, Ano68m, Ano69j, Ano69k, Ano69l, Ano69m, Ano69n, Ano70i, Ano70j, Ano70k, Ano70l, Ano71n, Ano71o, Ano71p, Ano71q, Ano71r, Ano71s]. Groups [GG13, Ano73y, Bel99, KD00, Koc01, Koc03a, Kru00, Lin99, Loe95, RLS06]. Groupware [Wal97]. growing [GR86]. Growth [Emi99, Man99b, Rob95, Wil77b]. Grunig [Chi89]. gruntled [Ken82a]. Gruwell [Flo07f]. Guatemala [GLCA22]. Guest [Ano68g, Ano69i, Ano70h, GR86]. Guidance [HSA21, HVS^{+}22, MGB89]. Guide [Ano79n, Ano97d, Bar98a, Bra01, Bra05, Fie68, Flo06d, GL16, Hol03, Kie83, Kir20, Kry00b, Lam05, LJ10, Lin04, Mal07, Mar08, OM12, Pet13, Ram05b, Rav06, Rou19, She06, Sto05, Tus06, Van62, AN85, BM00, BSR85, DBK^{+}93, Eck03, HPSR83, Hor91, JP86, Kae03, Ker86, Kie82, LC79, Moo02, Nai81, NC81, PM98, Ram89, Rat80, Rom81, Smi03, Szo82, Van83, Whi78a, Whi78e, Dou01, Ano07k, Ano71d, Fri85, Ril02b, Flo07a, Sta99]. Guided [Bar22]. Guideline [NNB^{+}22]. Guidelines [Ano86u, Bak82, Bri99, Dun81, GCMQ17, Hap73, Hof05, McG86, Whi82a, WRTBRL23, Bli76, Eva00, Har90, HB94, SWB^{+}05, Zap80]. guides [Jon73]. guiding [Kat98b]. Guinea [KNMM09]. gulf [Luf76]. Gun [Ste22]. Gurak [Flo06d]. Gutenberg [Per70]. Gwiasda [Whi78e].

H [Ano80f, Cha23, DR10, Fer12, Hou21d, Lof08, Mic65, Pic80, Rav06, Rou17a, Whi79d, You10, Zho08]. H. [Gar19]. Haas [Joh97b]. Habermas [Con98]. habit [Con78]. Habits [Ano68r, JC21, FC79, Ril99]. hackers [Bas97]. Hagan [Mic61a]. Haimes [Zho08]. Haimes-Korn [Zho08]. Half [Zha21]. halftone [Koc75]. Halfway [Sam70]. Hall [Die72, Nor97]. Haller [Bra99c]. Hamilton [LD20]. Hand [LBB18]. Handbook [Ano97d, Arn69, GB84a, Bro83a, CG83, Eck03, Flo01b, Mic81, You10, Car00, Lee62, Mic61a, Den83c, Flo01d, Den83b, Hin10, Waa09, Whi77c]. handbooks [Mac99b, Mac03b]. Handling [BIdJ^{+}21, Car89]. Hanganu [Zha21]. Hanganu-Bresch [Zha21]. Hanks [Mus78a]. harassment [RH92]. harbor [Mar06]. Hard [CNC20, Ada80, Mir91, Nag87a]. Hardware [Boe76]. Harmon [Wha04]. harmony [Lay88]. Harms [TVRR20]. Harnessing [Fer09a]. Hartley [Car16]. Hartwig [Rei79]. Harty [Cha80b]. Harvill [Whi79g]. Hashizume [Rok19]. Hauptman [Smi10]. haven’t [Pro94]. Hayhoe [Daw10, Hou21b]. Haynes [Erv99b]. hazard [McC92]. Hazardous [Flo04d]. HCI [HLY22]. head [NB04]. headings [HMF02, WS92]. Headquarter [Car73]. Heads [Sam58]. Health [FS09, GCMQ17, JW22, RWAK24, RRW^{+}17, Rui20, WD09, Man98a]. Healthcare [Ach24, AZ17]. Healthy [Vic03]. Healy [Cuf19, Cuf20]. hear [Gar80a]. Heard [Bro64, Pro94]. hearing [Cra85b]. hearing-impaired [Cra85b]. Heath [You10]. Hedges [YJY24]. Hegewald [Rei79]. held [Dob75, Ano80g]. Helen [Pen98b]. Hello [Per70]. Help [Kle03, LDKC23, Ano73y, Bak82, Bar89, Bar79, Bar82b, CJ89, Cor92, Geo83, JCF92, PDG^{+}88, Som89, Spe60, vdM96]. Helpdesks [Rob18, vSdJ07]. Helping [KSNF15, Moo86b, Moo87c]. Helplines [vSdJ07]. helps [Mic87]. Henderson [Hou21a]. Hennig [Flo07b]. Henry [Car16, Pot01b]. Herbert [Kre86]. Here [Ter65, Dor83]. Herman [Ano71c, Har80a, Sch75b, Mic85b]. Herrick [McC19]. Herta [Fre80a]. Heuristic [Fri15, NNB^{+}22, Rif10, Kre88, ZZ05]. Heuristics [MFL17]. Hew [Mig83]. hidden [Fra75, NC81]. hierarchical [BRL91]. High [Cor69, DSV10, Gir06, Hen22, Jin18b, Lev68, Til06, BB91, Mun02, Sau93, Tsu91]. High- [Jin18b]. High-Quality [Cor69]. High-Stakes [DSV10, Til06]. high-tech [Mun02]. high-technology [Tsu91]. Higher [Cho21, PD22, Bro79, Spa14]. Highlights [Ano59i]. highly [Ril99]. Hill [Hou22b, Kie83, Tay01]. Hillier [Ano63i]. Hilton [Sch20]. Hiltz [Sal80]. Hinting [Mac05a]. Hirokawa [All97]. Historical [HM94, Mal07, Sko79, Bat02, CC88, Har89, Teb80]. History [Ano14n, Ano14o, Con91, Ram00, Smi10, HH82, Nye99, PM84]. Hobgoblins [Mek72]. hoc [BCR00, LMJK14]. Hofmann [Fer12]. Hofstede [KPV08]. Holcombe [Riv85]. Holographic [Mau79]. Home [AZ17, dGKGN11]. Homepage [Jon15]. Homepages [JD11]. Homophily [ZLCZ17]. Hong [Flo03b]. Hopefully [Ano78j]. hopes [Gro95a]. horizons [Ham96]. Horizontal [ZZL22]. Horn [Sch78c]. hospital [Cox93, JW22]. host [Mar85]. hostile [Car89]. hot [Ano85y, Wei98]. Houman [Suo13]. Hours [CC19, Cas84]. House [Man18, Cra92, DG75, Koc75, Wha83]. Household [LdJK20]. Howell [Her95]. HTML [Kle03]. Huettner [Ano07k]. Hughes [Daw10]. Hulme [Ril02a]. Human [BJ21b, Bar09b, Cle87, Dud61b, Dud61c, Fro06, Luf65, McC19, Nyc97, Phe20, Sal80, Sch81c, Sim22, Cus93, Fea77, Ham72, Kal97, Mas08]. Human-Centered [Sim22]. Human-computer [Nyc97, Cus93]. Humanitarian [BP09, HW09, MVV09, TVRR20]. humanities [Lin77b]. humanizing [HN93].

Humez [Mal11]. humility [Hav79]. Humor [Gle82, OL98]. Hundleby [Eri11]. Hunting [Bli84, SSSC08]. Hurricane [Cos23b]. Hurry [KI06]. Hybrid [AME21, LNB^{+}04]. Hybridization [Bha12]. hybrids [Bow73]. Hyman [Whi79i, Mic81]. hyperbole [Gro95a]. Hypertext [Reh98, BR92, DS89, LS94, Sel95, vW00, Gir97]. Hyphenate [Ock71, Ock75]. hypothetical [Dre88].

IBM [Fre68]. Ibrar [LD20]. icon [KR95]. icons [BRL91]. ICT [DGPR11, GE13]. Idea [Bat80, Wil80, Wil82, Ano87k, Tay82]. Ideas [Ano63u, Rin63, Cra83c, Rau79, Ros83, Riv85, Whi78a]. identification [Die04]. identifications [ST99a]. Identify [GAGW15, Kry99]. Identifying [PMGW21, Sou85a, WBB13]. Identity [AW17, ACW19, HZX24, SWR21, WARAL20]. ideologies [BT95]. IEEE [Ano77a, Ano78d, Ano79b, Ano80c, Ano81c, Ano83c, Ano84b, Ano85d, Ano86c, Ano87b, Ano96, Ano08u, Ano09v, Ano09w, Ano14h, Rou19, Ano63h, Ano63j, Ano63m, Ano63n, Ano63u, Ano64s, Ano64t, Ano64u, Ano65g, Ano65h, Ano66g, Ano66h, Ano66i, Ano66j, Ano67b, Ano67i, Ano67j, Ano67k, Ano68h, Ano68i, Ano68j, Ano68k, Ano68l, Ano68m, Ano69j, Ano69k, Ano69l, Ano69m, Ano69n, Ano70i, Ano70j, Ano70k, Ano70l, Ano71n, Ano71o, Ano71p, Ano71q, Ano71r, Ano71s, Ano73y, Ano75m, Ano77m, Ano77q, Ano77n, Ano77o, Ano77p, Ano77r, Ano78l, Ano78m, Ano78o, Ano78n, Ano79o, Ano80t, Ano80u, Ano80v, Ano80w, Ano80b, Ano81k, Ano81a, Ano81m, Ano81l, Ano82k, Ano82a, Ano82b, Ano83v, Ano83a, Ano84x, Ano85s, Ano85y]. IEEE [Ano85z, Ano86t, Ano86a, Ano87t, Ano04m, Ano04n, Ano04o, Ano04p, Ano04q, Ano05j, Ano05k, Ano05l, Ano05m, Ano05n, Ano06j, Ano06k, Ano06l, Ano06m, Ano06n, Ano06o, Ano06p, Ano06q, Ano06r, Ano07l, Ano07n, Ano07o, Ano07a, Ano07p, Ano07q, Ano07r, Ano07s, Ano08k, Ano08n, Ano08o, Ano08p, Ano08q, Ano08r, Ano08s, Ano09m, Ano09n, Ano09o, Ano09p, Ano09q, Ano09r, Ano09s, Ano09u, Ano10m, Ano10n, Ano10o, Ano10p, Ano10q, Ano10r, Ano10s, Ano11e, Ano11f, Ano12o, Ano12p, Ano12a, Ano12b, Ano12q, Ano12r, Ano12s, Ano12t, Ano12u, Ano12v, Ano12w, Ano12x, Ano12y, Ano12z, Ano12-27, Ano13a, Ano13b, Ano13n, Ano13o, Ano13p, Ano13q, Ano13r, Ano13s, Ano13t, Ano13u, Ano14n, Ano14o, Ano14p, Ano14q, Ano14r]. IEEE [Ano14s, Ano14t, Ano14u, Ano14v, Ano14w, Ano15z, Ano15-27, Ano15p, Ano15q, Ano15r, Ano15s, Ano15t, Ano15u, Ano15v, Ano15w, Ano15x, Ano15y, Ano16t, Ano16u, Ano16h, Ano16n, Ano16i, Ano16j, Ano16o, Ano16k, Ano16l, Ano16p, Ano16q, Ano16m, Ano16r, Ano16s, Ano17y, Ano17z, Ano17-27, Ano17-28, Ano17m, Ano17q, Ano17r, Ano17n, Ano17s, Ano17t, Ano17o, Ano17u, Ano17v, Ano17p, Ano17w, Ano17x, Ano18v, Ano18w, Ano18j, Ano18n, Ano18o, Ano18k, Ano18p, Ano18q, Ano18l, Ano18r, Ano18s, Ano18m, Ano18t, Ano18u, Ano19a, Ano19d, Ano19b, Ano19e, Ano19c, Ano19f, Ano20c, Ano20h, Ano20d, Ano20i, Ano20e, Ano20j, Ano20f, Ano20g, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano22c]. IEEE [Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano24i, Bos64, Buc74, Bun71, CFMM14, Car73, Gan77, Gui89a, Ryd85, Wer71, Wil64, Ano63j, Ano63s, Ano70b, Ano71a, Ano73a, Ano74a, Ano75a, Ano76a, Ano78a, Ano79p, Ano79q, Ano79r, Ano79a, Ano80x, Ano80y, Ano80z, Ano80-27, Ano80a, Ano81n, Ano81o, Ano81p, Ano81b, Ano82l, Ano82m, Ano82n, Ano82c, Ano83w, Ano83x, Ano83y, Ano83b, Ano84y, Ano84z, Ano84a, Ano85t, Ano85a, Ano86u, Ano86b, Ano86l, Ano87a, Ano97c, Ano00, Ano04f, Ano04b, Ano04g, Ano04c, Ano04h, Ano04d, Ano04i, Ano04e, Ano05f, Ano05b, Ano05g, Ano05c, Ano05h, Ano05d]. IEEE [Ano05i, Ano05e, Ano06f, Ano06b, Ano06g, Ano06c, Ano06h, Ano06d, Ano06i, Ano06e, Ano07b, Ano07g, Ano07c, Ano07h, Ano07d, Ano07i, Ano07e, Ano07j, Ano07f, Ano08a, Ano08f, Ano08b, Ano08g, Ano08c, Ano08h, Ano08d, Ano08i, Ano08e, Ano09a, Ano09f, Ano09b, Ano09g, Ano09c, Ano09h, Ano09d, Ano09i, Ano09e, Ano10a, Ano10f, Ano10b, Ano10g, Ano10c, Ano10h, Ano10d, Ano10i, Ano10e, Ano11b, Ano11a, Ano12c, Ano12h, Ano12f, Ano12i, Ano12g, Ano12j, Ano12k, Ano13c, Ano13f, Ano13g, Ano13h, Ano13i, Ano14a, Ano14d, Ano14e, Ano14f, Ano14g, Ano15a, Ano15h, Ano15d, Ano15i, Ano15e, Ano15f, Ano15g, Ano16a, Ano17a, Ano18a, Ano20a, Ano22a, Ano22b, Ano23a, Ano23b, Ano23f, Ano23c, Ano23d, Ano23e, Ano24a]. IEEE [Ano24b, Ano24c, Bos64, BB75, BG07, Fry87]. IEEE-PC [Ano10m, Ano10n]. IEEE-TPC [Ano09s]. IEL [Ano06j, Ano06k, Ano07l, Ano08k, Ano09m, Ano09n]. If [Get86, Nob10, Oli58, Shu93, Bur84b, Wal97]. iFixit [GL16]. ignored [Ful80]. II [Car90b, Dud61c, Kat98b, MYV92, Sha94b, Whi77a]. III [Rho77, Som89]. Illustrated [BKSM82, Ano87y, BT95, Vog83]. Illustrating [ML78, Gre85, Whi78d]. illustration [Koc75, Hof80, Whi79f]. Illustrations [Roc59, Wal98, Mag77, RDRS04]. Illustruction [Con87]. Ilpo [Zim13]. I’m [Flo07e, Ano86-27, Nag86, San87b]. Image [DF70, Ama05, Bra03b, DF75, Lin80]. Images [AFH20, Lay88, Moe98, TL92]. Imagination [JW22, Ano14m, Ano15n, Ano15o, Gud82, Shu93]. Imitation [JF03]. immediacy [RH92]. immediate [Met73]. Immigrant [RRW^{+}17]. imminent [Sau92]. Impact [AZDA10, Bal18, CFMM14, Gra22, Jin18b, JW17, KCJC08, Kun73, MYT16, Moo16, RD16, Sut77, TJQN11, WFSC^{+}23, YTY12, Gri94, Lin99, LR85, LNCL05, Per96, RLS06, Woe80]. Impacting [ASHB16]. Impacts [PEG19]. Impaired [BNB21, Cra85b, Dur99]. Impartion [Ano63o]. impediments [Vis77]. Imperative [St.04, Dav93, Hou83, Mar93]. imperatives [Gor90]. imperfect [Fra81]. Implementation [And14, Cog15, LCX12, Wil64, RH01]. Implementing [Cho18, Bul72, LNB^{+}04]. Implications [CC19, CC22a, FMB18, Koc77, LL09a, LL09b, Lin72, Luf64, Tal77, Tha22, Tha24, Yu12, Cro95, DRG^{+}02, Gan04, Gri91, She73, WB92, vLUA02]. Importance [Gol58, MFL17, St.01b, GR86, Lin80, Swa05]. Important [Ren96]. Impress [Ben79]. Impression [FR13]. Improve [BH16, Coh69, KS14, KO20, LDKC23, Pit79, Web79, WP18, Gle82, LN03, Som89, St.00b]. Improved [Wes71, MP79, Pot01b]. Improvement [Ano63i, Ano63v, Ano63w, Ano64k, Mil64, Und67, Dul93, Hor89, Koc01, Mey64]. Improving [Ano67i, Cos23b, DGPR10, FC96, Fur90, GDTZ09, Hat66, Sei67, War88, WJ13, Coh60, Har81, Har84, Woe79, Per70]. In-Group [PHW13]. in-house [Cra92, Koc75, Wha83]. in-plant [Jen66]. inadvertent [Pad78]. inanimate [Sch81c]. Inarticulate [McK66]. Incentive [Joe79b]. Incident [WBB13, YFW15]. Inclusion [Bay22, BH21, Car24, Wal21]. Inclusive [Ach24, Wil24]. Income [Gei75, Har75b, Bak75, Van75]. Incomplete [Joh20]. Incorporating [Bae12]. Increase [Kil13]. Increasing [Con87, EA84, HBS08, Pal75]. incremental [Mic74]. incumbent [Wha85]. independent [Kre87b, Van79]. Index [And97, Ano61n, Ano71a, Ano73a, Ano75a, Ano76a, Ano78a, Ano79a, Ano80a, Ano81b, Ano82c, Ano83b, Ano84a, Ano85a, Ano86b, Ano87a, Ano97c, Ano00, Ano03a, Ano04a, Ano05a, Ano06a, Ano07b, Ano08a, Ano09a, Ano10a, Ano12c, Ano13c, Ano14a, Ano15a, Ano16a, Ano17a, Ano18a, Ano20a, Ano22a, Ano23a, The97, vdM97a, Ano61b, Ano99a, Ano99b, Ano01a, Ano01b, Ano02a, Ano02b, Ano03b, Mek72, Van63, Ver80, Ano64p]. indexed [Sno64]. Indexes [Ano64-28, Tuk62]. Indexing [Ano65i, Dio82, GH70, Ken64b, McC63, Sam70, Smi10, App01b, Mur00e]. India [Raj12, SM66]. Indian [Bed07, Shr20]. Indigenous [GLCA22]. Indirect [YK13, Mac05a, Wis91]. Individual [TJQN11, PY04]. inducting [MSC96]. inductively [FS00, SF02]. Industrial [Con70, HSA21, HVS^{+}22, Mic65, Vch94, Dou72, Jam84]. industries [Tsu91, Car16]. Industry [Bar82a, Bat18a, Boo17, Bro62, CCL24, Cos23a, Cro59, EBAS18, Far62, LR24, LLY18, Mar08, Yu10, Ano71d, BD84, CM05, GG84, Gor90, LS84, Mau79]. industry-oriented [Gor90]. Influence [HvMK^{+}12, HK07, KCLG13, Lof08, WPSB13, YJY24, ZLCZ17, vHK08, IP02, Koc05, SJ99, vW00]. Influenced [LO20, LS23]. Influences [Mun05, Ock05, Red10, Rei02]. Influencing [KNMM09, TYT91, XS24, ZB15, Par87]. info [NLK10]. Inform [AW22, Joh82]. Informal [Kei67, Hol74]. Information [Ano61o, Ano61p, Ano61q, Ano62h, Ano63p, Ano63q, Ano64v, Ano64w, Ano64x, Ano65j, Ano65k, Ano66k, Ano66l, Ano67l, Ano68n, Ano68o, Ano68p, Ano69o, Ano69p, Ano69q, Ano70m, Ano70n, Ano71t, Ano71u, Ano71v, Ano79p, Ano79q, Ano79r, Ano80x, Ano80y, Ano80z, Ano80-27, Ano81n, Ano81o, Ano81p, Ano82l, Ano82m, Ano82n, Ano83w, Ano83x, Ano83y, Ano84y, Ano84z, Ano85t, Ano86u, Ano87u, Ano04n, Ano04o, Ano04p, Ano04q, Ano04b, Ano04c, Ano04d, Ano04e, Ano05k, Ano05l, Ano05m, Ano05n, Ano05b, Ano05c, Ano05d, Ano05e, Ano06o, Ano06p, Ano06q, Ano06r, Ano06b, Ano06c, Ano06d, Ano06e, Ano07p, Ano07q, Ano07r, Ano07s, Ano07c, Ano07d, Ano07e, Ano07f, Ano08o, Ano08p, Ano08q, Ano08r, Ano08b, Ano08c, Ano08d]. Information [Ano09o, Ano09p, Ano09q, Ano09r, Ano09b, Ano09c, Ano09d, Ano09e, Ano10o, Ano10p, Ano10q, Ano10r, Ano10b, Ano10c, Ano10d, Ano10e, Ano11e, Ano11a, Ano12r, Ano12s, Ano12t, Ano12u, Ano12v, Ano12w, Ano13n, Ano13o, Ano13p, Ano13q, Ano13r, Ano13s, Ano13t, Ano13u, Ano14p, Ano14q, Ano14r, Ano14s, Ano14t, Ano14u, Ano14v, Ano14w, Ano15p, Ano15q, Ano15r, Ano15s, Ano15t, Ano15u, Ano15v, Ano15w, Ano16n, Ano16o, Ano16p, Ano16q, Ano16r, Ano16s, Ano17q, Ano17r, Ano17s, Ano17t, Ano17u, Ano17v, Ano17w, Ano17x, Ano18n, Ano18o, Ano18p, Ano18q, Ano18r, Ano18s, Ano18t, Ano18u, Ano19g, Ano19h, Ano19i, Ano19d, Ano19e, Ano19f, Ano20k, Ano20l, Ano20m, Ano20n, Ano20h, Ano20i, Ano20j, Ano21i]. Information [Ano21j, Ano21k, Ano21l, Ano22l, Ano22m, Ano22g, Ano22h, Ano22i, Ano22j, Ano22b, Ano22k, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano23b, Ano23f, Ano23c, Ano23d, Ano23e, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano24i, Ano24a, Ano24b, Ano24c, Bat14, Bha77, Bis70, Bot70, CR06, CBSM^{+}09, Cha06, Coi69, DM63, Dud61a, DJ82, DR10, Fer04b, Fie68, Fil06, FS09, Gar06, GYK14, GH70, Hac16, HW09, HSA21, Hof07, HK22, HK70, JJ10, Jos70, Jos75, Kae04, Kel61b, Key87, KPV08, Lie18, LHAL23, LE18, Mac68, Mac75a, Mac05b, Mai70, Mar70, McC63, Men23, Mic59, MKBF09, OM12, Pav94, Per63, Ram07, Reg70, Rin59, RRW^{+}17, Sab61, Sal68, Sch66, SA23, Sno64, SU09, TM15, Tom69, Web65, Wil64, WJ13]. Information [WH82, XZZ10, YK13, vHK08, Ano65i, Ano65n, Ano77m, Ano77q, Ano77n, Ano77o, Ano77p, Ano77r, Ano77-28, Ano78l, Ano78m, Ano78o, Ano78n, Ano85z, Ano04f, Ano04g, Ano04h, Ano04i, Ano05f, Ano05g, Ano05h, Ano05i, Ano06f, Ano06g, Ano06h, Ano06i, Ano07g, Ano07h, Ano07i, Ano07j, Ano08f, Ano08g, Ano08h, Ano08i, Ano08e, Ano09f, Ano09g, Ano09h, Ano09i, Ano10f, Ano10g, Ano10h, Ano10i, Ano11b, Ano12h, Ano12f, Ano12i, Ano12g, Ano12j, Ano12k, Ano13f, Ano13g, Ano13h, Ano13i, Ano14d, Ano14e, Ano14f, Ano14g, Ano15d, Ano15e, Ano15f, Ano15g, BBRT86, Bis96, BB91, BMRH05, BK03, Car90c, Cro95, DM81, EM78, Fes62, FC96, Fis98, Fis99, Flo01b, Foy96, Fry62, GKD77, Gol77, Gri91, Gro95a, HH93, Hol74, Hou83]. information [HMF02, IPC01, JVT01, Joe79a, Joe80a, Joe81b, Joe81c, JE95, KS04, Ken64a, KA03, Kos95, LN00, Men72, O’C62, Pot01b, Pri97, Que86, Rab92, RLG97, Rid87, Row75, Rub86b, Rub87, SSM74, Sch78c, Sch79, SD06, Sha89, She73, ST05a, Sko81, SB95, TDS95, Til79, TYT91, Vch94, WM87, Wei75, Wei98, Win90, Woe79, Yer74, And97, Ram03, Fil06, Kat84, Mac02b, Mur00c]. Information-Management [SU09]. Information-Retrieval [Mai70, Per63]. Information-Seeking [LE18]. informational [Sou85a]. Informative [Mog14, Ken64a]. Informativeness [MY14]. Informed [Izq24, SGM^{+}24]. infrastructure [Rob95, Web79]. Ingle [Flo07d]. Inglis [Cap93]. Ingroup [PHP16]. inherited [dALT92]. Initial [KL12b]. initiated [FS92]. initiating [MM00]. initiation [Dan77]. initiatives [Gal01]. Injustice [AD22]. inmates [Mur00f]. innovate [Joe79b]. Innovation [CLP08, DV09, Dia14, RD16, MOA02, Ste62, UC01, Whi79a]. Innovations [Par75, ZB15]. Innovative [Flo05b, BMRH05]. Input [Kun69, Var77]. Inquiry [MCC62, Mog09, Rud92]. Insertion [RDJM09]. Insider [Dre17, Flo04c, Ley08]. Insights [Ach24, Orr06, vMPKG15]. inspection [Cla79]. Instant [AZDA10, GTTX08, Hav79, HF08, HDKC08, Sta08, Ste08, Han80, Pas03, Zho05, ZZ05]. instinct [Man02]. Institute [Mic65]. Institutional [Ano58g, Ano59j, Ano59k, Ano59l, Ano61r, Ano61s, Ano61t, Wil10a]. Instruction [JD23, LKS13, Mil62, YTY12, Zem21, BP82, Bar90, Cha02, FBE^{+}03, Gol82a, Hor84, LS84, Mud04, Par87]. Instructional [MAK15, vdM97a, FG84, GW84, Har81]. Instructions [LdJK20, Mac05b, Roy07, TM15, vHK08, vdM07, Bag83, BS98, CH89, Cle87, Gan04, HH82, KS04, Pil81, Joh17]. instructive [ML99]. Instructor [WT09, CM02, DH04]. instructor-led [CM02]. Instructors [BJT^{+}19, CCL24]. instrumentation [Lyn83]. instruments [UN00]. Insurance [LLY18, RRW^{+}17]. Integrated [AZ17, Fer04c, JD23, Pla71, MYV^{+}91, MYV92, Pat08]. Integrating [BJ86, BS73, BH21, DL24, FC24, KC81, Met73, Mog14, NH05, NLK10, Yu11, ZLCZ17, ZFWX21]. Integration [HS18, Bar90, Mey87]. Integrative [AB15a, Bat14, Cla16, PD24, SV24, SCW20, Zol14]. Intel [SH93, Wil97]. Intellectual [Dia07, Flo05c, Gar06, Wis15, Bal08, SR83b]. Intelligence [Dui96, LRLM17]. Intelligent [Hou21c, Hul81]. intelligently [Hou83]. intensity [Pen99a]. Intention [ASHB16, LCX12]. Interact [SKZ23, Gla04, Har97]. Interacting [Rei02]. Interaction [CS06, Fer04a, Fer05b, FVB16, Gar19, Hel07, Ols96, RWAK24, vdHdJS06, Pas03, ST06a, Tru88, Nyc97]. interactionism [Ama03b]. Interactions [HLY22, Mac05a, Mac06]. Interactive [MP20, BP82, Cro95, Gol97, Jen02, Kru00, MMCL00, MCWK77]. interactive-graphics-system [BP82]. Interactivity [RD16, Zol14, ZB15]. Interchange [SP87]. Interconnecting [Noa03]. Intercultural [BESFW22, Lon14, Raj12, Sap08, VNC^{+}13, Wal18, Yu11, Yu12, Gol03, Yu24, Suo13]. interdependences [Kin77]. Interdisciplinary [WP18, FBE^{+}03, Goo05]. Interest [Ano63z]. Interface [SS10, Wei02b, Cas86, KA03, Koe77, Mar82, Ril02a, Ste04, SH91, Str99, Vel90]. Interfaces [Lan04, Har90, Vel90]. internal [KVRR06]. Internalization [SS10]. International [Ano60a, Ano67a, Ano85y, Ano07o, Bar09b, Bos87a, Bre10, DH92, Flo07b, Hac16, Heu75, Hov14, Joh10, MEZ12, PH24, Rav06, Sap08, St.07, Wer71, WARAL16, WARAL20, Wor87, Yua21, ZWX20, Ama01, Ama03b, Ama05, BPK94, EA84, Fre13, Gro01, LY98, MSC96, St.99b, St.01a, Suo13, ULK01, Wha01, ZVS01, Dob75, The97, Ril01]. Internationalising [Hof07]. Internet [Bas97, Bra01, Car00, CBSM^{+}09, GG96, Rav06, St.09, Cry02, CC22a, Flo03c, LN03, SCW20, SWB^{+}05, ZZL22, Mog09]. Internet-based [LN03, SWB^{+}05]. Internet-Mediated [SCW20]. interns [BD83]. internship [RC89]. Interpersonal [Ram05b, ZFWX21, Pot84, Smi85]. Interplay [Ock07]. interpret [Gal01]. interpretation [Luf72, SF99]. Interpreting [DL24, Joe81a, Joe82a, MML15]. interprofessional [LK72]. Intersectional [MSDF22]. intersubjective [McG72]. Intertwined [Red10]. Interventions [DR20, RWAK24]. Interview [Bar09a, Rea80]. interviewers [Car90a]. Interviewing [Wal86, Car90a, Car90b, Fla93]. Interviews [Fie68, vMPKG15, Boy81, Kle81]. intranet [Ama03a, Boc98, RH01]. intranets [RH01]. Introducing [Ano15z, Ano15-27, Ano16t, Ano16u, Ano17y, Ano17z, Ano17-27, Ano17-28, Ano18v, Ano18w, Ant18, KRM24, May19, Yu24]. Introduction [AB15b, Ano77y, BA16, Baz99, BI18, BTM12, CS06, ES99, Goo95, GZ97, Hil85b, Hou21d, IPC01, KHH^{+}08, LBE23, Lan94, MMW17, MR21, Orr03, Orr06, PM08, Pri97, Spi16, St.07, Sta08, UVB02, WS04, vdGRRv13, CM01, FD84, Gro05, HP86, Joe79b, Mal85b, PM04, Ste04, SP86, SB00, UC01, Cuf19, Ram03, Cuf20, Hyd18]. Introductions [Kan12, KT18, Ant99, CW90]. Introductory [Cho18, GG96, JF03]. Introverts [LP16]. Intuition [Fer09b]. inventing [Ano80j]. Invention [Mel09, Cot79, Foy96]. inventions [Hau79, Kim79]. inventor [Van79, Zim00]. investigate [McC94]. Investigation [ASHB16, DV09, GTTX08, KMR14, LE18, MKBF09, TJQN11, WHC^{+}12, XS24, vDVvL19, Her95, Koc01, RCS02, Zho05]. investigations [Dom91]. investigator [Hor75]. Investigators [Sch20]. Investor [KdJH14]. Invisible [Man98c, MB84]. IPCC [Ano07o, Ano04j]. iPhone [ZFWX21]. IRE [Ano58i, Ano59m, Ano59n, Ano59o, Ano61u, Ano61v, Ano61w, Ano62j, Cot59, Gan58, Pat59]. IRE-PGEWS [Pat59]. Ireland [Cle16]. Irish [Den98a, FCS10]. Irreproducible [Sch73a]. irresistible [Lin93]. Irvin [Whi78b]. Irving [Col65]. Isabel [Hou22b]. ISI(R) [GKD77]. island [Ano86-27]. Issue [AB15b, Ano63a, Ano63b, Ano69t, Ano13v, Ano15h, Ano15i, BA16, BI18, CS06, DH94, MMW17, MR21, Orr06, PM08, Spi16, Sta08, WS04, Ano64a, Ano77-30, Ano82e, Ano83l, Ano04s, Ano09x, Bra82a, DF87, ES99, Gro05, Hil85b, IPC01, Jen62, Joe80a, Joe81b, Joe81c, Joe81a, Joe82a, Joe82b, Joe83a, Mar06, Orr03, Pri97, Ste04, SB00, UVB02]. Issues [CS06, Goo96, SFM09, Wal87, WPSB13, Yea09, Bra93, CR73, DH92, GBF05, Goo95, Jen00a]. ISTE [vdM97b]. Italian [YB17, Yu20, Yu23]. iterated [HT77]. iteration [Koc03a]. iterative [Die89]. Ito [Fer07d]. I’ve [McN58a].

J [Ano64f, Ano80h, Bal18, Cha80b, Cha23, Cun10, Den83c, Den83b, DR10, Eri11, Fil06, Flo06d, Flo07b, Flo07e, Flo08, Gar19, Har80b, Hau80, Hof80, Kno80, LJ10, LK08, McE18, Nob10, Sch78b, Sto08, Whi77c, Xu07, Xu08, Zho08]. J. [Chi89]. Jackson [Yu24]. James [Ano80f, Flo03b, Whi78a, Whi79h, Bar98a, McC19, Saw99]. Jan [Sch78a, Ama00b]. Jane [Yu24]. Janet [Bra00]. Jansen [Wei02b]. Japanese [Fer07d, AFH20, Den88, Flo05e, FS99, FS00, Fur90, MG95, McC06, NKU03, Shi81, Shi82, SF02]. Japanese-English [Den88]. Jargon [Goe78]. Jason [Ana21, Flo03b, Joh20]. Je [Cam07]. Jean [Bea99, Cod09]. Jean-luc [Cod09]. Jefferson [Sch80b]. Jennifer [Dur22]. Jests [Cos78]. Jewel [Cle03]. Jimmie [And97]. Jo [Bra99a]. Joanna [Cha80a, GG22a]. Job [Bar09a, Bli84, Gol82a, Joh63, LLY18, vMPKG15, Cro86, Her64, JP86, Kat98a, Sch93, Vel88, Bro83b, EC01]. job-related [Sch93]. Jobs [GLL23]. Johan [Zim13]. John [Bro79, Fil05, Flo06d, Gon20, Kre87a, St.09, Zim13, Car16, DOHM96, FW90]. Johnny [Pen98c, Sch75c]. Johnson [Cun10, Dur22, Ano64j, Her64, Kni86]. Jon [Flo20]. Jonathan [St.09]. Jones [CD98, Mag03, Sta20, Whi77d]. Joseph [Wha04]. Joshua [Bar22]. Journal [Ano63j, FWJ77, Lie77, Lin77b, Wil77b, Axe73, Bow73, Cov75, Dan77, DH73, HTB80, Heu75, Kin77, Kun73, Lid75, MM73, Mir75, Pat76, Sac77, San91, Sar73, Shi75, Sta73, Ter73, Tho75, Tho73, Tho01, Urb77, Van75, Weh77, Wit77]. Journal-related [Lie77]. journalism [Gir60]. journalists [War80]. Journals [Ano75m, BL13, CFMM14, Hov08, Kun69, LHMN07, Smi20, Abe73, Abr77, Bal75, Bov73, Bra82c, Bro75, Cap78a, Dav85, DG75, Est82, Gar73, Gei77, HT77, Hoo77, Max73a, Max73b, Met73, Pal75, Pul86, Rus77, She73, Sub76, Sus85]. Joyce [Hou21d]. Jr. [Hau80]. Juan [Flo20].

Juárez [GLCA22]. judge [Bau62].

Judgment [Wee19]. Judith [Riv85, RA97]. Judson [Fre80b, Sch78d]. jumble [Von78]. June [Ano09u]. Jürgen [Con98]. Just [DR20, GC22, Win73, Ano87k, FHR87a, Gra13]. Justice [Ath22, BESFW22, Ben22, Car24, Che17, CC22b, Dur22, Flo20, Gan22, Gra22, JW17, PD22, Rea22, Sim22, Sta20, Wal23, Wel22, Cos78].

K. [Cuf19]. Kai [Rui20]. Kai-Yu [Rui20]. Kairos [Rif10]. Kalmbach [Saw99]. Karen [Mag03, Spy98]. Karin [LD20]. Karl [Whi78e]. Katharina [Rei79]. Kathleen [Sch78d]. Katz [Ber12]. keep [Hil89, Hug79]. Keeping [Tuk62]. Kelleen [Zha21]. Kelly [Sto15]. Ken [Pen98d]. Keven [Cha80b]. key [Bos87b, Bro84, Eva60, Hol74, MD84, Ode86, SSNJ05, SB91]. Keynote [Mey69, Buc74]. Keys [Car59, PP97]. Keywords [Cha23]. Kieran [Cuf20]. Killingsworth [And97]. Kim [Pen99b]. kind [RC89, Spi94]. Kinney [ES98]. Kirkman [Bro79]. Kitalong [Wis15]. KIVA [AA93]. Klischewski [Fer03]. Kmiec [Rou19]. Know [Luf69b, Wil10a, Hin72, Jen99a, Kru97, Luf75, Moo86a, Ste99]. Knowing [Fer09a]. Knowledge [BM00, BWPH12, DW07, Fer09a, For04, Fro06, HS18, HK07, Kil96, KCJC08, LD20, Mas08, PS12, SSNJ05, Sch70, SS10, SEN10, Yea10, App02, Bae12, DJ82, Fre13, Kru00, Lin72, MYV^{+}91, MYV92, Moo87c, Pal00, Pot00a, QK05, RH01, Ste04, Sch75d, Wal86, Zim99]. knowledge-producing [Kru00]. Knowledges [Ric09]. Known [ZJG17, Her95]. Kock [Ano80j]. Koestler [Ano80i]. Kohl [LJ10]. Koltay [Pen98h]. Korea [Rob96]. Korean [KS14, Yua21]. Kores [Her95]. Korn [Zho08]. Koskinen [Zim13]. Kostek [Pen98e]. Kostelnick [Cla12]. Kraft [Whi79g]. Kristen [Sta20]. Kudos [Moo87d]. Kukulska [Ril02a].

Kukulska-Hulme [Ril02a]. Kuniavsky [Pet13]. Kurosu [Rok19]. Kurt [Mus78a].

L [Bar98a, Den83c, Flo18, Hyd18, Mir99, Sal61, Tra12, Whi78a, Whi79g, Whi79h, You10]. Lab [KO20, Zim13, Bur84a, Bur78, Wal99, Coc04]. Labeling [MP79]. labels [RCB82, Vel87a]. laboratories [Jam84, Ano63i]. Laboratory [DM63, KO20, McC63, Sha61, Hug79]. Labs [Cle03]. Lacour [Whi78b]. Lancaster [Joe79a]. Landay [Flo03b]. Landmark [SV24]. Landscape [PM08]. Lane [Mag03]. Language [ACW19, AM07, BR68, Bro78, Car24, Che17, Cla12, Cli63, Cry02, Cun10, DSV10, Dre16, Dre17, DuB08, Fri19, Gou60, GCMQ17, JW17, LS61, Löc22, MMW17, Nob10, Ril00, RC59, Ros15, Sch63, Sch17, Sim22, ZWX20, vMPKG15, BJ79, Cra85a, Cro86, Dur99, Edd83, Fis84, FO06, Hal80, Kan78, Lip75, Mat77, MG82, McD94, MSC96, Nai81, Orr01, Orr03, Pen99a, Pri99, RL78, St.99b, TCW06, Whi82b, Woo90, Yu24, Ril02a]. Language-Analysis [Cli63]. language-impaired [Dur99]. languages [App01b, Luf65]. Lannon [Flo06d, Xu07]. Lap [HS79]. Lap-dissolve [HS79]. large [App01a, CH89, Cox93, SH91]. large-scale [App01a]. Larry [Mus78a]. Larson [Sch83]. last [FHR87a]. lately [Ano04k, Ano04l, Ano07m, Ano11d, Ver80]. Latent [KI06]. Later [Roc59]. Latest [SV24]. Lathrop [Whi78b]. Latin [Ano70b]. Latinx [PD22]. Launch [ZFWX21]. Launching [Rub66]. Laura [Hou23, Flo06d]. Lavishly [Ano87y]. Law [Flo05c, Ram07, Row78, Ste22, Ano77y, Cha81, Gan77, Koc77, Lie77, SS87, Tal77, Woo90, Zur75, Sub79]. Lawrence [Joh97b, Whi79g, FP79]. laws [Dou02b, PM04]. Lay [DGPR10, Cox80, Kan85]. Layout [Hur83, HTB80]. Lead [Cor83]. Leader [CL19, Flo04b, KDBC20, ZC24, Kol93, Ser88, SV06]. Leaders [SSS09]. Leadership [Ano63d, FCS10, Flo06c, KDBC20, Ram05a, SC12, Yu23, KD00, Ker86, Web79]. Leading [Ano08t, Ano09t, Kol93, Phi69b, St.02a]. Leads [Cho21]. lean [Koc01]. Leann [Pen98e]. learn [Kat98b, Kat98a, Mur00g, Red89]. Learned [VG16, WT09, Koc03a, MSW97]. learner [Gla04]. learners [Cra85a]. Learning [Ama04b, BC17, BSS13, BP15, BMS^{+}15, DW07, Hof05, HK07, KD84, KL12b, KSNF15, KCLG13, Lav06, Lie18, MP20, McC06, MFL17, MMVH10, SA23, SHB08, SY09, SAP20, WFSC^{+}23, ZWE21, vHK08, vdM97a, AL03, Bae12, CS96, Cod04, DH04, Del00, Dum96, GRS96, JDW96, Kre87b, LNB^{+}04, MMCL00, Mus78a, Par06, RKP00, Sou88, Ste99, VVL^{+}01, Woo81, Kat98a, Fer09a]. Least [LG14]. leaves [Gud82]. lectern [Kre88, Mon80a]. lecture [EE78, HS79, Jin86, Kuo93, ML78]. lecturer [Ano78p, FP79]. Lectures [Gra13, Rei79]. led [CM02]. Lee [Dur22, Ede87]. Leech [Sch78a]. Lefèvre [Fer06]. Legal [BH21, DF87, Flo05d, OL98, Pil87, Rin59, SFM09, Bro89, CEH87, Jen00a, Mar02, McC92, Ros94, SS87]. legalese [Ano78e, Sie78]. legible [Gou73]. legislation [Lin77a]. Legitimating [Lin19]. Lenart [Bar22]. length [HMF02]. Lesley [Pen98f]. Lessons [BS16, Bro65, Cho18, MSW97, Spi96, VG16, WT09, Dom91, Koc03a, Tra83, WR83, Dom95]. Let [Gar80a, Ker80, Sam58, Sch80d]. Letter [Ano64p, Ano71w, Bau62, Bro84, Loe68, Sch10, Fox81, Sal62, SO98, Ama00a]. Letters [Ano73w, Ano73x, HZX24, KI06, Tra74, Ega81, GB79, Mal01]. Leukefeld [Sch20]. Level [KZCR10, Rac59, Bri99, Gla04, PFM88, Ste99, TCW06]. levels [ACH^{+}98, Bar80, Bar81, GBF05]. leveraging [Par06]. Lewis [LK08]. lexical [BW90]. Leydens [Flo20]. Li [Car16]. Liability [WM87, Cle87, Con87, McC92, Woo90]. Librarian [Web75]. Librarians [Sor75]. Libraries [Nic59, Ano77-29, Fra75, Tal77]. Library [Ano12o, Ano12p, Ken64b, Sha61, Wil75, Bar98b, Ede77, Koe77, Tal77, Ano12x, Ano12y, Ano12z, Ano12-27, Ano15x, Ano15y]. library-publisher [Koe77]. Lie [GMT08]. Life [Ano71l, Cle03, Hov08, EC01, Fer07d]. lift [Sie78]. Light [ODF92]. Lightweight [Hou21c]. Like [Lay97, dGKGN11, Arm83, Fei79, Hin72]. Limerick [Wal23]. limitations [Red81]. Lincoln [Sch65]. Line [Flo07e, LS23, MM68, Ano85y, HHE97, Koc75, Lip75, RLG97]. linear [GR94]. Linearization [Mor66]. lingo [Ros79]. Lingua [Räi20]. linguist [Ken82a]. Linguistic [Har90, Mis05, Mac04b, Dur22]. Linguistics [Orr06, Con99]. Link [Dud61b, Dud61c, Gri81]. Linkage [BJ21b]. linkages [MOA02]. LinkedIn [BP15]. linking [Mar02]. links [PDK88]. Lisa [GG22a]. List [Ano64i, Loe68, Tel68]. listen [Pro94, Woo81]. listener [Hay80, JS81]. Listening [Bai81a, Con70, LL09b, Bah85, Con78, Fer07b]. listens [AB81]. listing [Ano87-28]. Listings [Cor69, Ano58g, Ano59j, Ano59k, Ano59l, Ano61r, Ano61s, Ano61t]. Lists [Car87, Smi10, BS98]. Literacies [Win97, BTM12]. Literacy [Gir06, GE13, GCMQ17, Rea22, Ree64, St.01b, Den96, Gri91, Mur00h, Joh97b]. Literature [AB15a, Ano64-28, Bar22, Bat14, Bot70, Bou62, Cla16, KK02, Mic68, PD24, SV24, SCW20, Van62, Zol14, Ede77, Hil89, Ken64a, Lac62, Mat77, SB79, Sno64, Sub79, Ter79, Tuk62]. Live [YJY24]. lives [FP95]. Living [PD22, Ris98]. Lizzie [Nag86]. Load [Che17]. Local [Boi18, Cho18, St.06b, vMPKG15, YJ01].

Localising [Hof07]. Localization

[MMVH10, PDD^{+}24]. Locating [Swa18]. location [MRE96]. logical [Hub85]. logistician [Fei79]. Logistics [JJ10]. Logos [Rif10]. Long [Van71]. Longitudinal [FL23, FVB16, Hov08, Ste99]. Longo [Rou19]. Look [Gra75a, Jos70, KPV08, Ano73y, Ora86]. Lookalike [VV16]. Looking [BG07, Phi72a, Phi72b, Vel88, Mur00g, Mur00d]. looming [Mar06]. Lost [Mel08]. Lotka [Sub79]. Lotus [Gur95]. Low [Jin18b, Dre75, Ese69]. Low-Impact [Jin18b]. Lower [Mac75b, Ort74, Mar77]. lowly [Ort74]. luc [Cod09]. Lucena [Flo20]. Lukens [Cun10]. Lutz [Bea99]. luxury [Mur58]. LVIII [Red98]. lying [Wei95]. Lysenko [Sta13].

M [Ano71c, Bar09b, Ber12, Bro83b, Cha65, Cha80a, Dav93, Daw10, DOHM96, DR10, Eri11, Fer07c, Fer07d, Fer08, Fer09a, Flo06d, Flo07b, Har80a, Hou21d, Mas08, Mek72, Mic65, Mic85b, Sal80, Xu07, Zap80]. M. [And97]. Ma [And97, Nyc97, Rok19]. Machine [Lev68, Sch66, AL03, DSC78, Luf65, Mar85, Var77]. machine-readable [Var77]. Machines [Doe69, Fer06, Rat61, SSS^{+}13]. Macnealy [Erv99a]. Macromedia [Ano07k]. Macrostructure [JLS20, Rau21]. mad [Von78]. Madelyn [Suo13]. Madness [Ano15-28]. magazine [Ant79, Bow73]. magazine/journal [Bow73]. Magazines [Mar64]. Magic [MD84, Hof87, Sch80a]. Magical [Dou02a]. Magnan [Ano71d]. Mahwah [Joh97b]. Maier [Ano63d]. Mail [Flo07e, Cho97a, EM94, HB94, Ros94, TDS95]. mainland [Che92]. Maintenance [HSA21, HVS^{+}22, LE18, Pla71, Ber92, Jon89]. Major [Con58a, Tus05, Ano63j, Ano87-28, Max88, RW86]. make [Bur84b, CH89, Cox80, Gou73, Hoy58, dALT92, Sch81d, Whi91, Whi79i]. makers [HSR84, Riv85]. makes [Aug80, GW84, Teb80]. Making [Ama03b, Ana21, Ano63r, Atl81, Baa07, DM81, FR86, Fri13, Gar79a, Joe80a, Kru00, Lyb78, MVV09, O’D07, Rab92, SNT^{+}14, Tho67, AV98, Ano04j, Koc75, All97]. Malaysia [BJ21b, MOA02]. Male [Bro93]. Male/female [Bro93]. Malone [Hou21d]. Man [Hat66, Jon78, Sau92]. man-computer [Jon78]. Manage [DR20]. Management [BJ21b, And14, AB15a, AB15b, Ano08s, Ano13v, Bat14, Bat18b, CL19, Cog15, Fon61, Fro06, GAGW15, GPHDM16, GYK14, Hac16, Mas08, MFL17, SU09, Ste05, Whi79a, Ack89, App02, Ben79, Ben85, BM00, Dou72, Fle85, Flo03a, Fre13, GH85, GHG91, Har75b, Hil85a, Hil85b, Hol74, Los85, MSW97, McA85, MYV^{+}91, MYV92, Sha94a, Sha94b, SR83a, Sou85a, Tho73, Top87, Cos23b, Bar09b, LK08, Pot00a]. management-systems [Fle85]. Manager [Ano63t, Lin04, Mey87]. Managerial [Goo97a, dGKNG06, Dou01]. Managers [Ale20, Räi20, PM04, Pro94, Whi80a]. Managing [Bak75, CCL24, DB24, Fro06, Gro99, KE86, Mas08, Xu08, Yor78, You10, Bas97, FO06, Gui94, JDW96, Kan94, SB91, Wha02b, Fer07c]. Manual [GW84, LKS13, Pla71, Ber92, Cra83b, Fee86, Fie83, LS84, SH93, Wat90, ZB84, vdM96, Man18, Fri85]. Manuals [Ano97d, Dor63, LdJK20, Mil62, SDB^{+}13, Atl81, Bar98b, Bis94, BT95, Cra92, FR86, GvdMdJP99, GW84, Gol82b, Gri83, KE86, Lop84, May82, PS96, Sha94a, Sha94b, Til79, vdM96]. manufacturing [MYV^{+}91, MYV92, Oli58]. Manuscript [MM68, Lin90a, Lou58]. Manuscripts [Kun69, Mic61b, JCJ^{+}75]. Mapping [PM08, RWAK24, ZNG22, ST05a, ST06a, Sch78c]. Maps [Cod09, Win90]. Marcus [Rok19]. Marginalia [Smi10]. Margot [Hou22a]. Maris [Flo03c]. Mark [Cor94, FP95, Pen98g]. mark-up [FP95]. Markel [Dav93]. Market [D’A64, Pot77, LJ10]. Marketability [Cro95]. Marketed [KI06]. Marketing [CNC20, Kil09, Mem77, EZ88, Gra75b, HC88, Sam88, War88, Wha02a, Zim99]. marketplace [AD88, Gur95]. markets [HM73]. Markham [Mog09]. markup [App01b]. Marshall [All97, Whi78b, Lev86a]. Mary [LD20, Dou01, Erv99a]. Marya [Riv85]. Masaaki [Rok19]. mass [Lin72]. master [Lem83, Mur00b]. masterpiece [And81]. Mate [LFT08]. Material [Rou17a, Tha04, MK84, Pic80]. Materiality [Joh97b]. Materials [Hof05, Jos81]. Math [Hou21a, Bur90]. Math-Based [Hou21a]. Mathematical [Ano68u, KG68, Pot77, WKSH68, Gre87]. Mathematics [Kor75, Pod84]. Mathes [Sch78b]. Matter [Ano71l, Car24, FDJW18, Oka64, Ano77-32, Wil80]. matters [Lip04, RK91, SS87]. maturity [Mac90]. Maximizing [Key87]. Maxwell [Flo03c]. May [Dob75, FBE^{+}03, Wal97]. maze [SS87]. McCormack [CD98]. McCulloch [LD20]. McD [Ano80g]. McGraw [Kie83]. McIntire [Wof09]. McKee [Gar19]. Mcluhan [Lev86a]. McLulan [Per70]. Me [Flo07e, Vel88]. mean [Mac05a]. Meaning [Gal01, Gui86]. Meanings [Sap08, ZFWX21, NC81]. Means [BP09, Cod04]. Measure [ZLCZ17]. measurement [FG84, Sid87]. measures [Fra81]. Measuring [GE13, RR93, SHS97, UN00, XS24, PY04]. Mechanical [Lev68, DRG^{+}02]. Mechanisms [KL12a, Wei77]. mechanization [Tel68]. Media [AZDA10, Ant18, CWSL19, Fer05b, GMT08, Goo97b, HDKC08, Koc05, KL12a, KCM15, LL09a, Lof10, MKBF09, MVV09, STP22, SvI18, Sto05, SWR21, SZLY22, TM08, Tom69, YFW15, Yua21, ZNG22, Bae12, DR94, ESM98, Lee00, Man99a, RD05, Spi94, St.01c, WH98, vLUA02]. Mediate [Rif10].

Mediated [BSS13, CB08, DB24, FVB16, GMG^{+}23, GG13, HK07, KL12b, SCW20, DR94, HFS94, Koc03a, Kru00, Kva94, RLS06, Str01]. Mediating [KL12a]. Mediators [BFH03]. mediatrix [Nee92]. Medical [Kir20, SDB^{+}13, BT95, She73, Swa05]. Medicine [Rui20]. medium [ES98, Gar79b, GN94, Joh90, Koc01, PY04, ULK01]. Meet [Ano58h, Ano63r, Bak82, Rob79b, Sam70]. Meeting [Hea69, Wha86, Wol06, Ano83-31, Got79]. Meetings [Ano63l, CHME07, KDBC20, Räi20, Bul72, Rau81, Fer07a]. Meets [BR16, TDS95]. Meloncon [GG22a]. Member [Ano12o, Ano12p, Ano15-29, Ano15-30, CL19, GG13, Ano63r, RLS06]. Members [Ano58j, Ano64p, SSS09, Sal62]. Membership [Tro66, Ano77q, Ano77r, Ano77-28, Ano78o, Ano07a, Ano12a, Ano12b, Ano13a, Ano13b]. Memo [Mig83]. memorandum [Fra89, Mac99a]. memory [Pot00a]. Memos [HSR84]. mendacity [Wei95]. Mental [JW22]. mentoring [Bro93]. Mentors [Kry99]. menu [BRL91, FS92, PDG^{+}88]. Meredith [Sch78d]. Merger [Hoo77]. Message [Ano64y, Ara59, GE13, LHAL23, SI06, WS23, GN94, Lin93, MRE96]. Messages [ZC24, Sch72c, Wis91]. Messaging [AZDA10, GTTX08, HF08, HDKC08, Sta08, Ste08, Los85, Pas03, Zho05, ZZ05]. Meta [NC81]. Meta-talk [NC81]. Metadata [Bar09c]. Metagenre [SSV^{+}23]. Metainvestigation [NP23]. Metaphor [Fer10, HZX24, SJYZ18, Yu20, Zim93, RLG97]. Metaphorical [Sel95]. Metaphors [WMPBA07, SRVR97]. Metaphors-in-Use [WMPBA07]. Method [Ano15-28, Coo10, Kin12, LDKC23, RB19, BCB84, BW90, CDL84, EE78, Hag90, JGH82, Mic74, Pit79, vW00, Mog09]. Methodological [Tha24, Gui89b].

Methodologies [Gan22]. methodology [Bat02, dJS00, Pak83]. Methods [Ano63d, Cla59, Daw10, Hou21b, Hun21, LBB18, NNB^{+}22, Per63, Rei99, RA97, SDB^{+}13, Smi85, Gre62, LD97, Mun02, Sch89, SR83b, Whi89, dJS00]. Metric [Ano74i]. Metrics [PMGW21]. Mexico [GLCA22]. Meyer [Rei79]. Meyer-Hartwig [Rei79]. MGM [Ano15-29, Ano15-30]. mHealth [Ach24, RWAK24]. Michael [Wha04, Ano78p, Fil06]. Michaelson [Kre86]. Micro [KNMM09, RDJM09]. Micro-Insertion [RDJM09]. microcomputer [Bra84]. microfiche [Bov73, HE77]. microfilms [Kun73]. microform [Gra75b]. Microformats [Sto09]. Microforms [Bur75, Hur77, Jos75, Web75]. micrographics [Ave75]. Microphone [Ano80-33, Tay59]. Microsoft [DR10, Wil24, DR10]. Microthesaurus [Rub66]. mid [Ben85, MG95]. mid-channel [Ben85]. mid-sized [MG95]. Midwest [ZS04]. Mike [Pet13]. Milestone [Ano17d, Kei67]. milestones [Wha86]. milieu [Zim01]. mill [Dob85]. Millennials [Shr20]. Millstone [Kei67]. Millwrights [Ram00]. mind [Bal08, Rid87, SR83b, Yee90]. minds [Cro80, Cod09]. mine [Reh94]. Mines [OW91]. minimal [SH93]. Minimalism [HVS^{+}22, Man98b, OC97, CvdM96]. Minimalist [CR06, vdM07, FW90]. Minimize [Che17]. Mining [KI06]. Miniprints [DBM73]. minus [Dou02a]. minute [Win73]. Minutes [Wol06]. Miracle [LS61]. Mirel [Wil03]. Miriam [Sch79]. Miscommunication [Bre10, Rou17b]. misconceptions [CvdM96]. Misinformation [TVRR20, Bas97]. Miss [Mek72]. Missing [LL09b, Gri81]. mission [Tar01]. misunderstood [Sch80b]. misused [Woo79]. mix [Mem77]. Mixed [SDB^{+}13]. Mixed-Methods [SDB^{+}13]. MNCs’ [XS24]. Mnemonic [Mor67b]. Mobile [BN07, Fer07d, GE13, KNMM09, LLY18, WBB13, Zet19]. Mobility [Rou17a, Del92]. Modalities [vHK08]. mode [Sau93, Vol82]. Model [AZ17, AAA19, AME21, GB16, HS18, KNMM09, KT20, KPV08, MV21, Wil24, And84, Ano81v, Ant99, Bis96, CR84, CE05, Dav74, Dri87, Evi04, Foy96, Gre87, HN93, Koc03b, Lin72, MSC96, Met73, Pot77, RD05, TSS^{+}01, Zap80]. Modeling [LG14, Sau93, TCW06]. Models [Gra96b, MSDF22, Ama01, BSR85, Chi89, Die04, FP95, HM73, JF03, Kry99]. Moderating [PFD^{+}21, HMF02, Ker86]. Modern [St.00a, Wof09, Tru88, Whi75]. Modernization [Phe20]. modes [GR89]. Module [Phe21]. Moed [Pet13]. Mogull [Kir20]. Momentum [Sch17]. Monitoring [HF08, DeT94]. Monkeywrenching [Ros15]. Monroe [Sch78d, Fre80b]. Monthly [EL84, Wen85]. Montmagny [Cap93]. Moore [Sta20]. Morale [Bro62]. Morris [Whi78d]. Morse [You11]. Mortal [Nob10]. Moses [Mig83]. Most [Fer07c, Ano06j, Ano06k, Ano07l, Ano08k, Ano09m, Ano09n, Bai81a, JS81]. Mother [Cun09]. Motion [Dwo69, Str60]. Motivating [ZNG22]. Motivation [LKS13]. motives [GR89, Fer10]. Motorcycle [Jon89]. Moulton [Ano80g]. move [Zim01]. Mover [AL03]. Moves [Flo05e, Kru00]. Moving [Mac90]. much [Cam96, Cha78, GN94]. Multi [MK84, Mat77, Koc03a]. multi-iteration [Koc03a]. Multi-language [Mat77]. Multi-use [MK84]. Multicase [Bre10]. Multicultural [Fer07a, Pin18]. Multidimensional [Jin18a, SKZ23, ZC24]. Multidisciplinary [AME21, PMGW21, VNC^{+}13]. Multifunctional [LCX12]. multijournal [Cru73]. Multilevel [Tri16]. Multilingual [Bat18b, Dur22, HS19, Hou23, SV24, LH00]. Multiliteracies [Moo05]. Multimedia [Bis70, FR13, HS78, LS94, Per70, TK04, MOA02]. Multimethod [DR94]. Multimodal [Joh17, Mou05, Räi20]. Multinational [PM04, YFW15, St.02c]. Multiplatform [Ant06, Ant18]. Multiple [DF70, DF75, DSC78, KZCR10, McC92, Met75, Dul92, Flo73, KSW78, ST99a]. Multiple-Image [DF70, DF75]. Multiple-Level [KZCR10]. Munter [Dou01]. murder [Jen02]. Muriel [Bar98a]. Murphy [Fre80a, Ano63-27]. Must [Pod84, Smi03, Whi78c]. Mutianyu [FDJW18]. Mutual [Red10]. Mutuality [FMT12]. My [Ser88, Cho97a, Rid87]. Myers [Den98a]. myIEEE [Ano14x]. Myron [Whi79h]. Myself [GL16]. mystery [Jen02]. myth [Dou02a]. Myths [Smi84, BD82].

N

[Ano77a, Ano78d, Ano79b, Ano80c, Ano81c, Ano83c, Ano84b, Ano85d, Ano86c, Ano87b, Ano96, Ano97a, Mal11, Mog09, Sta20]. N.J [Die72]. Nadziejka [Ano97a]. Nagelhout [Gre17]. Name [Ano64z, Cha84, Ibr85, JKW83]. names [Pad78, San83, Vel77]. Napoleon [St.02b]. Narain [Cle03]. Nardi [Nyc97]. Narrate [SKZ23]. Narrative [SZLY22, WARAL16, Moe98, EC00]. Narratives [Ath22, JW22, WARAL20]. Narratological [Baa07]. Narrow [JSZ07, Sul89]. NASA [PMGW21]. Natasha [Sta20]. Nathans [Sto15]. Nathans-Kelly [Sto15]. nation [Sal80]. National [Ano59i, Ano77c, Coc04, FR13, KPV08, Pal77, Sch66, SEN10, BGN^{+}95, Fry62, MOA02, Sha89, ULK01, VLA96, VVL^{+}01]. native [Den00, Kuo93]. Natural [BR68, Cli63, Tho70]. Naturalness [KCM15, Koc05]. Nature [Hin10, Hof87, KD00, Meh94]. Navigating [BTM12, Dia14]. Navigation [Bun71, SSSC08, App02, FS92]. Necessary [Mou75]. necessity [Mur58]. Need [Han67, Mud04, Mac68, Han75, KW93, Kru97, Mac75a, Ter72, WLMR82, Wit77, Wri81]. needed [Ano80e]. Needs [BJT^{+}19, Fil06, Bar82a, CM05, Dou72, Kus97, NKU03, Orr01, Pow81, Sar73, Sou85a, Tsu91]. Negative [Lin19]. neglected [Gar79b, GH85, Kur66, Min69]. Negotiating [Bat18b, Flo05e, Whi14, KD00]. Negotiation [WARAL20, Gri94, ULK01]. negotiations [PT05]. Neighborhood [Sel98]. Nested [Haw94]. Net [Joh10]. Netherlands [HvMK^{+}12]. Network [Ano17e, Ano17f, Ano17g, Ano17h, Ano18d, Ano18e, BWPH12, Fee86, Gar19, HS18, Noa03, Swa18, vDVvL19, Bel99, CS96, KL06, Sal80, Ano14n, Ano14o, Yea10]. Networking [LP16, KM94]. Networks [HLY22, Cho97b, Pfa86, Noa03]. neutral [Cor90]. neutrality [Wei93]. never [Bur84b]. newcomers [Kat98b, Kat98a]. News [Kon22, Ano78r, Wis91]. Newsletters [HvMK^{+}12, Pot82]. Next [Dav81b, JM01, Mun02, Rus77, SD06, Sho80]. next-generation [SD06]. Nick [Mon79]. Nicometo [Sto15]. Nineteenth [Fil05, Nee92, Dos06]. Nineteenth-Century [Dos06]. Nissen [Fer09a]. NJ [Joh97b, Nor97, vdM97a]. NNS [Fre03a]. No [SAP20, The97, dGKGN11, AB81]. Non [Ale20, KI06, KS14, KM94, ULK01, Kuo93, RK91, Sau92]. Non-commercial [KM94]. non-congruent [RK91]. Non-English-Speaking [KS14]. Non-face-to [ULK01]. non-native [Kuo93]. non-rational [Sau92]. Non-UX [Ale20]. Non-Writers [KI06]. nonacademic [Ede87, Cha96]. nonformulaic [Mac06]. Nonfunctional [KMR14]. nonlinear [Fri73]. nonnative [LL03, Orr03, RM03]. nonprofit [Ham96]. Nonprofits [KZCR10]. Nonsexist [BJ79]. Nonspecialist [Ano68-27]. nontechnical [JGH82]. Nontraditional [Sta99]. Nonverbal [GMG^{+}23, Zim93]. Norah [Dur22]. Normal [Gre17]. Normalized [YTY12]. Norman [Ano63d]. Normand [Erv99c]. North [Cos23b]. Norwegian [Zem21]. Nostalgic [Reh98]. Nostrand [Zim99]. note [HR60, JVT01]. notebooks [Hug79]. Notes [Coh60, Smi10, Ada81]. nothing [Gud82]. notices [Bea98]. Notification [ASHB16]. nouns [Sch81d]. Novel [FMB18, Mil64]. Novelty [MFL17]. Novice [Ley08, LNCL05, WBW89]. NTIS [Coy77, Urb77]. Nuclear [NB04, Til06]. numbers [Dou02a]. Nurnberg [FW90]. NVivo [HK11].

Oaks [All97]. Oaxaca [GLCA22]. Obfuscation [Arm82]. object [Bis96]. object-oriented [Bis96]. Objectives [Ano61x, Bai81a, Vel89]. objectivist [Bra93]. objectivity [San87b]. Objects [WJ13, Sch81c]. Observations [ELdJ12]. observed [Ibr89]. Observing [Pet13]. Obtaining [Gol77]. Obvious [Ano66m]. Occupational [Car12]. O’Connor [Zap80]. Odell [Ede87]. odor [Wei95]. Off [Nee80, EC01, Sie78]. Off-the-cuff [Nee80]. offer [Aic78, Ask73]. offered [vdM96]. Office [BS16, DR10, JP86]. Officers [Ano58h, Ano64r, Ano60k]. offices [Ama03b]. Official [ZWX20]. Officialese [McM61]. Offshore [WMPBA07]. Offshoring [O’D07]. O’Hair [You10]. Oil [YFW15]. Old [Ben61, DJ82, HN93]. Olde [Cun09]. Oliu [Den83c, Har80b, Den83b, Whi77c]. Oliva [Mas08]. Oliver [Sal61]. Olken [Mic81, Whi79i]. Olympic [Fla85]. On-demand [Hur77]. On-Line [MM68, Lip75, HHE97, RLG97]. on-product [Dri87]. on-site [Har97]. One [Eva60, Hun21, Luf69a, Luf71, AB81, Ano63r, Ano85z, Ano87k, Con78, Ken82a, Kre88, Luf68, San87a]. Ongoing [KC19, KW23, You10]. Online [Bar14, BIdJ^{+}21, Bat18a, CBSM^{+}09, CJ89, FP95, FS09, HBS08, JCF92, LR24, LFT08, Mog14, Rub86b, SA23, SHB08, St.06b, SSSC08, SWR21, Swa18, WT09, Zem21, ZLCZ17, AT88, Ama01, Ama03b, BFH03, BRL91, BJ86, BBRT86, EZ88, FC96, Hor90, JE95, Joh90, Ker86, Mac03b, MMCL00, PDG^{+}88, Que86, Rab92, Rid87, Sch86, STM09, SJ99, Vel87b, Wal87, Bar14]. only [Ano85z]. op [Kre00]. Open [Ano12-28, Ano14-27, Ano14y, Ano14z, Ano15-31, Ano15-32, O’D07, SLP^{+}79, Waa09, Ano12q, Sal62, SF99]. Opening [Fil82, Ora86, Phe21]. Operating [TM15, Fee86, Lop84]. Operation [Fie65]. Operationalization [vDVvL19]. Operations [Son65]. Opioids [Wel22]. Opportunities [Emi99, PH24, Cap92]. Opposing [You11]. opposition [Lay88]. Oppression [Sim22]. Optical [Lev68, Mac75b, Li80]. optical-fiber [Li80]. optimal [Jen02]. Optimization [Kil13, LMJK14, Koc03b]. Optimizing [Ste08]. options [Shi75]. Oral [Ano67i, VMM21, AAR84, EM78, Pot84, Rat80, Wal98, War88, Jen00b]. oratory [Mon78]. Orchestra [RC59]. Order [Ano07t, Ano06m, Ano06n]. O’Reilly [You11]. organic [BCB03]. Organization [BJ21b, Cov75, Hot06, SY09, App01a, BB99, Cha60, Del92, FS99, Kur66, Swa05, Rou17b]. Organizational [Ano63t, Cos23b, DC07, Fer09a, Gia66, Goo96, Lin07, RD16, TM08, ZZL22, FO06, Gal01, Goo95, KK90, RH01, ST99a, TSS^{+}01]. Organizational/Employee [DC07]. Organizations [Fer04b, ZWX20, Baa94, ESM98, ES99, Ham96, HP86, Hol74, Kat98a, PY04, Ske90, Fer08, Sch78b]. organize [Cra83c]. organized [FS00, SF02]. organizer [Car87]. Organizing [CHME07, Dro80, KM61, Smi60, Lin90b, LC79, Pot83]. Orientation [Ano69x, Ano70c, Ano71e, Ano71m, vdM07, Par86]. oriented [Bis96, DG75, Gor90, Tel68]. orientedness [Gla04]. original [Zim00]. orthography [RL78]. Other [Bos64, Fer07c, Met75, Pot82, Zha21, JVT01, RCK^{+}90, SSM74, Tru87, Vis77]. Our [Ano58h, Ano64z, Ano85z, Gra75b, SSV^{+}23, Sam58, Wis15, Koc05, Sch65, Von78]. ours [Reh94]. Out-Group [PHW13]. Outcomes [Cos23a, KCLG13, Wil10a]. outline [Rat86]. Outlining [Kin12, Pot00b]. Output [Cho21]. Outsourcing [Pad07, Raj12, St.07, Yea09]. overcome [Bai81b]. Overcoming [Hor91, Kor77, Sim22, Stu92]. overhead [Gou72]. Overlap [SU09]. Overload [KPV08]. Overlooking [Ano66m]. Overseas [XS24]. Overview [BL13, Tus05, Boy69]. Own [Ang59, Ano68w, KWLS17, Bat84, Gla04]. Owned [SZLY22, WS23]. Ownership [Ano77z, Reh94].

P [Bar22, Flo07d, Hin10, LK08, Luf64, Pic80, Wof09]. Pablo [Zet19]. Pacific [VMM21]. Packaging [Mel09, Ter73]. Page [Ano69x, Ano70c, Ano71e, Ano71m, Ano17b, Ano18b, Ano18c, Bur69, Mur00h, SM10, Wil98, Ano77-33, Ano78q, Ano83g, Ano83h, Ano83i, Ano84g, Ano84h, Ano84i, Ano84j, Ano85j, Ano85k, Ano85l, Ano86h, Ano86i, Ano86j, Ano86k, Ano87g, Ano87h, Ano87i, Ano87j, Ano12d, Ano12e, Ano13d, Ano13e, Ano14b, Ano14c, Ano15b, Ano15c, Ano16b, Ano16c, Ano17c, Gre87, Hur83, Rub86b]. Pages [Rav06, Mur00g, Wei98]. painless [Whi80b]. Pair [Cho21]. Palfrey [St.09]. Pam [Hou21b]. Panel [Bos64, Lou58, NLK10, Aic78, Dob75, Dre75]. Paper [CDM10, Hea69, Lan04, SM10, Bas72, Bli72, Bue76, Day77, De 72, Dou72, Gar72, Gar80b, Got79, Gre72, Ham72, Her64, Lin72, LK72, Mac72, McG72, Men72, Pri72, Rob72, Sar72, Shr72, Sum72, Ter72, Mic80]. paperless [Joe79a]. Papers [Ano63k, Ano63z, Ano08j, Ano09k, Ano09l, CNC20, DF70, Far64, Fer12, Fie65, Mic59, Ano75d, Ano77d, Ano77e, Ano77-30, Ano82d, Ano82e, Ano83l, Ano83j, Ano83k, Ano83-31, Ano84l, Ano85m, Ano86n, Ano86o, Ano87l, Ano87m, Ano04j, Ano09j, Ano10j, Ano10k, Ano10l, Ano11c, AL03, Bra82a, DF75, GG84, Har89, Hin72, HR60, Joe80a, Joe81b, Joe81c, Joe81a, Joe82a, Joe82b, Joe83a, Kie82, Kre86, Mat91, Mic74, MF83, Pot82, TS81, Van63, Whi79b]. papers-Tutorials [Ano10j, Ano10k, Ano10l]. Paradigm [Cas84, Hel07, AA93, Bra93]. Paradis [Bar98a]. Paradox [KL12b, RD05, Cas84, PS96, QK05, RH92]. paragraph [McC79]. Paragraphing [Col77]. Parallels [RP98]. Parker [DR10]. Part [Ber12, Ama04a, Dud61b, Dud61c, Kat98b, Kat98a]. Partial [LG14]. Partially [PHP16, PHW13]. Participants [Ano69r, Phe20]. Participation [KL12b, App02, BMH05, HB01]. Participatory [Fer05a, JSZ07, Lon14, Evi04]. partner [Has80]. partnering [HHE97]. Partners [CCL24, Sha61, Sor75]. Partnerships [Boo17, Gra22]. Parts [KdJH14]. Party [JJ10]. passage [Ora86]. passages [FS99]. Passenger [WJ13]. Passion [LKHC16]. passive [LR85]. Past [BG07, Mil61]. paste [Gri84]. Pasts [Tha22]. Patch [HBS08]. Patent [Ano79n, Dia14, Hau80, May79, You11, FC79, Noo79, Sko79, Doo79]. Patenting [Ano83-29, Ano83-30, Hau79, Kim79, Ano82o, Ano82p, Ano82q, Ano84-31, Wri79]. Patents [Ano82o, Ano82p, Ano82q, Ano83-29, Ano83-30, Ano84-31, Joe79b, Nus79, Ter79, Wol79]. Path [GYK14].

Pathos [Rif10]. Paths [LC17]. pathways [HFS94]. Patient [AZ17]. Patients [RRW^{+}17]. Patricia [Rob78]. Patrick [Ama00c]. pattern [HSV05, Pri99]. pattern-communication [HSV05]. Patterns [Bra04, CSSV20, KZCR10, KT18, Bel99, Nor00, PD05]. Patterns-Based [KZCR10]. Pau [Zet19]. Paul [Col65, Whi80a, Mur00i, Sac01]. Pauley [Flo03d]. Paying [St.01b]. PC [Ano73m, Ano73a, Ano74a, Ano75a, Ano76a, Ano78a, Ano79a, Ano80a, Ano81b, Ano82c, Ano83b, Ano84a, Ano85a, Ano86b, Ano87a, Ano10m, Ano10n, WBW89]. PC-16 [Ano73a]. PC-17 [Ano74a]. PC-18 [Ano75a]. PC-19 [Ano76a]. PC-21 [Ano78a]. PC-22 [Ano79a]. PC-23 [Ano80a]. PC-24 [Ano81b]. PC-25 [Ano82c]. PC-26 [Ano83b]. PC-27 [Ano84a]. PC-28 [Ano85a]. PC-29 [Ano86b]. PC-30 [Ano87a]. PC-based [WBW89]. PCC [Ano83-31, Ano84-32, Ano84-33]. PCS [Ano85y, Kre87b, Moo87d]. PDA [Swa05]. PDAs [Swa05]. Pearsall [Whi79d]. Peck [Fre80a]. pedagogical [Sel99]. Pedagogies [Gan22]. Pedagogy [BR16, Dur22, GG23, KO20, MPD10, RB19, SMP21, Rei99]. Pedersen [Hou22b]. Pedestrian [Fer07d]. Peer [CB08, VMM21, Wat20, WP18, YTY12, Gar80b]. Peers [Räi20]. Peircian [Cha02]. PENNTAP [MQ76]. Penrose [Ber12]. Pentium [Wil97]. People [Bar09c, Coh69, Fis92, Fro06, Gon20, LO20, LS23, CM01, Jos81, Kor77, Mas08]. People-Powered [Bar09c]. Perceived [CL19, KPV08, LHMN07, LKHC16, PHP16]. percept [Tru91]. Perception [DB24, HBM21, LWMR07, SMS96]. Perceptions [BJT^{+}19, CCL24, CI02, FDJW18, LC17, MY14, PEG19, TM08, WFSC^{+}23, Cho97a, Nel03, PY04, VLA96]. Perea [Zet19]. Perelman [Bra03a]. Performance [BH16, BWPH12, FMT12, GE13, KL12a, PDL16, Rat80, TCW06, WPSB13, ZWE21, CJ89, Cra98, Gor90, Kol93, MRE96, WBW89]. performing [Ama04a]. Periodic [Sus61]. Periodical [Ede77, Ask73]. periodicals [Ano85z, Pal77]. Perish [Hou83, Rem78, Sea73]. Perot [St.02b]. Perpetuating [BN07]. Persist [PD22]. persistent [Wei95]. Person [BJ21b]. persona [Bra84]. Personal [Coh69, Fes62, HdJK18, LKS13, LKHC16, SGM^{+}24, SC81, Baa94, Hil89, HE77, Kil82, Pfa86, Fer07d]. Personality [Rub86a]. Personalization [Tra12]. Personas [Fri15]. personification [MG82]. Perspective [BAB10, CWSL19, Che17, FDJW18, HS78, RD16, SC12, SSSC08, WARAL16, Adl00, Cla87, Goo94, Har89, VLTP95, Wol88]. Perspectives [AB15b, Bat18a, Boo17, Cle16, Fer08, KHH^{+}08, LC17, Ley08, SSW17, St.07, Tus05, WX06, Whi06, BPK94, Dul92, HM94, HCKJ03, Mar89a, Per93, Sel95, Sel99, Waa09]. Persuade [SAP20, Gil87b]. Persuasion [HvMK^{+}12, Mac61, Flo00a]. Persuasive [Rou17a, His78]. Persuasiveness [HvMK^{+}12]. Petroleum [Mar70]. Pettersson [Ram03]. Pfeiffer [Nor97]. Pflugfelder [Rou17a]. PGEWS [Ano60k, Ano61x, Ano62e, McN58a, Pat59, Sal62]. Ph.D. [KT20]. phantoms [Moo87b]. pharmaceutical [BM00]. phase [Rho77, Som89, Kan85]. Phases [WLM^{+}22]. philosophy [Dom95]. Phishing [WHC^{+}12]. Phone [KNMM09]. Phones [Fer07d, Lon14]. Photo [Whi78b]. Photocomposition [Bec68, Kor75, Mac75b, BS73]. Photocopying [Zur75, KF77]. photographer [Kow78]. Photographic [dGKNG06]. photographs [Sus85, Wal98]. photography [Bur78]. Phototypesetting [Cra83b, Lev68, McC65]. Physical [Dar82, Lid75]. physics [JF03]. Pictures [Dwo69, vHK08, Str60]. Picturing [Fer06].

Piet [Wei02b]. Pipeline [SvI18]. Pirsig [Jon89]. Pitch [BG16, CSSV20, MV21, PH24, SNT^{+}14, SNT^{+}15]. Pitches [CSSV20, LKHC16]. Pitching [BG16]. Pitfalls [Con58a, Shi81, Mic90, RO96]. Place [Jan08, ROW19, dGKGN11, Vel87a]. Place-Based [ROW19]. Places [Col65]. Plain [CAK^{+}17, Che17, Dre17, GCMQ17, JW17, MMW17, Ros15, Sch17, Sim22, Dre16]. Plain-Language [Dre17, GCMQ17, JW17]. Plain-Style [CAK^{+}17]. plan [Ano67b, Mar77]. Planning [And14, Bul72, FW23, Hof05, HK22, MV83, Moo16, Pot83, Roc59, Rua80, You10, BCR00, Cox93, Eng72, Hap73, St.85]. plant [Jen66, NB04]. platform [Nor58]. play [Fla05]. Playbook [SSV^{+}23]. playing [Rob96]. please [Dav81b]. pledge [Moo86c]. plenum [Wei98]. plight [Van79]. PLS [LG14]. Plurals [Ano70b]. plus [Dou02a, Mek72]. Pocketbook [Str84]. Podmajersky [Lie20]. Poe [Kie83, Ama00a]. Point [Zie69, Bal75, GD85, Mal01]. poise [Nor58]. polar [Luf73]. Polgar [Xu08]. Police [Ath22, Kni08]. Policies [Ano63j, Yea09, Gan58, Gol82b]. Policy [Ano63s, Ano70b, Kni08, St.09, Ste22, Lyb78, OW91, RK91, Rot78, TDS95, TYT91, UN00, Whi89]. polish [Ano81j]. politics [Bar93, Pod03, Coc04]. pollution [Bil73, Smi79]. polymer [Men72]. Poncini [Fer07a]. Poole [All97]. Pooling [Kre88]. Poor [Ort78]. popular [Den98b, JVT01]. Portable [Fer07d, Tay02]. Portals [Xu08]. Porter [Gar19]. Portfolio [LWMR07, Sto08, Wil10a]. Portfolios [Zho08]. portions [SSM74]. Posed [Bou62]. Positive [LR24]. positively [Gou73]. possibilities [Fis92]. Possible [Sha68, MC75a]. possibly [Sch80b]. Post [Cha06, BCR00]. poster [Ano79e, Mac78a]. Potential [KdJH14, Bas97, Bro93, FBE^{+}03, Key87, KL06, Man98a, PM98]. Potentially [GAGW15]. potentials [RO96]. Powell [Pen99b]. Power [Ano81r, Ano82r, CC22a, Kil96, Mac04c, NB04, ODF92, BCR00, Bra03b, Key87, MD84, Nye99, Swa80]. Powered [Bar09c, SV24]. PowerPoint [GH09]. POWR [HK22]. pp [All97, And97, Die72, Mek72, Nyc97, RA97, The97, vdM97a]. pp. [Joh97b]. Practical [Cuf19, Cuf20, Nor97, OM12, Phi69a, Ano81s, Ano84-32, Ano84-33, Flo02c, Hub89, Ken89, Mic90, Vel88, Sha94b, Sou96]. Practically [Whi80b]. Practice [BJ21a, Cle16, DL24, GG22c, Hel07, MCC62, NP23, Rei99, San86, St.09, TDG^{+}18, WX06, Wel22, Wer71, Wol06, Zim13, ABMF84, Bea99, BR00, Che92, HF84, Mas84, Mas85, Wan89, Flo20, Hen22, Kir20, Lof08, Fer03, Flo03a]. Practices [BJ21b, BIdJ^{+}21, BJT^{+}19, Cos23a, Fer08, GC22, HK11, Hun21, Jon15, LP16, St.07, SU09, Ste05, BPK94, BTM12, CM05, FWJ77, HP86, Nel03, Par06, PBK^{+}93, Hin10, Yea09]. Practitioner [Pet13, PT05]. Practitioners [Ach24, DGPR10, Kru97]. Pradeep [Pot01b]. pragmatics [Con98]. Pragmatist [SS10]. Precarious [Wel22]. Precis [Ano64-27]. precise [Edd83]. precision [Sch80b]. predesign [Wil80]. Preface [Ano68s, Ano69s, Ano73z, Ano75o, Ano77-27, Ano81t, Ano81u, Ano82s, Ano82t, Ano83-32, Ano83-33, Ano85-27, Arm85a, Arm85b, Arm86, Bra81, BJ81, Bra82b, Bra83, Joe77b, Joe78a, Joe78b, Joe78c, Joe78d, Joe79c, Joe79d, Joe80b, Joe80c, Joe80d, Joe80e, Joe82c, Joe83b, Joe84, JJ84, Mil84a, Mil84b]. preference [Rub86a]. Preferences [CAK^{+}17, GTTX08, LFT08, SSSC08, FS00]. Prefixes [Ano64-27]. Preliminary [Yu10]. prelingually [Mal87]. Prelude [Rho77]. Prentice [Die72, Nor97]. Preparation [Ano64p, Ano68u, Dor63, Gup73, Hat69, Mil62, Ack89, CH89, Gou60, HC88, MSD85, Yee88, Zap80]. prepare [Cla79, LD84, Rea80, Whi77d, Whi79g]. Prepared [ML78, Lou58, Var77]. Preparedness [FC24]. prepares [Sou88]. Preparing [Ano97d, BP09, Ede61, FW23, Fla05, Lee64, Lit90, Mac78a, Nel90, Pla71, VNC^{+}13, Hap73, MGB89, Pit79, Sko83, Woe75, Bec84, Wei01]. Prescriptions [Hay83]. prescriptivism [Mac99b]. Presence [TDG^{+}18, BFA05, RLS06]. Present [BG07, Rau79, Got79, His78, Ros83, Wha04]. Presentation [Bis70, BP15, DF70, Fer05b, Flo80, Flo05e, Gou70, ZFWX21, Aug80, Cam96, Cox80, DF75, EM78, Ker80, Ros83, Rua80, Sch82]. Presentations [Flo04a, Gre70, GH09, MV21, Sto15, Tus06, VMM21, Gle82, Hor89, Jen00b, LD84, Nee80, Pra87, Sko83, Wal98, Wan89, War88, Ano67i, Riv85, Flo01a, Fer12, Goo97a, Wei01]. Presented [Ano63k, Ano97a, Ano97b]. presenters [AW78]. Presenting [Dal81, Eal83, Gar80b, KR83, KA03, Rin59, Whi79b, BS98]. preserve [EM94]. Preserving [Dia07]. President [Moo86c]. Press [Ano08u, Ano09v, Ano09w, Nyc97, BFA05, Gir60]. Pressures [Gre17]. Preston [Whi80a]. Preteens [CBSM^{+}09]. prevent [Som89]. preventing [Cro86]. preview [Ano73m]. price [Kan85]. Primary [Boi18, SDB^{+}13, Met73, Mir75, Rus77]. Primer [Daw10, Flo05d, Hou21b]. primitive [Rot78]. Principled [Fer09a]. Principles [Ale20, Flo07c, JD83, Ram05a, ROW19, Ste05, Bag83, Cla75, Die89, Jon78, Wei02a]. Print [McC65, Sed69, Mal87, Sch86, Mur00d]. printed [Ano81r, Ano82r, Hur83]. Printer [Ano68w]. Printing [Hat69, Sho80, Ste77]. prints [Koc75]. Prioritizing [Ben22]. Privacy [BIdJ^{+}21, CBSM^{+}09, CC22a, Gar06, TDS95, Cho97a, HM94, Mar06, PM04]. Private [Wei75, Fre13, Wha01]. private-sector [Fre13]. probability [Mat91]. Probe [D’A64]. Probing [GMT08]. Problem [Ben61, Fer04a, Flo05e, KT18, KT20, PH24, Til06, Bri99, Hac84, Kap62b, KVRR06, KR95, Ora86, Par06, RK91, Sau92, Sko83, vdM96, Ano63d]. problem-based [Par06]. Problem-Solution [KT18]. Problem-Solution-Based [KT20]. problem-solving [Hac84, Sko83, vdM96, Ano63d]. Problems [Ana21, Axe73, Bou62, Cli63, JES13, Rin63, Str64, Sum72, Ter73, Win83, Die04, Flo02b, Gla04, Kor77, Mar86, Sau93, Wu88]. Procedural [GQ12, Roy07, vHK08, Die04, Gan04, RDRS04, Ste04, SHS97]. Procedure [Rub66, Gui89a, vdMG04]. Procedures [Men23, Axe73, Ber82, BKSM82, Cla75, Clo64, Gol82b, Hac85, Mor67a, Sau82]. Proceedings [Ano63h, Ano63s, Mic65]. Process [BNB21, Boe10, Hou83, JJ10, LeC01, AV98, Ano81v, BJ86, BMRH05, Cox93, Die89, Dul93, HP86, Hes75, Koc01, Koc03b, LNCL05, McG00, RY90, Ren96, RH94, Son65, TSS^{+}01, Wil77a, vLUA02, Ano80g, Kno80]. Processes [BH15, BR16, Joh17, LBE23, Red98, WOD^{+}16, Mas84, Mas85]. Processing [BR68, Cli63, Kun69, MM68, MKBF09, Pet69, Rho77, SGM^{+}24, WHC^{+}12, Web65, Bam73, BS73, DSF^{+}80, Gan04, HMF02, Ils80, Lip75, RP98, RL84, SP86, Woe80, Red98, Die72, Sch78e]. processors [GN94]. ProComm [Wal23]. produce [Vog83]. produced [GKD77]. Producing [Ano97d, Cra68, DBK^{+}93, AC85, Kru00, McG86]. Product [Bal18, Fer05b, MY14, Sty86, ZFWX21, Dou72, Dri87, Min69, PS96, SMS96, LeC01]. Production [GG22b, Van62, Ama01, Bro75, Cov75, DEBD83, FW78, KE86, Kro80, Rei02, Shi75, Tel68, ZB84]. Productive [Vic03]. Productivity [Cho21]. Products [Bat14, HdJK18, Rok19, Cle87, Wal74, Gon20]. Profession [GZ97, GG22c, WB92]. Professional [Ama00a, Ama04b, Ano58i, Ano59m, Ano59n, Ano59o, Ano61u, Ano61v, Ano61w, Ano62j, Ano63l, Ano63m, Ano63n, Ano64u, Ano73a, Ano74a, Ano75a, Ano76a, Ano77a, Ano77m, Ano77q, Ano77n, Ano77o, Ano77p, Ano77-28, Ano78d, Ano78a, Ano78l, Ano78m, Ano78o, Ano78n, Ano79b, Ano79p, Ano79q, Ano79r, Ano79a, Ano80c, Ano80x, Ano80y, Ano80z, Ano80-27, Ano80a, Ano80u, Ano80v, Ano80w, Ano80b, Ano81c, Ano81n, Ano81o, Ano81p, Ano81b, Ano81a, Ano81m, Ano81l, Ano82l, Ano82m, Ano82n, Ano82c, Ano82a, Ano82b, Ano83c, Ano83w, Ano83x, Ano83y, Ano83b, Ano83a, Ano84b, Ano84y, Ano84z, Ano84a, Ano85d, Ano85t, Ano86c, Ano86u, Ano86a, Ano86l, Ano87b, Ano96, Ano97c, Ano00, Ano04n, Ano04o, Ano04p, Ano04q, Ano04f, Ano04b]. Professional [Ano04g, Ano04c, Ano04h, Ano04d, Ano04i, Ano04e, Ano05k, Ano05l, Ano05m, Ano05n, Ano05f, Ano05b, Ano05g, Ano05c, Ano05h, Ano05d, Ano05i, Ano05e, Ano06o, Ano06p, Ano06q, Ano06r, Ano06f, Ano06b, Ano06g, Ano06c, Ano06h, Ano06d, Ano06i, Ano06e, Ano07o, Ano07b, Ano07p, Ano07q, Ano07r, Ano07s, Ano07g, Ano07c, Ano07h, Ano07d, Ano07i, Ano07e, Ano07j, Ano07f, Ano08a, Ano08o, Ano08p, Ano08q, Ano08r, Ano08f, Ano08b, Ano08g, Ano08c, Ano08h, Ano08d, Ano08i, Ano08e, Ano08u, Ano09a, Ano09o, Ano09p, Ano09q, Ano09r, Ano09f, Ano09b, Ano09g, Ano09c, Ano09h, Ano09d, Ano09i, Ano09e, Ano09v, Ano09w, Ano09x, Ano10a, Ano10o, Ano10p, Ano10q, Ano10r, Ano10f, Ano10b, Ano10g, Ano10c, Ano10h]. Professional [Ano10d, Ano10i, Ano10e, Ano11e, Ano11b, Ano11a, Ano12c, Ano12r, Ano12s, Ano12t, Ano12u, Ano12v, Ano12w, Ano12h, Ano12f, Ano12i, Ano12g, Ano12j, Ano12k, Ano13c, Ano13n, Ano13o, Ano13p, Ano13q, Ano13r, Ano13s, Ano13t, Ano13u, Ano13f, Ano13g, Ano13h, Ano13i, Ano14a, Ano14p, Ano14q, Ano14r, Ano14s, Ano14t, Ano14u, Ano14v, Ano14w, Ano14d, Ano14e, Ano14f, Ano14g, Ano15a, Ano15p, Ano15q, Ano15r, Ano15s, Ano15t, Ano15u, Ano15v, Ano15w, Ano15h, Ano15d, Ano15i, Ano15e, Ano15f, Ano15g, Ano16a, Ano16h, Ano16n, Ano16i, Ano16j, Ano16o, Ano16k, Ano16l, Ano16p, Ano16q, Ano16m, Ano16r, Ano16s, Ano17a, Ano17m, Ano17q, Ano17r, Ano17n, Ano17s, Ano17t, Ano17o, Ano17u, Ano17v, Ano17p]. Professional [Ano17w, Ano17x, Ano18a, Ano18j, Ano18n, Ano18o, Ano18k, Ano18p, Ano18q, Ano18l, Ano18r, Ano18s, Ano18m, Ano18t, Ano18u, Ano19a, Ano19d, Ano19b, Ano19e, Ano19c, Ano19f, Ano20a, Ano20c, Ano20h, Ano20d, Ano20i, Ano20e, Ano20j, Ano20f, Ano20g, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano22a, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22b, Ano22k, Ano23a, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano23b, Ano23f, Ano23c, Ano23d, Ano23e, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano24i, Ano24a, Ano24b, Ano24c, Bal08, Bay22, BH21, Ben22, BS16, BI18, BSS13, BG07, Buc74, Bur84a, Car20, Car24, Cha23, Cla12, De 72]. Professional [Doa19, Dom07, DR10, DGPR10, Eri11, Flo04c, Flo06b, Flo08, FC24, Fri13, FL23, Fry87, GLL23, GG22a, HS18, HK11, Lam16, Lam18, LBB18, LAWI24, Mar08, Mel08, Moo16, Orr06, Orr10, Rob96, SV24, SU09, Str21, Sur96, TDG^{+}18, Tha24, VV16, VG16, Wal18, WC15, WS04, Wol06, Zho08, ZM23, Ano85a, Ano86b, Ano87a, Axe73, Bea99, Bli72, Bra93, Buc89, Cha02, Cho97a, Cor92, Dav85, DSF^{+}80, DRG^{+}02, Dou02b, EC00, EC01, Flo02b, Goo94, HM95, JM01, Kry99, Lan94, Lid75, Mac99a, McC92, McD94, Ols91, Orr01, Pal00, RY90, RR93, Rob95, Sei72c, Sko83, Som89, Sou88, St.01a, Str84, Van83, Vel88, VVL^{+}01, Whi77d, Whi78e, Whi79c, Yee88, Yor78, Gar19, Wei95, Mac02a, Spi96]. professionalism [San86]. Professionalization [Cle16]. Professionals [Ale20, CAK^{+}17, DSV10, YK13, Yu10, vDVvL19, Ano87-28, Bas97, Mar06, SH91, Sou72, Sou75, SC81, Tsu91, Bro83a, Rob78]. Professions [Fer08, AAR84, BGN^{+}95, HPSR83, dALT92]. Proficiencies [KI06]. Proficiency [DSV10, Orr02, San86]. Profile [Ama00b, Ama00c, Bra99b, Bra99c, Bra00, Erv99b, Erv99c, Pen98a, Pen98b, Pen98c, Pen98d, Pen98e, Pen98f, Pen98g, Pen98h, Pen99b, Saw99, Tay01, RH94]. Profiles [BMF21]. Profiling [Gup83a]. profit [Ham96, Mem77]. Program [Ano15-29, Ano15-30, Ber68, CSSV20, KS14, Ock71, Ock75, Pie71, SSV^{+}23, SAP20, Wil10a, ZB17, Ack89, Ano87y, Kre87b, MV83, Mat77, PDK88, Pal77, You80, Ano64i]. Programmatic [KW23, Sel99]. programmed [Hau80]. Programmer [FR86]. Programming [BR68, Cho21, Rat61, Rea22, Hel88, Kau88]. Programs [Ano63v, Ano63w, Ano64k, Cos23a, GLL23, Gre17, And84, Gil87a, GHG91, Har84, How83, How84, Joe82b, Lay88, Mey64, SPR84, Zim01]. Progress [Dud61b, Dud61c, GR94, Pal77, Sen62]. Project [DR20, GAGW15, Lam18, LK08, Ree64, RCS02, VT21, WPSB13, XZZ10, YK13, Zem12, Cox93, Dro80, JDJ99, KK02, LL03, Loe91, Loe95, MYV^{+}91, MYV92, Ser88, Sha94a, Sha94b, VVL^{+}01, Web79, VMM21]. project-based [LL03]. projection [HS79]. Projections [St.07]. Projector [Ang59, Gou72]. Projects [CCL24, KSNF15, SJD20, IP02, KC81, Lyn83, Pot84, SH91]. Promises [Gro95a]. Promote [GCMQ17, ZWE21, Vel87b]. Promotional [LR24, SO98]. Pronoun [Coh69, Ort74]. Proofreading [Ano68w, Tur79]. Proper [Col65, Rac59, Col65]. Property [Dia07, Flo05c, Gar06, Bal08]. Proposal [Ano82e, Bra82a, Cra68, Ken63, MCC62, Wha86, Bar89, Bar91, BS89, DEBD83, Fre13, Goo05, Ham96, HC88, Jen66, KD00, Kil83, McC94, Ren96, Sar72, Smi60, Tra92, Wha85, Wha01, Flo00a]. proposal-writing [Jen66, KD00]. Proposals [Bec83, Fer12, Mac61, Bec84, BW90, CH89, Eng72, Eva60, Geo83, Gre87, Wha02b, Woe75, Mal01]. Proposed [Sus61]. Proposition [Cap78a]. Prose [Bil73, JGH82, Ort75, Sch67, Sch75a, Pot01a]. Prosecution [Dia14]. Protection [Rin63, Wor87, Mar06, Noo79]. protest [Gur95]. Protocol [Coo10, Lee00]. Protocols [ELdJ12, May19, KU04]. prototype [Ama05]. prototypes [Gil03, WJ13]. Prototyping [Lan04, SM10, MSW97, Wat90]. Proven [Hof05]. provide [KW93]. provided [Den98b]. providers [HCKJ03]. provides [Dre75]. providing [SWB^{+}05]. provision [Koe77]. proxemics [Fis92]. pseudonyms [Pad78]. Psychological [Baa94, Dev75, DR20]. psychologist [Rau81]. Psychology [Ano64f, Roh19]. PSYCHOM [Bas72, Bli72, De 72, Dou72, Gar72, Gre72, Ham72, Lin72, LK72, Mac72, McG72, Men72, Pri72, Rob72, Sar72, Shr72, Sum72, Ter72]. psychophysics [Shr72]. PTGEWS [Ano63r, Ano63t, Ano64p]. Pub [All97, vdM97a]. Public [Lof08, Lof10, MSDF22, Moo16, OW91, Til06, Tri16, Bai81b, Bar79, Bar82b, Bas97, Bol82, Chi89, Has80, MD84, RK91, Som89, TYT91, Wri79]. Publication [Ano23f, CFMM14, Dav85, Dre75, Mar77, ND83, Rot78, Sha68, Wri79, Ano04f, Ano04g, Ano04h, Ano04i, Ano05f, Ano05g, Ano05h, Ano05i, Ano06f, Ano06g, Ano06h, Ano06i, Ano07g, Ano07h, Ano07i, Ano07j, Ano08f, Ano08g, Ano08h, Ano08i, Ano09f, Ano09g, Ano09h, Ano09i, Ano10f, Ano10g, Ano10h, Ano10i, Ano11b, Ano12h, Ano12i, Ano12j, Ano12k, Ano13f, Ano13g, Ano13h, Ano13i, Ano14d, Ano14e, Ano14f, Ano14g, Ano15d, Ano15e, Ano15f, Ano15g, Bur75, Dun81, Kin77, MGB89, MM73, Pri97, Row75, She73, Sta73, Mic61b]. Publications [Ano63u, Bos64, Cot59, Cro59, Kai68, BD83, Bat02, Sam88, Stu69]. publicity [Arr74]. Publicize [Ano63l]. publish [Con83, Kre86, Mic80, MF83, Pat75, Rem78, Sea73]. publish-or-perish [Sea73]. Published [Ros64, Ant79, Bue76, Van77]. publisher [FWJ77, Har75b, Koe77, Shi75, Var77]. publisher-prepared [Var77]. Publishers [Fra75, Sor75]. publishes [Kre87b]. Publishing [Ano14-27, Hat69, LRM^{+}09, LO07, Met75, PM84, Rus77, Wil98, Ano62e, Ano12q, BJ86, Bro88a, Bro88b, Cov75, Coy77, DSC78, FHR87b, FD84, Gol97, HP86, Hur77, Key87, Lin90a, Mem77, PDK88, RFH87, Sut77, Van75]. Punctuation [Ber77, Sho80, ZZ05]. Punselie [Wei02b]. Purchasing [HBS08]. Purloined [Ama00a]. purpose [Fry87, Lin77a, Sty86]. Purposes [Izq24, DM81]. pursuit [Gro03]. Put [Mue80, Cam96, ML99, Str04, Vel87a]. Putt [Ram07, Cha81, Cha81, Ram07]. Putting [TC95]. Puzzle [Shr20]. pyramid [EL84].

Qualitative [DB24, HK11, Ock05, Tho01]. Quality [Ano04r, Bat14, Bat18b, Bli61, CL19, Cor69, Cra68, D’A65, Evi04, EBAS18, Hor89, KMR14, LRM^{+}09, LHMN07, MY14, Wil77a, Abr77, Cap92, Flo01b, Gui94, Jon89, LD97, Lin77b, LN00, Sch89, SMS96, St.00b]. Quantification [FMB18]. Quantitative [BP10, Kat84]. Quantities [Moe98]. quantity [Evi04]. Quasi [Baa07, BL13]. Quasi-Experimental [BL13]. Quasi-Rational [Baa07]. queries [SI06]. Query [Ano64-28, Per63, SD06]. quest [Ras86]. question [Dun82, Sar72, Sau93]. Questionnaire [GAGW15, Den99, Mac72]. Questions [Ano68-27, Bar09a, St.06b, Boy81, Dre88, Geo83, Rei88a]. Quetzaltenango [GLCA22]. Queue [Tho70]. quick [Dre83, Lin90b, Rei88b, Bra01]. quickly [Cra83c]. quiz [War80]. quoting [EM94].

R [Ano63d, Ano63-27, Ano80g, Dre16, Dur22, Flo07a, Fre13, Kat84, LJ10, Man18, Moe98, Rou17b, Sal80, Sch78e, Smi10, Sta20, Whi79g, You10, Lon14, Mic68]. R&D [DM63, Mic87]. R. [Fil05, Kre87a]. race [Sau93]. Rachel [Wil03]. Radio [Gar79b]. Rafal [St.09]. rail [Sau93]. Railroad [Fil05]. Rails [Fil05]. Ralf [Fer03]. Ramey [RA97]. Randlesome [Den98a]. Randy [All97]. Rapid [Rub66]. Rapport [CL19, BMH05]. Rapport-Management [CL19]. rarin [Hun81]. Rate [Pow80]. rates [Mar77]. Rating [BIdJ^{+}21]. ratio [Hof87]. Rational [Baa07, Cod09, Sau92]. Ray [Bra03a, Whi79c]. Raymond [Rui20]. RCA [Ano63i]. Re [GLCA22, Jon15, Foy96, Ano74i]. Re-Examination [Jon15]. re-invention [Foy96]. Reaching [Ano19j, Kan85]. Reactions [BJ21a, CDM10, Ken89]. reactor [NB04]. Read [Rid87, Dor83, Lou58, May79, Pil81]. Readability [Ano81h, Cla75, Doc61, DuB08, MY14, McC87, Plu81, ZJG17, Bar80, Bar81, BRS81, DGORM84, Dru85, KAOC81, Red81, Sha82, Whi87]. readable [DSC78, Lyb78, Var77]. readably [Tib82]. Reader [Bea59, Con99, Pet69, Ano86z, Ano87z, Ano87-27, Dre75, FS00, LD97, Mal85b, Olk87, Rub86b, Sch89, TC95, Wal85, Whi80c]. reader-focused [LD97, Sch89]. readers [Ano79p, Ano79q, Ano79r, Ano80x, Ano80y, Ano80z, Ano80-27, Ano81n, Ano81o, Ano81p, Ano82l, Ano82m, Ano82n, Ano83w, Ano83x, Ano83y, Ano84y, Ano84z, Ano85t, Ano86u, Con92, Dul80, MG95, Sou85a, SF02]. readership [Dat91]. Reading [Ano68r, Ano68v, JD21, Red89, Und67, AL03, Bah85, Cho97a, Gui86, LR85, Man86, MP79, SJ99, UN00, Flo07c, Whi79e]. reading-to-do [Gui86]. Ready [Sav88, Zap80]. Realities [Fer05a]. Reality [MMW^{+}22, TDG^{+}18, Tha24, Gro05, RVV^{+}02, TDS95]. Rearticulating [Reh98]. Reasonable [Cra68]. Reasons [vMPKG15]. Rebecca [Gan22, Sta20]. Reber [Lof08]. Rebirth [Sac77]. Recall [Mal87, Bea98]. recession [Sum72]. recipe [Tar01]. recipes [Cur80]. Recipient [SS10]. Recital [Ano64-29]. Recodification [Wil10b]. recognition [HSV05, Sei72c]. recommendation [Gat04, Wil84]. Recommendations [Phe20, VLA96]. Recommended [Wer71]. reconciling [BR00]. Record [Ano75m, SDB^{+}13]. Recruiting [AW22, Ano08u, Ano09v, Ano09w, Mun02]. Recruitment [Rif10]. recurs [Hof87]. Redefining [RL96]. redesign [Koc03b]. Rediscovery [Rau21]. Redstrom [Zim13]. Reducing [Cod04, BK03, Con87]. Reduction [FR13, San83]. reentry [Cor92]. Reep [Flo07c]. Reexamining [HDKC08]. referee [Cap78b, Man75]. Reference [Löc22, Oka64, WJ13, Fri85, Lid75]. References [Ano63e, HTB80, Mat91, Rei88b]. Refine [Lan04]. reflected [ULK01]. Reflections [Kva94]. reflective [PT05]. reflects [Den88]. Reflexivity [WLM^{+}22]. refuse [Ask73]. registration [Ano78k]. regularities [Baz99]. Regulation [Fer08]. Regulatory [KRM24, Wil10b]. Reidym [Wha04]. reinforcing [Evi04]. Reinventing [Bar14]. Reinvigorating [SM10]. rejection [JCJ^{+}75]. Rejections [KCM15]. Related [WPSB13, DBM73, Kol93, Lie77, Sch93, Ter73, Van63]. Relating [DJ82, VLTP95]. Relation [HK07]. Relational [AW22, MML15, MPD10, SC12]. Relations [Lof08, Lof10, Sto05, Bas97, Chi89, Som89, Suo13]. Relationship [AFH20, CL19, Dia14, LWMR07]. Relationships [Boo17, Sch72b, Lay73, Yor78]. relative [HR99]. release [Min69, Ris98]. releases [BFA05]. relevance [HMF02, WS89]. Relevant [Gir06]. reliability [WS89, dJS00]. Reliable [Boe10]. Relief [SY09]. relieving [MB84]. reluctant [Ger90]. Remaking [SNT^{+}15]. Remarks [Mey69]. Remley [Joh17]. Remote [SSS09]. Removing [Wil10b]. Renaissance [Ram98, BT95, TL92]. Renewal [Wha85]. Renner [Mur00i]. Repair [GL16, LE18]. repeatable [BMRH05]. Repetition [LHAL23, Zim83]. Report [Ano58b, Ano63t, Baa07, Bha12, Bre10, Bue77, Bur69, Flo03d, Flo05a, Izq24, KdJH14, KO20, ZB17, Zem21, dGKGN11, Ano71c, Bro89, Cap78b, EL84, Flo00b, Fre80b, Gat04, Hap73, JF03, MG95, Mey64, Moo87a, Moo85, SB84, Sen62, SF99, Sou85b, Ulr84, Wal99, Whi79g, Whi79h, Her95, Mic61a]. Report-Generator [Bur69]. Reporting [Hoh60, Lin19, WBB13, BR93, BCR00, Pot84, Mac02b]. Reports [EBAS18, Fil05, HZX24, Mic61b, Rav06, SJYZ18, Sus61, YB17, Yu20, Yu23, ZC24, dGKNG06, Ben79, Bre73, Cla79, Coh60, Ful80, GG84, His78, HSR84, KK02, Mal01, MF83, Pod03, Sch78b, Sei67, Whi78e, Kre86, Ano64e, Kie83, Whi79g]. Representation [GHR67, KL06]. representations [KVRR06]. represents [Dul93]. Reprints [Ano68t, D’A64, Ano06m, Ano06n, Ano07t]. Reproduction [Ano77-29]. Reputation [KdJH14]. Reputations [CFMM14]. Request [Ano63u, Joh10]. Requests [Sar72]. Required [Cas84, PFM88]. Requirements [Löc22, Adl00, Gan58, Per78, VLTP95, Yer74]. Research [AAA19, ABMF84, BESFW22, Bed12, BN10, BL13, BI18, Car20, Dav67, Daw10, DL17, DC07, Erv99a, FL23, GE13, Gro05, Han67, Hin10, HK22, HK11, Hou21b, Hun21, Jin18a, Jin18b, JLS20, Kan12, KT18, KT20, KL12a, KCLG13, KRM24, LS94, Lam14, LBB18, McC63, Pfa86, Rau21, Sen62, SC12, Sha61, Tha24, Waa09, WHC^{+}12, Whi06, WRTBRL23, Zim13, Ant99, Ant01, Bat02, Bro91b, Cho97b, CM05, CE05, Dud62, Eva00, Eva02, Fre80b, Han75, HH93, Hol74, Hor75, Ils80, Jam84, Koc01, Koc03a, Koc03b, Kru97, LL03, Lip04, Mac90, Mac97, MSC96, Meh94, Mir91, Sch73a, Son65, SWB^{+}05, SS87, Tuk62, VVL^{+}01, Wal74, dJS00, Pet13]. research-based [Eva00]. Researched [Hun81]. Researchers [Lam16, Phe20, St.01a]. Researching [Tha24]. Reshaping [MK84, Wil03]. Residence [HBS08]. Resistance [Lof08]. Resisting [Sta17]. Resolution [Flo05e]. Resolving [RM03, Rob18]. Resource [Bar09b, Ano06j, Ano06k, Ano07l, Ano08k, Ano09m, Ano09n, Kur66]. Resources [BJ21b, Fro06, JJ10, Mal07, Mas08, Räi20, Kat98b, Kre88, RCK^{+}90]. respond [Moo87b]. Responding [Fer07b, You10]. Response [GE13, Sch10, You11, dGKGN11, CSH84, Dul93]. Response-to-Complaint [Sch10]. Responses [CCL24, Dud61b, Dud61c, DLVE23, MG95]. Responsibility [Bha12, HZX24, Lin19, YFW15, Yu20, Yua21, ZC24, BS87, Sou85a]. Responsible [Wal64]. Responsive [KW23]. restricted [MCWK77]. restrictions [Tay73]. Restructuring [Pri72, Wan89, Rat86]. Results [Cho18, Com86, Den98a, KHV10, Sap08, BGN^{+}95, DBM73, GHG91, Gou60, HPSR83, Hub89, SJ99, VLA96]. Resume [BMF21, BP15]. resumed [Hun81]. Résumés [FMB18, Ega81]. Retention [Lie20]. rethinking [Spi94]. retrievable [Fra89]. Retrieval [DM63, GH70, Lam68, Mac68, Mai70, Per63, Sed69, Wil64, Ano65k, Ano65i, Die72, Kap62a, Ken64a, Mac75a, O’C62, SSM74, Sno64, Win90, Yer74].

Retrospection [Mal07]. retrospective [Ryd85]. Reuse [SNT^{+}15]. revelations [St.02b]. Reverse [Kin12, Rat86]. Revews [Lay97]. Review [Ale20, All97, Ana21, AB15a, And97, Ano63d, Ano64e, Ano64f, Ano69t, Bai24, Bal18, Bar14, Bar09a, Bar09b, Bar09c, Bar22, Bat14, Ber12, Bra01, Bra03a, Bra03b, CD98, CB08, Car00, Cha96, Cla16, Cla12, Con99, Cuf19, Cuf20, Cun09, Cun10, Daw10, DL17, Dou00, Dou01, Dre16, DOHM96, Dui96, Dur22, Dyr98, Eck03, EC00, EC01, Emi99, Eri11, Erv99a, Fer07b, Fer08, Fer09a, Fer09b, Fer10, Fer12, Flo00c, Flo00a, Flo00b, Flo01a, Flo01b, Flo01c, Flo01d, Flo02a, Flo02b, Flo03d, Flo03b, Flo03c, Flo03a, Flo06d, Flo07f, Flo08, Flo10, Flo18, Flo20, Gan22, Gar19, Gir97, Gon20, Goo97a, Gra96a, Gro99, Gro01, GG22a, GG22b, GG22c, Hen22, Hin10, Hol03, Hou21c, Hou21a, Hou21b, Hou21d, Hou22a, Hou22b, Hou23, Hun98, Hyd18, Jan08, Jen99b, Jen00b, Joh97a]. Review [Joh20, Kae03, Kam01, Kir20, Kle03, Kry00b, Kry00a, LM02, LJ10, LeC01, LK08, LD20, Lie20, Lof08, LN98, LN00, Mac02a, Mac02b, Mac03a, MY14, MYT16, Mag03, Mal01, Mal11, Man18, Mas08, McC19, McE18, Moe98, Mog09, Noa03, Nob10, Nor97, Nyc97, OM12, PD24, Pat59, Pin18, Pod03, Pot00a, Pot00b, Pot01b, Pot01a, Ram03, Ram98, Ram00, Red98, Reh98, Rei02, Rei99, Ril00, Ril01, Ril02b, Ris98, Roh19, Rok19, RA97, Rou19, Rui20, Sac01, SV24, Sch20, Sel98, SCW20, Sim96a, Smi03, Smi10, Spa14, Spi96, St.00a, St.01a, St.01b, St.01c, St.02a, St.02b, St.03, Sta99, Sta13, Sta20, Sto08, Sto15, Str01, Str03, Suo13, The97, Tra12, VMM21, Vic03, Waa09, Wat20, Wei01, Wei02a, Wha01, Wha02a, Wha02b, Wil03]. Review [Wil10b, Wil98, Wis15, Wof09, Xu08, Yea09, Yea10, YTY12, Zet19, Zha21, Zho08, Zim00, Zim01, Zol14, vdM97a, Boo17, Cha23, Far62, Fer07c, Flo07a, Flo07c, Gre17, GGS23, Hes75, Joh90, Jon73, Lee62, Mir91, Phi69b, RM03, Rou17a, Sch72a, Sch73b, Sch73c, Sch73d, Sei72a, Van62, Whi76a, Whi76b, Whi77b, Wil77a, Yu24, Cod09, Flo02c, Gen10, Hue02, Joh02, Moo02, St.09]. Reviewer [Ano68-27, Ano20b, Hay21]. Reviews [Ano63-27, Ano68t, Ano97d, Fre13, Kal97, Mur00a, Mur00b, Mur00c, Mur00f, Mur00d, Mur00i, Mur00h, Mur00e, Mur00k, Mur00j, Mur00g, Nad96, NNB^{+}22, Sim96b, Sou96, Win97, Ano62b, Ano84k, Bec84, BA83, Deb86, Den83a, DR10, Ede87, Fie67, Gup83b, Joh65, Kat84, Kre86, Kre87a, KK02, Mic85a, Mic85b, Riv85, Sch73e, You10]. revised [RL78]. Revising [GPHDM16, Lut84, Pot01a]. Revision [Dre17, Kin12, Bue76, PS96]. Revisions [CC22b, SNT^{+}14]. Revisiting [BESFW22]. revolution [Cap92]. rewards [AT88, Bro93]. rewiring [St.01c]. Rhetoric [AM07, Fil05, Flo04d, Jan08, Kni08, Kos07, Kos08, McC19, Rou17a, St.00a, Sta13, Tru88, Wee19, CR73, Dom95, Hea99, Heb97, Hor90, HWW04, KR95, McD94, Pod03, Rud92, St.99b, St.01a, Tru87, Wha02b, ZS04, Joh97a]. Rhetorical [Baa07, Boi18, DGPR10, ESPQ15, Fri13, Gar19, Gur95, HMF02, Jin18b, JW22, KI06, Lay97, Ley08, Lin07, Mel09, Rif10, Sch10, Ste22, VG16, Wol06, Cla87, Con92, Del92, Den88, For04, Her95, Meh94, Rob95, SF99, Spi94, Wei93]. Richard [ES98, Tay01]. Richards [Col65, Col65]. Richness [AZDA10, TM08, Bae12, ESM98, Koc05, RD05]. Right [KHV10, Den83a, Dor83, DV02, Dul93, Hug79, Kru00, Sar72, Szo82, Nob10]. Rightful [Fon61]. Righting [Cun09]. Rights [Dia07]. Riordan [Flo03d]. Ripe [Lam16]. Risk [Boi18, Cos23b, Flo04d, Hen22, LWMR07, SGM^{+}24, You10, Rou17b]. Risks [Dre88, AT88, Bro93]. Road [Flo04a]. Rob [Wei02b]. Robert [Mic80, Sch78c, Whi78c, Jon89, Sal61, Wha01]. Roberts [Cla12]. Robison [Hyd18]. robotics [MG82]. Robots [Kil09]. Rochester [Bra00]. Rohozinski [St.09]. Role [AD22, Ano63h, Ano63t, CHME07, Dia07, FR13, Fon61, Hag61, HW09, Joh20, KT14, KF77, KL12a, KSNF15, Lof08, MY14, MML15, Roy07, SS10, SC12, Wee19, Woe79, Bli72, Bow73, Boy87, Bre73, Cro86, DH04, DH73, FW78, Fis98, Fis99, Fla05, FO06, Fry62, Gui94, HTB80, Har84, HMF02, Lee00, MM00, McD94, PD05, Rub87, Sam88, San91, SH91, Ske90, Sur96, VVL^{+}01, Whi91, Win89, Woe75, Sch79]. Roles [FCS10, LLY18, Bur85, Con92, Sou88, TC95]. ROM [Moo02]. Ronald [St.09]. Rook [Kre87a]. RosE [Wil10a]. Rosenbaum [Ano97b]. Ross [McE61, St.02b, Sch83]. Ross-Larson [Sch83]. Rothschild [Hou21d]. rough [Bue76]. Roundup [Ano73-27]. Rowley [Bar09b]. Roy [Kie83]. Roze [Flo03c]. RSI [vHK08]. Rubric [Boe10]. Rudolf [Joh65]. Rule [Phe20]. Rules [KHV10, Sto15, Bro89, Hal80, Mac03b, Tib82]. rumors [Bas97]. Rune [Ram03]. Running [Ang59, Mur00f]. Rupert [Sal61]. Russell [Whi78e, McE61]. Russia [Goo94, Gre62, KM94, LS94, Lan94, Pav94, Ter94, Vch94]. Russian [PBK^{+}93]. Rutter [Whi78e].

S [Ano64e, Bal18, Ber12, Fer07b, Flo07e, Fre13, Mir99, Nor97, Sal80, Sch78e, Wha01, Wis15, Ama02]. S.D [GGS23]. sabotage [Bas97]. Sadler [Fer09b]. Sadler-Smith [Fer09b]. Sadri [Suo13]. Safe [Mar06, Vic03]. Safety [DR20, Vel87a, BS87, Con87]. Sage [All97]. said [Pro94]. sale [Wri79]. Sally [Whi80b]. Salton [Die72]. Same [Hay16]. Samuel [Ano80h, Kni86]. Satisfaction [BJ21b, vSdJ07]. SatisFAXtion [SH93]. Saucier [Bal18]. Say [Ano67e, McE58, Mon81a, Sti59, Den88]. Scale [Ven62, App01a]. Scandinavian [SSS09]. Scanning [Mac75b, DH92]. scenario [JCF92, DOHM96]. scenario-based [JCF92, DOHM96]. Scenarios [Bis99, VNC^{+}13]. Scenes [BS16]. schedule [Ano78r]. schema [KVRR06]. schemas [KVRR06]. Schizophrenic [Man86]. Schlesinger [Ano97b]. Scholarly [HLY22]. Scholars [HLY22]. Scholarship [Hov08, PD24]. School [FPKS^{+}21, Gir06, KS14, KSNF15, OW91]. Schreiber [Dur22, GG22a]. Schriver [Spy98]. Science [Bar98a, Bra59, CLP08, Dud71, Gir06, Gon20, GZ97, Jan08, Kil82, Rou17b, VNC^{+}13, Wha04, Ano87-28, Ant01, BM00, CI02, Caw75, CE05, Eng89, Fre03a, Fry62, HT77, Jen02, Kae03, Kuo93, Luf72, Mal87, Mir99, Nee92, Orr99, Orr03, Pas60, Pri72, SB79, Sch78d, Str60, Sub76, Sub79, Tho75, Wal74, Wei89, Sta99, Zha21, Mac02a]. Science-Relevant [Gir06]. Sciences [Ber12, Kae05]. Scientific [Ano63-27, Ano75m, Ara59, Arn69, Bed12, Ber12, BW19, Bov73, Fer10, Goo97a, Kae05, Kun69, LO07, Rei99, Reu65, Sch66, Sch70, Til06, Wha04, Zha21, Abe73, Abr77, ABMF84, And84, Bas72, Bat02, Bea99, BJ79, Blo87, BGN^{+}95, Bra78, Bre73, Bur75, Cha60, Che92, CR73, Day77, Dud62, EM78, Est82, Gar73, Gar79b, GV82, Gup83b, HM73, Har89, Hor75, HR60, JVT01, Kes60, Koc77, dALT92, Lin72, LM97, Man75, Mic80, O’C62, Pal75, Pfa86, Pul86, RCK^{+}90, Rei02, Rob73, Rot78, Row75, Rud92, Rus77, Sch78d, Sch80d, Sch75d, Sha89, Sut77, Tho73, TS81, Vis77, Wal74, Coc04, Pav94, Fer12, Kir20, Sch79, You11]. scientist [Her62, HS72]. scientist-engineer [HS72]. Scientists [Flo05c, Flo18, Ano82o, Ano82p, Ano82q, Ano83-29, Ano83-30, Ano84-31, BPK94, Bli87, Bro79, Cow80, PBK^{+}93, Rob72, Sch67, Sch75a, Smi03, WP82, Smi03, Whi78c]. Scientometric [LHMN07]. scorecard [EP83]. Scores [AFH20, Sha82]. Scotland [Dos06]. Scott [All97, Kir20]. Scranton [BS16]. Screen [GvdM02, GB84a, GvdMdJP99, KA03, Mur00h]. Screens [HSA21]. SCRIPT [MM68]. Scrum [Fri19]. SDD [DBM73]. Search [Kil09, Kil13, LMJK14, SD06, Tay82]. seasoning [Her89]. Second [Ano58i, Ano71x, BBRT86, Cla12, Pet13, Flo00c, Kre86, McD94, WLMR82]. secondary [Met73]. secret [Noo79]. Secretary [Ano68u, Lee62, Arn69]. Section [Cot59, Goo97b, St.07, vdGRRv13, BTM12, Goo94, Goo95, Lan94]. section-issues [Goo95]. Sections [Jin18a, Jin18b, SA23]. sector [Fre13, Wha01]. Security [GYK14, Sha89, Tay73]. See [Sti59, Ker80, Tru87]. Seeing [SSS09, MB84]. Seeking [KS14, Lan05, LE18, XZZ10]. selected [Ram89]. Selecting [BS98, Mac82, Mac78b, Rac59]. Selection [Men23, MKBF09, DR94]. Selective [HK70, BB75]. Self [AFH20, BP15, CHME07, GH85, NP23, SSS^{+}13, ZNG22, ZWE21, vdGRRv13, Gui94]. Self-Directed [NP23]. Self-Disclosure [ZNG22, GH85]. Self-Efficacy [ZWE21]. self-esteem [Gui94]. Self-Organizing [CHME07]. Self-Presentation [BP15]. Self-Service [SSS^{+}13, vdGRRv13]. Sell [CNC20]. Selling [Wei89]. Selves [Whi06]. Selves/Writing [Whi06]. SEM [LG14]. Semantic [CC88, KL06, Sab61, Son65, Sto09]. Semiconductor [Loe68]. Seminar [Ano63i, Ano63v, Ano63w, Fou59, Mey64, Eal83, Ano64k]. seminars [Bai81a]. semiology [Mar89b]. Semiotics [ZFWX21]. Senate [Pod03]. Senior [Car73, Bur84a, NH05]. Seniors [LKS13]. Sense [MVV09, Hoy58]. senses [Hoy58]. Sensitive [WD09]. sensitivity [Del92]. Sentence [Bra78, Hou21a]. Separate [Sta73]. Sequential [Roy07]. Series [Ano86m, Ano08u, Ano09v, Ano09w, Ber12, BS16, Red98]. Serve [SWR21]. Service [Ama04b, Bay22, Cho18, Doa19, HW09, JJ10, MP20, Mog14, SSS^{+}13, Yu11, vdGRRv13, Har97, Mat77]. Service-Learning [Ama04b]. Services [ASHB16, BIdJ^{+}21, GKD77, Koe77, ODF92, Tel68, Rok19]. Session [Ano73-27]. Sessions [Fri13, Ano79e, RM03]. Setting [vdHdJS06, vdM07, Jus72]. Settings [GE13, LNB^{+}04, Nor00, Swa05, Ede87]. Seven [Ril99, Dou02a, Rat85]. seven-plus-or-minus-two [Dou02a]. Sex [SSSC08]. sexism [Ber92, War80]. sexual [RH92]. SGML [Gol97, MSW97, Wei93]. SGML-based [MSW97]. Shades [Hun98]. Shape [Tru91, SD06]. Shaping [And14]. share [Moo87c]. Shared [Lam68, Sap08]. shareholder [Bro89]. Sharing [HK07, PS12, RH01, SEN10, Bae12, TYT91, Wis15]. Sharlat [Bal18]. Sharon [Bai24, LD20]. Sharpening [Bli84]. sheetfed [Ste77]. Ships [Fil05]. Shipwrights [Ram00]. shock [Cor92]. Shooting [Ath22, Ste22]. Shopping [HBS08, LFT08]. Short [Flo04a, GE13, Van71, BS89, Irw72, Rob78]. Short-Message [GE13]. shortcuts [FS92]. Should [Luf69b, Cam96, Fei79, Luf75, Nag87b, O’C62, Pat75, Lou58]. Show [LS23]. Showroom [Zim13]. sic [Sau93]. side [Den83a]. Sidney [Mic64]. sight [Sch72b]. Significance [Roy07, Dud62, SSM74]. significant [BR93]. signs [Von78]. silence [Loe91]. Silences [ELdJ12]. silent [Has80]. Silicon [Mur00k]. Siling [Car16]. Silvano [Sch80a]. Similar [Car12]. SIMM [Pla71]. Simon [Flo03c]. simple [Rei81]. simplified [ZB84, SHS97, Lin04]. Simplifies [Roc59]. Simplify [Sie78, Cra92, Cus93]. Simulated [KCJC08]. Simulation [BR16]. simulations [Fre03b]. since [Ano08t, Ano09t]. Single [Bis94, Mon81b]. Single-source [Bis94]. Site [VG16, vDVvL19, ACH^{+}98, App02, Boc98, Har97]. sites [Ama03a, App01a, Dur99, RL96, Ril02b, RK91, Wei98, Flo03b]. Situated [RKP00]. Situation [Lin07, Die04, Dav74]. situational [Nel90]. Situations [Boi18, Hen22]. Six [Hov08, Mon80a, Blo83]. sixth [Fre13]. Size [Hun21, BMH05, RLS06, Wal85]. sized [MG95]. Skees [Bro83a]. skilful [Wei95]. Skill [Tri16, GH85]. Skills [Bed12, BC17, Dud71, Flo06c, KS14, NP23, She06, vDVvL19, Bro79, EA84, Fla93, Ham96, HCKJ03, Pot84, Pow80, Rai86, Smi85, Tsu91, Wri81, Yor78, Ano63d]. sleeping [Cox93]. slide [Dav81b, Luf68, Sto15]. slides [Rei78, Rei79]. slip [CDL84]. Small [HSA21, Ven62, Fre13, KA03, Max88, Wha01]. Small-Scale [Ven62]. small-screen [KA03]. SMART [Die72, AZ17, DR10, HSA21]. Smartest [Hou22a]. Smith [Ano63-27, Bar09c, Fer09b, Zho08, Sch78e]. smooth [Sch93]. SNA [HLY22]. Snobs [Nob10]. Snow [Luf64]. Sobhani [Flo07b]. Soc [The97]. Social [AZDA10, Ant18, BAB10, BR16, BESFW22, Bar09c, Ben22, Bha12, Car24, Cha96, CWSL19, Che17, CC22b, DW07, Gan22, Gen10, Gra22, HZX24, HK07, JW17, Lin19, Lin77a, Per93, PD22, Rea22, Rob73, STP22, Sim22, SvI18, Sta20, SWR21, SZLY22, Waa09, Wal23, Wel22, YFW15, Yu20, Yua21, ZNG22, ZC24, vDVvL19, Adl00, BT95, Cro95, IPC01, KR95, Nye99, Per96, Rei02, RLS06, Fer03, Mir99]. Socially [Hov08]. societal [Ter72]. Societies [Tom69, Whi06, Ano73y, Koc77, Lid75]. Society [Ano80b, Ano81a, Ano82a, Ano82b, Ano83a, Ano22c, Par75, Reg70, IPC01, Moo87a, Pot77, Rot78, Wei89, Ano78d, Ano78l, Ano78m, Ano78o, Ano78n, Ano79b, Ano80c, Ano80u, Ano80v, Ano80w, Ano81c, Ano81m, Ano81l, Ano83c, Ano84b, Ano85d, Ano86c, Ano86a, Ano87b, Ano96, Ano04n, Ano04o, Ano04p, Ano04q, Ano05k, Ano05l, Ano05m, Ano05n, Ano06o, Ano06p, Ano06q, Ano06r, Ano07p, Ano07q, Ano07r, Ano07s, Ano08o, Ano08p, Ano08q, Ano08r, Ano09o, Ano09p, Ano09q, Ano09r, Ano10o, Ano10p, Ano10q, Ano10r, Ano11e, Ano12r, Ano12s, Ano12t, Ano12u, Ano12v, Ano12w, Ano13n, Ano13o, Ano13p, Ano13q, Ano13r, Ano13s, Ano13t, Ano13u, Ano14p, Ano14q, Ano14r, Ano14s, Ano14t, Ano14u, Ano14v, Ano14w, Ano15p, Ano15q]. Society [Ano15r, Ano15s, Ano15t, Ano15u, Ano15v, Ano15w, Ano16h, Ano16n, Ano16i, Ano16j, Ano16o, Ano16k, Ano16l, Ano16p, Ano16q, Ano16m, Ano16r, Ano16s, Ano17m, Ano17q, Ano17r, Ano17n, Ano17s, Ano17t, Ano17o, Ano17u, Ano17v, Ano17p, Ano17w, Ano17x, Ano18j, Ano18n, Ano18o, Ano18k, Ano18p, Ano18q, Ano18l, Ano18r, Ano18s, Ano18m, Ano18t, Ano18u, Ano19a, Ano19d, Ano19b, Ano19e, Ano19c, Ano19f, Ano20c, Ano20h, Ano20d, Ano20i, Ano20e, Ano20j, Ano20f, Ano20g, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano24i]. Society [Dar82]. Sociocultural [Fer04b]. Socioemotional [FCS10]. Sociotechnical [Tha24]. Soft [CNC20]. Software [Bar90, Boe76, DL17, DR10, Fer04a, Fri19, GQ12, HdJK18, JDJ99, MKBF09, RA97, SJD20, Waa09, ZB84, Ant99, CM02, Del00, Den98b, Duf87, Fer03, Gla04, Har97, NKU03, Pak81, PDK88, Pot01b, RLS06, SH91, WB92]. Solution [KT18, KT20]. solutions [Mar86]. Solving [Ana21, Fer04a, Gla04, JES13, Bri99, Flo02b, Fri73, Hac84, Sko83, vdM96, Shr20, Ano63d]. Some [Con58a, Fra75, Goe78, KS04, Lip83, MGB89, Mil62, Rin59, Ros83, Sha68, She73, BS98, Cor90, Gri91, KL06, STM09, Sou72, Sou75]. Sophisticated [Woo79]. Sophomore [KC81]. sound [Sch72b]. Soundspel [RL78]. Source [O’D07, Waa09, Ano85z, Bis94, Rob95]. Sources [Ano65n, Coi69, Pic80]. South [Ach24]. Souther [Whi79h]. SOVAM [Ter94]. Space [Hoh60, MPD10, Ano87y, Bar77, Cod04, ML99]. spacecraft [Mor67a]. spaces [JE95]. Spain [WARAL16]. Spaniards [Gri94]. Spanish [Izq24]. Spanner [PS12]. spatial [HTB80]. Speak [Ano63i, Sti59, AB81, Ken88, Ken89, Lub80, Mar85, Pei81, Shu93, Wol82, McE61, Flo10]. Speaker [YJY24, Bon83, Den00]. speakers [Has80, MO80, Orr03, Per78, RM03]. ’Speaking [Ken89, Gro95b, Ken88, KS14, Lam05, NP23, Bai81b, GB79, MD84, Win83, Dou01]. Spear [WHC^{+}12]. spec [Win99]. Special [AB15b, Ano77-30, Ano83l, Ano04s, Ano09x, Ano13v, Ano15h, Ano15i, BA16, BI18, CS06, Goo97b, Hat69, Lac63, Löc22, MMW17, MR21, Orr06, PM08, Spi16, St.07, Sta08, vdGRRv13, Ano82e, Bra82a, BTM12, DF87, ES99, Goo94, Goo95, Gro05, Hol74, IPC01, Joe80a, Joe81b, Joe81c, Joe81a, Joe82a, Joe82b, Joe83a, Lan94, Orr03, Pri97, Ste04, SB00, UVB02, WS04]. specialist [Aug80]. specialists [Bli87, Shi81, Shi82]. Specialization [Gei77]. Specialized [Mis05]. Specific [JD21, Orr01]. Specifications [Loe68, Str64, Lee64, Pil87]. Speck [Mal11]. Spectrum [Per63, Ano63j, Ano70b, Bos64]. Speech [Ada81, Ano61a, Ano64s, Ano65g, Ano66g, Ano66h, Ano66i, Ano66j, Ano67j, Ano67k, Ano68h, Ano68i, Ano68j, Ano68k, Ano68l, Ano68m, Ano69j, Ano69k, Ano69l, Ano69m, Ano69n, Ano70i, Ano70j, Ano70k, Ano70l, Ano71n, Ano71o, Ano71p, Ano71q, Ano71r, Ano71s, Ano77-31, Cla59, FS80, Gol58, Han67, MPD10, Sch81a, YK13, Bos87b, Cur80, Dec84, Dor83, Flo81, Gre69, Han75, Lin90b, Mic62, Pie58, Whi79c, Ano58i, Ano59m, Ano59n, Ano59o, Ano60a, Ano61u, Ano61v, Ano61w, Ano62j, Ano63m, Ano63n, Ano64t, Ano64u, Sal61, Ano71a]. Speech-Their [Gol58]. speeches [Cos78]. speechmaking [Shu93]. Speed [Ano68v, Lev68, MP79, Sau93]. speeds [MM73]. Spellbound [Hav79]. Spelling [GHR67, Fur90, MC75a, Cun09]. Spencer [OM12]. Spilka [Wil03]. Spill [YFW15]. Spinuzzi [Yea10]. Splatter [Car61]. Spoken [DSV10, Dre87, Lin80]. Spoken-Language [DSV10]. Sports [WC15]. Spring [Ano13v]. Squares [LG14]. St [Boo17, Hin10, Waa09]. staff [DSF^{+}80, JP86]. Staffing [Fie65]. Stage [PH24, SAP20, LD84, Sho80, Son65]. stakeholder [HCKJ03, ZS04]. Stakeholders [FDJW18, Gal01, MOA02]. Stakes [DSV10, Til06]. Stalinist [Sta13]. Stance [Dos06, HZX24, VMM21]. Stand [Bro64]. Standard [Ano77c, Bun71, Con58b, DR10, SP87, Ano65h, Ano65m, Ant99, Gui89a, Mou75]. Standardization [Bro75, HP86]. Standardized [Sal68]. Standards [Hac16, KRM24, RFH87, Str21, Til79, ZJG17, Dev75, Hac85, HP86, Whi79g]. Standchevici [Sta13]. Star [Mir99]. Starring [Jan08]. Start [SSV^{+}23, CW90, Pat76]. Start-Up [SSV^{+}23]. Starting [Mur00e]. Startups [WARAL20]. State [AB15a, GG23, GLL23, HK70, JW22, Mar64, WS23, Gre69, Kan78, Nag87c, Tra92]. State-of-the-Art [Mar64]. State-Owned [WS23]. Statements [PH24, Yu23, Tar01]. States [Con91, Dia07, Hov14, Mal07, Ros94, WARAL16, Ram00]. States/British [Hov14]. static [Man99a]. Stating [Kap62b]. Statistical [Arn62, BCR00, Lam16]. statistically [BR93]. Status [Ock07, Whi77a]. stay [Ano73y]. Steamboat [Fil05, Pod03]. steamboats [Pod03]. Steel [Cun10]. Stein [Riv85]. Stem [MAK15, AFH20, JD21].

Stem-Based [MAK15]. STEM-Specific [JD21]. step [Evi04]. Stephan [Zim13]. Stephanie [Ano97b]. Steps [LKS13, Gol82a, Mon80a, Pot82]. Stereotypes [Sta17]. Sternberg [Mic85a]. Steven [Flo03d]. Stevenson [Sch78b]. Still [BC70, Waa09, Ort78]. Stimulating [Rau83]. sting [Woo80]. Stop [Ste22, Tra83, Woo81, Tra92]. Storage [Kap62a, Sch70, Sed69, Sch75d, Yer74]. Stories [LKS13, WARAL16, EC01]. Storms [Bea99]. Story [Cun09, AA93, Cap93]. Storyboarding [Bar89, Eng72]. Strands [Red10]. Strange [BCB03]. Strategic [Lie20]. Strategies [AD88, AW22, Bea98, CSSV20, Cla12, Dia14, Fer07a, FCS10, Flo07c, Hen22, HS19, Lie18, Rem78, Rif10, Riv85, SHB08, SNT^{+}15, App01b, Bar91, Flo01a, For04, Fre13, Gre85, Hac84, LY98, Mac02a, Mag03, Pot00a, Spi94, St.02c, Wal87, Wha01, Car90a, Car90b, Erv99a, St.03, Cha80b, Flo06a, Jen99b]. Strategy [Ano67i, BA16, Cla16, GL16, GPHDM16, Gra75b, Lin90b, Nor00, RK94]. Strauss [Mek72]. Streamlining [Rei88b, Bat02]. Stress [Hen22]. Strike [BG16]. Striking [Ort78]. striving [Cam92]. Structural [LG14, ZLCZ17, Fes62]. Structure [Etz63, GH09, Kan12, Kin12, Koe77, PH24, YB17, Den88, Har89, LNCL05]. Structured [CFMM14, ESPQ15, Cra92, GN94]. Structures [Sab61, FWB86, Heb97]. Structuring [Joh10, DJ82, JD83, Pri97, Riv85]. Student [BW19, CL19, DPHF17, DB24, JD21, LC17, PEG19, SA23, She06, SAP20, YTY12, Yu12, Bar82a, JDJ99, KK02, Nel03, Ser88, Wol88, Zim88]. Student-Friendly [JD21]. Students [AFH20, BJ21a, CCL24, Dud71, FW23, FCS10, KT20, KS14, KSNF15, KRM24, LO07, Lie18, LDKC23, MP20, PEG19, Phe21, PD22, Sap08, VNC^{+}13, WP18, Wil10a, WFSC^{+}23, ZWE21, Cor92, DH92, EA84, Fla05, Jen00a, Ken88, Ken89, Kuo93, Mac05a, Mal87, McC94, MSC96, Mic84, Rai86, Sou88, Wal99, Yee88, Dur22, Smi03, Whi78e]. Studies [BH15, BH21, Ben22, BP09, CR06, CLP08, DuB08, GY68, Joh97b, KZCR10, LAWI24, Mal07, Moo16, PMGW21, Rou17b, Sch66, SCW20, SA10, WS04, vdM97a, Ada06, GBF05, Gro01, St.01c, Jan08, Bec84]. Studio [Wat20]. Study [Ano65n, BJ21a, BW19, Bre10, CBSM^{+}09, Cho18, Cla59, Cog15, Cos23b, DB24, DC07, Dre17, DGPR11, FR13, Fri15, Fri19, FL23, GAGW15, GYK14, GH70, HvMK^{+}12, Hov08, Hov14, HZX24, KI06, KC19, KCJC08, LHMN07, Mel09, Men23, Moo16, Raj12, Rei99, Roy07, Sam06, Sap08, SSS09, SDB^{+}13, She06, Tho70, Tri16, WC15, Wil24, WARAL20, WRTBRL23, YK13, Yu12, ZM23, BBRT86, Cho97a, Cod04, Con91, DR94, Gat04, Goo05, Gri94, HM95, Koc03a, Koc03b, Kus97, Lin90a, Lin93, Lut84, Mac97, Mac03b, MG95, Mal87, Mon78, NKU03, NB04, ODF92, PS96, Pit79, Por87, Rub86a, RH01, SB79, SSNJ05, Ste99, SH01, Tay73, TSS^{+}01, Zim93, BR16]. Studying [GQ12]. stuff [Bos84]. stuffy [Whi87]. Stull [Ale20]. Style [Bra05, CAK^{+}17, DGORM84, HvMK^{+}12, Hol03, LJ10, Mic61b, Rav06, WR83, Cam92, Car90c, CJ89, Jon73, MQ76, Ort78, PM98, Pil87, Plu84, Ras86, Rav88, Spi86, Woo80]. Styles [BH15, Del00, Gri94]. Stylistic [HB94]. Subassemblies [Roy07]. Subdisciplines [Kan12]. subgroups [PD05]. Subject [Ano77-32, Ano99b, Ano01b, Ano02b, Ano03b, DBK^{+}93, Hin72, Ter73]. Subjective [San87b]. subjects [Van63]. Sublime [Sim96a]. submission [SO98]. submissions [Bau62]. submittals [Ano86u]. Subscription [Flo75, Bak75, Mar77, Wil75]. substance [Woo80]. Success [TJQN11, Bak82, Cox93, Eva60, GP86, Gol82a, Jac81, Ken82b, MD84, Ode86, St.02c]. Successful [GMT08, GYK14, KB84, Ram07, Cha81, Eal83, Flo01a, Jin86, Tur79, Wei95, Wha01, Fre13, Rob78]. suck [Mur00g]. sudden [Cap93]. Suddenly [Flo04b]. Sue [Erv99a]. Suggestions [CH89, Yu10, AW78, Mac05a, PM84]. Suited [Joh10]. summaries [Bau62, Sek73]. Summary [FWJ77, Wen69, Ano73g, Bro89, Wen85]. Sunken [Fil05]. Sunwest [FO06]. Super [MOA02]. superhighway [Cro95]. supervisors [Ano80o]. supervisory [Bli76]. supplied [Hol77]. Support [CWSL19, FW23, Fie65, Joh20, Rob18, WP18, vSdJ07, BD83, BCB84, GvdM02, KW93, NKU03, San91, vdM96]. Supported [CHME07, CS96]. Supporting [BSS13, DSV10]. supports [Pul86]. Survey [Ano63z, BESFW22, CFMM14, KB17, MEZ12, Mil62, Rif10, BR92, Cho97b, Com86, Dav85, Dre75, FWJ77, GHG91, How83, How84, Kre00, Lac87, Rav89, VLA96, Tro66, Den98a]. Surveys [CDM10, DC07, BGN^{+}95, BK03, STM09]. Survival [Max73a, Max73b, She06, DH73, Eck03]. survive [Mic85a]. Susan [ES98]. Susceptibility [WHC^{+}12]. suspicion [Wei95]. Susskind [Mic61b]. Sustainability [Yu23]. Sustainable [LL09b]. Sustaining [GG22a]. Svartvik [Sch78a]. Swarts [Joh20]. Sweat [Ano67e]. sweet [Whi87]. switch [Kor77]. switching [BS98, GvdM02]. Symbiotic [JD23]. symbolic [Ama03b, Pla71]. Symbols [Ano68w, Coo65, Loe68, Ano65h, Bur90]. symmetrical [Chi89]. Symposia [Ano59i]. Symposium [Ano58i, Ano60a, Ano64-29, Phi69b]. Synchronicity [MVV09]. syndrome [Sea73]. Synopsis [Ano58j]. Synoptics [DBM73]. Syntax [Nob10]. synthesis [BCB03, Sch80a]. Synthesizing [KZCR10].

System [And14, BR68, Bec68, Boe76, CWSL19, Cog15, DM63, DOHM96, Fre68, Hou21a, Kun69, Lam68, MM68, Mar70, SDB^{+}13, Sus61, Wer71, YK13, Ack89, Bah85, BP82, Bar77, Cas86, CS96, Cru73, Die72, Doo79, FWB86, Fra89, FS92, HE77, KS04, KAOC81, Koc75, MSW97, MOA02, O’C62, Ora86, Rab92, Rav88, Row75, San91, Stu69, WP82, CD98, Fre68, Mak68, Wil10a]. system-initiated [FS92]. System/360 [Fre68]. Systematic [Fri73, ZNG22, TS81]. Systematization [MV21]. Systemic [Kin77]. Systems [AZ17, Ath22, BCJR87, Doe69, Lie18, MFL17, Mey87, Pla71, Sab61, Sch66, WMPBA07, WJ13, XZZ10, vdGRRv13, Ano77-31, BBRT86, Bha77, BR92, BMRH05, Deg80, DEBD83, Evi04, Fis98, Fis99, Fle85, Hac85, LS94, Los85, McG86, Mir99, SSNJ05, ST05a, Ske90, Sko79, Tho73, Til79, Yer74, Joe79a].

’t [Max88, Bar09a, Boo17, Col65, Den83c, Den83b, Dur22, ES98, Fer10, Har80b, Hen22, Mas08, Rob78, Sal61, Sta13, Sto15, Whi77c, Whi78b, Whi79f, Wis15, You10]. T. [Bro83b]. Table [Ano59p, Ano59q, Ano64-30, Ano64-31, Ano65o, Ano66n, Ano66o, Ano67n, Ano68x, Ano68y, Ano68z, Ano69u, Ano69v, Ano69w, Ano70o, Ano71y, Ano71z, Ano71-27, Ano04t, Ano04u, Ano04v, Ano04w, Ano05o, Ano05p, Ano05q, Ano05r, Ano06s, Ano06t, Ano06u, Ano06v, Ano07u, Ano07v, Ano07w, Ano07x, Ano08v, Ano08w, Ano08x, Ano08y, Ano09y, Ano09z, Ano09-27, Ano09-28, Ano10t, Ano10u, Ano10v, Ano10w, Ano11g, Ano12-29, Ano12-30, Ano12-31, Ano12-32, Ano13w, Ano13x, Ano13y, Ano13z, Ano14k, Ano14l, Ano14-28, Ano14-29, Ano14-30, Ano14-31, Ano15-33, Ano15-34, Ano15-35, Ano15-36, Ano16v, Ano16w, Ano16x, Ano16y, Ano17-29, Ano17-30, Ano17-31, Ano17-32, Ano18x, Ano18y, Ano18z, Ano18-27, Ano19k, Ano19l, Ano19m, Ano20o, Ano20p, Ano20q, Ano20r, Ano21m, Ano21n, Ano21o, Ano21p]. Table [Ano22n, Ano22o, Ano22p, Ano22q, Ano23n, Ano23o, Ano23p, Ano23q, Ano24j, Ano24k, Ano24l, Duk89, Get86, KA03, Ano63x, Ano63y, Ano12-33]. Tables [Arn62, Bue77, BS98]. Tabular [Boy68]. Tacit [App02]. Tackling [Kus97]. Tactics [Gra22, LO20, LS23, Bat80]. Tagging [Bar09c]. Taiwan [Che92, Lie18, Tsu91]. Take [Hin72]. Taken [SSW17]. Taking [Car61, VMM21, Ano86-27, Mur00h, Nag86]. Tale [LCX12, NNB^{+}22, TVRR20, PT05]. tales [St.02b]. Talk [Bah85, DV09, Van59, Fla93, Lou58, Luf68, NC81]. Talking [Bra59, Bol82, Flo01a]. Talks [YJY24, Ada80, Mac78a, Sch80d, Sei67]. Talmud [Wei98]. Tangled [Cun09]. Tapes [Wen69, Ese69]. Targeted [WHC^{+}12]. Targeting [Kil09]. Tarnow [Pen98a]. Task [FMT12, GG13, KL12a, Rat61, ES98, Fry87, JCF92, Par86, RLS06, SJ99, Str99, Wat90, vW00]. tasks [Cap93, Kru00, SI06]. taught [Sid87, Vel88]. Taxonomy [AAA19]. Teach [Bed12, BS16, Mil61, OL98, Wat20, BKSM82, GRS96, Jen99a, Jon89, Kry99, Par06, Pot84, Wal99]. teachers [Fey88, St.01a]. Teaching [Ano08j, Ano09k, Ano09j, Ano09l, Ano10j, Ano10k, Ano10l, Ano11c, Baa07, Bar82a, Bed12, BI18, BMS^{+}15, CLP08, Dum96, ESPQ15, FMB18, Flo05b, FO06, Gwi84, Hac84, Han67, HS19, HWW04, Kni08, KHV10, Kre89, Kuo93, Lee13, Mic84, Phe21, Rat61, Shi82, SMP21, Spi96, VG16, Ven62, WT09, Yu12, ZS04, Cox93, Hag90, Han75, JM01, Ken88, Ken89, LL03, McD94, Pal00, San87b, Sch75b, SF99, Sub92, Bro79]. Team [BN10, BMH05, BH16, DL17, DR20, FW23, Hov14, KL12a, Lam18, PDL16, PFD^{+}21, SSS09, SV06, WLM^{+}22, WPSB13, XZZ10, ZWE21, AC85, Dou72, Dro80, Fla05, HSV05, Kol93, Lee64, Loe95, Nor00, Pot84, RH94]. Team-Based [ZWE21]. team-writing [RH94]. Teams [AAA19, AG21, BN07, Bed07, BMS^{+}15, BH16, CL19, CHME07, DPHF17, Fer07c, FW23, FCS10, FMT12, FVB16, GDTZ09, KC19, KL12b, Ock07, PHP16, PHW13, RVvGS08, SSS09, SC12, VT21, WPSB13, Zem12, AC06, BM00, BMH05, CHR04, Gil03, Kol93, Loe91, LN03, LNCL05, MM00, Ock05, PD05, RKP00, SSNJ05, SV06, SB00, SH01, TWHN00]. Teamwork [Flo01c]. Tears [Ano67e]. Tech [DV09, The97, Eck03, Mun02]. Technical [Ada80, AD22, Ama04b, Ana21, And81, AME21, Ano63c, Ano63l, Ano63m, Ano63n, Ano63-27, Ano63z, Ano64u, Ano67a, Ano97d, Ant06, Arn69, Arr74, Baa07, Bar14, BPK94, Bat18a, Bay22, Bed07, Bed12, BH21, Ben22, Ber12, Bis70, Boc98, Boe76, BP10, Boe10, BL13, BI18, BW19, Boo17, Bos84, BB91, Bou62, Bra05, BSR85, Bro62, Car12, Car20, Car61, Car80, Car24, CG83, Cha23, Cho18, Cle16, CC22b, Cog15, Coh69, Con92, Con58a, Coo10, Cos23b, Cox80, Cun10, Dav67, Dav10, Daw10, DF70, Dia07, Dia14, Dom07, DL24, Dor63, Dre16, Dwo69, Eri11, Far64, FG84, Fer10, Fie65, Fle85, Flo02d, Flo04d, Flo05b, Flo06d, Flo06a, Fon61, FL23, Gan22, GLL23, Gir60, Gir06, GLCA22, Goo97a, Gre17, Gre70, Gre85, GZ97, GG22a, GG22c, Han67]. Technical [Han75, HM61, His77, Hou21b, Hou23, Hue02, Hun98, Hun21, Jen99b, Joe82b, JES13, Joh17, Joh20, Kae04, Kal97, Kau88, Kil09, KM61, KHV10, Kni86, KB17, LM02, Lam14, Lam16, Lam18, LBB18, LP16, LMJK14, Löc22, Lof10, LDKC23, LAWI24, LHMN07, LN98, Mac04a, Mag77, Mal07, Max88, McC65, McC94, McC63, Mic61b, Mic65, Mil64, Mil61, MSD85, Mog14, Mun02, Mur58, Nad96, Nic59, Nor00, Pad07, PD24, Pav94, Phi69a, PDD^{+}24, Pod84, RM61, Räi20, Ram98, Ram00, Rat61, Rav06, Red10, Ric09, Rin59, Rob18, Ros64, Ros15, ROW19, Rou19, SM66, SV24, Sch68, Sch66, Sch70, Sha68, Shr72, ST05a, SSW17, Sou96, SPR84, Ste05, SA10, Sto15, Str60, Swa18, Tha22, Tha24, The97, Tom69, UC01, Van71, Van62, VG16]. Technical [Ven62, VMM21, WMPBA07, Wea85, Wen69, Wil03, WS04, Yea09, Yee88, Yu10, ZWE21, Zho08, ZM23, Adl00, Ama02, Ama04a, ABMF84, And84, Ano71d, Ant79, AL03, Axe73, Bar98b, Bar82a, BJ86, Bat02, Bau62, Bea99, BW90, BS87, BJ79, Bis99, Bis90, Bis96, Blo87, BGN^{+}95, Bov73, Bow73, Bra93, Bra82c, Bre73, Bro80, Bro65, Bro83b, Bur85, Bus91, Cam95, Cap78a, Cas84, Cha80a, Cha80b, CW90, Che92, CS96, Cla75, Cla87, Coh60, Col77, Com86, CC88, Con91, CE05, Cor83, Cor94, Dav93, DGORM84, DH73, Del92, DF75, Den88, Den00, Den78, Den83b, Dob75, DF87, DBK^{+}93, Eal83, EP83, Eck03, EC01, Eng89, EM78, Est82, FS80, Fes62, FS85, FD84, Fis98, Fis99, Fis92, Fla93, Fla05, Flo01b, Flo01d, For04, Fra89, Fry81, GRS96, GG84, Gil87a]. technical [Gle82, GG96, Gre62, Gro01, Gup73, Gup83b, Hac84, Hag90, HN93, Ham72, Har80a, HH93, Har00, Hel88, Hin72, Hol76, Hol74, Hor89, How83, How84, Jen00a, Jen00b, JM01, JGH82, Jon89, Kan85, Kap62a, Ken64a, Ken88, Ken89, Ker77, Kie82, KE86, KR95, Kre89, KC81, dALT92, Lay88, Lip04, Lit90, Luz05, Mac90, Mac02a, Mac02b, Mac06, MGB89, MK84, Man99b, MM00, Mar93, Mar97, Mar84, MV83, Mas84, Mas85, Max73a, Max73b, McG00, Mey87, Mic81, Mic85b, MO80, Mil81, Min69, Mir99, Moo02, MC75a, OW91, Ort78, Par87, Pat76, Per70, Per78, PP97, PBK^{+}93, Pit79, Plu81, Plu84, Pod03, Por87, Pot77, Pow82a, Pow81, RCK^{+}90, RFH87, Rat85, RR93, Ren96, RC89, Rub87, Sac01, Sar73, SSM74, Sch81a, Sch72c, Sch75b, Sch78b]. technical [SB84, Sch80d, Sch75d, Sel99, Shi81, Shi82, Ske90, Sko81, Smi60, Smi85, SB95, Sou72, Sou75, St.99b, St.00b, SS87, Str84, Sus85, Tay73, Tay02, Teb80, TL92, Ter79, Tha04, Tho01, Ulr84, Van83, Vel77, Vol82, Wei01, Wha83, Whi91, Whi77c, Whi79e, Whi79g, Win88, Woe79, Wri81, Zim83, Zim93, Zim01, ZS04, Flo03d, Flo03c, Kre87a, Mag03, Nor97, Ano63-27, Ano71c, Dyr98, Flo00b, Flo02b, Flo02c, Flo05a, Flo07b, Flo07c, GGS23, Hof80, Hol03, Hou21d, Joh02, Kry00a, LeC01, Mic64, Sta20, Wha02a, Whi78d, Whi79f, Whi79g, Whi79h, Xu07]. Technical-Support [Rob18]. technical/scientific [BGN^{+}95]. technicalese [Ano78e]. Technically [BA83, Ano73y, DG75]. Technically-write [BA83]. Technicians [LE18]. Technique [Bro62, Ken64b, Lam16, Mil64, SM10, Tay59, TWHN00]. Techniques [AW78, Cla59, Hin72, Kan94, Kil13, MCC62, Mor66, RM61, SR83b, Bli87, Fri73, Hac85, HC88, ST06a]. Technocrat [Ram07, Cha81]. Technological [CC19, Hat69, McC19, Sim96a, Waa09, Sau93]. Technologies [LP16, LMJK14, RWAK24, Ric09, Win97, GS85, Gro95a, HP88]. technologized [Dom95]. Technology [Ach24, Ano08s, BH15, BN10, CWSL19, CHME07, DV09, DB24, Dia14, DOHM96, FS85, FCS10, Gra96a, HK22, HK07, JJ10, Joh97b, JSZ07, Jos70, KNMM09, KL12b, Lam05, LLY18, Mac68, Mac75a, MV21, Mar70, MP20, Mer05, Mou05, Mun05, Rou17a, Rou17b, Rui20, SV24, SA23, Sha68, WPSB13, WARAL20, Wil98, vdM97a, AA93, Ano04s, Ano14-32, Ano14-33, Ano14-34, Ave75, BMH05, Fie77, HM94, Hoh60, JM01, Joe81a, Joe82a, JD83, Kil82, LF95, Lay88, Luf72, Mac02a, Max88, MQ76, MYV^{+}91, MYV92, Mic81, Nee92, Nel03, ODF92, PDK88, Per93, Per96, Phi69b, Pod03, Pri72, Rob95, Ste04, Sau93, ST99a, SV06, Sut77, Tsu91, War88, WH82, Zur75, Cha96, Whi78b, Zim99, Pic80, St.01b, Whi79i]. Technology-Based [Gra96a, War88]. Technology-Mediated [DB24, KL12b]. Technology-Powered [SV24]. Technology-Related [WPSB13]. Technology-Supported [CHME07]. Ted [Whi79c]. Teens [CBSM^{+}09]. telecommunications [RL84, Ter94, Yea10]. Telecommuting [HR99, TJQN11, Pry95]. Teleconnections [Lev86b]. telephone [Com86]. TELEPORT [Ter94]. Television [BS16, Bar93, Cra85b, GS85]. Teleworkers [DLVE23, ST99a]. tell [Gar73, Hin72, dJS00]. template [St.85]. templates [SD06]. temporal [KVRR06]. Temporary [KC19, Lin99]. Ten [CvdM96, RVvGS08, Tib82, WR83]. Tenth [Xu07]. Teradata [St.85]. term [GG84]. Terminal [Fie77, Fre68, Lut84]. terminals [Ano78r]. terminological [Lip75]. Terminology [Mai70, Lip75]. Terms [Ano70b, Bun71, Ano65i, Ano74i, Che92]. Territory [KT18]. Test [AFH20, Coo10, DSV10, Mor67a, HR99, Hub89, Kan94, Str04, TC95]. Testimonios [PD22]. Testing [Bar98b, Cho18, Doc61, JSZ07, LG14, May19, MEZ12, Sch86, Sou96, SM10, Wil24, Eva02, KU04, Ram89, Ros89, SH93, SRVR97, Sul89, WS89, vW00, vdM97b]. tests [RK94]. Texas [Gat04]. Text [AM07, BR68, Cha77, Cli63, Cor69, DSF^{+}80, GY68, Jan08, KG68, Red98, Roy07, Sam70, WKSH68, dGKNG06, vHK08, BS73, Bur85, Car86, Die04, Dia77, FS00, Har81, Har84, HC88, HMF02, Ils80, JD83, LD97, Rav88, SP86, Sch89, Spi94, SJ99, SF02, Swa05, WBW89, WS92, Zie69, Flo03a]. text-critiquing [Rav88]. text-editing [Har84]. text-focused [Sch89]. Text-Graphical [Roy07]. Text/360 [Zie69]. Textbook [KSNF15]. Textbooks [JM01, Wol06, CW90]. Texts [RCK^{+}90, dGKGN11, Bra78, DWP85, JVT01, LH00]. Textual [Pet69, TM15, GvdMdJP99, RP98]. Tham [Ana21]. Thanks [Ano20b, Cam07, Hay21, Hay22, Hay23, Hay24]. Thatcher [Yea09]. Their [BJ21a, Gol58, KCJC08, LDKC23, Mar86, Räi20, Rau21, Rub87, Str64, WFSC^{+}23, Con92, Cro80, HS72, HMF02, Moo87c, NKU03, Ort75, Sch67, Sch75a, VLA96]. theirs [Reh94]. Them [PHW13, Ada81, BD82, Gou73, Ste99, Vel87a]. Thematic [AAA19]. theme [San87a]. Themes [dGKNG06]. Theodore [Cha65, Mek72]. Theorems [Cod09]. Theoretical [Mar89a, Sha94a]. Theories [Gan22, HDKC08]. Theorizing [Yea10]. Theory [AW22, Cha96, Dud61a, Flo03a, HF84, Hea69, Hel07, Kel61b, LCX12, LG14, Mas84, Mas85, Nyc97, WX06, Ama05, ABMF84, BR00, Bos92, Del92, McG72, Son62, Tra83, Wal99, Win89]. There [Jus72, Lon14, dGKGN11]. They’re [Ada80]. Things [GH09, Rob96]. Think [Coo10, ELdJ12, Gon20, May19, MEZ12, Ada80, Flo07d]. Think-Aloud [Coo10, ELdJ12, May19]. Think-Alouds [MEZ12]. Thinking [Ana21, BR00, ESPQ15, KU04, MD83, McN58a, Tha22, vW00, Bis99, Fer03, Kre89, Mus78a, Rau83]. thinking-software [Fer03]. Third [JJ10, Hof80, Whi78b, Whi79f, Sch63]. Third-Party [JJ10]. Thistlebottom [Mek72]. Thom [Bra99c]. Thomas [Hau80, Sch20, Whi79d, Whi79g, Zim13, Whi79f]. those [CH89]. thought [Kno80, MS02]. thoughts [Cor90, Mil81]. Thousand [All97, Dec84]. Thralls [Fer08]. Threats [KCJC08]. Three [Boi18, Cas84, CLP08, Cur80, FDJW18, Kan12, MMW^{+}22, Mur00j, Dou02b, HCKJ03]. Three-Dimensional [MMW^{+}22]. threes [Mon81a]. throughout [Hof87, McG00]. Throw [Flo07e]. Throwing [BG16]. thy [Ama04a]. Tibor [Pen98h]. Ticket [SSS^{+}13]. Ties [BWPH12]. Tillery [Gre17]. Tim [Bec84]. Time [Has80, Lam68, Ste99, Whi91, WT09, AACF81, Wha86]. Time-Shared [Lam68]. Times [Gei77, Dul93]. TIP [GQ12]. TIP-EXE [GQ12]. Tips [Bol82, Edd83, Ker77, Blo83, Lip83]. Title [Ano77-33, Ano78q]. Titles [Bot70, Ant01, Ken64a, SSM74]. Tjarks [Flo07b]. Tjarks-Sobhani [Flo07b]. Tobacco [Her95]. Today [Flo07b, Mil61, PM98, Flo03d, Flo05a, Flo00b]. Together [Ano14-32, Ano14-33, Ano14-34, Eng89]. Toil [Ano67e]. Tong [Rui20]. Tongue [Cun09, Sch65]. Too [Cha78, KL12b]. Tool [Ant18, GQ12, LDKC23, WPSB13, AL03, BS89, DH94, Gou60, NKU03, PM98, Rat86, Spr97, St.85]. Toolkit [Ant06, Wil24, Flo08]. Tools [Boe10, Bur78, Cro59, Löc22, SV24, Sim22, SR83b, EZ88, HSV05, Koc05, LN03, Sha94a, Sha94b, Fer07c]. Top [Ben85, ZZL22, Ben79, De 81]. Top-Down [ZZL22, Ben85]. topic [Cox80]. Topics [BS16, Car20, DBM73]. Topping [Car86]. Torres [Mas08]. Torres-Coronas [Mas08]. Torrey [Lie20]. total [Cod04, Row75, Hel07]. tough [Whi87]. TPC [Ano09s, Ano09u, Ano10s, Ano11f, BESFW22, Cos23a]. Traceability [MKBF09]. Traces [Wis15]. Tracing [Fer04b]. Tracking [ELdJ12, LS23, Men23, For04]. Tracy [Flo10]. Trade [Mar64, Noo79, Cha58]. trademark [Dia77]. Tradition [Ram98]. Traditional [TM08, App01b, Cod04, LNB^{+}04, Rub86b]. Tragic [Bal18]. Tragico [Luf69a]. trainer [Jen99a]. trainers [BD82]. Training [Bis90, Car12, DR20, Flo06c, GB16, Hof05, Lof10, Mel08, Raj12, Und67, WP18, Wen69, Zem12, Zem21, Bae12, BP82, BTM12, BD84, CM02, Fre03a, Fri85, Fry81, Jos81, Orr03, Rai86, RCK^{+}90, San87c, Smi85, VLTP95, VLA96, Whi91, Win89, Str03]. Trajectories [BJ21a]. Trans [MMVH10, VMM21]. Trans-Atlantic [MMVH10, VMM21]. Transactions [Ano71a, Ano73a, Ano74a, Ano75a, Ano76a, Ano78a, Ano79p, Ano79q, Ano79r, Ano79a, Ano80x, Ano80y, Ano80z, Ano80-27, Ano80a, Ano81n, Ano81o, Ano81p, Ano81b, Ano82l, Ano82m, Ano82n, Ano82c, Ano83w, Ano83x, Ano83y, Ano83b, Ano84y, Ano84z, Ano84a, Ano85t, Ano86u, Ano86l, Ano97c, Ano00, Ano04f, Ano04b, Ano04g, Ano04c, Ano04h, Ano04d, Ano04i, Ano04e, Ano05f, Ano05b, Ano05g, Ano05c, Ano05h, Ano05d, Ano05i, Ano05e, Ano06f, Ano06b, Ano06g, Ano06c, Ano06h, Ano06d, Ano06i, Ano06e, Ano07b, Ano07g, Ano07c, Ano07h, Ano07d, Ano07i, Ano07e, Ano07j, Ano07f, Ano08a, Ano08f, Ano08b, Ano08g, Ano08c, Ano08h, Ano08d, Ano08i, Ano08e, Ano09a, Ano09f, Ano09b, Ano09g, Ano09c, Ano09h, Ano09d, Ano09i]. Transactions [Ano09e, Ano10a, Ano10f, Ano10b, Ano10g, Ano10c, Ano10h, Ano10d, Ano10i, Ano10e, Ano11b, Ano11a, Ano12c, Ano12h, Ano12f, Ano12i, Ano12g, Ano12j, Ano12k, Ano13c, Ano13f, Ano13g, Ano13h, Ano13i, Ano14a, Ano14d, Ano14e, Ano14f, Ano14g, Ano15a, Ano15h, Ano15d, Ano15i, Ano15e, Ano15f, Ano15g, Ano16a, Ano17a, Ano18a, Ano20a, Ano22a, Ano22b, Ano22k, Ano23a, Ano23b, Ano23f, Ano23c, Ano23d, Ano23e, Ano24a, Ano24b, Ano24c, BG07, Fry87, Wal21, Ano85a, Ano86b, Ano87a, Hay16]. Transcendence [McC19]. Transcript [Dob75]. Transfer [Dia14, FPKS^{+}21, KO20, AA93, Ano04s, For04, Lin72, MYV^{+}91, MYV92, Per93, Rob95, SSNJ05, Sch79, Whi79i, Mic81]. Transfer-Focused [KO20]. transferable [BMRH05]. Transferring [MQ76, Ham96]. transformations [BFA05]. Transforming [Flo20]. Transition [KSNF15, WLM^{+}22, Boc98, Sch93]. Transitional [Mar75, Wis91]. translatability [SHS97]. Translatable [LJ10]. Translate [Mil81]. translated [FC96]. Translating [Che92, Lav06, SI06, St.02c, JGH82]. Translation [CCL24, DL24, LR24, LMJK14, Lip71, LAWI24, Luf69b, Mel08, PD24, PDD^{+}24, Ama03a, Dat91, Fla05, Lip75, Luf65, Luf75, St.00b]. translations [KSW78]. Transliteracies [Wal18]. transmission [Li80]. Transnational [AW17, ACW19, HS19, Evi04]. Transnationally [Gro99]. Transportation [Moo16, Rou17a]. Traps [Mic59]. Travel [Wes71]. treasure [Sch65]. Treasury [Her95]. Treatise [HN93]. treatment [Vel77]. trees [DH94, Cod09]. Trenches [AB15b, STM09]. Trends [JD11, St.04, KE86]. trespassing [Vel87b]. Tribe [Flo18]. Tricks [Cha58]. troubleshooting [Tho67]. Trust [FMT12, Hou22a, MPD10, PDL16, PFD^{+}21, CHR04, HR99]. trusted [PM98]. Trustworthy [Hou22a]. Truth [Coc04, Por87]. Try [Hol76, Get86]. Tufte [Kat84, Moe98]. Turmoil [BJ21a]. Turn [Ano69x, Ano70c, Ano71e, Ano71m, Sta20, Jos81, Ter65]. Turner [Mir99]. Turning [Ken82a, Mur59, Bro93]. Turoff [Sal80]. Tusting [LD20]. Tutor [WP18, Mac04b, Mac05a]. Tutorial [AM07, Baa07, Car12, Dav10, Dia07, Flo73, Gir06, HK11, LO07, Lof10, Luf69a, Luf71, McC06, Pad07, Wil10a, Wol06, Yu11, Bae12, Luf68, STM09]. Tutorials [AACF81, Ano08j, Ano09k, Ano09j, Ano09l, Ano10j, Ano10k, Ano10l, Ano11c, Gla04]. TV [BB91, Cha78, Cro95, Rea80]. Twain [Cor94]. Twain-technical [Cor94]. Twelve [Pad07, Sch75a, Sch67]. Twenty [Dre16, Ram00, Sim96b, St.01b, MMW17]. Twenty-First [Dre16, Sim96b, Ram00, St.01b, MMW17]. Twisted [Fil05]. Twitter [HLY22, LHAL23, SvI18, SKZ23, WS23]. Two [Ano63-27, BH15, DL24, GTTX08, JSZ07, KG68, LC17, McC65, Mur00k, NNB^{+}22, Sim96b, TVRR20, Ano63r, Ano87n, BGN^{+}95, Bus91, Dou02a, Dul93, Evi04, GvdMdJP99, Kos95, KU04, Son65, SPR84, Str04, UN00, Luf64]. Two-Dimensional [KG68]. two-stage [Son65]. two-step [Evi04]. two-way [Ano87n, Dul93]. two-year [SPR84]. Type [KG68, Cra83a, Flo73, Jus72, Szo82, Mac03a]. Types [SEN10]. typescripts [Zap80]. Typesetting [Boe76, Boy68, Cor69, Kun69, Pie71, Boy69, Jus72]. Typewriter [MM73]. Typewritten [Mic61b]. Typing [Ano68u]. typists [Zap80]. typo [MC75a]. Typographic [BCJR87, HTB80]. Typographical [MC75b]. typographics [Got78]. typography [Cra83a, Tru88, Mur00i, Mac03a].

U [Lon14]. U.S [Hau80]. U.S.

[Ano77-34, BPK94, Doo79, Gre62, How83, How84, Kim79]. UK [Den98a, Wor87, dGKGN11]. UK-Based [dGKGN11]. Ulf [Rei79]. ultimate [Car87, Zim00]. unbearable [Bro91a]. Uncertain [Joh20, Dul93]. Uncertainties [DB24]. uncharacteristically [Tho67]. unclassified [Sha89]. Uncovering [KL12b]. Undergraduate [FC24, VNC^{+}13]. Undergraduates [KO20]. Underlying [Ano15-28]. understand [SS87]. understandability [Gui89a]. Understanding [Ath22, GQ12, GMG^{+}23, Gro99, Gui94, HK07, KL12b, Man98c, PKLS17, Pot02, Red81, STP22, TDG^{+}18, Ano77-31, Hau80, Luf76, Man99b, McG72, CNC20]. understood [Bur84b]. unemployed [Gre72]. Unfortunate [Ano64z]. Unfriendly [Bro87]. unidentified [Dul80]. Unifying [BA16]. Unintended [HF08]. union [FO06]. United [Ram00, Con91, Dia07, Hov14, Mal07, Ros94, WARAL16]. Units [Wer71, Ano65h]. universal [Con98, Pin18]. Universities [How83, How84, Rai86]. University [CCL24, Gat04, Sap08, SLP^{+}79, Yu23, Zem21, GG84, Hor84]. Unix [PDG^{+}88]. Unjust [CC22b]. unlock [Bro84]. Unlocking [DuB08]. unmanned [Ano87y]. Unorthographic [GY68]. Unplanned [Men72]. unrestricted [MCWK77].

Unsupervised [Tri16]. Update [Sed69]. Updates [Phe20]. Upgrading [Bat84]. upward [Del92]. Urban [WBB13, Cox93]. Usability [Cho18, Coo10, Eva02, FL23, Gui89b, HdJK18, JSZ07, LKS13, May19, MEZ12, NNB^{+}22, Red10, RRW^{+}17, Ros89, Sap08, SH93, Sou96, SM10, SA10, vdHdJS06, CM05, Die89, FC96, GR89, Kan94, KU04, Mir91, Ram89, RK94, SRVR97, Sul89, Vel89, WS89, vW00, vdM97b, Pot01b]. Usable [Fer04a, DWP85, GW84, Joe80a, Joe81b, Joe81c]. Usage [DB24, Cho97a, MCWK77, Whi75]. Use [ASHB16, AZDA10, Bea59, BP10, Boe10, CWSL19, Doe69, DSV10, FCS10, Fri19, GQ12, GMG^{+}23, GCMQ17, HSA21, KNMM09, Kil13, Löc22, LG14, LLY18, Mac05b, MG82, Men23, Met75, Mun05, Räi20, Sab61, Sal68, Sam58, Ste08, Ste22, SJYZ18, SKZ23, WMPBA07, WPSB13, Wen69, YJY24, Ada81, Ano77h, Atl81, BD83, BRL91, BJ79, Blo87, Dia77, DSC78, Edd83, EM94, Flo73, FS92, HSV05, HFS94, HE77, Hor91, JF03, KS04, MK84, RLS06, ST99a, ST05a, Sop75, St.00b, Wal74, Wri79, Zim83]. Used [Ano81h, Car24, McG00]. Useful [Fer04a, McC87, Bos84, Hub85]. Usefulness [AZDA10, GAGW15]. useless [McC87]. User [Ale20, BIdJ^{+}21, CS06, Fer04c, Fil06, FL23, GL16, HdJK18, HBM21, HVS^{+}22, JSZ07, Joh20, Lan04, LdJK20, Mac05b, MY14, Pet13, Pin18, Que86, Roh19, RRW^{+}17, Sap08, SDB^{+}13, Sha94a, Sha94b, SSS^{+}13, Str99, Swa18, Tho70, Tri16, WC15, Wil24, WBB13, WJ13, ZLCZ17, ZM23, Zie69, vSdJ07, vdGRRv13, vdM07, AACF81, App02, Atl81, Bae12, BMRH05, Bro86, BB86, CJ89, Dou72, Evi04, FR86, Gri81, HB01, Mar86, May82, MRE96, Ode86, Pak81, PDG^{+}88, PFM88, Ril02a, SH91, Sch86, SRVR97, ST06a, Tel68, Vel90, Pot01b, Tra12]. User- [SDB^{+}13]. User-Centered [BIdJ^{+}21, Fer04c, HVS^{+}22, JSZ07, SSS^{+}13, Pot01b, Tra12]. user-database [ST06a]. user-developed [PFM88]. user-driven [May82]. User-Experience [ZM23]. User-Generated [GL16, WC15, SRVR97]. user-oriented [Tel68]. Users [HdJK18, Izq24, Ber82, Blo83, Com86, Fur90, Rid87, Tal77, Wea85, DR10, Gri83]. uses [Ano77f, Den96]. Using [AW22, Bed12, BCJR87, BS16, CB08, Coh69, Cra92, Cra98, D’A64, Del00, DC07, EZ88, Gat04, GRS96, GHR67, Gol97, Gon20, Hac85, Hil89, HK11, HDKC08, KZCR10, KT18, KT20, KO20, Lam05, LMJK14, Lee13, LN03, MPD10, Men23, Pot84, Räi20, Ric09, Rif10, RB19, Sid87, Som89, Sto09, Str21, Wal99, Wat20, Wil10a, YTY12, ZWE21, vMPKG15, Bau62, BM00, Cor92, Fre13, Gol77, Hay83, HC88, Kry99, Per70, Rei88b, SD06, WBW89, Whi89, Ano71d]. Utility [Wal74, Zim01]. utilization [Lin72]. utopianism [Sau93]. UX [Ale20, Lie20, Pin18, Roh19, Cos23a, GG23, GB16, GLL23, JD23, KW23, ZM23, Ale20, Zet19].

v [You11, Ano73-27]. Va [The97]. Valid [Boe10]. validate [Sid87]. validity [WS89, dJS00]. Valley [Mur00k]. Value [Ano68-27, BAB10, Ram05a, WD09, Fis99, Fra89, Nor00, RR93, Sch82, Wol79]. Value-Sensitive [WD09]. values [HR99, Max73a, Max73b]. vanquishing [QK05]. vantage [Bal75]. Vardaman [Rob78, Rob78]. Variation [YJY24, San87a, vW00]. varied [ST99a]. Variety [KL12a]. Vector [GHR67]. Vehicles [Arn84, GR94]. Veil [Wil10b]. velocity [Der99]. Velsen [Tra12]. Vending [SSS^{+}13]. Vendor [SDB^{+}13]. Vendor-Generated [SDB^{+}13]. Venture [LKHC16]. Verbal [Sch81b, Smi79, ZFWX21, vdHdJS06, KU04]. Verbalizations [ELdJ12]. verge [Mau79]. Verification [LKS13]. verifying [BW90]. Vernon [Her95]. versatile [BS89]. versatility [Ese69]. Versus [LBB18, NNB^{+}22, SA23, Ano68v, Coc04, Fra75, FS00, JM01, LFT08, ML78, PHW13, Ros89, Sha89, SF02, Ste77, Wol88, Zim88]. verve [Pow82b]. Very [Lev68, Bos84]. Vestigiality [Sop75]. via [Bas97, KG68, RDJM09, Sal80, VMM21, Wil24]. Video [GDTZ09, Wen69, WRTBRL23, DBK^{+}93, Mal87, Whi91, Str01]. Video-Conferencing [GDTZ09]. Video-mediated [Str01]. Videoconferencing [Spr97]. Videos [Mog14, STP22]. View [SS10, Zie69, Ano63r, Ben85, Rau81, Web75]. Viewpoint [Ano58d, Bha77, Dal81, Har75b, Hor75]. Views [Sim96b, Kos95]. Vincent [Hen22]. Virginia [JW22]. Virtual [AAA19, BN07, BN10, BSS13, Bre10, BMS^{+}15, DPHF17, Fer07c, FCS10, FMT12, FVB16, GDTZ09, Hin10, Hot06, KL12a, LRLM17, Ock07, PDL16, RD16, RVvGS08, SEN10, TDG^{+}18, Tha24, VT21, WPSB13, Zem12, Ama03b, AC06, BB99, CHR04, ES99, JDW96, Lee00, Lin99, LNCL05, Ock05, PD05, RKP00, SSNJ05, SV06, SH01, TWHN00]. Virtuality [PFD^{+}21, RVV^{+}02, ST99a]. Visibility [Kil13]. Visio [DR10, DR10]. Vision [BA16, Tar01]. Visions [McC19]. visited [Ano04k, Ano04l, Ano07m, Ano11d]. Visual [AM07, Aug78, Cla12, Gou73, Gri91, HP88, Hof07, Hor90, Kos07, Kos08, STP22, TM15, WS92, ZFWX21, Ano77-30, Bos92, GvdMdJP99, Kat84, MG95, Mus78a, Ols91, RP98, WH98, Moe98, St.02b, Wof09]. Visualization [Cuf19, Cuf20, Wel22]. Visualizations [MMW^{+}22, Str21]. Visualizing [BIdJ^{+}21, Kae04, DR10]. Visually [BNB21, MD83]. Visuals [AM07, Hof05]. Vital [Dud61b, Dud61c]. Vitalizing [HS79]. Vizio [DR10]. vocabularies [MCWK77]. Vocabulary [Mil64, Wal85]. Vocabulary-Improvement [Mil64]. voice [Bra03b, LR85, Loe91, RLS06, Rob79a]. Vol [Ano71a, Ano73a, Ano74a, Ano75a, Ano78a, Ano79a, Ano80a, Ano81b, Ano82c, Ano83b, Ano97c, Ano00, Ano07b, Ano08a, Ano09a, Ano10a, Ano12c, Ano13c, Ano14a, Ano15a, Ano16a, Ano17a, Ano18a, Ano20a, Ano22a, Ano23a, Red98, Ano76a, Ano84a, Ano85a, Ano86b, Ano87a]. Volume [BWPH12]. Volumes [Ano61n]. volunteers [Moo87d]. VR [Tha24]. vs [LKHC16, Luf65, Noo79].

W

[Ano63-27, Ano80g, Den83c, Hyd18, Joe79a, Kie83, Mir99, Riv85, Sch78b, Whi79h, You10]. Waiting [Ano68w]. Waking [Cox93]. Walker [DR10, Whi87]. Wall [FDJW18]. Wallace [Whi80b]. Walter [Den83c, Den83b, Har80b, Rei79, Whi77c]. Walton [Gan22, Sta20]. want [Jen99a, Rid87, Shu93]. wanted [Boc98]. War [Ros79]. Warner [Bar09b, Pen98f]. warning [RCB82]. Warnings [GMT08, Cle87, Dri87]. Waste [Til06]. watching [DeT94, Mon80b, Mon81b, Mon81c]. Water [Kon22]. Watson [ES98]. Way [BC70, Lan04, Ano87n, Dio82, Dul93, Eng72, Hug79, Von78]. Wayne [Her95]. ways [Har81]. weaknesses [Rav88]. Wearable [Rui20]. Weather [MMW^{+}22]. Weaving [Zim00]. Web [BFA05, BK03, CD98, Dur99, Fil06, Gen10, Mur00b, St.03, Ste77, Wei02a, App01a, Ril02b, SWB^{+}05, Zim00, ACH^{+}98, App02, CM01, CM02, Evi04, Eva00, KCJC08, MMCL00, Pet14, RLG97, RL96, Ric09, RO96, Sto09, Str21, Wei98, WOD^{+}16, ZVS01, Bar09c, Mur00c, Wof09, Zim00, LN00, LN00, Mur00g, Sto08, Str03]. Web-Based [CD98, BK03, KCJC08, Pet14, CM02, Evi04, MMCL00, Str03]. web-design [SWB^{+}05]. Web.studies [St.01c]. Weblogs [DW07]. Website [XS24, Joh10, Kil13, Lin07, SSSC08]. Websites [Kil09, XS24, YJ01, Cod04, ZWX20, Zol14, ZB15, vW00]. Webster [Sch63]. Weeks [Bat18a]. Weisman [Mic85b, Ano71c, Har80a, Mic65]. Weiss [Rav06]. Welcome [McN58b]. Well [Gud82, HM95, Ama04a, Joh82]. Well-written [Gud82]. Wen [Car16]. Wensveen [Zim13]. We’re [BN07]. Westendorp [Wei02b]. Western [LdJK20]. Whalen [Bec84, Flo08, Gon20]. Where [BN07, Lam14, Phi69b, Rob79b, Ter72, Lip04, TDS95, Vel87a, Woo81]. Which [Mac03b]. while [Ano68w]. whistle [HWW04]. whistle-blowing [HWW04]. White [CNC20, Bro75, Whi79h]. whiz [Bro88a]. Who [Mon81c, Pow81, Sar73, Wal64, CH89, Gre72, Mar85, Cho97a]. ’whole [Bal08, GH09]. whom [Mon81c]. WI [Pat08]. Wicked [Joh20]. Wide [Mur00c, RLG97, RL96, Xu08, Zim00]. Wider [Mar77]. Wiedersehen [Per70]. Wiki [Tri16]. Wikis [Fer07c]. Wilcox [Mic64]. Wild [Joh20]. Wiley [RA97]. Will [Bar09a, Doe69, Nob10]. Willerton [Dre16]. William [GG22b, Bro83a, Nor97]. Williams [GGS23]. Williamson [Wis15]. Wilson [Whi75, vdM97a]. Win [HVS^{+}22, Ram07, Bec83, Fre13, GB84b, Lay73, Flo10]. Winning [Ano17d, Bar91, Bec84, Flo00a, Geo83, Gre85, LD84, Wha01]. Winston [Ano80j, Mon78]. Wired [Sel98]. wisdom [LN00, STM09]. Within [FC24, GB16, Hyd18, PD22, AC06, Dav10, Del92, IPC01, Lay88, Reg70]. without [Ano04r, Coh69, Fuj58]. Wixon [RA97]. Wolfgang [Fer06]. Wolman [Cun09]. Woman [Nee92]. Women [BGN^{+}95, JW22, WARAL20, Nee92, TL92]. won’t [Ful80]. Word [McC63, MCWK77, Mon80b, Mon81b, Mon81c, Pet59, Ano63o, Ano81r, Ano82r, Dre87, JC78, Pro94, Woe80, Lie20]. wordless [HH82]. Words [Ano63-28, Bai24, Col65, Ock71, Ock75, Sch72c, Dec84, ML99, MC75a, Swa80, Woo79]. WordStar [Fri85, Fri85]. Work [AW17, BH21, CC19, Con99, Cun10, DOHM96, FPKS^{+}21, Gan22, Hin73, KSNF15, LLY18, Ric09, Swa18, TJQN11, VG16, Vic03, WOD^{+}16, Yea10, Bel99, FR86, Hag90, KB84, Lin99, Loe95, Mal85a, Mic87, Mir99, Nag87a, Rei88a, SR83b, Zim88, Spi96, Vic03, Flo02a, Flo02b, Whi79d, Whi79i]. work-world [Hag90]. Workbook [Bar22, Hau80]. Worker [BWPH12, DM81]. workforce [NH05]. Working [Fou59, Phe20, Die89, Heu75, Gro99]. Workplace [BR16, Bed12, Dou00, Fer05a, Gra96b, HDKC08, Hun21, JC21, Ley08, Sta08, Sta17, Ste08, Win97, Wol06, Ama03b, Cho97a, Den96, FP95, GD85, Lee00, Nel90, Par06, RH92, ST99a, Str89, VLTP95, Mag03, Bai24, Emi99, Hou21d, Bra99a]. Workplaces [Bre10, Fer08, Hin10, HS19]. Works [Ano77-34, Mac04a, Smi10, Har80b]. Works-Cited [Smi10]. worksheet [Nel90]. workshop [Bec85, Buc74, Jen66]. Workshops [Ano59r]. World [Baa07, Cha06, Coh69, GPHDM16, Mur00c, Zim00, Ano87-28, Bas72, Fle85, Flo02a, Hag90, Rem78, Sno64, Whi79d, RLG97, RL96, Mal11]. World-Wide-Web [RLG97]. Worlds [BSS13, Spi96, Dou00]. worldwide [Dat91]. worth [Dec84, Mac03b]. Write [Bah85, Bar22, Bec83, GD85, His78, LMJK14, McM61, Ros64, Ulr84, Van59, Whi80c, And83, Arm83, BA83, Bou58, Bro83b, CH89, Day77, Est82, Fei79, Flo02a, Ful80, Geo83, Get86, Gri83, Gwi84, Hin73, Kat98b, Kat98a, Kna84, Kre86, Mic80, MF83, Nel90, Oli58, Sav88, Sch67, Sch75a, Sch78c, Sch80b, Smi03, Sou72, Sou75, Van77, Vel87a, Wel82, Whi79d, Whi79g, Wil84, Flo07d, Whi78c]. Writer [BP15, Car00, Dia07, Hag61, Mic61a, Mil61, Pet69, Rat61, Ama04a, Ano81s, BS87, Bro65, CG83, Cor94, DH94, Den83c, Eck03, FW78, FS85, Flo00c, FR86, Fry87, Ivo80, Kni86, Kow78, Mar75, Mey87, Mil65, Ort72, Tay02, Cha65]. Writers [Ale20, Flo07f, KI06, WC15, Bis90, Bur85, FP95, Flo01b, Kus97, LL03, Mas84, McC92, McG00, Nee92, Ril99, SS87, Sur96, Wor87, Wri81, Flo18, Eck03, Mur00a, Str84]. Writers-in-a-Hurry [KI06]. writes [Car80]. Writing [Ano58i, Ano59m, Ano59n, Ano59o, Ano60a, Ano61a, Ano61u, Ano61v, Ano61w, Ano62j, Ano63c, Ano63i, Ano63m, Ano63n, Ano63v, Ano63w, Ano64k, Ano64s, Ano64t, Ano64u, Ano65g, Ano66g, Ano66h, Ano66i, Ano66j, Ano67j, Ano67k, Ano68h, Ano68i, Ano68j, Ano68k, Ano68l, Ano68m, Ano68v, Ano69j, Ano69k, Ano69l, Ano69m, Ano69n, Ano70i, Ano70j, Ano70k, Ano70l, Ano71n, Ano71o, Ano71p, Ano71q, Ano71r, Ano71s, Ano71a, Ano77c, Ano97d, Ant99, Ant06, Arn61, Baa07, BH15, Ben61, BC70, Bli61, BW19, Bra59, Bra05, BB86, Bro62, BJT^{+}19, Car59, Car61, Car24, Cha96, Cla59, Clo64, Coh69, Con58a, Con58b, Cos84, Cot79, Cra85a, Cra85b, Dav67, Dia14, Doc61, Dou00, Dud71, Dul80, DGPR10, Ega81, Emi99, Far64, Fei59]. Writing [Fer12, Flo02c, Flo03d, Flo05a, Flo07c, Fuj58, Gil87b, Gir06, GP86, Gol82b, Gol58, GB79, HS19, HSR84, HK22, Hol03, Hou21a, Jac81, Joh90, Jos81, Kae05, Kei67, Ken64a, KT20, KMR14, KO20, Kin58, Kry00b, KRM24, Lam05, LJ10, LMJK14, Lee13, LS84, Ley08, LD20, Lie20, Lin04, Lip58, Mac04c, Mal85b, Mal01, MCC62, Mar08, McC65, Mic61e, Mor66, Mor67b, Nor97, Pet14, Phi69a, Pow82b, Rac59, RM61, Ram98, Rav06, Reu65, Ric09, Roc59, Rom81, Rou19, SM66, SV24, Sau82, Sch78b, Spi96, Spi86, Van71, WP18, Wha85, Whi06, Wil10b, WFSC^{+}23, Wol06, YTY12, ZB17, ZWE21, Abs82, All96, Ama03a, And81, Ano81j, Ano82e, AL03, BCB03, Bar82a, Bea99, BM00, BJ79, Bis99, BD82, Bos84, Bos87b, Bra82a, BSR85, Bro86, Bro80]. writing [Bro83b, Bro84, Bus91, Cam95, Car87, CI02, CW90, CEH87, CJ89, Cla75, Col77, Cor90, Cor83, Cor92, Day83, DGORM84, Deg80, DS89, Den88, Den78, Dou01, Dun81, EP83, Eng72, Eng89, FG84, Fis92, Fle85, Flo00a, Flo00b, Flo02b, Fre80b, Goo05, Gre85, Ham96, Har00, Hat84, Hel88, Hin72, His77, Hol76, Hor84, Hub85, Hul80, HR60, Jen66, JM01, JF03, Kat98a, Kau88, KD00, Ken82b, Ker77, KD84, Kie82, Kil83, KB84, KW93, Kre89, Kre00, LeC01, LL03, Lip83, Lop84, LN03, LNB^{+}04, LNCL05, Mac99b, Mac02a, Mac03b, Mac04b, Mac05a, Mac06, MB84, Mas84, MMCL00, Mey64, Mic62, Mic74, Mic87, Mil81, Moo85, Mor67a, Mue80, NKU03, Nel03, Ort78, PS96, Pie58, PP97, Plu84, Pot02, Pow80, Rat85, Rei81, RH94, Rob78]. writing [RC89, Sch72c, Sch75c, Sch93, Shi81, Shi82, Sid87, Smi84, Smi03, Ste99, Str89, Teb80, Tib82, Tra92, Van83, Van63, Wal99, WP82, Wei95, Wha83, Wu88, You80, Zim93, Ano63-27, Bro83a, Ede87, Erv99a, Joh97b, Bra99a, Mic85b, Ano63-27, Flo06a, GG22b, Hue02, Jen99b, Reh98, Win97, Ano71c, Cha80a, Cha80b, Den83b, Har80a, Rei02, Sch75b, Sch78d, Whi77c, Whi79h, Ber12, EC01, Har80b, Hau80, Hou22b, Lay97, Sch80b, Wha02b, Whi78e, Zim01]. Writing-Across-the-Curriculum [ZB17]. writing-for-translation [Ama03a]. Writing-Improvement [Ano63i, Ano63v, Ano63w, Mey64]. writing-principles [Cla75]. Writings [Gra96b, Rin63, Abs82]. Written [CSSV20, Bat84, Ben79, Dec84, DM81, Flo81, Gil03, Gud82, HS72, Pot84, Sch78e]. Wrong [KHV10, Gwi84]. wrongs [Car80]. WSN [AZ17]. WSN-Based [AZ17]. www.ieee.org [Ano04k, Ano04l, Ano07m, Ano11d].

Xiaojuan [Rok19]. XML [Mur00j]. Xplore [Ano12x, Ano12y, Ano12z, Ano12-27, Ano15x, Ano15y, Ano09u, Ano14h].

Year [Mal07, Gol77, HN93, SPR84]. Years [RVvGS08, FHR87a, Her89]. York [Mek72, RA97]. You’ll [Bur84b]. You’re [Flo04b, Nob10]. yourself [Bur84b, KR83, Sch83]. YouTube [May19, STP22]. Yu [Cha23, Rui20]. Yvonne [GG22c, Fer03].

Z [JC21]. Zachry [Fer08]. Zealand [Gra13, KPV08]. Zemliansky [Hin10]. Zen [BR93, Jon89]. Zhihu [YJY24]. Zimmerman [Bar98a]. Zittrain [St.09].

References

Amsden:1993:KSP

[AA93] D. C. Amsden and A. A. Amsden. The KIVA story: a paradigm of technology transfer. IEEE Transactions on Professional Communication, 36 (4):190–195, December 1993. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Alaiad:2019:VTT

[AAA19] Ahmad Alaiad, Yazan Alnsour, and Mohammad Alsharo. Virtual teams: Thematic taxonomy, constructs model, and future research directions. IEEE Transactions on Professional Communication, 62(3):211– 238, September 2019. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Al-Awar:1981:TFT

[AACF81] Janan Al-Awar, Alphonse Chapanis, and W. Randolph Ford. Tutorials for the firsttime computer user. IEEE Transactions on Professional Communication, PC-24(1):30– 37, March 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Applbaum:1984:EOC

[AAR84] Ronald L. Applbaum, Karl W. E. Anatol, and Diana C. Reep. Effective oral communication for business and the professions. IEEE Transactions on Professional Communication, PC-27(1):46–47, March 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Addeo:1981:ESW

[AB81] Edmond G. Addeo and Robert E. Burger. Ego-speak, why no one listens to you. IEEE Transactions on Professional Communication, PC-24(2):104–105, June 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Andersen:2015:CSC

[AB15a] Rebekka Andersen and Tatiana Batova. The current state of component content management: an integrative literature review. IEEE Transactions on Professional Communication, 58(3):247–270, September 2015. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Andersen:2015:ISI

[AB15b] Rebekka Andersen and Tatiana Batova. Introduction to the special issue: Content management perspectives from the trenches. IEEE Transactions on Professional Communication, 58(3):242–246, September 2015. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Abelson:1973:CES

[Abe73] Philip H. Abelson. The coming evolution of scientific journals. IEEE Transactions on Professional Communication, PC-16(3):69–70, September 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anderson:1984:NET

[ABMF84] Paul V. Anderson, R. John Brockmann, Carolyn R. Miller, and Charles R. Fenno. New essays in technical and scientific communication: Research, theory, practice. IEEE Transactions on Professional Communication, PC-27(1):45– 46, March 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Abrahams:1977:FEQ

[Abr77]  S. G. Abrahams.  Frame-

work for estimating the quality of scientific journals. IEEE Transactions on Professional Communication, PC-20(2): 133–136, September 1977. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Abshire:1982:BWW

[Abs82] Gary M. Abshire. A bibliography of writings on writing. IEEE Transactions on Professional Communication, PC-25 (4):211–219, December 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Abshire:1985:TAP

[AC85] Gary M. Abshire and Dan Culberson. A team approach to producing good documentation. IEEE Transactions on Professional Communication, PC-28(4):38–41, December 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anawati:2006:BAW

[AC06] D. Anawati and A. Craig. Behavioral adaptation within cross-cultural virtual teams. IEEE Transactions on Professional Communication, 49(1): 44–56, March 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anderson:1998:EWS

[ACH^{+}98] S. L. Anderson, C. P. Campbell, N. Hindle, J. Price, and R. Scasny. Editing a Web site: extending the levels of edit. IEEE Transactions on Professional Communication, 41(1): 47–57, March 1998. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Acharya:2024:DEI

[Ach24] Keshab Raj Acharya. Designing equitable and inclusive mHealth technology: Insights from Global South healthcare practitioners. IEEE Transactions on Professional Communication, 67(2):229–245, June 2024. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ackenhusen:1989:DMS

[Ack89] J. G. Ackenhusen. Database management system for conference program preparation. IEEE Transactions on Professional Communication, 32(3): 194–199, September 1989. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ai:2019:LIT

[ACW19] Bin Ai, Can Cui, and Lifei Wang. Language, identity, and transnational communication: Chinese business expatriates in Africa. IEEE Transactions on Professional Communication, 62(2):178–191, June 2019. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Adkins:1988:SBD

[AD88] B. Adkins and J. Daley. Strategies for business development in the federal marketplace. IEEE Transactions on Professional Communication, 31(2): 57–64, June 1988. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Agboka:2022:RTC

[AD22] Godwin Y. Agboka and Isidore K. Dorpenyo. The role of technical communicators in confronting injustice everywhere. IEEE Transactions on Professional Communication, 65(1):5–10, March 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Adamy:1980:TTT

[Ada80] David L. Adamy. Technical talks: They’re not as hard as you think. IEEE Transactions on Professional Communication, PC-23(1):22–25, March 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Adamy:1981:SNH

[Ada81] David L. Adamy. Speech notes: How and when to use them. IEEE Transactions on Professional Communication, PC-24 (3):130–132, September 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Adams:2006:ACS

[Ada06] Mary Alice Adams. Authorship in composition studies. IEEE Transactions on Professional Communication, 49(4): 375–376, December 2006. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Adler:2000:ESP

[Adl00] T. R. Adler. An evaluation of the social perspective in the development of technical requirements. IEEE Transactions on Professional Communication, 43(4):386–396, December 2000. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Apple:2020:RBF

[AFH20] Matthew T. Apple, Joseph Falout, and Glen Hill. The relationship between future career self images and English achievement test scores of Japanese STEM students. IEEE Transactions on Professional Communication, 63(4): 372–385, December 2020. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Alzoubi:2021:CCC

[AG21] Yehia Alzoubi and Asif Gill.

The critical communication challenges between geographically distributed agile development teams: Empirical findings. IEEE Transactions on Professional Communication, 64(4):322–337, December 2021. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Aichroth:1978:FPD

[Aic78] Joseph W. Aichroth. Flatpanel displays offer graphics alternatives. IEEE Transactions on Professional Communication, PC-21(1):14–17, March 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anthony:2003:MML

[AL03] L. Anthony and G. V. Lashkia. Mover: a machine learning tool to assist in the reading and writing of technical papers. IEEE Transactions on Professional Communication, 46(3): 185–193, September 2003. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Alexander:2020:UFN

[Ale20] Delainey A. Alexander. UX Fundamentals for Non-UX Professionals: User Experience Principles for Managers, Writers, Designers, and Developers: Edward Stull [Book Review]. IEEE Transactions on Professional Communication, 63(2):203–204, June 2020. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Allen:1996:DEW

[All96] N. Allen. Designing an electronic writing classroom. IEEE Transactions on Professional Communication, 39(4):232– 238, December 1996. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Allen:1997:RHM

[All97]  N. Allen.  Randy Y. Hi-

rokawa and Marshall Scott Poole, Eds., Communication and Group Decision Making, Thousand Oaks, Ca: Sage Pub., 1996, 488 pp. [Book Review]. IEEE Transactions on Professional Communication, 40(4):308–309, December 1997. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Amare:2007:LVT

[AM07] N. Amare and Alan Manning.

The language of visuals: Text + graphics = visual rhetoric tutorial. IEEE Transactions on Professional Communication, 50(1):57–70, March 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Amare:2000:PCP

[Ama00a] N. E. Amare. Professional communication and Poe’s “The Purloined Letter”. IEEE Transactions on Professional Communication, 43(2):206– 209, June 2000. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Amare:2000:PJA

[Ama00b] N. E. Amare. Profile of Jan D’Arcy. IEEE Transactions on Professional Communication, 43(3):331–335, September 2000. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Amare:2000:PPD

[Ama00c]  N. E. Amare.  Profile of

Patrick Dean. IEEE Transactions on Professional Communication, 43(2):213–214, June 2000. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Amant:2001:CCC

[Ama01]  K. St Amant.  Cultures,

computers, and communication: evaluating models of international online production. IEEE Transactions on Professional Communication, 44(4): 291–295, December 2001. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Amare:2002:CTC

[Ama02] N. Amare. The culture(s) of the technical communicator. IEEE Transactions on Professional Communication, 45(2): 128–132, June 2002. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Amant:2003:DEW

[Ama03a] K. S. Amant. Designing effective writing-for-translation intranet sites. IEEE Transactions on Professional Communication, 46(1):55–62, March 2003. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Amant:2003:MCI

[Ama03b] K. S. Amant. Making contact in international virtual offices: an application of symbolic interactionism to online workplace discourse. IEEE Transactions on Professional Communication, 46(3):236– 240, September 2003. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Amare:2004:AWT

[Ama04a] N. Amare. Act well thy part: performing technical writer and engineer. IEEE Transactions on Professional Communication, 47(3):211– 215, September 2004. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Amare:2004:SLT

[Ama04b] N. Amare. Service-learning in technical and professional communication. IEEE Transactions on Professional Communication, 47(2):148–149, June 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Amant:2005:PTA

[Ama05] K. St Amant. A prototype theory approach to international image design. IEEE Transactions on Professional Communication, 48(2):219–222, June 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Andrews:2021:HMM

[AME21] Christopher D. M. Andrews, Mehrube Mehrubeoglu, and Charles Etheridge. Hybrid model for multidisciplinary collaborations for technical communication education in engineering. IEEE Transactions on Professional Communication, 64(1):52–65, March 2021. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Arthur:1985:EGB

[AN85] Richard Arthur and Rose Norman. The engineer’s guide to better communication. IEEE Transactions on Professional Communication, PC-28(1):54– 55, March 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anabire:2021:DTT

[Ana21] Dorcas Asenyeni Anabire. Design Thinking in Technical Communication: Solving Problems Through Making and Collaboration: Jason C. K. Tham: [Book Review]. IEEE Transactions on Professional Communication, 64(3):302– 303, September 2021. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Andrews:1981:TWE

[And81] Alex M. Andrews. Technical writing: The engineer’s masterpiece. IEEE Transactions on Professional Communication, PC-24(3):141–145, September 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Andersen:1983:GEW

[And83] Dan Andersen. Getting engineers to write. IEEE Transactions on Professional Communication, PC-26(4):170–171, December 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anderson:1984:WTS

[And84] Paul V. Anderson. What technical and scientific communicators do: a comprehensive model for developing academic programs. IEEE Transactions on Professional Communication, PC-27(3):161–167, September 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Andrews:1997:MJK

[And97] D. C. Andrews. M. Jimmie Killingsworth, Information In Action. Boston, Ma: Allyn & Bacon, 1996, 556 pp., Index. [Book Review]. IEEE Transactions on Professional Communication, 40(3): 235–236, September 1997. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Andersen:2014:PSF

[And14] Rebekka Andersen. Planning for the shaping force of cultural dynamics in a component content-management system implementation. IEEE Transactions on Professional Communication, 57(3):216– 234, September 2014. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Angell:1959:RYO

[Ang59] James B. Angell. Running your own projector. IRE Transactions on Engineering Writing and Speech, EWS-2(1):33, January 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1958:A

[Ano58a] Anonymous. Abstracts. IRE Transactions on Engineering Writing and Speech, EWS-1(1): 37, March 1958. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1958:AR

[Ano58b] Anonymous. Annual report. IRE Transactions on Engineering Writing and Speech, EWS-1(2):18–19, August 1958. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1958:B

[Ano58c] Anonymous. The bootstrap. IRE Transactions on Engineering Writing and Speech, EWS-1(2):1, August 1958. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1958:CV

[Ano58d]  Anonymous.  The chapter

viewpoint. IRE Transactions on Engineering Writing and Speech, EWS-1(2):4–5, August 1958. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1958:CB

[Ano58e]  Anonymous.  Contributors’

biographies. IRE Transactions on Engineering Writing and Speech, EWS-1(2):32a– 32a, August 1958. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1958:FC

[Ano58f] Anonymous. [Front cover]. IRE Transactions on Engineering Writing and Speech, EWS-1(1): c1, March 1958. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1958:IL

[Ano58g] Anonymous. Institutional listings. IRE Transactions on Engineering Writing and Speech, EWS-1(2):32b–32b, August 1958. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1958:MOO

[Ano58h] Anonymous. Meet our 1958– 1959 officers. IRE Transactions on Engineering Writing and Speech, EWS-1(2):2–3, August 1958. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1958:SAS

[Ano58i] Anonymous. Second annual symposium IRE — Professional Group on Engineering Writing and Speech. IRE Transactions on Engineering Writing and Speech, EWS-1(2): 30, August 1958. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1958:SMC

[Ano58j] Anonymous. Synopsis of members comments. IRE Transactions on Engineering Writing and Speech, EWS-1(2):29, August 1958. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1959:A

[Ano59a] Anonymous. [Advertisement]. IRE Transactions on Engineering Writing and Speech, EWS-2(2):64, June 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1959:C

[Ano59b] Anonymous. Contents. IRE Transactions on Engineering Writing and Speech, EWS-2(1): c1, January 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1959:CBa

[Ano59c] Anonymous. Contributors’ biographies. IRE Transactions on Engineering Writing and Speech, EWS-2(1):40, January 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1959:CBb

[Ano59d] Anonymous. Contributors’ biographies. IRE Transactions on Engineering Writing and Speech, EWS-2(2):62–63, June 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1959:CBc

[Ano59e] Anonymous. Contributors’ biographies. IRE Transactions on Engineering Writing and Speech, EWS-2(3):94–96, December 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1959:E

[Ano59f] Anonymous. Editorial. IRE Transactions on Engineering Writing and Speech, EWS-2(3): 68, December 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1959:FCa

[Ano59g] Anonymous. [Front cover].

IRE Transactions on Engineering Writing and Speech, EWS-2(2):c1, June 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1959:FCb

[Ano59h] Anonymous. [Front cover].

IRE Transactions on Engineering Writing and Speech, EWS-2(3):c1, December 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1959:HDN

[Ano59i] Anonymous. Highlights of the 1959 dual national symposia. IRE Transactions on Engineering Writing and Speech, EWS-2(3):66–67, December 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1959:ILa

[Ano59j] Anonymous. Institutional listings. IRE Transactions on Engineering Writing and Speech, EWS-2(1):40–a–40–a, January 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1959:ILb

[Ano59k]  Anonymous.  Institutional

listings. IRE Transactions on Engineering Writing and Speech, EWS-2(2):64–a–64–a, June 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1959:ILc

[Ano59l] Anonymous. Institutional listings. IRE Transactions on Engineering Writing and Speech, EWS-2(3):96–a–96–a, December 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1959:IPGa

[Ano59m] Anonymous. IRE Professional Group on Engineering Writing and Speech. IRE Transactions on Engineering Writing and Speech, EWS-2(1):c2, January 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1959:IPGb

[Ano59n] Anonymous. IRE Professional Group on Engineering Writing and Speech. IRE Transactions on Engineering Writing and Speech, EWS-2(2):c2, June 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1959:IPGc

[Ano59o] Anonymous. IRE Professional Group on Engineering Writing and Speech. IRE Transactions on Engineering Writing and Speech, EWS-2(3):c2, December 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1959:TCa

[Ano59p] Anonymous. Table of contents.

IRE Transactions on Engineering Writing and Speech, EWS-2(2):41, June 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1959:TCb

[Ano59q] Anonymous. Table of contents.

IRE Transactions on Engineering Writing and Speech, EWS-2(3):65, December 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1959:W

[Ano59r] Anonymous. Workshops. IRE Transactions on Engineering Writing and Speech, EWS-2(1): 1, January 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1960:ISE

[Ano60a] Anonymous. The 1960 International Symposium on Engineering Writing and Speech. IRE Transactions on Engineering Writing and Speech, EWS-3(2):66, July 1960. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1960:ABa

[Ano60b]  Anonymous.  Authors’ bi-

ographies. IRE Transactions on Engineering Writing and Speech, EWS-3(1):31–32, April 1960. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1960:ABb

[Ano60c]  Anonymous.  Authors’ bi-

ographies. IRE Transactions on Engineering Writing and Speech, EWS-3(2):67–68, July 1960. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1960:Ca

[Ano60d] Anonymous. Contents. IRE Transactions on Engineering Writing and Speech, EWS-3(1): 1, April 1960. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1960:Cb

[Ano60e] Anonymous. Contents. IRE Transactions on Engineering Writing and Speech, EWS-3(2): 1, July 1960. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1960:Cc

[Ano60f] Anonymous. The cover. IRE Transactions on Engineering Writing and Speech, EWS-3(2): 34, July 1960. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1960:FCa

[Ano60g] Anonymous. [Front cover].

IRE Transactions on Engineering Writing and Speech, EWS-3(1):c1, April 1960. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1960:FCb

[Ano60h] Anonymous. [Front cover].

IRE Transactions on Engineering Writing and Speech, EWS-3(2):c1, July 1960. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1960:FIC

[Ano60i]  Anonymous.  [Front inside

cover]. IRE Transactions on Engineering Writing and Speech, EWS-3(1):c2, April 1960. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1960:IBC

[Ano60j] Anonymous. Inside back cover.

IRE Transactions on Engineering Writing and Speech, EWS-3(2):c3, July 1960. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1960:PO

[Ano60k] Anonymous. PGEWS 1960– 61 officers. IRE Transactions on Engineering Writing and Speech, EWS-3(2):35, July 1960. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1961:ARA

[Ano61a] Anonymous. Abstracts of recent articles on engineering writing and speech. IRE Transactions on Engineering Writing and Speech, EWS-4(3):101, December 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1961:AI

[Ano61b] Anonymous. Author index. IRE Transactions on Engineering Writing and Speech, EWS-4(1):29, January 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1961:Aa

[Ano61c] Anonymous. The authors. IRE Transactions on Engineering Writing and Speech, EWS-4(1): 30–32, January 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1961:Ab

[Ano61d] Anonymous. The authors. IRE Transactions on Engineering Writing and Speech, EWS-4(2): 63–64, May 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1961:Ac

[Ano61e] Anonymous. The authors. IRE Transactions on Engineering Writing and Speech, EWS-4(3):103–104, December 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1961:Ca

[Ano61f] Anonymous. Contents. IRE Transactions on Engineering Writing and Speech, EWS-4(1): 1, January 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1961:Cb

[Ano61g] Anonymous. Contents. IRE Transactions on Engineering Writing and Speech, EWS-4(2): 33, May 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1961:Cc

[Ano61h] Anonymous. Contents. IRE Transactions on Engineering Writing and Speech, EWS-4(3): 65, December 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1961:Ce

[Ano61i] Anonymous. Correction. IRE Transactions on Engineering Writing and Speech, EWS-4(3): 100, December 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1961:Cd

[Ano61j] Anonymous. The cover. IRE Transactions on Engineering Writing and Speech, EWS-4(3): 66–67, December 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1961:FCa

[Ano61k] Anonymous. [Front cover]. IRE Transactions on Engineering Writing and Speech, EWS-4(1):c1, January 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1961:FCb

[Ano61l] Anonymous. [Front cover]. IRE Transactions on Engineering Writing and Speech, EWS-4(2):c1, May 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1961:FCc

[Ano61m] Anonymous. [Front cover]. IRE Transactions on Engineering Writing and Speech, EWS-4(3):c1, December 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1961:IVE

[Ano61n] Anonymous. Index, volumes EWS-1 to EWS-3. IRE Transactions on Engineering Writing and Speech, EWS-4(1):28– 29, January 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1961:IAa

[Ano61o]  Anonymous.  Information

for authors. IRE Transactions on Engineering Writing and Speech, EWS-4(1):32a– 32a, January 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1961:IAb

[Ano61p]  Anonymous.  Information

for authors. IRE Transactions on Engineering Writing and Speech, EWS-4(2):64a– 64a, May 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1961:IAc

[Ano61q] Anonymous. Information for authors. IRE Transactions on Engineering Writing and Speech, EWS-4(3):104a–104a, December 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1961:ILa

[Ano61r] Anonymous. Institutional listings. IRE Transactions on Engineering Writing and Speech, EWS-4(1):32b–32b, January 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1961:ILb

[Ano61s] Anonymous. Institutional listings. IRE Transactions on Engineering Writing and Speech, EWS-4(2):64b–64b, May 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1961:ILc

[Ano61t] Anonymous. Institutional listings. IRE Transactions on Engineering Writing and Speech, EWS-4(3):104b–104b, December 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1961:IPGa

[Ano61u] Anonymous. IRE Professional Group on Engineering Writing and Speech. IRE Transactions on Engineering Writing and Speech, EWS-4(1):c2, January 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1961:IPGb

[Ano61v] Anonymous. IRE Professional Group on Engineering Writing and Speech. IRE Transactions on Engineering Writing and Speech, EWS-4(2):c2, May 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1961:IPGc

[Ano61w] Anonymous. IRE Professional Group on Engineering Writing and Speech. IRE Transactions on Engineering Writing and Speech, EWS-4(3):c2, December 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1961:OP

[Ano61x] Anonymous. The objectives of PGEWS. IRE Transactions on Engineering Writing and Speech, EWS-4(3):68, December 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1962:A

[Ano62a] Anonymous. The authors. IRE Transactions on Engineering Writing and Speech, EWS-5(1):34–35, August 1962. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1962:BR

[Ano62b] Anonymous. Book reviews.

IRE Transactions on Engineering Writing and Speech, EWS-5(2):100, December 1962. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1962:Ca

[Ano62c] Anonymous. Contents. IRE Transactions on Engineering Writing and Speech, EWS-5(1): c1, August 1962. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1962:Cb

[Ano62d] Anonymous. Correspondence.

IRE Transactions on Engineering Writing and Speech, EWS-5(2):100, December 1962. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1962:EPF

[Ano62e] Anonymous. Editorial [PGEWS and future publishing]. IRE Transactions on Engineering Writing and Speech, EWS-5(1): 1, August 1962. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1962:FC

[Ano62f] Anonymous. [Front cover].

IRE Transactions on Engineering Writing and Speech, EWS-5(2):c1, December 1962. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1962:FIC

[Ano62g]  Anonymous.  [Front inside

cover]. IRE Transactions on Engineering Writing and Speech, EWS-5(2):c2, December 1962. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1962:IA

[Ano62h]  Anonymous.  Information

for authors. IRE Transactions on Engineering Writing and Speech, EWS-5(1):35a– 35a, August 1962. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1962:IBC

[Ano62i] Anonymous. Inside back cover. IRE Transactions on Engineering Writing and Speech, EWS-5(2):c4, December 1962. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1962:IPG

[Ano62j] Anonymous. IRE Professional Group on Engineering Writing and Speech. IRE Transactions on Engineering Writing and Speech, EWS-5(1):c2, August 1962. ISSN 0893-7869 (print), 2331-3692 (electronic).

Anonymous:1963:AILa

[Ano63a] Anonymous. About this issue . . .. IEEE Transactions on Engineering Writing and Speech, EWS-6(1):1, September 1963. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1963:AILb

[Ano63b] Anonymous. About this issue . . .. IEEE Transactions on Engineering Writing and Speech, EWS-6(2):65, December 1963. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1963:BAT

[Ano63c] Anonymous. A behavioural analysis of technical writing. IEEE Transactions on Engineering Writing and Speech, EWS-6(1):60, September 1963. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1963:BRP

[Ano63d] Anonymous. Book review — Problem-Solving Discussions and Conferences: Leadership Methods and Skills, by Norman R. F. Maier. IEEE Transactions on Engineering Writing and Speech, EWS-6(1):62, September 1963. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1963:CR

[Ano63e] Anonymous. Coherent references. IEEE Transactions on Engineering Writing and Speech, EWS-6(1):60–61, September 1963. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1963:Ca

[Ano63f]  Anonymous.  Contributors.

IEEE Transactions on Engineering Writing and Speech, EWS-6(1):63–64, September 1963. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1963:Cb

[Ano63g]  Anonymous.  Contributors.

IEEE Transactions on Engineering Writing and Speech, EWS-6(2):88, December 1963. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1963:DRP

[Ano63h] Anonymous. A debate on the role of the “Proceedings” in the IEEE. IEEE Transactions on Engineering Writing and Speech, EWS-6(1):2–6, September 1963. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1963:DHR

[Ano63i] Anonymous. Dr. Hillier of RCA Laboratories to speak at writing-improvement seminar. IEEE Transactions on Engineering Writing and Speech, EWS-6(2):87, December 1963. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1963:EPN

[Ano63j] Anonymous. Editorial policies of the new major IEEE journal — the “IEEE Spectrum”. IEEE Transactions on Engineering Writing and Speech, EWS-6(1):59, September 1963.

CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1963:FEP

[Ano63k] Anonymous. Five EWS papers to be presented at aerospace conference. IEEE Transactions on Engineering Writing and Speech, EWS-6(2):85, December 1963. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1963:HPP

[Ano63l] Anonymous. How to publicize professional technical group chapter meetings. IEEE Transactions on Engineering Writing and Speech, EWS-6(1):61– 62, September 1963. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1963:IPTa

[Ano63m] Anonymous. IEEE Professional Technical Group on Engineering Writing and Speech. IEEE Transactions on Engineering Writing and Speech, EWS-6(1):c2, September 1963. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1963:IPTb

[Ano63n] Anonymous. IEEE Professional Technical Group on Engineering Writing and Speech. IEEE Transactions on Engineering Writing and Speech, EWS-6(2):c2, December 1963.

CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1963:ICN

[Ano63o] Anonymous. “Impartion” — to coin a new word. IEEE Transactions on Engineering Writing and Speech, EWS-6(1):62, September 1963. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1963:IAa

[Ano63p] Anonymous. Information for authors. IEEE Transactions on Engineering Writing and Speech, EWS-6(1):64a– 64a, September 1963. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1963:IAb

[Ano63q] Anonymous. Information for authors. IEEE Transactions on Engineering Writing and Speech, EWS-6(2):88a– 88a, December 1963. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1963:MTE

[Ano63r] Anonymous. Making two ends meet — one member’s view of PTGEWS activities. IEEE Transactions on Engineering Writing and Speech, EWS-6(2): 85–86, December 1963. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1963:NPP

[Ano63s] Anonymous. A new policy for the Proceedings of the IEEE. IEEE Transactions on Engineering Writing and Speech, EWS-6(1):59, September 1963. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1963:RPC

[Ano63t] Anonymous. A report on the PTGEWS Colloquium on “The Manager’s Role in Organizational Communications”. IEEE Transactions on Engineering Writing and Speech, EWS-6(1):60, September 1963. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1963:RIC

[Ano63u] Anonymous. A request for ideas and comments on IEEE publications. IEEE Transactions on Engineering Writing and Speech, EWS-6(1):59, September 1963. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1963:SWIa

[Ano63v] Anonymous. A seminar on writing-improvement programs for engineers. IEEE Transactions on Engineering Writing and Speech, EWS-6(1):64b– 64b, September 1963. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1963:SWIb

[Ano63w] Anonymous. A seminar on writing-improvement programs for engineers. IEEE Transactions on Engineering Writing and Speech, EWS-6(2):88b– 88b, December 1963. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1963:TCa

[Ano63x] Anonymous. [Table of contents]. IEEE Transactions on Engineering Writing and Speech, EWS-6(1):c1, September 1963. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1963:TCb

[Ano63y] Anonymous. [Table of contents]. IEEE Transactions on Engineering Writing and Speech, EWS-6(2):c1, December 1963. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1963:TPI

[Ano63z] Anonymous. Technical papers interest survey. IEEE Transactions on Engineering Writing and Speech, EWS-6(2):86, December 1963. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1963:TBR

[Ano63-27] Anonymous. Two book reviews: Technical Writing by R. W. Smith; and The Art of Technical and Scientific Writing by E. Ehrlich and D. Murphy. IEEE Transactions on Engineering Writing and Speech, EWS-6(2):86–87, December 1963. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1963:W

[Ano63-28] Anonymous. Words. IEEE Transactions on Engineering Writing and Speech, EWS-6 (2):65, December 1963. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1964:AI

[Ano64a] Anonymous. About this issue. IEEE Transactions on Engineering Writing and Speech, EWS-7(1):1, March 1964. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1964:Aa

[Ano64b] Anonymous. [Advertisement].

IEEE Transactions on Engineering Writing and Speech, EWS-7(1):44–b–44–b, March 1964. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1964:Ab

[Ano64c] Anonymous. [Advertisement].

IEEE Transactions on Engineering Writing and Speech, EWS-7(2):29b–29b, September 1964. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1964:Ac

[Ano64d] Anonymous. [Advertisement]. IEEE Transactions on Engineering Writing and Speech, EWS-7(3):46a–46c, December 1964. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1964:BRE

[Ano64e] Anonymous. Book review: Elements of Engineering Reports, by D. S. Davis. IEEE Transactions on Engineering Writing and Speech, EWS-7(2):27, September 1964. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1964:BRP

[Ano64f] Anonymous. Book review: The Psychology of Communication, by J. Eisenson, et. al. IEEE Transactions on Engineering Writing and Speech, EWS-7 (1):40–43, March 1964. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1964:CCa

[Ano64g] Anonymous. Catalog cards. IEEE Transactions on Engineering Writing and Speech, EWS-7(1):44a–44a, March 1964. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1964:CCb

[Ano64h] Anonymous. Catalog cards.

IEEE Transactions on Engineering Writing and Speech, EWS-7(2):28a–28a, September 1964. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1964:CLP

[Ano64i] Anonymous. Check list for Program Chairman. IEEE Transactions on Engineering Writing and Speech, EWS-7(1):40, March 1964. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1964:CFJ

[Ano64j] Anonymous. Comment on Farrell and Johnson in EWS 6-2. IEEE Transactions on Engineering Writing and Speech, EWS-7(1):40, March 1964. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic). See [Joh63].

Anonymous:1964:CSW

[Ano64k] Anonymous. Comment on the Seminar on Writing Improvement Programs for Engineers. IEEE Transactions on Engineering Writing and Speech, EWS-7(1):39–40, March 1964. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1964:Cd

[Ano64l] Anonymous. Contents. IEEE Transactions on Engineering Writing and Speech, EWS-7 (3):1, December 1964. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1964:Ca

[Ano64m]  Anonymous.  Contributors.

IEEE Transactions on Engineering Writing and Speech, EWS-7(1):44, March 1964. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1964:Cb

[Ano64n]  Anonymous.  Contributors.

IEEE Transactions on Engineering Writing and Speech, EWS-7(2):28, September 1964. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1964:Ce

[Ano64o]  Anonymous.  Contributors.

IEEE Transactions on Engineering Writing and Speech, EWS-7(3):45–46, December 1964. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1964:ECE

[Ano64p] Anonymous. Editor’s comment on Eimbinder letter: a call to PTGEWS members for preparation of 1963–64 EWS Index. IEEE Transactions on Engineering Writing and Speech, EWS-7(1):40, March 1964. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1964:FC

[Ano64q] Anonymous. [Front cover].

IEEE Transactions on Engineering Writing and Speech, EWS-7(3):c1, December 1964. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1964:GEO

[Ano64r] Anonymous. G-EWS Officers, 1964–1965. IEEE Transactions on Engineering Writing and Speech, EWS-7(3):44, December 1964. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1964:IEW

[Ano64s] Anonymous. IEEE Engineering Writing and Speech Group. IEEE Transactions on Engineering Writing and Speech, EWS-7(3):c2, December 1964. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1964:IGE

[Ano64t]  Anonymous.  IEEE Group

on Engineering Writing and Speech. IEEE Transactions on Engineering Writing and Speech, EWS-7(2):c2, September 1964. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1964:IPT

[Ano64u] Anonymous. IEEE Professional Technical Group on Engineering Writing and Speech. IEEE Transactions on Engineering Writing and Speech, EWS-7(1):c2, March 1964. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1964:IAa

[Ano64v] Anonymous. Information for authors. IEEE Transactions on Engineering Writing and Speech, EWS-7(1):45c– 45c, March 1964. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1964:IAb

[Ano64w] Anonymous. Information for authors. IEEE Transactions on Engineering Writing and Speech, EWS-7(2):29c– 29c, September 1964. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1964:IAc

[Ano64x] Anonymous. Information for authors. IEEE Transactions on Engineering Writing and Speech, EWS-7(3):46d–46d, December 1964. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1964:MC

[Ano64y] Anonymous. A message from the Chairman. IEEE Transactions on Engineering Writing and Speech, EWS-7(3):2, December 1964. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1964:OUN

[Ano64z] Anonymous. Our unfortunate name. IEEE Transactions on Engineering Writing and Speech, EWS-7(1):1, March 1964. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1964:PP

[Ano64-27] Anonymous. A precis on prefixes. IEEE Transactions on Engineering Writing and Speech, EWS-7(2):27, September 1964. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1964:QIE

[Ano64-28] Anonymous. A query on indexes to EWS literature. IEEE Transactions on Engineering Writing and Speech, EWS-7(1): 40, March 1964. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1964:SRL

[Ano64-29] Anonymous. The symposium recital … a dialogue. IEEE Transactions on Engineering Writing and Speech, EWS-7 (2):1, September 1964. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1964:TCa

[Ano64-30] Anonymous. Table of contents. IEEE Transactions on Engineering Writing and Speech, EWS-7(1):c1, March 1964. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1964:TCb

[Ano64-31] Anonymous. Table of contents. IEEE Transactions on Engineering Writing and Speech, EWS-7(2):c1, September 1964. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1965:A

[Ano65a] Anonymous. [Advertisement]. IEEE Transactions on Engineering Writing and Speech, EWS-8(2):83–85, December 1965. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1965:Ca

[Ano65b] Anonymous. Contents. IEEE Transactions on Engineering Writing and Speech, EWS-8(1): 1, June 1965. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1965:Cb

[Ano65c]  Anonymous.  Contributors.

IEEE Transactions on Engineering Writing and Speech, EWS-8(1):39, June 1965. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1965:Cc

[Ano65d]  Anonymous.  Contributors.

IEEE Transactions on Engineering Writing and Speech, EWS-8(2):82, December 1965. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1965:FCa

[Ano65e] Anonymous. [Front cover].

IEEE Transactions on Engineering Writing and Speech, EWS-8(1):c1, June 1965. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1965:FCb

[Ano65f] Anonymous. [Front cover].

IEEE Transactions on Engineering Writing and Speech, EWS-8(2):c1, December 1965. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1965:IEW

[Ano65g] Anonymous. IEEE Engineering Writing and Speech Group. IEEE Transactions on Engineering Writing and Speech, EWS-8(2):c2, December 1965. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1965:ISS

[Ano65h] Anonymous. IEEE standard symbols for units. IEEE Transactions on Engineering Writing and Speech, EWS-8(1):18– 26, June 1965. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1965:ITG

[Ano65i] Anonymous. Indexing terms — G-EWS information retrieval. IEEE Transactions on Engineering Writing and Speech, EWS-8(2):47, December 1965. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1965:IA

[Ano65j] Anonymous. Information for authors. IEEE Transactions on Engineering Writing and Speech, EWS-8(2):85a– 85a, December 1965. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1965:IRC

[Ano65k] Anonymous. Information retrieval catalog cards. IEEE Transactions on Engineering Writing and Speech, EWS-8 (1):41–43, June 1965. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1965:IFC

[Ano65l]  Anonymous.  [Inside front

cover]. IEEE Transactions on Engineering Writing and Speech, EWS-8(1):c2, June 1965. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1965:NS

[Ano65m] Anonymous. A new standard.

IEEE Transactions on Engineering Writing and Speech, EWS-8(1):2, June 1965. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1965:SIS

[Ano65n] Anonymous. Sources of information — a study. IEEE Transactions on Engineering Writing and Speech, EWS-8 (2):46, December 1965. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1965:TC

[Ano65o] Anonymous. Table of contents.

IEEE Transactions on Engineering Writing and Speech, EWS-8(2):45, December 1965. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1966:A

[Ano66a] Anonymous. [Advertisement].

IEEE Transactions on Engineering Writing and Speech, EWS-9(1):27, July 1966. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1966:ANA

[Ano66b] Anonymous. Appointment of new Associate Editor. IEEE Transactions on Engineering Writing and Speech, EWS-9(1):

2, July 1966. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1966:Ca

[Ano66c]  Anonymous.  Contributors.

IEEE Transactions on Engineering Writing and Speech, EWS-9(1):26, July 1966. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1966:Cb

[Ano66d]  Anonymous.  Contributors.

IEEE Transactions on Engineering Writing and Speech, EWS-9(2):51, December 1966. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1966:FCa

[Ano66e] Anonymous. [Front cover]. IEEE Transactions on Engineering Writing and Speech, EWS-9(1):c1, July 1966. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1966:FCb

[Ano66f] Anonymous. [Front cover]. IEEE Transactions on Engineering Writing and Speech, EWS-9(2):c1, December 1966. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1966:IEWa

[Ano66g] Anonymous. IEEE Engineering Writing and Speech Group.

IEEE Transactions on Engineering Writing and Speech, EWS-9(1):c2, July 1966. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1966:IEWb

[Ano66h] Anonymous. IEEE Engineering Writing and Speech Group. IEEE Transactions on Engineering Writing and Speech, EWS-9(1):27a–27a, July 1966. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1966:IEWc

[Ano66i] Anonymous. IEEE Engineering Writing and Speech Group. IEEE Transactions on Engineering Writing and Speech, EWS-9(2):c2, December 1966. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1966:IEWd

[Ano66j] Anonymous. IEEE Engineering Writing and Speech Group. IEEE Transactions on Engineering Writing and Speech, EWS-9(2):51a–51a, December 1966. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1966:IAa

[Ano66k] Anonymous. Information for authors. IEEE Transactions on Engineering Writing and Speech, EWS-9(1):27b–27b, July 1966. CODEN IEEWAY.

ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1966:IAb

[Ano66l] Anonymous. Information for authors. IEEE Transactions on Engineering Writing and Speech, EWS-9(2):51b–51b, December 1966. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1966:OO

[Ano66m] Anonymous. Overlooking the obvious? IEEE Transactions on Engineering Writing and Speech, EWS-9(1):3, July 1966. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1966:TCa

[Ano66n] Anonymous. Table of contents. IEEE Transactions on Engineering Writing and Speech, EWS-9(1):1, July 1966. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1966:TCb

[Ano66o] Anonymous. Table of contents. IEEE Transactions on Engineering Writing and Speech, EWS-9(2):nil1, December 1966. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1967:ITC

[Ano67a] Anonymous. 1968 International Technical Communications Conference. IEEE Transactions on Engineering Writing and Speech, EWS-10(2):64, December 1967. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1967:API

[Ano67b] Anonymous. Affiliate plan of the IEEE. IEEE Transactions on Engineering Writing and Speech, EWS-10(1):24, July 1967. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1967:Ca

[Ano67c]  Anonymous.  Contributors.

IEEE Transactions on Engineering Writing and Speech, EWS-10(1):22, July 1967. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1967:Cb

[Ano67d]  Anonymous.  Contributors.

IEEE Transactions on Engineering Writing and Speech, EWS-10(2):63, December 1967. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1967:DEA

[Ano67e] Anonymous. Don’t ever again say “Blood, Toil, Tears and Sweat”. IEEE Transactions on Engineering Writing and Speech, EWS-10(2):46–47, December 1967. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1967:E

[Ano67f] Anonymous. Editorial. IEEE Transactions on Engineering Writing and Speech, EWS-10 (1):2, July 1967. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1967:FCa

[Ano67g] Anonymous. [Front cover]. IEEE Transactions on Engineering Writing and Speech, EWS-10(1):c1, July 1967. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1967:FCb

[Ano67h] Anonymous. [Front cover]. IEEE Transactions on Engineering Writing and Speech, EWS-10(2):c1, December 1967. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1967:IEC

[Ano67i] Anonymous. IEEE–EWS Conference on Improving the Strategy of Oral Presentations. IEEE Transactions on Engineering Writing and Speech, EWS-10(1):21–22, July 1967. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1967:IEWa

[Ano67j] Anonymous. IEEE Engineering Writing and Speech Group. IEEE Transactions on Engineering Writing and Speech, EWS-10(2):c2, December 1967. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1967:IEWb

[Ano67k] Anonymous. IEEE Engineering Writing and Speech Group. IEEE Transactions on Engineering Writing and Speech, EWS-10(2):64a–64a, December 1967. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1967:IA

[Ano67l] Anonymous. Information for authors. IEEE Transactions on Engineering Writing and Speech, EWS-10(2):64b–64b, December 1967. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1967:IFC

[Ano67m]  Anonymous.  [Inside front

cover]. IEEE Transactions on Engineering Writing and Speech, EWS-10(1):c2, July 1967. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1967:TC

[Ano67n] Anonymous. Table of contents. IEEE Transactions on Engineering Writing and Speech, EWS-10(2):25, December 1967. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1967:WEG

[Ano67o] Anonymous. Why an EWS Group? IEEE Transactions on Engineering Writing and Speech, EWS-10(2):26, December 1967. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1968:Ca

[Ano68a]  Anonymous.  Contributors.

IEEE Transactions on Engineering Writing and Speech, EWS-11(2):145–148, August 1968. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1968:Cb

[Ano68b]  Anonymous.  Contributors.

IEEE Transactions on Engineering Writing and Speech, EWS-11(3):165, December 1968. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1968:E

[Ano68c] Anonymous. Editorial. IEEE Transactions on Engineering Writing and Speech, EWS-11 (2):22, August 1968. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1968:FCa

[Ano68d] Anonymous. [Front cover]. IEEE Transactions on Engineering Writing and Speech, EWS-11(1):c1, July 1968. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1968:FCb

[Ano68e] Anonymous. [Front cover].

IEEE Transactions on Engineering Writing and Speech, EWS-11(2):c1, August 1968. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1968:FCc

[Ano68f] Anonymous. [Front cover].

IEEE Transactions on Engineering Writing and Speech, EWS-11(3):c1, December 1968. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1968:GE

[Ano68g]  Anonymous.  Guest edito-

rial. IEEE Transactions on Engineering Writing and Speech, EWS-11(3):150, December 1968. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1968:IEWa

[Ano68h] Anonymous. IEEE Engineering Writing and Speech Group. IEEE Transactions on Engineering Writing and Speech, EWS-11(1):c2, July 1968. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1968:IEWb

[Ano68i]  Anonymous.  IEEE Engi-

neering Writing and Speech Group. IEEE Transactions on Engineering Writing and Speech, EWS-11(1):20a–20a, July 1968. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1968:IEWc

[Ano68j] Anonymous. IEEE Engineering Writing and Speech Group. IEEE Transactions on Engineering Writing and Speech, EWS-11(2):c2, August 1968. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1968:IEWd

[Ano68k] Anonymous. IEEE Engineering Writing and Speech Group. IEEE Transactions on Engineering Writing and Speech, EWS-11(2):148a–148a, August 1968. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1968:IEWe

[Ano68l] Anonymous. IEEE Engineering Writing and Speech Group. IEEE Transactions on Engineering Writing and Speech, EWS-11(3):c2, December 1968. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1968:IEWf

[Ano68m] Anonymous. IEEE Engineering Writing and Speech Group. IEEE Transactions on Engineering Writing and Speech, EWS-11(3):168a–168a, December 1968. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1968:IAa

[Ano68n] Anonymous. Information for authors. IEEE Transactions on Engineering Writing and Speech, EWS-11(1):20b–20b, July 1968. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1968:IAb

[Ano68o] Anonymous. Information for authors. IEEE Transactions on Engineering Writing and Speech, EWS-11(2):148b–148b, August 1968. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1968:IAc

[Ano68p] Anonymous. Information for authors. IEEE Transactions on Engineering Writing and Speech, EWS-11(3):168b–168b, December 1968. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1968:NE

[Ano68q]  Anonymous.  New Editor.

IEEE Transactions on Engineering Writing and Speech, EWS-11(3):151, December 1968. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1968:RHE

[Ano68r] Anonymous. On the reading habits of engineer–authors. IEEE Transactions on Engineering Writing and Speech, EWS-11(1):17, July 1968. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1968:P

[Ano68s] Anonymous. Preface. IEEE Transactions on Engineering Writing and Speech, EWS-11 (2):22, August 1968. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1968:RRC

[Ano68t] Anonymous. Reprints, reviews, and current awareness. IEEE Transactions on Engineering Writing and Speech, EWS-11 (1):2, July 1968. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1968:SFF

[Ano68u] Anonymous. The secretary’s friend, or a form for the preparation of mathematical expressions for typing. IEEE Transactions on Engineering Writing and Speech, EWS-11(1):17– 19, July 1968. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1968:SRV

[Ano68v] Anonymous. Speed reading versus speed writing. IEEE Transactions on Engineering Writing and Speech, EWS-11 (1):19–20, July 1968. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1968:SPY

[Ano68w]  Anonymous.  Symbols for

proofreading your own copy or what to do while waiting for the printer. IEEE Transactions on Engineering Writing and Speech, EWS-11(3): 166, December 1968. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1968:TCa

[Ano68x] Anonymous. Table of contents.

IEEE Transactions on Engineering Writing and Speech, EWS-11(1):1, July 1968. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1968:TCb

[Ano68y] Anonymous. Table of contents.

IEEE Transactions on Engineering Writing and Speech, EWS-11(2):21, August 1968. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1968:TCc

[Ano68z] Anonymous. Table of contents. IEEE Transactions on Engineering Writing and Speech, EWS-11(3):149, December 1968. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1968:VQN

[Ano68-27] Anonymous. The value of questions from the nonspecialist reviewer. IEEE Transactions on Engineering Writing and Speech, EWS-11(1):19, July 1968. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1969:A

[Ano69a]  Anonymous.  [Advertise-

ment]. IEEE Transactions on Engineering Writing and Speech, EWS-12(2):66a–66a, May 1969. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1969:Cb

[Ano69b]  Anonymous.  Contributors.

IEEE Transactions on Engineering Writing and Speech, EWS-12(1):20, May 1969. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1969:Cc

[Ano69c]  Anonymous.  Contributors.

IEEE Transactions on Engineering Writing and Speech, EWS-12(3):90, October 1969. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1969:Ca

[Ano69d] Anonymous. Correction. IEEE Transactions on Engineering Writing and Speech, EWS-12 (1):19, May 1969. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1969:E

[Ano69e] Anonymous. Editorial. IEEE Transactions on Engineering Writing and Speech, EWS-12 (1):2, May 1969. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1969:FCa

[Ano69f] Anonymous. [Front cover].

IEEE Transactions on Engineering Writing and Speech, EWS-12(1):c1, May 1969. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1969:FCb

[Ano69g] Anonymous. [Front cover].

IEEE Transactions on Engineering Writing and Speech, EWS-12(2):c1, May 1969. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1969:FCc

[Ano69h] Anonymous. [Front cover].

IEEE Transactions on Engineering Writing and Speech, EWS-12(3):c1, October 1969. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1969:GE

[Ano69i] Anonymous. Guest editorial.

IEEE Transactions on Engineering Writing and Speech, EWS-12(2):66, May 1969. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1969:IEWa

[Ano69j] Anonymous. IEEE Engineering Writing and Speech Group. IEEE Transactions on Engineering Writing and Speech, EWS-12(1):c2, May 1969. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1969:IEWb

[Ano69k]  Anonymous.  IEEE Engi-

neering Writing and Speech Group. IEEE Transactions on Engineering Writing and Speech, EWS-12(1):20a–20a, May 1969. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1969:IEWc

[Ano69l] Anonymous. IEEE Engineering Writing and Speech Group. IEEE Transactions on Engineering Writing and Speech, EWS-12(2):c2, May 1969. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1969:IEWd

[Ano69m] Anonymous. IEEE Engineering Writing and Speech Group. IEEE Transactions on Engineering Writing and Speech, EWS-12(3):c2, October 1969.

CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1969:IEWe

[Ano69n] Anonymous. IEEE Engineering Writing and Speech Group. IEEE Transactions on Engineering Writing and Speech, EWS-12(3):91a–91a, October 1969. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1969:IAa

[Ano69o] Anonymous. Information for authors. IEEE Transactions on Engineering Writing and Speech, EWS-12(1):20b–20b, May 1969. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1969:IAb

[Ano69p] Anonymous. Information for authors. IEEE Transactions on Engineering Writing and Speech, EWS-12(2):66b–66b, May 1969. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1969:IAc

[Ano69q] Anonymous. Information for authors. IEEE Transactions on Engineering Writing and Speech, EWS-12(3):91b–91b, October 1969. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1969:Pb

[Ano69r]  Anonymous.  Participants.

IEEE Transactions on Engineering Writing and Speech, EWS-12(2):64–65, May 1969. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1969:Pa

[Ano69s] Anonymous. Preface. IEEE Transactions on Engineering Writing and Speech, EWS-12 (2):22, May 1969. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1969:RI

[Ano69t] Anonymous. A review of the issue. IEEE Transactions on Engineering Writing and Speech, EWS-12(3):68, October 1969. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1969:TCa

[Ano69u] Anonymous. Table of contents. IEEE Transactions on Engineering Writing and Speech, EWS-12(1):1, May 1969. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1969:TCb

[Ano69v] Anonymous. Table of contents. IEEE Transactions on Engineering Writing and Speech, EWS-12(2):21, May 1969. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1969:TCc

[Ano69w] Anonymous. Table of contents.

IEEE Transactions on Engineering Writing and Speech, EWS-12(3):67, October 1969. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1969:TPO

[Ano69x] Anonymous. Turn page orientation. IEEE Transactions on Engineering Writing and Speech, EWS-12(3):89, October 1969. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic). See comments [Ano70c, Ano71e, Ano71m].

Anonymous:1970:A

[Ano70a] Anonymous. Announcement.

IEEE Transactions on Engineering Writing and Speech, EWS-13(2):92, September 1970. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1970:CIS

[Ano70b] Anonymous. Comments of IEEE Spectrum policy for plurals of Latin terms. IEEE Transactions on Engineering Writing and Speech, EWS-13 (1):28–29, May 1970. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1970:CTP

[Ano70c] Anonymous. Comments on “Turn Page Orientation”.

IEEE Transactions on Engineering Writing and Speech, EWS-13(1):28, May 1970. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic). See [Ano69x].

Anonymous:1970:Ca

[Ano70d]  Anonymous.  Contributors.

IEEE Transactions on Engineering Writing and Speech, EWS-13(1):30–31, May 1970. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1970:Cb

[Ano70e]  Anonymous.  Contributors.

IEEE Transactions on Engineering Writing and Speech, EWS-13(2):90–91, September 1970. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1970:E

[Ano70f] Anonymous. Editorial. IEEE Transactions on Engineering Writing and Speech, EWS-13 (2):33, September 1970. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1970:FC

[Ano70g] Anonymous. [Front cover]. IEEE Transactions on Engineering Writing and Speech, EWS-13(1):c1, May 1970. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1970:GE

[Ano70h] Anonymous. Guest editorial.

IEEE Transactions on Engineering Writing and Speech, EWS-13(1):1, May 1970. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1970:IEWa

[Ano70i] Anonymous. IEEE Engineering Writing and Speech Group. IEEE Transactions on Engineering Writing and Speech, EWS-13(1):c2, May 1970. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1970:IEWb

[Ano70j]  Anonymous.  IEEE Engi-

neering Writing and Speech Group. IEEE Transactions on Engineering Writing and Speech, EWS-13(1):31a–31a, May 1970. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1970:IEWc

[Ano70k]  Anonymous.  IEEE Engi-

neering Writing and Speech Group. IEEE Transactions on Engineering Writing and Speech, EWS-13(2):c2, September 1970. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1970:IEWd

[Ano70l] Anonymous. IEEE Engineering Writing and Speech Group.

IEEE Transactions on Engineering Writing and Speech, EWS-13(2):92a–92a, September 1970. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1970:IAa

[Ano70m] Anonymous. Information for authors. IEEE Transactions on Engineering Writing and Speech, EWS-13(1):31b–31b, May 1970. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1970:IAb

[Ano70n] Anonymous. Information for authors. IEEE Transactions on Engineering Writing and Speech, EWS-13(2):92b–92b, September 1970. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1970:TC

[Ano70o] Anonymous. Table of contents. IEEE Transactions on Engineering Writing and Speech, EWS-13(2):c1, September 1970. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1971:IIT

[Ano71a] Anonymous. 1971 index IEEE Transactions on Engineering Writing and Speech vol. EWS-14. IEEE Transactions on Engineering Writing and Speech, EWS-14(3):80a–80b, December 1971. CODEN IEEWAY.

ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1971:A

[Ano71b] Anonymous. Acknowledgment.

IEEE Transactions on Engineering Writing and Speech, EWS-14(1):38, February 1971. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1971:TRW

[Ano71c] Anonymous. Technical report writing — Herman M. Weisman. IEEE Transactions on Engineering Writing and Speech, EWS-14(2):61, June 1971. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1971:UTA

[Ano71d] Anonymous. Using technical art, an industry guide — George A. Magnan. IEEE Transactions on Engineering Writing and Speech, EWS-14 (2):60, June 1971. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1971:CTP

[Ano71e] Anonymous. Comments on “Turn Page Orientation”. IEEE Transactions on Engineering Writing and Speech, EWS-14(1):38, February 1971. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic). See [Ano69x].

Anonymous:1971:Cb

[Ano71f]  Anonymous.  Contributors.

IEEE Transactions on Engineering Writing and Speech, EWS-14(1):39, February 1971. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1971:Cc

[Ano71g]  Anonymous.  Contributors.

IEEE Transactions on Engineering Writing and Speech, EWS-14(2):62, June 1971. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1971:Cd

[Ano71h]  Anonymous.  Contributors.

IEEE Transactions on Engineering Writing and Speech, EWS-14(3):80, December 1971. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1971:Ca

[Ano71i] Anonymous. Correction. IEEE Transactions on Engineering Writing and Speech, EWS-14 (1):38, February 1971. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1971:Ea

[Ano71j] Anonymous. Editorial. IEEE Transactions on Engineering Writing and Speech, EWS-14 (1):1–2, February 1971. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1971:Eb

[Ano71k] Anonymous. Editorial. IEEE Transactions on Engineering Writing and Speech, EWS-14 (3):63, December 1971. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1971:EML

[Ano71l] Anonymous. Editorial: a matter of life and death. IEEE Transactions on Engineering Writing and Speech, EWS-14 (2):41, June 1971. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1971:FCT

[Ano71m] Anonymous. Further comments on “Turn Page Orientation”. IEEE Transactions on Engineering Writing and Speech, EWS-14(1):38, February 1971. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic). See [Ano69x].

Anonymous:1971:IEWa

[Ano71n] Anonymous. IEEE Engineering Writing and Speech Group. IEEE Transactions on Engineering Writing and Speech, EWS-14(1):c2, February 1971. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1971:IEWb

[Ano71o] Anonymous. IEEE Engineering Writing and Speech Group. IEEE Transactions on Engineering Writing and Speech, EWS-14(1):39a–39a, February 1971. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1971:IEWc

[Ano71p] Anonymous. IEEE Engineering Writing and Speech Group. IEEE Transactions on Engineering Writing and Speech, EWS-14(2):c2, June 1971. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1971:IEWd

[Ano71q]  Anonymous.  IEEE Engi-

neering Writing and Speech Group. IEEE Transactions on Engineering Writing and Speech, EWS-14(2):62a–62a, June 1971. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1971:IEWe

[Ano71r] Anonymous. IEEE Engineering Writing and Speech Group. IEEE Transactions on Engineering Writing and Speech, EWS-14(3):c2, December 1971. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1971:IEWf

[Ano71s] Anonymous. IEEE Engineering Writing and Speech Group.

IEEE Transactions on Engineering Writing and Speech, EWS-14(3):80c–80c, December 1971. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1971:IAa

[Ano71t] Anonymous. Information for authors. IEEE Transactions on Engineering Writing and Speech, EWS-14(1):39b–39b, February 1971. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1971:IAb

[Ano71u] Anonymous. Information for authors. IEEE Transactions on Engineering Writing and Speech, EWS-14(2):62b–62b, June 1971. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1971:IAc

[Ano71v] Anonymous. Information for authors. IEEE Transactions on Engineering Writing and Speech, EWS-14(3):80d–80d, December 1971. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1971:LE

[Ano71w] Anonymous. Letter to the editor. IEEE Transactions on Engineering Writing and Speech, EWS-14(2):60, June 1971. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1971:SD

[Ano71x] Anonymous. Second death. IEEE Transactions on Engineering Writing and Speech, EWS-14(2):60, June 1971. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1971:TCa

[Ano71y] Anonymous. Table of contents. IEEE Transactions on Engineering Writing and Speech, EWS-14(1):c1, February 1971. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1971:TCb

[Ano71z] Anonymous. Table of contents. IEEE Transactions on Engineering Writing and Speech, EWS-14(2):c1, June 1971. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1971:TCc

[Ano71-27] Anonymous. Table of contents. IEEE Transactions on Engineering Writing and Speech, EWS-14(3):c1, December 1971. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Anonymous:1972:BC

[Ano72a] Anonymous. [Back cover]. IEEE Transactions on Professional Communication, PC-15 (4):c4, December 1972. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1972:Ca

[Ano72b]  Anonymous.  Contributors.

IEEE Transactions on Professional Communication, PC-15 (1):18, March 1972. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1972:Cb

[Ano72c]  Anonymous.  Contributors.

IEEE Transactions on Professional Communication, PC-15 (2):68–70, June 1972. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1972:Cc

[Ano72d]  Anonymous.  Contributors.

IEEE Transactions on Professional Communication, PC-15 (3):105–106, September 1972. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1972:Cd

[Ano72e]  Anonymous.  Contributors.

IEEE Transactions on Professional Communication, PC-15 (4):124, December 1972. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1972:FCa

[Ano72f] Anonymous. [Front cover].

IEEE Transactions on Professional Communication, PC-15 (1):c1, March 1972. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1972:FCb

[Ano72g] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, PC-15 (2):c1, June 1972. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1972:FCc

[Ano72h] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, PC-15 (3):c1, September 1972. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1972:FCd

[Ano72i] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, PC-15 (4):c1, December 1972. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1972:FIC

[Ano72j]  Anonymous.  [Front inside

cover]. IEEE Transactions on Professional Communication, PC-15(4):c2, December 1972. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1972:IBC

[Ano72k] Anonymous. Inside back cover. IEEE Transactions on Professional Communication, PC-15 (4):c3, December 1972. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1972:IFCa

[Ano72l] Anonymous. Inside front cover.

IEEE Transactions on Professional Communication, PC-15 (1):c2, March 1972. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1972:IFCb

[Ano72m] Anonymous. Inside front cover.

IEEE Transactions on Professional Communication, PC-15 (2):c2, June 1972. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1972:IFCc

[Ano72n] Anonymous. Inside front cover.

IEEE Transactions on Professional Communication, PC-15 (3):c2, September 1972. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1973:IIT

[Ano73a] Anonymous. 1973 index IEEE Transactions on Professional Communication vol. PC-16. IEEE Transactions on Professional Communication, PC-16 (4):1–6, December 1973. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1973:Aa

[Ano73b] Anonymous. Advertisements. IEEE Transactions on Professional Communication, PC-16 (1):24, March 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1973:Ab

[Ano73c] Anonymous. Advertisements. IEEE Transactions on Professional Communication, PC-16 (2):1, June 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1973:Ac

[Ano73d] Anonymous. Advertisements. IEEE Transactions on Professional Communication, PC-16 (3):191, September 1973. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1973:Ad

[Ano73e] Anonymous. Advertisements. IEEE Transactions on Professional Communication, PC-16 (4):1, December 1973. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1973:BC

[Ano73f] Anonymous. [Back cover]. IEEE Transactions on Professional Communication, PC-16 (4):c4, December 1973. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1973:CS

[Ano73g] Anonymous. Conference summary. IEEE Transactions on Professional Communication, PC-16(3):50–53, September 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1973:Ca

[Ano73h]  Anonymous.  Contributors.

IEEE Transactions on Professional Communication, PC-16 (1):22–23, March 1973. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1973:Cb

[Ano73i]  Anonymous.  Contributors.

IEEE Transactions on Professional Communication, PC-16 (2):44, June 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1973:Cc

[Ano73j]  Anonymous.  Contributors.

IEEE Transactions on Professional Communication, PC-16 (3):183–189, September 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1973:Cd

[Ano73k]  Anonymous.  Contributors.

IEEE Transactions on Professional Communication, PC-16 (4):217, December 1973. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1973:E

[Ano73l] Anonymous. Editorial. IEEE Transactions on Professional Communication, PC-16(4): 193, December 1973. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1973:EGP

[Ano73m] Anonymous. Editorial: G-PC in ’73. A preview of ’74. IEEE Transactions on Professional Communication, PC-16 (1):1, March 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1973:FCa

[Ano73n] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, PC-16 (1):c1, March 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1973:FCb

[Ano73o] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, PC-16 (2):c1, June 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1973:FCc

[Ano73p] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, PC-16 (3):c1, September 1973. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1973:FCd

[Ano73q] Anonymous. [Front cover].

IEEE Transactions on Professional Communication, PC-16 (4):c1, December 1973. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1973:IBC

[Ano73r] Anonymous. Inside back cover.

IEEE Transactions on Professional Communication, PC-16 (4):c3, December 1973. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1973:IFCa

[Ano73s] Anonymous. Inside front cover.

IEEE Transactions on Professional Communication, PC-16 (1):c2, March 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1973:IFCb

[Ano73t] Anonymous. Inside front cover.

IEEE Transactions on Professional Communication, PC-16 (2):c2, June 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1973:IFCc

[Ano73u] Anonymous. Inside front cover.

IEEE Transactions on Professional Communication, PC-16 (3):c2, September 1973. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1973:IFCd

[Ano73v] Anonymous. Inside front cover. IEEE Transactions on Professional Communication, PC-16 (4):c2, December 1973. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1973:LEa

[Ano73w] Anonymous. Letters to the Editor. IEEE Transactions on Professional Communication, PC-16(2):41–43, June 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1973:LEb

[Ano73x] Anonymous. Letters to the Editor. IEEE Transactions on Professional Communication, PC-16(4):215, December 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1973:LYF

[Ano73y]  Anonymous.  A look to

your future IEEE can help you stay current technically through groups and societies. IEEE Transactions on Professional Communication, PC-16 (3):190, September 1973. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1973:P

[Ano73z] Anonymous. Preface. IEEE Transactions on Professional Communication, PC-16(3):49, September 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1973:SVR

[Ano73-27] Anonymous. Session V.

Roundup and critique. IEEE Transactions on Professional Communication, PC-16(3): 156–182, September 1973. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1974:ITP

[Ano74a] Anonymous. 1974 IEEE Transactions on Professional Communication vol. PC-17. IEEE Transactions on Professional Communication, PC-17(3/4): 1–3, September/December 1974. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1974:A

[Ano74b] Anonymous. Advertisements.

IEEE Transactions on Professional Communication, PC-17 (1):30, March 1974. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1974:Ca

[Ano74c]  Anonymous.  Conributors.

IEEE Transactions on Professional Communication, PC-17 (1):28–29, March 1974. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1974:Cb

[Ano74d]  Anonymous.  Contributors.

IEEE Transactions on Professional Communication, PC-17 (2):42, June 1974. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1974:FCa

[Ano74e] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, PC-17 (1):c1, March 1974. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1974:FCb

[Ano74f] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, PC-17 (2):c1, June 1974. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1974:FCc

[Ano74g] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, PC-17 (3/4):c1, September/December 1974. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1974:IFC

[Ano74h] Anonymous. Inside front cover. IEEE Transactions on Professional Communication, PC-17 (1):c2, March 1974. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1974:MTR

[Ano74i] Anonymous. Metric terms: - er or -re? IEEE Transactions on Professional Communication, PC-17(3/4):54– 62, September/December 1974. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1975:IIT

[Ano75a] Anonymous. 1975 index IEEE Transactions on Professional Communication vol. PC-18. IEEE Transactions on Professional Communication, PC-18 (4):1–4, December 1975. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1975:A

[Ano75b] Anonymous. Announcement.

IEEE Transactions on Professional Communication, PC-18 (4):336, December 1975. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1975:BC

[Ano75c] Anonymous. [Back cover].

IEEE Transactions on Professional Communication, PC-18 (3):c4, September 1975. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1975:CP

[Ano75d] Anonymous. Call for papers.

IEEE Transactions on Professional Communication, PC-18 (4):335, December 1975. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1975:Ca

[Ano75e]  Anonymous.  Contributors.

IEEE Transactions on Professional Communication, PC-18 (1):49, March 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1975:Cb

[Ano75f]  Anonymous.  Contributors.

IEEE Transactions on Professional Communication, PC-18 (1):56, March 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1975:Cd

[Ano75g]  Anonymous.  Contributors.

IEEE Transactions on Professional Communication, PC-18 (2):77, June 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1975:Cc

[Ano75h] Anonymous. Correspondence. IEEE Transactions on Professional Communication, PC-18 (2):74–76, June 1975. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1975:FCa

[Ano75i] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, PC-18 (1):c1, March 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1975:FCb

[Ano75j] Anonymous. [Front cover].

IEEE Transactions on Professional Communication, PC-18 (2):c1, June 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1975:FCc

[Ano75k] Anonymous. [Front cover].

IEEE Transactions on Professional Communication, PC-18 (3):c1, September 1975. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1975:FCd

[Ano75l] Anonymous. [Front cover].

IEEE Transactions on Professional Communication, PC-18 (4):c1, December 1975. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1975:ICS

[Ano75m] Anonymous. IEEE Conference on Scientific Journals, Record, 1975. IEEE Transactions on Professional Communication, PC-18(3):??, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1975:IBC

[Ano75n] Anonymous. Inside back cover.

IEEE Transactions on Professional Communication, PC-18 (4):c3, December 1975. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1975:P

[Ano75o] Anonymous. Preface. IEEE Transactions on Professional Communication, PC-18(3):85– 92, September 1975. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1976:IIT

[Ano76a] Anonymous. 1976 index IEEE Transactions on Professional Communication vol. PC-19. IEEE Transactions on Professional Communication, PC-19 (2):1, June 1976. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1976:BCa

[Ano76b] Anonymous. [Back cover]. IEEE Transactions on Professional Communication, PC-19 (1):c4, March 1976. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1976:BCb

[Ano76c] Anonymous. [Back cover]. IEEE Transactions on Professional Communication, PC-19 (2):c4, June 1976. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1976:FCa

[Ano76d] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, PC-19 (1):c1, March 1976. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1976:FCb

[Ano76e] Anonymous. [Front cover].

IEEE Transactions on Professional Communication, PC-19 (2):c1, June 1976. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1976:FIC

[Ano76f]  Anonymous.  [Front inside

cover]. IEEE Transactions on Professional Communication, PC-19(1):c2, March 1976. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1977:AGA

[Ano77a] Anonymous. Alfred N. Goldsmith award of the IEEE Group on Professional Communication. IEEE Transactions on Professional Communication, PC-20(1):45–46, June 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1977:BC

[Ano77b] Anonymous. [Back cover].

IEEE Transactions on Professional Communication, PC-20 (2):c4, September 1977. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1977:ANS

[Ano77c] Anonymous. American National Standard for Writing Abstracts. IEEE Transactions on Professional Communication, PC-20(4):252–254, December 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1977:CPa

[Ano77d] Anonymous. Call for papers. IEEE Transactions on Professional Communication, PC-20 (2):138, September 1977. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1977:CPb

[Ano77e] Anonymous. Call for papers. IEEE Transactions on Professional Communication, PC-20 (3):203, November 1977. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1977:CU

[Ano77f] Anonymous. Computer uses. IEEE Transactions on Professional Communication, PC-20 (3):199, November 1977. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1977:CCC

[Ano77g] Anonymous. Copyright clearance center. IEEE Transactions on Professional Communication, PC-20(3):174–176, November 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1977:FU

[Ano77h] Anonymous. Fair use. IEEE Transactions on Professional Communication, PC-20(3): 190–194, November 1977. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1977:FCa

[Ano77i] Anonymous. [Front cover].

IEEE Transactions on Professional Communication, PC-20 (1):c1, June 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1977:FCb

[Ano77j] Anonymous. [Front cover].

IEEE Transactions on Professional Communication, PC-20 (2):c1, September 1977. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1977:FCc

[Ano77k] Anonymous. [Front cover].

IEEE Transactions on Professional Communication, PC-20 (3):c1, November 1977. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1977:FCd

[Ano77l] Anonymous. [Front cover].

IEEE Transactions on Professional Communication, PC-20 (4):c1, December 1977. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1977:IGPa

[Ano77m]  Anonymous.  IEEE Group

on Professional Communication information for authors. IEEE Transactions on Professional Communication, PC-20 (1):46, June 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1977:IGPc

[Ano77n]  Anonymous.  IEEE Group

on Professional Communication information for authors. IEEE Transactions on Professional Communication, PC-20 (2):139, September 1977. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1977:IGPd

[Ano77o]  Anonymous.  IEEE Group

on Professional Communication information for authors. IEEE Transactions on Professional Communication, PC-20 (3):204, November 1977. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1977:IGPe

[Ano77p]  Anonymous.  IEEE Group

on Professional Communication information for authors. IEEE Transactions on Professional Communication, PC-20 (4):257, December 1977. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1977:IGPb

[Ano77q]  Anonymous.  IEEE Group

on Professional Communication membership information. IEEE Transactions on Professional Communication, PC-20 (1):47, June 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1977:IMI

[Ano77r] Anonymous. IEEE membership information. IEEE Transactions on Professional Communication, PC-20(1):48, June 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1977:IBCa

[Ano77s] Anonymous. Inside back cover.

IEEE Transactions on Professional Communication, PC-20 (1):c3, June 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1977:IBCb

[Ano77t] Anonymous. Inside back cover.

IEEE Transactions on Professional Communication, PC-20 (2):c3, September 1977. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1977:IBCc

[Ano77u] Anonymous. Inside back cover. IEEE Transactions on Professional Communication, PC-20 (3):c3, November 1977. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1977:IFCa

[Ano77v] Anonymous. Inside front cover. IEEE Transactions on Professional Communication, PC-20 (2):c2, September 1977. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1977:IFCb

[Ano77w] Anonymous. Inside front cover. IEEE Transactions on Professional Communication, PC-20 (3):c2, November 1977. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1977:IFCc

[Ano77x] Anonymous. Inside front cover. IEEE Transactions on Professional Communication, PC-20 (4):c2, December 1977. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1977:INC

[Ano77y]  Anonymous.  Introduction

to the new copyright law. IEEE Transactions on Professional Communication, PC-20 (3):142–146, November 1977.

CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1977:OC

[Ano77z] Anonymous. Ownership of copyright. IEEE Transactions on Professional Communication, PC-20(3):199–202, November 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1977:P

[Ano77-27] Anonymous. Preface. IEEE Transactions on Professional Communication, PC-20(2):49– 51, September 1977. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1977:PCG

[Ano77-28] Anonymous. Professional communication group membership information. IEEE Transactions on Professional Communication, PC-20(2): 140, September 1977. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1977:RLA

[Ano77-29] Anonymous. Reproduction by libraries and archives. IEEE Transactions on Professional Communication, PC-20 (3):194–198, November 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1977:SIA

[Ano77-30] Anonymous. Special issue: Audio/visual communication call for papers. IEEE Transactions on Professional Communication, PC-20(4):205, December 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1977:SUS

[Ano77-31] Anonymous. Speech understanding systems. IEEE Transactions on Professional Communication, PC-20(4):221–225, December 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1977:SMC

[Ano77-32] Anonymous. Subject matter of copyright. IEEE Transactions on Professional Communication, PC-20(3):184–188, November 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1977:TP

[Ano77-33] Anonymous. [Title page]. IEEE Transactions on Professional Communication, PC-20 (1):c1, June 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1977:WUG

[Ano77-34] Anonymous. Works of the U.S. government. IEEE Transactions on Professional Communication, PC-20(3):188–189, November 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1978:IIT

[Ano78a] Anonymous. 1978 index IEEE Transactions on Professional Communication vol. PC-21. IEEE Transactions on Professional Communication, PC-21 (4):1–4, December 1978. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1978:Aa

[Ano78b] Anonymous. Advertisements.

IEEE Transactions on Professional Communication, PC-21 (3):134–136, September 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1978:Ab

[Ano78c] Anonymous. Advertisements.

IEEE Transactions on Professional Communication, PC-21 (4):172, December 1978. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1978:AGA

[Ano78d] Anonymous. Alfred N. Goldsmith award of the IEEE Professional Communication Society. IEEE Transactions on Professional Communication, PC-21(2):88, June 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1978:FLT

[Ano78e]  Anonymous.  Federalese,

legalese, technicalese. IEEE Transactions on Professional Communication, PC-21(1):30– 31, March 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1978:FCa

[Ano78f] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, PC-21 (1):c1, March 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1978:FCb

[Ano78g] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, PC-21 (2):c1, June 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1978:FCc

[Ano78h] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, PC-21 (3):c1, September 1978. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1978:FCd

[Ano78i] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, PC-21 (4):c1, December 1978. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1978:H

[Ano78j] Anonymous. Hopefully. IEEE Transactions on Professional Communication, PC-21(1):29– 30, March 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1978:HAC

[Ano78k]  Anonymous.  How to ap-

ply for copyright registration. IEEE Transactions on Professional Communication, PC-21 (2):79–85, June 1978. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1978:IPCa

[Ano78l] Anonymous. IEEE Professional Communication Society information for authors. IEEE Transactions on Professional Communication, PC-21 (1):40, March 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1978:IPCb

[Ano78m] Anonymous. IEEE Professional Communication Society information for authors. IEEE Transactions on Professional Communication, PC-21 (3):131, September 1978. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1978:IPCd

[Ano78n] Anonymous. IEEE Professional Communication Society information for authors.

IEEE Transactions on Professional Communication, PC-21 (4):171, December 1978. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1978:IPCc

[Ano78o] Anonymous. IEEE Professional Communication Society membership information. IEEE Transactions on Professional Communication, PC-21 (3):132–133, September 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1978:MFA

[Ano78p] Anonymous. Michael Faraday’s advice to a lecturer. IEEE Transactions on Professional Communication, PC-21 (1):2–5, March 1978. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1978:TP

[Ano78q]  Anonymous.  [Title page].

IEEE Transactions on Professional Communication, PC-21 (1):1, March 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1978:UND

[Ano78r] Anonymous. US news data terminals compress deadline schedule. IEEE Transactions on Professional Communication, PC-21(1):23–24, March 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1979:IIT

[Ano79a] Anonymous. 1979 index IEEE Transactions on Professional Communication vol. PC-22. IEEE Transactions on Professional Communication, PC-22 (4):1–5, December 1979. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1979:AGA

[Ano79b] Anonymous. Alfred N. Goldsmith award of the IEEE Professional Communication Society. IEEE Transactions on Professional Communication, PC-22(1):42, March 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1979:BCa

[Ano79c] Anonymous. [Back cover].

IEEE Transactions on Professional Communication, PC-22 (2):c4, June 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1979:BCb

[Ano79d] Anonymous. [Back cover].

IEEE Transactions on Professional Communication, PC-22 (4):c4, December 1979. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1979:CTP

[Ano79e]  Anonymous.  Communicat-

ing through poster sessions. IEEE Transactions on Professional Communication, PC-22 (3):137–140, September 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1979:Ca

[Ano79f] Anonymous. Contents. IEEE Transactions on Professional Communication, PC-22(1): 1, March 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1979:Cb

[Ano79g] Anonymous. Contents. IEEE Transactions on Professional Communication, PC-22(2):1, June 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1979:Cc

[Ano79h] Anonymous. Contents. IEEE Transactions on Professional Communication, PC-22(3):1, September 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1979:Cd

[Ano79i] Anonymous. Contents. IEEE Transactions on Professional Communication, PC-22(4):1, December 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1979:FCa

[Ano79j] Anonymous. [Front cover].

IEEE Transactions on Professional Communication, PC-22 (1):c1, March 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1979:FCb

[Ano79k] Anonymous. [Front cover].

IEEE Transactions on Professional Communication, PC-22 (2):c1, June 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1979:FCc

[Ano79l] Anonymous. [Front cover].

IEEE Transactions on Professional Communication, PC-22 (3):c1, September 1979. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1979:FCd

[Ano79m] Anonymous. [Front cover].

IEEE Transactions on Professional Communication, PC-22 (4):c1, December 1979. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1979:GPD

[Ano79n] Anonymous. Guide for patent drawings. IEEE Transactions on Professional Communication, PC-22(2):109–111, June 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1979:ICF

[Ano79o] Anonymous. IEEE copyright form. IEEE Transactions on Professional Communication, PC-22(1):43–44, March 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1979:IARa

[Ano79p]  Anonymous.  Information

for authors and readers of the IEEE Transactions on Professional Communication. IEEE Transactions on Professional Communication, PC-22(1):45, March 1979. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1979:IARb

[Ano79q]  Anonymous.  Information

for authors and readers of the IEEE Transactions on Professional Communication. IEEE Transactions on Professional Communication, PC-22(2):128, June 1979. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1979:IARc

[Ano79r]  Anonymous.  Information

for authors and readers of the IEEE Transactions on Professional Communication. IEEE Transactions on Professional Communication, PC-22 (3):177, September 1979. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1979:IFCa

[Ano79s]  Anonymous.  [Inside front

cover]. IEEE Transactions on Professional Communication, PC-22(1):c2, March 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1979:IFCb

[Ano79t]  Anonymous.  [Inside front

cover]. IEEE Transactions on Professional Communication, PC-22(2):c2, June 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1979:IFCc

[Ano79u]  Anonymous.  [Inside front

cover]. IEEE Transactions on Professional Communication, PC-22(3):c2, September 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1979:IFCd

[Ano79v]  Anonymous.  [Inside front

cover]. IEEE Transactions on Professional Communication, PC-22(4):c2, December 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1980:IIT

[Ano80a] Anonymous. 1980 index IEEE Transactions on Professional Communication vol. PC-23.

IEEE Transactions on Professional Communication, PC-23 (4):1–5, December 1980. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1980:IPCd

[Ano80b]  Anonymous.  1981 IEEE

Professional Communication Society Conference. IEEE Transactions on Professional Communication, PC-23(4): 205, December 1980. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1980:AGA

[Ano80c] Anonymous. Alfred N. Goldsmith award of the IEEE Professional Communication Society. IEEE Transactions on Professional Communication, PC-23(2):111, June 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1980:BC

[Ano80d] Anonymous. [Back cover]. IEEE Transactions on Professional Communication, PC-23 (4):c4, December 1980. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1980:BCN

[Ano80e]  Anonymous.  Better com-

munication needed. IEEE Transactions on Professional Communication, PC-23(3):

115, September 1980. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1980:CCC

[Ano80f] Anonymous. Chase, chance, and creativity, James H. Austin. IEEE Transactions on Professional Communication, PC-23(2):104–105, June 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1980:CCP

[Ano80g] Anonymous. Communication:

a creative process, Eugene R. Moulton and McD. W. Held. IEEE Transactions on Professional Communication, PC-23 (2):101, June 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1980:CCE

[Ano80h]  Anonymous.  Communica-

tion as creative experience, J. Samuel Bois. IEEE Transactions on Professional Communication, PC-23(2):101, June 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1980:ACA

[Ano80i] Anonymous. The act of creation, Arthur Koestler. IEEE Transactions on Professional Communication, PC-23(2): 103–104, June 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1980:CEA

[Ano80j]  Anonymous.  The creative

engineer: The art of inventing, Winston E. Kock. IEEE Transactions on Professional Communication, PC-23 (2):105, June 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1980:Ca

[Ano80k] Anonymous. Contents. IEEE Transactions on Professional Communication, PC-23(1): 1, March 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1980:Cb

[Ano80l] Anonymous. Contents. IEEE Transactions on Professional Communication, PC-23(2):1, June 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1980:Cc

[Ano80m] Anonymous. Contents. IEEE Transactions on Professional Communication, PC-23(3):1, September 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1980:Cd

[Ano80n] Anonymous. Contents. IEEE Transactions on Professional Communication, PC-23(4):1, December 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1980:DCS

[Ano80o] Anonymous. Documentation for construction supervisors. IEEE Transactions on Professional Communication, PC-23 (4):198–200, December 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1980:FCa

[Ano80p] Anonymous. [Front cover].

IEEE Transactions on Professional Communication, PC-23 (1):c1, March 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1980:FCb

[Ano80q] Anonymous. [Front cover].

IEEE Transactions on Professional Communication, PC-23 (2):c1, June 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1980:FCc

[Ano80r] Anonymous. [Front cover].

IEEE Transactions on Professional Communication, PC-23 (3):c1, September 1980. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1980:FCd

[Ano80s] Anonymous. [Front cover].

IEEE Transactions on Professional Communication, PC-23 (4):c1, December 1980. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1980:ICF

[Ano80t] Anonymous. IEEE copyright form. IEEE Transactions on Professional Communication, PC-23(1):63–64, March 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1980:IPCa

[Ano80u] Anonymous. IEEE Professional Communication Society. IEEE Transactions on Professional Communication, PC-23 (1):1, March 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1980:IPCb

[Ano80v] Anonymous. IEEE Professional Communication Society. IEEE Transactions on Professional Communication, PC-23 (2):1, June 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1980:IPCc

[Ano80w] Anonymous. IEEE Professional Communication Society. IEEE Transactions on Professional Communication, PC-23 (3):1, September 1980. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1980:IARa

[Ano80x]  Anonymous.  Information

for authors and readers of the IEEE Transactions on Professional Communication.

IEEE Transactions on Professional Communication, PC-23(1):62, March 1980. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1980:IARb

[Ano80y]  Anonymous.  Information

for authors and readers of the IEEE Transactions on Professional Communication. IEEE Transactions on Professional Communication, PC-23(2):112, June 1980. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1980:IARc

[Ano80z]  Anonymous.  Information

for authors and readers of the IEEE Transactions on Professional Communication. IEEE Transactions on Professional Communication, PC-23 (3):155, September 1980. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1980:IARd

[Ano80-27] Anonymous. Information for authors and readers of the IEEE Transactions on Professional Communication. IEEE Transactions on Professional Communication, PC-23 (4):206, December 1980. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1980:IBC

[Ano80-28] Anonymous. [Inside back cover]. IEEE Transactions on Professional Communication, PC-23(4):c3, December 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1980:IFCa

[Ano80-29] Anonymous. [Inside front cover]. IEEE Transactions on Professional Communication, PC-23(1):c2, March 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1980:IFCb

[Ano80-30] Anonymous. [Inside front cover]. IEEE Transactions on Professional Communication, PC-23(2):c2, June 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1980:IFCc

[Ano80-31] Anonymous. [Inside front cover]. IEEE Transactions on Professional Communication, PC-23(3):c2, September 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1980:IFCd

[Ano80-32] Anonymous. [Inside front cover]. IEEE Transactions on Professional Communication, PC-23(4):c2, December 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1980:MFF

[Ano80-33] Anonymous. Microphone:

Friend or foe? IEEE Transactions on Professional Communication, PC-23(1):52– 53, March 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1981:IPCa

[Ano81a] Anonymous. 1981 IEEE Professional Communication Society Conference. IEEE Transactions on Professional Communication, PC-24(1):57, March 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1981:IIT

[Ano81b] Anonymous. 1981 index IEEE Transactions on Professional Communication vol. PC-24. IEEE Transactions on Professional Communication, PC-24 (4):1–4, December 1981. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1981:AGA

[Ano81c] Anonymous. Alfred N. Goldsmith award of the IEEE Professional Communication Society. IEEE Transactions on Professional Communication, PC-24(1):55, March 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1981:Ca

[Ano81d] Anonymous. Contents. IEEE Transactions on Professional Communication, PC-24(1): 1, March 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1981:Cb

[Ano81e] Anonymous. Contents. IEEE Transactions on Professional Communication, PC-24(2):1, June 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1981:Cc

[Ano81f] Anonymous. Contents. IEEE Transactions on Professional Communication, PC-24(3):1, September 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1981:Cd

[Ano81g] Anonymous. Contents. IEEE Transactions on Professional Communication, PC-24(4):1, December 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1981:FRF

[Ano81h] Anonymous. Forum: Readability formulas: Used or abused? IEEE Transactions on Professional Communication, PC-24 (1):43–45, March 1981. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1981:FC

[Ano81i] Anonymous. [Front cover].

IEEE Transactions on Professional Communication, PC-24 (4):c1, December 1981. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1981:HPY

[Ano81j] Anonymous. How to polish your writing. IEEE Transactions on Professional Communication, PC-24(4):179–181, December 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1981:ICF

[Ano81k] Anonymous. IEEE copyright form. IEEE Transactions on Professional Communication, PC-24(1):59–60, March 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1981:IPCc

[Ano81l] Anonymous. IEEE Professional Communication Society. IEEE Transactions on Professional Communication, PC-24 (4):1, December 1981. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1981:IPCb

[Ano81m] Anonymous. IEEE Professional Communication Society 1981 conference. IEEE Transactions on Professional Communication, PC-24(2): 109, June 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1981:IARa

[Ano81n]  Anonymous.  Information

for authors and readers of the IEEE Transactions on Professional Communication. IEEE Transactions on Professional Communication, PC-24(1):58, March 1981. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1981:IARb

[Ano81o]  Anonymous.  Information

for authors and readers of the IEEE Transactions on Professional Communication. IEEE Transactions on Professional Communication, PC-24 (2):110–112, June 1981. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1981:IARc

[Ano81p]  Anonymous.  Information

for authors and readers of the IEEE Transactions on Professional Communication. IEEE Transactions on Professional Communication, PC-24 (4):204, December 1981. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1981:IFC

[Ano81q]  Anonymous.  [Inside front

cover]. IEEE Transactions on Professional Communication, PC-24(4):c2, December 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1981:PPW

[Ano81r] Anonymous. Power of the printed word. IEEE Transactions on Professional Communication, PC-24(2):65–71, June 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1981:PW

[Ano81s] Anonymous. The practical writer. IEEE Transactions on Professional Communication, PC-24(2):63–65, June 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1981:Pa

[Ano81t] Anonymous. Preface. IEEE Transactions on Professional Communication, PC-24(1):2– 4, March 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1981:Pb

[Ano81u] Anonymous. Preface. IEEE Transactions on Professional Communication, PC-24(2):62, June 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1981:PMD

[Ano81v]  Anonymous.  The process

model of document design. IEEE Transactions on Professional Communication, PC-24 (4):176–178, December 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1982:IPCa

[Ano82a] Anonymous. 1982 IEEE Professional Communication Society Conference. IEEE Transactions on Professional Communication, PC-25(2):105, June 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1982:IPCb

[Ano82b]  Anonymous.  1982 IEEE

Professional Communication Society Conference. IEEE Transactions on Professional Communication, PC-25(3): 165, September 1982. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1982:IIT

[Ano82c] Anonymous. 1982 index IEEE Transactions on Professional Communication vol. PC-25. IEEE Transactions on Professional Communication, PC-25 (4):1–4, December 1982. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1982:CP

[Ano82d] Anonymous. Call for papers.

IEEE Transactions on Professional Communication, PC-25 (1):52, March 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1982:CPS

[Ano82e] Anonymous. Call for papers special issue: Proposal writing. IEEE Transactions on Professional Communication, PC-25 (3):166, September 1982. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1982:CYA

[Ano82f] Anonymous. Can you afford not to communicate effectively? IEEE Transactions on Professional Communication, PC-25(4):222, December 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1982:Ca

[Ano82g] Anonymous. Contents. IEEE Transactions on Professional Communication, PC-25(1): 1, March 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1982:Cb

[Ano82h] Anonymous. Contents. IEEE Transactions on Professional Communication, PC-25(2):1, June 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1982:Cc

[Ano82i] Anonymous. Contents. IEEE Transactions on Professional Communication, PC-25(3):1, September 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1982:Cd

[Ano82j] Anonymous. Contents. IEEE Transactions on Professional Communication, PC-25(4):1, December 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1982:ICF

[Ano82k] Anonymous. IEEE copyright form. IEEE Transactions on Professional Communication, PC-25(1):55–56, March 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1982:IARa

[Ano82l]  Anonymous.  Information

for authors and readers of the IEEE Transactions on Professional Communication. IEEE Transactions on Professional Communication, PC-25(2):108, June 1982. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1982:IARb

[Ano82m]  Anonymous.  Information

for authors and readers of the IEEE Transactions on Professional Communication.

IEEE Transactions on Professional Communication, PC-25 (3):168, September 1982. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1982:IARc

[Ano82n]  Anonymous.  Information

for authors and readers of the IEEE Transactions on Professional Communication. IEEE Transactions on Professional Communication, PC-25 (4):224, December 1982. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1982:PPEa

[Ano82o]  Anonymous.  Patents and

patenting for engineers and scientists. IEEE Transactions on Professional Communication, PC-25(1):54, March 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1982:PPEb

[Ano82p]  Anonymous.  Patents and

patenting for engineers and scientists. IEEE Transactions on Professional Communication, PC-25(2):107, June 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1982:PPEc

[Ano82q]  Anonymous.  Patents and

patenting for engineers and scientists. IEEE Transactions on Professional Communication, PC-25(3):167, September 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1982:PPW

[Ano82r] Anonymous. Power of the printed word. IEEE Transactions on Professional Communication, PC-25(3):157–161, September 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1982:Pa

[Ano82s] Anonymous. Preface. IEEE Transactions on Professional Communication, PC-25(2):58– 60, June 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1982:Pb

[Ano82t] Anonymous. Preface. IEEE Transactions on Professional Communication, PC-25(4): 170–171, December 1982. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1983:IPC

[Ano83a] Anonymous. 1983 IEEE Professional Communication Society Conference. IEEE Transactions on Professional Communication, PC-26(2):93, June 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1983:IIT

[Ano83b] Anonymous. 1983 index IEEE Transactions on Professional Communication vol. PC-26. IEEE Transactions on Professional Communication, PC-26 (4):1–5, December 1983. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1983:AGA

[Ano83c] Anonymous. Alfred N. Goldsmith award of the IEEE Professional Communication Society. IEEE Transactions on Professional Communication, PC-26(1):47, March 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1983:BICa

[Ano83d]  Anonymous.  [Back inside

cover]. IEEE Transactions on Professional Communication, PC-26(1):c3, March 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1983:BICb

[Ano83e]  Anonymous.  [Back inside

cover]. IEEE Transactions on Professional Communication, PC-26(2):c3, June 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1983:BICc

[Ano83f]  Anonymous.  [Back inside

cover]. IEEE Transactions on Professional Communication, PC-26(3):c3, September 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1983:BPBa

[Ano83g] Anonymous. [Blank page — back cover]. IEEE Transactions on Professional Communication, PC-26(1):c4, March 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1983:BPBb

[Ano83h] Anonymous. [Blank page — back cover]. IEEE Transactions on Professional Communication, PC-26(2):c4, June 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1983:BPBc

[Ano83i] Anonymous. [Blank page — back cover]. IEEE Transactions on Professional Communication, PC-26(3):c4, September 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1983:CPa

[Ano83j] Anonymous. Call for papers. IEEE Transactions on Professional Communication, PC-26 (1):48, March 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1983:CPb

[Ano83k] Anonymous. Call for papers.

IEEE Transactions on Professional Communication, PC-26 (1):49, March 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1983:CPS

[Ano83l] Anonymous. Call for papers:

Special issue: Developing the ability to communicate. IEEE Transactions on Professional Communication, PC-26(2):92, June 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1983:CYAa

[Ano83m] Anonymous. Cam you afford not to communicate effectively? IEEE Transactions on Professional Communication, PC-26(2):96, June 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1983:CYAb

[Ano83n] Anonymous. Can you afford not to communicate effectively? IEEE Transactions on Professional Communication, PC-26(3):155, September 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1983:Ca

[Ano83o] Anonymous. Contents. IEEE Transactions on Professional Communication, PC-26(1):

1, March 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1983:Cb

[Ano83p] Anonymous. Contents. IEEE Transactions on Professional Communication, PC-26(2):1, June 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1983:Cc

[Ano83q] Anonymous. Contents. IEEE Transactions on Professional Communication, PC-26(3):1, September 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1983:Cd

[Ano83r] Anonymous. Contents. IEEE Transactions on Professional Communication, PC-26(4):1, December 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1983:FCa

[Ano83s] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, PC-26 (1):c1, March 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1983:FCb

[Ano83t] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, PC-26 (2):c1, June 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1983:FCc

[Ano83u] Anonymous. [Front cover].

IEEE Transactions on Professional Communication, PC-26 (3):c1, September 1983. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1983:ICF

[Ano83v] Anonymous. IEEE copyright form. IEEE Transactions on Professional Communication, PC-26(1):51–52, March 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1983:IARa

[Ano83w]  Anonymous.  Information

for authors and readers of the IEEE Transactions on Professional Communication. IEEE Transactions on Professional Communication, PC-26(1):50, March 1983. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1983:IARb

[Ano83x] Anonymous. Information for authors and readers of the IEEE Transactions on Professional Communication. IEEE Transactions on Professional Communication, PC-26(2):94, June 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1983:IARc

[Ano83y]  Anonymous.  Information

for authors and readers of the IEEE Transactions on Professional Communication. IEEE Transactions on Professional Communication, PC-26 (3):156, September 1983. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1983:IFCa

[Ano83z] Anonymous. Inside front cover. IEEE Transactions on Professional Communication, PC-26 (1):c2, March 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1983:IFCb

[Ano83-27] Anonymous. Inside front cover. IEEE Transactions on Professional Communication, PC-26 (2):c2, June 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1983:IFCc

[Ano83-28] Anonymous. Inside front cover. IEEE Transactions on Professional Communication, PC-26 (3):c2, September 1983. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1983:PPEa

[Ano83-29] Anonymous. Patents and patenting for engineers and scientists. IEEE Transactions on Professional Communication, PC-26(2):95, June 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1983:PPEb

[Ano83-30] Anonymous. Patents and patenting for engineers and scientists. IEEE Transactions on Professional Communication, PC-26(3):153, September 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1983:PCP

[Ano83-31] Anonymous. PCC 84: Call for papers & meeting announcement. IEEE Transactions on Professional Communication, PC-26(4):215, December 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1983:Pa

[Ano83-32] Anonymous. Preface. IEEE Transactions on Professional Communication, PC-26(3):98, September 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1983:Pb

[Ano83-33] Anonymous. Preface. IEEE Transactions on Professional Communication, PC-26(4): 158–159, December 1983. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1984:IIT

[Ano84a] Anonymous. 1984 index IEEE Transactions on Professional Communication vol. PC-27. IEEE Transactions on Professional Communication, PC-27 (4):1–5, December 1984. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1984:AGA

[Ano84b] Anonymous. Alfred N. Goldsmith award of the IEEE Professional Communication Society. IEEE Transactions on Professional Communication, PC-27(1):50, March 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1984:BICa

[Ano84c]  Anonymous.  [Back inside

cover]. IEEE Transactions on Professional Communication, PC-27(1):c3, March 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1984:BICb

[Ano84d]  Anonymous.  [Back inside

cover]. IEEE Transactions on Professional Communication, PC-27(2):c3, June 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1984:BICc

[Ano84e]  Anonymous.  [Back inside

cover]. IEEE Transactions on Professional Communication, PC-27(3):c3, September 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1984:BICd

[Ano84f]  Anonymous.  [Back inside

cover]. IEEE Transactions on Professional Communication, PC-27(4):c3, December 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1984:BPBa

[Ano84g] Anonymous. [Blank page — back cover]. IEEE Transactions on Professional Communication, PC-27(1):c4, March 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1984:BPBb

[Ano84h] Anonymous. [Blank page — back cover]. IEEE Transactions on Professional Communication, PC-27(2):c4, June 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1984:BPBc

[Ano84i] Anonymous. [Blank page — back cover]. IEEE Transactions on Professional Communication, PC-27(3):c4, September 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1984:BPBd

[Ano84j] Anonymous. [Blank page — back cover]. IEEE Transactions on Professional Communication, PC-27(4):c4, December 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1984:BR

[Ano84k] Anonymous. Book reviews. IEEE Transactions on Professional Communication, PC-27 (1):45, March 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1984:CP

[Ano84l] Anonymous. Call for papers. IEEE Transactions on Professional Communication, PC-27 (4):235, December 1984. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1984:CYA

[Ano84m] Anonymous. Can you afford not to communicate effectively? IEEE Transactions on Professional Communication, PC-27(1):55, March 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1984:CF

[Ano84n] Anonymous. Centennial foreword. IEEE Transactions on Professional Communication, PC-27(1):2–5, March 1984.

CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1984:Ca

[Ano84o] Anonymous. Contents. IEEE Transactions on Professional Communication, PC-27(1): 1, March 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1984:Cb

[Ano84p] Anonymous. Contents. IEEE Transactions on Professional Communication, PC-27(2):1, June 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1984:Cc

[Ano84q] Anonymous. Contents. IEEE Transactions on Professional Communication, PC-27(3):1, September 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1984:Cd

[Ano84r] Anonymous. Contents. IEEE Transactions on Professional Communication, PC-27(4):1, December 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1984:EE

[Ano84s]  Anonymous.  Essentials of

English. IEEE Transactions on Professional Communication, PC-27(2):105–107, June 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1984:FCa

[Ano84t] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, PC-27 (1):c1, March 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1984:FCb

[Ano84u] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, PC-27 (2):c1, June 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1984:FCc

[Ano84v] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, PC-27 (3):c1, September 1984. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1984:FCd

[Ano84w] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, PC-27 (4):c1, December 1984. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1984:ICF

[Ano84x] Anonymous. IEEE copyright form. IEEE Transactions on Professional Communication, PC-27(1):53–54, March 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1984:IARa

[Ano84y]  Anonymous.  Information

for authors and readers of the IEEE Transactions on Professional Communication. IEEE Transactions on Professional Communication, PC-27(1):52, March 1984. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1984:IARb

[Ano84z]  Anonymous.  Information

for authors and readers of the IEEE Transactions on Professional Communication. IEEE Transactions on Professional Communication, PC-27(2):112, June 1984. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1984:IFCa

[Ano84-27] Anonymous. Inside front cover.

IEEE Transactions on Professional Communication, PC-27 (1):c2, March 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1984:IFCb

[Ano84-28] Anonymous. Inside front cover.

IEEE Transactions on Professional Communication, PC-27 (2):c2, June 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1984:IFCc

[Ano84-29] Anonymous. Inside front cover. IEEE Transactions on Professional Communication, PC-27 (3):c2, September 1984. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1984:IFCd

[Ano84-30] Anonymous. Inside front cover. IEEE Transactions on Professional Communication, PC-27 (4):c2, December 1984. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1984:PPE

[Ano84-31] Anonymous. Patents and patenting for engineers and scientists. IEEE Transactions on Professional Communication, PC-27(1):56, March 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1984:PPAa

[Ano84-32] Anonymous. PCC 84: The practical aspects of engineering communication. IEEE Transactions on Professional Communication, PC-27(1): 51, March 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1984:PPAb

[Ano84-33] Anonymous. PCC 84: The practical aspects of engineering communication. IEEE Transactions on Professional Communication, PC-27(2): 111, June 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1985:IIT

[Ano85a] Anonymous. 1985 index IEEE transactions on professional communication vol. PC-28. IEEE Transactions on Professional Communication, PC-28 (4):1–4, December 1985. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1985:CFa

[Ano85b] Anonymous. 1986–1987 Congressional fellowships. IEEE Transactions on Professional Communication, PC-28(3):46, September 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1985:CFb

[Ano85c] Anonymous. 1986–1987 Congressional fellowships. IEEE Transactions on Professional Communication, PC-28(4):52, December 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1985:AGA

[Ano85d] Anonymous. Alfred N. Goldsmith award of the IEEE Professional Communication Society. IEEE Transactions on Professional Communication, PC-28(2):47, June 1985.

CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1985:BC

[Ano85e] Anonymous. [Back cover]. IEEE Transactions on Professional Communication, PC-28 (3):c4, September 1985. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1985:BICa

[Ano85f]  Anonymous.  [Back inside

cover]. IEEE Transactions on Professional Communication, PC-28(1):c3, March 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1985:BICb

[Ano85g]  Anonymous.  [Back inside

cover]. IEEE Transactions on Professional Communication, PC-28(2):c3, June 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1985:BICc

[Ano85h]  Anonymous.  [Back inside

cover]. IEEE Transactions on Professional Communication, PC-28(3):c3, September 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1985:BICd

[Ano85i]  Anonymous.  [Back inside

cover]. IEEE Transactions on Professional Communication, PC-28(4):c3, December 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1985:BPBa

[Ano85j] Anonymous. [Blank page — back cover]. IEEE Transactions on Professional Communication, PC-28(1):c4, March 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1985:BPBb

[Ano85k] Anonymous. [Blank page — back cover]. IEEE Transactions on Professional Communication, PC-28(2):c4, June 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1985:BPBc

[Ano85l] Anonymous. [Blank page — back cover]. IEEE Transactions on Professional Communication, PC-28(4):c4, December 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1985:CP

[Ano85m] Anonymous. Call for papers.

IEEE Transactions on Professional Communication, PC-28(4):50–51, December 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1985:C

[Ano85n] Anonymous. Contents. IEEE Transactions on Professional Communication, PC-28(1): 1, March 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1985:FCa

[Ano85o] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, PC-28 (1):c1, March 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1985:FCb

[Ano85p] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, PC-28 (2):c1, June 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1985:FCc

[Ano85q] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, PC-28 (3):c1, September 1985. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1985:FCd

[Ano85r] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, PC-28 (4):c1, December 1985. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1985:ICF

[Ano85s] Anonymous. IEEE copyright form. IEEE Transactions on Professional Communication, PC-28(1):57–58, March 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1985:IAR

[Ano85t]  Anonymous.  Information

for authors and readers of the IEEE Transactions on Professional Communication. IEEE Transactions on Professional Communication, PC-28(1):56, March 1985. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1985:IFCa

[Ano85u] Anonymous. Inside front cover.

IEEE Transactions on Professional Communication, PC-28 (1):c2, March 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1985:IFCb

[Ano85v] Anonymous. Inside front cover.

IEEE Transactions on Professional Communication, PC-28 (2):c2, June 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1985:IFCc

[Ano85w] Anonymous. Inside front cover.

IEEE Transactions on Professional Communication, PC-28 (3):c2, September 1985. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1985:IFCd

[Ano85x] Anonymous. Inside front cover. IEEE Transactions on Professional Communication, PC-28 (4):c2, December 1985. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1985:IHL

[Ano85y] Anonymous. International hot line now available through IEEE PCS. IEEE Transactions on Professional Communication, PC-28(2):48, June 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1985:OIP

[Ano85z] Anonymous. Our IEEE periodicals have only one competitor as an information source in electrotechnology. IEEE Transactions on Professional Communication, PC-28(1):59– 60, March 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1985:P

[Ano85-27] Anonymous. Preface. IEEE Transactions on Professional Communication, PC-28(1): 2, March 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1986:IPC

[Ano86a] Anonymous. 1986 IEEE Professional Communication Society conference. IEEE Transactions on Professional Communication, PC-29(3):38, September 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1986:IIT

[Ano86b] Anonymous. 1986 index IEEE transactions on professional communication vol. PC-29. IEEE Transactions on Professional Communication, PC-29 (4):101–104, December 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1986:AGA

[Ano86c] Anonymous. Alfred N. Goldsmith award of the IEEE Professional Communication Society. IEEE Transactions on Professional Communication, PC-29(2):40, June 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1986:BICa

[Ano86d]  Anonymous.  [Back inside

cover]. IEEE Transactions on Professional Communication, PC-29(1):c3, March 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1986:BICb

[Ano86e]  Anonymous.  [Back inside

cover]. IEEE Transactions on Professional Communication, PC-29(2):c3, June 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1986:BICc

[Ano86f]  Anonymous.  [Back inside

cover]. IEEE Transactions on Professional Communication, PC-29(3):c3, September 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1986:BICd

[Ano86g]  Anonymous.  [Back inside

cover]. IEEE Transactions on Professional Communication, PC-29(4):c3, December 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1986:BPBa

[Ano86h] Anonymous. [Blank page — back cover]. IEEE Transactions on Professional Communication, PC-29(1):c4, March 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1986:BPBb

[Ano86i] Anonymous. [Blank page — back cover]. IEEE Transactions on Professional Communication, PC-29(2):c4, June 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1986:BPBc

[Ano86j] Anonymous. [Blank page — back cover]. IEEE Transactions on Professional Communication, PC-29(3):c4, September 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1986:BPBd

[Ano86k] Anonymous. [Blank page — back cover]. IEEE Transactions on Professional Communication, PC-29(4):c4, December 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1986:ITP

[Ano86l] Anonymous. IEEE Transactions on Professional Communication. IEEE Transactions on Professional Communication, PC-29(3):39, September 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1986:CBS

[Ano86m] Anonymous. The CommuGuide\begin{matrix}c \\ ◯ \end{matrix} booklet series. IEEE Transactions on Professional Communication, PC-29(4):100, December 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1986:CPa

[Ano86n] Anonymous. Call for papers.

IEEE Transactions on Professional Communication, PC-29 (1):48, March 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1986:CPb

[Ano86o] Anonymous. Call for papers. IEEE Transactions on Professional Communication, PC-29 (3):40, September 1986. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1986:FCa

[Ano86p] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, PC-29 (1):c1, March 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1986:FCb

[Ano86q] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, PC-29 (2):c1, June 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1986:FCc

[Ano86r] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, PC-29 (3):c1, September 1986. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1986:FCd

[Ano86s] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, PC-29 (4):c1, December 1986. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1986:ICF

[Ano86t] Anonymous. IEEE copyright form. IEEE Transactions on Professional Communication, PC-29(1):46–47, March 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1986:IAR

[Ano86u] Anonymous. Information for authors and readers of the IEEE Transactions on Professional Communication: Guidelines for submittals. IEEE Transactions on Professional Communication, PC-29(1): 45, March 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1986:IFCa

[Ano86v] Anonymous. Inside front cover.

IEEE Transactions on Professional Communication, PC-29 (1):c2, March 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1986:IFCb

[Ano86w] Anonymous. Inside front cover.

IEEE Transactions on Professional Communication, PC-29 (2):c2, June 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1986:IFCc

[Ano86x] Anonymous. Inside front cover. IEEE Transactions on Professional Communication, PC-29 (3):c2, September 1986. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1986:IFCd

[Ano86y] Anonymous. Inside front cover. IEEE Transactions on Professional Communication, PC-29 (4):c2, December 1986. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1986:RF

[Ano86z]  Anonymous.  The readers’

forum. IEEE Transactions on Professional Communication, PC-29(3):23–24, September 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1986:WIT

[Ano86-27] Anonymous. What I’m taking to the desert island. IEEE Transactions on Professional Communication, PC-29(3):2, September 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1987:IIT

[Ano87a] Anonymous. 1987 index IEEE transactions on professional communication vol. PC-30. IEEE Transactions on Professional Communication, PC-30 (4):1–4, December 1987. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1987:AGA

[Ano87b] Anonymous. Alfred N. Goldsmith award of the IEEE Professional Communication Society. IEEE Transactions on Professional Communication, PC-30(2):115, June 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1987:BICa

[Ano87c]  Anonymous.  [Back inside

cover]. IEEE Transactions on Professional Communication, PC-30(1):c3, March 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1987:BICb

[Ano87d]  Anonymous.  [Back inside

cover]. IEEE Transactions on Professional Communication, PC-30(2):c3, June 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1987:BICc

[Ano87e]  Anonymous.  [Back inside

cover]. IEEE Transactions on Professional Communication, PC-30(3):c3, September 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1987:BICd

[Ano87f]  Anonymous.  [Back inside

cover]. IEEE Transactions on Professional Communication, PC-30(4):c3, December 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1987:BPBa

[Ano87g] Anonymous. [Blank page — back cover]. IEEE Transactions on Professional Communication, PC-30(1):c4, March 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1987:BPBb

[Ano87h] Anonymous. [Blank page — back cover]. IEEE Transactions on Professional Communication, PC-30(2):c4, June 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1987:BPBc

[Ano87i] Anonymous. [Blank page — back cover]. IEEE Transactions on Professional Communication, PC-30(3):c4, September 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1987:BPBd

[Ano87j] Anonymous. [Blank page — back cover]. IEEE Transactions on Professional Communication, PC-30(4):c4, December 1987. CODEN IEPCBU.

ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1987:BBJ

[Ano87k] Anonymous. Bring back just one good idea. IEEE Transactions on Professional Communication, PC-30(1):58, March 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1987:CPa

[Ano87l] Anonymous. Call for papers.

IEEE Transactions on Professional Communication, PC-30 (1):61–62, March 1987. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1987:CPb

[Ano87m] Anonymous. Call for papers.

IEEE Transactions on Professional Communication, PC-30 (4):253, December 1987. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1987:CTW

[Ano87n] Anonymous. Concerning twoway communication. IEEE Transactions on Professional Communication, PC-30(1): 64, March 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1987:FCa

[Ano87o] Anonymous. [Front cover].

IEEE Transactions on Professional Communication, PC-30 (1):c1, March 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1987:FCb

[Ano87p] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, PC-30 (2):c1, June 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1987:FCc

[Ano87q] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, PC-30 (3):c1, September 1987. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1987:FCd

[Ano87r] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, PC-30 (4):c1, December 1987. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1987:FIC

[Ano87s]  Anonymous.  [Front inside

cover]. IEEE Transactions on Professional Communication, PC-30(1):c2, March 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1987:ICF

[Ano87t] Anonymous. IEEE copyright form. IEEE Transactions on Professional Communication, PC-30(1):59–60, March 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1987:IA

[Ano87u] Anonymous. Information for authors. IEEE Transactions on Professional Communication, PC-30(1):57, March 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1987:IFCa

[Ano87v] Anonymous. Inside front cover.

IEEE Transactions on Professional Communication, PC-30 (2):c2, June 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1987:IFCb

[Ano87w] Anonymous. Inside front cover.

IEEE Transactions on Professional Communication, PC-30 (3):c2, September 1987. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1987:IFCc

[Ano87x] Anonymous. Inside front cover.

IEEE Transactions on Professional Communication, PC-30 (4):c2, December 1987. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1987:NCC

[Ano87y]  Anonymous.  Now a con-

cise, comprehensive, and lavishly illustrated account of the unmanned space program. IEEE Transactions on Professional Communication, PC-30 (2):116, June 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1987:RFa

[Ano87z]  Anonymous.  The readers’

forum. IEEE Transactions on Professional Communication, PC-30(1):28–29, March 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1987:RFb

[Ano87-27] Anonymous. The readers’ forum. IEEE Transactions on Professional Communication, PC-30(2):84, June 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1987:WML

[Ano87-28] Anonymous. The world’s major listing of professionals in electrical/electronics/ computer science and engineering. IEEE Transactions on Professional Communication, PC-30(1):63, March 1987. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1996:AGA

[Ano96] Anonymous. Alfred N. Goldsmith Award of the IEEE Professional Communication Society. IEEE Transactions on Professional Communication, 39(2):70, June 1996. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1997:AGA

[Ano97a]  Anonymous.  1996 Alfred

N. Goldsmith Award Presented To David E. Nadziejka. IEEE Transactions on Professional Communication, 40 (2):148, June 1997. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1997:EKS

[Ano97b]  Anonymous.  1996 Emily

K. Schlesinger Award Presented To Stephanie Rosenbaum. IEEE Transactions on Professional Communication, 40(2):149, June 1997. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1997:IIT

[Ano97c]  Anonymous.  1997 index

IEEE Transactions on Professional Communication vol. 40. IEEE Transactions on Professional Communication, 40(4): 1–5, December 1997. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1997:CGW

[Ano97d] Anonymous. The Complete Guide to Writing and Producing Technical Manuals and Handbook for Preparing Engineering Documents: From Concept to Completion [Book Reviews]. IEEE Transactions on Professional Communication, 40(1):56–57, March 1997. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1999:AI

[Ano99a] Anonymous. Author index. IEEE Transactions on Professional Communication, 42(4): 1–2, December 1999. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1999:SI

[Ano99b] Anonymous. Subject index. IEEE Transactions on Professional Communication, 42(4): 2–4, December 1999. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2000:IIT

[Ano00] Anonymous. 2000 index IEEE Transactions on Professional Communication vol. 43. IEEE Transactions on Professional Communication, 43(4):422– 427, December 2000. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2001:AI

[Ano01a] Anonymous. Author index.

IEEE Transactions on Professional Communication, 44(4): 308–309, December 2001. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2001:SI

[Ano01b] Anonymous. Subject index.

IEEE Transactions on Professional Communication, 44(4): 309–312, December 2001. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2002:AI

[Ano02a] Anonymous. Author index.

IEEE Transactions on Professional Communication, 45(4): 321–322, December 2002. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2002:SI

[Ano02b] Anonymous. Subject index.

IEEE Transactions on Professional Communication, 45(4): 322–326, December 2002. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2003:AI

[Ano03a] Anonymous. Author index.

IEEE Transactions on Professional Communication, 46(4): 346–347, December 2003. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2003:SI

[Ano03b] Anonymous. Subject index. IEEE Transactions on Professional Communication, 46(4): 347–351, December 2003. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2004:I

[Ano04a]  Anonymous.  2004 index.

IEEE Transactions on Professional Communication, 47(4): 339–344, December 2004. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2004:ITPb

[Ano04b] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 47(1):c3, March 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2004:ITPd

[Ano04c] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 47(2):c3, June 2004. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2004:ITPf

[Ano04d] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 47(3):c3, September 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2004:ITPh

[Ano04e] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 47(4):c3, December 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2004:ITPa

[Ano04f] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 47(1):c2, March 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2004:ITPc

[Ano04g] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 47(2):c2, June 2004. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2004:ITPe

[Ano04h] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 47(3):c2, September 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2004:ITPg

[Ano04i] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 47(4):c2, December 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2004:CPM

[Ano04j] Anonymous. Call for papers on making connections, IPCC 2005. IEEE Transactions on Professional Communication, 47(3):226, September 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2004:YVLa

[Ano04k] Anonymous. Have you visited lately? www.ieee.org [advertisement]. IEEE Transactions on Professional Communication, 47(3):228, September 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2004:YVLb

[Ano04l] Anonymous. Have you visited lately? www.ieee.org [advertisement]. IEEE Transactions on Professional Communication, 47(4):338, December 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2004:ICF

[Ano04m] Anonymous. IEEE copyright form. IEEE Transactions on Professional Communication, 47(1):83–84, March 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2004:IPCa

[Ano04n] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 47(1):c4, March 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2004:IPCb

[Ano04o] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 47(2):c4, June 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2004:IPCc

[Ano04p] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 47(3):c4, September 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2004:IPCd

[Ano04q] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 47(4):c4, December 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2004:QCA

[Ano04r] Anonymous. Quality without compromise [advertisement]. IEEE Transactions on Professional Communication, 47 (1):82, March 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2004:SIC

[Ano04s]  Anonymous.  Special issue

on communication in technology transfer and diffusion. IEEE Transactions on Professional Communication, 47(3): 227, September 2004. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2004:TCa

[Ano04t] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 47(1): c1, March 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2004:TCb

[Ano04u] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 47 (2):c1, June 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2004:TCc

[Ano04v] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 47(3): c1, September 2004. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2004:TCd

[Ano04w] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 47(4): c1, December 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2005:I

[Ano05a]  Anonymous.  2005 index.

IEEE Transactions on Professional Communication, 48(4): 441–444, December 2005. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2005:ITPb

[Ano05b] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 48(1):c3, March 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2005:ITPd

[Ano05c] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 48(2):c3, June 2005. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2005:ITPf

[Ano05d] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 48(3):c3, September 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2005:ITPh

[Ano05e] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 48(4):c3, December 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2005:ITPa

[Ano05f] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 48(1):c2, March 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2005:ITPc

[Ano05g] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 48(2):c2, June 2005. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2005:ITPe

[Ano05h] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 48(3):c2, September 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2005:ITPg

[Ano05i] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 48(4):c2, December 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2005:ICF

[Ano05j]  Anonymous.  IEEE copy-

right form. IEEE Transactions on Professional Communication, 48(1):115–116, March 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2005:IPCa

[Ano05k] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 48(1):c4, March 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2005:IPCb

[Ano05l] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 48(2):c4, June 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2005:IPCc

[Ano05m] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 48(3):c4, September 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2005:IPCd

[Ano05n] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 48(4):c4, December 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2005:TCa

[Ano05o] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 48(1): c1, March 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2005:TCb

[Ano05p] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 48 (2):c1, June 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2005:TCc

[Ano05q] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 48(3): c1, September 2005. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2005:TCd

[Ano05r] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 48(4): c1, December 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2006:I

[Ano06a]  Anonymous.  2006 index.

IEEE Transactions on Professional Communication, 49(4): 377–380, December 2006. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2006:ITPb

[Ano06b] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 49(1):c3, March 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2006:ITPd

[Ano06c] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 49(2):c3, June 2006. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2006:ITPf

[Ano06d] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 49(3):c3, September 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2006:ITPh

[Ano06e] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 49(4):C3, December 2006.

CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2006:ITPa

[Ano06f] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 49(1):c2, March 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2006:ITPc

[Ano06g] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 49(2):c2, June 2006. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2006:ITPe

[Ano06h] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 49(3):c2, September 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2006:ITPg

[Ano06i] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 49(4):C2, December 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2006:EIIa

[Ano06j]  Anonymous.  Explore IEL

IEEE’s most comprehensive resource [advertisement]. IEEE Transactions on Professional Communication, 49(2):211, June 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2006:EIIb

[Ano06k]  Anonymous.  Explore IEL

IEEE’s most comprehensive resource [advertisement]. IEEE Transactions on Professional Communication, 49(3):295, September 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2006:ICF

[Ano06l] Anonymous. IEEE copyright form. IEEE Transactions on Professional Communication, 49(1):91–92, March 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2006:IOFa

[Ano06m] Anonymous. IEEE order form for reprints. IEEE Transactions on Professional Communication, 49(2):212, June 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2006:IOFb

[Ano06n] Anonymous. IEEE order form for reprints. IEEE Transactions on Professional Communication, 49(3):296, September 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2006:IPCa

[Ano06o] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 49(1):c4, March 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2006:IPCb

[Ano06p] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 49(2):c4, June 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2006:IPCc

[Ano06q] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 49(3):c4, September 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2006:IPCd

[Ano06r] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 49(4):C4, December 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2006:TCa

[Ano06s] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 49(1): c1, March 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2006:TCb

[Ano06t] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 49 (2):c1, June 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2006:TCc

[Ano06u] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 49(3): c1, September 2006. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2006:TCd

[Ano06v] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 49(4): C1, December 2006. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2007:IMA

[Ano07a] Anonymous. 2007 IEEE membership application. IEEE Transactions on Professional Communication, 50(3):271– 272, September 2007. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2007:IIT

[Ano07b] Anonymous. 2007 index IEEE Transactions on Professional Communication vol. 50. IEEE Transactions on Professional Communication, 50(4):387– 392, December 2007. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2007:ITPb

[Ano07c] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 50(1):C3, March 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2007:ITPd

[Ano07d] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 50(2):C3, June 2007. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2007:ITPf

[Ano07e] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 50(3):C3, September 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2007:ITPh

[Ano07f] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 50(4):C3, December 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2007:ITPa

[Ano07g] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 50(1):C2, March 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2007:ITPc

[Ano07h] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 50(2):C2, June 2007. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2007:ITPe

[Ano07i] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 50(3):C2, September 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2007:ITPg

[Ano07j] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 50(4):C2, December 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2007:MCD

[Ano07k] Anonymous. Macromedia Captivate: The Definite Guide (Huettner, B.; 2005). IEEE Transactions on Professional Communication, 50(2):182– 183, June 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2007:EII

[Ano07l]  Anonymous.  Explore IEL

IEEE’s most comprehensive resource [advertisement]. IEEE Transactions on Professional Communication, 50(3):270, September 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2007:YVL

[Ano07m] Anonymous. Have you visited lately? www.ieee.org [advertisement]. IEEE Transactions on Professional Communication, 50(4):384, December 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2007:ICF

[Ano07n] Anonymous. IEEE copyright form. IEEE Transactions on Professional Communication, 50(4):385–386, December 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2007:IIP

[Ano07o]  Anonymous.  IEEE Inter-

national Professional Communication Conference 2008 (IPCC). IEEE Transactions on Professional Communication, 50(4):383, December 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2007:IPCa

[Ano07p] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 50(1):C4, March 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2007:IPCb

[Ano07q] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 50(2):C4, June 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2007:IPCc

[Ano07r] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 50(3):C4, September 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2007:IPCd

[Ano07s] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 50(4):C4, December 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2007:OFR

[Ano07t] Anonymous. Order form for reprints. IEEE Transactions on Professional Communication, 50(2):184, June 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2007:TCa

[Ano07u] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 50(1): C1, March 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2007:TCb

[Ano07v] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 50 (2):C1, June 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2007:TCc

[Ano07w] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 50(3): C1, September 2007. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2007:TCd

[Ano07x] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 50(4): C1, December 2007. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2008:IIT

[Ano08a] Anonymous. 2008 index IEEE Transactions on Professional Communication vol. 51. IEEE Transactions on Professional Communication, 51(4):421– 424, December 2008. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2008:ITPb

[Ano08b] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 51(1):C3, March 2008. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2008:ITPd

[Ano08c] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 51(2):C3, June 2008. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2008:ITPf

[Ano08d] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 51(3):C3, September 2008. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2008:ITPh

[Ano08e] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 51(4):C3, December 2008. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2008:ITPa

[Ano08f] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 51(1):C2, March 2008. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2008:ITPc

[Ano08g] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 51(2):C2, June 2008. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2008:ITPe

[Ano08h] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 51(3):C2, September 2008. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2008:ITPg

[Ano08i] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 51(4):C2, December 2008. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2008:CPT

[Ano08j] Anonymous. Call for papers: Tutorials & teaching cases.

IEEE Transactions on Professional Communication, 51(3): 345, September 2008. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2008:EII

[Ano08k]  Anonymous.  Explore IEL

IEEE’s most comprehensive resource [advertisement]. IEEE Transactions on Professional Communication, 51(3):348, September 2008. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2008:FCa

[Ano08l] Anonymous. [Front cover].

IEEE Transactions on Professional Communication, 51(3): C1, September 2008. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2008:FCb

[Ano08m] Anonymous. [Front cover].

IEEE Transactions on Professional Communication, 51(4): C1, December 2008. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2008:ICF

[Ano08n]  Anonymous.  IEEE copy-

right form. IEEE Transactions on Professional Communication, 51(1):131–132, March 2008. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2008:IPCa

[Ano08o] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 51(1):C4, March 2008. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2008:IPCb

[Ano08p] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 51(2):C4, June 2008. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2008:IPCc

[Ano08q] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 51(3):C4, September 2008. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2008:IPCd

[Ano08r] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 51(4):C4, December 2008. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2008:ITM

[Ano08s]  Anonymous.  IEEE Tech-

nology Management Council.

IEEE Transactions on Professional Communication, 51(3): 346, September 2008. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2008:LFS

[Ano08t] Anonymous. Leading the field since 1884 [advertisement]. IEEE Transactions on Professional Communication, 51(3): 347, September 2008. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2008:REI

[Ano08u] Anonymous. Recruiting an Editor for IEEE Press Series in Professional Engineering Communication. IEEE Transactions on Professional Communication, 51(4):420, December 2008. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2008:TCa

[Ano08v] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 51(1): C1, March 2008. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2008:TCb

[Ano08w] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 51 (2):C1, June 2008. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2008:TCc

[Ano08x] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 51(3): 237, September 2008. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2008:TCd

[Ano08y] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 51(4): 349, December 2008. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2009:IIT

[Ano09a] Anonymous. 2009 index IEEE Transactions on Professional Communication vol. 52. IEEE Transactions on Professional Communication, 52(4):416– 420, December 2009. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2009:ITPb

[Ano09b] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 52(1):C3, March 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2009:ITPd

[Ano09c] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 52(2):C3, June 2009. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2009:ITPf

[Ano09d] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 52(3):C3, September 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2009:ITPh

[Ano09e] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 52(4):C3, December 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2009:ITPa

[Ano09f] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 52(1):C2, March 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2009:ITPc

[Ano09g] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 52(2):C2, June 2009. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2009:ITPe

[Ano09h] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 52(3):C2, September 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2009:ITPg

[Ano09i] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 52(4):C2, December 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2009:CPTb

[Ano09j] Anonymous. Call for papers on tutorials and teaching cases. IEEE Transactions on Professional Communication, 52(2): 224, June 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2009:CPTa

[Ano09k] Anonymous. Call for papers:

Tutorials & teaching cases. IEEE Transactions on Professional Communication, 52(1): 119, March 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2009:CPTc

[Ano09l] Anonymous. Calls for papers on tutorials and teaching cases. IEEE Transactions on Professional Communication, 52(3): 324, September 2009. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2009:EIIa

[Ano09m]  Anonymous.  Explore IEL

IEEE’s most comprehensive resource [advertisement]. IEEE Transactions on Professional Communication, 52(1):120, March 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2009:EIIb

[Ano09n]  Anonymous.  Explore IEL

IEEE’s most comprehensive resource [advertisement]. IEEE Transactions on Professional Communication, 52(2):223, June 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2009:IPCa

[Ano09o] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 52(1):C4, March 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2009:IPCb

[Ano09p] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 52(2):C4, June 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2009:IPCc

[Ano09q] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 52(3):C4, September 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2009:IPCd

[Ano09r] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 52(4):C4, December 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2009:ITA

[Ano09s] Anonymous. IEEE-TPC article abstracts. IEEE Transactions on Professional Communication, 52(3):323, September 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2009:LFS

[Ano09t]  Anonymous.  Leading the

field since 1884 [advertisement]. IEEE Transactions on Professional Communication, 52(2):222, June 2009. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2009:NIX

[Ano09u] Anonymous. New IEEE Xplore content for IEEE TPC in June 2009. IEEE Transactions on Professional Communication, 52(1):118, March 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2009:REIa

[Ano09v] Anonymous. Recruiting an Editor for IEEE Press Series in Professional Engineering Communication. IEEE Transactions on Professional Communication, 52(1):117, March 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2009:REIb

[Ano09w] Anonymous. Recruiting an Editor for IEEE Press Series in Professional Engineering Communication. IEEE Transactions on Professional Communication, 52(2):221, June 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2009:SIF

[Ano09x]  Anonymous.  Special issue

focus: Professional communication in global contexts. IEEE Transactions on Professional Communication, 52(4): 415, December 2009. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2009:TCa

[Ano09y] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 52(1): C1, March 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2009:TCb

[Ano09z] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 52 (2):C1, June 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2009:TCc

[Ano09-27] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 52(3): C1, September 2009. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2009:TCd

[Ano09-28] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 52(4):

C1, December 2009. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2010:IIT

[Ano10a] Anonymous. 2010 index IEEE Transactions on Professional Communication vol. 53. IEEE Transactions on Professional Communication, 53(4):428– 434, December 2010. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2010:ITPb

[Ano10b] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 53(1):C3, March 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2010:ITPd

[Ano10c] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 53(2):C3, June 2010. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2010:ITPf

[Ano10d] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 53(3):C3, September 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2010:ITPh

[Ano10e] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 53(4):C3, December 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2010:ITPa

[Ano10f] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 53(1):C2, March 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2010:ITPc

[Ano10g] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 53(2):C2, June 2010. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2010:ITPe

[Ano10h] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 53(3):C2, September 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2010:ITPg

[Ano10i] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 53(4):C2, December 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2010:CPTa

[Ano10j] Anonymous. Call for paperstutorials and teaching cases. IEEE Transactions on Professional Communication, 53(2): 187, June 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2010:CPTb

[Ano10k] Anonymous. Call for paperstutorials and teaching cases. IEEE Transactions on Professional Communication, 53(3): 328, September 2010. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2010:CPTc

[Ano10l] Anonymous. Call for paperstutorials and teaching cases. IEEE Transactions on Professional Communication, 53(4):

436, December 2010. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2010:IPAa

[Ano10m] Anonymous. IEEE-PC article abstracts. IEEE Transactions on Professional Communication, 53(2):188, June 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2010:IPAb

[Ano10n] Anonymous. IEEE-PC article abstracts. IEEE Transactions on Professional Communication, 53(3):327, September 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2010:IPCa

[Ano10o] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 53(1):C4, March 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2010:IPCb

[Ano10p] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 53(2):C4, June 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2010:IPCc

[Ano10q] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 53(3):C4, September 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2010:IPCd

[Ano10r] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 53(4):C4, December 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2010:ITA

[Ano10s] Anonymous. IEEE TPC article abstracts. IEEE Transactions on Professional Communication, 53(4):435, December 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2010:TCa

[Ano10t] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 53(1): C1, March 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2010:TCb

[Ano10u] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 53 (2):C1, June 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2010:TCc

[Ano10v] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 53(3): C1, September 2010. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2010:TCd

[Ano10w] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 53(4): C1, December 2010. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2011:ITPb

[Ano11a] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 54(1):C3, March 2011. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2011:ITPa

[Ano11b] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 54(1):C2, March 2011. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2011:CPT

[Ano11c] Anonymous. Call for papers — tutorials and teaching cases. IEEE Transactions on Professional Communication, 54(1): 103, March 2011. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2011:YVL

[Ano11d] Anonymous. Have you visited lately? www.ieee.org [advertisement]. IEEE Transactions on Professional Communication, 54(1):102, March 2011. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2011:IPC

[Ano11e] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 54(1):C4, March 2011. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2011:ITA

[Ano11f] Anonymous. IEEE TPC article abstracts. IEEE Transactions on Professional Communication, 54(1):104, March 2011. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2011:TC

[Ano11g] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 54(1):

C1, March 2011. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2012:IMFa

[Ano12a] Anonymous. 2012 IEEE membership form. IEEE Transactions on Professional Communication, 55(1):106–107, March 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2012:IMFb

[Ano12b] Anonymous. 2012 IEEE membership form. IEEE Transactions on Professional Communication, 55(2):203–204, June 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2012:IIT

[Ano12c] Anonymous. 2012 index IEEE Transactions on Professional Communication vol. 55. IEEE Transactions on Professional Communication, 55(4):381– 385, December 2012. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2012:BPa

[Ano12d] Anonymous. [Blank page]. IEEE Transactions on Professional Communication, 55(1): 104, March 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2012:BPb

[Ano12e] Anonymous. [Blank page].

IEEE Transactions on Professional Communication, 55(3): 292, September 2012. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2012:ITPb

[Ano12f] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 55(1):C3, March 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2012:ITPd

[Ano12g] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 55(2):C3, June 2012. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2012:ITPa

[Ano12h] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 55(1):C2, March 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2012:ITPc

[Ano12i] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 55(2):C2, June 2012. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2012:ITPe

[Ano12j] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 55(3):C2, September 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2012:ITPf

[Ano12k] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 55(4):C2, December 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2012:FCa

[Ano12l] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, 55(1): C1, March 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2012:FCb

[Ano12m] Anonymous. [Front cover].

IEEE Transactions on Professional Communication, 55 (2):C1, June 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2012:FCc

[Ano12n] Anonymous. [Front cover].

IEEE Transactions on Professional Communication, 55(3): C1, September 2012. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2012:IMDa

[Ano12o] Anonymous. IEEE member digital library [advertisement]. IEEE Transactions on Professional Communication, 55(3): 295, September 2012. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2012:IMDb

[Ano12p] Anonymous. IEEE member digital library [advertisement]. IEEE Transactions on Professional Communication, 55(4): 380, December 2012. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2012:IOA

[Ano12q]  Anonymous.  IEEE open

access publishing [advertisement]. IEEE Transactions on Professional Communication, 55(4):378, December 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2012:IPCa

[Ano12r] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 55(1):C4, March 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2012:IPCb

[Ano12s] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 55(2):C4, June 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2012:IPCc

[Ano12t] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 55(3):C3, September 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2012:IPCd

[Ano12u] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 55(3):C4, September 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2012:IPCe

[Ano12v] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 55(4):C3, December 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2012:IPCf

[Ano12w] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 55(4):C4, December 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2012:IXDa

[Ano12x] Anonymous. IEEE Xplore Digital Library [advertisement]. IEEE Transactions on Professional Communication, 55(1): 105, March 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2012:IXDb

[Ano12y] Anonymous. IEEE Xplore Digital Library [advertisement]. IEEE Transactions on Professional Communication, 55(2): 202, June 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2012:IXDc

[Ano12z] Anonymous. IEEE Xplore Digital Library [advertisement]. IEEE Transactions on Professional Communication, 55(3): 294, September 2012. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2012:IXDd

[Ano12-27] Anonymous. IEEE Xplore Digital Library [advertisement]. IEEE Transactions on Professional Communication, 55(4): 379, December 2012. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2012:OAA

[Ano12-28] Anonymous. Open access [advertisement]. IEEE Transactions on Professional Communication, 55(3):293, September 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2012:TCa

[Ano12-29] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 55 (1):1, March 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2012:TCb

[Ano12-30] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 55(2): 105, June 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2012:TCc

[Ano12-31] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 55(3): 205, September 2012. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2012:TCd

[Ano12-32] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 55(4): 293, December 2012. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2012:TCF

[Ano12-33] Anonymous. [Table of contents — front cover]. IEEE Transactions on Professional Communication, 55(4):C1, December 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2013:IMAa

[Ano13a] Anonymous. 2013 IEEE membership application. IEEE Transactions on Professional Communication, 56(1):91– 92, March 2013. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2013:IMAb

[Ano13b] Anonymous. 2013 IEEE membership application. IEEE Transactions on Professional Communication, 56(2):191– 192, June 2013. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2013:IIT

[Ano13c] Anonymous. 2013 index IEEE Transactions on Professional Communication vol. 56. IEEE Transactions on Professional Communication, 56(4):372– 376, December 2013. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2013:BPa

[Ano13d] Anonymous. [Blank page]. IEEE Transactions on Professional Communication, 56(1): 89–90, March 2013. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2013:BPb

[Ano13e] Anonymous. [Blank page]. IEEE Transactions on Professional Communication, 56(3): 269–270, September 2013. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2013:ITPa

[Ano13f] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 56(1):C2, March 2013. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2013:ITPb

[Ano13g] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 56(2):C2, June 2013. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2013:ITPc

[Ano13h] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 56(3):C2, September 2013. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2013:ITPd

[Ano13i] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 56(4):C2, December 2013. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2013:FCa

[Ano13j] Anonymous. [Front cover].

IEEE Transactions on Professional Communication, 56(1): C1, March 2013. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2013:FCb

[Ano13k] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, 56 (2):C1, June 2013. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2013:FCc

[Ano13l] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, 56(3): C1, September 2013. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2013:FCd

[Ano13m] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, 56(4): C1, December 2013. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2013:IPCa

[Ano13n] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 56(1):C3, March 2013. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2013:IPCb

[Ano13o] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 56(1):C4, March 2013. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2013:IPCc

[Ano13p] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 56(2):C3, June 2013. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2013:IPCd

[Ano13q] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 56(2):C4, June 2013. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2013:IPCe

[Ano13r] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 56(3):C3, September 2013. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2013:IPCf

[Ano13s] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 56(3):C4, September 2013. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2013:IPCg

[Ano13t] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 56(4):C3, December 2013. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2013:IPCh

[Ano13u] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 56(4):C4, December 2013. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2013:SSI

[Ano13v] Anonymous. Spring 2015 special issue on content management. IEEE Transactions on Professional Communication, 56(3):264–268, September 2013. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2013:TCa

[Ano13w] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 56 (1):1, March 2013. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2013:TCb

[Ano13x] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 56 (2):91, June 2013. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2013:TCc

[Ano13y] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 56(3): 193, September 2013. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2013:TCd

[Ano13z] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 56(4): 271, December 2013. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2014:IIT

[Ano14a] Anonymous. 2014 index IEEE Transactions on Professional Communication vol. 57. IEEE Transactions on Professional Communication, 57(4):364– 367, December 2014. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2014:BPa

[Ano14b] Anonymous. [Blank page].

IEEE Transactions on Professional Communication, 57(1): B75–B76, March 2014. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2014:BPb

[Ano14c] Anonymous. [Blank page]. IEEE Transactions on Professional Communication, 57(3): B240, September 2014. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2014:ITPa

[Ano14d] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 57(1):C2, March 2014. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2014:ITPb

[Ano14e] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 57(2):C2, June 2014. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2014:ITPc

[Ano14f] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 57(3):C2, September 2014. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2014:ITPd

[Ano14g] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 57(4):C2, December 2014. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2014:DWY

[Ano14h] Anonymous. Do what you do better with What’s New @ IEEE Xplore. IEEE Transactions on Professional Communication, 57(4):369, December 2014. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2014:FCa

[Ano14i] Anonymous. [Front cover].

IEEE Transactions on Professional Communication, 57(1): C1, March 2014. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2014:FCb

[Ano14j] Anonymous. [Front cover].

IEEE Transactions on Professional Communication, 57 (2):C1, June 2014. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2014:FCTa

[Ano14k] Anonymous. [Front cover / table of contents]. IEEE Transactions on Professional Communication, 57(3):C1, September 2014. CODEN IEPCBU.

ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2014:FCTb

[Ano14l] Anonymous. [Front cover / table of contents]. IEEE Transactions on Professional Communication, 57(4):C1, December 2014. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2014:FYI

[Ano14m] Anonymous. Fuel your imagination [advertisement]. IEEE Transactions on Professional Communication, 57(4):370, December 2014. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2014:IGHa

[Ano14n] Anonymous. IEEE Global History Network. IEEE Transactions on Professional Communication, 57(1):78, March 2014. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2014:IGHb

[Ano14o] Anonymous. IEEE Global History Network. IEEE Transactions on Professional Communication, 57(2):155, June 2014. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2014:IPCa

[Ano14p] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 57(1):C3, March 2014. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2014:IPCb

[Ano14q] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 57(1):C4, March 2014. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2014:IPCc

[Ano14r] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 57(2):C3, June 2014. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2014:IPCd

[Ano14s] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 57(2):C4, June 2014. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2014:IPCe

[Ano14t] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 57(3):C3, September 2014. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2014:IPCf

[Ano14u] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 57(3):C4, September 2014. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2014:IPCg

[Ano14v] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 57(4):C3, December 2014. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2014:IPCh

[Ano14w] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 57(4):C4, December 2014. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2014:M

[Ano14x] Anonymous. myIEEE. IEEE Transactions on Professional Communication, 57(4):368, December 2014. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2014:OAa

[Ano14y] Anonymous. Open access.

IEEE Transactions on Professional Communication, 57(1): 77, March 2014. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2014:OAb

[Ano14z] Anonymous. Open access.

IEEE Transactions on Professional Communication, 57(2): 154, June 2014. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2014:OAP

[Ano14-27] Anonymous. Open access publishing. IEEE Transactions on Professional Communication, 57(4):371, December 2014. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2014:TCa

[Ano14-28] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 57 (1):1, March 2014. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2014:TCb

[Ano14-29] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 57 (2):77, June 2014. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2014:TCc

[Ano14-30] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 57(3): 157, September 2014. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2014:TCd

[Ano14-31] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 57(4): 241, December 2014. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2014:TWAa

[Ano14-32] Anonymous. Together, we are advancing technology. IEEE Transactions on Professional Communication, 57 (1):79, March 2014. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2014:TWAb

[Ano14-33] Anonymous. Together, we are advancing technology. IEEE Transactions on Professional Communication, 57(2):156, June 2014. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2014:TWAc

[Ano14-34] Anonymous. Together, we are advancing technology. IEEE Transactions on Professional Communication, 57(4):372, December 2014. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2015:IIT

[Ano15a] Anonymous. 2015 index IEEE Transactions on Professional Communication vol. 58. IEEE Transactions on Professional Communication, 58(4):413– 416, December 2015. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2015:BPa

[Ano15b] Anonymous. [Blank page].

IEEE Transactions on Professional Communication, 58(1): B123–B124, March 2015. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2015:BPb

[Ano15c]  Anonymous.  Blank page.

IEEE Transactions on Professional Communication, 58(3): 344, September 2015. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2015:ITPb

[Ano15d] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 58(1):C2, March 2015. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2015:ITPd

[Ano15e] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 58(2):C2, June 2015. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2015:ITPe

[Ano15f] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 58(3):C2, September 2015. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2015:ITPf

[Ano15g] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 58(4):C2, December 2015. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2015:ITPa

[Ano15h] Anonymous. IEEE Transactions on Professional Communication special issue on entrepreneurship communication. IEEE Transactions on Professional Communication, 58(1):125–127, March 2015. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2015:ITPc

[Ano15i] Anonymous. IEEE Transactions on Professional Communication special issue on entrepreneurship communication. IEEE Transactions on Professional Communication, 58(2):236–238, June 2015. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2015:FCa

[Ano15j] Anonymous. [Front cover].

IEEE Transactions on Professional Communication, 58(1): C1, March 2015. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2015:FCb

[Ano15k] Anonymous. [Front cover].

IEEE Transactions on Professional Communication, 58 (2):C1, June 2015. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2015:FCc

[Ano15l] Anonymous. [Front cover].

IEEE Transactions on Professional Communication, 58(3): C1, September 2015. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2015:FCd

[Ano15m] Anonymous. [Front cover].

IEEE Transactions on Professional Communication, 58(4): C1, December 2015. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2015:FYIa

[Ano15n] Anonymous. Fuel your imagination. IEEE Transactions on Professional Communication, 58(3):347, September 2015. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2015:FYIb

[Ano15o] Anonymous. Fuel your imagination. IEEE Transactions on Professional Communication, 58(4):412, December 2015. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2015:IPCa

[Ano15p] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 58(1):C3, March 2015. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2015:IPCb

[Ano15q] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 58(1):C4, March 2015. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2015:IPCc

[Ano15r] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 58(2):C3, June 2015. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2015:IPCd

[Ano15s] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 58(2):C4, June 2015. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2015:IPCe

[Ano15t] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 58(3):C3, September 2015. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2015:IPCf

[Ano15u] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 58(3):C4, September 2015. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2015:IPCg

[Ano15v] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 58(4):C3, December 2015. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2015:IPCh

[Ano15w] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 58(4):C4, December 2015. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2015:IXDa

[Ano15x] Anonymous. IEEE Xplore Digital Library. IEEE Transactions on Professional Communication, 58(1):129, March 2015. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2015:IXDb

[Ano15y] Anonymous. IEEE Xplore Digital Library. IEEE Transactions on Professional Communication, 58(2):240, June 2015. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2015:IICa

[Ano15z] Anonymous. Introducing IEEE Collabratec. IEEE Transactions on Professional Communication, 58(3):345, September 2015. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2015:IICb

[Ano15-27] Anonymous. Introducing IEEE Collabratec. IEEE Transactions on Professional Communication, 58(4):410, December 2015. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2015:MUM

[Ano15-28] Anonymous. The madness underlying the method. IEEE Transactions on Professional Communication, 58(2):126– 134, June 2015. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2015:MGMa

[Ano15-29] Anonymous. Member geta-member (MGM) program. IEEE Transactions on Professional Communication, 58(3): 346, September 2015. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2015:MGMb

[Ano15-30] Anonymous. Member geta-member (MGM) program. IEEE Transactions on Professional Communication, 58(4): 411, December 2015. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2015:OAa

[Ano15-31] Anonymous. Open access.

IEEE Transactions on Professional Communication, 58(1):

128, March 2015. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2015:OAb

[Ano15-32] Anonymous. Open access. IEEE Transactions on Professional Communication, 58(2): 239, June 2015. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2015:TCa

[Ano15-33] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 58 (1):1, March 2015. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2015:TCb

[Ano15-34] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 58(2): 125, June 2015. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2015:TCc

[Ano15-35] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 58(3): 241, September 2015. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2015:TCd

[Ano15-36] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 58(4): 345, December 2015. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2016:IIT

[Ano16a] Anonymous. 2016 index IEEE Transactions on Professional Communication vol. 59. IEEE Transactions on Professional Communication, 59(4):431– 435, December 2016. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2016:BPa

[Ano16b] Anonymous. [Blank page].

IEEE Transactions on Professional Communication, 59(1): B68, March 2016. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2016:BPb

[Ano16c] Anonymous. Blank page. IEEE Transactions on Professional Communication, 59(3):B313– B314, September 2016. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2016:FCa

[Ano16d] Anonymous. [Front cover].

IEEE Transactions on Professional Communication, 59(1): C1, March 2016. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2016:FCb

[Ano16e] Anonymous. [Front cover].

IEEE Transactions on Professional Communication, 59 (2):C1, June 2016. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2016:FCc

[Ano16f] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, 59(3): C1, September 2016. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2016:FCd

[Ano16g] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, 59(4): C1, December 2016. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2016:IPCa

[Ano16h] Anonymous. IEEE Professional Communication Society. IEEE Transactions on Professional Communication, 59(1): C2, March 2016. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2016:IPCc

[Ano16i] Anonymous. IEEE Professional Communication Society. IEEE Transactions on Professional Communication, 59(1): C4, March 2016. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2016:IPCd

[Ano16j] Anonymous. IEEE Professional Communication Society.

IEEE Transactions on Professional Communication, 59 (2):C2, June 2016. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2016:IPCf

[Ano16k] Anonymous. IEEE Professional Communication Society. IEEE Transactions on Professional Communication, 59 (2):C4, June 2016. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2016:IPCg

[Ano16l] Anonymous. IEEE Professional Communication Society. IEEE Transactions on Professional Communication, 59(3): C2, September 2016. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2016:IPCj

[Ano16m] Anonymous. IEEE Professional Communication Society. IEEE Transactions on Professional Communication, 59(4): C2, December 2016. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2016:IPCb

[Ano16n] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 59(1):C3, March 2016. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2016:IPCe

[Ano16o] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 59(2):C3, June 2016. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2016:IPCh

[Ano16p] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 59(3):C3, September 2016. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2016:IPCi

[Ano16q] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 59(3):C4, September 2016. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2016:IPCk

[Ano16r] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 59(4):C3, December 2016. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2016:IPCl

[Ano16s] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 59(4):C4, December 2016. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2016:IICa

[Ano16t] Anonymous. Introducing IEEE Collabratec. IEEE Transactions on Professional Communication, 59(3):315, September 2016. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2016:IICb

[Ano16u] Anonymous. Introducing IEEE Collabratec. IEEE Transactions on Professional Communication, 59(4):436, December 2016. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2016:TCa

[Ano16v] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 59 (1):1, March 2016. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2016:TCb

[Ano16w] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 59 (2):69, June 2016. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2016:TCc

[Ano16x] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 59(3): 165, September 2016. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2016:TCd

[Ano16y] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 59(4): 345, December 2016. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2017:IIT

[Ano17a]  Anonymous.  2017 index

IEEE Transactions on Professional Communication vol. 60. IEEE Transactions on Professional Communication, 60(4): 1–5, December 2017. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2017:BPa

[Ano17b] Anonymous. Blank page. IEEE Transactions on Professional Communication, 60(1):B115– B116, March 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2017:BPb

[Ano17c] Anonymous. [Blank page]. IEEE Transactions on Professional Communication, 60(3):

B334, September 2017. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2017:DMW

[Ano17d] Anonymous. A diamond milestone for a winning combination. IEEE Transactions on Professional Communication, 60(1):2–3, March 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2017:EYNa

[Ano17e] Anonymous. Expand your network. IEEE Transactions on Professional Communication, 60(3):335, September 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2017:EYNb

[Ano17f] Anonymous. Expand your network. IEEE Transactions on Professional Communication, 60(3):336, September 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2017:EYNc

[Ano17g] Anonymous. Expand your network. IEEE Transactions on Professional Communication, 60(4):458, December 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2017:EYNd

[Ano17h] Anonymous. Expand your network. IEEE Transactions on Professional Communication, 60(4):459, December 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2017:FCa

[Ano17i] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, 60(1): C1, March 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2017:FCb

[Ano17j] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, 60 (2):C1, June 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2017:FCc

[Ano17k] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, 60(3): C1, September 2017. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2017:FCd

[Ano17l] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, 60(4): C1, December 2017. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2017:IPCa

[Ano17m] Anonymous. IEEE Professional Communication Society. IEEE Transactions on Professional Communication, 60(1): C2, March 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2017:IPCd

[Ano17n] Anonymous. IEEE Professional Communication Society. IEEE Transactions on Professional Communication, 60 (2):C2, June 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2017:IPCg

[Ano17o] Anonymous. IEEE Professional Communication Society. IEEE Transactions on Professional Communication, 60(3): C2, September 2017. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2017:IPCj

[Ano17p] Anonymous. IEEE Professional Communication Society. IEEE Transactions on Professional Communication, 60(4): C2, December 2017. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2017:IPCb

[Ano17q] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 60(1):C3, March 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2017:IPCc

[Ano17r] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 60(1):C4, March 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2017:IPCe

[Ano17s] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 60(2):C3, June 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2017:IPCf

[Ano17t] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 60(2):C4, June 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2017:IPCh

[Ano17u] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 60(3):C3, September 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2017:IPCi

[Ano17v] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 60(3):C4, September 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2017:IPCk

[Ano17w] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 60(4):C3, December 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2017:IPCl

[Ano17x] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 60(4):C4, December 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2017:IICa

[Ano17y] Anonymous. Introducing IEEE Collabratec. IEEE Transactions on Professional Communication, 60(1):117, March 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2017:IICb

[Ano17z] Anonymous. Introducing IEEE Collabratec. IEEE Transactions on Professional Communication, 60(2):236, June 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2017:IICc

[Ano17-27] Anonymous. Introducing IEEE Collabratec. IEEE Transactions on Professional Communication, 60(3):337, September 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2017:IICd

[Ano17-28] Anonymous. Introducing IEEE Collabratec. IEEE Transactions on Professional Communication, 60(4):460, December 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2017:TCa

[Ano17-29] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 60 (1):1, March 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2017:TCb

[Ano17-30] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 60(2): 117, June 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2017:TCc

[Ano17-31] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 60(3): 237, September 2017. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2017:TCd

[Ano17-32] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 60(4): 335, December 2017. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2018:IIT

[Ano18a]  Anonymous.  2018 index

IEEE Transactions on Professional Communication vol. 61. IEEE Transactions on Professional Communication, 61(4): 1–6, December 2018. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2018:BPa

[Ano18b]  Anonymous.  Blank page.

IEEE Transactions on Professional Communication, 61(1): B114, March 2018. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2018:BPb

[Ano18c]  Anonymous.  Blank page.

IEEE Transactions on Professional Communication, 61(3):

B350, September 2018. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2018:EYNa

[Ano18d] Anonymous. Expand your network. IEEE Transactions on Professional Communication, 61(3):352, September 2018. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2018:EYNb

[Ano18e] Anonymous. Expand your network. IEEE Transactions on Professional Communication, 61(4):444, December 2018. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2018:FCa

[Ano18f] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, 61(1): C1, March 2018. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2018:FCb

[Ano18g] Anonymous. [Front cover]. IEEE Transactions on Professional Communication, 61 (2):C1, June 2018. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2018:FCc

[Ano18h]  Anonymous.  Front cover.

IEEE Transactions on Professional Communication, 61(3):

C1, September 2018. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2018:FCd

[Ano18i]  Anonymous.  Front cover.

IEEE Transactions on Professional Communication, 61(4): C1, December 2018. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2018:IPCa

[Ano18j] Anonymous. IEEE Professional Communication Society. IEEE Transactions on Professional Communication, 61(1): C2, March 2018. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2018:IPCd

[Ano18k] Anonymous. IEEE Professional Communication Society. IEEE Transactions on Professional Communication, 61 (2):C2, June 2018. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2018:IPCg

[Ano18l] Anonymous. IEEE Professional Communication Society. IEEE Transactions on Professional Communication, 61(3): C2, September 2018. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2018:IPCj

[Ano18m] Anonymous. IEEE Professional Communication Society. IEEE Transactions on Professional Communication, 61(4): C2, December 2018. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2018:IPCb

[Ano18n] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 61(1):C3, March 2018. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2018:IPCc

[Ano18o] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 61(1):C4, March 2018. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2018:IPCe

[Ano18p] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 61(2):C3, June 2018. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2018:IPCf

[Ano18q] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 61(2):C4, June 2018. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2018:IPCh

[Ano18r] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 61(3):C3, September 2018. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2018:IPCi

[Ano18s] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 61(3):C4, September 2018. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2018:IPCk

[Ano18t] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 61(4):C3, December 2018. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2018:IPCl

[Ano18u] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 61(4):C4, December 2018. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2018:IICa

[Ano18v] Anonymous. Introducing IEEE Collabratec. IEEE Transactions on Professional Communication, 61(3):351, September 2018. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2018:IICb

[Ano18w] Anonymous. Introducing IEEE Collabratec. IEEE Transactions on Professional Communication, 61(4):443, December 2018. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2018:TCa

[Ano18x] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 61 (1):1, March 2018. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2018:TCb

[Ano18y] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 61(2): 115, June 2018. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2018:TCc

[Ano18z] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 61(3):

225, September 2018. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2018:TCd

[Ano18-27] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 61(4): 351, December 2018. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2019:IPCa

[Ano19a] Anonymous. IEEE Professional Communication Society. IEEE Transactions on Professional Communication, 62(1): C2, March 2019. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2019:IPCc

[Ano19b] Anonymous. IEEE Professional Communication Society. IEEE Transactions on Professional Communication, 62 (2):C2, June 2019. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2019:IPCe

[Ano19c] Anonymous. IEEE Professional Communication Society. IEEE Transactions on Professional Communication, 62(3): C2, September 2019. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2019:IPCb

[Ano19d] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 62(1):C4, March 2019. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2019:IPCd

[Ano19e] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 62(2):C4, June 2019. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2019:IPCf

[Ano19f] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 62(3):C3, September 2019. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2019:IAa

[Ano19g] Anonymous. Information for authors. IEEE Transactions on Professional Communication, 62(1):114, March 2019. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2019:IAb

[Ano19h] Anonymous. Information for authors. IEEE Transactions on Professional Communication, 62(2):210, June 2019. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2019:IAc

[Ano19i] Anonymous. Information for authors. IEEE Transactions on Professional Communication, 62(3):314, September 2019. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2019:RG

[Ano19j] Anonymous. Reaching out globally. IEEE Transactions on Professional Communication, 62(1):1–3, March 2019. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2019:TCa

[Ano19k] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 62(1): C1, March 2019. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2019:TCb

[Ano19l] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 62 (2):C1, June 2019. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2019:TCc

[Ano19m] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 62(3): C1, September 2019. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2020:IIT

[Ano20a]  Anonymous.  2020 index

IEEE Transactions on Professional Communication vol. 63. IEEE Transactions on Professional Communication, 63(4): 1–6, December 2020. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2020:ETR

[Ano20b]  Anonymous.  An Editor’s

thanks for reviewer excellence. IEEE Transactions on Professional Communication, 63(1): 1–2, March 2020. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2020:IPCa

[Ano20c] Anonymous. IEEE Professional Communication Society. IEEE Transactions on Professional Communication, 63(1): C2, March 2020. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2020:IPCc

[Ano20d] Anonymous. IEEE Professional Communication Society. IEEE Transactions on Professional Communication, 63 (2):C2, June 2020. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2020:IPCe

[Ano20e] Anonymous. IEEE Professional Communication Society. IEEE Transactions on Professional Communication, 63(3): C2, September 2020. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2020:IPCg

[Ano20f] Anonymous. IEEE Professional Communication Society. IEEE Transactions on Professional Communication, 63(4): C2, December 2020. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2020:IPCh

[Ano20g] Anonymous. IEEE Professional Communication Society. IEEE Transactions on Professional Communication, 63(4): C4, December 2020. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2020:IPCb

[Ano20h] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 63(1):C4, March 2020. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2020:IPCd

[Ano20i] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 63(2):C4, June 2020. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2020:IPCf

[Ano20j] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 63(3):C4, September 2020. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2020:IAa

[Ano20k] Anonymous. Information for authors. IEEE Transactions on Professional Communication, 63(1):C3, March 2020. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2020:IAb

[Ano20l] Anonymous. Information for authors. IEEE Transactions on Professional Communication, 63(2):C3, June 2020. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2020:IAc

[Ano20m] Anonymous. Information for authors. IEEE Transactions on Professional Communication, 63(3):C3, September 2020. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2020:IAd

[Ano20n] Anonymous. Information for authors. IEEE Transactions on Professional Communication, 63(4):C3, December 2020. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2020:TCa

[Ano20o] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 63(1): C1, March 2020. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2020:TCb

[Ano20p] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 63 (2):C1, June 2020. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2020:TCc

[Ano20q] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 63(3): C1, September 2020. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2020:TCd

[Ano20r] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 63(4): C1, December 2020. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2021:IPCa

[Ano21a] Anonymous. IEEE Professional Communication Society. IEEE Transactions on Professional Communication, 64(1): C2, March 2021. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2021:IPCb

[Ano21b] Anonymous. IEEE Professional Communication Society. IEEE Transactions on Professional Communication, 64(1): C4, March 2021. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2021:IPCc

[Ano21c] Anonymous. IEEE Professional Communication Society. IEEE Transactions on Professional Communication, 64 (2):C2, June 2021. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2021:IPCd

[Ano21d] Anonymous. IEEE Professional Communication Society. IEEE Transactions on Professional Communication, 64 (2):C4, June 2021. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2021:IPCe

[Ano21e] Anonymous. IEEE Professional Communication Society. IEEE Transactions on Professional Communication, 64(3):

C2, September 2021. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2021:IPCf

[Ano21f] Anonymous. IEEE Professional Communication Society. IEEE Transactions on Professional Communication, 64(3): C4, September 2021. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2021:IPCg

[Ano21g] Anonymous. IEEE Professional Communication Society. IEEE Transactions on Professional Communication, 64(4): C2, December 2021. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2021:IPCh

[Ano21h] Anonymous. IEEE Professional Communication Society. IEEE Transactions on Professional Communication, 64(4): C4, December 2021. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2021:IAa

[Ano21i] Anonymous. Information for authors. IEEE Transactions on Professional Communication, 64(1):C3, March 2021. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2021:IAb

[Ano21j] Anonymous. Information for authors. IEEE Transactions on Professional Communication, 64(2):C3, June 2021. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2021:IAc

[Ano21k] Anonymous. Information for authors. IEEE Transactions on Professional Communication, 64(3):C3, September 2021. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2021:IAd

[Ano21l] Anonymous. Information for authors. IEEE Transactions on Professional Communication, 64(4):C3, December 2021. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2021:TCa

[Ano21m] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 64(1): C1, March 2021. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2021:TCb

[Ano21n] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 64 (2):C1, June 2021. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2021:TCc

[Ano21o] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 64(3): C1, September 2021. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2021:TCd

[Ano21p] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 64(4): C1, December 2021. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2022:IIT

[Ano22a]  Anonymous.  2022 index

IEEE Transactions on Professional Communication vol. 65. IEEE Transactions on Professional Communication, 65(4): 1–7, December 2022. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2022:ITPa

[Ano22b] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 65(3):C3, September 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2022:IPCa

[Ano22c] Anonymous. IEEE professional communication society.

IEEE Transactions on Professional Communication, 65(1): C2, March 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2022:IPCb

[Ano22d] Anonymous. IEEE Professional Communication Society. IEEE Transactions on Professional Communication, 65(1): C4, March 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2022:IPCc

[Ano22e] Anonymous. IEEE Professional Communication Society. IEEE Transactions on Professional Communication, 65 (2):C2, June 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2022:IPCd

[Ano22f] Anonymous. IEEE Professional Communication Society. IEEE Transactions on Professional Communication, 65 (2):C4, June 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2022:IPCe

[Ano22g] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 65(3):C2, September 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2022:IPCf

[Ano22h] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 65(3):C4, September 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2022:IPCg

[Ano22i] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 65(4):C2, December 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2022:IPCh

[Ano22j] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 65(4):C4, December 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2022:ITPb

[Ano22k] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 65(4):C3, December 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2022:IAa

[Ano22l] Anonymous. Information for authors. IEEE Transactions on Professional Communication, 65(1):C3, March 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2022:IAb

[Ano22m] Anonymous. Information for authors. IEEE Transactions on Professional Communication, 65(2):C3, June 2022. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2022:TCa

[Ano22n] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 65(1): C1, March 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2022:TCb

[Ano22o] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 65 (2):C1, June 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2022:TCc

[Ano22p] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 65(3): C1, September 2022. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2022:TCd

[Ano22q] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 65(4): C1, December 2022. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2023:IIT

[Ano23a]  Anonymous.  2023 index

IEEE Transactions on Professional Communication vol. 66. IEEE Transactions on Professional Communication, 66(4): 1–6, December 2023. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2023:ITPa

[Ano23b] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 66(1):C3, March 2023. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2023:ITPc

[Ano23c] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 66(2):C3, June 2023. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2023:ITPd

[Ano23d] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 66(3):C3, September 2023. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2023:ITPe

[Ano23e] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 66(4):C3, December 2023. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2023:ITPb

[Ano23f] Anonymous. IEEE Transactions on Professional Communication publication information. IEEE Transactions on Professional Communication, 66(2):C2, June 2023. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2023:IPCa

[Ano23g] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 66(1):C2, March 2023. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2023:IPCb

[Ano23h] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 66(1):C4, March 2023. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2023:IPCc

[Ano23i] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 66(2):C4, June 2023. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2023:IPCd

[Ano23j] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 66(3):C2, September 2023. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2023:IPCe

[Ano23k] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 66(3):C4, September 2023. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2023:IPCf

[Ano23l] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 66(4):C2, December 2023. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2023:IPCg

[Ano23m] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 66(4):C4, December 2023. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2023:TCa

[Ano23n] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 66(1): C1, March 2023. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2023:TCb

[Ano23o] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 66 (2):C1, June 2023. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2023:TCc

[Ano23p] Anonymous. Table of contents. IEEE Transactions on Professional Communication, 66(3): C1, September 2023. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2023:TCd

[Ano23q] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 66(4): C1, December 2023. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2024:ITPa

[Ano24a] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 67(1):C3, March 2024. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2024:ITPb

[Ano24b] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 67(2):C3, June 2024. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2024:ITPc

[Ano24c] Anonymous. IEEE Transactions on Professional Communication information for authors. IEEE Transactions on Professional Communication, 67(3):C3, September 2024. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2024:IPCa

[Ano24d] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 67(1):C2, March 2024. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2024:IPCb

[Ano24e] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 67(1):C4, March 2024. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2024:IPCc

[Ano24f] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 67(2):C2, June 2024. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2024:IPCd

[Ano24g] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 67(2):C4, June 2024. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2024:IPCe

[Ano24h] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 67(3):C2, September 2024. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2024:IPCf

[Ano24i] Anonymous. IEEE Professional Communication Society information. IEEE Transactions on Professional Communication, 67(3):C4, September 2024. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2024:TCa

[Ano24j] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 67(1): C1, March 2024. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2024:TCb

[Ano24k] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 67 (2):C1, June 2024. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:2024:TCc

[Ano24l] Anonymous. Table of contents.

IEEE Transactions on Professional Communication, 67(3): C1, September 2024. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anthony:1979:HGA

[Ant79] James P. Anthony. How to get an article published in a technical magazine. IEEE Transactions on Professional Communication, PC-22(4):190–193, December 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anthony:1999:WRA

[Ant99] L. Anthony. Writing research article introductions in software engineering: how accurate is a standard model? IEEE Transactions on Professional Communication, 42(1): 38–46, March 1999. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anthony:2001:CFR

[Ant01]  L. Anthony.  Characteris-

tic features of research article titles in computer science. IEEE Transactions on Professional Communication, 44(3): 187–194, September 2001. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anthony:2006:DFM

[Ant06] L. Anthony. Developing a freeware, multiplatform corpus analysis toolkit for the technical writing classroom. IEEE Transactions on Professional Communication, 49(3): 275–286, September 2006. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anthony:2018:IFF

[Ant18]  Laurence Anthony.  Intro-

ducing FireAnt: a freeware, multiplatform social media data-analysis tool. IEEE Transactions on Professional Communication, 61(4):428– 442, December 2018. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Applen:2001:DCO

[App01a] J. D. Applen. Disease classification and the organization of large-scale web sites. IEEE Transactions on Professional Communication, 44(4): 286–290, December 2001. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Applen:2001:EML

[App01b] J. D. Applen. Extensible markup languages and traditional abstracting and indexing strategies. IEEE Transactions on Professional Communication, 44(3):202–206, September 2001. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Applen:2002:TKK

[App02] J. D. Applen. Tacit knowledge, knowledge management, and active user participation in Web site navigation. IEEE Transactions on Professional Communication, 45(4):302– 306, December 2002. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Arader:1959:FDS

[Ara59] Harry F. Arader. The function and design of the scientific message. IRE Transactions on Engineering Writing and Speech, EWS-2(1):2–5, January 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Armstrong:1982:CO

[Arm82] J. Scott Armstrong. Creative obfuscation. IEEE Transactions on Professional Communication, PC-25(1):30– 32, March 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Arms:1983:ELW

[Arm83] Valarie M. Arms. Engineers like to write on a computer! IEEE Transactions on Professional Communication, PC-26 (4):175–177, December 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Arms:1985:Pa

[Arm85a] Valarie M. Arms. Preface. IEEE Transactions on Professional Communication, PC-28 (2):1, June 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Arms:1985:Pb

[Arm85b] Valarie M. Arms. Preface. IEEE Transactions on Professional Communication, PC-28 (4):1, December 1985. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Arms:1986:P

[Arm86] Valarie M. Arms. Preface.

IEEE Transactions on Professional Communication, PC-29 (1):1, March 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Arnold:1961:WA

[Arn61] Christian K. Arnold. The writing of abstracts. IRE Transactions on Engineering Writing and Speech, EWS-4(3):80– 82, December 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Arnold:1962:CST

[Arn62] Christian K. Arnold. The construction of statistical tables. IRE Transactions on Engineering Writing and Speech, EWS-5(1):9–14, August 1962. ISSN 0893-7869 (print), 2331-3692 (electronic).

Arnold:1969:HST

[Arn69] Judith F. Arnold. A Handbook for the Scientific and Technical Secretary — George Freedman. IEEE Transactions on Engineering Writing and Speech, EWS-12(3):91, October 1969. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Arnold:1984:VD

[Arn84] William R. Arnold. Vehicles for documentation. IEEE Transactions on Professional Communication, PC-27(4):222–229, December 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Arrott:1974:TP

[Arr74] William Arrott. Technical publicity. IEEE Transactions on Professional Communication, PC-17(2):36–38, June 1974. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ada:2016:FII

[ASHB16] Serkan Ada, Raj Sharman, Wencui Han, and Joseph A. Brennan. Factors impacting the intention to use emergency notification services in campus emergencies: An empirical investigation. IEEE Transactions on Professional Communication, 59(2):89–109, June 2016. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Askew:1973:OYC

[Ask73] William G. Askew. The offer you can’t refuse the engineering periodical. IEEE Transactions on Professional Communication, PC-16(3):97– 99, September 1973. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ackerman:1988:RRO

[AT88] S. S. Ackerman and W. W.

Turechek. The risks and rewards of online editing. IEEE Transactions on Professional Communication, 31(3):122– 123, September 1988. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Atherton:2022:TDJ

[Ath22]  Rachel Atherton.  Toward

data justice: Understanding police shooting data systems and narratives. IEEE Transactions on Professional Communication, 65(1):118–134, March 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Atlas:1981:UEM

[Atl81] Marshall A. Atlas. The user edit: Making manuals easier to use. IEEE Transactions on Professional Communication, PC-24(1):28–29, March 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Auger:1978:VAC

[Aug78] B. Y. Auger. Visual aids to clarity. IEEE Transactions on Professional Communication, PC-21(2):71–76, June 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Auger:1980:SMP

[Aug80]  B. Y. Auger.  The spe-

cialist makes a presentation. IEEE Transactions on Professional Communication, PC-23 (3):122–125, September 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Allen:1998:EEA

[AV98] L. Allen and D. Voss. Ethics for editors: an analytical decision-making process. IEEE Transactions on Professional Communication, 41 (1):58–65, March 1998. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Avedon:1975:TM

[Ave75] Don M. Avedon. The technology of micrographics. IEEE Transactions on Professional Communication, PC-18(3): 154–159, September 1975. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Abshire:1978:TSP

[AW78] Gary M. Abshire and Kenneth L. White. Techniques and suggestions for presenters. IEEE Transactions on Professional Communication, PC-21 (3):97–100, September 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ai:2017:TBC

[AW17] Bin Ai and Lifei Wang.

Transnational business communication and identity work in Australia. IEEE Transactions on Professional Communication, 60(2):201–213, June 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Alexander:2022:RRU

[AW22] Jamal-Jared Alexander and Rebecca Walton. Relational recruiting: Using black feminist theory to inform graduate recruiting strategies. IEEE Transactions on Professional Communication, 65(1):164– 178, March 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Axelby:1973:PPE

[Axe73] George S. Axelby. Problems and procedures in editing a professional technical journal. IEEE Transactions on Professional Communication, PC-16(3):54–56, September 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Alaiad:2017:PAW

[AZ17] Ahmad Alaiad and Lina Zhou.

Patients’ adoption of WSNbased smart home healthcare systems: an integrated model of facilitators and barriers. IEEE Transactions on Professional Communication, 60(1):

4–23, March 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anandarajan:2010:GAI

[AZDA10] Murugan Anandarajan, Maliha Zaman, Qizhi Dai, and Bay Arinze. Generation Y adoption of instant messaging: an examination of the impact of social usefulness and media richness on use richness. IEEE Transactions on Professional Communication, 53(2):132–143, June 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Blicq:1983:BRT

[BA83] R. S. Blicq and Gerald J. Alred. Book reviews: Technicallywrite! IEEE Transactions on Professional Communication, PC-26(1):42, March 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Batova:2016:ISI

[BA16] Tatiana Batova and Rebekka Andersen. Introduction to the special issue: Content strategy a unifying vision. IEEE Transactions on Professional Communication, 59(1):2–6, March 2016. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Baard:1994:PFP

[Baa94] P. P. Baard. Psychological fusion and personal conflict in organizations. IEEE Transactions on Professional Communication, 37(1):14–17, March 1994. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Baake:2007:DMQ

[Baa07] Ken Baake. Decision-making in a quasi-rational world: Teaching technical, narratological, and rhetorical discourse in report writing tutorial. IEEE Transactions on Professional Communication, 50(2):163– 171, June 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Baehr:2010:AVC

[BAB10] Craig Baehr and Konstanze Alex-Brown. Assessing the value of corporate blogs: a social capital perspective. IEEE Transactions on Professional Communication, 53(4):358– 369, December 2010. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Baehr:2012:IUA

[Bae12]  Craig Baehr.  Incorporat-

ing user appropriation, media richness, and collaborative knowledge sharing into blended e-learning training tutorial. IEEE Transactions on Professional Communication, 55(2):175–184, June 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Baggett:1983:FPD

[Bag83] Patricia Baggett. Four principles for designing instructions. IEEE Transactions on Professional Communication, PC-26 (3):99–105, September 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bahill:1985:WRS

[Bah85] A. Terry Bahill. Write for the reading system; talk to the listening system. IEEE Transactions on Professional Communication, PC-28(2):44–45, June 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Baird:1981:GMS

[Bai81a] John E. Baird. Getting the most out of seminars: Listening by objectives. IEEE Transactions on Professional Communication, PC-24(4):189–191, December 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Baird:1981:HOE

[Bai81b] John E. Baird. How to overcome errors in public speaking. IEEE Transactions on Professional Communication, PC-24 (2):94–98, June 1981. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bailie:2024:EWC

[Bai24] Rahel Anne Bailie. Engineering Words: Communicating Clearly in the Workplace: Sharon Burton and Bonni Graham Gonzalez: [book review]. IEEE Transactions on Professional Communication, 67(1): 152–153, March 2024. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Baker:1975:MSI

[Bak75] Houston Baker. Managing subscription income. IEEE Transactions on Professional Communication, PC-18(3):93– 96, September 1975. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Baker:1982:GGH

[Bak82] H. Kent Baker. Guidelines for gatherings help you meet with success. IEEE Transactions on Professional Communication, PC-25(3):152–154, September 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Balaban:1975:CAJ

[Bal75] Miriam Balaban. Cooperation among journals: a European vantage point. IEEE Transactions on Professional Communication, PC-18(3):116–120, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ballentine:2008:PCW

[Bal08] Brian D. Ballentine. Professional communication and a ’whole new mind’: Engaging with ethics, intellectual property, design, and globalization. IEEE Transactions on Professional Communication, 51(3): 328–340, September 2008. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Balagia:2018:TDI

[Bal18] Michelle Balagia. Tragic Design: the Impact of Bad Product Design and How to Fix It (Sharlat, J. and Saucier, S.) [Book Review]. IEEE Transactions on Professional Communication, 61(2):223–224, June 2018. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bamford:1973:EPC

[Bam73] Harold E. Bamford. The editorial processing center. IEEE Transactions on Professional Communication, PC-16(3):82– 83, September 1973. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Barry:1977:CSS

[Bar77] James D. Barry. Communication system for space. IEEE Transactions on Professional Communication, PC-20 (4):206, December 1977. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bartlett:1979:HEC

[Bar79] S. C. Bartlett. How engineers can help educate the public. IEEE Transactions on Professional Communication, PC-22 (3):171–173, September 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Barry:1980:CRL

[Bar80] Jeanne G. Barry. Computerized readability levels. IEEE Transactions on Professional Communication, PC-23(2):88– 90, June 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Barry:1981:CRL

[Bar81] Jeanne G. Barry. Computerized readability levels. IEEE Transactions on Professional Communication, PC-24(1):45– 46, March 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Barnum:1982:TTW

[Bar82a] Carol M. Barnum. Teaching technical writing to the engineering student: Industry’s needs, the students’ expectations. IEEE Transactions on Professional Communication, PC-25(3):136–139, September 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bartlett:1982:HEC

[Bar82b] S. C. Bartlett. How engineers can help educate the public.

IEEE Transactions on Professional Communication, PC-25 (4):199–201, December 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Barakat:1989:SCH

[Bar89] R. A. Barakat. Storyboarding can help your proposal. IEEE Transactions on Professional Communication, 32(1): 20–25, March 1989. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Barker:1990:SDI

[Bar90] T. T. Barker. Software documentation: from instruction to integration. IEEE Transactions on Professional Communication, 33(4):172– 177, December 1990. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Barakat:1991:DWP

[Bar91] R. A. Barakat. Developing winning proposal strategies. IEEE Transactions on Professional Communication, 34(3): 130–139, September 1991. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Baranowski:1993:ETP

[Bar93] M. Baranowski. European television: politics and broadcasting in the European Community. IEEE Transactions on Professional Communication, 36(1):35–47, March 1993. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Barchilon:1998:JGP

[Bar98a] M. G. Barchilon. James G.

Paradis and Muriel L. Zimmerman, The MIT Guide To Science and Engineering Communication. IEEE Transactions on Professional Communication, 41(2):154–155, June 1998. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Barnett:1998:TDL

[Bar98b] M. Barnett. Testing a digital library of technical manuals. IEEE Transactions on Professional Communication, 41(2): 116–122, June 1998. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Barclay:2009:AIH

[Bar09a] Lizabeth A. Barclay. Acing the Interview: How to Ask and Answer the Questions that Will Get You the Job! (Beshara, T.; 2008) [Book Review]. IEEE Transactions on Professional Communication, 52(1):113–114, March 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Barclay:2009:GIH

[Bar09b] Lizabeth A. Barclay. Globalizing International Human Resource Management (Rowley, C. and Warner, M., Eds.; 2008) [Book Review]. IEEE Transactions on Professional Communication, 52(2):217–218, June 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Barclay:2009:TPP

[Bar09c] Lizabeth A. Barclay. Tagging: People-Powered Metadata for the Social Web (Smith, G.; 2008) [Book Review]. IEEE Transactions on Professional Communication, 52(3):321– 322, September 2009. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Barchilon:2014:OEE

[Bar14] Marian Barchilon. Online Education 2.0: Evolving, Adapting, and Reinventing Online Technical Education [Book Review]. IEEE Transactions on Professional Communication, 57(2): 147–149, June 2014. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Barr:2022:YWL

[Bar22] Nancy Barr. So, You Have to Write a Literature Review: a Guided Workbook for Engineers: Catherine G. P. Berdanier and Joshua B.

Lenart: [Book Review]. IEEE Transactions on Professional Communication, 65(4):530– 531, December 2022. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Basu:1972:BEC

[Bas72] R. N. Basu. Barriers to effective communication in the scientific world: a PSYCHOM ’72 paper. IEEE Transactions on Professional Communication, PC-15(2):30–33, June 1972. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Basso:1997:HPR

[Bas97] J. Basso. How public relations professionals are managing the potential for sabotage, rumors, and misinformation disseminated via the internet by computer hackers. IEEE Transactions on Professional Communication, 40(1): 28–33, March 1997. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bates:1980:IT

[Bat80] Marcia J. Bates. Idea tactics.

IEEE Transactions on Professional Communication, PC-23 (2):95–100, June 1980. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bates:1984:UWC

[Bat84] Gary D. Bates. Upgrading written communication: Your firm’s and your own. IEEE Transactions on Professional Communication, PC-27(2):89– 92, June 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Battalio:2002:MSH

[Bat02] J. T. Battalio. A methodology for streamlining historical research: the analysis of technical and scientific publications. IEEE Transactions on Professional Communication, 45(1):21–39, March 2002. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Batova:2014:CCM

[Bat14]  Tatiana Batova.  Compo-

nent content management and quality of information products for global audiences: an integrative literature review. IEEE Transactions on Professional Communication, 57(4): 325–339, December 2014. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Batova:2018:GTC

[Bat18a] Tatiana Batova. Global technical communication in 7.5 weeks online: Combining industry and academic perspectives. IEEE Transactions on Professional Communication, 61(3):311–329, September 2018. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Batova:2018:NMQ

[Bat18b] Tatiana Batova. Negotiating multilingual quality in component content-management environments. IEEE Transactions on Professional Communication, 61(1):77–100, March 2018. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Baum:1962:LEU

[Bau62] Harry Baum. Letter to the editor [on using summaries to judge technical article submissions]. IRE Transactions on Engineering Writing and Speech, EWS-5(1):33, August 1962. ISSN 0893-7869 (print), 2331-3692 (electronic).

Bay:2022:FDE

[Bay22] Jennifer Bay. Fostering diversity, equity, and inclusion in the technical and professional communication service course. IEEE Transactions on Professional Communication, 65(1): 213–225, March 2022. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bazerman:1999:ICR

[Baz99] C. Bazerman. Introduction:

Changing regularities of genre [commentary]. IEEE Transactions on Professional Communication, 42(1):1–2, March 1999. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Baruch:1975:IAE

[BB75] Jordan J. Baruch and Nazir Bhagat. The IEEE Annals: an experiment in selective dissemination. IEEE Transactions on Professional Communication, PC-18(3):296–308, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Brockmann:1986:WBC

[BB86] R. John Brockmann and Annette Norris Bradford. Writing better computer user documentation. IEEE Transactions on Professional Communication, PC-29(3):37, September 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bostian:1991:TID

[BB91] C. W. Bostian and F. F. Bostian. Technical information and the 1989 debate about high definition TV. IEEE Transactions on Professional Communication, 34(4):206– 210, December 1991. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Burn:1999:CAV

[BB99] J. Burn and M. Barnett. Communicating for advantage in the virtual organization. IEEE Transactions on Professional Communication, 42(4):215– 222, December 1999. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Basara:1986:CSO

[BBRT86] Diane Basara, Darrell Burgin, Gabrielle Ryan, and Paul Trummel. A case study of online information: Second generation systems design. IEEE Transactions on Professional Communication, PC-29(4):81– 86, December 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Biscardi:1970:WCS

[BC70] Louis E. Biscardi and Winston N. Coburn. Writing can still be the better way. IEEE Transactions on Engineering Writing and Speech, EWS-13 (1):13–17, May 1970. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Bodnar:2017:CGB

[BC17] Cheryl A. Bodnar and Renee M. Clark. Can gamebased learning enhance engineering communication skills? IEEE Transactions on Professional Communication, 60(1): 24–41, March 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Beard:1984:MDC

[BCB84] Richard E. Beard, Peter V. Callamaras, and R. John Brockmann. A method for designing computer support documentation. IEEE Transactions on Professional Communication, PC-27(2):108–109, June 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

BHATTACHARYYA:2003:SBO

[BCB03] G. Bhattacharyya, S. Calimsiz, and G. M. Bodner. Strange bedfellows: organic synthesis and essay-writing. IEEE Transactions on Professional Communication, 46(4): 320–326, December 2003. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Biles:1987:UES

[BCJR87] Al Biles, Frank Cost, Guy Johnson, and Ken Reek. Using expert systems in typographic design. IEEE Transactions on Professional Communication, PC-30(2):102–111, June 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Battle:2000:SPP

[BCR00] M. V. Battle, G. S. M. Cowan, and E. A. Rakow. Statistical power: planning for it, estimating it post hoc, and reporting it. IEEE Transactions on Professional Communication, 43 (4):345–354, December 2000. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Booher:1982:EMA

[BD82] E. Kathleen Booher and Wilma Davidson. Eleven myths about writing and how trainers can debunk them. IEEE Transactions on Professional Communication, PC-25(3):133–135, September 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bagby:1983:HUG

[BD83] Susan A. Bagby and Sherry A.

Dell. How to use graduate interns to support your publications department. IEEE Transactions on Professional Communication, PC-26(4):184–186, December 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Buchholz:1984:CTC

[BD84] William J. Buchholz and L. W.

Denton. Communication training and consulting in business, industry, and government. IEEE Transactions on Professional Communication, PC-27(1):47–48, March 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Beatts:1959:UYR

[Bea59] P. M. Beatts. Use your reader’s eyes. IRE Transactions on Engineering Writing and Speech, EWS-2(1):6–11, January 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Beauvais:1998:SCR

[Bea98] P. J. Beauvais. Strategies of containment in recall notices. IEEE Transactions on Professional Communication, 41(3): 205–208, September 1998. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Beason:1999:JLC

[Bea99] Larry Beason. Jean A. Lutz and C. Gilbert Storms, The practice of technical and scientific communication: writing in professional contexts. IEEE Transactions on Professional Communication, 42(3): 190–192, September 1999. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bechard:1968:DCA

[Bec68]  Henry L. Bechard.  Digi-

tally coded alphanumeric photocomposition system. IEEE Transactions on Engineering Writing and Speech, EWS-11 (2):23–30, August 1968. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Beck:1983:PWW

[Bec83] Clark E. Beck. Proposals: Write to win. IEEE Transactions on Professional Communication, PC-26(2):56–60, June 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Beck:1984:BRT

[Bec84] Clark E. Beck. Book reviews: Tim Whalen, Preparing contract-winning proposals and feasibility studies. IEEE Transactions on Professional Communication, PC-27 (4):230, December 1984. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Beck:1985:CEW

[Bec85] Charles E. Beck. Conducting an editing workshop. IEEE Transactions on Professional Communication, PC-28(1):43– 45, March 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bedi:2007:EIT

[Bed07] Melanie Doulton Bedi. The evolution of Indian technical communication teams commentary. IEEE Transactions on Professional Communication, 50(2):88–90, June 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bednar:2012:TCU

[Bed12] Lucy Bednar. Teaching case using a research in technical and scientific communication class to teach essential workplace skills. IEEE Transactions on Professional Communication, 55(4):363–377, December 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Belanger:1999:CPD

[Bel99] F. Belanger. Communication patterns in distributed work groups: a network analysis. IEEE Transactions on Professional Communication, 42(4): 261–275, December 1999. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bennett:1961:BWN

[Ben61] John B. Bennett. Better writing: New answers for an old problem. IRE Transactions on Engineering Writing and Speech, EWS-4(2):44–46, May 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Benjamin:1979:ITM

[Ben79] Flissy Benjamin. Impress top management with your written reports. IEEE Transactions on Professional Communication, PC-22(4):202–203, December 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Benrey:1985:TMC

[Ben85]  Ronald M. Benrey.  Top-

down management communication: The view from midchannel. IEEE Transactions on Professional Communication, PC-28(3):17–20, September 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bennett:2022:PAS

[Ben22] Kristin C. Bennett. Prioritizing access as a social justice concern: Advocating for ableism studies and disability justice in technical and professional communication. IEEE Transactions on Professional Communication, 65(1): 226–240, March 2022. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Berns:1968:FP

[Ber68] Gerald M. Berns. The FOR-MAT program. IEEE Transactions on Engineering Writing and Speech, EWS-11(2):85– 91, August 1968. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Bernstein:1977:P

[Ber77] Theodore M. Bernstein. Punctuation. IEEE Transactions on Professional Communication, PC-20(1):38–44, June 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Berry:1982:HGU

[Ber82]  Elizabeth Berry.  How to

get users to follow procedures. IEEE Transactions on Professional Communication, PC-25 (1):22–25, March 1982. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bernhardt:1992:DSC

[Ber92] S. A. Bernhardt. The design of sexism: the case of an army maintenance manual. IEEE Transactions on Professional Communication, 35(4): 217–221, December 1992. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Berrett:2012:WSE

[Ber12] Jolynne Berrett. Writing in the Sciences: Exploring Conventions of Scientific Discourse (Part of the Allyn & Bacon Series in Technical Communication), 3/E (Penrose, A. M. and Katz, S. B.; 2010) [Book Review]. IEEE Transactions on Professional Communication, 55(1):97–99, March 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Baniya:2022:RST

[BESFW22] Sweta Baniya, Jessica Edwards, Jennifer Sano-Franchini, and Josephine Walwema. Revisiting SL in TPC through social justice and intercultural frameworks: Findings from survey research. IEEE Transactions on Professional Communication, 65(3):353– 369, September 2022. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Benbunan-Fich:2005:WPT

[BFA05] R. Benbunan-Fich and S. Altschuller.

Web presence transformations in the 1990s: an analysis of press releases. IEEE Transactions on Professional Communication, 48(2):131–146, June 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

BENBUNAN-FICH:2003:MEO

[BFH03] R. Benbunan-Fich and S. R.

Hiltz. Mediators of the effectiveness of online courses. IEEE Transactions on Professional Communication, 46(4): 298–312, December 2003. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Brammer:2007:ITP

[BG07] Charlotte Brammer and Ryan Galloway. IEEE Transactions on Professional Communication: Looking to the past to discover the present. IEEE Transactions on Professional Communication, 50(4): 275–279, December 2007. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Belinsky:2016:TCP

[BG16] Stacy J. Belinsky and Brian Gogan. Throwing a changeup, pitching a strike: an autoethnography of frame acquisition, application, and fit in a pitch development and delivery experience. IEEE Transactions on Professional Communication, 59(4):323–341, December 2016. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Boiarsky:1995:WTS

[BGN^{+}95] C. Boiarsky, L. Grove, B. Northrop, M. Phillips, F. Myers, and P. Earnest. Women in technical/scientific professions: results of two national surveys. IEEE Transactions on Professional Communication, 38 (2):68–76, June 1995. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bailie:2015:ECT

[BH15] Rahel Anne Bailie and Jeffrey Huset. The effect of CMS technology on writing styles and processes: Two case studies. IEEE Transactions on Professional Communication, 58(3): 309–327, September 2015. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Brewer:2016:BCB

[BH16] Edward C. Brewer and Terence L. Holmes. Better communication = better teams: a communication exercise to improve team performance. IEEE Transactions on Professional Communication, 59(3): 288–298, September 2016. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bennett:2021:GFI

[BH21] Kristin C. Bennett and Mark A.

Hannah. Generative fusions: Integrating technical and professional communication, disability studies, and legal studies in the work of disability inclusion and access. IEEE Transactions on Professional Communication, 64(3):235– 249, September 2021. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bhagat:1977:IDS

[Bha77] Nazir A. Bhagat. Information dissemination a systems viewpoint. IEEE Transactions on Professional Communication, PC-20(2):76–79, September 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bhatia:2012:CSR

[Bha12] Aditi Bhatia. The corporate social responsibility report: The hybridization of a confused genre (2007–2011). IEEE Transactions on Professional Communication, 55(3): 221–238, September 2012. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Boettger:2018:ISI

[BI18] Ryan K. Boettger and Suguru Ishizaki. Introduction to the special issue: Datadriven approaches to research and teaching in professional and technical communication. IEEE Transactions on Professional Communication, 61(4): 352–355, December 2018. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Barth:2021:PRU

[BIdJ^{+}21] Susanne Barth, Dan Ionita, Menno D. T. de Jong, Pieter H. Hartel, and Marianne Junger. Privacy rating: a user-centered approach for visualizing data handling practices of online services. IEEE Transactions on Professional Communication, 64(4):354–373, December 2021. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Billings:1973:PP

[Bil73] Millard F. Billings. Prose pollution. IEEE Transactions on Professional Communication, PC-16(1):11–14, March 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bishop:1970:MPT

[Bis70] Anthony C. L. Bishop. The multimedia presentation of technical information. IEEE Transactions on Engineering Writing and Speech, EWS-13 (1):24–27, May 1970. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Bist:1990:TNT

[Bis90] G. Bist. Training new technical writers. IEEE Transactions on Professional Communication, 33(3):130–132, September 1990. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bist:1994:SSM

[Bis94] G. Bist. Single-source manuals.

IEEE Transactions on Professional Communication, 37(2): 81–87, June 1994. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bist:1996:AOO

[Bis96] G. Bist. Applying the objectoriented model to technical information. IEEE Transactions on Professional Communication, 39(1):49–57, March 1996. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bishop:1999:STC

[Bis99] K. Bishop. Scenarios for technical communication: critical thinking and writing. IEEE Transactions on Professional Communication, 42(3):198– 199, September 1999. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Billingsley:1979:NUL

[BJ79] Patricia A. Billingsley and Neil A. Johnson. Nonsexist use of language in scientific and technical writing. IEEE Transactions on Professional Communication, PC-22 (4):193–198, December 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Brand:1981:Pb

[BJ81] Jeff Brand and R. J. Joenk. Preface. IEEE Transactions on Professional Communication, PC-24(4):154–155, December 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Barstow:1986:IOD

[BJ86] Thomas R. Barstow and Joseph T. Jaynes. Integrating online documentation into the technical publishing process. IEEE Transactions on Professional Communication, PC-29(4):37–41, December 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Barr:2021:TTC

[BJ21a] Nancy B. Barr and Jaclyn E. Johnson. Trajectories in turmoil: a case study of engineering students reactions to disruptions in their community of practice. IEEE Transactions on Professional Communication, 64(1):38–51, March 2021.

CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ahmad:2021:LBC

[BJ21b] Kamarul Zaman Bin Ahmad and Sajjad M. Jasimuddin. The linkage between communication satisfaction, human resources management practices, person–organization fit, and commitment: Evidence from Malaysia. IEEE Transactions on Professional Communication, 64(1):23–37, March 2021. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Buswell:2019:EIW

[BJT^{+}19] Natascha Trellinger Buswell, Brent K. Jesiek, Cary D. Troy, Rebecca R. Essig, and Joshua Boyd. Engineering instructors on writing: Perceptions, practices, and needs. IEEE Transactions on Professional Communication, 62(1): 55–74, March 2019. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Burkey:2003:WBS

[BK03] J. Burkey and W. L. Kuechler. Web-based surveys for corporate information gathering: a bias-reducing design framework. IEEE Transactions on Professional Communication, 46(2):81–93, June 2003. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Braby:1982:IFT

[BKSM82] Richard Braby, J. Peter Kincaid, Paul Scott, and William C. McDaniel. Illustrated formats to teach procedures. IEEE Transactions on Professional Communication, PC-25(2):61–66, June 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Boettger:2013:OEQ

[BL13] Ryan K. Boettger and Chris Lam. An overview of experimental and quasi-experimental research in technical communication journals (1992–2011). IEEE Transactions on Professional Communication, 56(4): 272–293, December 2013. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Blicq:1961:AQC

[Bli61] Ronald S. Blicq. Apply quality control to engineering writing. IRE Transactions on Engineering Writing and Speech, EWS-4(2):47–51, May 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Blicq:1972:CRC

[Bli72] Ronald S. Blicq. Changing the role of the communications professional: a PSYCHOM ’72 paper. IEEE Transactions on Professional Communication, PC-15(3):86–87, September 1972. CODEN IEPCBU.

ISSN 0361-1434 (print), 1558-1500 (electronic).

Blicq:1976:EGS

[Bli76] Ron Blicq. Editing guidelines for supervisory engineers. IEEE Transactions on Professional Communication, PC-19 (2):33–37, June 1976. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Blicq:1984:JHS

[Bli84] Ron S. Blicq. Job hunting: Sharpening your competitive edge. IEEE Transactions on Professional Communication, PC-27(4):201–210, December 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Blicq:1987:CTE

[Bli87] Ron S. Blicq. Communication techniques for engineers, scientists, and computer specialists. IEEE Transactions on Professional Communication, PC-30 (4):216–217, December 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bloomwell:1983:FST

[Blo83]  A. E. Bloomwell.  Forty-

six tips for flip-chart users. IEEE Transactions on Professional Communication, PC-26 (3):106–107, September 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bloch:1987:HEC

[Blo87] Carolyn Bloch. How engineers can find and use federal scientific and technical data. IEEE Transactions on Professional Communication, PC-30 (1):19–27, March 1987. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bernhardt:2000:KMP

[BM00] S. A. Bernhardt and G. A. McCulley. Knowledge management and pharmaceutical development teams: using writing to guide science. IEEE Transactions on Professional Communication, 43 (1):22–34, March 2000. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Berdanier:2021:CDC

[BMF21] Catherine G. P. Berdanier, Mary McCall, and Gracemarie Mike Fillenwarth. Characterizing disciplinarity and conventions in engineering resume profiles. IEEE Transactions on Professional Communication, 64(4):390–406, December 2021. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bradner:2005:TST

[BMH05] E. Bradner, G. Mark, and T. D. Hertel. Team size and technology fit: participation, awareness,and rapport in distributed teams. IEEE Transactions on Professional Communication, 48(1):68–77, March 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bragge:2005:GIE

[BMRH05] J. Bragge, H. Merisalo-Rantanen, and P. Hallikainen. Gathering innovative end-user feedback for continuous development of information systems: a repeatable and transferable e-collaboration process. IEEE Transactions on Professional Communication, 48 (1):55–67, March 2005. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Brewer:2015:TLC

[BMS^{+}15] Pam Estes Brewer, Alanah Mitchell, Robert Sanders, Paul Wallace, and David D. Wood. Teaching and learning in crossdisciplinary virtual teams. IEEE Transactions on Professional Communication, 58 (2):208–229, June 2015. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Baskerville:2007:APV

[BN07] Richard Baskerville and Joe Nandhakumar. Activating and perpetuating virtual teams: Now that we’re mobile, where do we go? IEEE Transactions on Professional Communication, 50(1):17–34, March 2007.

CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bjorn:2010:TAN

[BN10] Pernille Bjørn and Ojelanki Ngwenyama. Technology alignment: a new area in virtual team research. IEEE Transactions on Professional Communication, 53(4):382– 400, December 2010. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Baker:2021:EPB

[BNB21] Matthew J. Baker, E. M. Nightingale, and Suzy Bills. An editing process for blind or visually impaired editors. IEEE Transactions on Professional Communication, 64(3): 275–287, September 2021. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bocchi:1998:TET

[Boc98] J. Bocchi. Technical editing in transition: editors wanted for intranet site development. IEEE Transactions on Professional Communication, 41(1): 5–15, March 1998. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Boehm:1976:SHC

[Boe76] Peter J. Boehm. Software and hardware considerations for a technical typesetting system.

IEEE Transactions on Professional Communication, PC-19 (1):15–19, March 1976. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Boettger:2010:RUT

[Boe10] Ryan K. Boettger. Rubric use in technical communication: Exploring the process of creating valid and reliable assessment tools. IEEE Transactions on Professional Communication, 53(1):4–17, March 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Boiarsky:2018:CBG

[Boi18] Carolyn Boiarsky. Communication between government agencies and local communities: Rhetorical analyses of primary documents in three environmental risk situations. IEEE Transactions on Professional Communication, 61(1): 2–21, March 2018. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bolmer:1982:TTP

[Bol82] Joan Bolmer. Tips on talking in public. IEEE Transactions on Professional Communication, PC-25(1):40–42, March 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bone:1983:CFE

[Bon83] John R. Bone. The care and feeding of the executive speaker. IEEE Transactions on Professional Communication, PC-26(1):20–24, March 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Book:2017:AIR

[Boo17] Cassandra Book. Academy Industry Relationships and Partnerships: Perspectives for Technical Communicators [by Bridgeford, T. and St. Amant, K.; Book review]. IEEE Transactions on Professional Communication, 60(2):232– 233, June 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

BCIEWSG:1964:CIS

[Bos64] Boston Chapter IEEE Engineering Writing and Speech Group. Critique of the IEEE Spectrum and other IEEE publications — a panel discussion. IEEE Transactions on Engineering Writing and Speech, EWS-7(2):21–26, September 1964. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Bostian:1984:TWV

[Bos84] Frieda F. Bostian. Technical writing very useful stuff. IEEE Transactions on Professional Communication, PC-27 (3):121–125, September 1984.

CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Boschen:1987:ICG

[Bos87a] Allan C. Boschen. International communication: Getting down to basics. IEEE Transactions on Professional Communication, PC-30(2):74–76, June 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Boston:1987:CKE

[Bos87b] Bruce O. Boston. Cadence:

The key to effective speech writing. IEEE Transactions on Professional Communication, PC-30(4):227–228, December 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bosley:1992:GVC

[Bos92] D. S. Bosley. Gender and visual communication: toward a feminist theory of design. IEEE Transactions on Professional Communication, 35(4): 222–229, December 1992. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bottle:1970:ICT

[Bot70] Robert T. Bottle. The information content of titles in engineering literature. IEEE Transactions on Engineering Writing and Speech, EWS-13(2):41–45, September 1970.

CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Boushey:1958:WYW

[Bou58] Homer A. Boushey. When you write for the Air Force. IRE Transactions on Engineering Writing and Speech, EWS-1(1):28–30, March 1958. ISSN 0893-7869 (print), 2331-3692 (electronic).

Bourne:1962:PPE

[Bou62] Charles P. Bourne. Problems posed by an expanding technical literature. IRE Transactions on Engineering Writing and Speech, EWS-5(1):2–8, August 1962. ISSN 0893-7869 (print), 2331-3692 (electronic).

Bovee:1973:STJ

[Bov73]  Warren G. Bovee.  Scien-

tific and technical journals on microfiche. IEEE Transactions on Professional Communication, PC-16(3):113–116, September 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bowen:1973:RTM

[Bow73] D. H. Michael Bowen. The role of technical magazine/journal hybrids. IEEE Transactions on Professional Communication, PC-16(3):66–67, September 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Boyle:1968:ATE

[Boy68] John J. Boyle. Automated typesetting of existing computer tabular data. IEEE Transactions on Engineering Writing and Speech, EWS-11 (2):129–133, August 1968. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Boyle:1969:CTO

[Boy69] John J. Boyle. Computerized typesetting — an overview. IEEE Transactions on Engineering Writing and Speech, EWS-12(2):28–29, May 1969. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Boyd:1981:BIF

[Boy81] David P. Boyd. Business interviews frequent questions, acceptable answers. IEEE Transactions on Professional Communication, PC-24(4):165–168, December 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Boyell:1987:ERF

[Boy87] Roger L. Boyell. The emerging role of the forensic engineer. IEEE Transactions on Professional Communication, PC-30 (1):16–18, March 1987. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bailey:1982:GBI

[BP82] M. C. Bailey and W. Pferd. Graphics-based instruction for interactive-graphics-system training. IEEE Transactions on Professional Communication, PC-25(2):67–73, June 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Berndt:2009:CBC

[BP09] Annette Berndt and Carla Paterson. Complementing business case studies with humanitarian case studies: a means of preparing global engineers. IEEE Transactions on Professional Communication, 52(4): 398–410, December 2009. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Boettger:2010:QCA

[BP10] Ryan K. Boettger and Laura A. Palmer. Quantitative content analysis: Its use in technical communication. IEEE Transactions on Professional Communication, 53(4):346– 357, December 2010. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bremner:2015:LEA

[BP15] Stephen Bremner and Ban Phung. Learning from the experts: an analysis of résumé writers’ self-presentation on LinkedIn. IEEE Transactions on Professional Communication, 58(4):367–380, December 2015. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Barclay:1994:TCP

[BPK94] R. O. Barclay, T. E. Pinelli, and J. M. Kennedy. Technical communication practices of Dutch and U.S. aerospace engineers and scientists: international perspectives on aerospace. IEEE Transactions on Professional Communication, 37(2):97–107, June 1994. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Barnett:1968:NLP

[BR68] M. P. Barnett and W. M. Ruhsam. A natural language programming system for text processing. IEEE Transactions on Engineering Writing and Speech, EWS-11(2):45–52, August 1968. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Boyle:1992:SCH

[BR92] C. Boyle and K. Ratliff. A survey and classification of hypertext documentation systems. IEEE Transactions on Professional Communication, 35(2): 98–111, June 1992. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Battle:1993:ZAR

[BR93] M. V. Battle and E. A. Rakow. Zen and the art of reporting differences in data that are not statistically significant. IEEE Transactions on Professional Communication, 36(2):75–80, June 1993. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Boren:2000:TAR

[BR00] T. Boren and J. Ramey. Thinking aloud: reconciling theory and practice. IEEE Transactions on Professional Communication, 43(3):261– 278, September 2000. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Balzotti:2016:CBP

[BR16] Jonathan Balzotti and Jacob D. Rawlins. Client-based pedagogy meets workplace simulation: Developing social processes in the Arisoph Case Study. IEEE Transactions on Professional Communication, 59(2):140–152, June 2016. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bragg:1959:TWA

[Bra59] Sir Lawrence Bragg. Talking and writing about science. IRE Transactions on Engineering Writing and Speech, EWS-2(3): 69–72, December 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Bram:1978:SCS

[Bra78] V. A. Bram. Sentence construction in scientific and engineering texts. IEEE Transactions on Professional Communication, PC-21(4):162–164, December 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Brand:1981:Pa

[Bra81] Jeff Brand. Preface. IEEE Transactions on Professional Communication, PC-24(3): 114–115, September 1981. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Brand:1982:CPS

[Bra82a] Jeff Brand. Call for papers special issue: Proposal writing. IEEE Transactions on Professional Communication, PC-25 (2):106, June 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Brand:1982:P

[Bra82b] Jeff Brand. Preface. IEEE Transactions on Professional Communication, PC-25(1):2– 4, March 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Braunstein:1982:BTJ

[Bra82c] Leslie A. Braunstein. Beyond the technical journals. IEEE Transactions on Professional Communication, PC-25 (4):192–195, December 1982.

CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Brand:1983:P

[Bra83] Jeff Brand. Preface. IEEE Transactions on Professional Communication, PC-26(2):54, June 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bradford:1984:PMD

[Bra84] David Bradford. The persona in microcomputer documentation. IEEE Transactions on Professional Communication, PC-27(2):65–68, June 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Brasseur:1993:COP

[Bra93] L. E. Brasseur. Contesting the objectivist paradigm: gender issues in the technical and professional communication curriculum. IEEE Transactions on Professional Communication, 36(3):114–123, September 1993. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Brammer:1999:JAW

[Bra99a] C. Brammer. Jo Allen, writing in the workplace. IEEE Transactions on Professional Communication, 42(3):193– 195, September 1999. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Brammer:1999:PDF

[Bra99b] C. Brammer. Profile of David F. Beer. IEEE Transactions on Professional Communication, 42(4):306–308, December 1999. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Brammer:1999:PTH

[Bra99c] C. Brammer. Profile of Thom Haller. IEEE Transactions on Professional Communication, 42(3):181–184, September 1999. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Brammer:2000:PJR

[Bra00] C. Brammer. Profile of Janet Rochester. IEEE Transactions on Professional Communication, 43(2):215–216, June 2000. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Brammer:2001:QGI

[Bra01] C. Brammer. Quick Guide To The Internet For Composition, 2000 Edition [Book Review]. IEEE Transactions on Professional Communication, 44(1): 65–66, March 2001. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Brammer:2003:AGG

[Bra03a]  C. Brammer.  Alan G.

Gross and Ray D. Dearin, Chaim Perelman [Book Review]. IEEE Transactions on Professional Communication, 46(4):340–341, December 2003. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Brammer:2003:PCC

[Bra03b] C. Brammer. The power of corporate communication: Crafting the voice and image of your business [Book Review]. IEEE Transactions on Professional Communication, 46(2): 149–151, June 2003. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Brammer:2004:CPE

[Bra04] C. Brammer. Communication patterns of engineers. IEEE Transactions on Professional Communication, 47(3): 216–217, September 2004. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Brammer:2005:TWB

[Bra05] C. Brammer. Technical writing basics: a guide to style & form, 3rd ed. IEEE Transactions on Professional Communication, 48(2):232–233, June 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Brearley:1973:RTR

[Bre73] Neil Brearley. The role of technical reports in scientific and technical communication.

IEEE Transactions on Professional Communication, PC-16 (3):117–119, September 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Brewer:2010:MIV

[Bre10] Pamela Estes Brewer. Miscommunication in international virtual workplaces: a report on a multicase study. IEEE Transactions on Professional Communication, 53(4):329– 345, December 2010. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bridgwood:1999:GCE

[Bri99] M. A. Bridgwood. Guidelines for communication and engineering problem solving at the basic level. IEEE Transactions on Professional Communication, 42(3):156–165, September 1999. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Barfield:1991:UIE

[BRL91] W. Barfield, C. Rosenberg, and G. Levasseur. The use of icons, earcons, and commands in the design of an online hierarchical menu. IEEE Transactions on Professional Communication, 34(2):101–108, June 1991. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Brookes:1962:TWI

[Bro62] B. C. Brookes. Technical writing in industry: Technique and morale. IRE Transactions on Engineering Writing and Speech, EWS-5(1):15–20, August 1962. ISSN 0893-7869 (print), 2331-3692 (electronic).

Broner:1964:SH

[Bro64] Maurice A. Broner. Stand up and be heard. IEEE Transactions on Engineering Writing and Speech, EWS-7(3):25–30, December 1964. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Brogan:1965:LBF

[Bro65] John A. Brogan. Lessons from Benjamin Franklin, America’s first great technical writer. IEEE Transactions on Engineering Writing and Speech, EWS-8(1):3–7, June 1965. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Broer:1971:ABD

[Bro71] J. W. Broer. Abstracts in block diagram form. IEEE Transactions on Engineering Writing and Speech, EWS-14(3):64–67, December 1971. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Broadbent:1975:SPJ

[Bro75] Margaret Broadbent. Standardization in production of journals: a black and white case? IEEE Transactions on Professional Communication, PC-18(3):123–126, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Broadbent:1978:LE

[Bro78]  D. E. Broadbent.  Lan-

guage and ergonomics. IEEE Transactions on Professional Communication, PC-21(1):34– 37, March 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Broughton:1979:TCS

[Bro79] Bradford B. Broughton. Teaching communication skills to engineers and scientists in higher education, John Kirkman, Ed. IEEE Transactions on Professional Communication, PC-22 (4):219–220, December 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Broer:1980:BTW

[Bro80] Jan W. Broer. Breadboarding for technical writing? IEEE Transactions on Professional Communication, PC-23(2):82– 84, June 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Brockmann:1983:WDS

[Bro83a] R. John Brockmann. William D. Skees, Writing handbook for computer professionals. IEEE Transactions on Professional Communication, PC-26(1):45– 46, March 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bronson:1983:MGB

[Bro83b] Judith Gunn Bronson. T. M. Georges, Business and technical writing Cookbook: How to write coherently on the job. IEEE Transactions on Professional Communication, PC-26 (2):90, June 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Broughton:1984:KCU

[Bro84] Bradford B. Broughton. A key course to unlock communication: Letter writing. IEEE Transactions on Professional Communication, PC-27 (4):193–196, December 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Brockmann:1986:NGB

[Bro86] R. John Brockmann. The new generation of books on writing computer user documentation. IEEE Transactions on Professional Communication, PC-29 (2):35–39, June 1986. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bronson:1987:UE

[Bro87] Judith Gunn Bronson. Unfriendly eyes. IEEE Transactions on Professional Communication, PC-30(3):173–178, September 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Brockmann:1988:DPB

[Bro88a] R. J. Brockmann. Desktop publishing — beyond gee whiz. IEEE Transactions on Professional Communication, 31(1): 21–29, March 1988. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Brockmann:1988:DPC

[Bro88b] R. J. Brockmann. Desktop publishing: a critical bibliography. IEEE Transactions on Professional Communication, 31(1):30–36, March 1988. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Brown:1989:CCL

[Bro89] J. R. Brown. Corporate communications and the legal environment: the summary annual report and the shareholder communication rules. IEEE Transactions on Professional Communication, 32(2): 120–124, June 1989. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Brockmann:1991:UDC

[Bro91a] R. J. Brockmann. The unbearable distraction of color. IEEE Transactions on Professional Communication, 34(3): 153–159, September 1991. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Brown:1991:RC

[Bro91b] J. H. U. Brown. The research consortium. IEEE Transactions on Professional Communication, 34(2):75–78, June 1991. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Brown:1993:MFM

[Bro93] C. D. Brown. Male/female mentoring: turning potential risks into rewards. IEEE Transactions on Professional Communication, 36(4):197– 200, December 1993. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bruce:1981:WRF

[BRS81] Bertram Bruce, Andee Rubin, and Kathleen Starr. Why readability formulas fail. IEEE Transactions on Professional Communication, PC-24(1):50– 52, March 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bemer:1973:ICT

[BS73] Robert W. Bemer and A. Richard Shriver. Integrating computer text processing with photocomposition. IEEE Transactions on Professional Communication, PC-16(3):92–96, September 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bedford:1987:TWR

[BS87] Marilyn S. Bedford and F. Cole Stearns. The technical writer’s responsibility for safety. IEEE Transactions on Professional Communication, PC-30(3): 127–132, September 1987. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Budish:1989:SPV

[BS89] B. E. Budish and R. L. Sandhusen. The short proposal: versatile tool for communicating corporate culture in competitive climates. IEEE Transactions on Professional Communication, 32(2):81–85, June 1989. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Boekelder:1998:SSS

[BS98] A. Boekelder and M. Steehouder. Selecting and switching: some advantages of diagrams over tables and lists for presenting instructions. IEEE Transactions on Professional Communication, 41(4): 229–241, December 1998. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bloch:2016:LSU

[BS16] Janel Bloch and Sandra E.

Spataro. Lessons from Scranton: Using scenes from the television series the office to teach topics in professional communication. IEEE Transactions on Professional Communication, 59(3):274–287, September 2016. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Brinegar:1985:TWG

[BSR85] Bonnie Carter Brinegar, Craig Barnwell Skates, and Cynthia H. Rogers. Technical writing: a guide with models. IEEE Transactions on Professional Communication, PC-28(1):52– 53, March 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bosch-Sijtsema:2013:PVW

[BSS13] Petra M. Bosch-Sijtsema and Anu Sivunen. Professional virtual worlds supporting computer-mediated communication, collaboration, and learning in geographically distributed contexts. IEEE Transactions on Professional Communication, 56(2):160– 175, June 2013. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Brasseur:1995:GIC

[BT95] L. E. Brasseur and T. L. Thompson. Gendered ideologies: cultural and social contexts for illustrated medical manuals in Renaissance England. IEEE Transactions on Professional Communication, 38(4):204–215, December 1995. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Brewer:2012:ISS

[BTM12] Pam Estes Brewer, Joo Seng Tan, and James Melton. Introduction to the special section: Navigating the boundaries in global training and education: new literacies, competencies, and practices. IEEE Transactions on Professional Communication, 55(2):156–159, June 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Buckley:1974:DKA

[Buc74] Robert J. Buckley. Digest of keynote address IEEE Professional Communications conference/workshop. IEEE Transactions on Professional Communication, PC-17(2):31–32, June 1974. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Buchholz:1989:DPC

[Buc89] W. J. Buchholz. Deciphering professional codes of ethics. IEEE Transactions on Professional Communication, 32(2): 62–68, June 1989. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Buehler:1976:CRR

[Bue76] Mary Fran Buehler. Creative revision: From rough draft to published paper. IEEE Transactions on Professional Communication, PC-19(2):26– 32, June 1976. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Buehler:1977:RCT

[Bue77] Mary Fran Buehler. Report construction: Tables. IEEE Transactions on Professional Communication, PC-20(1):29– 32, June 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bulwer:1972:PIE

[Bul72]  Lytton Bulwer.  Planning

and implementing effective engineering meetings. IEEE Transactions on Professional Communication, PC-15(1):9– 12, March 1972. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Buntschuh:1971:ISD

[Bun71] R. Buntschuh. IEEE standard definitions: Navigation aid terms. IEEE Transactions on Engineering Writing and Speech, EWS-14(3):79, December 1971. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Burns:1969:RGA

[Bur69]  J. R. Burns.  A report-

generator approach to automated page composition. IEEE Transactions on Engineering Writing and Speech, EWS-12 (2):52–56, May 1969. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Burchinal:1975:MEP

[Bur75] Lee G. Burchinal. Microforms and electronics publication emerging bases for scientific communication. IEEE Transactions on Professional Communication, PC-18(3):174–176, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Burrier:1978:TLP

[Bur78] Dale G. Burrier. Tools of lab photography. IEEE Transactions on Professional Communication, PC-21(2):66–69, June 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Burchard:1984:PAS

[Bur84a] Gina M. Burchard. Professional accountability in the senior engineering lab. IEEE Transactions on Professional Communication, PC-27(2):93– 96, June 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Burke:1984:YNG

[Bur84b] Thomas E. Burke. You’ll never get ahead in engineering if you can’t make yourself understood. IEEE Transactions on Professional Communication, PC-27(1):25–28, March 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Burchard:1985:CDN

[Bur85] Gina M. Burchard. Courseware design: New roles for text and technical writers. IEEE Transactions on Professional Communication, PC-28(1):20– 26, March 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Burton:1990:EMW

[Bur90] B. W. Burton. Editing math: what to do with the symbols. IEEE Transactions on Professional Communication, 33(2): 62–65, June 1990. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bush:1991:CTC

[Bus91] D. Bush. Comparing the two cultures in technical writing. IEEE Transactions on Professional Communication, 34(2): 67–69, June 1991. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Beck:1990:ETP

[BW90] C. E. Beck and K. A. Wegner. Enthusiasm in technical proposals: verifying a method of lexical analysis. IEEE Transactions on Professional Communication, 33(3): 118–123, September 1990. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Boyer:1992:EDC

[BW92] E. P. Boyer and T. G. Webb. Ethics and diversity: a correlation enhanced through corporate communication. IEEE Transactions on Professional Communication, 35(1):38–43, March 1992. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Boettger:2019:GES

[BW19] Ryan K. Boettger and Stefanie Wulff. Gender effects in student technical and scientific writing a corpus-based study. IEEE Transactions on Professional Communication, 62(3): 239–252, September 2019. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Burton:2012:DEV

[BWPH12] Paul Burton, Yu Wu, Victor R. Prybutok, and Gina Harden. Differential effects of the volume and diversity of communication network ties on knowledge workers’ performance. IEEE Transactions on Professional Communication, 55(3):239–253, September 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Campbell:2017:PSP

[CAK^{+}17] Kim Sydow Campbell, Nicole Amare, Erin Kane, Alan D. Manning, and Jefrey S. Naidoo. Plain-style preferences of US professionals. IEEE Transactions on Professional Communication, 60(4): 401–411, December 2017. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Calkins:1974:CAF

[Cal74] Ken Calkins. Cognate the averation fraught. IEEE Transactions on Professional Communication, PC-17(1):18– 20, March 1974. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Campbell:1992:ESS

[Cam92] C. P. Campbell. Engineering style: striving for efficiency. IEEE Transactions on Professional Communication, 35(3): 130–137, September 1992. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Campbell:1995:ECE

[Cam95] C. P. Campbell. Ethos: character and ethics in technical writing. IEEE Transactions on Professional Communication, 38(3):132–138, September 1995. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Campbell:1996:HMB

[Cam96] C. P. Campbell. How much B.S. should I put into this presentation? IEEE Transactions on Professional Communication, 39(2):103–104, June 1996. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Campbell:2007:DJD

[Cam07] Kim Sydow Campbell. Dank je, domo arigato, and thanks.

IEEE Transactions on Professional Communication, 50(4): 273–274, December 2007. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Capen:1978:PTJ

[Cap78a] Edgar C. Capen. Proposition 13 for technical journals. IEEE Transactions on Professional Communication, PC-21 (4):139, December 1978. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Capen:1978:RR

[Cap78b] Edgar C. Capen. A referee’s report. IEEE Transactions on Professional Communication, PC-21(4):144–145, December 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Cappello:1992:QCR

[Cap92] J. R. Cappello. The quality control revolution: new opportunities for corporate communication. IEEE Transactions on Professional Communication, 35(1):26–30, March 1992. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Capps:1993:CSC

[Cap93] R. Capps. Communicating sudden change in tasks and culture: The Inglis Montmagny story. IEEE Transactions on Professional Communication, 36(1):30–34, March 1993. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Carroll:1959:KGA

[Car59] John M. Carroll. Keys to good article writing. IRE Transactions on Engineering Writing and Speech, EWS-2(3):78–82, December 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Carroll:1961:TSY

[Car61] John M. Carroll. Taking the splatter out of your technical writing. IRE Transactions on Engineering Writing and Speech, EWS-4(2):52–55, May 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Carter:1973:IHS

[Car73] H. J. Carter. The IEEE Headquarters’ Senior Editors. IEEE Transactions on Professional Communication, PC-16 (2):40–41, June 1973. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Carroll:1980:TWW

[Car80] John E. Carroll. Technical writes and wrongs. IEEE Transactions on Professional Communication, PC-23(2):68– 71, June 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Carliner:1986:TT

[Car86] Saul Carliner. Topping the text. IEEE Transactions on Professional Communication, PC-29(4):72–74, December 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Carliner:1987:LUO

[Car87] Saul Carliner. Lists: The ultimate organizer for engineering writing. IEEE Transactions on Professional Communication, PC-30(4):218–221, December 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Carle:1989:HHA

[Car89] G. Carle. Handling a hostile audience-with your eyes. IEEE Transactions on Professional Communication, 32(1): 29–31, March 1989. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Carle:1990:GIGa

[Car90a] G. Carle. Good interviewing is good business. I. Strategies for interviewers. IEEE Transactions on Professional Communication, 33(1):19–22, March 1990. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Carle:1990:GIGb

[Car90b] G. Carle. Good interviewing is good business. II. Strategies for applicants. IEEE Transactions on Professional Communication, 33(2):67–69, June 1990. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Carliner:1990:EES

[Car90c]  S. Carliner.  Elements of

editorial style for computerdelivered information. IEEE Transactions on Professional Communication, 33(1):38– 45, March 1990. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Carnegie:2000:IWH

[Car00] T. A. M. Carnegie. The Internet Writer’s Handbook [Book Review]. IEEE Transactions on Professional Communication, 43(2):219–221, June 2000. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Carliner:2012:DAS

[Car12] Saul Carliner. Different approaches to similar challenges: an analysis of the occupational cultures of the disciplines of technical communication and training tutorial. IEEE Transactions on Professional Communication, 55(2): 160–174, June 2012. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Carradini:2016:JHW

[Car16] Stephen Carradini. John Hartley, Wen Wen, and Henry Siling Li, Creative Economy and Culture: Challenges, Changes, and Futures for the Creative Industries. IEEE Transactions on Professional Communication, 59(4):428–430, December 2016. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Carradini:2020:CRT

[Car20] Stephen Carradini. A comparison of research topics associated with technical communication, business communication, and professional communication, 1963–2017. IEEE Transactions on Professional Communication, 63(2):118– 138, June 2020. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Carthell:2024:ABL

[Car24] Alicia J. Carthell. An analysis of bias in language content in books used in technical and professional writing courses: a diversity, equity, inclusion, and social justice matter. IEEE Transactions on Professional Communication, 67(1):26–46, March 2024. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Casari:1984:RTH

[Cas84] Laura E. Casari. Required:

Three hours in technical communications paradigm for a paradox. IEEE Transactions on Professional Communication, PC-27(3):116–120, September 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Casey:1986:GAS

[Cas86] Bernice E. Casey. A games approach to system interface design. IEEE Transactions on Professional Communication, PC-29(4):65–71, December 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Cawkell:1975:CBE

[Caw75] A. E. Cawkell. Connections between engineering and science. IEEE Transactions on Professional Communication, PC-18 (2):71–73, June 1975. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Carlson:2008:UCM

[CB08] Patricia A. Carlson and Frederick C. Berry. Using computermediated peer review in an engineering design course. IEEE Transactions on Professional Communication, 51(3):264– 279, September 2008. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Chai:2009:IOI

[CBSM^{+}09] Sangmi Chai, Sharmistha Bagchi-Sen, Claudia Morrell, H. R. Rao, and Shambhu J. Upadhyaya. Internet and online information privacy: an exploratory study of preteens and early teens. IEEE Transactions on Professional Communication, 52(2):167–182, June 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Connor:1988:SBT

[CC88] J. J. Connor and J. T. H. Connor. Semantic bypassing in technical communication: the historical case of antiseptics. IEEE Transactions on Professional Communication, 31(1): 13–17, March 1988. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Chen:2019:AHW

[CC19] Adela Chen and Gretchen Irwin Casterella. After-hours work connectivity: Technological antecedents and implications. IEEE Transactions on Professional Communication, 62(1):75–93, March 2019. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Cheung:2022:PFP

[CC22a] Ming Cheung and Zhen Troy Chen. Power, freedom, and privacy on a discipline-and-control Facebook, and the implications for Internet governance. IEEE Transactions on Professional Communication, 65(4):467–484, December 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Clem:2022:URS

[CC22b] Sam Clem and Ryan Cheek. Unjust revisions: a social justice framework for technical editing. IEEE Transactions on Professional Communication, 65(1):135–150, March 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Cao:2024:UIC

[CCL24] Daqin Cao, Jianfen Chen, and Shijie Liu. University industry collaboration in managing translation projects: Perceptions and responses from students, instructors, and industry partners. IEEE Transactions on Professional Communication, 67(3):352–368, September 2024. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Callery:1998:CMD

[CD98] Bea Callery and Matt Doyle. Colin McCormack and David Jones, Building A Web-Based Education System [Book Review]. IEEE Transactions on Professional Communication, 41(4):286–288, December 1998. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Crawford:1984:CSM

[CDL84] C. C. Crawford, John W.

Demidovich, and William N. Ledbetter. The Crawford slip method. IEEE Transactions on Professional Communication, PC-27(2):109–110, June 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Croteau:2010:ERP

[CDM10] Anne-Marie Croteau, Linda Dyer, and Marco Miguel. Employee reactions to paper and electronic surveys: an experimental comparison. IEEE Transactions on Professional Communication, 53(3):249– 259, September 2010. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Coppola:2005:BSB

[CE05] N. W. Coppola and N. Elliot. Big science or bricolage: an alternative model for research in technical communication. IEEE Transactions on Professional Communication, 48(3):261–268, September 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Charrow:1987:CEL

[CEH87] Veda R. Charrow, Myra K.

Erhardt, and James W. Hill. Clear and effective legal writing. IEEE Transactions on Professional Communication, PC-30(2):113–114, June 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Canavero:2014:IJA

[CFMM14] Flavio Canavero, Fiorenzo Franceschini, Domenico Maisano, and Luca Mastrogiacomo. Impact of journals and academic reputations of authors: a structured bibliometric survey of the IEEE publication galaxy. IEEE Transactions on Professional Communication, 57 (1):17–40, March 2014. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Chandler:1983:TWH

[CG83] Harry E. Chandler and William Giesecke. Technical writer’s handbook. IEEE Transactions on Professional Communication, PC-26(2):90, June 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Clauser:1989:SWW

[CH89] J. K. Clauser and J. W. Hill. Suggestions for those who write instructions for the preparation of large government proposals or why do you make it so difficult for us to make it easy for you? IEEE Transactions on Professional Communication, 32(1):17–19, March 1989. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Chapline:1958:TT

[Cha58] J. D. Chapline. Tricks of the trade. IRE Transactions on Engineering Writing and Speech, EWS-1(1):6–11, March 1958. ISSN 0893-7869 (print), 2331-3692 (electronic).

Chapline:1959:HBD

[Cha59] J. D. Chapline. How to become a dictator. IRE Transactions on Engineering Writing and Speech, EWS-2(2):42– 45, June 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Chapline:1960:EFS

[Cha60] J. D. Chapline. The editorial function in scientific organization. IRE Transactions on Engineering Writing and Speech, EWS-3(2):48–53, July 1960. ISSN 0893-7869 (print), 2331-3692 (electronic).

Chadwell:1965:CWT

[Cha65] E. A. Chadwell. The Careful Writer, by Theodore M. Bernstein. IEEE Transactions on Engineering Writing and Speech, EWS-8(2):81, December 1965. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Chamberlin:1977:TE

[Cha77] Hal Chamberlin. Text editing.

IEEE Transactions on Professional Communication, PC-20 (1):13–15, June 1977. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Chandler:1978:TMT

[Cha78] Harry E. Chandler. Too much TV or? IEEE Transactions on Professional Communication, PC-21(4):149, December 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Chaffee:1980:BTB

[Cha80a] Patricia Chaffee. Basic technical and business writing, Joanna M. Freeman. IEEE Transactions on Professional Communication, PC-23(3): 153, September 1980. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Chaffee:1980:SBT

[Cha80b] Patricia Chaffee. Strategies for business and technical writing, Keven J. Harty. IEEE Transactions on Professional Communication, PC-23(4):200–201, December 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Chaffee:1981:PLS

[Cha81] Patricia Chaffee. Putt’s law and the successful technocrat — Archibald Putt. IEEE Transactions on Professional Communication, PC-24(4): 202, December 1981. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Changxin:1984:WCN

[Cha84] Fan Changxin. What’s in a Chinese name? IEEE Transactions on Professional Communication, PC-27(4):215–216, December 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Charney:1996:NWS

[Cha96] D. Charney. Nonacademic Writing: Social Theory and Technology [Book Review]. IEEE Transactions on Professional Communication, 39(4): 239, December 1996. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Chambers:2002:PAP

[Cha02] T. L. Chambers. A Peircian approach to professional ethics instruction. IEEE Transactions on Professional Communication, 45(1):45–49, March 2002. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Chase:2006:FIP

[Cha06] P. Chase. Freedom of information in a post 9-11 world. IEEE Transactions on Professional Communication, 49(3): 289–290, September 2006. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Chase:2023:KTP

[Cha23] Josh Chase. Keywords in Technical and Professional Communication (Yu, H. and Buehl, J., Eds.) [book review]. IEEE Transactions on Professional Communication, 66(4):407– 409, December 2023. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Chengzhao:1992:TES

[Che92] Lou Chengzhao. Translating English scientific and technical terms into Chinese: comparing the practice in mainland China and Taiwan. IEEE Transactions on Professional Communication, 35(1):19–23, March 1992. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Cheung:2017:PLM

[Che17] Iva W. Cheung. Plain language to minimize cognitive load: a social justice perspective. IEEE Transactions on Professional Communication, 60(4):448–457, December 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Childers:1989:JGA

[Chi89] L. Childers. J. Grunig’s asymmetrical and symmetrical models of public relations: contrasting features and ethical dimensions. IEEE Transactions on Professional Communication, 32(2):86–93, June 1989. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Crowston:2007:RFF

[CHME07] Kevin Crowston, James Howison, Chengetai Masango, and U. Yeliz Eseryel. The role of face-to-face meetings in technology-supported selforganizing distributed teams. IEEE Transactions on Professional Communication, 50(3): 185–203, September 2007. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Chociey:1997:WRM

[Cho97a] P. A. Chociey. “Who’s reading my e-mail?”: a study of professionals’ e-mail usage and privacy perceptions in the workplace. IEEE Transactions on Professional Communication, 40(1):34–40, March 1997. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Chou:1997:CNC

[Cho97b] Chien Chou. Computer networks in communication survey research. IEEE Transactions on Professional Communication, 40(3):197–208, September 1997. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Chong:2018:IUT

[Cho18] Felicia Chong. Implementing usability testing in introductory technical communication service courses: Results and lessons from a local study.

IEEE Transactions on Professional Communication, 61(2): 196–205, June 2018. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Choi:2021:BCL

[Cho21] Stephen Choi. Better communication leads to a higher output? An analysis of pair communication on pair programming productivity. IEEE Transactions on Professional Communication, 64(4):338– 353, December 2021. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Coppola:2004:BTV

[CHR04] N. W. Coppola, S. R. Hiltz, and N. G. Rotter. Building trust in virtual teams. IEEE Transactions on Professional Communication, 47(2):95–104, June 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Carsten:2002:PAS

[CI02] L. D. Carsten and D. L. Illman. Perceptions of accuracy in science writing. IEEE Transactions on Professional Communication, 45(3):153– 156, September 2002. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Cherry:1989:OHE

[CJ89] J. M. Cherry and S.-R. Jackson. Online help: effects of content and writing style on user performance and attitudes. IEEE Transactions on Professional Communication, 32(4):294–299, December 1989. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Campbell:2019:ELR

[CL19] Kim Sydow Campbell and Chris Lam. The effect of leader rapport-management feedback on leader member relationship quality and perceived group effectiveness in student teams. IEEE Transactions on Professional Communication, 62(3): 253–262, September 2019. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Clarke:1959:MSW

[Cla59] Ralph Emerson Clarke. Methods for the study of writing and speech techniques. IRE Transactions on Engineering Writing and Speech, EWS-2(1):12– 14, January 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Clark:1975:RTW

[Cla75] Andrew K. Clark. Readability in technical writing-principles and procedures. IEEE Transactions on Professional Communication, PC-18(2):67–70, June 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Clark:1979:HPI

[Cla79] Richard L. Clark. How to prepare inspection reports. IEEE Transactions on Professional Communication, PC-22 (4):210–213, December 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Clark:1987:ETC

[Cla87] Gregory Clark. Ethics in technical communication: a rhetorical perspective. IEEE Transactions on Professional Communication, PC-30(3):190–195, September 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Clayton:2012:DVL

[Cla12] Garth Clayton. Designing Visual Language: Strategies for Professional Communicators, Second Edition (Kostelnick, C. and Roberts, D. D.; 2011) [Book Review]. IEEE Transactions on Professional Communication, 55(1):100–101, March 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Clark:2016:CSI

[Cla16] Dave Clark. Content strategy: an integrative literature review. IEEE Transactions on Professional Communication, 59(1):7–23, March 2016.

CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Clement:1987:HFI

[Cle87] David E. Clement. Human factors, instructions and warnings, and products liability. IEEE Transactions on Professional Communication, PC-30 (3):149–156, September 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Cleary:2003:NGB

[Cle03] K. Cleary. Narain Gehani, Bell Labs: Life in the Crown Jewel. IEEE Transactions on Professional Communication, 46(4): 331–332, December 2003. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Cleary:2016:CPP

[Cle16] Yvonne Cleary. Community of practice and professionalization perspectives on technical communication in Ireland. IEEE Transactions on Professional Communication, 59(2): 126–139, June 2016. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Climenson:1963:LAP

[Cli63] W. D. Climenson. Languageanalysis problems in the computer processing of natural text. IEEE Transactions on Engineering Writing and Speech, EWS-6(2):72–78, December 1963. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Clothier:1964:WCP

[Clo64] John W. Clothier. Writing calibration procedures. IEEE Transactions on Engineering Writing and Speech, EWS-7 (1):6–9, March 1964. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Craig:2008:IAC

[CLP08] Jennifer L. Craig, Neal Lerner, and Mya Poe. Innovation across the curriculum: Three case studies in teaching science and engineering communication. IEEE Transactions on Professional Communication, 51(3):280–301, September 2008. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Carter:2001:WAP

[CM01] J. Carter and M. Markel. Web accessibility for people with disabilities: an introduction for Web developers. IEEE Transactions on Professional Communication, 44(4): 225–233, December 2001. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Coppola:2002:CST

[CM02] N. W. Coppola and R. Myre. Corporate software training: is Web-based training as effective as instructor-led training? IEEE Transactions on Professional Communication, 45(3): 170–186, September 2002. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Cooke:2005:CUE

[CM05] L. Cooke and S. Mings. Connecting usability education and research with industry needs and practices. IEEE Transactions on Professional Communication, 48(3):296– 312, September 2005. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Campbell:2020:HSS

[CNC20] Kim Sydow Campbell, Jefrey S. Naidoo, and Sean M. Campbell. Hard or soft sell? Understanding white papers as content marketing. IEEE Transactions on Professional Communication, 63(1): 21–38, March 2020. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Cochran:2004:CPV

[Coc04]  M. E. Cochran.  “Cyni-

cal Politics” versus “Scientific Truth”: communication crisis at Brookhaven National Lab. IEEE Transactions on Professional Communication, 47(4): 256–267, December 2004. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Codone:2004:RDS

[Cod04] S. Codone. Reducing the distance: a study of course websites as a means to create a total learning space in traditional courses. IEEE Transactions on Professional Communication, 47(3):190–199, September 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Codone:2009:JLD

[Cod09]  Susan Codone.  Jean-luc

Doumont, Trees, Maps, and Theorems: Effective Communication for Rational Minds — Book Review. IEEE Transactions on Professional Communication, 52(4):411– 412, December 2009. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Coggio:2015:TCA

[Cog15] Grace Leinbach Coggio. Technical communicators as agents and adopters of change: a case study of the implementation of an early content-management system. IEEE Transactions on Professional Communication, 58(3):271–288, September 2015. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Cohen:1960:NTI

[Coh60] Sol Cohen. Notes toward improving technical reports. IRE Transactions on Engineering Writing and Speech, EWS-3(2): 62–65, July 1960. ISSN 0893-7869 (print), 2331-3692 (electronic).

Cohen:1969:WPU

[Coh69] Gerald I. Cohen. A world without people using the personal pronoun to improve technical writing. IEEE Transactions on Engineering Writing and Speech, EWS-12(3):79–82, October 1969. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Coile:1969:ISE

[Coi69] Russell C. Coile. Information sources for electrical and electronics engineers. IEEE Transactions on Engineering Writing and Speech, EWS-12(3): 71–78, October 1969. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Cole:1965:PWP

[Col65] L. M. Cole. Proper Words in Proper Places, by Paul I. Richards and Irving T. Richards. IEEE Transactions on Engineering Writing and Speech, EWS-8(2):81, December 1965. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Colby:1977:PTW

[Col77] John B. Colby. Paragraphing in technical writing. IEEE Transactions on Professional Communication, PC-20(1):20– 23, June 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Compton:1986:RTS

[Com86] Joan Callahan Compton. Results of a telephone survey of technical documentation users. IEEE Transactions on Professional Communication, PC-29(4):93–100, December 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Connolly:1958:SMP

[Con58a] Thomas E. Connolly. Some major pitfalls in technical writing. IRE Transactions on Engineering Writing and Speech, EWS-1(2):22–25, August 1958. ISSN 0893-7869 (print), 2331-3692 (electronic).

Connors:1958:DSE

[Con58b] Thomas Connors. The double standard in engineering writing. IRE Transactions on Engineering Writing and Speech, EWS-1(2):6–8, August 1958. ISSN 0893-7869 (print), 2331-3692 (electronic).

Connelly:1970:LDI

[Con70] J. Campbell Connelly. Listening development in an industrial environment. IEEE Transactions on Engineering Writing and Speech, EWS-13(2):84–89, September 1970. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Connelly:1978:OBL

[Con78] J. Campbell Connelly. The one bad listening habit. IEEE Transactions on Professional Communication, PC-21(2):53– 56, June 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Conway:1983:PP

[Con83] Richard A. Conway. To publish or not to publish. IEEE Transactions on Professional Communication, PC-26(1):17– 19, March 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Conrads:1987:IIS

[Con87]  John A. Conrads.  “Illus-

truction”: Increasing safety and reducing liability exposure. IEEE Transactions on Professional Communication, PC-30 (3):133–135, September 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Connor:1991:HST

[Con91] J. J. Connor. History and the study of technical communication in Canada and the United States. IEEE Transactions on Professional Communication, 34(1):3–6, March 1991. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Coney:1992:TRT

[Con92] M. B. Coney. Technical readers and their rhetorical roles. IEEE Transactions on Professional Communication, 35(2): 58–63, June 1992. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Constantinides:1998:JHW

[Con98] H. Constantinides. Jürgen Habermas “What is universal pragmatics?”. IEEE Transactions on Professional Communication, 41(2):143–145, June 1998. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Constantinides:1999:LWR

[Con99] H. Constantinides. Linguistics At Work: a Reader of Applications [Book Review]. IEEE Transactions on Professional Communication, 42(1): 67–68, March 1999. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Cook:1965:SEE

[Coo65] Howard L. Cook. Symbols for electricity and electronics. IEEE Transactions on Engineering Writing and Speech, EWS-8(2):60–66, December 1965. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Cooke:2010:ACT

[Coo10] Lynne Cooke. Assessing concurrent think-aloud protocol as a usability test method: a technical communication approach. IEEE Transactions on Professional Communication, 53(3): 202–215, September 2010. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Corley:1969:HQC

[Cor69] F. C. Corley. High-quality computer typesetting for text, formulas, and listings. IEEE Transactions on Engineering Writing and Speech, EWS-12 (2):46–49, May 1969. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Corey:1983:LYA

[Cor83] Jim Corey. Lead your ACE:

Accuracy, clarity, and effectiveness in technical writing. IEEE Transactions on Professional Communication, PC-26 (1):13–14, March 1983. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Corbett:1990:CAS

[Cor90] M. Z. Corbett. Clearing the air: some thoughts on genderneutral writing. IEEE Transactions on Professional Communication, 33(1):2–6, March 1990. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Corey:1992:CRS

[Cor92] J. R. Corey. Cultural reentry shock: using the professional writing class to help foreign students. IEEE Transactions on Professional Communication, 35(3):169–171, September 1992. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Corey:1994:MTT

[Cor94] J. Corey. Mark twain-technical writer. IEEE Transactions on Professional Communication, 37(2):108, June 1994. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Costello:1978:JCD

[Cos78]  John Costello.  Jests can

do justice to your speeches. IEEE Transactions on Professional Communication, PC-21 (2):60–63, June 1978. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Cosgrove:1984:WA

[Cos84] M. C. Cosgrove. Writing for the audience. IEEE Transactions on Professional Communication, PC-27(1):38–41, March 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Cosgrove:2023:FGC

[Cos23a] Samantha Jo Cosgrove. Finding the gap: a comparison of UX industry practices and UX course outcomes in TPC programs. IEEE Transactions on Professional Communication, 66(4):356–365, December 2023. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Cosgrove:2023:ITR

[Cos23b] Samantha Jo Cosgrove. Improving technical and risk communication: an organizational study of North Carolina Emergency Management and Hurricane Florence. IEEE Transactions on Professional Communication, 66(3):284– 299, September 2023. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Cotcher:1959:IPS

[Cot59]  Alfred L. Cotcher.  IRE

publications — the section bulletin. IRE Transactions on Engineering Writing and Speech, EWS-2(1):34–39, January 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Cottone:1979:WID

[Cot79]  James F. Cottone.  Writ-

ing an invention disclosure. IEEE Transactions on Professional Communication, PC-22 (2):105–108, June 1979. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Coviello:1975:OAJ

[Cov75] Roselle Coviello. Organization and administration of journal production in commercial publishing. IEEE Transactions on Professional Communication, PC-18(3):139–140, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Cowen:1980:GUG

[Cow80] Robert C. Cowen. Garbage under glass: What are scientists dishing out? IEEE Transactions on Professional Communication, PC-23(3):116–117, September 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Cox:1980:TTL

[Cox80] John L. Cox. Technical topic, lay audience: How to make a good presentation. IEEE Transactions on Professional Communication, PC-23(1):26– 28, March 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Cox:1993:WSG

[Cox93] J. P. Cox. Waking a sleeping giant — the planning for success of a large project process at a large urban teaching hospital. IEEE Transactions on Professional Communication, 36 (1):20–23, March 1993. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Coyne:1977:DBP

[Coy77] Joseph G. Coyne. Data base publishing at NTIS. IEEE Transactions on Professional Communication, PC-20(2):94– 95, September 1977. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Cox:1973:HRC

[CR73] Barbara G. Cox and Charles G.

Roland. How rhetoric confuses scientific issues. IEEE Transactions on Professional Communication, PC-16(3):140–142, September 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Coney:1984:CCE

[CR84] Mary B. Coney and Judith A.

Ramey. A communication curriculum in engineering education: an alternative model. IEEE Transactions on Professional Communication, PC-27 (3):138–143, September 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Carroll:2006:CSM

[CR06] John M. Carroll and Mary Beth Rosson. Case studies as minimalist information. IEEE Transactions on Professional Communication, 49(4):297– 310, December 2006. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Crane:1968:PQP

[Cra68] Glenn B. Crane. Producing a quality proposal at reasonable cost. IEEE Transactions on Engineering Writing and Speech, EWS-11(3): 160–165, December 1968. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Craig:1983:DTB

[Cra83a] James Craig. Designing with type: a basic course in typography. IEEE Transactions on Professional Communication, PC-26(4):213, December 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Craig:1983:PDM

[Cra83b]  James Craig.  Phototype-

setting: a design manual. IEEE Transactions on Professional Communication, PC-26 (4):213–214, December 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Crawford:1983:HGO

[Cra83c] C. C. Crawford. How to gather and organize ideas quickly. IEEE Transactions on Professional Communication, PC-26 (4):187–190, December 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Crandall:1985:WAE

[Cra85a] Kathleen Eilers Crandall.

Writing for adult English language learners. IEEE Transactions on Professional Communication, PC-28(4):3–10, December 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Crane:1985:WCC

[Cra85b]  Dave Crane.  Writing for

closed-captioned television for the hearing-impaired. IEEE Transactions on Professional Communication, PC-28(4):15– 18, December 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Craig:1992:USD

[Cra92] J. S. Craig. Using a structured design analysis to simplify complex in-house computer manuals. IEEE Transactions on Professional Communication, 35(1):7–12, March 1992. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Crane:1998:UBP

[Cra98] T. B. Crane. Using business performance to communicate change. IEEE Transactions on Professional Communication, 41(4):266–269, December 1998. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Crede:1995:EEC

[Cre95] K. L. Crede. Environmental effects of the computer age. IEEE Transactions on Professional Communication, 38(1): 33–40, March 1995. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Cross:1959:PTE

[Cro59] Alexander H. Cross. Publications as tools of the electronics industry. IRE Transactions on Engineering Writing and Speech, EWS-2(2):49– 51, June 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Crow:1980:HCT

[Cro80] Porter J. Crow. How to change their minds. IEEE Transactions on Professional Communication, PC-23(1):48– 49, March 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Cronin:1986:RLP

[Cro86]  Michael W. Cronin.  The

role of language in preventing or ameliorating job burnout. IEEE Transactions on Professional Communication, PC-29 (3):3–9, September 1986. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Cronin:1995:MSI

[Cro95] G. Cronin. Marketability and social implications of interactive TV and the information superhighway. IEEE Transactions on Professional Communication, 38(1):24–32, March 1995. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Crum:1973:FMS

[Cru73] John K. Crum. Financing a multijournal system in the 1970’s. IEEE Transactions on Professional Communication, PC-16(3):71–72, September 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Crystal:2002:LI

[Cry02] D. Crystal. Language and the internet. IEEE Transactions on Professional Communication, 45(2):142–144, June 2002. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Chou:1996:CNS

[CS96] Chien Chou and Chuen-Tsai Sun. A computer-network-supported cooperative distance learning system for technical communication education. IEEE Transactions on Professional Communication, 39(4): 205–214, December 1996. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Chan:2006:ISI

[CS06] H. C. Chan and K. Siau. Introduction to the special issue on communication issues in user database interaction. IEEE Transactions on Professional Communication, 49(2): 93–95, June 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Caernarven-Smith:1984:AAR

[CSH84] Patricia Caernarven-Smith and John Hoffman. Audience analysis and response. IEEE Transactions on Professional Communication, PC-27(4):233–234, December 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Cabezas:2020:EPP

[CSSV20] Paula Cabezas, Clay Spinuzzi, Omar Sabaj, and Germán Varas. Editing the pitch: Patterns of editing strategies of written pitches in a Chilean accelerator program. IEEE Transactions on Professional Communication, 63(4): 296–310, December 2020. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Cuffe:2019:KHD

[Cuf19] Paul Cuffe. K. Healy: Data Visualization: a Practical Introduction [Book Review]. IEEE Transactions on Professional Communication, 62(3):310– 311, September 2019. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Cuffe:2020:DVP

[Cuf20] Paul Cuffe. Data Visualization: a Practical Introduction: Kieran Healy [Book Review]. IEEE Transactions on Professional Communication, 63(4): 400–401, December 2020. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Cunningham:2009:RMT

[Cun09] Donald H. Cunningham. Righting the Mother Tongue: From Olde English to Email, the Tangled Story of English Spelling (Wolman, D.; 2008) [Book Review]. IEEE Transactions on Professional Communication, 52(2):219–220, June 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Cunningham:2010:LWT

[Cun10] Donald H. Cunningham. The Language of Work: Technical Communication at Lukens Steel 1810 to 1925 (Johnson, C. J.; 2009) [Book Review]. IEEE Transactions on Professional Communication, 53(2): 185–186, June 2010. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Curtis:1980:TBR

[Cur80] Ethell Curtis. Three basic recipes for a speech. IEEE Transactions on Professional Communication, PC-23(1):28– 31, March 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Cusack:1993:ESH

[Cus93] M. W. Cusack. Efforts to simplify human-computer communication. IEEE Transactions on Professional Communication, 36(1):14–19, March 1993. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Carroll:1996:TMA

[CvdM96] J. M. Carroll and H. van der Meij. Ten misconceptions about minimalism. IEEE Transactions on Professional Communication, 39(2):72–86, June 1996. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Charney:1990:SFA

[CW90] M. J. Charney and J. H. Williams. From start to finish: approaches to introductions and conclusions in technical writing textbooks. IEEE Transactions on Professional Communication, 33(4): 220–225, December 1990. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Chen:2019:HTS

[CWSL19] Xiayu Chen, Shaobo Wei, Chunhua Sun, and Yezheng Liu. How technology support for contextualization affects enterprise social media use: a media system dependency perspective. IEEE Transactions on Professional Communication, 62(3):279–297, September 2019. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

DAprix:1964:URP

[D’A64]  Roger M. D’Aprix.  Us-

ing reprints to probe the market. IEEE Transactions on Engineering Writing and Speech, EWS-7(2):5–7, September 1964. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

DAprix:1965:QCE

[D’A65] Roger M. D’Aprix. Quality control for engineering authorship. IEEE Transactions on Engineering Writing and Speech, EWS-8(1):15–17, June 1965. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Daley:1981:PYV

[Dal81] Kevin R. Daley. Presenting your viewpoint. IEEE Transactions on Professional Communication, PC-24(2):91–94, June 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ladd:1992:WDD

[dALT92] M. de Armas Ladd and M. Tangum. What difference does inherited difference make? Exploring culture and gender in scientific and technical professions. IEEE Transactions on Professional Communication, 35(3):183–190, September 1992. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Daniel:1977:BJI

[Dan77] Robert S. Daniel. Birth of a journal: The initiation. IEEE Transactions on Professional Communication, PC-20(2):82– 84, September 1977. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Darrow:1982:HAA

[Dar82] Karl K. Darrow. How to address the American Physical Society. IEEE Transactions on Professional Communication, PC-25(3):154–157, September 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Datta:1991:AWR

[Dat91] J. Datta. Addressing a worldwide readership through the filter of translation. IEEE Transactions on Professional Communication, 34(3):147– 150, September 1991. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Davis:1967:ERE

[Dav67] Richard M. Davis. Experimental research in the effectiveness of technical writing. IEEE Transactions on Engineering Writing and Speech, EWS-10(2):33–38, December 1967. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Davis:1974:CSM

[Dav74] Richard M. Davis. The communication situation a model and discussion. IEEE Transactions on Professional Communication, PC-17(1):9–17, March 1974. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Davey:1981:DCC

[Dav81a] Terrance M. Davey. Drawings can communicate better. IEEE Transactions on Professional Communication, PC-24 (3):148–150, September 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Davies:1981:NSP

[Dav81b] David Davies. Next slide please. IEEE Transactions on Professional Communication, PC-24(4):196–197, December 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Davis:1985:PPJ

[Dav85] Richard M. Davis. Publication in professional journals: a survey of editors. IEEE Transactions on Professional Communication, PC-28(2):34–43, June 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Davis:1993:CEI

[Dav93] W. A. Davis. Comments on “An ethical imperative for technical communicators” by M. Markel. IEEE Transactions on Professional Communication, 36(4):215–216, December 1993. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic). See [Mar93].

Davis:2010:ATC

[Dav10] Marjorie T. Davis. Assessing technical communication within engineering contexts tutorial. IEEE Transactions on Professional Communication, 53(1):33–45, March 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dawson:2010:RPT

[Daw10] Joseph A. Dawson. -A Research Primer for Technical Communication: Methods, Exemplars, and Analyses (Hughes, M. A. and Hayhoe, G. F.) [Book Review]. IEEE Transactions on Professional Communication, 53 (2):181–182, June 2010. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Day:1977:HWS

[Day77]  Robert A. Day.  How to

write a scientific paper. IEEE Transactions on Professional Communication, PC-20(1):32– 37, June 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Day:1983:EW

[Day83] Yvonne Lewis Day. The economics of writing. IEEE Transactions on Professional Communication, PC-26(1):4– 8, March 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

De:2024:MUT

[DB24] Papiya De and Medha Bakhshi.

Managing uncertainties in technology-mediated communication: a qualitative study of business students’ perception of emoji/emoticon usage in a business context. IEEE Transactions on Professional Communication, 67 (2):211–228, June 2024. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dowhal:1993:PVT

[DBK^{+}93] D. Dowhal, G. Bist, P. Kohlmann, S. Musker, and H. Rogers. Producing a video on a technical subject: a guide. IEEE Transactions on Professional Communication, 36(2):62–69, June 1993. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dugger:1973:AER

[DBM73] Gordon L. Dugger, Ruth F.

Bryans, and William T. Morris. AIAA experiments and results on SDD, synoptics, miniprints, and related topics. IEEE Transactions on Professional Communication, PC-16 (3):100–106, September 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dressel:1987:ABE

[DC87] Susan Dressel and Joe Chew. Authenticity beats eloquence. IEEE Transactions on Professional Communication, PC-30 (2):82–83, June 1987. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Downing:2007:UES

[DC07] Joe R. Downing and Russell S. Clark. Using electronic surveys in organizational/employee communication research: a study at GE’s Global Research Center. IEEE Transactions on Professional Communication, 50(3): 249–262, September 2007. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

DeWitt:1972:EPC

[De 72] David De Witt. Effective professional communication: a PSYCHOM ’72 paper. IEEE Transactions on Professional Communication, PC-15(3):85– 86, September 1972. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

DeMare:1981:CT

[De 81] George De Mare. Communicating at the top. IEEE Transactions on Professional Communication, PC-24(2):99–101, June 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Debs:1986:BR

[Deb86] Mary Beth Debs. Book reviews. IEEE Transactions on Professional Communication, PC-29(2):35, June 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dressel:1983:AAS

[DEBD83] Susan Dressel, James S. Euler, Susan A. Bagby, and Sherry A. Dell. ASAPP: Automated systems approach to proposal production. IEEE Transactions on Professional Communication, PC-26(2):63–68, June 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Decker:1984:GSW

[Dec84] Bert Decker. A good speech is worth a thousand (written) words. IEEE Transactions on Professional Communication, PC-27(1):32–34, March 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Degise:1980:SAB

[Deg80] Robert F. Degise. A systems approach to business writing. IEEE Transactions on Professional Communication, PC-23 (2):77–78, June 1980. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Delaney:1981:WHA

[Del81] W. A. Delaney. When and how to argue. IEEE Transactions on Professional Communication, PC-24(4):186–188, December 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dell:1992:CBT

[Del92] Sherry Dell. A communicationbased theory of the glass ceiling: rhetorical sensitivity and upward mobility within the technical organization. IEEE Transactions on Professional Communication, 35(4):230– 235, December 1992. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Delanghe:2000:ULS

[Del00] S. Delanghe. Using learning styles in software documentation. IEEE Transactions on Professional Communication, 43(2):201–205, June 2000. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Denton:1978:CEE

[Den78] L. W. Denton. Continuing engineering education in technical writing. IEEE Transactions on Professional Communication, PC-21(4):167–169, December 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Denton:1983:BRB

[Den83a] L. W. Denton. Book reviews:

Betty Edwards, Drawing on the right side of the brain. IEEE Transactions on Professional Communication, PC-26(2):85, June 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Denton:1983:HTW

[Den83b] L. W. Denton. Handbook of technical writing, 2nd edition, Charles T. Brusaw and Gerald J. Alred and Walter E. Oliu. IEEE Transactions on Professional Communication, PC-26 (2):87, June 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Denton:1983:BWH

[Den83c] L. W. Denton. The business writer’s handbook, Charles T. Brusaw and Gerald J. Alred and Walter E. Oliu and L. W. Denton. IEEE Transactions on Professional Communication, PC-26(2):87, June 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dennett:1988:SBT

[Den88] J. T. Dennett. Not to say is better than to say: how rhetorical structure reflects cultural context in Japanese-English technical writing. IEEE Transactions on Professional Communication, 31(3):116– 119, September 1988. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dennett:1996:AUG

[Den96] J. Temple Dennett. Australia uses genre analysis to address workplace literacy. IEEE Transactions on Professional Communication, 39(3):115– 116, September 1996. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dennett:1998:CRA

[Den98a] J. T. Dennett. Collin Randlesome and Andrew Myers, “Cultural Fluency: Results From A UK and Irish Survey”. IEEE Transactions on Professional Communication, 41(1):70–71, March 1998. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dennett:1998:ABC

[Den98b]  J. Temple Dennett.  Ad-

dressing bias in clip art provided with popular software. IEEE Transactions on Professional Communication, 41(4):

270–273, December 1998. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dennett:1999:CQ

[Den99] J. Temple Dennett. A copyright questionnaire. IEEE Transactions on Professional Communication, 42(2):114– 116, June 1999. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dennett:2000:GBN

[Den00]  J. Temple Dennett.  Go-

ing beyond the native speaker in technical communication. IEEE Transactions on Professional Communication, 43(3): 327–330, September 2000. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dery:1999:EVC

[Der99] M. Dery. Escape velocity: cyberculture at the end of the century. IEEE Transactions on Professional Communication, 42(2):141–142, June 1999. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

DeTienne:1994:BBW

[DeT94] K. B. DeTienne. Big Brother is watching: computer monitoring and communication. IEEE Transactions on Professional Communication, 37 (1):5–10, March 1994. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Devivo:1975:CES

[Dev75] Anita Devivo. Copy editing standards at the American Psychological Association. IEEE Transactions on Professional Communication, PC-18 (3):141–144, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Demer:1970:MIP

[DF70] Louis J. Demer and Louis H. Fentnor. Multiple-image presentation of technical papers — the concept. IEEE Transactions on Engineering Writing and Speech, EWS-13(1):2– 8, May 1970. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Demer:1975:MIP

[DF75] Louis J. Demer and Louis H. Fentnor. Multiple-image presentation of technical papers the concept. IEEE Transactions on Professional Communication, PC-18(1):11– 17, March 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Doheny-Farina:1987:LEA

[DF87] Stephen Doheny-Farina. Legal and ethical aspects of technical communication: a special issue. IEEE Transactions on Professional Communication, PC-30(3):119–120, September 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Drott:1975:CTO

[DG75] M. Carl Drott and Belver C.

Griffith. Characteristics of technically oriented house journals. IEEE Transactions on Professional Communication, PC-18(1):45–49, March 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

deGroot:2011:TNP

[dGKGN11] Elizabeth de Groot, Hubert Korzilius, Marinel Gerritsen, and Catherine Nickerson. There’s no place like home: UK-based financial analysts’ response to Dutch-English and British-English annual report texts. IEEE Transactions on Professional Communication, 54(1):1–17, March 2011. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

deGroot:2006:CAT

[dGKNG06] E. B. de Groot, H. Korzilius, C. Nickerson, and M. Gerritsen. A corpus analysis of text themes and photographic themes in managerial forewords of Dutch-English and British annual general reports. IEEE Transactions on Professional Communication, 49(3): 217–235, September 2006. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

De-George:1984:SRT

[DGORM84] James De-George, Gary O. Olson, Richard Ray, and Roger H. Muenier. Style and readability in technical writing. IEEE Transactions on Professional Communication, PC-27 (4):233, December 1984. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dysart-Gale:2010:IPW

[DGPR10] D. Dysart-Gale, K. Pitula, and T. Radhakrishnan. Improving professional writing for lay practitioners: a rhetorical approach. IEEE Transactions on Professional Communication, 53(3):293–303, September 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dysart-Gale:2011:CCI

[DGPR11] Deborah Dysart-Gale, Kristina Pitula, and Thiruvengadam Radhakrishnan. Culture, communication, and ICT for development: a Caribbean study. IEEE Transactions on Professional Communication, 54(1): 43–55, March 2011. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Deacon:1973:ERS

[DH73] Bill Deacon and David Hamilton. The editor’s role in the survival of the broad technical journal. IEEE Transactions on Professional Communication, PC-16(3):60–63, September 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dahl:1992:ISA

[DH92] R. D. Dahl and R. B. Horowitz.

International students and awareness of digital scanning issues. IEEE Transactions on Professional Communication, 35(3):172–175, September 1992. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dennett:1994:ITT

[DH94] J. Temple Dennett and M. Hseih.

Issue trees: a tool to aid the engineering writer. IEEE Transactions on Professional Communication, 37(2):88–96, June 1994. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Danchak:2004:DCR

[DH04] M. M. Danchak and M.-P.

Huguet. Designing for the changing role of the instructor in blended learning. IEEE Transactions on Professional Communication, 47(3):200– 210, September 2004. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Diamond:1977:HUT

[Dia77] Sidney A. Diamond. How to use a trademark in text. IEEE Transactions on Professional Communication, PC-20 (1):27–29, June 1977. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Diaz:2007:TWR

[Dia07] Charlsye Smith Diaz. The technical writer’s role in preserving intellectual property rights outside the United States tutorial. IEEE Transactions on Professional Communication, 50(2):121–129, June 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Diaz:2014:SWA

[Dia14] Charlsye Smith Diaz. Strategies for writing about innovation: Navigating the relationship between technical documentation, patent prosecution, and technology transfer. IEEE Transactions on Professional Communication, 57(2): 113–122, June 2014. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dierk:1972:SRS

[Die72] S. F. Dierk. The SMART retrieval system: Experiments in automatic document processing, Gerard Salton, Ed. (Englewood Cliffs, N.J.: Prentice– Hall, 1971, 556 pp., $15.00).

IEEE Transactions on Professional Communication, PC-15 (1):17, March 1972. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dieli:1989:UPW

[Die89] M. Dieli. The usability process:

working with iterative design principles. IEEE Transactions on Professional Communication, 32(4):272–278, December 1989. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

DIEHL:2004:AAD

[Die04] V. A. Diehl. Access to affordances, development of situation models, and identification of procedural text problems. IEEE Transactions on Professional Communication, 47(1): 54–64, March 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Diodato:1982:IAW

[Dio82]  Virgil Diodato.  Indexing:

Another way for authors to communicate. IEEE Transactions on Professional Communication, PC-25(3):120–122, September 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dunkle:1982:ISR

[DJ82] Susan B. Dunkle and Purvis M.

Jackson. Information structuring: Relating old and new knowledge. IEEE Transactions on Professional Communication, PC-25(4):175–177, December 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

deJong:2000:TDE

[dJS00] M. de Jong and P. J. Schellens. Toward a document evaluation methodology: what does research tell us about the validity and reliability of evaluation methods? IEEE Transactions on Professional Communication, 43(3):242–260, September 2000. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

DeFranco:2017:RAS

[DL17] Joanna F. DeFranco and Philip A. Laplante. Review and analysis of software development team communication research. IEEE Transactions on Professional Communication, 60(2):165–182, June 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dong:2024:ITC

[DL24] Lin Dong and Shuangyan Li. Integrating technical communication into China’s translation and interpreting curriculum: Course design, practice, and evaluation of two graduate classes. IEEE Transactions on Professional Communication, 67(3):369–385, September 2024. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Duong:2023:DCR

[DLVE23] Bao Duong, Jaeung Lee, Craig Van Slyke, and T. Selwyn Ellis. Distress coping responses among teleworkers. IEEE Transactions on Professional Communication, 66(3): 259–283, September 2023. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Demlinger:1963:DIR

[DM63] Marlin Demlinger and William F.

Marsden. Development of an information retrieval system for an electronics R&D laboratory. IEEE Transactions on Engineering Writing and Speech, EWS-6(1):10–19, September 1963. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Diehl:1981:MWI

[DM81] William Diehl and Larry Mikulecky. Making written information fit workers’ purposes. IEEE Transactions on Professional Communication, PC-24(1):5–9, March 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Doan:2019:CCF

[Doa19] Sara Doan. Contradictory comments: Feedback in professional communication service courses. IEEE Transactions on Professional Communication, 62(2):115–129, June 2019.

CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dobson:1975:TPT

[Dob75] David B. Dobson. Transcript of panel on technical composition — held at the International Technical Communications Conference, Anaheim, Calif., May 1975. IEEE Transactions on Professional Communication, PC-18(4):327–334, December 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dobrin:1985:DGA

[Dob85] David Dobrin. Do not grind armadillo armor in this mill. IEEE Transactions on Professional Communication, PC-28(4):30–37, December 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Docter:1961:TRE

[Doc61] Stewart Docter. Testing the readability of engineering writing. IRE Transactions on Engineering Writing and Speech, EWS-4(3):91–96, December 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Doebler:1969:CSC

[Doe69] Paul D. Doebler. The coming systems concept: How editors will use machines. IEEE Transactions on Engineering Writing and Speech, EWS-12(2):61– 63, May 1969. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Duffy:1996:JMC

[DOHM96] Thomas M. Duffy, Dick Osgood, Doug Holyoak, and David Monson. John M. Carroll, Ed., Scenario-Based Design: Envisioning Work and Technology in System Development [Book Review]. IEEE Transactions on Professional Communication, 39(4): 241, December 1996. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dombrowski:1991:LCI

[Dom91]  P. M. Dombrowski.  The

lessons of the Challenger investigations. IEEE Transactions on Professional Communication, 34(4):211–216, December 1991. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dombrowski:1995:CET

[Dom95] P. M. Dombrowski. Can ethics be technologized? Lessons from Challenger, philosophy, and rhetoric. IEEE Transactions on Professional Communication, 38(3):146–150, September 1995. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dombrowski:2007:EFE

[Dom07] Paul M. Dombrowski. The evolving face of ethics in technical and professional communication: Challenger to Columbia. IEEE Transactions on Professional Communication, 50(4):306–319, December 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dood:1979:UPC

[Doo79] Kendall J. Dood. The U.S. Patent classification system. IEEE Transactions on Professional Communication, PC-22 (2):95–100, June 1979. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dornfest:1963:BPT

[Dor63] A. Dornfest. Blueprint for the preparation of technical manuals for complex electronic equipment. IEEE Transactions on Engineering Writing and Speech, EWS-6(1):36–49, September 1963. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Dornbusch:1983:ARR

[Dor83] Joan F. Dornbusch. It’s all right to read your speech and here’s how. IEEE Transactions on Professional Communication, PC-26(1):25–29, March 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dossena:2006:DBN

[Dos06] M. Dossena. Doing business in Nineteenth-Century Scotland:

Expressing authority, conveying stance. IEEE Transactions on Professional Communication, 49(3):246–253, September 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Douglass:1972:CIU

[Dou72] Wade A. Douglass. Communicating industrial user and management needs to the new product engineering team: a PSYCHOM ’72 paper. IEEE Transactions on Professional Communication, PC-15(2):52– 54, June 1972. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Doumont:2000:WAA

[Dou00] J. Doumont. Worlds Apart:

Acting and Writing In Academic and Workplace Contexts [Book Review]. IEEE Transactions on Professional Communication, 43(2):226–227, June 2000. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Doumont:2001:MMG

[Dou01] J.-L. Doumont. Mary Munter, Guide to managerial communication: effective business writing and speaking, 5th ed. [Book Review]. IEEE Transactions on Professional Communication, 44(2):151–152, June 2001. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Doumont:2002:MNS

[Dou02a] J.-L. Doumont. Magical numbers: the seven-plus-or-minus-two myth. IEEE Transactions on Professional Communication, 45(2):123–127, June 2002. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Doumont:2002:TLP

[Dou02b] J.-L. Doumont. The three laws of professional communication. IEEE Transactions on Professional Communication, 45(4): 291–296, December 2002. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Davison:2017:EEG

[DPHF17] Robert M. Davison, Niki Panteli, Andrew M. Hardin, and Mark A. Fuller. Establishing effective global virtual student teams. IEEE Transactions on Professional Communication, 60(3):317–329, September 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

DAmbra:1994:MAS

[DR94] J. D’Ambra and R. E. Rice. Multimethod approaches for the study of computer-mediated communication, equivocality, and media selection. IEEE Transactions on Professional Communication, 37(4):231–239, December 1994. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dutta-Roy:2010:MOV

[DR10]  Amitava Dutta-Roy.  Mi-

crosoft Office Visio 2007: Inside Out (Walker, M. H.); Visio 2008 Bible (Biafore, B.); and Visualizing Information with Microsoft Office Visio 2007: Smart Diagrams for Business Users — reviews of “Microsoft Vizio 2007 Professional Edition” and “Microsoft Vizio 2007 Standard Edition” (Parker, D. J.) [Software Reviews]. IEEE Transactions on Professional Communication, 53(1):84–86, March 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dusenberry:2020:BPS

[DR20] Lisa Dusenberry and Joy Robinson. Building psychological safety through training interventions: Manage the team, not just the project. IEEE Transactions on Professional Communication, 63(3): 207–226, September 2020. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dresia:1975:PSP

[Dre75] David R. Dresia. Publication survey panel provides low cost reader feedback. IEEE Transactions on Professional Communication, PC-18(3):292–295, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dressel:1983:GQF

[Dre83] Susan Dressel. Generating a quick first draft. IEEE Transactions on Professional Communication, PC-26(4):172–174, December 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dressel:1987:SW

[Dre87]  Susan Dressel.  The spo-

ken word. IEEE Transactions on Professional Communication, PC-30(1):38, March 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dressel:1988:RAH

[Dre88]  S. Dressel.  Risks in an-

swering hypothetical questions. IEEE Transactions on Professional Communication, 31(1): 8–9, March 1988. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dreher:2016:PLE

[Dre16] Kira Dreher. Plain Language and Ethical Action: a Dialogic Approach to Technical Content in the Twenty-First Century [by Willerton, R.; Book Review]. IEEE Transactions on Professional Communication, 59(3):311–312, September 2016. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dreher:2017:IAP

[Dre17] Kira Dreher. Insider audiences and plain-language revision: a city charter case study. IEEE Transactions on Professional Communication, 60(4): 430–447, December 2017. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Delinchant:2002:CDA

[DRG^{+}02] B. Delinchant, V. Riboulet, L. Gerbaud, P. Marin, F. Noel, and F. Wurtz. E-cooperative design among mechanical and electrical engineers: implications for communication between professional cultures. IEEE Transactions on Professional Communication, 45(4): 231–249, December 2002. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Driver:1987:CMD

[Dri87] Russell W. Driver. A communication model for determining the appropriateness of on-product warnings. IEEE Transactions on Professional Communication, PC-30(3): 157–163, September 1987. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Drossel:1980:OPT

[Dro80] Margaret Drossel. Organizing a project team. IEEE Transactions on Professional Communication, PC-23(3):148–150, September 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Drury:1985:ER

[Dru85] Alinda Drury. Evaluating readability. IEEE Transactions on Professional Communication, PC-28(4):11–14, December 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Delisle:1989:CWH

[DS89] N. M. Delisle and M. D. Schwartz. Collaborative writing with hypertext. IEEE Transactions on Professional Communication, 32(3):183– 188, September 1989. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Drott:1978:MUM

[DSC78] M. Carl Drott, K. Subramanyam, and Harold W. Cheney. Multiple use of machine readable data in publishing. IEEE Transactions on Professional Communication, PC-21 (4):145–149, December 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

DeVolpi:1980:TPP

[DSF^{+}80] A. DeVolpi, G. S. Stanford, C. L. Fink, E. A. Rhodes, and M. R. Fenrick. Text processing for the professional staff. IEEE Transactions on Professional Communication, PC-23 (4):164–167, December 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Downey:2010:HSE

[DSV10] Ryan Downey, Masanori Suzuki, and Alistair Van Moere. Highstakes English-language assessments for aviation professionals: Supporting the use of a fully automated test of spoken-language proficiency. IEEE Transactions on Professional Communication, 53 (1):18–32, March 2010. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

DuBay:2008:ULC

[DuB08] William H. DuBay. Unlocking language: The classic readability studies. IEEE Transactions on Professional Communication, 51(4):416–417, December 2008. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dudley:1961:CIT

[Dud61a] B. Dudley. A comment on information theory! IRE Transactions on Engineering Writing and Speech, EWS-4(1):25– 26, January 1961. ISSN 0893-7869 (print), 2331-3692 (electronic). See comment [Kel61a].

Dudley:1961:HRVa

[Dud61b] Beverly Dudley. Human responses — a vital link in communications progress. Part I.

IRE Transactions on Engineering Writing and Speech, EWS-4(1):5–13, January 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Dudley:1961:HRVb

[Dud61c] Beverly Dudley. Human responses — a vital link in communications progress. Part II. IRE Transactions on Engineering Writing and Speech, EWS-4(2):35–41, May 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Dudley:1962:CSS

[Dud62] B. Dudley. Communicating the significance of scientific research. IRE Transactions on Engineering Writing and Speech, EWS-5(2):65–70, December 1962. ISSN 0893-7869 (print), 2331-3692 (electronic).

Dudley:1971:WSE

[Dud71] Juanita Williams Dudley. Writing skills of engineering and science students. IEEE Transactions on Engineering Writing and Speech, EWS-14 (2):42–46, June 1971. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Duffy:1987:EES

[Duf87] Thomas M. Duffy. Engineering effective software. IEEE Transactions on Professional Communication, PC-30(1):41– 42, March 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Duin:1996:CIC

[Dui96] A. H. Duin. Collective Intelligence in Computer-Based Collaboration [Book Review]. IEEE Transactions on Professional Communication, 39(3): 167, September 1996. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dukes:1989:TCD

[Duk89]  E. Dukes.  Table con-

struction: do’s and don’ts. IEEE Transactions on Professional Communication, 32 (1):36–40, March 1989. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dulek:1980:WUR

[Dul80] Ron Dulek. Writing to unidentified readers. IEEE Transactions on Professional Communication, PC-23(3):125–127, September 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dulek:1992:CYC

[Dul92] R. E. Dulek. Could you be clearer? An examination of the multiple perspectives of clarity. IEEE Transactions on Professional Communication, 35 (2):84–87, June 1992. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dulye:1993:TBT

[Dul93] L. M. Dulye. Toward better two-way: why communications process improvement represents the right response during uncertain times. IEEE Transactions on Professional Communication, 36 (1):24–29, March 1993. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dumont:1996:TLC

[Dum96] R. A. Dumont. Teaching and learning in cyberspace. IEEE Transactions on Professional Communication, 39(4): 192–204, December 1996. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dunn:1981:GWP

[Dun81] George R. Dunn. Guidelines for writing for publication. IEEE Transactions on Professional Communication, PC-24 (4):172–175, December 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dunwoody:1982:QA

[Dun82] Sharon Dunwoody. A question of accuracy. IEEE Transactions on Professional Communication, PC-25(4):196–199, December 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Durham:1999:WSL

[Dur99] M. Durham. Web sites for the language-impaired. IEEE Transactions on Professional Communication, 42(3):175– 177, September 1999. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Durazzi:2022:LJC

[Dur22] Allison Durazzi. Linguistic Justice on Campus: Pedagogy and Advocacy for Multilingual Students: Brooke R. Schreiber, Eunjeong Lee, Jennifer T. Johnson, and Norah Fahim: [Book Review]. IEEE Transactions on Professional Communication, 65(1): 258–260, March 2022. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Doumont:2002:CRG

[DV02] J.-L. Doumont and P. Vandenbroeck. Choosing the right graph. IEEE Transactions on Professional Communication, 45(1):1–6, March 2002. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Davidson:2009:TTI

[DV09] Elizabeth Davidson and Emmanuelle Vaast. Tech talk: an investigation of blogging in technology innovation discourse. IEEE Transactions on Professional Communication, 52(1):40–60, March 2009.

CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Du:2007:LWE

[DW07] Helen S. Du and Christian Wagner. Learning with Weblogs: Enhancing cognitive and social knowledge construction. IEEE Transactions on Professional Communication, 50(1): 1–16, March 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dworkin:1969:TCM

[Dwo69] Solomon Dworkin. Technical communications with motion pictures. IEEE Transactions on Engineering Writing and Speech, EWS-12(2):39–42, May 1969. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Duffy:1985:DUT

[DWP85] Thomas M. Duffy, Robert Waller, and Angelina Piccoli. Designing usable texts. IEEE Transactions on Professional Communication, PC-28(4):47– 49, December 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Dyrud:1998:TCG

[Dyr98] D. L. Dyrud. Technical Communication In The Global Community [Book Review]. IEEE Transactions on Professional Communication, 41(3): 226–227, September 1998. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Evans:1984:IAS

[EA84] Eston E. Evans and Charles K.

Alexander. Increasing aural skills of international graduate students. IEEE Transactions on Professional Communication, PC-27(3):168–171, September 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ealey:1983:PST

[Eal83] Thomas Ealey. Presenting the successful technical seminar. IEEE Transactions on Professional Communication, PC-26 (1):35–37, March 1983. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Eito-Brun:2018:AQR

[EBAS18] Ricardo Eito-Brun and Antonio Amescua-Seco. Automation of quality reports in the aerospace industry. IEEE Transactions on Professional Communication, 61(2):166– 177, June 2018. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Eicher-Catt:2000:NPC

[EC00]  D. Eicher-Catt.  Narrative

and professional communication [Book Review]. IEEE Transactions on Professional Communication, 43(4):409– 411, December 2000. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Eicher-Catt:2001:WPL

[EC01] D. Eicher-Catt. Writing a professional life: stories of technical communicators on and off the job [Book Review]. IEEE Transactions on Professional Communication, 44(4): 300–301, December 2001. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ecker:2003:TWS

[Eck03] P. S. Ecker. The tech writer’s survival guide: a comprehensive handbook for aspiring technical writers [Book Review]. IEEE Transactions on Professional Communication, 46(3): 251–253, September 2003. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Eddington:1983:TMP

[Edd83] Ruth Eddington. Tips for more precise language use. IEEE Transactions on Professional Communication, PC-26 (4):201–203, December 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Eden:1961:PB

[Ede61] Brian F. Eden. Preparing a bibliography. IRE Transactions on Engineering Writing and Speech, EWS-4(3):97–100, December 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Edelman:1977:PLC

[Ede77] Hendrik Edelman. Periodical literature in a changing library environment. IEEE Transactions on Professional Communication, PC-20(2):70– 72, September 1977. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Eden:1987:BRL

[Ede87] Rick A. Eden. Book reviews: Lee Odell and Dixie Goswami, Writing in nonacademic settings. IEEE Transactions on Professional Communication, PC-30(2):112–113, June 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Engin:1978:LGE

[EE78] Ann W. Engin and Ali E. Engin. The lecture: Greater effectiveness for a familiar method. IEEE Transactions on Professional Communication, PC-21 (1):8–11, March 1978. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Egan:1981:WRC

[Ega81]  Christine Egan.  Writing

resumes and cover letters. IEEE Transactions on Professional Communication, PC-24 (4):156–160, December 1981.

CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Evans:1984:MRP

[EL84] Joyce M. Evans and Roseann Litzinger. Monthly report pyramid. IEEE Transactions on Professional Communication, PC-27(2):82–84, June 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Elling:2012:CCT

[ELdJ12] Sanne Elling, Leo Lentz, and Menno de Jong. Combining concurrent think-aloud protocols and eye-tracking observations: an analysis of verbalizations and silences. IEEE Transactions on Professional Communication, 55(3):206– 220, September 2012. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Estrin:1978:EOP

[EM78] Herman A. Estrin and Edward J. Monahan. Effective oral presentation of scientific and technical information. IEEE Transactions on Professional Communication, PC-21 (2):49–53, June 1978. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Eklundh:1994:UQP

[EM94] K. S. Eklundh and C. MacDonald. The use of quoting to preserve context in electronic mail dialogues. IEEE Transactions on Professional Communication, 37(4):197–202, December 1994. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Emig:1999:DWR

[Emi99] I. W. Emig. The Dynamics of Writing Review: Opportunities For Growth and Change In The Workplace [Book Review]. IEEE Transactions on Professional Communication, 42(4): 311–312, December 1999. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Engelbret:1972:SBW

[Eng72] David Engelbret. Storyboarding a better way of planning and writing proposals. IEEE Transactions on Professional Communication, PC-15 (4):115–118, December 1972. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Engle:1989:TWC

[Eng89]  C. L. Engle.  A techni-

cal writing curriculum that brings art and science together. IEEE Transactions on Professional Communication, 32(3): 159–161, September 1989. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Echols:1983:STW

[EP83] Carla Echols and Lois Pryor. A scorecard for technical writing. IEEE Transactions on Professional Communication, PC-26 (4):162–165, December 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Erickson:2011:ATP

[Eri11]  Joe Erickson.  Assessment

in Technical and Professional Communication (Hundleby, M. and Allen, J., Eds.; 2010) [Book Review]. IEEE Transactions on Professional Communication, 54(1):97–99, March 2011. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ervin:1999:MSM

[Erv99a] N. Ervin. Mary Sue Macnealy, Strategies For Empirical Research In Writing [Book Review]. IEEE Transactions on Professional Communication, 42(1):64–66, March 1999. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ervin:1999:PCD

[Erv99b] N. Ervin. Profile of Charles D. Haynes. IEEE Transactions on Professional Communication, 42(3):185–187, September 1999. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ervin:1999:PNC

[Erv99c] N. Ervin. Profile of Normand Croteau. IEEE Transactions on Professional Communication, 42(4):309–310, December 1999. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

El-Shinnawy:1998:SKR

[ES98] M. El-Shinnawy. Susan T.

Kinney and Richard T. Watson, The effect of medium and task on dyadic communication. IEEE Transactions on Professional Communication, 41(2): 140–142, June 1998. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

El-Shinnawy:1999:ISI

[ES99] M. El-Shinnawy. Introduction to the special issue: communication in virtual organizations. IEEE Transactions on Professional Communication, 42(4): 213–214, December 1999. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Eselgroth:1969:ATV

[Ese69] Theodore C. Eselgroth. Audio tapes for versatility and low cost. IEEE Transactions on Engineering Writing and Speech, EWS-12(2):42–46, May 1969. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

El-Shinnawy:1998:ACM

[ESM98] M. El-Shinnawy and M. L. Markus. Acceptance of communication media in organizations: richness or features? IEEE Transactions on Professional Communication, 41(4): 242–253, December 1998. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Evia:2015:TSA

[ESPQ15] Carlos Evia, Matthew R. Sharp, and Manuel A. Pérez-Quiñones. Teaching structured authoring and DITA through rhetorical and computational thinking. IEEE Transactions on Professional Communication, 58(3):328–343, September 2015. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Estrin:1982:HWS

[Est82]  Herman A. Estrin.  How

to write for scientific and technical journals. IEEE Transactions on Professional Communication, PC-25(1):32– 33, March 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Etz:1963:SC

[Etz63] Donald V. Etz. The structure of communication. IEEE Transactions on Engineering Writing and Speech, EWS-6(2): 69–71, December 1963. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Evans:1960:MEE

[Eva60] Frank W. Evans. More effective engineering proposals — one key to success. IRE Transactions on Engineering Writing and Speech, EWS-3(2):54– 58, July 1960. ISSN 0893-7869 (print), 2331-3692 (electronic).

Evans:2000:CDR

[Eva00]  M. B. Evans.  Challenges

in developing research-based Web design guidelines. IEEE Transactions on Professional Communication, 43(3):302– 312, September 2000. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Evans:2002:UTR

[Eva02] R. T. Evans. Usability testing and research. IEEE Transactions on Professional Communication, 45(2):151–152, June 2002. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

EVIA:2004:QQT

[Evi04] C. Evia. Quality over quantity: a two-step model for reinforcing user feedback in transnational Web-based systems through participatory design. IEEE Transactions on Professional Communication, 47(1):71–74, March 2004. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

English-Zemke:1988:UCO

[EZ88]  P. English-Zemke.  Using

color in online marketing tools. IEEE Transactions on Professional Communication, 31 (2):70–74, June 1988. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Farrell:1962:CBI

[Far62] Thomas Farrell. Communication in business and industry [book review]. IRE Transactions on Engineering Writing and Speech, EWS-5(1):31, August 1962. ISSN 0893-7869 (print), 2331-3692 (electronic).

Farkas:1964:WBT

[Far64] L. L. Farkas. Writing better technical papers. IEEE Transactions on Engineering Writing and Speech, EWS-7(3):31– 34, December 1964. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Fisher:2003:CMP

[FBE^{+}03] D. Fisher, T. Bowers, A. Ellerton, T. J. Brumm, and S. K. Mickelson. As the case may be: the potential of electronic cases for interdisciplinary communication instruction. IEEE Transactions on Professional Communication, 46(4):313– 319, December 2003. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Franz:1979:GHB

[FC79] Warren L. Franz and John S.

Child. Good habits before filing a patent application. IEEE Transactions on Professional Communication, PC-22(2):86– 88, June 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Fisher:1996:IUO

[FC96] J. Fisher and J. Chong. Improving the usability of online information when translated from English to Chinese. IEEE Transactions on Professional Communication, 39(3): 122–128, September 1996. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ford:2024:IPP

[FC24] Julie Dyke Ford and Destiny Crawford. Integrating professional preparedness ePortfolios within an undergraduate engineering curriculum. IEEE Transactions on Professional Communication, 67(2): 246–259, June 2024. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Flammia:2010:LRS

[FCS10] Madelyn Flammia, Yvonne Cleary, and Darina M. Slattery. Leadership roles, socioemotional communication strategies, and technology use of Irish and US students in virtual teams. IEEE Transactions on Professional Communication, 53(2):89–101, June 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Firman:1984:ITP

[FD84] Anthony H. Firman and L. W. Denton. An introduction to technical publishing. IEEE Transactions on Professional Communication, PC-27(2): 109, June 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Feng:2018:MPD

[FDJW18] Jieyun Feng, Luyao Dai, Jinlin Jiang, and Rining Wei. A matter of perspective: a discursive analysis of the perceptions of three stakeholders of the Mutianyu Great Wall. IEEE Transactions on Professional Communication, 61(1): 22–47, March 2018. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Featheringham:1977:CCH

[Fea77] Tom R. Featheringham. Computerized conferencing and human communication. IEEE Transactions on Professional Communication, PC-20(4): 207–213, December 1977. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Feenberg:1986:NDO

[Fee86] Andrew Feenberg. Network design: an operating manual for computer conferencing. IEEE Transactions on Professional Communication, PC-29 (1):2–7, March 1986. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Feistman:1959:EW

[Fei59] M. L. Feistman. Engineering writing “Up or down”. IRE Transactions on Engineering Writing and Speech, EWS-2(1): 22–23, January 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Feineman:1979:HWG

[Fei79] George Feineman. How to write good like a logistician should. IEEE Transactions on Professional Communication, PC-22(4):200–201, December 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ferris:2003:YDC

[Fer03] T. L. J. Ferris. Yvonne Dittrich, Christiane Floyd, and Ralf Klischewski, Eds. Social thinking-software practice. IEEE Transactions on Professional Communication, 46(4): 337–339, December 2003. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ferris:2004:IDC

[Fer04a]  T. L. J. Ferris.  Interac-

tion design for complex problem solving: Developing useful and usable software. IEEE Transactions on Professional Communication, 47(3):218– 219, September 2004. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ferris:2004:TGT

[Fer04b] T. L. J. Ferris. Tracing genres through organizations: a sociocultural approach to information. IEEE Transactions on Professional Communication, 47(4):329–330, December 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ferris:2004:UCD

[Fer04c] T. L. J. Ferris. User-centered design: an integrated approach. IEEE Transactions on Professional Communication, 47(1):75–77, March 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ferris:2005:PID

[Fer05a] T. L. J. Ferris. Participatory IT design: Designing for business and workplace realities. IEEE Transactions on Professional Communication, 48(3): 327–328, September 2005. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ferris:2005:PMP

[Fer05b] T. L. J. Ferris. Presentation media for product interaction. IEEE Transactions on Professional Communication, 48(2): 230–231, June 2005. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ferris:2006:WLP

[Fer06] Timothy L. J. Ferris. Wolfgang Lefèvre, Picturing Machines 1400–1700. IEEE Transactions on Professional Communication, 49(4):365–367, December 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ferris:2007:DSM

[Fer07a] Timothy L. J. Ferris. Discursive Strategies in Multicultural Business Meetings (Poncini, G.; 2004). IEEE Transactions on Professional Communication, 50(2):177–179, June 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ferris:2007:LRC

[Fer07b] Timothy L. J. Ferris. Listening and Responding (Collins, S. D.) [Book Review]. IEEE Transactions on Professional Communication, 50(4):379– 380, December 2007. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ferris:2007:MVT

[Fer07c] Timothy L. J. Ferris. Managing Virtual Teams: Getting the Most From Wikis, Blogs, and Other Collaborative Tools [Book review; Brown, M. K. et al.; 2007]. IEEE Transactions on Professional Communication, 50(3):263–265, September 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ferris:2007:PPP

[Fer07d] Timothy L. J. Ferris. Personal, Portable, Pedestrian: Mobile Phones in Japanese Life (Ito, M. et al., Eds.; 2005). IEEE Transactions on Professional Communication, 50(1): 71–73, March 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ferris:2008:CPW

[Fer08] Timothy L. J. Ferris. Communicative Practices in Workplaces and the Professions: Cultural Perspectives on the Regulation of Discourse and Organizations (Zachry, M. and Thralls, C.) [Book Review]. IEEE Transactions on Professional Communication, 51(2): 232–234, June 2008. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ferris:2009:HKD

[Fer09a] Timothy L. J. Ferris. Harnessing Knowledge Dynamics: Principled Organizational Knowing and Learning (Nissen, M. E.; 2006) [Book Review]. IEEE Transactions on Professional Communication, 52(2):215–216, June 2009. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ferris:2009:IIS

[Fer09b] Timothy L. J. Ferris. Inside Intuition (Sadler-Smith, E.; 2008) [Book Review]. IEEE Transactions on Professional Communication, 52(1):115– 116, March 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ferris:2010:MMS

[Fer10] Timothy L. J. Ferris. Motives for Metaphor in Scientific and Technical Communication, (Giles, T. D.; 2008) [Book Review]. IEEE Transactions on Professional Communication, 53(2):179–180, June 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ferris:2012:SWC

[Fer12] Timothy L. J. Ferris. Scientific Writing and Communication Papers, Proposals, and Presentations (Hofmann, A. H.; 2010) [Book Review]. IEEE Transactions on Professional Communication, 55(3): 289–291, September 2012. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Fest:1962:PSF

[Fes62] Thorrel B. Fest. Personal and structural factors in the communication of technical information. IRE Transactions on Engineering Writing and Speech, EWS-5(2):55–60, December 1962. ISSN 0893-7869 (print), 2331-3692 (electronic).

Feyerherm:1988:EEE

[Fey88] J. Feyerherm. Errors, editors, and English teachers. IEEE Transactions on Professional Communication, 31(1): 10–12, March 1988. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Feinberg:1984:TWA

[FG84] Susan Feinberg and Jerry I. Goldman. Technical writing attitude measurement and instructional goals. IEEE Transactions on Professional Communication, PC-27(3):155–160, September 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Farkas:1987:JLY

[FHR87a] David Farkas, Mark Haselkorn, and Judith Ramey. In just the last 10 years. IEEE Transactions on Professional Communication, PC-30(1):43– 44, March 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Farkas:1987:DPA

[FHR87b] David K. Farkas, Mark Haselkorn, and Judith Ramey. Desktop publishing: The author as compositor. IEEE Transactions on Professional Communication, PC-30(2):101, June 1987. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Fields:1965:SOT

[Fie65] C. W. Fields. The staffing and operation of a technical papers support group. IEEE Transactions on Engineering Writing and Speech, EWS-8(2):55–60, December 1965. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Field:1967:BR

[Fie67] P. A. Field. Book reviews.

IEEE Transactions on Engineering Writing and Speech, EWS-10(1):23, July 1967. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Fields:1968:GGI

[Fie68] C. W. Fields. Guide to gathering information in face-to-face interviews. IEEE Transactions on Engineering Writing and Speech, EWS-11(3): 168, December 1968. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Fields:1977:TAT

[Fie77] Craig Fields. Terminal access technology of the 1990’s. IEEE Transactions on Professional Communication, PC-20 (1):2–6, June 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Field:1983:GAM

[Fie83] Janet N. Field. Graphic arts manual. IEEE Transactions on Professional Communication, PC-26(4):212, December 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Filley:1982:ODC

[Fil82] Richard D. Filley. Opening the door to communication through graphics. IEEE Transactions on Professional Communication, PC-25(2):91– 94, June 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Filimon:2005:RJB

[Fil05] Stephany Filimon. R. John Brockmann, Twisted Rails, Sunken Ships: The Rhetoric of Nineteenth Century Steamboat and Railroad Accident Reports, 1833–1879. IEEE Transactions on Professional Communication, 48(4):432–433, December 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Filimon:2006:MJA

[Fil06] Stephany Filimon. Michael J. Albers, Communication of Complex Information: User Goals and Information Needs for Dynamic Web Information.

IEEE Transactions on Professional Communication, 49(1): 82–84, March 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Fischborn:1984:DL

[Fis84] Herb Fischborn. In defense of the language. IEEE Transactions on Professional Communication, PC-27(1):7–9, March 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Fisk:1992:PPP

[Fis92] M. L. Fisk. People, proxemics, and possibilities for technical writing. IEEE Transactions on Professional Communication, 35(3):176–182, September 1992. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Fisher:1998:DRT

[Fis98] J. Fisher. Defining the role of a technical communicator in the development of information systems. IEEE Transactions on Professional Communication, 41(3):186–199, September 1998. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Fisher:1999:VTC

[Fis99] J. Fisher. The value of the technical communicator’s role in the development of information systems. IEEE Transactions on Professional Communication, 42(3):145–155, September 1999. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Friess:2023:LSU

[FL23] Erin Friess and Christina Liles. Longitudinal study of usability and user experience in technical and professional communication research. IEEE Transactions on Professional Communication, 66(4):338– 355, December 2023. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Flaherty:1985:OE

[Fla85] Deborah L. Flaherty. The Olympic experience. IEEE Transactions on Professional Communication, PC-28(2):21– 23, June 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Flammia:1993:CGT

[Fla93] M. Flammia. The challenge of getting technical experts to talk: why interviewing skills are crucial to the technical communication curriculum. IEEE Transactions on Professional Communication, 36(3): 124–129, September 1993. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Flammia:2005:PTC

[Fla05] M. Flammia. Preparing technical communication students to play a role on the translation team. IEEE Transactions on Professional Communication, 48(4):401–412, December 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Fleischhauer:1985:TWM

[Fle85] F. Wm. Fleischhauer. Technical writing in the managementsystems world. IEEE Transactions on Professional Communication, PC-28(3):31–34, September 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Flores:1973:TUM

[Flo73] Ivan Flores. Tutorial use of multiple type faces. IEEE Transactions on Professional Communication, PC-16(1):2– 6, March 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Florentine:1975:SF

[Flo75] Harry Florentine. Subscription fulfillment. IEEE Transactions on Professional Communication, PC-18(3):145–149, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:1980:PF

[Flo80] Raymond E. Floyd. Presentation fundamentals. IEEE Transactions on Professional Communication, PC-23(1):40– 42, March 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:1981:WS

[Flo81] Raymond E. Floyd. The written speech. IEEE Transactions on Professional Communication, PC-24(4):195–196, December 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2000:PWA

[Flo00a] R. E. Floyd. Proposal writing: the art of friendly and winning persuasion [Book Review]. IEEE Transactions on Professional Communication, 43(4): 420–421, December 2000. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2000:TRW

[Flo00b] R. E. Floyd. Technical report writing today, 7th edition [Book Review]. IEEE Transactions on Professional Communication, 43(3):339– 340, September 2000. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2000:BWC

[Flo00c]  R. E. Floyd.  The busi-

ness writer’s companion, second edition [Book Review]. IEEE Transactions on Professional Communication, 43(2): 222–223, June 2000. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2001:ACT

[Flo01a] R. E. Floyd. Ava Cross, Talking business-strategies for successful presentations [Book Review]. IEEE Transactions on Professional Communication, 44(2):153–154, June 2001. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2001:DQT

[Flo01b]  R. E. Floyd.  Developing

quality technical information — a handbook for writers and editors [Book Review]. IEEE Transactions on Professional Communication, 44(3): 212–213, September 2001. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2001:DTB

[Flo01c] R. E. Floyd. The Discipline of Teamwork [Book Review]. IEEE Transactions on Professional Communication, 44(1): 71–72, March 2001. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2001:TCH

[Flo01d]  R. E. Floyd.  The tech-

nical communicator’s handbook [Book Review]. IEEE Transactions on Professional Communication, 44(4):296– 297, December 2001. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2002:HWW

[Flo02a] R. E. Floyd. How to write for the world of work — 6th ed. [Book Review]. IEEE Transactions on Professional Communication, 45(4):309– 310, December 2002. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2002:TPW

[Flo02b] R. E. Floyd. Technical and professional writing — solving problems at work [Book Review]. IEEE Transactions on Professional Communication, 45(3):209–210, September 2002. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2002:TWP

[Flo02c] R. E. Floyd. Technical Writing — a practical approach (4th Ed.) — Book Review. IEEE Transactions on Professional Communication, 45(1): 65–66, March 2002. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2002:TCA

[Flo02d] R. E. Floyd. Technical communication and its applications, 2nd ed. IEEE Transactions on Professional Communication, 45(2):145–146, June 2002. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2003:TDD

[Flo03a] R. E. Floyd. Text databases and document management: Theory and practice [Book Review]. IEEE Transactions on Professional Communication, 46(2):147–148, June 2003. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2003:DKV

[Flo03b] R. E. Floyd. Douglas K.

Van Duyne, James A. Landay, and Jason I. Hong, The design of sites [Book Review]. IEEE Transactions on Professional Communication, 46(4): 335–336, December 2003. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2003:MRS

[Flo03c] R. E. Floyd. Maris Roze and Simon Maxwell, Technical communication in the age of the Internet — fourth ed. [Book Review]. IEEE Transactions on Professional Communication, 46(4):342– 343, December 2003. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2003:DGR

[Flo03d] Raymond Floyd. Daniel G. Riordan and Steven E. Pauley, Technical Report Writing Today — 8th ed. [Book Review]. IEEE Transactions on Professional Communication, 46(1): 75–76, March 2003. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2004:SRG

[Flo04a] R. Floyd. The short road to great presentations. IEEE Transactions on Professional Communication, 47(2):150– 151, June 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2004:ALW

[Flo04b] R. E. Floyd. The accidental leader: What to do when you’re suddenly in charge. IEEE Transactions on Professional Communication, 47(4): 333–334, December 2004. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2004:PCC

[Flo04c] R. E. Floyd. Professional communication: The corporate insider’s approach. IEEE Transactions on Professional Communication, 47(3):220– 221, September 2004. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2004:RRT

[Flo04d] R. E. Floyd. The rhetoric of risk technical documentation in hazardous environments.

IEEE Transactions on Professional Communication, 47(3): 222–223, September 2004. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2005:TRW

[Flo05a] R. E. Floyd. Technical Report Writing Today, 9th Edition. IEEE Transactions on Professional Communication, 48(4): 434–435, December 2005. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2005:IAT

[Flo05b] R. E. Floyd. Innovative approaches to teaching technical communication. IEEE Transactions on Professional Communication, 48(3):334– 336, September 2005. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2005:IPL

[Flo05c] R. E. Floyd. Intellectual property law for engineers and scientists. IEEE Transactions on Professional Communication, 48(1):113–114, March 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2005:LPD

[Flo05d] R. E. Floyd. A legal primer for the digital age. IEEE Transactions on Professional Communication, 48(2):228–229, June 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2005:NMP

[Flo05e] R. E. Floyd. Negotiating moves: Problem presentation and resolution in Japanese business discourse. IEEE Transactions on Professional Communication, 48(3):339– 340, September 2005. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2006:SBT

[Flo06a] R. E. Floyd. Strategies for Business and Technical Writing, 5th Edition. IEEE Transactions on Professional Communication, 49(3):293– 294, September 2006. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2006:EPC

[Flo06b] R. E. Floyd. Effective professional communication. IEEE Transactions on Professional Communication, 49(2):208– 210, June 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2006:TDE

[Flo06c] R. E. Floyd. Training and development: Enhancing communication and leadership skills. IEEE Transactions on Professional Communication, 49(1): 87–88, March 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2006:LJG

[Flo06d] Raymond E. Floyd. Laura J. Gurak and John M. Lannon, A Concise Guide to Technical Communication — 2nd Edition [Book Review]. IEEE Transactions on Professional Communication, 49(4): 373–374, December 2006. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2007:GBE

[Flo07a] Raymond E. Floyd. Guide to Business Etiquette [Book review; Cook, R. A. et al.; 2005]. IEEE Transactions on Professional Communication, 50(3): 266–267, September 2007. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2007:TCI

[Flo07b] Raymond E. Floyd. Technical Communication International: Today and in the Future (Hennig, J. and Tjarks-Sobhani, M., Eds.; 2005). IEEE Transactions on Professional Communication, 50(1):76–77, March 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2007:TWP

[Flo07c] Raymond E. Floyd. Technical Writing: Principles, Strategies, and Reading, 6th Edition [Book review; Reep, D. C.; 2006]. IEEE Transactions on Professional Communication, 50(3):268–269, September 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2007:TBY

[Flo07d] Raymond E. Floyd. Think Before You Write (Ingle, P.; 2006). IEEE Transactions on Professional Communication, 50(2):175–176, June 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2007:TMB

[Flo07e] Raymond E. Floyd. Throw Me the Bottom Line I’m Drowning in E-Mail! (Fielden, J. S. et al.; 2003). IEEE Transactions on Professional Communication, 50(1):74–75, March 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2007:EGF

[Flo07f] Raymond E. Floyd. Erin Gruwell and the Freedom Writers, The Freedom Writers Diary [Book Review]. IEEE Transactions on Professional Communication, 50(4):381– 382, December 2007. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2008:PCT

[Flo08]  Raymond E. Floyd.  The

Professional Communications Toolkit (Whalen, D. J. et al.) [Book Review]. IEEE Transactions on Professional Communication, 51(2):230–231, June 2008. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2010:SWT

[Flo10] Raymond E. Floyd. Speak to Win (Tracy, B.; 2008) [Book Review]. IEEE Transactions on Professional Communication, 53(2):183–184, June 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2018:FTS

[Flo18] Raymond Floyd. The Forgotten Tribe: Scientists as Writers (Emerson, L.) [Book Review]. IEEE Transactions on Professional Communication, 61(3):346–347, September 2018. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Floyd:2020:EJT

[Flo20] Raymond Floyd. Engineering Justice: Transforming Engineering Education and Practice: Jon A. Leydens and Juan C. Lucena [Book Review]. IEEE Transactions on Professional Communication, 63(1): 96–97, March 2020. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Fillenwarth:2018:QED

[FMB18] Gracemarie Mike Fillenwarth, Mary McCall, and Catherine Berdanier. Quantification of engineering disciplinary discourse in résumés: a novel genre analysis with teaching implications. IEEE Transactions on Professional Communication, 61(1):48–64, March 2018. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Fuller:2012:EDV

[FMT12] Christie M. Fuller, Kent Marett, and Douglas P. Twitchell. An examination of deception in virtual teams: Effects of deception on task performance, mutuality, and trust. IEEE Transactions on Professional Communication, 55 (1):20–35, March 2012. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ford:2006:TRL

[FO06] J. D. Ford and L. Ortiz. Teaching the role of language in managing organizational change: the Sunwest anti-union campaign case. IEEE Transactions on Professional Communication, 49(2):176–188, June 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Fong:1961:RRM

[Fon61] Irving J. Fong. The rightful role of management in technical communications. IRE Transactions on Engineering Writing and Speech, EWS-4(1): 14–16, January 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Ford:2004:KTA

[For04] J. D. Ford. Knowledge transfer across disciplines: tracking rhetorical strategies from a technical communication classroom to an engineering classroom. IEEE Transactions on Professional Communication, 47(4):301–315, December 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Fountain:1959:WEE

[Fou59] A. M. Fountain. Working with electrical engineers in seminar. IRE Transactions on Engineering Writing and Speech, EWS-2(2):46–48, June 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Fox:1981:CL

[Fox81] Marcia R. Fox. The cover letter. IEEE Transactions on Professional Communication, PC-24(4):163–165, December 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Foy:1996:RIC

[Foy96] P. S. Foy. The re-invention of the corporate information model. IEEE Transactions on Professional Communication, 39(1):23–29, March 1996. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Friday:1979:LBA

[FP79] J. Friday and G. Porter. Lawrence Braqq’s advice to a lecturer. IEEE Transactions on Professional Communication, PC-22(1):4–8, March 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Farkas:1995:OEM

[FP95] D. K. Farkas and S. E. Poltrock. Online editing, markup models, and the workplace lives of editors and writers. IEEE Transactions on Professional Communication, 38 (2):110–117, June 1995. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ford:2021:NET

[FPKS^{+}21] Julie Dyke Ford, Marie Paretti, Daria Kotys-Schwartz, Susannah Howe, and Robin Ott. New engineers transfer of communication activities from school to work. IEEE Transactions on Professional Communication, 64(2):105–120, June 2021. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Fowler:1986:PWC

[FR86] Susan L. Fowler and David Roeger. Programmer and writer collaboration: Making user manuals that work. IEEE Transactions on Professional Communication, PC-29(4):21– 25, December 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Fang:2013:RNC

[FR13] Xiang Fang and T. M. Rajkumar. The role of national culture and multimedia on first impression bias reduction: an experimental study in US and China. IEEE Transactions on Professional Communication, 56(4):354–371, December 2013. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Fraser:1975:PVL

[Fra75] Walter J. Fraser. Publishers versus libraries: Some hidden dimensions in the current debate. IEEE Transactions on Professional Communication, PC-18(3):200–206, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Frase:1981:EIM

[Fra81] Lawrence T. Frase. Ethics of imperfect measures. IEEE Transactions on Professional Communication, PC-24(1):48– 50, March 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Franke:1989:VRT

[Fra89] E. A. Franke. The value of the retrievable technical memorandum system to an engineering company. IEEE Transactions on Professional Communication, 32(1):12–16, March 1989. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Freeman:1968:ATS

[Fre68] Donald E. Freeman. Administrative Terminal System for IBM System/360. IEEE Transactions on Engineering Writing and Speech, EWS-11(3): 152–156, December 1968. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Freeman:1980:EBC

[Fre80a] Diann E. Freeman. Effective business communications, 3rd ed., Herta A. Murphy and Charles E. Peck. IEEE Transactions on Professional Communication, PC-23(4):200–201, December 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Freeman:1980:ERR

[Fre80b] Diann E. Freeman. Effective research and report writing in government, Judson Monroe. IEEE Transactions on Professional Communication, PC-23 (4):201–202, December 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Freeman:2003:SCT

[Fre03a] J. Freeman. The science of conversation: training in dialogue for NNS in engineering. IEEE Transactions on Professional Communication, 46(3): 157–167, September 2003. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Freiermuth:2003:CBS

[Fre03b]  M. R. Freiermuth.  Case-

based simulations in the EST classroom. IEEE Transactions on Professional Communication, 46(3):221–230, September 2003. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Frey:2013:SPS

[Fre13] Robert S. Frey. Successful proposal strategies for small businesses: using knowledge management to win government, private-sector, and international contracts. sixth edition (Frey, R. S.) [Book Reviews]. IEEE Transactions on Professional Communication, 56(1):86–88, March 2013. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Friedman:1973:STS

[Fri73] J. E. Friedman. Systematic techniques for solving nonlinear communications. IEEE Transactions on Professional Communication, PC-16(1):15– 17, March 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Friedmann:1985:WTG

[Fri85]  D. C. Friedmann.  Word-

star training guide and Word-Star reference manual. IEEE Transactions on Professional Communication, PC-28(1): 55, March 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Friess:2013:RAP

[Fri13] Erin Friess. Rhetorical appeals of professional designers in decision-making sessions. IEEE Transactions on Professional Communication, 56(4): 313–331, December 2013. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Friess:2015:PHE

[Fri15] Erin Friess. Personas in heuristic evaluation: an exploratory study. IEEE Transactions on Professional Communication, 58(2):176–191, June 2015. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Friess:2019:SLU

[Fri19] Erin Friess. Scrum language use in a software engineering firm: an exploratory study. IEEE Transactions on Professional Communication, 62(2): 130–147, June 2019. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Froehlich:2006:HRM

[Fro06] T. Froehlich. e-human resources management: Managing knowledge people. IEEE Transactions on Professional Communication, 49(2):199– 201, June 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Fry:1962:RNS

[Fry62] Bernard M. Fry. The role of the national science foundation in science information activities. IRE Transactions on Engineering Writing and Speech, EWS-5(2):87–93, December 1962. ISSN 0893-7869 (print), 2331-3692 (electronic).

Frye:1981:ATT

[Fry81] Robert H. Frye. Artistic technical training. IEEE Transactions on Professional Communication, PC-24(2):86–89, June 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Fry:1987:ECW

[Fry87] Robert Fry. Expanding concepts of the writer’s purpose, audience, and task: The IEEE Transactions on Professional Communication, 1981– 86. IEEE Transactions on Professional Communication, PC-30(1):4–11, March 1987. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Fearing:1980:STC

[FS80] Bertie E. Fearing and Thomas M. Sawyer. Speech for technical communicators: a bibliography. IEEE Transactions on Professional Communication, PC-23(1):53–61, March 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Field:1985:TCT

[FS85] Ronald M. Field and David Strong. Technology considerations for the technical writer. IEEE Transactions on Professional Communication, PC-28 (1):51–52, March 1985. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Furman:1992:ESI

[FS92] A. M. Furman and J. H. Spyridakis. The effect of systeminitiated advice on the use of menu navigation shortcuts. IEEE Transactions on Professional Communication, 35(2): 112–119, June 1992. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Fukuoka:1999:OJE

[FS99] W. Fukuoka and J. H. Spyridakis. The organization of Japanese expository passages. IEEE Transactions on Professional Communication, 42(3):

166–174, September 1999. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Fukuoka:2000:JRC

[FS00] W. Fukuoka and J. H. Spyridakis. Japanese readers’ comprehension of and preferences for inductively versus deductively organized text. IEEE Transactions on Professional Communication, 43(4):355– 367, December 2000. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Freeman:2009:ECI

[FS09] Kris S. Freeman and Jan H.

Spyridakis. Effect of contact information on the credibility of online health information. IEEE Transactions on Professional Communication, 52(2): 152–166, June 2009. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Fujii:1958:EWG

[Fuj58] Ted Fujii. Engineering writing without gobbledegook. IRE Transactions on Engineering Writing and Speech, EWS-1(2): 31, August 1958. ISSN 0893-7869 (print), 2331-3692 (electronic).

Fuller:1980:HWR

[Ful80] Don Fuller. How to write reports that won’t be ignored.

IEEE Transactions on Professional Communication, PC-23 (2):79–81, June 1980. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Furugori:1990:ISC

[Fur90]  T. Furugori.  Improving

spelling checkers for Japanese users of English. IEEE Transactions on Professional Communication, 33(3):138– 142, September 1990. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Fuller:2016:LEC

[FVB16] Robert M. Fuller, Chelley M. Vician, and Susan A. Brown. Longitudinal effects of computer-mediated communication anxiety on interaction in virtual teams. IEEE Transactions on Professional Communication, 59(3):166– 185, September 2016. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Fader:1978:WRC

[FW78] Sunny Fader and Ron Whittaker. The writer’s role in CCTV production. IEEE Transactions on Professional Communication, PC-21(3): 101–102, September 1978. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Farkas:1990:JCN

[FW90] D. K. Farkas and T. R.

Williams. John Carroll’s The Nurnberg Funnel and minimalist documentation. IEEE Transactions on Professional Communication, 33(4):182– 187, December 1990. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic). The Nurnberg Funnel: Designing Minimalist Instruction for Practical Computer Skill, — John M. Carroll, MIT Press, Cambridge, MA 1990.

Feuer:2023:PDP

[FW23] Maria Feuer and Joanna Wolfe. Planning for difference: Preparing students to create flexible and elaborated team charters that can adapt to support diverse teams. IEEE Transactions on Professional Communication, 66 (1):78–93, March 2023. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Flavin:1986:CCD

[FWB86] R. A. Flavin, J. D. Williford, and H. Barzilai. Computer conferencing data structures in the GRANDiose system. IEEE Transactions on Professional Communication, PC-29(1):34– 44, March 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Fry:1977:JPP

[FWJ77] Bernard M. Fry, Herbert S. White, and Elizabeth L. Johnson. Journal publisher practices and attitudes on article copying: Summary of a survey. IEEE Transactions on Professional Communication, PC-20 (3):151–154, November 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gollner:2015:SUD

[GAGW15] Joe Gollner, Rebekka Andersen, Kathleen Gollner, and Tenny Webster. A study of the usefulness of deploying a questionnaire to identify cultural dynamics potentially affecting a content-management project. IEEE Transactions on Professional Communication, 58(3): 289–308, September 2015. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gallivan:2001:MCH

[Gal01] M. J. Gallivan. Meaning to change: how diverse stakeholders interpret organizational communication about change initiatives. IEEE Transactions on Professional Communication, 44(4):243–266, December 2001. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gannett:1958:IEP

[Gan58] E. K. Gannett. IRE editorial policies and requirements.

IRE Transactions on Engineering Writing and Speech, EWS-1(1):20, March 1958. ISSN 0893-7869 (print), 2331-3692 (electronic).

Gannett:1977:INC

[Gan77] E. K. Gannett. IEEE and the new copyright law. IEEE Transactions on Professional Communication, PC-20(3): 165–166, November 1977. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

GANIER:2004:FAP

[Gan04] F. Ganier. Factors affecting the processing of procedural instructions: implications for document design. IEEE Transactions on Professional Communication, 47(1): 15–26, March 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ganguly:2022:ETC

[Gan22] Priyanka Ganguly. Equipping Technical Communicators for Social Justice Work: Theories, Methodologies, and Pedagogies: Rebecca Walton and Godwin Y. Agboka: [Book Review]. IEEE Transactions on Professional Communication, 65(1): 255–257, March 2022. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Garfield:1972:CEE

[Gar72] Eugene Garfield. Communication, engineering, and engineers: a PSYCHOM ’72 paper. IEEE Transactions on Professional Communication, PC-15 (2):49–51, June 1972. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Garfield:1973:WSJ

[Gar73] Eugene Garfield. What scientific journals can tell us about scientific journals. IEEE Transactions on Professional Communication, PC-16(4): 200–203, December 1973. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Garfield:1979:MCC

[Gar79a] Eugene Garfield. Making contacts at conferences. IEEE Transactions on Professional Communication, PC-22(3): 131–133, September 1979. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Garfield:1979:RNM

[Gar79b] Eugene Garfield. Radio: The neglected medium for scientific communication. IEEE Transactions on Professional Communication, PC-22(1):14– 15, March 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Garrett:1980:LHI

[Gar80a] Nancy Garrett. Let’s hear it for the audience. IEEE Transactions on Professional Communication, PC-23(1):7–8, March 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Garver:1980:PPP

[Gar80b] Robert V. Garver. Presenting the peer paper. IEEE Transactions on Professional Communication, PC-23(1):18– 22, March 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Garg:2006:IEP

[Gar06] N. Garg. Information ethics:

Privacy and intellectual property. IEEE Transactions on Professional Communication, 49(2):202–203, June 2006. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Garza:2019:HMJ

[Gar19] Alyssa Garza. H. A. McKee and J. E. Porter: Professional Communication and Network Interaction: a Rhetorical and Ethical Approach [Book Review]. IEEE Transactions on Professional Communication, 62(3):312–313, September 2019. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gattis:2004:UCR

[Gat04] L. Gattis. Using criteria in the recommendation report: a case study of the Texas A&M University Bonfire Commission findings. IEEE Transactions on Professional Communication, 47(4):281–289, December 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gould:1979:WDS

[GB79] John D. Gould and Stephen J. Boies. Writing, dictating, and speaking letters. IEEE Transactions on Professional Communication, PC-22(1):16– 18, March 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Galitz:1984:HSF

[GB84a] Wilbert O. Galitz and R. John Brockmann. Handbook of screen format design. IEEE Transactions on Professional Communication, PC-27(4): 232, December 1984. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Greenly:1984:HWG

[GB84b] Robert B. Greenly and Clark E. Beck. How to win government contracts. IEEE Transactions on Professional Communication, PC-27(4):230–231, December 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Getto:2016:TMU

[GB16] Guiseppe Getto and Fred Beecher. Toward a model of UX education: Training UX designers within the academy. IEEE Transactions on Professional Communication, 59 (2):153–164, June 2016. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gallivan:2005:FAL

[GBF05] M. J. Gallivan and R. Benunan-Fich. A framework for analyzing levels of analysis issues in studies of e-collaboration. IEEE Transactions on Professional Communication, 48(1): 87–104, March 2005. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gouge:2022:BTM

[GC22] Catherine C. Gouge and Erin Brock Carlson. Building toward more just data practices. IEEE Transactions on Professional Communication, 65(1):241–254, March 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Grene:2017:UPL

[GCMQ17] Margaret Grene, Yvonne Cleary, and Ann Marcus-Quinn. Use of plain-language guidelines to promote health literacy. IEEE Transactions on Professional Communication, 60(4):384–400, December 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Goodman:1985:WPE

[GD85] Michael B. Goodman and Bonnie Devet. Write to the point: Effective communication in the workplace. IEEE Transactions on Professional Communication, PC-28(1):53–54, March 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Guo:2009:IEV

[GDTZ09] Zixiu Guo, John D’Ambra, Tim Turner, and Huiying Zhang. Improving the effectiveness of virtual teams: a comparison of video-conferencing and face-to-face communication in China. IEEE Transactions on Professional Communication, 52(1):1–16, March 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gomez:2013:RAM

[GE13] Elizabeth Avery Gomez and Norbert Elliot. Research article measuring mobile ICT literacy: Short-message performance assessment in emergency response settings. IEEE Transactions on Professional Communication, 56(1):16– 32, March 2013. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Geiger:1975:IA

[Gei75] Stephen R. Geiger. Income from authors. IEEE Transactions on Professional Communication, PC-18(3):97–102, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Geiger:1977:STC

[Gei77] Stephen R. Geiger. Specialization: Times change, journals change. IEEE Transactions on Professional Communication, PC-20(2):87–89, September 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gentle:2010:CCS

[Gen10] Anne Gentle. Conversation and Community: The Social Web for Documentation Book Review. IEEE Transactions on Professional Communication, 53(4):423–425, December 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Georges:1983:FQH

[Geo83] T. M. Georges. Fifteen questions to help you write winning proposals. IEEE Transactions on Professional Communication, PC-26(2):84, June 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gerrard:1990:CRC

[Ger90]  M. Gerrard.  Confessions

of a reluctant consultant.

IEEE Transactions on Professional Communication, 33 (1):25–27, March 1990. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gettys:1986:IYW

[Get86] Deborah Gettys. If you write documentation, then try a decision table. IEEE Transactions on Professional Communication, PC-29(4):61–64, December 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Georgopoulos:1984:UTP

[GG84] Chris J. Georgopoulos and Voula C. Georgopoulos. From university term papers to industry technical reports an attempt to bridge the existing gap. IEEE Transactions on Professional Communication, PC-27(3):144–148, September 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Goubil-Gambrell:1996:DEI

[GG96] P. Goubil-Gambrell. Designing effective internet assignments in introductory technical communication courses. IEEE Transactions on Professional Communication, 39(4): 224–231, December 1996. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Giordano:2013:ETC

[GG13] Gabriel Giordano and Joey F.

George. The effects of task complexity and group member experience on computermediated groups facing deception. IEEE Transactions on Professional Communication, 56(3):210–225, September 2013. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gupta:2022:ACC

[GG22a] Varun Gupta and Chetna Gupta. Assembling Critical Components: a Framework for Sustaining Technical and Professional Communication: Joanna Schreiber and Lisa Melonçon: [book review]. IEEE Transactions on Professional Communication, 65(4): 534–535, December 2022. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gupta:2022:BCC

[GG22b] Varun Gupta and Chetna Gupta. Beyond Conversation: Collaboration and the Production of Writing: William Duffy: [book review]. IEEE Transactions on Professional Communication, 65(3):445– 447, September 2022. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gupta:2022:PPT

[GG22c] Varun Gupta and Chetna Gupta. The Profession and Practice of Technical Communication: Yvonne Cleary: [book review]. IEEE Transactions on Professional Communication, 65(3):442–444, September 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gallagher:2023:SUP

[GG23] Philip B. Gallagher and Guiseppe Getto. The state of UX pedagogy. IEEE Transactions on Professional Communication, 66(4):322– 337, December 2023. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gupta:2023:TCE

[GGS23] Varun Gupta, Chetna Gupta, and Lutz Sommer. Technical Communication for Environmental Action (Williams, S.D., Ed.) [book review]. IEEE Transactions on Professional Communication, 66(4):410– 412, December 2023. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Graves:1970:FSA

[GH70] Roy W. Graves and Donald P. Helander. A feasibility study of automatic indexing and information retrieval. IEEE Transactions on Engineering Writing and Speech, EWS-13(2):

58–59, September 1970. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Gibson:1985:SDN

[GH85] Jane Whitney Gibson and Richard M. Hodgetts. Selfdisclosure: a neglected management skill. IEEE Transactions on Professional Communication, PC-28(3):41–45, September 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gross:2009:SPP

[GH09] Alan G. Gross and Joseph E.

Harmon. The structure of PowerPoint presentations: The art of grasping things whole. IEEE Transactions on Professional Communication, 52(2): 121–137, June 2009. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Goodman:1991:CGM

[GHG91] M. B. Goodman, J. W. Hill, and K. R. Greene. Communication in graduate management programs: results of a survey. IEEE Transactions on Professional Communication, 34(1):24–35, March 1991. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Giangardella:1967:SCV

[GHR67] John J. Giangardella, James F.

Hudson, and Richard S. Roper.

Spelling correction by vector representation using a digital computer. IEEE Transactions on Engineering Writing and Speech, EWS-10(2):57–62, December 1967. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Giacoletto:1966:COC

[Gia66] L. J. Giacoletto. Computers and organizational communications. IEEE Transactions on Engineering Writing and Speech, EWS-9(1):24–25, July 1966. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Gilbertson:1987:BTC

[Gil87a] Michael Gilbertson. Balancing technical communication programs. IEEE Transactions on Professional Communication, PC-30(4):250–253, December 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gilsdorf:1987:WP

[Gil87b] J. W. Gilsdorf. Writing to persuade. IEEE Transactions on Professional Communication, PC-30(2):68–73, June 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

GILBERT:2003:DPD

[Gil03] C. C. Gilbert. Documents as prototypes: designing written drafts for communication across cross-disciplinary teams.

IEEE Transactions on Professional Communication, 46(4): 327–330, December 2003. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Girdwood:1960:TJB

[Gir60]  James Girdwood.  Techni-

cal journalism and the business press. IRE Transactions on Engineering Writing and Speech, EWS-3(1):20–23, April 1960. ISSN 0893-7869 (print), 2331-3692 (electronic).

Girill:1997:HCB

[Gir97]  T. R. Girill.  Hypertext

and Cognition [Book Review]. IEEE Transactions on Professional Communication, 40 (2):144–145, June 1997. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Girill:2006:BSR

[Gir06] T. R. Girill. Building sciencerelevant literacy with technical writing in high school tutorial. IEEE Transactions on Professional Communication, 49(4): 346–353, December 2006. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Garfield:1977:IDB

[GKD77] E. Garfield, M. Koenig, and T. Direnzo. ISI(R) data-base-produced information services. IEEE Transactions on Professional Communication, PC-20(2):95–99, September 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Getto:2016:IMU

[GL16] Guiseppe Getto and Jack T. Labriola. iFixit myself: Usergenerated content strategy in “The Free Repair Guide for Everything”. IEEE Transactions on Professional Communication, 59(1):37–55, March 2016. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

GLASBEEK:2004:SPY

[Gla04] H. Glasbeek. Solving problems on your own: how do exercises in tutorials interact with software learners’ level of goalorientedness? IEEE Transactions on Professional Communication, 47(1):44–53, March 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gonzales:2022:RDT

[GLCA22] Laura Gonzales, Robin Lewy, Erika Hernández Cuevas, and Vianna Lucía González Ajiataz. (Re)Designing technical documentation about COVID-19 with and for indigenous communities in Gainesville, Florida, Oaxaca de Juárez, Mexico, and Quetzaltenango, Guatemala. IEEE Transactions on Professional Communication, 65(1):34–49, March 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gleason:1982:HCI

[Gle82] James P. Gleason. Humor can improve your technical presentations. IEEE Transactions on Professional Communication, PC-25(2):86–90, June 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Getto:2023:SUT

[GLL23] Guiseppe Getto, Jack T. Labriola, and Amber Lancaster. The state of UX in technical and professional communication: Courses, programs, and jobs. IEEE Transactions on Professional Communication, 66(4):317–321, December 2023. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

George:2023:TGU

[GMG^{+}23] Joey F. George, Annette M.

Mills, Gabriel Giordano, Manjul Gupta, Vanesa M. Tennant, and Carmen C. Lewis. Toward a greater understanding of the use of nonverbal cues to deception in computer-mediated communication. IEEE Transactions on Professional Communication, 66(2):131–149, June 2023. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

George:2008:EWC

[GMT08] Joey F. George, Kent Marett, and Patti A. Tilley. The effects of warnings, computerbased media, and probing activity on successful lie detection. IEEE Transactions on Professional Communication, 51(1):1–17, March 2008. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Grider:1994:WMM

[GN94] D. Grider and R. Norman. When the medium is (much of) the message: the domain of structured document processors. IEEE Transactions on Professional Communication, 37(3):175–178, September 1994. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Goeller:1978:SCJ

[Goe78] Leo F. Goeller. Some communications jargon. IEEE Transactions on Professional Communication, PC-21(1):32– 33, March 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Goldsmith:1958:GWS

[Gol58] Alfred N. Goldsmith. Good writing and speech-their importance to the engineer. IRE Transactions on Engineering Writing and Speech, EWS-1(2): 9, August 1958. ISSN 0893-7869 (print), 2331-3692 (electronic).

Goldhar:1977:OUI

[Gol77] Joel D. Goldhar. Obtaining and using information in the year 2002. IEEE Transactions on Professional Communication, PC-20(2):126–129, September 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gold:1982:JIF

[Gol82a] Leon S. Gold. Job instruction: Four steps to success. IEEE Transactions on Professional Communication, PC-25 (1):26–30, March 1982. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Goldfarb:1982:WPP

[Gol82b] Stephen M. Goldfarb. Writing policies and procedures manuals. IEEE Transactions on Professional Communication, PC-25(1):14–15, March 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Goldie:1997:USC

[Gol97] P. Goldie. Using SGML to create complex interactive documents for electronic publishing. IEEE Transactions on Professional Communication, 40(2): 129–138, June 1997. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Golemon:2003:CIC

[Gol03] P. Golemon. Communicating in the intercultural classroom. IEEE Transactions on Professional Communication, 46(3): 231–235, September 2003. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gonzales:2020:DHP

[Gon20] Toni M. Gonzales. Design for How People Think: Using Brain Science to Build Better Products: John Whalen [Book Review]. IEEE Transactions on Professional Communication, 63(4):402–403, December 2020. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Goodman:1994:SSP

[Goo94] M. B. Goodman. The special section on professional communication in Russia: an American perspective. IEEE Transactions on Professional Communication, 37(2):60–61, June 1994. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Goodman:1995:ISS

[Goo95] M. B. Goodman. Introduction to the special sectionissues in corporate and organizational communication: communication and change. IEEE Transactions on Professional Communication, 38 (1):18–19, March 1995. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Goodman:1996:ICO

[Goo96] M. B. Goodman. Issues in corporate and organizational communication: Communication and change. IEEE Transactions on Professional Communication, 39(1):2, March 1996. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Goodman:1997:DDS

[Goo97a] M. B. Goodman. Designing & Delivering Scientific, Technical, and Managerial Presentations [Book Review]. IEEE Transactions on Professional Communication, 40(4):313– 315, December 1997. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Goodman:1997:SSE

[Goo97b] M. B. Goodman. Special section on electronic media and corporate communication. IEEE Transactions on Professional Communication, 40(1): 26–27, March 1997. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gooch:2005:DCI

[Goo05] J. C. Gooch. The dynamics and challenges of interdisciplinary collaboration: a case study of “Cortical depth of Bench” in group proposal writing. IEEE Transactions on Professional Communication, 48(2):177–190, June 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gordon:1990:CII

[Gor90] G. G. Gordon. Cultural imperatives for industry-oriented performance. IEEE Transactions on Professional Communication, 33(2):89–93, June 1990. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gottschall:1978:CT

[Got78] Edward M. Gottschall. Communications typographics. IEEE Transactions on Professional Communication, PC-21(1):18– 23, March 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gottlieb:1979:HPP

[Got79] Herbert H. Gottlieb. How to present a paper at an AAPT meeting. IEEE Transactions on Professional Communication, PC-22(3):134–137, September 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gould:1960:LET

[Gou60] Jay R. Gould. Language as an engineering tool — from preparation to results. IRE Transactions on Engineering Writing and Speech, EWS-3(1):24– 27, April 1960. ISSN 0893-7869 (print), 2331-3692 (electronic).

Gould:1970:AP

[Gou70] Calvin R. Gould. Anatomy of a presentation. IEEE Transactions on Engineering Writing and Speech, EWS-13(1):17–24, May 1970. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Gould:1972:OP

[Gou72] Calvin R. Gould. The overhead projector. IEEE Transactions on Professional Communication, PC-15(1):2–6, March 1972. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gould:1973:VAH

[Gou73]  Calvin R. Gould.  Visual

aids how to make them positively legible. IEEE Transactions on Professional Communication, PC-16(2):35–40, June 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Glatthorn:1986:WS

[GP86] Allan A. Glatthorn and Lynda Price. Writing for success. IEEE Transactions on Professional Communication, PC-29(3):35–36, September 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gonzales:2016:RCM

[GPHDM16] Laura Gonzales, Liza Potts, Bill Hart-Davidson, and Mike McLeod. Revising a contentmanagement course for a content strategy world. IEEE Transactions on Professional Communication, 59(1):56– 67, March 2016. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ganier:2012:TES

[GQ12] Franck Ganier and Ronan Querrec. TIP-EXE: a software tool for studying the use and understanding of procedural documents. IEEE Transactions on Professional Communication, 55(2):106–121, June 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Grice:1986:GEG

[GR86] Roger A. Grice and Daniel Rosich. From the guest editors the growing importance of computer documentation. IEEE Transactions on Professional Communication, PC-29 (4):2, December 1986. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Grice:1989:DMM

[GR89] R. A. Grice and L. S. Ridgway. A discussion of modes and motives for usability evaluation. IEEE Transactions on Professional Communication, 32(4):

230–237, December 1989. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Grice:1994:ECV

[GR94] R. A. Grice and L. S. Ridgway. The evolution of communication vehicles: linear progress and cyclical progress. IEEE Transactions on Professional Communication, 37(3): 179–182, September 1994. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Graham:1975:NLC

[Gra75a]  Joan Graham.  New look

at cold comp. IEEE Transactions on Professional Communication, PC-18(3):135–138, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gray:1975:OMM

[Gra75b]  Edward Gray.  Our mi-

croform marketing strategy. IEEE Transactions on Professional Communication, PC-18 (3):160–163, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Graves:1996:BCT

[Gra96a] R. Graves. Business Communication: a Technology-Based Approach [Book Review]. IEEE Transactions on Professional Communication, 39(4):244, December 1996.

CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Graves:1996:WWD

[Gra96b] R. Graves. Writings from the workplace: Documents, models, cases. IEEE Transactions on Professional Communication, 39(4):244, December 1996. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Grant:2013:FFJ

[Gra13] Lynn Grant. The frequency and function of Just in British and New Zealand engineering lectures. IEEE Transactions on Professional Communication, 56(2):176–190, June 2013. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Grant:2022:CTE

[Gra22] Carrie Grant. Collaborative tactics for equitable community partnerships toward social justice impact. IEEE Transactions on Professional Communication, 65(1):151–163, March 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Grey:1958:EED

[Gre58] Lennox Grey. Engineering English is different. IRE Transactions on Engineering Writing and Speech, EWS-1(1):2–5, March 1958. ISSN 0893-7869 (print), 2331-3692 (electronic).

Green:1962:CMT

[Gre62] John C. Green. Comparison of methods of technical communication in the U.S. and Russia. IRE Transactions on Engineering Writing and Speech, EWS-5(2):94–98, December 1962. ISSN 0893-7869 (print), 2331-3692 (electronic).

Gregory:1969:CSS

[Gre69] Dana S. Gregory. Compressed speech — the state of the art. IEEE Transactions on Engineering Writing and Speech, EWS-12(1):12–17, May 1969. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Greene:1970:EBT

[Gre70] Thomas G. Greene. Effective business and technical presentations. IEEE Transactions on Engineering Writing and Speech, EWS-13(1):29, May 1970. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Greene:1972:CEW

[Gre72] Thomas G. Greene. Communicating with the engineer who becomes unemployed: a PSYCHOM ’72 paper. IEEE Transactions on Professional Communication, PC-15(2):55– 59, June 1972. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Greenly:1985:TWI

[Gre85] Robert B. Greenly. Technical writing and illustrating strategies for winning government contracts. IEEE Transactions on Professional Communication, PC-28(2):7–12, June 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Greenly:1987:MMA

[Gre87] Robert B. Greenly. A mathematical model for assigning author page budgets in proposals. IEEE Transactions on Professional Communication, PC-30 (4):229–232, December 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Green:2017:NNP

[Gre17] Elise A. Green. The New Normal: Pressures on Technical Communication Programs in the Age of Austerity [by Tillery, D. and Nagelhout, E., Eds.; Book review]. IEEE Transactions on Professional Communication, 60(2):234–235, June 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Grimm:1981:EUD

[Gri81] Susan J. Grimm. EDP user documentation: The missing link. IEEE Transactions on Professional Communication, PC-24(2):79–83, June 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Grimm:1983:HWC

[Gri83] Susan J. Grimm. How to write computer manuals for users. IEEE Transactions on Professional Communication, PC-26 (1):45, March 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Grice:1984:CPE

[Gri84] Roger A. Grice. Cut and paste enters the computer age. IEEE Transactions on Professional Communication, PC-27 (2):78–81, June 1984. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gribbons:1991:VLC

[Gri91]  W. M. Gribbons.  Visual

literacy in corporate communication: some implications for information design. IEEE Transactions on Professional Communication, 34 (1):42–50, March 1991. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Grindsted:1994:ICS

[Gri94] A. Grindsted. The impact of cultural styles on negotiation: a case study of Spaniards and Danes. IEEE Transactions on Professional Communication, 37(1):34–38, March 1994. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Grove:1994:WBA

[Gro94]  L. K. Grove.  When the

basics aren’t enough: finding a comprehensive editor. IEEE Transactions on Professional Communication, 37(3): 171–174, September 1994. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gross:1995:PHH

[Gro95a] H. Gross. Promises, hopes, and hyperbole: adoption of new information technologies. IEEE Transactions on Professional Communication, 38(2): 118–122, June 1995. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Grove:1995:SE

[Gro95b] L. K. Grove. Speaking for the environment. IEEE Transactions on Professional Communication, 38(3):165–168, September 1995. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Grove:1999:WGE

[Gro99] L. K. Grove. Working In A Global Environment: Understanding, Communicating, and Managing Transnationally [Book Review]. IEEE Transactions on Professional Communication, 42(1):69–70, March 1999. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Grove:2001:GCC

[Gro01] L. K. Grove. Global contexts:

case studies in international technical communication [Book Review]. IEEE Transactions on Professional Communication, 44(4):302–303, December 2001. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

GROVE:2003:PCC

[Gro03] L. K. Grove. In pursuit of constructive criticism. IEEE Transactions on Professional Communication, 46(4):257– 264, December 2003. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Grove:2005:RRC

[Gro05] L. K. Grove. Research reality check: introduction to the special issue. IEEE Transactions on Professional Communication, 48(3):237–241, September 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gelinas:1996:ULC

[GRS96] U. J. Gelinas, D. V. Rama, and T. M. Skelton. Using learning communities to teach technical communication in accounting. IEEE Transactions on Professional Communication, 39(3): 135–145, September 1996. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gross:1985:NTT

[GS85] Lynne S. Gross and James H. Stone. The new television technologies. IEEE Transactions on Professional Communication, PC-28(1):52, March 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Guo:2008:EII

[GTTX08] Zixiu Guo, Felix B. Tan, Tim Turner, and Huizhong Xu. An exploratory investigation into instant messaging preferences in two distinct cultures. IEEE Transactions on Professional Communication, 51(4): 396–415, December 2008. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gudknecht:1982:WWD

[Gud82] Alan R. Gudknecht. Wellwritten documentation leaves nothing to imagination. IEEE Transactions on Professional Communication, PC-25(3): 112–119, September 1982. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Guillemette:1986:MRD

[Gui86] Ronald A. Guillemette. The meaning of reading-to-do documentation. IEEE Transactions on Professional Communication, PC-29(4):33–36, December 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Guillemette:1989:CPA

[Gui89a] R. A. Guillemette. The cloze procedure: assessing the understandability of an IEEE standard. IEEE Transactions on Professional Communication, 32(1):41–47, March 1989. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Guillemette:1989:UCD

[Gui89b] R. A. Guillemette. Usability in computer documentation design: conceptual and methodological considerations. IEEE Transactions on Professional Communication, 32(4): 217–229, December 1989. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Guindon:1994:URS

[Gui94] M. H. Guindon. Understanding the role of self-esteem in managing communication quality. IEEE Transactions on Professional Communication, 37(1): 21–27, March 1994. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gupta:1973:PTB

[Gup73] Madhu-Sudain Gupta. Preparation of a technical bibliography. IEEE Transactions on Professional Communication, PC-16(1):7–10, March 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gupta:1983:PFA

[Gup83a] Ashis Gupta. Profiling: a functional aid to effective communication. IEEE Transactions on Professional Communication, PC-26(4):197–200, December 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gupta:1983:BRS

[Gup83b] Madhu S. Gupta. Book reviews of scientific and technical books. IEEE Transactions on Professional Communication, PC-26(3):117–120, September 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gurak:1995:RDC

[Gur95] L. J. Gurak. Rhetorical dynamics of corporate communication in cyberspace: the protest over Lotus Market-Place. IEEE Transactions on Professional Communication, 38(1):2–10, March 1995. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

George:1982:DES

[GV82] James E. George and Anders Vinberg. The display of engineering and scientific data. IEEE Transactions on Professional Communication, PC-25 (2):95–97, June 1982. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gellevij:2002:SCS

[GvdM02] M. Gellevij and H. van der Meij. Screen captures to support switching attention. IEEE Transactions on Professional Communication, 45(2): 115–122, June 2002. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gellevij:1999:ESC

[GvdMdJP99] M. Gellevij, H. van der Meij, T. de Jong, and J. Pieters. The effects of screen captures in manuals: a textual and two visual manuals compared. IEEE Transactions on Professional Communication, 42(2): 77–91, June 1999. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gleason:1984:MDW

[GW84] James P. Gleason and Joan P.

Wackerman. Manual dexterity what makes instructional manuals usable. IEEE Transactions on Professional Communication, PC-27(2):59–61, June 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Gwiasda:1984:CCT

[Gwi84] Karl E. Gwiasda. Of classrooms and contexts: Teaching engineers to write wrong.

IEEE Transactions on Professional Communication, PC-27 (3):149–151, September 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Galli:1968:ESC

[GY68] Enrico J. Galli and Hisao M. Yamada. Experimental studies in computer-assisted correction of unorthographic text. IEEE Transactions on Engineering Writing and Speech, EWS-11 (2):75–84, August 1968. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Goo:2014:PSM

[GYK14] Jahyun Goo, Myung-Seong Yim, and Dan J. Kim. A path to successful management of employee security compliance: an empirical study of information security climate. IEEE Transactions on Professional Communication, 57(4): 286–308, December 2014. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Grove:1997:IBC

[GZ97] L. K. Grove and D. E. Zimmerman. Introduction: Bringing communication science to technical communication- advancing the profession. IEEE Transactions on Professional Communication, 40(3):157– 167, September 1997. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hackos:1984:TPS

[Hac84] Joann T. Hackos. Teaching problem-solving strategies in the technical communication classroom. IEEE Transactions on Professional Communication, PC-27(4):178–179, December 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hackos:1985:USA

[Hac85] Joann T. Hackos. Using systems analysis techniques in the development of standards and procedures. IEEE Transactions on Professional Communication, PC-28(3):25–30, September 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hackos:2016:ISI

[Hac16] JoAnn T. Hackos. International standards for information development and content management. IEEE Transactions on Professional Communication, 59(1):24–36, March 2016. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hagemann:1961:RDE

[Hag61] E. R. Hagemann. The role and duty of the engineer– writer. IRE Transactions on Engineering Writing and Speech, EWS-4(1):17–18, January 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Hager:1990:TAA

[Hag90] P. J. Hager. The technical article assignment: a work-world method for teaching audience analysis. IEEE Transactions on Professional Communication, 33(2):71–76, June 1990. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Halle:1980:RL

[Hal80] Morris Halle. The rules of language. IEEE Transactions on Professional Communication, PC-23(4):184–189, December 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hammet:1972:HET

[Ham72] Frank Hammet. The human equation in technical communication: a PSYCHOM ’72 paper. IEEE Transactions on Professional Communication, PC-15(3):94–99, September 1972. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hamilton:1996:BEH

[Ham96] S. S. Hamilton. Broadening employment horizons: transferring proposal writing skills from for-profit to nonprofit organizations. IEEE Transactions on Professional Communication, 39(2):99–102, June 1996. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hand:1967:TSN

[Han67] Harry E. Hand. Technical speech: a need for teaching and research. IEEE Transactions on Engineering Writing and Speech, EWS-10(2):48–51, December 1967. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Hand:1975:TSN

[Han75] Harry E. Hand. Technical speech: a need for teaching and research. IEEE Transactions on Professional Communication, PC-18(1):18–21, March 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hanley:1980:ECI

[Han80] John W. Hanley. Encountering and countering the instant experts. IEEE Transactions on Professional Communication, PC-23(3):117–120, September 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Haness:1983:HCD

[Han83] Joel Haness. How to critique a document. IEEE Transactions on Professional Communication, PC-26(1):15–17, March 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Happ:1973:GCP

[Hap73] W. W. Happ. Guidelines and checklists for preparing the planning report. IEEE Transactions on Professional Communication, PC-16(4):209–214, December 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Harman:1975:CEC

[Har75a] Eleanor T. Harman. Copy editing: Can we afford it? IEEE Transactions on Professional Communication, PC-18 (3):133–134, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Harris:1975:IMV

[Har75b] Michael S. Harris. Income management from the viewpoint of the commercial publisher. IEEE Transactions on Professional Communication, PC-18(3):108–112, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Harbaugh:1980:BTW

[Har80a] Annette D. Harbaugh. Basic technical writing, 4th ed., Herman M. Weisman. IEEE Transactions on Professional Communication, PC-23(4): 202–203, December 1980. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Harbaugh:1980:WWW

[Har80b] Annette D. Harbaugh. Writing that works, Walter E. Oliu, Charles T. Brusaw, and Gerald J. Alred. IEEE Transactions on Professional Communication, PC-23(4):202, December 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hartley:1981:EWI

[Har81] James Hartley. Eighty ways of improving instructional text. IEEE Transactions on Professional Communication, PC-24 (1):17–27, March 1981. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hartley:1984:RCT

[Har84] James Hartley. The role of colleagues and text-editing programs in improving text. IEEE Transactions on Professional Communication, PC-27(1):42– 44, March 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Harmon:1989:SSE

[Har89] J. E. Harmon. The structure of scientific and engineering papers: a historical perspective. IEEE Transactions on Professional Communication, 32(3): 132–138, September 1989. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Harris:1990:LGG

[Har90] R. A. Harris. Linguistic guidelines for graphic interfaces. IEEE Transactions on Professional Communication, 33(1): 46–53, March 1990. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Harmison:1997:CED

[Har97] M. Harmison. Creating electronic documents that interact with diagnostic software for onsite service. IEEE Transactions on Professional Communication, 40(2):92–101, June 1997. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Harney:2000:TWE

[Har00] M. Harney. Is technical writing an engineering discipline? IEEE Transactions on Professional Communication, 43(2): 210–212, June 2000. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Haskins:1980:TSP

[Has80] William A. Haskins. Time: a silent partner for public speakers. IEEE Transactions on Professional Communication, PC-23(1):38–39, March 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hatch:1966:IBM

[Hat66] A. S. Hatch. Improving boss– man and man–boss communication. IEEE Transactions on Engineering Writing and Speech, EWS-9(1):15–23, July 1966. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Hattery:1969:TCF

[Hat69] Lowell H. Hattery. Technological change — forecasting, appraisal, preparation, and accommodation with special attention to printing and publishing. IEEE Transactions on Engineering Writing and Speech, EWS-12(2):35–38, May 1969. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Hatch:1984:BW

[Hat84] Richard Hatch. Business writing. IEEE Transactions on Professional Communication, PC-27(1):48, March 1984. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hauptman:1979:PIB

[Hau79] Gunter A. Hauptman. Patenting inventions based on algorithms. IEEE Transactions on Professional Communication, PC-22(2):92–94, June 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hauptman:1980:WUU

[Hau80] Gunter A. Hauptman. Writing and understanding U.S. Patent claims: a programmed workbook, by Thomas J. Greer, Jr. IEEE Transactions on Professional Communication, PC-23(2):107, June 1980. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Havens:1979:SIH

[Hav79] H. Gordon Havens. Spellbound: Instant humility. IEEE Transactions on Professional Communication, PC-22(1): 21, March 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hawkes:1994:ND

[Haw94] L. Hawkes. Nested dialogues. IEEE Transactions on Professional Communication, 37(4): 240–244, December 1994. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Haynes:1980:BEL

[Hay80] Marion E. Haynes. Becoming an effective listener. IEEE Transactions on Professional Communication, PC-23(2):91– 94, June 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hays:1983:PUB

[Hay83] Robert Hays. Prescriptions for using boilerplate. IEEE Transactions on Professional Communication, PC-26(2):60– 62, June 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hayhoe:2016:NES

[Hay16] George F. Hayhoe. A new Editor, The same Transactions. IEEE Transactions on Professional Communication, 59 (2):70, June 2016. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hayhoe:2021:ETR

[Hay21] George F. Hayhoe. An Editor’s thanks for reviewer excellence. IEEE Transactions on Professional Communication, 64(1):1–2, March 2021. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hayhoe:2022:ET

[Hay22] George F. Hayhoe. An Editor’s thanks. IEEE Transactions on Professional Communication, 65(1):2–4, March 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hayhoe:2023:ET

[Hay23] George F. Hayhoe. An editor’s thanks. IEEE Transactions on Professional Communication, 66(1):1–2, March 2023. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hayhoe:2024:ET

[Hay24] George F. Hayhoe. An Editor’s thanks. IEEE Transactions on Professional Communication, 67(1):1–3, March 2024. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Horowitz:1994:SGM

[HB94] R. B. Horowitz and M. G. Barchilon. Stylistic guidelines for e-mail. IEEE Transactions on Professional Communication, 37(4):207–212, December 1994. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hartwick:2001:CDU

[HB01] J. Hartwick and H. Barki. Communication as a dimension of user participation. IEEE Transactions on Professional Communication, 44 (1):21–36, March 2001. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hassanzadeh:2021:UPD

[HBM21] Zahra Hassanzadeh, Robert Biddle, and Sky Marsen. User perception of data breaches. IEEE Transactions on Professional Communication, 64(4): 374–389, December 2021. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hantula:2008:OSF

[HBS08] Donald A. Hantula, Diane Di-Clemente Brockman, and Carter L. Smith. Online shopping as foraging: The effects of increasing delays on purchasing and patch residence. IEEE Transactions on Professional Communication, 51 (2):147–154, June 2008. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hubbard:1988:APP

[HC88] K. Hubbard and J. Campbell.

Automated proposal preparation techniques using a marketing text database. IEEE Transactions on Professional Communication, 31(2):65–69, June 1988. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hornik:2003:CSP

[HCKJ03] S. Hornik, Houn-Gee Chen, G. Klein, and J. J. Jiang. Communication skills of IS providers: an expectation gap analysis from three stakeholder perspectives. IEEE Transactions on Professional Communication, 46(1):17–34, March 2003. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Haaksma:2018:UPC

[HdJK18] Tim R. Haaksma, Menno D. T.

de Jong, and Joyce Karreman. Users personal conceptions of usability and user experience of electronic and software products. IEEE Transactions on Professional Communication, 61(2):116–132, June 2018. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hung:2008:RMC

[HDKC08] Yu-Ting Caisy Hung, Nguyen Thi Thao Duyen, Wei-Chang Kong, and Ai-Ling Chua. Reexamining media capacity theories using workplace instant messaging. IEEE Transactions on Professional Communication, 51(4):352–368, December 2008. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hoekstra:1977:AMF

[HE77] W. Hoekstra and V. H. C. Evers. Automated microfiche filing system for personal use. IEEE Transactions on Professional Communication, PC-20 (4):228–233, December 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Heald:1969:ETM

[Hea69] J. H. Heald. An equilibrium theory for the meeting paper. IEEE Transactions on Engineering Writing and Speech, EWS-12(1):18–19, May 1969. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Hea:1999:REC

[Hea99] A. C. K. Hea. A rhetoric of electronic communities. IEEE Transactions on Professional Communication, 42(3):200– 202, September 1999. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Heba:1997:DAR

[Heb97] G. Heba. Digital architectures: a rhetoric of electronic document structures. IEEE Transactions on Professional Communication, 40(4): 275–283, December 1997. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Heinle:1958:AB

[Hei58]  Chas. E. Heinle.  About

that book. IRE Transactions on Engineering Writing and Speech, EWS-1(2):20–21, August 1958. ISSN 0893-7869 (print), 2331-3692 (electronic).

Helm:1988:PTW

[Hel88]  L. B. Helm.  From pro-

gramming to technical writing. IEEE Transactions on Professional Communication, 31(4): 175–176, December 1988. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Helms:2007:TIT

[Hel07] Jason Helms. Total Interaction: Theory and Practice of a New Paradigm for the Design Disciplines (Buurman, G., Ed.; 2005). IEEE Transactions on Professional Communication, 50(2):180–181, June 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hennes:2022:CRC

[Hen22] Jack Hennes. Communicating in Risk, Crisis, and High Stress Situations: Evidence-Based Strategies and Practice: Vincent T. Covello: [book review]. IEEE Transactions on Professional Communication, 65(2):351–352, June 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hersey:1962:SC

[Her62] Irwin Hersey. The scientist as communicator. IRE Transactions on Engineering Writing and Speech, EWS-5(2):71–74, December 1962. ISSN 0893-7869 (print), 2331-3692 (electronic).

Hertzberg:1964:CEJ

[Her64] Martha S. Hertzberg. Communication — the engineer’s job? … A commentary on the Johnson paper in EWS 6-2. IEEE Transactions on Engineering Writing and Speech, EWS-7(2):2–4, September 1964. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic). See [Joh63].

Herbert:1989:CCA

[Her89] E. Herbert. Corporate culture: add 99 years of seasoning. IEEE Transactions on Professional Communication, 32(2):

69–75, June 1989. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Herrington:1995:EGD

[Her95] T. K. Herrington. Ethics and graphic design: a rhetorical analysis of the document design in the “Report of the Department of the Treasury on the Bureau of Alcohol, Tobacco, and Firearms investigation of Vernon Wayne Howell also known as David Kores”. IEEE Transactions on Professional Communication, 38(3): 151–157, September 1995. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hess:1975:ERP

[Hes75] Eugene L. Hess. Effects of the review process. IEEE Transactions on Professional Communication, PC-18(3):196–199, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Heumann:1975:IWG

[Heu75] Karl F. Heumann. An international working group in journal cooperation. IEEE Transactions on Professional Communication, PC-18(3):121–122, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hatch:1984:BCT

[HF84] Richard Hatch and Cheryl B.

Fenno. Business communication: Theory and practice. IEEE Transactions on Professional Communication, PC-27 (1):48–49, March 1984. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Holton:2008:UCE

[HF08] Carolyn F. Holton and Robert M. Fuller. Unintended consequences of electronic monitoring of instant messaging. IEEE Transactions on Professional Communication, 51(4): 381–395, December 2008. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Held:1994:EPF

[HFS94] J. A. Held, C. Fitzpatrick, and J. E. Sharp. Empty pathways: faculty use of computer-mediated communication. IEEE Transactions on Professional Communication, 37(4):203–206, December 1994. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hodgkinson:1982:DWI

[HH82] Richard Hodgkinson and John Hughes. Developing wordless instructions: a case history. IEEE Transactions on Professional Communication, PC-25 (2):74–79, June 1982. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Harmon:1993:BCT

[HH93] J. E. Harmon and D. R. Hamrin. Bibliography on communicating technical research information. IEEE Transactions on Professional Communication, 36(1):2–6, March 1993. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hackos:1997:CPD

[HHE97] J. T. Hackos, M. Hammar, and A. Elser. Customer partnering: data gathering for complex on-line documentation. IEEE Transactions on Professional Communication, 40(2): 102–110, June 1997. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hill:1985:BMC

[Hil85a] James W. Hill. Bibliography on management communications. IEEE Transactions on Professional Communication, PC-28(3):47–48, September 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hill:1985:IIM

[Hil85b] James W. Hill. Introduction to this issue on management communications. IEEE Transactions on Professional Communication, PC-28(3):1– 4, September 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hill:1989:UPB

[Hil89] R. J. Hill. Using personal bibliographic databases to keep up with the engineering literature. IEEE Transactions on Professional Communication, 32(3): 189–193, September 1989. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hine:1972:TSY

[Hin72] Edward A. Hine. Take a subject you know, then tell it like it is (techniques for writing better technical papers). IEEE Transactions on Professional Communication, PC-15 (1):6–8, March 1972. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hine:1973:WIB

[Hin73]  Edward A. Hine.  Work

it out before you write it. IEEE Transactions on Professional Communication, PC-16 (4):205–206, December 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hinson:2010:HRV

[Hin10] K. Hinson. Handbook of Research on Virtual Workplaces and the New Nature of Business Practices (Zemliansky, P. and St. Amant, K.) [Book Review]. IEEE Transactions on Professional Communication, 53(3):322–324, September 2010. CODEN IEPCBU.

ISSN 0361-1434 (print), 1558-1500 (electronic).

Hische:1977:TWD

[His77] Gerard Hische. Technical writing is different. IEEE Transactions on Professional Communication, PC-20(1):15–17, June 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hissong:1978:WPP

[His78] Douglas W. Hissong. Write and present persuasive reports. IEEE Transactions on Professional Communication, PC-21 (4):150–152, December 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Housman:1970:SAS

[HK70] Edward M. Housman and Elaine D. Kaskela. State of the art in selective dissemination of information. IEEE Transactions on Engineering Writing and Speech, EWS-13(2):78–83, September 1970. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Hwang:2007:UAC

[HK07] Yujong Hwang and Dan J.

Kim. Understanding affective commitment, collectivist culture, and social influence in relation to knowledge sharing in technology mediated learning. IEEE Transactions on Professional Communication, 50(3):

232–248, September 2007. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hoover:2011:UNA

[HK11] Ryan S. Hoover and Amy L. Koerber. Using NVivo to answer the challenges of qualitative research in professional communication: Benefits and best practices tutorial. IEEE Transactions on Professional Communication, 54(1):68–82, March 2011. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Holden:2022:PFB

[HK22] Julie Holden and Li Lian Khaw. POWR: a framework to bridge research planning and dissertation writing in engineering and information technology. IEEE Transactions on Professional Communication, 65(3):427–441, September 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hsu:2022:CTS

[HLY22] Long-Jing Hsu, Weijane Lin, and Hsiu-Ping Yueh. Connecting Twitter with scholarly networks: Exploring HCI scholars interactions from an SNA approach. IEEE Transactions on Professional Communication, 65(4):502–515, December 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Harwell:1961:TC

[HM61] George Harwell and H. B. M.

Technical communication. IRE Transactions on Engineering Writing and Speech, EWS-4(1): 27, January 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Hamelman:1973:MSM

[HM73] Paul W. Hamelman and Edward M. Mazze. Of models and scientific markets. IEEE Transactions on Professional Communication, PC-16 (3):120–125, September 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hitch:1994:HPT

[HM94] L. P. Hitch and J. P. Miller.

Historical perspectives on technology, ethics, and privacy. IEEE Transactions on Professional Communication, 37(1): 11–13, March 1994. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hynds:1995:AWC

[HM95] P. Hynds and W. Martin.

Atrisco Well #5: a case study of failure in professional communication. IEEE Transactions on Professional Communication, 38(3):139– 145, September 1995. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Huhmann:2002:RFH

[HMF02] B. A. Huhmann, D. L. Mothersbaugh, and G. R. Franke. Rhetorical figures in headings and their effect on text processing: the moderating role of information relevance and text length. IEEE Transactions on Professional Communication, 45(3):157–169, September 2002. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hager:1993:CTA

[HN93] P. J. Hager and R. J. Nelson. Chaucer’s ‘A Treatise on the Astrolabe’: a 600-year-old model for humanizing technical documents. IEEE Transactions on Professional Communication, 36(2):87–94, June 1993. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hofland:1980:TIT

[Hof80] Robert M. Hofland. Technical illustration, third edition J. C. Gibby. IEEE Transactions on Professional Communication, PC-23(2):108, June 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hoffer:1987:MRR

[Hof87]  William Hoffer.  A magic

ratio recurs throughout art and nature. IEEE Transactions on Professional Communication, PC-30(2):77–81, June 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hofmann:2005:GLP

[Hof05] P. L. Hofmann. Graphics for learning: Proven guidelines for planning, designing, and evaluating visuals in training materials. IEEE Transactions on Professional Communication, 48(2):234–236, June 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hofmann:2007:LIG

[Hof07] Patrick Hofmann. Localising and internationalising graphics and visual information commentary. IEEE Transactions on Professional Communication, 50(2):91–92, June 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hohmann:1960:STR

[Hoh60] Robert E. Hohmann. Space technology: Reporting the new dimension. IRE Transactions on Engineering Writing and Speech, EWS-3(1):7–15, April 1960. ISSN 0893-7869 (print), 2331-3692 (electronic).

Holland:1974:SCH

[Hol74] Winford E. Holland. The special communicator and his behavior in research organizations: a key to the management of informal technical information flow. IEEE Transactions on Professional Communication, PC-17(3/4): 48–53, September/December 1974. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Holder:1976:TWT

[Hol76] Fred W. Holder. Try writing a technical article. IEEE Transactions on Professional Communication, PC-19(2):38– 41, June 1976. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Holoviak:1977:ASC

[Hol77] Judy C. Holoviak. Authorsupplied composition. IEEE Transactions on Professional Communication, PC-20(2):62– 65, September 1977. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Holloway:2003:TWB

[Hol03] B. R. Holloway. Technical Writing Basics: a Guide to Style and Form-2nd ed. [Book Review]. IEEE Transactions on Professional Communication, 46(1):73–74, March 2003. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hoover:1977:MJ

[Hoo77] Dick M. Hoover. Merger of journals. IEEE Transactions on Professional Communication, PC-20(2):86–87, September 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Horton:1975:ESC

[Hor75] Derek Horton. Effective scientific communication: The viewpoint of a research investigator. IEEE Transactions on Professional Communication, PC-18 (3):234–238, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Horn:1984:CAW

[Hor84] William Dennis Horn. Computerassisted writing instruction at Clarkson University. IEEE Transactions on Professional Communication, PC-27(4): 197–200, December 1984. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Horowitz:1989:QIT

[Hor89] R. B. Horowitz. Quality improvement in technical documents and presentations: application in the classroom. IEEE Transactions on Professional Communication, 32(3): 162–170, September 1989. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Horton:1990:VRO

[Hor90] W. Horton. Visual rhetoric for online documents. IEEE Transactions on Professional Communication, 33(3):108– 114, September 1990. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Horton:1991:OCG

[Hor91] W. Horton. Overcoming chromophobia: a guide to the confident and appropriate use of color. IEEE Transactions on Professional Communication, 34(3):160–171, September 1991. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hota:2006:CVO

[Hot06] S. R. Hota. Communication in a virtual organization. IEEE Transactions on Professional Communication, 49(2): 204–205, June 2006. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Houze:1983:EIP

[Hou83]  William C. Houze.  The

electronic imperative: Process information intelligently or perish. IEEE Transactions on Professional Communication, PC-26(1):38–41, March 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Houser:2021:MBW

[Hou21a] Alan Houser. A Math-Based Writing System for Engineers: Sentence Algebra & Document Algorithms: Brad Henderson: [Book Review]. IEEE Transactions on Professional Communication, 64(1):98–99, March 2021. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Houser:2021:RPT

[Hou21b] Alan Houser. A Research Primer for Technical Communication: Methods, Exemplars, and Analyses, 2nd Edition: George F. Hayhoe and Pam Estes Brewer [Book Review]. IEEE Transactions on Professional Communication, 64(4): 470–471, December 2021. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Houser:2021:CIC

[Hou21c] Alan Houser. Creating Intelligent Content With Lightweight DITA: Carlos Evia: [Book Review]. IEEE Transactions on Professional Communication, 64(1):100–101, March 2021. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Houser:2021:TEI

[Hou21d] Alan Houser. Technical Editing: an Introduction to Editing in the Workplace: Donald H. Cunningham, Edward A.

Malone, and Joyce M. Rothschild [Book Review]. IEEE Transactions on Professional Communication, 64(4):468– 469, December 2021. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Houser:2022:THS

[Hou22a] Alan Houser. Trustworthy: How the Smartest Brands Beat Cynicism and Bridge the Trust Gap: Margot Bloomstein: [book review]. IEEE Transactions on Professional Communication, 65(3):448– 449, September 2022. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Houser:2022:WFC

[Hou22b] Alan Houser. Writing Futures: Collaborative, Algorithmic, Autonomous: Ann Hill Duin and Isabel Pedersen: [book review]. IEEE Transactions on Professional Communication, 65(4):532– 533, December 2022. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Houser:2023:DME

[Hou23] Alan Houser. Designing Multilingual Experiences in Technical Communication: Laura Gonzales: [book review]. IEEE Transactions on Professional Communication, 66(1):114– 116, March 2023. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hovav:2008:SDL

[Hov08] Anat Hovav. The socially driven life cycle of academic scholarship: a longitudinal study of six electronic journals. IEEE Transactions on Professional Communication, 51(1): 79–94, March 2008. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hovde:2014:FEC

[Hov14] Marjorie Rush Hovde. Factors that enable and challenge international engineering communication: a case study of a United States/British design team. IEEE Transactions on Professional Communication, 57(4):242–265, December 2014. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Howard:1983:STC

[How83] Lionel A. Howard. A survey of technical communication programs in U.S. colleges and universities. IEEE Transactions on Professional Communication, PC-26(3):113–116, September 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Howard:1984:STC

[How84] Lionel A. Howard. A survey of technical communication programs in U.S. colleges and universities 1984. IEEE Transactions on Professional Communication, PC-27(3):172–176, September 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hoyler:1958:MSM

[Hoy58] Cyril N. Hoyler. More senses make more sense. IRE Transactions on Engineering Writing and Speech, EWS-1(1):25, March 1958. ISSN 0893-7869 (print), 2331-3692 (electronic).

Harris:1986:SPP

[HP86] Patricia R. Harris and Thomas E. Pinelli. The standardization of publishing practices: an introduction to organizations and the standards process. IEEE Transactions on Professional Communication, PC-29(3):31– 34, September 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hamblen:1988:VCT

[HP88] J. O. Hamblen and A. Parker. Visual communication technologies for computer graphics. IEEE Transactions on Professional Communication, 31(3): 135–141, September 1988. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hamilton:1983:CRG

[HPSR83] Cheryl Hamilton, Cordell Parker, Doyle D. Smith, and Michael J. Rossi. Communicating for results: a guide for business and the professions. IEEE Transactions on Professional Communication, PC-26(2):91, June 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hume-Rothery:1960:NWS

[HR60] W. Hume-Rothery. A note on the writing of scientific papers. IRE Transactions on Engineering Writing and Speech, EWS-3(2):43–44, July 1960. ISSN 0893-7869 (print), 2331-3692 (electronic).

Harrington:1999:TTT

[HR99] S. J. Harrington and C. P.

Ruppel. Telecommuting: a test of trust, competing values, and relative advantage. IEEE Transactions on Professional Communication, 42(4): 223–239, December 1999. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Holland:1972:ESE

[HS72] Winford E. Holland and Bette A. Stead. Exploring the scientist-engineer conflict: a form and content analysis of their written communication. IEEE Transactions on Professional Communication, PC-15 (4):119–123, December 1972. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Harkins:1978:PM

[HS78] Craig Harkins and Emily Schlesinger. Perspective on multimedia. IEEE Transactions on Professional Communication, PC-21(3):118–128, September 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Harpp:1979:VLL

[HS79] David N. Harpp and James P. Snyder. Vitalizing the lecture: Lap-dissolve projection. IEEE Transactions on Professional Communication, PC-22 (1):9–13, March 1979. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hannah:2018:EEB

[HS18] Mark A. Hannah and Michael Simeone. Exploring an ethnography-based knowledge network model for professional communication analysis of knowledge integration. IEEE Transactions on Professional Communication, 61(4): 372–388, December 2018. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hodges:2019:WTW

[HS19] Amy Hodges and Leslie Seawright. Writing in transnational workplaces: Teaching strategies for multilingual engineers. IEEE Transactions on Professional Communication, 62(3):298–309, September 2019. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Heinonen:2021:IDS

[HSA21] Hanna Heinonen, Sanni Siltanen, and Petri Ahola. Information design for small screens: Toward smart glass use in guidance for industrial maintenance. IEEE Transactions on Professional Communication, 64(4):407–426, December 2021. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Holcombe:1984:WDM

[HSR84] Marya W. Holcombe, Judith K. Stein, and William E. Rivers. Writing for decision makers: Memos and reports with a competitive edge. IEEE Transactions on Professional Communication, PC-27(1):47, March 1984. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hayne:2005:UPC

[HSV05] S. C. Hayne, C. A. P. Smith, and L. R. Vijayasarathy. The use of pattern-communication tools and team pattern recognition. IEEE Transactions on Professional Communication, 48(4):377–390, December 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hirst:1977:CSJ

[HT77] Graeme Hirst and Nadia Talent. Computer science journals an iterated citation analysis. IEEE Transactions on Professional Communication, PC-20(4):233–238, December 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hartley:1980:RST

[HTB80] James Hartley, Mark Trueman, and Peter Burnhill. The role of spatial and typographic cues in the layout of journal references. IEEE Transactions on Professional Communication, PC-23 (3):138–142, September 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Huber:1985:LAW

[Hub85] Carol E. Huber. The logical art of writing useful comparisons. IEEE Transactions on Professional Communication, PC-28 (1):27–28, March 1985. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hubbard:1989:PAE

[Hub89] S. E. Hubbard. A practical approach to evaluating test results. IEEE Transactions on Professional Communication, 32(4):283–288, December 1989. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Huettner:2002:ETW

[Hue02] B. Huettner. The Elements of Technical Writing (2nd ed.) — Book Review. IEEE Transactions on Professional Communication, 45(1):59–60, March 2002. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hughson:1979:RWK

[Hug79] Roy V. Hughson. The right way to keep laboratory notebooks. IEEE Transactions on Professional Communication, PC-22(2):83–85, June 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hulbert:1980:EBW

[Hul80] Jack E. Hulbert. Effective business writing. IEEE Transactions on Professional Communication, PC-23(3):128–129, September 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hulbert:1981:CIB

[Hul81] Jack E. Hulbert. Conducting intelligent business dialog. IEEE Transactions on Professional Communication, PC-24 (3):132–135, September 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hunt:1981:NYR

[Hun81] Bridgford Hunt. The new you: Researched, resumed, and rarin’ to go. IEEE Transactions on Professional Communication, PC-24(4):160–162, December 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hundleby:1998:ETC

[Hun98]  M. N. Hundleby.  Ethics

In Technical Communication: Shades of Gray [Book Review]. IEEE Transactions on Professional Communication, 41(3): 218–220, September 1998. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Huntsman:2021:AWA

[Hun21]  Sherena Huntsman.  Ad-

dressing workplace accessibility practices through technical communication research methods: One size does not fit all. IEEE Transactions on Professional Communication, 64(3): 221–234, September 2021. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hurst:1977:DPM

[Hur77] Marlene Hurst. On-demand publishing from microforms. IEEE Transactions on Professional Communication, PC-20 (2):100–101, September 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hurlburt:1983:LDP

[Hur83] Allen Hurlburt. Layout: The design of the printed page. IEEE Transactions on Professional Communication, PC-26 (4):213, December 1983. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Huth:1975:CAE

[Hut75] Edward J. Huth. Cooperation among editors: The council of biology editors. IEEE Transactions on Professional Communication, PC-18(3):113–115, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hendriks:2012:SCP

[HvMK^{+}12] Berna Hendriks, Frank van Meurs, Hubert Korzilius, Rob le Pair, and Stephanie le Blanc-Damen. Style congruency and persuasion: a cross-cultural study into the influence of differences in style dimensions on the persuasiveness of business newsletters in Great Britain and The Netherlands. IEEE Transactions on Professional Communication, 55 (2):122–141, June 2012. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Heinonen:2022:MWU

[HVS^{+}22] Hanna Heinonen, Jenni Virtaluoto, Tiia Suomivuori, Kristian Forsman, Tuomas Kangas, and Sanni Siltanen. Minimalism for the win: Usercentered design for guidance in industrial maintenance. IEEE Transactions on Professional Communication, 65(4):485– 501, December 2022. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Haselkorn:2009:RIC

[HW09] Mark Haselkorn and Rebecca Walton. The role of information and communication in the context of humanitarian service. IEEE Transactions on Professional Communication, 52(4):325–328, December 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

House:2004:TER

[HWW04] R. House, A. Watt, and J. M. Williams. Teaching Enron: the rhetoric and ethics of whistleblowing. IEEE Transactions on Professional Communication, 47(4):244–255, December 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hyde:2018:EWE

[Hyd18] Charlotte Hyde. Ethics Within Engineering: an Introduction (Robison, W. L.) [Book Review]. IEEE Transactions on Professional Communication, 61(1):110–111, March 2018. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Hu:2024:MSI

[HZX24] Chunyu Hu, Aoran Zhang, and Yilin Xu. Metaphor, stance, and identity: a corpus-based study of CEO letters in Chinese and American corporate social responsibility reports. IEEE Transactions on Professional Communication, 67(1): 47–62, March 2024. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ibrahim:1985:WAN

[Ibr85] A. M. Ibrahim. What’s in an Arabic name? IEEE Transactions on Professional Communication, PC-28(4):28– 29, December 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ibrahim:1989:AO

[Ibr89] A. M. Ibrahim. Acronyms observed. IEEE Transactions on Professional Communication, 32(1):27–28, March 1989. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ilson:1980:RRT

[Ils80] Richard Ilson. Recent research in text processing. IEEE Transactions on Professional Communication, PC-23 (4):168–170, December 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ingram:2002:IGC

[IP02] S. Ingram and A. Parker. The influence of gender on collaborative projects in an engineering classroom. IEEE Transactions on Professional Communication, 45(1):7–20, March 2002. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Irani:2001:ISI

[IPC01] Z. Irani, A. Pouloudi, and W. Currie. Introduction to the special issue: social construct within an information society. IEEE Transactions on Professional Communication, 44(1): 4–5, March 2001. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Irwin:1972:CCS

[Irw72] J. D. Irwin. The conduct of conferences and short courses. IEEE Transactions on Professional Communication, PC-15 (2):45–48, June 1972. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ivory:1980:FAC

[Ivo80] Carol Ivory. First aid for the curriculum writer. IEEE Transactions on Professional Communication, PC-23(4): 182–184, December 1980. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Izquierdo:2024:ARC

[Izq24] Marlén Izquierdo. The audit report in contrast: Developing corpus-informed applications for Spanish users of English for business purposes.

IEEE Transactions on Professional Communication, 67(3): 333–351, September 2024. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Jacob:1981:WSB

[Jac81] Nora B. Jacob. Writing for success in business. IEEE Transactions on Professional Communication, PC-24(3):121–122, September 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Jamerson:1984:CIR

[Jam84] Frank E. Jamerson. Communication in industrial research laboratories. IEEE Transactions on Professional Communication, PC-27(1):35– 37, March 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Janas:2008:STP

[Jan08] Michael Janas. Starring the Text: The Place of Rhetoric in Science Studies (Gross, A. G.; 2006) [Book Review]. IEEE Transactions on Professional Communication, 51 (2):235–236, June 2008. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Johnstone:1978:WW

[JC78] Alex Johnstone and James Cassels. What’s in a word? IEEE Transactions on Professional Communication, PC-21 (4):165–167, December 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Janssen:2021:GZW

[JC21] Dawn Janssen and Stephen Carradini. Generation Z workplace communication habits and expectations. IEEE Transactions on Professional Communication, 64(2):137–153, June 2021. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Jackson:1992:OSB

[JCF92] S.-R. Jackson, J. Cherry, and B. Fryer. Online scenario-based task help. IEEE Transactions on Professional Communication, 35(2):91–97, June 1992. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Juhasz:1975:ARM

[JCJ^{+}75] Stephen Juhasz, Earl Calvert, Tom Jackson, David A. Kronick, and Joseph Shipman. Acceptance and rejection of manuscripts. IEEE Transactions on Professional Communication, PC-18(3):177–185, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Jonassen:1983:TTP

[JD83] David H. Jonassen and L. W. Denton. The technology of text: Principles for structuring, designing and displaying text. IEEE Transactions on Professional Communication, PC-26(1):44–45, March 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Jones:2011:FHD

[JD11] Scott L. Jones and Dianna De-Grow. Fortune 500 homepages: Design trends. IEEE Transactions on Professional Communication, 54(1):18–30, March 2011. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic). See comments [Jon15].

John:2021:DSS

[JD21] Divya John and G. Sandhiya Devi. Designing STEM-specific student-friendly reading content for the engineering English classroom. IEEE Transactions on Professional Communication, 64(4):444–455, December 2021. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Jacobsen:2023:TIU

[JD23] Kylie M. Jacobsen and Danielle DeVasto. Toward integrated UX instruction with symbiotic classrooms. IEEE Transactions on Professional Communication, 66(4):397–406, December 2023. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Johansson:1999:SEA

[JDJ99] C. Johansson, Y. Dittrich, and A. Juustila. Software engineering across boundaries: student project in distributed collaboration. IEEE Transactions on Professional Communication, 42(4):286–296, December 1999. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Jorn:1996:DMV

[JDW96] L. A. Jorn, A. H. Duin, and B. J. Wahlstrom. Designing and managing virtual learning communities. IEEE Transactions on Professional Communication, 39(4):183– 191, December 1996. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Johnson-Eilola:1995:ACA

[JE95] J. Johnson-Eilola. Accumulation, circulation, association: economies of information in online spaces. IEEE Transactions on Professional Communication, 38(4):228–238, December 1995. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Jennings:1962:AI

[Jen62] E. R. Jennings. About this issue. IRE Transactions on Engineering Writing and Speech, EWS-5(2):38, December 1962. ISSN 0893-7869 (print), 2331-3692 (electronic).

Jensen:1966:PPW

[Jen66] Paul E. T. Jensen. An inplant proposal-writing workshop. IEEE Transactions on Engineering Writing and Speech, EWS-9(2):46–47, December 1966. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Jenson:1999:ATY

[Jen99a] J. D. Jenson. The accidental trainer: you know computers, so they want you to teach everyone else. IEEE Transactions on Professional Communication, 42(3):196–197, September 1999. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Jenson:1999:SBT

[Jen99b] J. D. Jenson. Strategies For Business and Technical Writing, 4th Ed. [Book Review]. IEEE Transactions on Professional Communication, 42(4): 318, December 1999. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Jennings:2000:ESE

[Jen00a] A. S. Jennings. Employed students: ethical and legal issues in the technical communication classroom. IEEE Transactions on Professional Communication, 43(4):368–385, December 2000. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Jenson:2000:OPT

[Jen00b] J. D. Jenson. Oral presentations for technical communication [Book Review]. IEEE Transactions on Professional Communication, 43(4):415– 417, December 2000. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Jennings:2002:CIS

[Jen02]  A. S. Jennings.  Creating

an interactive science murder mystery game: the optimal experience of flow. IEEE Transactions on Professional Communication, 45(4):297– 301, December 2002. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Johnson-Eilola:2013:SPT

[JES13] Johndan Johnson-Eilola and Stuart A. Selber. Solving problems in technical communication. IEEE Transactions on Professional Communication, 56(3):256–259, September 2013. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Jones:2003:ICU

[JF03] A. A. Jones and T. E. Freeman. Imitation, copying, and the use of models: report writing in an introductory physics course. IEEE Transactions on Professional Communication, 46(3): 168–184, September 2003. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Johnson:1982:MTT

[JGH82] Parker H. Johnson, Judith Garrard, and William Hausman. A method for translating technical to nontechnical prose. IEEE Transactions on Professional Communication, PC-25 (4):182–185, December 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Jingjuan:1986:HDS

[Jin86] Madame Yu Jingjuan. How to deliver a successful lecture in China. IEEE Transactions on Professional Communication, PC-29(2):19–22, June 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Jin:2018:MAR

[Jin18a]  Bixi Jin.  A multidimen-

sional analysis of research article discussion sections in the field of chemical engineering. IEEE Transactions on Professional Communication, 61(3): 242–256, September 2018. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Jin:2018:RDR

[Jin18b]  Bixi Jin.  Rhetorical dif-

ferences in research article discussion sections of highand low-impact articles in the field of chemical engineering. IEEE Transactions on Professional Communication, 61(1): 65–76, March 2018. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Jones:1984:P

[JJ84] Edwin C. Jones and R. J. Joenk. Preface. IEEE Transactions on Professional Communication, PC-27(3):114–115, September 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Jeffers:2010:ITR

[JJ10] Patrick I. Jeffers and Rhoda C. Joseph. Information technology resources and customerservice process coordination in third-party logistics. IEEE Transactions on Professional Communication, 53(1):69– 83, March 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Juhasz:1983:WN

[JKW83] Stephen Juhasz, Theodore A. Klein, Jr., and James L. Wood. What’s in a name? IEEE Transactions on Professional Communication, PC-26 (3):121–122, September 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Jin:2020:EMR

[JLS20] Guangsa Jin, Chenle Li, and Ya Sun. Exploring the macrostructure of research articles in economics. IEEE Transactions on Professional Communication, 63(3):227– 243, September 2020. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Jenson:2001:TVT

[JM01] J. D. Jenson and D. Morrison. Textbooks versus technology: teaching professional writing to the next generation of technical communicators. IEEE Transactions on Professional Communication, 44(4):234–242, December 2001. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Joenk:1977:F

[Joe77a] R. J. Joenk. Foreword. IEEE Transactions on Professional Communication, PC-20(1):1, June 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Joenk:1977:P

[Joe77b] R. J. Joenk. Preface. IEEE Transactions on Professional Communication, PC-20(3): 141, November 1977. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Joenk:1978:Pa

[Joe78a] R. J. Joenk. Preface. IEEE Transactions on Professional Communication, PC-21(1):

1, March 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Joenk:1978:Pb

[Joe78b] R. J. Joenk. Preface. IEEE Transactions on Professional Communication, PC-21(2):42, June 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Joenk:1978:Pc

[Joe78c] R. J. Joenk. Preface. IEEE Transactions on Professional Communication, PC-21(3):90, September 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Joenk:1978:Pd

[Joe78d] R. J. Joenk. Preface. IEEE Transactions on Professional Communication, PC-21(4): 138, December 1978. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Joenk:1979:TPI

[Joe79a] R. J. Joenk. Toward paperless information systems, F. W. Lancaster. IEEE Transactions on Professional Communication, PC-22(1):41, March 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Joenk:1979:PII

[Joe79b] R. J. Joenk. Patents: Incentive to innovate and communicate an introduction. IEEE Transactions on Professional Communication, PC-22(2):46– 59, June 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Joenk:1979:Pa

[Joe79c] R. J. Joenk. Preface. IEEE Transactions on Professional Communication, PC-22(1):2– 3, March 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Joenk:1979:Pb

[Joe79d] R. J. Joenk. Preface. IEEE Transactions on Professional Communication, PC-22(4): 178, December 1979. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Joenk:1980:CPS

[Joe80a] R. J. Joenk. Call for papers special issue: Making information usable. IEEE Transactions on Professional Communication, PC-23(1):2, March 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Joenk:1980:Pa

[Joe80b] R. J. Joenk. Preface. IEEE Transactions on Professional Communication, PC-23(1):3– 2, March 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Joenk:1980:Pb

[Joe80c] R. J. Joenk. Preface. IEEE Transactions on Professional Communication, PC-23(2):66, June 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Joenk:1980:Pc

[Joe80d] R. J. Joenk. Preface. IEEE Transactions on Professional Communication, PC-23(3): 114, September 1980. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Joenk:1980:Pd

[Joe80e] R. J. Joenk. Preface. IEEE Transactions on Professional Communication, PC-23(4): 158, December 1980. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Joenk:1981:CPSc

[Joe81a] R. J. Joenk. Call for papers special issue: Interpreting technology. IEEE Transactions on Professional Communication, PC-24(4):203, December 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Joenk:1981:CPSa

[Joe81b] R. J. Joenk. Call for papers special issue: More usable information through graphics. IEEE Transactions on Professional Communication, PC-24 (1):56, March 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Joenk:1981:CPSb

[Joe81c] R. J. Joenk. Call for papers special issue: More usable information through graphics. IEEE Transactions on Professional Communication, PC-24 (2):108, June 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Joenk:1982:CPSa

[Joe82a] R. J. Joenk. Call for papers special issue: Interpreting technology. IEEE Transactions on Professional Communication, PC-25(1):53, March 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Joenk:1982:CPSb

[Joe82b] R. J. Joenk. Call for papers special issue: Technical documentation programs. IEEE Transactions on Professional Communication, PC-25 (4):223, December 1982. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Joenk:1982:P

[Joe82c] R. J. Joenk. Preface. IEEE Transactions on Professional Communication, PC-25(3): 110–111, September 1982. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Joenk:1983:CPS

[Joe83a] R. J. Joenk. Call for papers special issue: Developing the ability to communicate. IEEE Transactions on Professional Communication, PC-26 (3):154, September 1983. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Joenk:1983:P

[Joe83b] R. J. Joenk. Preface. IEEE Transactions on Professional Communication, PC-26(1):2– 3, March 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Joenk:1984:P

[Joe84] R. J. Joenk. Preface. IEEE Transactions on Professional Communication, PC-27(4): 178–179, December 1984. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Johnson:1963:CEJ

[Joh63] J. Richard Johnson. Communication — the engineer’s job. IEEE Transactions on Engineering Writing and Speech, EWS-6(2):82–84, December 1963. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic). See comment [Ano64j, Her64].

Johnson:1965:BRH

[Joh65] J. Richard Johnson. Book reviews: How to be brief, by Rudolf Flesch. IEEE Transactions on Engineering Writing and Speech, EWS-8(1):36, June 1965. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Johnson:1982:HWD

[Joh82] Thomas P. Johnson. How well do you inform? IEEE Transactions on Professional Communication, PC-25(1):5–9, March 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Johnson:1990:WOM

[Joh90] B. Johnson. Writing for the online medium: a critical review. IEEE Transactions on Professional Communication, 33(4): 178–181, December 1990. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Johnson:1997:RAD

[Joh97a] R. D. Johnson. Rhetoric and The Arts of Design [Book Review]. IEEE Transactions on Professional Communication, 40(2):146–147, June 1997. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Johnson:1997:CHW

[Joh97b] R. R. Johnson. Christina Haas, Writing Technology: Studies In The Materiality of Literacy. Mahwah, NJ: Lawrence Erlbaum Assoc., 1996. 279 pp.

IEEE Transactions on Professional Communication, 40(3): 244–245, September 1997. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Johnston:2002:TCE

[Joh02] C. Johnston. Technical Communication (6th, ed.) — Book Review. IEEE Transactions on Professional Communication, 45(1):57–58, March 2002. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Johansen:2010:SCR

[Joh10] J. Johansen. Structuring a corporate request: an international Website critique suited to the net generation. IEEE Transactions on Professional Communication, 53(3): 304–316, September 2010. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Johnson:2017:DRH

[Joh17] Sarah Johnson. Dirk Remley, How the Brain Processes Multimodal Technical Instructions. IEEE Transactions on Professional Communication, 60(3): 330–331, September 2017. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Johnson:2020:WIU

[Joh20] Meagan Johnson. Wicked, Incomplete, and Uncertain: User Support in the Wild and the Role of Technical Communication: Jason Swarts [Book Review]. IEEE Transactions on Professional Communication, 63(3):272–274, September 2020. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Jones:1973:RSG

[Jon73] Dorothy D. Jones. A review of style guides. IEEE Transactions on Professional Communication, PC-16(1):18– 22, March 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Jones:1978:FPM

[Jon78] P. F. Jones. Four principles of man-computer dialog. IEEE Transactions on Professional Communication, PC-21 (4):154–159, December 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Jones:1980:BBE

[Jon80] Hilary Jones. I believe in basic English. IEEE Transactions on Professional Communication, PC-23(1):34–37, March 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Jones:1989:EQW

[Jon89] D. R. Jones. Exploring quality: what Robert Pirsig’s ‘Zen and the Art of Motorcycle Maintenance’ can teach us about technical communication. IEEE Transactions on Professional Communication, 32(3):154–158, September 1989. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Jones:2015:REF

[Jon15]  Scott L. Jones.  A re-

examination of Fortune 500 homepage design practices. IEEE Transactions on Professional Communication, 58 (1):20–44, March 2015. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic). See [JD11].

Joseph:1970:LFC

[Jos70] Earl C. Joseph. A look at the future of computerized information technology. IEEE Transactions on Engineering Writing and Speech, EWS-13(2):37–40, September 1970. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Josephs:1975:IDM

[Jos75] Melvin J. Josephs. Information dissemination with microforms. IEEE Transactions on Professional Communication, PC-18(3):164–167, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Joseph:1981:WTM

[Jos81] Albert Joseph. Writing training materials that turn people on. IEEE Transactions on Professional Communication, PC-24(4):169–171, December 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Johnson:1986:GJD

[JP86] Bonnie McDaniel Johnson and Lynda Price. Getting the job done a guide to better communication for office staff. IEEE Transactions on Professional Communication, PC-29 (3):36, September 1986. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Johnson:1981:YME

[JS81] Wendell Johnson and Emily Schlesinger. Your most enchanted listener. IEEE Transactions on Professional Communication, PC-24(2):105–106, June 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Johnson:2007:UCT

[JSZ07] Robert R. Johnson, Michael J.

Salvo, and Meredith W. Zoetewey. User-centered technology in participatory culture: Two decades beyond a narrow conception of usability testing. IEEE Transactions on Professional Communication, 50(4): 320–332, December 2007. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Justus:1972:TMT

[Jus72] Paul E. Justus. There is more to typesetting than setting type. IEEE Transactions on Professional Communication, PC-15(1):13–16, March 1972. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Jansen:2001:NEF

[JVT01] F. Jansen, A. Van Lijf, and E. Toussaint. A note on the evaluation of footnotes and other devices for background information in popular scientific texts. IEEE Transactions on Professional Communication, 44(3):195–201, September 2001. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Jones:2017:SJI

[JW17] Natasha N. Jones and Miriam F. Williams. The social justice impact of plain language: a critical approach to plainlanguage analysis. IEEE Transactions on Professional Communication, 60(4):412– 429, December 2017. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Jones:2022:ARA

[JW22] Natasha N. Jones and Miriam F. Williams. Archives, rhetorical absence, and critical imagination: Examining Black women’s mental health narratives at Virginia’s Central State Hospital. IEEE Transactions on Professional Communication, 65(1):179–196, March 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kim:2003:PIS

[KA03] L. Kim and M. J. Albers.

Presenting information on the small-screen interface: effects of table formatting. IEEE Transactions on Professional Communication, 46(2):94–104, June 2003. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kaempf:2003:GSE

[Kae03] C. Kaempf. The MIT guide to science and engineering communication — 2nd ed. [Book Review]. IEEE Transactions on Professional Communication, 46(2):144–146, June 2003. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kaempf:2004:VTI

[Kae04] C. Kaempf. Visualizing technical information a cultural critique. IEEE Transactions on Professional Communication, 47(2):152–154, June 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kaempf:2005:WSE

[Kae05] C. Kaempf. Writing in the sciences exploring conventions of scientific discourse. IEEE Transactions on Professional Communication, 48(2):223– 225, June 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kaiman:1968:CAP

[Kai68] Arthur Kaiman. Computer-Aided Publications Editor. IEEE Transactions on Engineering Writing and Speech, EWS-11(2):65–75, August 1968. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Kalmbach:1997:HFT

[Kal97] J. Kalmbach. Human Factors for Technical Communicators [Book Reviews]. IEEE Transactions on Professional Communication, 40(1):55– 56, March 1997. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kampf:2001:CCB

[Kam01] C. E. Kampf. The Cultural Context In Business Communication [Book Review]. IEEE Transactions on Professional Communication, 44 (1):58–59, March 2001. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kanfer:1978:SL

[Kan78] Stefan Kanfer. The state of the language, 1977. IEEE Transactions on Professional Communication, PC-21(2):43–44, June 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kantrowitz:1985:WPT

[Kan85] B. Michael Kantrowitz. What price technical editing? Phase I: Reaching a lay audience. IEEE Transactions on Professional Communication, PC-28 (1):13–19, March 1985. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kantner:1994:TMU

[Kan94] L. Kantner. Techniques for managing a usability test. IEEE Transactions on Professional Communication, 37(3): 143–148, September 1994. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kanoksilapatham:2012:RAS

[Kan12] Budsaba Kanoksilapatham.

Research article structure of research article introductions in three engineering subdisciplines. IEEE Transactions on Professional Communication, 55(4):294–309, December 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kincaid:1981:CRE

[KAOC81] J. Peter Kincaid, James A. Aagard, John W. O’Hara, and Larry K. Cottrell. Computer readability editing system. IEEE Transactions on Professional Communication, PC-24(1):38–42, March 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kaplan:1962:SRT

[Kap62a] Sidney Kaplan. Storage and retrieval of technical documentation. IRE Transactions on Engineering Writing and Speech, EWS-5(2):83–86, December 1962. ISSN 0893-7869 (print), 2331-3692 (electronic).

Kapp:1962:SP

[Kap62b] Reginald O. Kapp. Stating the problem. IRE Transactions on Engineering Writing and Speech, EWS-5(2):61–64, December 1962. ISSN 0893-7869 (print), 2331-3692 (electronic).

Katter:1984:BRE

[Kat84] O. E. Katter. Book reviews: Edward R. Tufte, The visual display of quantitative information. IEEE Transactions on Professional Communication, PC-27(2):108, June 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Katz:1998:PLW

[Kat98a] S. M. Katz. Part I — Learning to write in organizations: what newcomers learn about writing on the job. IEEE Transactions on Professional Communication, 41(2):107–115, June 1998. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Katz:1998:PIH

[Kat98b] S. M. Katz. Part II — How newcomers learn to write: resources for guiding newcomers. IEEE Transactions on Professional Communication, 41(3): 165–174, September 1998. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kaufman:1988:TWC

[Kau88] J. Kaufman. Technical writing and computer programming. IEEE Transactions on Professional Communication, 31(4): 171–174, December 1988. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kolin:1984:SWW

[KB84] Philip C. Kolin and Judith Gunn Bronson. Successful writing at work. IEEE Transactions on Professional Communication, PC-27(1): 49, March 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kreth:2017:DST

[KB17] Melinda L. Kreth and Elizabeth Bowen. A descriptive survey of technical editors. IEEE Transactions on Professional Communication, 60(3): 238–255, September 2017. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Krowne:1981:IET

[KC81] C. M. Krowne and D. H. Covington. Integrating engineering and technical communications in sophomore electrical engineering projects. IEEE Transactions on Professional Communication, PC-24(3):145–147, September 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kauffmann:2019:CST

[KC19] David Kauffmann and Golan Carmi. A comparative study of temporary and ongoing teams on e-environment. IEEE Transactions on Professional Communication, 62(2):148– 164, June 2019. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kock:2008:ESS

[KCJC08] Ned Kock, Ruth Chatelain-Jardon, and Jesus Carmona. An experimental study of simulated Web-based threats and their impact on knowledge communication effectiveness. IEEE Transactions on Professional Communication, 51 (2):183–197, June 2008. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kumi:2013:RAL

[KCLG13] Richard Kumi, Christopher M. Conway, Moez Limayem, and Sandeep Goyal. Research article learning in color: How color and affect influence learning outcomes. IEEE Transactions on Professional Communication, 56(1):2–15, March 2013. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kock:2015:MNC

[KCM15] Ned Kock, Jesus Carmona, and Murad Moqbel. Media naturalness and compensatory adaptation: Counterintuitive effects on correct rejections of deceitful contract clauses. IEEE Transactions on Professional Communication, 58(4): 381–395, December 2015. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Keyser:1984:LTW

[KD84] George F. Keyser and Eugene M. De Loatch. Learning through writing in an engineering course. IEEE Transactions on Professional Communication, PC-27(3):126–129, September 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kent-Drury:2000:BBN

[KD00]  R. Kent-Drury.  Bridging

boundaries, negotiating differences: the nature of leadership in cross-functional proposalwriting groups. IEEE Transactions on Professional Communication, 43(1):90–98, March 2000. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kim:2020:FLA

[KDBC20] Hyun Woo Kim, Bertha Du-Babcock, and Hyejung Chang. Following the leader: an analysis of leadership and conformity in business meetings. IEEE Transactions on Professional Communication, 63(4): 311–326, December 2020. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Karreman:2014:EDP

[KdJH14] Joyce Karreman, Menno de Jong, and Stefan Hofmans. The effects of different parts of the annual report on potential investors’ attitudes towards the company and on the corporate reputation. IEEE Transactions on Professional Communication, 57(2):78–97, June 2014. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Killingsworth:1986:MPT

[KE86] M. Jimmie Killingsworth and Kimberly Eiland. Managing the production of technical manuals: Recent trends. IEEE Transactions on Professional Communication, PC-29 (2):23–26, June 1986. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Keith:1967:IWM

[Kei67]  Marshall Keith.  Informal

writing: Milestone or millstone? IEEE Transactions on Engineering Writing and Speech, EWS-10(2):39–40, December 1967. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Kelton:1961:CDC

[Kel61a] G. Kelton. Comment on Dudley’s comment. IRE Transactions on Engineering Writing and Speech, EWS-4(1):26, January 1961. ISSN 0893-7869 (print), 2331-3692 (electronic). See [Dud61a].

Kelton:1961:IT

[Kel61b] Gilbert Kelton. Information theory! IRE Transactions on Engineering Writing and Speech, EWS-4(1):21–25, January 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Kendall:1963:PD

[Ken63] R. Kendall. The proposal digest. IEEE Transactions on Engineering Writing and Speech, EWS-6(2):79–81, December 1963. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Kennedy:1964:WIT

[Ken64a] Robert A. Kennedy. Writing informative titles for technical literature — an aid to efficient information retrieval. IEEE Transactions on Engineering Writing and Speech, EWS-7 (1):4–5, March 1964. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Kent:1964:TFC

[Ken64b] George Kent. A technique for facilitation of card-catalog indexing in the conventional library. IEEE Transactions on Engineering Writing and Speech, EWS-7(1):2–3, March 1964. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Kenneth:1982:TDL

[Ken82a]  R. G. Kenneth.  Turning

the disgruntled linguist into a gruntled one. IEEE Transactions on Professional Communication, PC-25(3):150–151, September 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kent:1982:YBW

[Ken82b] John L. Kent. Your business writing affects your success. IEEE Transactions on Professional Communication, PC-25 (4):220–221, December 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kennedy:1988:SAC

[Ken88] W. A. Kennedy. Speaking across the curriculum: teaching technical students to speak effectively. IEEE Transactions on Professional Communication, 31(3):130–134, September 1988. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kennedy:1989:RSA

[Ken89] W. A. Kennedy. Reactions to ’speaking across the curriculum: teaching technical students to speak effectively’: practical concerns and considerations. IEEE Transactions on Professional Communication, 32(1):32–35, March 1989. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kerfoot:1977:TTW

[Ker77]  Glenn W. Kerfoot.  Tips

on technical writing. IEEE Transactions on Professional Communication, PC-20(1):17– 19, June 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kerfoot:1980:LAS

[Ker80] Glenn Kerfoot. Let the audience see your presentation. IEEE Transactions on Professional Communication, PC-23 (1):50–52, March 1980. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kerr:1986:ELG

[Ker86] Elaine B. Kerr. Electronic leadership: a guide to moderating online conferences. IEEE Transactions on Professional Communication, PC-29(1):12– 18, March 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kessler:1960:BSC

[Kes60] M. M. Kessler. Background to scientific communication. IRE Transactions on Engineering Writing and Speech, EWS-3(1): 3–6, April 1960. ISSN 0893-7869 (print), 2331-3692 (electronic).

Keyes:1987:IDM

[Key87] Elizabeth Keyes. Information design: Maximizing the power and potential of electronic publishing equipment. IEEE Transactions on Professional Communication, PC-30 (1):32–37, March 1987. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Keplinger:1977:RCC

[KF77] Michael S. Keplinger and Robert W. Frase. Role of CONTU in computers and photocopying. IEEE Transactions on Professional Communication, PC-20(3):167–170, November 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Klerer:1968:ETC

[KG68] Melvin Klerer and Fred Grossman. Editing and type composition of two-dimensional mathematical text via computer. IEEE Transactions on Engineering Writing and Speech, EWS-11(2):53–64, August 1968. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Kock:2008:IDP

[KHH^{+}08] Ned Kock, Donald A. Hantula, Stephen C. Hayne, Gad Saad, Peter M. Todd, and Richard T. Watson. Introduction to Darwinian perspectives on electronic communication. IEEE Transactions on Professional Communication, 51(2): 133–146, June 2008. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Knievel:2010:BRT

[KHV10] Michael Knievel, April Heaney, and Meg Van Baalen-Wood. Breaking the rules: Teaching grammar wrong for the right results in technical communication consulting for engineers. IEEE Transactions on Professional Communication, 53(1): 58–68, March 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kaufer:2006:CSC

[KI06] D. Kaufer and S. Ishizaki. A corpus study of canned letters: Mining the latent rhetorical proficiencies marketed to writers-in-a-hurry and nonwriters. IEEE Transactions on Professional Communication, 49(3):254–266, September 2006. CODEN IEPCBU.

ISSN 0361-1434 (print), 1558-1500 (electronic).

Kiefer:1982:GWB

[Kie82] Kate Kiefer. A guide for writing better technical papers. IEEE Transactions on Professional Communication, PC-25 (3):162–163, September 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kiefer:1983:RWP

[Kie83] Kate Kiefer. Roy W. Poe, The McGraw–Hill Guide to Effective Business Reports. IEEE Transactions on Professional Communication, PC-26(1):43– 44, March 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Killingsworth:1982:STG

[Kil82] M. Jimmie Killingsworth. Science and technology for a general audience: The personal essay. IEEE Transactions on Professional Communication, PC-25(4):186–188, December 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Killingsworth:1983:BPW

[Kil83] M. Jimmie Killingsworth. A bibliography on proposal writing. IEEE Transactions on Professional Communication, PC-26(2):79–83, June 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Killingsworth:1996:GKD

[Kil96] M. J. Killingsworth. Genre knowledge in disciplinary communication: Cognition/ culture/power. IEEE Transactions on Professional Communication, 39(2):107, June 1996. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Killoran:2009:TAR

[Kil09] John B. Killoran. Targeting an audience of robots: Search engines and the marketing of technical communication business websites. IEEE Transactions on Professional Communication, 52(3):254– 271, September 2009. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Killoran:2013:HUS

[Kil13] John B. Killoran. How to use search engine optimization techniques to increase Website visibility. IEEE Transactions on Professional Communication, 56(1):50–66, March 2013. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kimball:1979:PUI

[Kim79] Albert B. Kimball. Patenting U.S. inventions abroad. IEEE Transactions on Professional Communication, PC-22 (2):119–122, June 1979. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kinn:1958:WE

[Kin58] John M. Kinn. Writing for “electronics”. IRE Transactions on Engineering Writing and Speech, EWS-1(1):21–22, March 1958. ISSN 0893-7869 (print), 2331-3692 (electronic).

King:1977:SEI

[Kin77]  Donald W. King.  Sys-

temic and economic interdependences in journal publication. IEEE Transactions on Professional Communication, PC-20(2):106–113, September 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

King:2012:ROM

[Kin12] Cynthia L. King. Reverse outlining: a method for effective revision of document structure. IEEE Transactions on Professional Communication, 55(3): 254–261, September 2012. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kirkscey:2020:SMC

[Kir20] Russell Kirkscey. Scientific and Medical Communication: a Guide for Effective Practice: Scott A. Mogull [Book Review]. IEEE Transactions on Professional Communication, 63(2):201–202, June 2020. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kallendorf:1990:COC

[KK90] Craig Kallendorf and Carol Kallendorf. The case for organizational communication. IEEE Transactions on Professional Communication, 33(4): 188–192, December 1990. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Krishnan:2002:LRS

[KK02] L. A. Krishnan and S. S.

Kathpalia. Literature reviews in student project reports. IEEE Transactions on Professional Communication, 45(3): 187–197, September 2002. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Khalifa:2006:SND

[KL06] M. Khalifa and Vanessa Liu.

Semantic network discussion representation: applicability and some potential benefits. IEEE Transactions on Professional Communication, 49(1): 69–81, March 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kock:2012:RAE

[KL12a] Ned Kock and Gary S. Lynn.

Research article electronic media variety and virtual team performance: The mediating role of task complexity coping mechanisms. IEEE Transactions on Professional Communication, 55(4):325– 344, December 2012. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Koh:2012:TET

[KL12b] Elizabeth Koh and John Lim. Too early, too bad: Uncovering and understanding the initial participation paradox in technology-mediated learning teams. IEEE Transactions on Professional Communication, 55(1):55–84, March 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Klepper:1981:EID

[Kle81] Michael M. Klepper. Executive interviews: Do you or don’t you? IEEE Transactions on Professional Communication, PC-24(4):192–194, December 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Klein:2003:BEH

[Kle03] J. M. E. Klein. Building Enhanced HTML Help with DHTML & CSS [Book Review]. IEEE Transactions on Professional Communication, 46(1):68–69, March 2003. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Klimann:1979:GE

[Kli79] A. H. R. Klimann. Grammatical engineering. IEEE Transactions on Professional Communication, PC-22(1):37– 38, March 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kindler:1961:OTC

[KM61] Herbert S. Kindler and Herbert B. Michaelson. Organizing the technical conference. IRE Transactions on Engineering Writing and Speech, EWS-4(2):62, May 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Kuzminsky:1994:NCC

[KM94] M. Kuzminsky and A. Mendkovich. Non-commercial computer networking in Russia. IEEE Transactions on Professional Communication, 37(2): 70–72, June 1994. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kim:2014:FNQ

[KMR14] Jongwoo Kim, Kannan Mohan, and Balasubramaniam Ramesh. Functional and nonfunctional quality in cloudbased collaborative writing: an empirical investigation. IEEE Transactions on Professional Communication, 57(3):182– 203, September 2014. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kock:2005:EBC

[KN05] N. Kock and J. Nosek. Expanding the boundaries of ecollaboration. IEEE Transactions on Professional Communication, 48(1):1–9, March 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Knapp:1984:CEW

[Kna84] Joan Knapp. Can engineers write? IEEE Transactions on Professional Communication, PC-27(1):10–13, March 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kniskern:1986:SJT

[Kni86] William Kniskern. Samuel Johnson: Technical writer. IEEE Transactions on Professional Communication, PC-29 (2):3–6, June 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Knievel:2008:TDF

[Kni08] Michael Knievel. Teaching deadly force: Collaborative dynamics and the rhetoric of police policy. IEEE Transactions on Professional Communication, 51(1):95–109, March 2008. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kaba:2009:MFI

[KNMM09] Bangaly Kaba, Koffi N’Da, Peter Meso, and Victor Wacham A. Mbarika. Micro factors influencing the attitudes toward and the use of a mobile technology: a model of cell-phone use in Guinea. IEEE Transactions on Professional Communication, 52(3): 272–290, September 2009. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Knoll:1980:PTC

[Kno80] Henry Knoll. Process and thought in composition, 2nd ed., Frank J. D’Angelo. IEEE Transactions on Professional Communication, PC-23(4): 203–204, December 1980. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kim:2020:UTF

[KO20] Dave Kim and Wendy M.

Olson. Using a transferfocused writing pedagogy to improve undergraduates lab report writing in gateway engineering laboratory courses. IEEE Transactions on Professional Communication, 63(1): 64–84, March 2020. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Koch:1975:HCS

[Koc75] Mitchell Koch. An in-house camera system for making line and halftone illustration prints. IEEE Transactions on Professional Communication, PC-18 (3):288–291, September 1975.

CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Koch:1977:INC

[Koc77] H. William Koch. Implications of the new copyright law for scientific societies. IEEE Transactions on Professional Communication, PC-20(3):159–164, November 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kock:2001:CAL

[Koc01] N. Kock. Compensatory adaptation to a lean medium: an action research investigation of electronic communication in process improvement groups. IEEE Transactions on Professional Communication, 44(4): 267–285, December 2001. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kock:2003:ARL

[Koc03a] N. Kock. Action research: lessons learned from a multiiteration study of computermediated communication in groups. IEEE Transactions on Professional Communication, 46(2):105–128, June 2003. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kock:2003:CFB

[Koc03b] N. Kock. Communicationfocused business process redesign: assessing a communication flow optimization model through an action research study at a defense contractor. IEEE Transactions on Professional Communication, 46(1): 35–54, March 2003. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kock:2005:MRM

[Koc05] N. Kock. Media richness or media naturalness? The evolution of our biological communication apparatus and its influence on our behavior toward e-communication tools. IEEE Transactions on Professional Communication, 48(2): 117–130, June 2005. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Koenig:1977:SDP

[Koe77] Michael E. D. Koenig. Structure of document provision services in the library-publisher interface. IEEE Transactions on Professional Communication, PC-20(2):74–75, September 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kolb:1993:LET

[Kol93] J. A. Kolb. Leading engineering teams: leader behaviors related to team performance. IEEE Transactions on Professional Communication, 36(4): 206–211, December 1993. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kong:2022:CAC

[Kon22]  Yeqing Kong.  A corpus-

assisted critical discourse analysis of news construction of the Flint water crisis. IEEE Transactions on Professional Communication, 65(4):450– 466, December 2022. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Korbuly:1975:NAC

[Kor75] Dorothy K. Korbuly. New approach to coding displayed mathematics for photocomposition. IEEE Transactions on Professional Communication, PC-18(3):283–287, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Korbuly:1977:OPP

[Kor77] Dorothy K. Korbuly. Overcoming people problems in the switch to automated composition. IEEE Transactions on Professional Communication, PC-20(2):58–61, September 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kostelnick:1995:CAI

[Kos95] C. Kostelnick. Cultural adaptation and information design: two contrasting views. IEEE Transactions on Professional Communication, 38(4):

182–196, December 1995. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kostelnick:2007:VRD

[Kos07] Charles Kostelnick. The visual rhetoric of data displays: The conundrum of clarity. IEEE Transactions on Professional Communication, 50(4): 280–294, December 2007. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kostelnick:2008:VRD

[Kos08] Charles Kostelnick. The visual rhetoric of data displays: The conundrum of clarity*. IEEE Transactions on Professional Communication, 51(1): 116–130, March 2008. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kowal:1978:PW

[Kow78] John Paul Kowal. The photographer as writer. IEEE Transactions on Professional Communication, PC-21(2):69– 71, June 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kock:2008:CHM

[KPV08] Ned Kock, Ronaldo Parente, and Jacques Verville. Can Hofstede’s model explain national differences in perceived information overload? A look at data from the US and New Zealand. IEEE Transactions on Professional Communication, 51(1):33–49, March 2008. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kenny:1983:PY

[KR83] Michael Kenny and Jeanine Roadhouse. Presenting yourself. IEEE Transactions on Professional Communication, PC-26(2):89–90, June 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Killingsworth:1995:IPC

[KR95] M. J. Killingsworth and M. E. Rosenberg. The icon as a problem in cognition and social construction: complexity and consensual domains in technical rhetoric. IEEE Transactions on Professional Communication, 38(4):216–227, December 1995. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kremers:1986:BRH

[Kre86] Marshall Kremers. Book reviews: Herbert B. Michaelson, How to write and publish engineering papers and reports, second edition. IEEE Transactions on Professional Communication, PC-29(3):35, September 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kremers:1987:BRR

[Kre87a] Marshall Kremers. Book reviews: R. John Brockmann and Fern Rook, ed., Technical communication and ethics. IEEE Transactions on Professional Communication, PC-30 (3):212, September 1987. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kremers:1987:PPI

[Kre87b] Marshall Kremers. PCS publishes independent learning program. IEEE Transactions on Professional Communication, PC-30(4):216, December 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kreppel:1988:PRA

[Kre88] M. C. Kreppel. Pooling resources around the lectern: one heuristic approach. IEEE Transactions on Professional Communication, 31(4):177– 180, December 1988. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kremers:1989:TET

[Kre89] M. Kremers. Teaching ethical thinking in a technical writing course. IEEE Transactions on Professional Communication, 32(2):58–61, June 1989. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kreth:2000:SCO

[Kre00] M. L. Kreth. A survey of the co-op writing experiences of recent engineering graduates. IEEE Transactions on Professional Communication, 43(2): 137–152, June 2000. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kuryloski:2024:IES

[KRM24] Lauren Kuryloski, Erin Rowley, and Kristen R. Moore. Introducing engineering students to standards and regulatory research and writing. IEEE Transactions on Professional Communication, 67(1): 133–151, March 2024. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Krouse:1980:CCB

[Kro80] John K. Krouse. CAD/CAM bridging the gap from design to production. IEEE Transactions on Professional Communication, PC-23(4):189–196, December 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Krull:1997:WPN

[Kru97] R. Krull. What practitioners need to know to evaluate research. IEEE Transactions on Professional Communication, 40(3):168–181, September 1997. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kruempel:2000:MRI

[Kru00] K. Kruempel. Making the right (interactive) moves for knowledge-producing tasks in computer-mediated groups. IEEE Transactions on Professional Communication, 43 (2):185–195, June 2000. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kryder:1999:MMC

[Kry99] L. G. Kryder. Mentors, models and clients: using the professional engineering community to identify and teach engineering genres. IEEE Transactions on Professional Communication, 42(1):3–11, March 1999. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kryder:2000:TCE

[Kry00a] L. Kryder. Technical communication, 2nd edition [Book Review]. IEEE Transactions on Professional Communication, 43(4):418–419, December 2000. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kryder:2000:GWE

[Kry00b] L. G. Kryder. A Guide To Writing As An Engineer [Book Review]. IEEE Transactions on Professional Communication, 43(2):228–229, June 2000. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

KARREMAN:2004:SES

[KS04] J. Karreman and M. Steehouder. Some effects of system information in instructions for use. IEEE Transactions on Professional Communication, 47(1):34–43, March 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kim:2014:SEP

[KS14] Eun Gyong Kim and Abigail Shin. Seeking an effective program to improve communication skills of non-English-speaking graduate engineering students: The case of a Korean engineering school. IEEE Transactions on Professional Communication, 57(1): 41–55, March 2014. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kramer-Simpson:2015:LBC

[KSNF15] Elisabeth Kramer-Simpson, Julianne Newmark, and Julie Dyke Ford. Learning beyond the classroom and textbook: Client projects role in helping students transition from school to work. IEEE Transactions on Professional Communication, 58(1):106–122, March 2015. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kirkman:1978:CEA

[KSW78] John Kirkman, Christine Snow, and Ian Watson. Controlled English as an alternative to multiple translations. IEEE Transactions on Professional Communication, PC-21 (4):159–161, December 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Katz:2014:RCC

[KT14] Adi Katz and Dov Te’eni. The role of communication complexity in adaptive contextualization. IEEE Transactions on Professional Communication, 57(2):98–112, June 2014. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Khaw:2018:ETI

[KT18] Li Lian Khaw and Wei Wu Tan. Establishing a territory in the introductions of engineering research articles using a problem-solution patterns approach. IEEE Transactions on Professional Communication, 61(2):133–150, June 2018. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Khaw:2020:CCE

[KT20] Li Lian Khaw and Wei Wu Tan. Creating contexts in engineering research writing using a problem-solution-based writing model: Experience of Ph.D. students. IEEE Transactions on Professional Communication, 63(2):155–171, June 2020. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Krahmer:2004:TAT

[KU04] E. Krahmer and N. Ummelen. Thinking about thinking aloud: a comparison of two verbal protocols for usability testing. IEEE Transactions on Professional Communication, 47(2):105–117, June 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kuney:1969:PMI

[Kun69]  J. H. Kuney.  Processing

manuscripts for input to a computerized typesetting system for scientific journals. IEEE Transactions on Engineering Writing and Speech, EWS-12 (2):49–52, May 1969. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Kuney:1973:IMJ

[Kun73] Joseph H. Kuney. Impact of microfilms on journal costs. IEEE Transactions on Professional Communication, PC-16(3):80–81, September 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kuo:1993:TLC

[Kuo93] Chih-Hua Kuo. Teaching lecture comprehension to nonnative science students. IEEE Transactions on Professional Communication, 36(2):70–74, June 1993. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kuriloff:1966:CON

[Kur66] Arthur H. Kuriloff. Communication with the organization — a neglected resource. IEEE Transactions on Engineering Writing and Speech, EWS-9(1): 4–7, July 1966. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Kushner:1997:TNF

[Kus97] S. Kushner. Tackling the needs of foreign academic writers: a case study. IEEE Transactions on Professional Communication, 40(1):20–25, March 1997. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kvan:1994:RCM

[Kva94]  T. Kvan.  Reflections on

computer-mediated architectural design. IEEE Transactions on Professional Communication, 37(4):226–230, December 1994. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Khatri:2006:CFB

[KVRR06] V. Khatri, I. Vessey, S. Ram, and V. Ramesh. Cognitive fit between conceptual schemas and internal problem representations: the case of geospatio-temporal conceptual schema comprehension.

IEEE Transactions on Professional Communication, 49(2): 109–127, June 2006. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kramberg-Walker:1993:NPW

[KW93] C. Kramberg-Walker. The need to provide writing support for academic engineers. IEEE Transactions on Professional Communication, 36(3): 130–136, September 1993. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kowalewski:2023:FAD

[KW23] Scott J. Kowalewski and Bill Williamson. Fostering advocacy, developing empathetic UX bricoleurs: Ongoing programmatic assessment and responsive curriculum design. IEEE Transactions on Professional Communication, 66(4): 382–396, December 2023. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kamel:2017:EGB

[KWLS17] Moataz M. Kamel, Mohamed K. Watfa, Brian Lobo, and Daniel Sobh. Is enterprise gamification being cannibalized by its own brand? IEEE Transactions on Professional Communication, 60 (2):147–164, June 2017. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kase:2010:SIC

[KZCR10] Sue E. Kase, Yang Zhang, John M. Carroll, and Mary Beth Rosson. Synthesizing IT case studies of nonprofits using a multiple-level patterns-based framework. IEEE Transactions on Professional Communication, 53(3):216–232, September 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lacy:1962:ARL

[Lac62] Edward A. Lacy. Abstracts of recent literature. IRE Transactions on Engineering Writing and Speech, EWS-5(2):99, December 1962. ISSN 0893-7869 (print), 2331-3692 (electronic).

Lacy:1963:SDE

[Lac63] Edward A. Lacy. Special dictionaries for the electronics engineer. IEEE Transactions on Engineering Writing and Speech, EWS-6(1):31–35, September 1963. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Lacy:1987:EDS

[Lac87] Edward A. Lacy. Electronics dictionaries: a survey. IEEE Transactions on Professional Communication, PC-30(1):54– 56, March 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lambert:1968:TSA

[Lam68] G. H. Lambert. A time-shared automatic data retrieval and composition system. IEEE Transactions on Engineering Writing and Speech, EWS-11 (2):134–140, August 1968. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Lam:2005:GEC

[Lam05]  C. Lam.  Guide to elec-

tronic communication: Using technology for effective business writing and speaking. IEEE Transactions on Professional Communication, 48(3): 332–333, September 2005. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lam:2014:WDW

[Lam14] Chris Lam. Where did we come from and where are we going? Examining authorship characteristics in technical communication research. IEEE Transactions on Professional Communication, 57(4): 266–285, December 2014. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lam:2016:CAS

[Lam16] Chris Lam. Correspondence analysis: a statistical technique ripe for technical and professional communication researchers. IEEE Transactions on Professional Communication, 59(3):299–310, September 2016. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lam:2018:MTF

[Lam18] Chris Lam. More than a feeling: Applying a data-driven framework in the technical and professional communication team project. IEEE Transactions on Professional Communication, 61(4):409– 427, December 2018. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lantsberg:1994:ISS

[Lan94] H. S. Lantsberg. Introduction to the special section: professional communication in Russia. IEEE Transactions on Professional Communication, 37(2):58–59, June 1994. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lancaster:2004:PPF

[Lan04] A. Lancaster. Paper prototyping: The fast and easy way to design and refine user interfaces. IEEE Transactions on Professional Communication, 47(4):335–336, December 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lancaster:2005:GSE

[Lan05] A. Lancaster. Grant seeking in an electronic age. IEEE Transactions on Professional Communication, 48(1):111– 112, March 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Laviosa:2006:DDL

[Lav06] S. Laviosa. Data-driven learning for translating Anglicisms in business communication. IEEE Transactions on Professional Communication, 49(3): 267–274, September 2006. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lopez-Arroyo:2024:BBB

[LAWI24] Belén López-Arroyo, Xiaobo Wang, and Suguru Ishizaki. Building bridges between technical and professional communication and translation studies. IEEE Transactions on Professional Communication, 67 (3):260–265, September 2024. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Layton:1973:EAR

[Lay73] Edward Layton. Editor-author relationships: Both can win. IEEE Transactions on Professional Communication, PC-16(3):57–59, September 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lay:1988:ITC

[Lay88] M. M. Lay. Images of technology and creativity: in opposition or harmony within technical communications programs? IEEE Transactions on Professional Communication, 31(4): 184–188, December 1988. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lay:1997:WLE

[Lay97] M. M. Lay. Writing Like an Engineer: a Rhetorical Education [Book Revews]. IEEE Transactions on Professional Communication, 40 (1):58–59, March 1997. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lauer:2018:HCC

[LBB18] Claire Lauer, Eva Brumberger, and Aaron Beveridge. Hand collecting and coding versus data-driven methods in technical and professional communication research. IEEE Transactions on Professional Communication, 61(4):389– 408, December 2018. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lang:2023:CAC

[LBE23] Susan Lang, Duncan A. Buell, and Norbert Elliot. Computerassisted corpus analysis: an introduction to concepts, processes, and decisions. IEEE Transactions on Professional Communication, 66(1):94– 113, March 2023. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lowry-Coqrry:1979:BGC

[LC79] Elizabeth Lowry-Coqrry. A basic guide to conference organizing. IEEE Transactions on Professional Communication, PC-22(3):140–145, September 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lee:2017:ESP

[LC17] Chien-Ching Lee and Soo-Fun Chin. Engineering students’ perceptions of graduate attributes: Perspectives from two educational paths. IEEE Transactions on Professional Communication, 60(1): 42–55, March 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lin:2012:TFF

[LCX12] Julian Lin, Hock Chan, and Lingling Xu. A tale of four functions in a multifunctional device: Extending implementation intention theory. IEEE Transactions on Professional Communication, 55(1): 36–54, March 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Leech:1984:HPS

[LD84] Thomas Leech and L. W. Denton. How to prepare, stage, and deliver winning presentations. IEEE Transactions on Professional Communication, PC-27 (4):234, December 1984. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lentz:1997:ETQ

[LD97] L. Lentz and M. De Jong.

The evaluation of text quality: expert-focused and readerfocused methods compared. IEEE Transactions on Professional Communication, 40(3): 224–234, September 1997. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Li:2020:AWD

[LD20] Qiong Li and Jianying Du.

Academics Writing: The Dynamics of Knowledge Creation: Karin Tusting, Sharon Mc-Culloch, Ibrar Bhatt, Mary Hamilton, David Barton [Book Review]. IEEE Transactions on Professional Communication, 63(3):277–278, September 2020. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Li:2020:CDB

[LdJK20] Qian Li, Menno D. T. de Jong, and Joyce Karreman. Cultural differences between Chinese and Western user instructions: a content analysis of user manuals for household appliances. IEEE Transactions on Professional Communication, 63(1):3–20, March 2020. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Londner:2023:CCM

[LDKC23] Edward Londner, Matthew Dabkowski, Ian Kloo, and J. D. Caddell. The communication coefficient method: a new faculty grading tool designed to help engineering students improve their technical communication. IEEE Transactions on Professional Communication, 66(2):202–219, June 2023. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lundin:2018:IMT

[LE18] Jonatan Lundin and Yvonne Eriksson. An investigation of maintenance technicians information-seeking behavior in a repair center. IEEE Transactions on Professional Communication, 61(3):257– 274, September 2018. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

LeComte:2001:TWP

[LeC01] C. LeComte. Technical writing — Process and product, 3rd ed. [Book Review]. IEEE Transactions on Professional Communication, 44(4):306– 307, December 2001. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lee:1962:ESC

[Lee62] LaVern G. Lee. The Engineering Secretary’s Complete Handbook [book review]. IRE Transactions on Engineering Writing and Speech, EWS-5(1): 32, August 1962. ISSN 0893-7869 (print), 2331-3692 (electronic).

Lee:1964:PES

[Lee64] George C. Lee. Preparing effective specifications — a team approach. IEEE Transactions on Engineering Writing and Speech, EWS-7(1):20–21, March 1964. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Lee:2000:RCP

[Lee00] O. Lee. The role of cultural protocol in media choice in a Confucian virtual workplace. IEEE Transactions on Professional Communication, 43(2): 196–200, June 2000. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lee:2013:TEB

[Lee13]  Chien-Ching Lee.  Teach-

ing evidence-based writing using corporate blogs. IEEE Transactions on Professional Communication, 56(3):242– 255, September 2013. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lem:1983:GM

[Lem83] Dean Phillip Lem. Graphics master 2. IEEE Transactions on Professional Communication, PC-26(4):212–213, December 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Levine:1968:VHS

[Lev68] Samuel W. Levine. A very high speed electro-optical-mechanical phototypesetting machine. IEEE Transactions on Engineering Writing and Speech, EWS-11(2):31–35, August 1968. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Levinson:1986:MMC

[Lev86a]  Paul Levinson.  Marshall

Mcluhan and computer conferencing. IEEE Transactions on Professional Communication, PC-29(1):9–11, March 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Levinson:1986:T

[Lev86b] Paul Levinson. Teleconnections. IEEE Transactions on Professional Communication, PC-29(1):8, March 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Leydens:2008:NIP

[Ley08] Jon A. Leydens. Novice and insider perspectives on academic and workplace writing: Toward a continuum of rhetorical awareness. IEEE Transactions on Professional Communication, 51(3):242–263, September 2008. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Laplante:1995:CTC

[LF95] P. Laplante and H. Flaxman.

The convergence of technology and creativity in the corporate environment. IEEE Transactions on Professional Communication, 38(1):20–23, March 1995. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lenton:2008:SME

[LFT08] Alison P. Lenton, Barbara Fasolo, and Peter M. Todd. Shopping for a mate: Expected versus experienced preferences in online mate choice. IEEE Transactions on Professional Communication, 51(2): 169–182, June 2008. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lowry:2014:PLS

[LG14] Paul Benjamin Lowry and James Gaskin. Partial least squares (PLS) structural equation modeling (SEM) for building and testing behavioral causal theory: When to choose it and how to use it. IEEE Transactions on Professional Communication, 57(2):123– 146, June 2014. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lentz:2000:BDD

[LH00] L. Lentz and J. Hulst. Babel in document design: the evaluation of multilingual texts. IEEE Transactions on Professional Communication, 43(3): 313–322, September 2000. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lopez:2023:EMR

[LHAL23] Manuela López, Carmen Hidalgo-Alcázar, and Paul Leger. The effect of message repetition on information diffusion on Twitter: an agent-based approach. IEEE Transactions on Professional Communication, 66 (2):150–169, June 2023. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lowry:2007:SSP

[LHMN07] Paul Benjamin Lowry, Sean LaMarc Humpherys, Jason Malwitz, and Joshua Nix. A scientometric study of the perceived quality of business and technical communication journals. IEEE Transactions on Professional Communication, 50(4): 352–378, December 2007. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Li:1980:CAO

[Li80] Tingye Li. The coming of age of optical-fiber transmission. IEEE Transactions on Professional Communication, PC-23 (3):151–152, September 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lide:1975:CBP

[Lid75] David R. Lide. Cooperation between professional societies and a government agency: The journal of physical and chemical reference data. IEEE Transactions on Professional Communication, PC-18(3): 127–129, September 1975. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lieb:1977:JRA

[Lie77] Charles H. Lieb. Journalrelated aspects of the new copyright law. IEEE Transactions on Professional Communication, PC-20(3):155–159, November 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lie:2018:EAL

[Lie18] Ting Lie. The effect of active learning strategies on communication apprehension in information systems students in Taiwan. IEEE Transactions on Professional Communication, 61(1):101–109, March 2018. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lies:2020:SWU

[Lie20] Erica Lies. Strategic Writing for UX: Drive Engagement, Conversion, and Retention With Every Word: Torrey Podmajersky [Book Review]. IEEE Transactions on Professional Communication, 63(4): 404–405, December 2020. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lin:1972:CBS

[Lin72] Nan Lin. A comparison between the scientific communication model and the mass communication model: Implications for the transfer and utilization of scientific knowledge: a PSYCHOM ’72 paper. IEEE Transactions on Professional Communication, PC-15 (2):34–39, June 1972. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Linden:1977:SPC

[Lin77a] Bella L. Linden. Social purpose of copyright legislation. IEEE Transactions on Professional Communication, PC-20 (3):147–149, November 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lineback:1977:JQH

[Lin77b] Richard H. Lineback. Journal quality in the humanities. IEEE Transactions on Professional Communication, PC-20 (2):129–130, September 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Linver:1980:ISI

[Lin80] Sandy Linver. The importance of the spoken image. IEEE Transactions on Professional Communication, PC-23 (1):8–10, March 1980. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lindeborg:1990:FMP

[Lin90a]  R. A. Lindeborg.  Faster

manuscript publishing: a case study. IEEE Transactions on Professional Communication, 33(1):7–11, March 1990. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lindeborg:1990:QES

[Lin90b] R. A. Lindeborg. A quick and easy strategy for organizing a speech. IEEE Transactions on Professional Communication, 33(3):133–136, September 1990. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lindeborg:1993:IEM

[Lin93] R. A. Lindeborg. The irresistible electronic message of the 1990s: a case study. IEEE Transactions on Professional Communication, 36(3): 152–157, September 1993. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lind:1999:GIT

[Lin99] M. R. Lind. The gender impact of temporary virtual work groups. IEEE Transactions on Professional Communication, 42(4):276–285, December 1999. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lingaiah:2004:MGB

[Lin04] D. Lingaiah. The Manager’s Guide to Business Writing: Business Writing Simplified. IEEE Transactions on Professional Communication, 47(1): 80–81, March 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lin:2007:OWD

[Lin07] Canchu Lin. Organizational Website design as a rhetorical situation. IEEE Transactions on Professional Communication, 50(1):35–44, March 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lin:2019:LNA

[Lin19] Yuting Lin. Legitimating negative aspects in corporate social responsibility reporting: Evidence from China. IEEE Transactions on Professional Communication, 62(3):263– 278, September 2019. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lippke:1958:WED

[Lip58] James A. Lippke. Writing for “electronic design”. IRE Transactions on Engineering Writing and Speech, EWS-1(1): 26–27, March 1958. ISSN 0893-7869 (print), 2331-3692 (electronic).

Lippmann:1971:ACA

[Lip71] Erhard O. Lippmann. An approach to computer-aided translation. IEEE Transactions on Engineering Writing and Speech, EWS-14(1):10–33, February 1971. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Lippmann:1975:LGT

[Lip75] Erhard O. Lippmann. Online generation of terminological digests in language translation. An aid in terminology processing. IEEE Transactions on Professional Communication, PC-18(4):309–319, December 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lippincott:1983:STC

[Lip83]  H. F. Lippincott.  Some

tips for clear writing. IEEE Transactions on Professional Communication, PC-26(1):11– 12, March 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lippincott:2004:GMW

[Lip04] G. Lippincott. Gray matters: where are the technical communicators in research and design for aging audiences? IEEE Transactions on Professional Communication, 47(3): 157–170, September 2004. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Little:1990:PTC

[Lit90] S. B. Little. Preparing the technical communicator of the future. IEEE Transactions on Professional Communication, 33(1):28–31, March 1990. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lantz-Jones:2010:GES

[LJ10] Amy Lantz-Jones. The Global English Style Guide: Writing Clear, Translatable Documentation for a Global Market (Kohl, J. R.) [Book Review]. IEEE Transactions on Professional Communication, 53(3): 325–326, September 2010. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lufkin:1972:CBI

[LK72] James M. Lufkin and Steven C. Krantz. Cultural barriers to interprofessional communication: a PSYCHOM ’72 paper. IEEE Transactions on Professional Communication, PC-15(2):26– 29, June 1972. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lee-Kim:2008:FPM

[LK08] Julia Lee-Kim. Fundamentals of Project Management, 3rd ed. (Lewis, J. P.; 2006) [Book Review]. IEEE Transactions on Professional Communication, 51(3):341–342, September 2008. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lucas:2016:CEP

[LKHC16] Kristen Lucas, Sharon A.

Kerrick, Jenna Haugen, and Cole J. Crider. Communicating entrepreneurial passion: Personal passion vs. perceived passion in venture pitches. IEEE Transactions on Professional Communication, 59(4): 363–378, December 2016. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Loorbach:2013:VSP

[LKS13] Nicole Loorbach, Joyce Karreman, and Michaël Steehouder. Verification steps and personal stories in an instruction manual for seniors: Effects on confidence, motivation, and usability. IEEE Transactions on Professional Communication, 56 (4):294–312, December 2013.

CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Levis:2003:PBA

[LL03] J. M. Levis and G. M. Levis. A project-based approach to teaching research writing to nonnative writers. IEEE Transactions on Professional Communication, 46(3):210– 220, September 2003. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lee:2009:EBC

[LL09a] Zoonky Lee and Younghwa Lee. Emailing the boss: Cultural implications of media choice. IEEE Transactions on Professional Communication, 52(1):61–74, March 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Leydens:2009:LMD

[LL09b] Jon A. Leydens and Juan C. Lucena. Listening as a missing dimension in engineering education: Implications for sustainable community development efforts. IEEE Transactions on Professional Communication, 52(4):359– 376, December 2009. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lu:2018:HDE

[LLY18] Ming-Hsin Lu, Weijane Lin, and Hsiu-Ping Yueh. How do employees in different job roles in the insurance industry use mobile technology differently at work? IEEE Transactions on Professional Communication, 61(2):151–165, June 2018. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lincoln:1997:DSG

[LM97] J. E. Lincoln and D. L. Monk.

Displaying scientific graphics on computer. IEEE Transactions on Professional Communication, 40(2):78–91, June 1997. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

LaRussa-Martin:2002:TCE

[LM02] C. LaRussa-Martin. Technical communication (8th ed.) — book review. IEEE Transactions on Professional Communication, 45(1):63–64, March 2002. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Laursen:2014:UAH

[LMJK14] Anne Lise Laursen, Birthe Mousten, Vigdis Jensen, and Constance Kampf. Using an AD-HOC corpus to write about emerging technologies for technical writing and translation: The case of search engine optimization. IEEE Transactions on Professional Communication, 57(1):56–74, March 2014. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lozano-Nieto:1998:ATD

[LN98] A. Lozano-Nieto. The Art of Technical Documentation, 2nd Ed. [Book Review]. IEEE Transactions on Professional Communication, 41(4):289– 290, December 1998. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lozano-Nieto:2000:WWH

[LN00] A. Lozano-Nieto. Web wisdom: how to evaluate and create information quality on the web [Book Review]. IEEE Transactions on Professional Communication, 43(3):341– 342, September 2000. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lowry:2003:UIB

[LN03] P. B. Lowry and J. F. Nunamaker. Using Internet-based, distributed collaborative writing tools to improve coordination and group awareness in writing teams. IEEE Transactions on Professional Communication, 46(4):277– 297, December 2003. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lozano-Nieto:2004:EEE

[LN04]  A. Lozano-Nieto.  Electri-

cal and electronics engineering dictionary. IEEE Transactions on Professional Communication, 47(4):337, December 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lowry:2004:CHD

[LNB^{+}04] P. B. Lowry, J. F. Nunamaker, Q. E. Booker, A. Curtis, and M. R. Lowry. Creating hybrid distributed learning environments by implementing distributed collaborative writing in traditional educational settings. IEEE Transactions on Professional Communication, 47(3):171–189, September 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lowry:2005:IPS

[LNCL05] P. B. Lowry, J. F. Nunamaker, A. Curtis, and M. R. Lowry. The impact of process structure on novice, virtual collaborative writing teams. IEEE Transactions on Professional Communication, 48(4): 341–364, December 2005. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Leydens:2007:PSE

[LO07] J. A. Leydens and B. M. Olds.

Publishing in scientific and engineering contexts: a course for graduate students tutorial. IEEE Transactions on Professional Communication, 50(1): 45–56, March 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lauer:2020:HPI

[LO20] Claire Lauer and Shaun O’Brien. How people are influenced by deceptive tactics in everyday charts and graphs. IEEE Transactions on Professional Communication, 63(4): 327–340, December 2020. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lockinger:2022:TCU

[Löc22] Georg Löckinger. Technical communicators’ use of and requirements for special language reference tools. IEEE Transactions on Professional Communication, 65(2):294–320, June 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Loebenstein:1968:LSL

[Loe68] Julian Loebenstein. List of specifications for letter and graphic symbols, definitions and abbreviations for discrete semiconductor devices. IEEE Transactions on Engineering Writing and Speech, EWS-11 (1):15–16, July 1968. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Loehr:1991:BSV

[Loe91] L. Loehr. Between silence and voice: communicating in cross-functional project teams. IEEE Transactions on Professional Communication, 34(1): 51–56, March 1991. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Loehr:1995:CGC

[Loe95]  L. Loehr.  Composing in

groups: the concept of authority in cross functional project team work. IEEE Transactions on Professional Communication, 38(2):83–94, June 1995. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lofstrom:2008:GIP

[Lof08] Joyce Lofstrom. Gaining Influence in Public Relations: The Role of Resistance in Practice (Berger, B. K. and Reber, B. H.; 2006) [Book Review]. IEEE Transactions on Professional Communication, 51(1): 114–115, March 2008. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lofstrom:2010:CAM

[Lof10] Joyce Lofstrom. A collaborative approach for media training between technical communication and public relations tutorial. IEEE Transactions on Professional Communication, 53(2):164–173, June 2010.

CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Longo:1995:CCE

[Lon95] B. Longo. Corporate culture and its effect on audience awareness. IEEE Transactions on Professional Communication, 38(4):239–243, December 1995. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Longo:2014:RUT

[Lon14] Bernadette Longo. R U there? Cell phones, participatory design, and intercultural dialogue. IEEE Transactions on Professional Communication, 57(3):204–215, September 2014. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lopinto:1984:DWO

[Lop84]  Lidia Lopinto.  Designing

and writing operating manuals. IEEE Transactions on Professional Communication, PC-27 (1):29–31, March 1984. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Losey:1985:EMS

[Los85] Cheryl L. Losey. Electronic messaging systems for more effective management. IEEE Transactions on Professional Communication, PC-28(3):35– 39, September 1985. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Loughren:1958:PDS

[Lou58] Arthur V. Loughren. Panel discussion — Should a talk be read from a prepared manuscript? IRE Transactions on Engineering Writing and Speech, EWS-1(1):14–17, March 1958. ISSN 0893-7869 (print), 2331-3692 (electronic).

Lauren:2016:NFI

[LP16] Benjamin Lauren and Stacey Pigg. Networking in a field of introverts: The egonets, networking practices, and networking technologies of technical communication entrepreneurs. IEEE Transactions on Professional Communication, 59(4):342–362, December 2016. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

LoMaglio:1985:IPV

[LR85] Larry J. LoMaglio and Victoria J. Robinson. The impact of passive voice on reading comprehension. IEEE Transactions on Professional Communication, PC-28(4):26– 27, December 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Labrador:2024:PET

[LR24] Belén Labrador and Noelia Ramón. Positive evaluation in the translation of online promotional discourse in the cheese industry. IEEE Transactions on Professional Communication, 67(3):316– 332, September 2024. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Li:2017:EDC

[LRLM17] Ye Li, Pei-Luen Patrick Rau, Hui Li, and Alexander Maedche. Effects of a dyad’s cultural intelligence on global virtual collaboration. IEEE Transactions on Professional Communication, 60(1):56–75, March 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Laplante:2009:QCP

[LRM^{+}09] Phil Laplante, Jon Rockne, Paolo Montuschi, Thomas Baldwin, Mike Hinchey, Linda Shafer, Jeffrey Voas, and Wenping Wang. Quality in conference publishing. IEEE Transactions on Professional Communication, 52(2):183– 196, June 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Laird:1961:ML

[LS61] Chariton Laird and Chester W. Sall. The miracle of language. IRE Transactions on Engineering Writing and Speech, EWS-4(2):62, May 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Lehman:1984:WII

[LS84] Anita J. Lehman and David Strong. Writing for industry: an instruction manual. IEEE Transactions on Professional Communication, PC-27 (4):232, December 1984. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lakayev:1994:RDH

[LS94] A. S. Lakayev and M. M. Subbotin. Research and development of hypertext and multimedia systems in Russia. IEEE Transactions on Professional Communication, 37(2):73–74, June 1994. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lauer:2023:WET

[LS23] Claire Lauer and Christopher A. Sanchez. What eye tracking can show us about how people are influenced by deceptive tactics in line graphs. IEEE Transactions on Professional Communication, 66(3): 220–235, September 2023. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lubetkin:1980:HS

[Lub80] Maurice Lubetkin. How to speak up or down. IEEE Transactions on Professional Communication, PC-23(1):43– 44, March 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lufkin:1964:IEC

[Luf64] James M. Lufkin. The implications for engineers of C. P. Snow’s Two Cultures. IEEE Transactions on Engineering Writing and Speech, EWS-7(3): 8–15, December 1964. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Lufkin:1965:HVM

[Luf65] James M. Lufkin. Human vs machine translation of foreign languages. IEEE Transactions on Engineering Writing and Speech, EWS-8(1):8– 14, June 1965. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Lufkin:1968:STT

[Luf68] James M. Lufkin. The slide talk: a tutorial drama in one act. IEEE Transactions on Engineering Writing and Speech, EWS-11(1):7–14, July 1968. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Lufkin:1969:CCT

[Luf69a]  James M. Lufkin.  The

coauthors: a comico-tragico-tutorial drama in one act. IEEE Transactions on Engineering Writing and Speech, EWS-12(3):83–88, October 1969. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Lufkin:1969:WES

[Luf69b] James M. Lufkin. What everybody should know about translation. IEEE Transactions on Engineering Writing and Speech, EWS-12(1):3–9, May 1969. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Lufkin:1971:FAT

[Luf71] James M. Lufkin. The fatal abstract: a tutorial farce in one act. IEEE Transactions on Engineering Writing and Speech, EWS-14(1):3–9, February 1971. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Lufkin:1972:GIS

[Luf72] James M. Lufkin. Generalization and the interpretation of science and technology. IEEE Transactions on Professional Communication, PC-15 (4):108–111, December 1972. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lufkin:1973:BPB

[Luf73] James M. Lufkin. Barbara’s polar bears. IEEE Transactions on Professional Communication, PC-16(2):27–34, June 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lufkin:1975:WES

[Luf75] James M. Lufkin. What everybody should know about translation. IEEE Transactions on Professional Communication, PC-18(1):22–28, March 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lufkin:1976:GBC

[Luf76] James Lufkin. The gulf between correctness and understanding. IEEE Transactions on Professional Communication, PC-19(1):4–6, March 1976. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lulis:2006:GAC

[Lul06] E. Lulis. Genre across the curriculum. IEEE Transactions on Professional Communication, 49(2):206–207, June 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lutz:1984:SRE

[Lut84] Jean A. Lutz. A study of revising and editing at the terminal. IEEE Transactions on Professional Communication, PC-27 (2):73–77, June 1984. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Luzon:2005:GAT

[Luz05] M. J. Luzon. Genre analysis in technical communication. IEEE Transactions on Professional Communication, 48(3): 285–295, September 2005. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lee:2007:ERB

[LWMR07] Chei Sian Lee, Mary Beth Watson-Manheim, and Arkalgud Ramaprasad. Exploring the relationship between communication risk perception and communication portfolio. IEEE Transactions on Professional Communication, 50(2):130– 146, June 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Leininger:1998:AIE

[LY98] C. Leininger and Rue Yuan.

Aligning international editing efforts with global business strategies. IEEE Transactions on Professional Communication, 41(1):16–23, March 1998. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lybbert:1978:MPR

[Lyb78] E. K. Lybbert. Making policy readable. IEEE Transactions on Professional Communication, PC-21(4):152–154, December 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Lynch:1983:DEI

[Lyn83]  Gary Lynch.  Document-

ing engineering instrumentation projects. IEEE Transactions on Professional Communication, PC-26(4):178–183, December 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

MacAskill:1961:PEP

[Mac61] Robert B. MacAskill. Persuasion in engineering proposals. IRE Transactions on Engineering Writing and Speech, EWS-4(2):56–57, May 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

MacAdam:1968:TGN

[Mac68] Walter K. MacAdam. Technology’s great need: Information retrieval. IEEE Transactions on Engineering Writing and Speech, EWS-11(1):3– 6, July 1968. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Macek:1972:EQE

[Mac72] A. J. Macek. The evolution of a questionnaire for engineers: a PSYCHOM ’72 paper. IEEE Transactions on Professional Communication, PC-15(3):80– 84, September 1972. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Macadam:1975:TGN

[Mac75a] Walter K. Macadam. Technology’s great need: Information retrieval. IEEE Transactions on Professional Communication, PC-18(1):50–53, March 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Mack:1975:LCC

[Mac75b] Paul F. Mack. Lower composition costs through optical scanning and photocomposition. IEEE Transactions on Professional Communication, PC-18 (3):279–282, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Macgregor:1978:PPT

[Mac78a] A. J. Macgregor. Preparing poster talks. IEEE Transactions on Professional Communication, PC-21(3):103–105, September 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Macgregor:1978:SAC

[Mac78b]  A. J. Macgregor.  Select-

ing the appropriate chart. IEEE Transactions on Professional Communication, PC-21 (3):106–107, September 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

MacGregor:1982:SAC

[Mac82]  A. Jean MacGregor.  Se-

lecting the appropriate chart. IEEE Transactions on Professional Communication, PC-25 (2):102–104, June 1982. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

MacNealy:1990:MTM

[Mac90] M. S. MacNealy. Moving toward maturity: research in technical communication. IEEE Transactions on Professional Communication, 33(4): 197–204, December 1990. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

MacNealy:1997:TBC

[Mac97] M. S. MacNealy. Toward better case study research. IEEE Transactions on Professional Communication, 40(3): 182–196, September 1997. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Macari:1999:MCP

[Mac99a] N. Macari. A memorandum conundrum [professional communication]. IEEE Transactions on Professional Communication, 42(3):178–180, September 1999. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Mackiewicz:1999:GPW

[Mac99b] J. Mackiewicz. Gauging prescriptivism in writing handbooks. IEEE Transactions on Professional Communication, 42(1):55–58, March 1999. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Mackiewicz:2002:PTW

[Mac02a] J. Mackiewicz. Professional and technical writing strategies: communicating in technology and science, 5th edition [Book Review]. IEEE Transactions on Professional Communication, 45(3):203–205, September 2002. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Mackiewicz:2002:RTI

[Mac02b]  J. Mackiewicz.  Reporting

technical information — 10th ed. [Book Review]. IEEE Transactions on Professional Communication, 45(4):314– 316, December 2002. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Mackiewicz:2003:TTB

[Mac03a] J. Mackiewicz. Type and typography [Book Review]. IEEE Transactions on Professional Communication, 46(3):246– 247, September 2003. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Mackiewicz:2003:WRO

[Mac03b] J. Mackiewicz. Which rules for online writing are worth following?: a study of eight rules in eleven handbooks. IEEE Transactions on Professional Communication, 46(2): 129–137, June 2003. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Mackiewicz:2004:CWT

[Mac04a] J. Mackiewicz. Central works in technical communication.

IEEE Transactions on Professional Communication, 47(4): 331–332, December 2004. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Mackiewicz:2004:ETE

[Mac04b] J. Mackiewicz. The effects of tutor expertise in engineering writing: a linguistic analysis of writing tutors’ comments. IEEE Transactions on Professional Communication, 47(4): 316–328, December 2004. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Mackiewicz:2004:WPC

[Mac04c] J. Mackiewicz. Writing power: Communication in an engineering center. IEEE Transactions on Professional Communication, 47(2):155–156, June 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Mackiewicz:2005:HWT

[Mac05a] J. Mackiewicz. Hinting at what they mean: indirect suggestions in writing tutors’ interactions with engineering students. IEEE Transactions on Professional Communication, 48(4):365–376, December 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Mackiewicz:2005:UED

[Mac05b] J. Mackiewicz. Use and effect of declarative information in user instructions. IEEE Transactions on Professional Communication, 48(1):105– 106, March 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Mackiewicz:2006:FFN

[Mac06] J. Mackiewicz. The functions of formulaic and nonformulaic compliments in interactions about technical writing. IEEE Transactions on Professional Communication, 49(1): 12–27, March 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Magnan:1977:TDI

[Mag77] George A. Magnan. Technical drawings and illustrations. IEEE Transactions on Professional Communication, PC-20 (4):239–245, December 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Magyar:2003:DJK

[Mag03] M. Magyar. Dan Jones and Karen Lane, Technical communication: strategies for college and the workplace [Book Review]. IEEE Transactions on Professional Communication, 46(4):344–345, December 2003. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Maier:1970:GIR

[Mai70] Robert J. Maier. Glossary of information-retrieval terminology. IEEE Transactions on Engineering Writing and Speech, EWS-13(2):34–36, September 1970. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Makris:1968:AS

[Mak68] Constantine J. Makris. The ARCOL/System. IEEE Transactions on Engineering Writing and Speech, EWS-11(2):35– 44, August 1968. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Mehlenbacher:2015:IDS

[MAK15] Brad Mehlenbacher, Meagan Kittle Autry, and Ashley Rose Kelly. Instructional design for stem-based collaborative, colocated classroom composition. IEEE Transactions on Professional Communication, 58(4):396– 409, December 2015. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Malcolm:1985:HCC

[Mal85a] Andrew Malcolm. How closed captions work. IEEE Transactions on Professional Communication, PC-28(4):19–20, December 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Malcolm:1985:WDR

[Mal85b]  Andrew Malcolm.  Writ-

ing for the disadvantaged reader: an introduction. IEEE Transactions on Professional Communication, PC-28(4):1– 2, December 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Malcolm:1987:RFC

[Mal87] Andrew Malcolm. Recall of flashed capitalized video print: a study of prelingually deaf science and engineering college students. IEEE Transactions on Professional Communication, PC-30(1):50–53, March 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Malkinson:2001:GPW

[Mal01] T. J. Malkinson. Get to the point! Writing effective email, letters, reports and proposals [Book Review]. IEEE Transactions on Professional Communication, 44(3):214– 215, September 2001. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Malone:2007:HST

[Mal07] Edward A. Malone. Historical studies of technical communication in the United States and England: a fifteen-year retrospection and guide to resources. IEEE Transactions on Professional Communication, 50(4):333–351, December 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Malone:2011:DSC

[Mal11] Edward A. Malone. On the Dot: The Speck that Changed the World (Humez, A. and Humez, N.; 2008) [Book Review]. IEEE Transactions on Professional Communication, 54(1):100–101, March 2011. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Manheim:1975:SR

[Man75] Frank T. Manheim. The scientific referee. IEEE Transactions on Professional Communication, PC-18(3):190–195, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Manschreck:1986:SDR

[Man86] Theo C. Manschreck. Schizophrenic disorders and reading disturbance. IEEE Transactions on Professional Communication, PC-29(2):27–30, June 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Mann:1998:AFC

[Man98a] S. Mann. Achieving frontline communication excellence: the potential cost to health. IEEE Transactions on Professional Communication, 41(4): 254–265, December 1998. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Manning:1998:MBC

[Man98b] A. D. Manning. Minimalism beyond computer documentation. IEEE Transactions on Professional Communication, 41(3):200–204, September 1998. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Manning:1998:UCI

[Man98c] A. D. Manning. Understanding comics: The invisible art. IEEE Transactions on Professional Communication, 41(1): 66–69, March 1998. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Manning:1999:DSC

[Man99a] A. D. Manning. Dynamic and static communication in electronic media. IEEE Transactions on Professional Communication, 42(4):301–305, December 1999. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Manning:1999:EGT

[Man99b] A. D. Manning. Error and the growth of technical understanding. IEEE Transactions on Professional Communication, 42(2):123–127, June 1999. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Manning:2001:WE

[Man01] A. D. Manning. What is explanation? IEEE Transactions on Professional Communication, 44(1):53–57, March 2001. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Manning:2002:GI

[Man02] A. D. Manning. The grammar instinct. IEEE Transactions on Professional Communication, 45(2):133–137, June 2002. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Manz:2018:ECM

[Man18] Thomas A. Manz. The Engineering Communication Manual (House, R., et al) [Book Review]. IEEE Transactions on Professional Communication, 61(2):221–222, June 2018. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Marcott:1964:CSA

[Mar64] Creighton M. Marcott. The changing state-of-the-art in trade magazines. IEEE Transactions on Engineering Writing and Speech, EWS-7(3):23– 24, December 1964. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Martinez:1970:ISP

[Mar70] Samuel J. Martinez. An information system for petroleum engineering technology. IEEE Transactions on Engineering Writing and Speech, EWS-13(2):60–64, September 1970. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Marshek:1975:TDW

[Mar75] Kurt M. Marshek. Transitional devices for the writer. IEEE Transactions on Professional Communication, PC-18 (4):320–322, December 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Marks:1977:PCP

[Mar77] Robert H. Marks. Publication charge plan: Wider circulation at lower subscription rates. IEEE Transactions on Professional Communication, PC-20 (2):114–117, September 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Marcus:1982:DFI

[Mar82] Aaron Marcus. Designing the face of an interface. IEEE Transactions on Professional Communication, PC-25(3): 122–126, September 1982. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Martin:1984:ETC

[Mar84] Peter Martin. Ergonomics in technical communication.

IEEE Transactions on Professional Communication, PC-27 (2):62–64, June 1984. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Marcus:1985:HMD

[Mar85] Stephen Marcus. The host in the machine: Decorum in computers who speak. IEEE Transactions on Professional Communication, PC-28(2):29– 33, June 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Martin:1986:CUD

[Mar86]  Peter Martin.  Computer

user documentation problems: Their causes and solutions. IEEE Transactions on Professional Communication, PC-29(4):19–20, December 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Marra:1989:TPE

[Mar89a] F. J. Marra. Theoretical perspectives on emergency communications. IEEE Transactions on Professional Communication, 32(2):94–100, June 1989. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Martin:1989:SD

[Mar89b] M. Martin. The semiology of documents. IEEE Transactions on Professional Communication, 32(3):171–177, September 1989. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Markel:1993:EIT

[Mar93] M. Markel. An ethical imperative for technical communicators. IEEE Transactions on Professional Communication, 36(2):81–86, June 1993. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic). See comments [Dav93].

Markel:1997:ETC

[Mar97] M. Markel. Ethics and technical communication: a case for foundational approaches. IEEE Transactions on Professional Communication, 40(4): 284–298, December 1997. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Markel:2002:DLE

[Mar02] M. Markel. Deep linking:

an ethical and legal analysis. IEEE Transactions on Professional Communication, 45(2): 77–83, June 2002. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Markel:2006:SHP

[Mar06] M. Markel. Safe harbor and privacy protection: a looming issue for IT professionals. IEEE Transactions on Professional Communication, 49(1): 1–11, March 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Marsen:2008:PWC

[Mar08] Sky Marsen. Professional writing: The complete guide for business, industry, and IT, 2nd ed. IEEE Transactions on Professional Communication, 51 (4):418–419, December 2008. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Masse:1984:TPW

[Mas84]  Roger E. Masse.  Theory

and practice of writing processes for technical writers. IEEE Transactions on Professional Communication, PC-27 (4):185–192, December 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Masse:1985:TPE

[Mas85] Roger E. Masse. Theory and practice of editing processes in technical communication. IEEE Transactions on Professional Communication, PC-28 (1):34–42, March 1985. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Mastarone:2008:HRM

[Mas08] Ginnifer L. Mastarone. E-Human Resources Management: Managing Knowledge People (Torres-Coronas, T. and Arias-Oliva, M., Eds.; 2005) [Book Review]. IEEE Transactions on Professional Communication, 51(3):343– 344, September 2008. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Matthews:1977:MLS

[Mat77] Paul A. Matthews. Multilanguage service literature program. IEEE Transactions on Professional Communication, PC-20(4):225–228, December 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Matricciani:1991:PDA

[Mat91] E. Matricciani. The probability distribution of the age of references in engineering papers. IEEE Transactions on Professional Communication, 34(1): 7–12, March 1991. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Maugh:1979:HFI

[Mau79] Thomas H. Maugh. Holographic filing: an industry on the verge of birth. IEEE Transactions on Professional Communication, PC-22(1):34– 36, March 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Maxwell:1973:SVTa

[Max73a] Robert Maxwell. Survival values in technical journals. IEEE Transactions on Professional Communication, PC-16(3):64– 65, September 1973. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Maxwell:1973:SVTb

[Max73b]  Robert Maxwell.  Survival

values in technical journals. IEEE Transactions on Professional Communication, PC-16 (4):204, December 1973. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Maxwell:1988:TST

[Max88] W. Maxwell. Technical with a small ’t’: technology and the technical communication major. IEEE Transactions on Professional Communication, 31(4):189–190, December 1988. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Maynard:1979:HRP

[May79] John T. Maynard. How to read a patent. IEEE Transactions on Professional Communication, PC-22(2):112–118, June 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Maynard:1982:UDA

[May82]  John Maynard.  A user-

driven approach to better user manuals. IEEE Transactions on Professional Communication, PC-25(1):16–19, March 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

May:2019:YGT

[May19] Jamie May. YouTube gamers and think-aloud protocols: Introducing usability testing. IEEE Transactions on Professional Communication, 62(1): 94–103, March 2019. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Marcus:1984:SRI

[MB84] Stephen Marcus and Sheridan Blau. Not seeing is relieving: Invisible writing with computers. IEEE Transactions on Professional Communication, PC-27(2):69–72, June 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Morris:1975:ACT

[MC75a] Robert Morris and Lorinda L.

Cherry. Appendix A. Common technical English words, and Appendix B. Possible typo’s and spelling errors. IEEE Transactions on Professional Communication, PC-18(1):57– 64, March 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic). See [MC75b].

Morris:1975:CDT

[MC75b] Robert Morris and Lorinda L.

Cherry. Computer detection of typographical errors. IEEE Transactions on Professional Communication, PC-18 (1):54–56, March 1975. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic). See supplemental word list [MC75a].

McAlister:1985:BYC

[McA85] James McAlister. Build your case to management. IEEE Transactions on Professional Communication, PC-28(3):21– 24, September 1985. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Mandel:1962:PIW

[MCC62] Siegfried Mandel, David L. Caldwell, and William E. Collins. Proposal and inquiry writing — analysis, techniques, practice. IRE Transactions on Engineering Writing and Speech, EWS-5(1):31–32, August 1962. ISSN 0893-7869 (print), 2331-3692 (electronic).

McCreary:1963:CWI

[McC63] Bruce McCreary. Clue-word indexing of technical information in a research laboratory. IEEE Transactions on Engineering Writing and Speech, EWS-6(1): 28–30, September 1963. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

McClure:1965:TBT

[McC65] Lucille McClure. Two bibliographies: Technical writing books in print; phototypesetting. IEEE Transactions on Engineering Writing and Speech, EWS-8(2):67–70, December 1965. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

McCanless:1979:EP

[McC79] George F. McCanless. The engineer and the paragraph. IEEE Transactions on Professional Communication, PC-22 (4):198–199, December 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

McClure:1987:RFU

[McC87] Glenda M. McClure. Readability formulas: Useful or useless? IEEE Transactions on Professional Communication, PC-30 (1):12–15, March 1987. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

McCord:1992:MDL

[McC92] E. A. McCord. Multiple drafts and legal liability: a hazard for professional writers. IEEE Transactions on Professional Communication, 35(3): 138–142, September 1992. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

McCoy:1994:TSI

[McC94]  J. D. McCoy.  Technical

students investigate technical proposal development. IEEE Transactions on Professional Communication, 37 (1):39–41, March 1994. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

McCool:2006:ALJ

[McC06]  M. McCool.  Adapting e-

learning for Japanese audiences tutorial. IEEE Transactions on Professional Communication, 49(4):335–345, December 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

McClellan:2019:JHV

[McC19] John McClellan. James A. Herrick: Visions of Technological Transcendence: Human Enhancement and Rhetoric of the Future [Book Review]. IEEE Transactions on Professional Communication, 62(1): 104–106, March 2019. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

McDaniel:1994:RCR

[McD94] B. McDaniel. The role of contrastive rhetoric in teaching professional communication in English as a second or foreign language. IEEE Transactions on Professional Communication, 37(1):29–33, March 1994. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

McElwee:1958:WDT

[McE58] Eleanor M. McElwee. What did they say? IRE Transactions on Engineering Writing and Speech, EWS-1(2):10– 12, August 1958. ISSN 0893-7869 (print), 2331-3692 (electronic).

McElwee:1961:RCR

[McE61] Eleanor M. McElwee. Russell Conwell Ross, Speak with Ease. IRE Transactions on Engineering Writing and Speech, EWS-4(3):102, December 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

McElroy:2018:AAB

[McE18] Stephen J. McElroy. Assembling Arguments (Buehl, J.) [Book Review]. IEEE Transactions on Professional Communication, 61(1):112–113, March 2018. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

McGrath:1972:CTI

[McG72] Edward F. McGrath. Communication theory and intersubjective understanding: a PSYCHOM ’72 paper. IEEE Transactions on Professional Communication, PC-15(2):23– 26, June 1972. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

McGraw:1986:GPD

[McG86] Karen L. McGraw. Guidelines for producing documentation for expert systems. IEEE Transactions on Professional Communication, PC-29(4):42– 47, December 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

McGee:2000:CCU

[McG00] L. McGee. Communication channels used by technical writers throughout the documentation process. IEEE Transactions on Professional Communication, 43(1):35– 50, March 2000. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

McKeehan:1966:IEC

[McK66] John H. McKeehan. The inarticulate engineer (a communications dilemma). IEEE Transactions on Engineering Writing and Speech, EWS-9(2):48– 50, December 1966. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

McMillan:1961:HWO

[McM61] Brockway McMillan. How to write officialese. IRE Transactions on Engineering Writing and Speech, EWS-4(2):42– 43, May 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

McNamara:1958:IBT

[McN58a] Daniel J. McNamara. I’ve been thinking about PGEWS. IRE Transactions on Engineering Writing and Speech, EWS-1(2):26–28, August 1958. ISSN 0893-7869 (print), 2331-3692 (electronic).

McNamara:1958:W

[McN58b] Daniel J. McNamara. Welcome. IRE Transactions on Engineering Writing and Speech, EWS-1(1):1, March 1958. ISSN 0893-7869 (print), 2331-3692 (electronic).

Michaelis:1977:WUI

[MCWK77] Paul Roller Michaelis, Alphonse Chapanis, Gerald D. Weeks, and Michael J. Kelly. Word usage in interactive dialog with restricted and unrestricted vocabularies. IEEE Transactions on Professional Communication, PC-20(4):214–221, December 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

McKim:1983:TV

[MD83] Robert H. McKim and L. W.

Denton. Thinking visually. IEEE Transactions on Professional Communication, PC-26 (2):85–86, June 1983. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Montalbo:1984:PEM

[MD84] Thomas Montalbo and L. W.

Denton. The power of eloquence: Magic key to success in public speaking. IEEE Transactions on Professional Communication, PC-27(4): 234, December 1984. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Mehlenbacher:1994:RNA

[Meh94] B. Mehlenbacher. The rhetorical nature of academic research funding. IEEE Transactions on Professional Communication, 37(3):157–162, September 1994. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Mektz:1972:MTH

[Mek72]  Harvey W. Mektz.  Miss

Thistlebottom’s Hobgoblins, Theodore M. Bernstein (New York: Farrar, Strauss & Giroux, 1971, 253 pp. plus index, $6.95). IEEE Transactions on Professional Communication, PC-15(1):17, March 1972. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Melton:2008:LTP

[Mel08] James H. Melton. Lost in translation: Professional communication competencies in global training contexts. IEEE Transactions on Professional Communication, 51(2):198– 214, June 2008. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Melton:2009:GGC

[Mel09]  James H. Melton.  Going

global: a case study of rhetorical invention, packaging, delivery, and feedback collection. IEEE Transactions on Professional Communication, 52(3): 229–242, September 2009. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Memmolo:1977:MMP

[Mem77] John A. Memmolo. Marketing mix for publishing profit. IEEE Transactions on Professional Communication, PC-20 (2):117–120, September 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Menzel:1972:UAI

[Men72] Herbert Menzel. Unplanned acquisition of information in the experience of polymer chemists: a PSYCHOM ’72 paper. IEEE Transactions on Professional Communication, PC-15(2):40–43, June 1972. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Meng:2023:UET

[Men23] Michael Meng. Using eye tracking to study information selection and use in procedures. IEEE Transactions on Professional Communication, 66(1): 7–25, March 2023. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Mercuri:2005:TE

[Mer05] R. I. A. Mercuri. Technology as experience. IEEE Transactions on Professional Communication, 48(4):436–438, December 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Metzner:1973:IPS

[Met73] A. W. Kenneth Metzner. Integrating primary and secondary journals: a model for the immediate future. IEEE Transactions on Professional Communication, PC-16(3):84– 91, September 1973. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Metzner:1975:MUO

[Met75] A. W. Kenneth Metzner. Multiple use and other benefits of computerized publishing. IEEE Transactions on Professional Communication, PC-18 (3):274–278, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Meyer:1964:SWI

[Mey64] Charles A. Meyer. Seminar on writing-improvement programs for engineers — a report. IEEE Transactions on Engineering Writing and Speech, EWS-7(3): 18–22, December 1964. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Meyer:1969:KR

[Mey69] C. A. Meyer. Keynote remarks. IEEE Transactions on Engineering Writing and Speech, EWS-12(2):27, May 1969. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Meyer:1987:SIT

[Mey87] Benjamin D. Meyer. Systems integration: The technical writer as data manager. IEEE Transactions on Professional Communication, PC-30 (1):47–49, March 1987. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

McDonald:2012:ETA

[MEZ12] Sharon McDonald, Helen M.

Edwards, and Tingting Zhao. Exploring think-alouds in usability testing: an international survey. IEEE Transactions on Professional Communication, 55(1):2–19, March 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Michaelson:1983:HWP

[MF83] Herbert B. Michaelson and Charles R. Fenno. How to write and publish engineering papers and reports. IEEE Transactions on Professional Communication, PC-26(2):86–87, June 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

McDaniel:2017:CCM

[MFL17] Rudy McDaniel, Joseph R.

Fanfarelli, and Robb Lindgren. Creative content management: Importance, novelty, and affect as design heuristics for learning management systems. IEEE Transactions on Professional Communication, 60(2): 183–200, June 2017. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

McDaniel:1982:LRU

[MG82] Ellen McDaniel and Gwendolyn Gong. The language of robotics: Use and abuse of personification. IEEE Transactions on Professional Communication, PC-25(4):178–181, December 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Maitra:1995:RAR

[MG95] K. Maitra and D. Goswami. Responses of American readers to visual aspects of a midsized Japanese company’s annual report: a case study. IEEE Transactions on Professional Communication, 38(4): 197–203, December 1995. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Manley:1989:SGP

[MGB89] R. Manley, J. H. Graham, and R. C. Baxter. Some guidance on preparing technical articles for publication. IEEE Transactions on Professional Communication, 32(1):5–11, March 1989. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Michaelson:1959:IGT

[Mic59] Herbert B. Michaelson. Information gaps and traps in engineering papers. IRE Transactions on Engineering Writing and Speech, EWS-2(3):89– 92, December 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Michaelson:1960:C

[Mic60] Herbert B. Michaelson. The cover. IRE Transactions on Engineering Writing and Speech, EWS-3(1):2, April 1960. ISSN 0893-7869 (print), 2331-3692 (electronic).

Michaelson:1961:CVH

[Mic61a] Herbert B. Michaelson. Charles E. Van Hagan, Report Writers’ Handbook. IRE Transactions on Engineering Writing and Speech, EWS-4(3):102, December 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Michaelson:1961:CSD

[Mic61b] Herbert B. Michaelson. Charles Susskind, Dictionary of Style for Typewritten Technical Reports and Manuscripts for Publication. IRE Transactions on Engineering Writing and Speech, EWS-4(3):102, December 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Michaelson:1961:Ca

[Mic61c] Herbert B. Michaelson. The cover. IRE Transactions on Engineering Writing and Speech, EWS-4(1):2, January 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Michaelson:1961:Cb

[Mic61d] Herbert B. Michaelson. The cover. IRE Transactions on Engineering Writing and Speech, EWS-4(2):34, May 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Michaelson:1961:CAE

[Mic61e] Herbert B. Michaelson. Creative aspects of engineering writing. IRE Transactions on Engineering Writing and Speech, EWS-4(3):77–79, December 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Michaelson:1962:ARA

[Mic62] H. B. Michaelson. Abstracts of recent articles on engineering writing and speech. IRE Transactions on Engineering Writing and Speech, EWS-5(1):30, August 1962. ISSN 0893-7869 (print), 2331-3692 (electronic).

Michaelson:1964:TCS

[Mic64] Herbert B. Michaelson. Technical Communication, by Sidney Wilcox. IEEE Transactions on Engineering Writing and Speech, EWS-7(3):45, December 1964. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Michaelson:1965:PIT

[Mic65] Herbert B. Michaelson. Proceedings of the 1964 Institute in Technical and Industrial Communications, Ed. H. M. Weisman. IEEE Transactions on Engineering Writing and Speech, EWS-8(1):37–38, June 1965. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Michaelson:1968:AMD

[Mic68] Herbert B. Michaelson. Achieving a more disciplined R and D literature. IEEE Transactions on Engineering Writing and Speech, EWS-11(3): 156–159, December 1968. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Michaelson:1974:IMW

[Mic74] Herbert B. Michaelson. The incremental method of writing engineering papers. IEEE Transactions on Professional Communication, PC-17(1):21– 22, March 1974. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Michaelson:1980:HWP

[Mic80] Herbert B. Michaelson. How to write and publish a scientific paper, Robert A. Day. IEEE Transactions on Professional Communication, PC-23 (2):109, June 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Michaelson:1981:HOT

[Mic81] Herbert B. Michaelson. Hyman Olken, The technical communicator’s handbook of technology transfer. IEEE Transactions on Professional Communication, PC-24(2):106, June 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Michaelson:1984:TES

[Mic84] Herbert B. Michaelson. Teaching engineering students to communicate. IEEE Transactions on Professional Communication, PC-27(3):152–155, September 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Michaelson:1985:BRD

[Mic85a] Herbert B. Michaelson. Book reviews: David Sternberg, How to complete and survive a doctoral dissertation. IEEE Transactions on Professional Communication, PC-28(4):47, December 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Michaelson:1985:BRH

[Mic85b] Herbert B. Michaelson. Book reviews: Herman M. Weisman, Basic technical writing, fifth edition. IEEE Transactions on Professional Communication, PC-28(2):46, June 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Michaelson:1987:HWH

[Mic87] Herbert B. Michaelson. How writing helps R&D work. IEEE Transactions on Professional Communication, PC-30(2):85– 86, June 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Michaelson:1990:HAC

[Mic90] H. Michaelson. How an author can avoid the pitfalls of practical ethics. IEEE Transactions on Professional Communication, 33(2):58–61, June 1990. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Mignard:1983:MMH

[Mig83] James E. Mignard. Memo to Moses from Hew. IEEE Transactions on Professional Communication, PC-26(2):55– 56, June 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Miller:1961:WCT

[Mil61] Walter James Miller. What can the technical writer of the past teach the technical writer of today? IRE Transactions on Engineering Writing and Speech, EWS-4(3):69–76, December 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Miller:1962:PCI

[Mil62] Ephraim M. Miller. Preparation of commercial instruction manuals: a survey and some comments. IRE Transactions on Engineering Writing and Speech, EWS-5(1):26– 29, August 1962. ISSN 0893-7869 (print), 2331-3692 (electronic).

Milek:1964:NTV

[Mil64] John T. Milek. A novel technical vocabulary-improvement technique. IEEE Transactions on Engineering Writing and Speech, EWS-7(3):35–36, December 1964. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Miller:1965:EWC

[Mil65] Nathaniel Miller. The engineering writer Circa ’70: an extrapolation. IEEE Transactions on Engineering Writing and Speech, EWS-8(1):30–33, June 1965. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Miller:1981:TYT

[Mil81] Ryle L. Miller. Translate your thoughts into good technical writing. IEEE Transactions on Professional Communication, PC-24(2):73–74, June 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Milley:1984:Pa

[Mil84a] David Milley. Preface. IEEE Transactions on Professional Communication, PC-27(1): 6, March 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Milley:1984:Pb

[Mil84b] David Milley. Preface. IEEE Transactions on Professional Communication, PC-27(2):58, June 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Mintz:1969:NFT

[Min69] Harold K. Mintz. A neglected form of technical communication — the new product release. IEEE Transactions on Engineering Writing and Speech, EWS-12(3):69–70, October 1969. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Miranda:1975:PJ

[Mir75] Robert N. Miranda. The primary journal. IEEE Transactions on Professional Communication, PC-18(3):239–243, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Mirel:1991:CRE

[Mir91] B. Mirel. Critical review of experimental research on the usability of hard copy documentation. IEEE Transactions on Professional Communication, 34(2):109–122, June 1991. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Mireil:1999:GCB

[Mir99] Barbara Mireil. G. C. Bowker, S. L. Star, W. Turner, and L. Gasser, Social science, technical systems, and cooperative work: beyond the great divide. IEEE Transactions on Professional Communication, 42(2): 132–134, June 1999. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Mishra:2005:SDL

[Mis05] R. K. Mishra. Specialized discourse: Linguistic features and changing conventions. IEEE Transactions on Professional Communication, 48(3):337– 338, September 2005. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Mann:1984:MUR

[MK84] James A. Mann and John B. Ketchum. Multi-use: Reshaping technical material. IEEE Transactions on Professional Communication, PC-27(2):85– 88, June 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Mohan:2009:ECM

[MKBF09] Kannan Mohan, Nanda Kumar, and Raquel Benbunan-Fich. Examining communication media selection and information processing in software development traceability: an empirical investigation. IEEE Transactions on Professional Communication, 52(1): 17–39, March 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

McCarthy:1981:G

[MKM81] E. J. McCarthy, J. J. Kilpatrick, and J. MacNelly. The gobbledegook. IEEE Transactions on Professional Communication, PC-24(3): 116, September 1981. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Maddox:1978:ILP

[ML78] H. Maddox and R. J. Loughran.

Illustrating the lecture: Prepared diagrams versus built-up diagrams. IEEE Transactions on Professional Communication, PC-21(1):12–13, March 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Maes:1999:HPI

[ML99] A. A. Maes and H. Lenting. How to put the instructive space into words. IEEE Transactions on Professional Communication, 42(2): 100–113, June 1999. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Madnick:1968:SLM

[MM68] Stuart E. Madnick and Allen Moulton. SCRIPT, an on-line manuscript processing system. IEEE Transactions on Engineering Writing and Speech, EWS-11(2):92–100, August 1968. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Marks:1973:TCC

[MM73] Robert H. Marks and A. W. Kenneth Metzner. Typewriter composition cuts journal costs, speeds publication. IEEE Transactions on Professional Communication, PC-16(3):73– 79, September 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Marchwinski:2000:TCR

[MM00] T. Marchwinski and K. Mandziuk. The technical communicator’s role in initiating cross-functional teams. IEEE Transactions on Professional Communication, 43 (1):67–76, March 2000. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Mehlenbacher:2000:AIL

[MMCL00] B. Mehlenbacher, C. R. Miller, D. Covington, and J. S. Larsen. Active and interactive learning online: a comparison of Webbased and conventional writing classes. IEEE Transactions on Professional Communication, 43(2):166–184, June 2000. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Marett:2015:RRF

[MML15] Kent Marett, Emily Garrigues Marett, and Shanan R. Litchfield. The role of relational familiarity when interpreting family business communication. IEEE Transactions on Professional Communication, 58(1):69–85, March 2015. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Mousten:2010:LLT

[MMVH10] Birthe Mousten, Bruce Maylath, Sonia Vandepitte, and John Humbley. Learning localization through trans-Atlantic collaboration: Bridging the gap. IEEE Transactions on Professional Communication, 53(4):401–411, December 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Matveeva:2017:PLT

[MMW17] Natalya Matveeva, Michelle Moosally, and Russell Willerton. Plain language in the Twenty-First Century: Introduction to the special issue on plain language. IEEE Transactions on Professional Communication, 60(4):336– 342, December 2017. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Meister:2022:DTD

[MMW^{+}22] Philippe Meister, Jack Miller, Kexin Wang, Michael C. Dorneich, Eliot Winer, Lori J. Brown, and Geoffrey Whitehurst. Designing threedimensional augmented reality weather visualizations to enhance general aviation weather education. IEEE Transactions on Professional Communication, 65(2):321–336, June 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Miller:1980:CDT

[MO80] Joseph B. Miller and William F. O’Hearn. Crucial decisions for technical speakers. IEEE Transactions on Professional Communication, PC-23(1):14– 17, March 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Mohan:2002:MMS

[MOA02] A. V. Mohan, A. A. Omar, and K. A. Aziz. Malaysia’s Multimedia Super Corridor Cluster: communication linkages among stakeholders in a national system of innovation. IEEE Transactions on Professional Communication, 45(4): 265–275, December 2002. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Moeller:1998:ERT

[Moe98]  E. W. Moeller.  Edward

R. Tufte, Visual Explanations: Images and Quantities, Evidence and Narrative [Book Review]. IEEE Transactions on Professional Communication, 41(2):156–157, June 1998. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Mogull:2009:IIC

[Mog09]  Scott A. Mogull.  Inter-

net Inquiry: Conversations about Method (Markham, A. N. and Baym, N. K., Eds.; 2009) [Book Review]. IEEE Transactions on Professional Communication, 52(3):318– 320, September 2009. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Mogull:2014:IOI

[Mog14] Scott A. Mogull. Integrating online informative videos into technical communication service courses. IEEE Transactions on Professional Communication, 57(4):340– 363, December 2014. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Montalbo:1978:WCS

[Mon78] Thomas Montalbo. Winston Churchill: a study in oratory. IEEE Transactions on Professional Communication, PC-21 (1):5–8, March 1978. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Montalbo:1979:CEA

[Mon79] Thomas Montalbo. Communicate effectively! Arnold Nick Carter. IEEE Transactions on Professional Communication, PC-22(3):174, September 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Montalbo:1980:SSL

[Mon80a] Thomas Montalbo. Six steps to the lectern. IEEE Transactions on Professional Communication, PC-23(1):5–6, March 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Montoya:1980:WW

[Mon80b] Sarah Montoya. Word watching. IEEE Transactions on Professional Communication, PC-23(3):133–135, September 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Montalbo:1981:SIT

[Mon81a] Thomas Montalbo. Say it in threes. IEEE Transactions on Professional Communication, PC-24(3):128–129, September 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Montoya:1981:WWS

[Mon81b] Sarah Montoya. Word watching: Single or double consonant? IEEE Transactions on Professional Communication, PC-24(4):199–201, December 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Montoya:1981:WWW

[Mon81c] Sarah Montoya. Word watching: Who or whom? IEEE Transactions on Professional Communication, PC-24(2):84– 85, June 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Moorhead:1985:CAE

[Moo85]  Alice E. Moorhead.  The

conference approach to engineers’ report writing. IEEE Transactions on Professional Communication, PC-28(3):13– 16, September 1985. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Moore:1986:GKY

[Moo86a] Lois K. Moore. Getting to know you. IEEE Transactions on Professional Communication, PC-29(4):1, December 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Moore:1986:HEC

[Moo86b] Lois K. Moore. Helping engineers communicate. IEEE Transactions on Professional Communication, PC-29(3):1, September 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Moore:1986:PP

[Moo86c] Lois K. Moore. The President’s pledge. IEEE Transactions on Professional Communication, PC-29(2):1, June 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Moore:1987:ARS

[Moo87a]  Lois K. Moore.  Annual

report on society activities. IEEE Transactions on Professional Communication, PC-30 (3):117–118, September 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Moore:1987:EPR

[Moo87b] Lois K. Moore. Engineering phantoms respond. IEEE Transactions on Professional Communication, PC-30(2):65– 66, June 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Moore:1987:HES

[Moo87c] Lois K. Moore. Helping engineers share their knowledge. IEEE Transactions on Professional Communication, PC-30 (1):1, March 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Moore:1987:KPV

[Moo87d] Lois K. Moore. Kudos to PCS volunteers. IEEE Transactions on Professional Communication, PC-30(4):213, December 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Moore:2002:CGT

[Moo02] P. Moore. A concise guide to technical communication (with CD-ROM) — Book Review. IEEE Transactions on Professional Communication, 45(1): 61–62, March 2002. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Moore:2005:MDA

[Moo05] M. R. Moore. Multiliteracies for a digital age. IEEE Transactions on Professional Communication, 48(4):427– 429, December 2005. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Moore:2016:PEE

[Moo16]  Kristen R. Moore.  Pub-

lic engagement in environmental impact studies: a case study of professional communication in transportation planning. IEEE Transactions on Professional Communication, 59(3):245–260, September 2016. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Morris:1966:LTE

[Mor66] Jackson E. Morris. Linearization techniques for engineering writing. IEEE Transactions on Engineering Writing and Speech, EWS-9(2):33–35, December 1966. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Morris:1967:CAE

[Mor67a] Jackson E. Morris. Computeraided engineering writing: Test procedures for Apollo spacecraft. IEEE Transactions on Engineering Writing and Speech, EWS-10(1):10–13, July 1967. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Morris:1967:CMD

[Mor67b] Jackson E. Morris. Construction of a mnemonic dictionary for computer-aided engineering writing. IEEE Transactions on Engineering Writing and Speech, EWS-10(2):41–45, December 1967. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Mount:1975:SCN

[Mou75]  Ellis Mount.  A standard

for citations: Necessary cooperation. IEEE Transactions on Professional Communication, PC-18(3):130–132, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Moura:2005:DTM

[Mou05] H. Moura. Discourse & technology: Multimodal discourse analysis. IEEE Transactions on Professional Communication, 48(3):329–331, September 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Milroy:1979:LGI

[MP79] R. Milroy and E. C. Poulton. Labeling graphs for improved reading speed. IEEE Transactions on Professional Communication, PC-22(1):30– 33, March 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Matheson:2020:EUS

[MP20] Breeanne Matheson and Emily January Petersen. Engaging US students in culturally aware content creation and interactive technology design through service learning. IEEE Transactions on Professional Communication, 63(2):188–200, June 2020. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

McNair:2010:TPR

[MPD10] Lisa D. McNair, Marie C.

Paretti, and Marcia Davitt. Towards a pedagogy of relational space and trust: Analyzing distributed collaboration using discourse and speech act analysis. IEEE Transactions on Professional Communication, 53(3):233–248, September 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

McCormick:1976:TTP

[MQ76] Robert H. McCormick and Donna S. Queeney. Transferring technology PENNTAP style. IEEE Transactions on Professional Communication, PC-19(1):1–3, March 1976. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Meloncon:2021:ISI

[MR21] Lisa Melonçon and Nupoor Ranade. Introduction to special issue on accessibility. IEEE Transactions on Professional Communication, 64(3):215– 220, September 2021. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

McLellan:1996:EAM

[MRE96] S. G. McLellan, A. W. Roesler, and A. L. Elliot. The effect of advice message location on user performance. IEEE Transactions on Professional Communication, 39(1):43–48, March 1996. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Manning:2002:WT

[MS02] A. Manning and P. Skagestad. What is thought? IEEE Transactions on Professional Communication, 45 (1):50–53, March 2002. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

McGowan:1996:CCM

[MSC96] U. McGowan, J. Seton, and M. Cargill. A collaborating colleague model for inducting international engineering students into the language and culture of a foreign research environment. IEEE Transactions on Professional Communication, 39(3):117–121, September 1996. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Moffett:1985:TBP

[MSD85] John B. Moffett, Stephen G.

Smith, and S. J. De Amicis. Technical brochure preparation. IEEE Transactions on Professional Communication, PC-28(2):17–20, June 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Mckoy:2022:EPF

[MSDF22] Temptaous Mckoy, Cecilia D.

Shelton, Carleigh Davis, and Erin A. Frost. Embodying public feminisms: Collaborative intersectional models for engagement. IEEE Transactions on Professional Communication, 65(1):70–86, March 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Madigan:1997:LLP

[MSW97] C. Madigan, M. K. Silber, and S. Wilson. Lessons learned prototyping an SGML-based computerized document management system. IEEE Transactions on Professional Communication, 40(2):139–143, June 1997. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

MUDRAYA:2004:NDD

[Mud04] O. V. Mudraya. Need for datadriven instruction of engineering English. IEEE Transactions on Professional Communication, 47(1):65–70, March 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Mueller:1980:PCY

[Mue80] Douglas Mueller. Put clarity in your writing. IEEE Transactions on Professional Communication, PC-23(4):173–178, December 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Munger:2002:TCB

[Mun02] R. Munger. Technical communicators beware: the next generation of high-tech recruiting methods. IEEE Transactions on Professional Communication, 45(4):276–290, December 2002. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Munkvold:2005:EGC

[Mun05] B. E. Munkvold. Experiences from global e-collaboration: Contextual influences on technology adoption and use. IEEE Transactions on Professional Communication, 48 (1):78–86, March 2005. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Murray:1958:TFL

[Mur58] Ronald M. Murray. Technical films — a luxury or a necessity. IRE Transactions on Engineering Writing and Speech, EWS-1(1):31–36, March 1958. ISSN 0893-7869 (print), 2331-3692 (electronic).

Murphy:1959:TEA

[Mur59] Walter A. Murphy. Turning engineers into authors. IRE Transactions on Engineering Writing and Speech, EWS-2(3): 75–77, December 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Murphy:2000:EWB

[Mur00a] A. J. Murphy. Editing for writers [Book Reviews]. IEEE Transactions on Professional Communication, 43(1):102– 103, March 2000. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Murphy:2000:EWD

[Mur00b] A. J. Murphy. Effective Web design: master the essential [Book Reviews]. IEEE Transactions on Professional Communication, 43(1):110, March 2000. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Murphy:2000:IAW

[Mur00c]  A. J. Murphy.  Informa-

tion Architecture for the World Wide Web [Book Reviews]. IEEE Transactions on Professional Communication, 43(1): 104–106, March 2000. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Murphy:2000:LGP

[Mur00d] A. J. Murphy. Looking good in print [Book Reviews]. IEEE Transactions on Professional Communication, 43(1):100– 102, March 2000. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Murphy:2000:SIB

[Mur00e] A. J. Murphy. Starting an indexing business [Book Reviews]. IEEE Transactions on Professional Communication, 43(1):102, March 2000. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Murphy:2000:IRA

[Mur00f] A. J. Murphy. The inmates are running the asylum [Book Reviews]. IEEE Transactions on Professional Communication, 43(1):106–107, March 2000. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Murphy:2000:WPS

[Mur00g] A. J. Murphy. Web pages that suck: learn good design by looking at bad design [Book Reviews]. IEEE Transactions on Professional Communication, 43(1):107–108, March 2000. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Murphy:2000:PST

[Mur00h]  A. J. Murphy.  Page to

screen: taking literacy into the electronic [book reviews]. IEEE Transactions on Professional Communication, 43(1): 103–104, March 2000. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Murphy:2000:PRA

[Mur00i] A. J. Murphy. Paul Renner.

The art of typography [Book Reviews]. IEEE Transactions on Professional Communication, 43(1):99–100, March 2000. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Murphy:2000:TBX

[Mur00j] A. J. Murphy. Three books on XML [book reviews]. IEEE Transactions on Professional Communication, 43(1):108– 109, March 2000. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Murphy:2000:TBS

[Mur00k] A. J. Murphy. Two books on Silicon Valley [book reviews]. IEEE Transactions on Professional Communication, 43(1): 110–115, March 2000. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Musgrave:1978:DVA

[Mus78a] Joan F. Musgrave. Draw! A visual approach to thinking, learning, and communicating, Kurt Hanks and Larry Belliston. IEEE Transactions on Professional Communication, PC-21(3):129, September 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Musgrave:1978:ECA

[Mus78b]  Joan F. Musgrave.  Ex-

periments in computer-aided graphic expression. IEEE Transactions on Professional Communication, PC-21(3): 110–117, September 1978. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Mashburn:1983:PCP

[MV83] John Mashburn and Janice Vantrease. Planning a company program for technical documentation. IEEE Transactions on Professional Communication, PC-26(4):166–169, December 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Margherita:2021:EPA

[MV21] Alessandro Margherita and David Verrill. Elevator pitch assessment model: a systematization of dimensions in technology entrepreneurship presentations. IEEE Transactions on Professional Communication, 64(4):304–321, December 2021. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Muhren:2009:MSM

[MVV09] Willem J. Muhren, Gerd Van Den Eede, and Bartel Van de Walle. Making sense of media synchronicity in humanitarian crises. IEEE Transactions on Professional Communication, 52(4):377–397, December 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Mackiewicz:2014:PRU

[MY14] Jo Mackiewicz and Dave Yeats.

Product review users’ perceptions of review quality: The role of credibility, informativeness, and readability. IEEE Transactions on Professional Communication, 57(4): 309–324, December 2014. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Mackiewicz:2016:IRE

[MYT16] Jo Mackiewicz, Dave Yeats, and Thomas Thornton. The impact of review environment on review credibility. IEEE Transactions on Professional Communication, 59(2):71–88, June 2016. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

McDaniel:1991:DDM

[MYV^{+}91] E. McDaniel, R. E. Young, J. Vesterager, K. Bergsson, S. Jensen, and E. Tvedt. Document-driven management of knowledge and technology transfer: Denmark’s CIM/ GEMS project in computerintegrated manufacturing. IEEE Transactions on Professional Communication, 34(2):83–93, June 1991. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

McDaniel:1992:DDM

[MYV92] E. McDaniel, R. E. Young, and J. Vesterager. Documentdriven management of knowledge and technology transfer: Denmark’s CIM/GEMS project in computer-integrated manufacturing. II. IEEE Transactions on Professional Communication, 35(2):71–83, June 1992. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Nadziejka:1996:HET

[Nad96] D. E. Nadziejka. How to Edit Technical Documents [Book Reviews]. IEEE Transactions on Professional Communication, 39(1):59, March 1996. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Nagle:1986:EWI

[Nag86] Joan G. Nagle. From the Editor: what I’m taking to Aunt Lizzie’s. IEEE Transactions on Professional Communication, PC-29(2):2, June 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Nagle:1987:CHW

[Nag87a] Joan G. Nagle. Communication as hard work. IEEE Transactions on Professional Communication, PC-30(2):67, June 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Nagle:1987:FYS

[Nag87b] Joan G. Nagle. Funny you should ask! IEEE Transactions on Professional Communication, PC-30(4):214–215, December 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Nagle:1987:SCA

[Nag87c] Joan G. Nagle. The state of the communication arts. IEEE Transactions on Professional Communication, PC-30 (1):2–3, March 1987. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Nailen:1981:EGC

[Nai81] Richard L. Nailen. An engineer’s guide to clear language. IEEE Transactions on Professional Communication, PC-24 (3):117–119, September 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Nelson-Burns:2004:DBN

[NB04] C. Nelson-Burns. The Davis– Besse Nuclear Power Plant eroded reactor head: a case study. IEEE Transactions on Professional Communication, 47(4):268–280, December 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Nierenberg:1981:MTG

[NC81] Gerard I. Nierenberg and Henry H. Calero. Meta-talk, a guide to hidden meanings in conversation. IEEE Transactions on Professional Communication, PC-24(2):103–104, June 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Nelson:1983:PD

[ND83] Roy Paul Nelson and L. W. Denton. Publication design. IEEE Transactions on Professional Communication, PC-26 (4):214, December 1983. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Neeck:1980:CP

[Nee80]  Dale A. Neeck.  Off-the-

cuff presentations. IEEE Transactions on Professional Communication, PC-23(1):17– 18, March 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Neeley:1992:WMW

[Nee92] K. A. Neeley. Woman as mediatrix: women as writers on science and technology in the eighteenth and nineteenth centuries. IEEE Transactions on Professional Communication, 35(4):208–216, December 1992. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Nelson:1990:PWD

[Nel90] R. J. Nelson. Preparing to write the document: a worksheet for situational analysis in the workplace. IEEE Transactions on Professional Communication, 33(1):12–18, March 1990. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Nelson:2003:ETS

[Nel03] S. Nelson. Engineering and technology student perceptions of collaborative writing practices. IEEE Transactions on Professional Communication, 46(4):265–276, December 2003. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Norback:2005:IWC

[NH05] J. S. Norback and J. R. Hardin.

Integrating workforce communication into senior design. IEEE Transactions on Professional Communication, 48(4): 413–426, December 2005. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Nicholson:1959:TLB

[Nic59] Natalie N. Nicholson. Technical libraries in the Boston area. IRE Transactions on Engineering Writing and Speech, EWS-2(2):52–55, June 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Narita:2003:CSD

[NKU03] M. Narita, K. Kurokawa, and T. Utsuro. Case study on the development of a computerbased support tool for assisting Japanese software engineers with their English writing needs. IEEE Transactions on Professional Communication, 46(3):194–209, September 2003. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Norback:2010:IEP

[NLK10] J. Norback, E. Leeds, and K. Kulkarni. Integrating an executive panel on communication info an engineering curriculum. IEEE Transactions on Professional Communication, 53(4):412–422, December 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Nizamani:2022:HEV

[NNB^{+}22] Sehrish Nizamani, Saad Nizamani, Nazish Basir, Gulsher Laghari, Khalil Khoumbati, and Sarwat Nizamani. Heuristic evaluation versus guideline reviews: a tale of comparing two domain usability expert’s evaluation methods. IEEE Transactions on Professional Communication, 65(4): 516–529, December 2022. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Noam:2003:INN

[Noa03] E. M. Noam. Interconnecting the Network of Networks [Book Review]. IEEE Transactions on Professional Communication, 46(1):66–67, March 2003. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Noble:2010:MSL

[Nob10] Allan Noble. Mortal Syntax:

101 Language Choices That Will Get You Clobbered by the Grammar Snobs Even If You’re Right (Casagrande, J.) [Book Review]. IEEE Transactions on Professional Communication, 53(1):87–88, March 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Noone:1979:TSV

[Noo79] Thomas M. Noone. Trade secret vs patent protection. IEEE Transactions on Professional Communication, PC-22 (2):123–127, June 1979. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Norko:1958:FPP

[Nor58] R. J. Norko. Formula for platform poise. IRE Transactions on Engineering Writing and Speech, EWS-1(1):18– 19, March 1958. ISSN 0893-7869 (print), 2331-3692 (electronic).

Northrop:1997:WPT

[Nor97] M. J. Northrop. William S. Pfeiffer, Technical Writing: a Practical Approach (3rd Ed.). Englewood Cliffs, NJ: Prentice–Hall, 1997 [Book Review]. IEEE Transactions on Professional Communication, 40(3):237–238, September 1997. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Norton:2000:TCB

[Nor00] D. W. Norton. Technical communication as business strategy: how changes in discursive patterns affect the value of technical communication in cross-functional team settings. IEEE Transactions on Professional Communication, 43(1): 77–89, March 2000. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Nandagopal:2023:MSS

[NP23] Shoba Kanare Nandagopal and Ruth Sara Philip. A metainvestigation of speaking skills:

Practice, feedback, and selfdirected efforts. IEEE Transactions on Professional Communication, 66(2):170–185, June 2023. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Nussbaumer:1979:PE

[Nus79] Henri J. Nussbaumer. Patents and the engineer. IEEE Transactions on Professional Communication, PC-22(2):68– 72, June 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Nyce:1997:BNE

[Nyc97] J. M. Nyce. Bonnie A. Nardi, Ed., Context and Consciousness: Activity Theory and Human-computer Interaction. Cambridge, Ma: Mit Press, 1996, 400 pp. [Book Review]. IEEE Transactions on Professional Communication, 40(3): 242–243, September 1997. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Nye:1999:CPS

[Nye99] D. E. Nye. Consuming power:

a social history of American energies. IEEE Transactions on Professional Communication, 42(2):137–138, June 1999. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

OConnor:1962:WSR

[O’C62] John O’Connor. What should a retrieval system for scientific information do? IRE Transactions on Engineering Writing and Speech, EWS-5(2):75– 77, December 1962. ISSN 0893-7869 (print), 2331-3692 (electronic).

Oatey:1997:MF

[OC97] M. Oatey and M. Cawood. Minimalism as a framework. IEEE Transactions on Professional Communication, 40(4): 265–274, December 1997. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ocker:1971:PHE

[Ock71] Wolfgang A. Ocker. A program to hyphenate English words. IEEE Transactions on Engineering Writing and Speech, EWS-14(2):53–59, June 1971. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Ocker:1975:PHE

[Ock75] Wolfgang A. Ocker. Program to hyphenate English words. IEEE Transactions on Professional Communication, PC-18 (2):78–84, June 1975. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ocker:2005:ICA

[Ock05] R. J. Ocker. Influences on creativity in asynchronous virtual teams: a qualitative analysis of experimental teams. IEEE Transactions on Professional Communication, 48(1): 22–39, March 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ocker:2007:BAI

[Ock07] Rosalie J. Ocker. A balancing act: The interplay of status effects on dominance in virtual teams. IEEE Transactions on Professional Communication, 50(3):204–218, September 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

ODonnell:2007:MOS

[O’D07] Casey O’Donnell. Making an open source case for offshoring commentary. IEEE Transactions on Professional Communication, 50(2):85–87, June 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Odescalchi:1986:DKU

[Ode86] Esther Kando Odescalchi. Documentation is the key to user success. IEEE Transactions on Professional Communication, PC-29(4):16–18, December 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ostroff:1992:BED

[ODF92] D. H. Ostroff, D. Donnelly, and A. Fried. The business environment, demographics and technology: a case study of Florida Power and Light’s electronic employee communication services. IEEE Transactions on Professional Communication, 35(1):31–37, March 1992. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Okada:1964:MR

[Oka64] S. Okada. On the matter of reference. IEEE Transactions on Engineering Writing and Speech, EWS-7(3):15–18, December 1964. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Oaks:1998:WCH

[OL98] D. D. Oaks and J. Lewis. What can humor and advertising teach us about legal drafting? IEEE Transactions on Professional Communication, 41(4):277–279, December 1998. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Oliver:1958:IYW

[Oli58] Frank J. Oliver. If you write for “electrical manufacturing”. IRE Transactions on Engineering Writing and Speech, EWS-1(1):23–24, March 1958. ISSN 0893-7869 (print), 2331-3692 (electronic).

Olken:1987:RF

[Olk87] Hyman Olken. The readers’ forum. IEEE Transactions on Professional Communication, PC-30(4):233–234, December 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Olsen:1991:EPC

[Ols91]  G. R. Olsen.  Eidetecker:

the professional communicator in the new visual culture. IEEE Transactions on Professional Communication, 34(1): 13–19, March 1991. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Olsen:1996:EIC

[Ols96] L. A. Olsen. Explanation and interaction: The computer generation of explanatory discourse. IEEE Transactions on Professional Communication, 39(2):106, June 1996. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Osaka:2012:PGI

[OM12] Noriko Osaka and Arun J.

Martin. A Practical Guide to Information Architecture (Spencer. Donna; 2010) [Book Review]. IEEE Transactions on Professional Communication, 55(3):287–288, September 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Oram:1986:OPN

[Ora86] Andrew Oram. Opening passage: a new look at the system documentation problem.

IEEE Transactions on Professional Communication, PC-29 (4):3–10, December 1986. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Orr:1999:GFC

[Orr99] T. Orr. Genre in the field of computer science and computer engineering. IEEE Transactions on Professional Communication, 42(1):32– 37, March 1999. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Orr:2001:ELE

[Orr01] T. Orr. English language education for specific professional needs. IEEE Transactions on Professional Communication, 44(3):207–211, September 2001. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Orr:2002:APE

[Orr02]  T. Orr.  Assessing profi-

ciency in engineering English. IEEE Transactions on Professional Communication, 45(1): 40–44, March 2002. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Orr:2003:ISI

[Orr03] T. Orr. Introduction to the special issue [English language training for nonnative speakers of English in science or engineering]. IEEE Transactions on Professional Communication, 46(3):153–156, September 2003. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Orr:2006:ISI

[Orr06] T. Orr. Introduction to the special issue: Insights from corpus linguistics for professional communication. IEEE Transactions on Professional Communication, 49(3):213– 216, September 2006. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Orr:2010:APC

[Orr10]  Thomas Orr.  Assessment

in professional communication. IEEE Transactions on Professional Communication, 53(1): 1–3, March 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Orth:1972:AW

[Ort72] Melvin F. Orth. Abstracting for the writer. IEEE Transactions on Professional Communication, PC-15(2):43–44, June 1972. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Orth:1974:LBL

[Ort74] Melvin F. Orth. Lower the boom on the lowly pronoun. IEEE Transactions on Professional Communication, PC-17 (2):39–41, June 1974. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Orth:1975:CTP

[Ort75] Melvin F. Orth. Color their prose gray. IEEE Transactions on Professional Communication, PC-18(2):65–66, June 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Orth:1978:SPS

[Ort78] Mel F. Orth. Striking out: Poor style and grammar still abound in technical writing. IEEE Transactions on Professional Communication, PC-21 (2):44–47, June 1978. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Olds:1991:PPT

[OW91] B. M. Olds and K. B. Wiley. Public policy and technical communication across the curriculum at the Colorado School of Mines. IEEE Transactions on Professional Communication, 34(4):240–246, December 1991. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Padin:1978:ANI

[Pad78]  Mary Ellen Padin.  Au-

thor names and inadvertent pseudonyms in bibliographic data bases. IEEE Transactions on Professional Communication, PC-21(2):77–78, June 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Padmanabhan:2007:TCO

[Pad07] Poornima Padmanabhan. Technical communication outsourcing: The twelve driver framework tutorial. IEEE Transactions on Professional Communication, 50(2):109–120, June 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Pakin:1981:EUD

[Pak81] Sandra Pakin. Evaluate user documentation before you buy the software. IEEE Transactions on Professional Communication, PC-24(2):75–78, June 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Pakin:1983:DDM

[Pak83] Sandra Pakin. Documentation development methodology. IEEE Transactions on Professional Communication, PC-26 (1):45, March 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Palmour:1975:AIA

[Pal75]  Vernon E. Palmour.  Al-

ternatives for increasing access to scientific journals. IEEE Transactions on Professional Communication, PC-18 (3):210–215, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Palmour:1977:NPA

[Pal77] Vernon E. Palmour. National program for access to periodicals: Recent progress. IEEE Transactions on Professional Communication, PC-20(2):72– 74, September 1977. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Paltridge:2000:GKT

[Pal00] B. Paltridge. Genre knowledge and teaching professional communication. IEEE Transactions on Professional Communication, 43(4):397– 401, December 2000. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Parisi:1975:CIA

[Par75] Paul A. Parisi. Composition innovations at the American Society of Civil Engineers. IEEE Transactions on Professional Communication, PC-18 (3):244–273, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Partridge:1986:WTO

[Par86] Stephen Kurt Partridge. So what is task orientation, anyway? IEEE Transactions on Professional Communication, PC-29(4):26–32, December 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Parsons:1987:EFI

[Par87] Gerald M. Parsons. Ethical factors influencing curriculum design and instruction in technical communication. IEEE Transactions on Professional Communication, PC-30 (3):202–207, September 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Paretti:2006:AAL

[Par06] Marie C. Paretti. Audience awareness: leveraging problembased learning to teach workplace communication practices. IEEE Transactions on Professional Communication, 49(2): 189–198, June 2006. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Paskusz:1960:CAS

[Pas60] G. F. Paskusz. Communication: Art or science? IRE Transactions on Engineering Writing and Speech, EWS-3(2): 59–61, July 1960. ISSN 0893-7869 (print), 2331-3692 (electronic).

Pascal:2003:ECI

[Pas03] C. L. Pascal. Enabling chance interaction through instant messaging. IEEE Transactions on Professional Communication, 46(2):138–141, June 2003. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Patterson:1959:RIP

[Pat59] T. T. Patterson. A review of the IRE-PGEWS. IRE Transactions on Engineering Writing and Speech, EWS-2(3):83, December 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Patterson:1975:WYS

[Pat75] T. T. Patterson. Why you should publish. IEEE Transactions on Professional Communication, PC-18(1):2–4, March 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Patterson:1976:HST

[Pat76] T. T. Patterson. How to start a technical journal. IEEE Transactions on Professional Communication, PC-19(1):7– 14, March 1976. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Patton:2008:BWB

[Pat08] Martha D. Patton. Beyond WI: Building an integrated communication. IEEE Transactions on Professional Communication, 51(3):313–327, September 2008. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Pavlov:1994:STI

[Pav94] A. N. Pavlov. The Scientific and Technical Information Center of Russia. IEEE Transactions on Professional Communication, 37(2):62–65, June 1994. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Pinelli:1993:TCP

[PBK^{+}93] T. E. Pinelli, R. O. Barclay, M. L. Kenne, M. Flammia, and J. M. Kennedy. The technical communication practices of Russian and US aerospace engineers and scientists. IEEE Transactions on Professional Communication, 36(2):95–104, June 1993. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Panteli:2005:RSC

[PD05] N. Panteli and R. M. Davison. The role of subgroups in the communication patterns of global virtual teams. IEEE Transactions on Professional Communication, 48(2): 191–200, June 2005. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Phillips:2022:LTH

[PD22] Lisa L. Phillips and Raquel L.

Deleon. Living testimonios: How Latinx graduate students persist and enact social justice within higher education. IEEE Transactions on Professional Communication, 65(1): 197–212, March 2022. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Palumbo:2024:ECT

[PD24] Giuseppe Palumbo and Ann Hill Duin. The evolutionary convergence of technical communication and translation: an integrative literature review of scholarship from 2000 to 2022. IEEE Transactions on Professional Communication, 67(3): 266–284, September 2024. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Pihlaja:2024:TLG

[PDD^{+}24] Beau Pihlaja, Meghalee Das, Laurena Davis, Aliethia Dean, Danyela M. Fonseca, Elizabeth Hughes, Regan Joswiak, Laura Koleva, Min Yang, and Jiaxin Zhang. Translation and localization in global technical communication. IEEE Transactions on Professional Communication, 67(3):386– 402, September 2024. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Palmer:1988:DEO

[PDG^{+}88] J. Palmer, T. Duffy, K. Gomoll, T. Gomoll, J. Richards-Palmquist, and J. A. Trumble. The design and evaluation of online help for Unix EMACS: capturing the user in menu design. IEEE Transactions on Professional Communication, 31(1):44–51, March 1988. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Palmer:1988:CPG

[PDK88] J. Palmer, T. M. Duffy, and K. Knabe. Conference program generator software: a program that links desktop publishing and data base technology. IEEE Transactions on Professional Communication, 31(1):37–43, March 1988. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Paul:2016:GVT

[PDL16] Ravi Paul, John R. Drake, and Huigang Liang. Global virtual team performance: The effect of coordination effectiveness, trust, and team cohesion. IEEE Transactions on Professional Communication, 59(3): 186–202, September 2016. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Paretti:2019:FSP

[PEG19] Marie C. Paretti, Andreas Eriksson, and Magnus Gustafsson. Faculty and student perceptions of the impacts of communication in the disciplines (CID) on students development as engineers. IEEE Transactions on Professional Communication, 62(1):27–42, March 2019. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Pei:1981:DSA

[Pei81] Mario Pei. Double speak in America. IEEE Transactions on Professional Communication, PC-24(2):102–103, June 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Peneguy:1998:PET

[Pen98a] L. D. Peneguy. Profile of Eugen Tarnow. IEEE Transactions on Professional Communication, 41(3):216–217, September 1998. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Peneguy:1998:PHC

[Pen98b] L. D. Peneguy. Profile of Helen Constantinides. IEEE Transactions on Professional Communication, 41(1):74– 76, March 1998. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Peneguy:1998:PJG

[Pen98c] L. D. Peneguy. Profile of Johnny Grissom. IEEE Transactions on Professional Communication, 41(4):280–282, December 1998. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Peneguy:1998:PKD

[Pen98d] L. D. Peneguy. Profile of Ken Dawson. IEEE Transactions on Professional Communication, 41(3):213–215, September 1998. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Peneguy:1998:PLK

[Pen98e] L. D. Peneguy. Profile of Leann Kostek. IEEE Transactions on Professional Communication, 41(4):283–285, December 1998. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Peneguy:1998:PLW

[Pen98f] L. D. Peneguy. Profile of Lesley Warner. IEEE Transactions on Professional Communication, 41(2):146–148, June 1998. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Peneguy:1998:PMB

[Pen98g] L. D. Peneguy. Profile of Mark Barnett. IEEE Transactions on Professional Communication, 41(1):72–73, March 1998. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Peneguy:1998:PTK

[Pen98h] L. D. Peneguy. Profile of Tibor Koltay. IEEE Transactions on Professional Communication, 41(2):149–151, June 1998. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Peneguy:1999:CLI

[Pen99a] L. D. Peneguy. Curbing language intensity. IEEE Transactions on Professional Communication, 42(1):52–54, March 1999. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Peneguy:1999:PKP

[Pen99b] L. D. Peneguy. Profile of Kim Powell. IEEE Transactions on Professional Communication, 42(1):59–60, March 1999. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Perry:1963:DQS

[Per63] James W. Perry. Defining the query spectrum — the basis for developing and evaluating information-retrieval methods. IEEE Transactions on Engineering Writing and Speech, EWS-6(1):20–27, September 1963. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Perica:1970:WGH

[Per70] Louis Perica. Auf Wiedersehen, Gutenberg — Hello, McLulan! Improving technical communications using the multimedia approach. IEEE Transactions on Engineering Writing and Speech, EWS-13 (1):8–12, May 1970. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Perry:1978:ART

[Per78] Robert E. Perry. Audience requirements for technical speakers. IEEE Transactions on Professional Communication, PC-21(3):91–96, September 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Perkins:1993:SPT

[Per93] J. M. Perkins. Social perspectives on technology transfer. IEEE Transactions on Professional Communication, 36(4): 185–189, December 1993. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Perugini:1996:AAS

[Per96] V. Perugini. Anytime, anywhere: the social impact of emerging communication technology. IEEE Transactions on Professional Communication, 39(1):4–15, March 1996. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Peterson:1959:GWI

[Pet59] Arnold P. G. Peterson. The Greeks had a word for it. IRE Transactions on Engineering Writing and Speech, EWS-2(3): 73–74, December 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Peterson:1969:TPC

[Pet69] Norman D. Peterson. Textual processing, a computerized aid for the writer, editor, and reader. IEEE Transactions on Engineering Writing and Speech, EWS-12(1):9– 12, May 1969. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Petit:2013:EGM

[Pet13] Angela Petit. Elizabeth Goodman, Mike Kuniavsky, Andrea Moed, Observing the User Experience: a Practitioner’s Guide to User Research (Second Edition). IEEE Transactions on Professional Communication, 56(3):260– 261, September 2013. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Petit:2014:DWB

[Pet14] Angela Petit. Designing Webbased applications for 21st century writing classrooms. IEEE Transactions on Professional Communication, 57(3): 235–237, September 2014. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Pfaffenberger:1986:RNS

[Pfa86] Bryan Pfaffenberger. Research networks, scientific communication, and the personal computer. IEEE Transactions on Professional Communication, PC-29(1):30–33, March 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Paul:2021:MEV

[PFD^{+}21] Ravi C. Paul, Christopher P. Furner, John R. Drake, Richard D. Hauser, and Eric Kisling. The moderating effect of virtuality on team trust and effectiveness. IEEE Transactions on Professional Communication, 64(2):185–200, June 2021. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Pierson:1988:FAL

[PFM88] J. K. Pierson, K. A. Forcht, and W. Moates. Factors affecting level of documentation required for user-developed applications. IEEE Transactions on Professional Communication, 31(3):142–148, September 1988. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Pellegrini:2024:CEO

[PH24] Mason Pellegrini and Paul Thompson Hunter. Constructing entrepreneurial opportunities: The argumentative structure of early-stage business pitch problem statements at an international business accelerator. IEEE Transactions on Professional Communication, 67(2):156–172, June 2024. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Phelps:2020:MUC

[Phe20] Johanna Phelps. Modernization updates to the common rule: Recommendations for researchers working with human participants. IEEE Transactions on Professional Communication, 63(1):85–95, March 2020. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Phelps:2021:OCD

[Phe21] Christin Phelps. Opening the conversation: The development of a faculty-facing module on teaching students with dyslexia. IEEE Transactions on Professional Communication, 64(3):250–263, September 2021. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Phillips:1969:PTW

[Phi69a] John Phillips. Practical technical writing. IEEE Transactions on Engineering Writing and Speech, EWS-12(3):91, October 1969. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Phillips:1969:WTL

[Phi69b]  John Phillips.  Where is

technology leading communications? — a review of the February 1969 EWS Symposium. IEEE Transactions on Engineering Writing and Speech, EWS-12(2):23–26, May 1969. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Phillips:1972:ELAa

[Phi72a]  John C. Phillips.  Edito-

rial: Looking ahead to ’72. IEEE Transactions on Professional Communication, PC-15 (1):1, March 1972. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Phillips:1972:ELAb

[Phi72b]  John C. Phillips.  Edito-

rial: Looking ahead to ’73. IEEE Transactions on Professional Communication, PC-15 (3):71, September 1972. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Plotnick:2016:IDP

[PHP16] Linda Plotnick, Starr Roxanne Hiltz, and Robin Privman. Ingroup dynamics and perceived effectiveness of partially distributed teams. IEEE Transactions on Professional Communication, 59(3):203– 229, September 2016. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Privman:2013:GUV

[PHW13] Robin Privman, Starr Roxanne Hiltz, and Yiran Wang. Ingroup (us) versus out-group (them) dynamics and effectiveness in partially distributed teams. IEEE Transactions on Professional Communication, 56(1):33–49, March 2013. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Pickus:1980:TCS

[Pic80] Robert M. Pickus. Technology and copyright: Sources and material, George P. Bush and Robert H. Dreyfuss, Eds. IEEE Transactions on Professional Communication, PC-23 (3):153–154, September 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Pierce:1958:CFE

[Pie58] J. R. Pierce. The challenging field of engineering writing and speech. IRE Transactions on Engineering Writing and Speech, EWS-1(1):12– 13, March 1958. ISSN 0893-7869 (print), 2331-3692 (electronic).

Pierson:1971:CPE

[Pie71] John L. Pierson. A computer program for electronic typesetting. IEEE Transactions on Engineering Writing and Speech, EWS-14(2):46–52, June 1971. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Pilditch:1981:DYR

[Pil81]  James Pilditch.  Did you

read the instructions carefully? IEEE Transactions on Professional Communication, PC-24 (4):185–186, December 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Pilfold:1987:LSS

[Pil87] David Pilfold. Legal style in specifications. IEEE Transactions on Professional Communication, PC-30(3):168–172, September 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Pinkerton:2018:UUD

[Pin18] R. Jason Pinkerton. Universal UX Design: Building Multicultural User Experience (Ferreira, A.) [Book Review]. IEEE Transactions on Professional Communication, 61(3): 348–349, September 2018. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Pitner:1979:IYM

[Pit79]  Stephen L. Pitner.  Im-

prove your method of preparing a technical study. IEEE Transactions on Professional Communication, PC-22(1):22– 23, March 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Passera:2017:DCF

[PKLS17] Stefania Passera, Anne Kankaanranta, and Leena Louhiala-Salminen. Diagrams in contracts: Fostering understanding in global business communication. IEEE Transactions on Professional Communication, 60(2):118–146, June 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Plata:1971:FPS

[Pla71] Errol F. Plata. Feasibility of preparing a Symbolic Integrated Maintenance Manual (SIMM) for complex new electronic systems. IEEE Transactions on Engineering Writing and Speech, EWS-14(1):34–37, February 1971. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Plung:1981:RFT

[Plu81] Daniel L. Plung. Readability formulas and technical communication. IEEE Transactions on Professional Communication, PC-24(1):52–54, March 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Plung:1984:ASY

[Plu84] Daniel L. Plung. Add style to your technical writing. IEEE Transactions on Professional Communication, PC-27(1):20– 24, March 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Plung:1984:PAC

[PM84] Daniel L. Plung and Lois K. Moore. Publishing an anthology a case history with suggestions. IEEE Transactions on Professional Communication, PC-27(4):217–221, December 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Perkins:1998:TSG

[PM98] J. Perkins and C. Maloney. Today’s style guide: trusted tool with added potential.

IEEE Transactions on Professional Communication, 41 (1):24–32, March 1998. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Perkins:2004:MDP

[PM04] E. Perkins and M. Markel.

Multinational data-privacy laws: an introduction for IT managers. IEEE Transactions on Professional Communication, 47(2):85–94, June 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Paretti:2008:ISI

[PM08] Marie C. Paretti and Lisa D.

McNair. Introduction to the special issue on communication in engineering curricula: Mapping the landscape. IEEE Transactions on Professional Communication, 51(3): 238–241, September 2008. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Palma:2021:IMM

[PMGW21] Giulia Palma, Bryan Mesmer, Amy Guerin, and Kristin Weger. Identifying multidisciplinary metrics to analyze NASA case studies. IEEE Transactions on Professional Communication, 64(2):170– 184, June 2021. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Podell:1984:MME

[Pod84] Edwin J. Podell. Mathematics must be effective in technical communication. IEEE Transactions on Professional Communication, PC-27(2):97–100, June 1984. CODEN IEPCBU. ISSN 0099-9474.

Podolsky:2003:ESS

[Pod03] J. Podolsky. Exploding steamboats, Senate debates, and technical reports: The convergence of technology, politics, and rhetoric in the Steamboat Bill of 1838 [Book Review]. IEEE Transactions on Professional Communication, 46(3): 244–245, September 2003. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Porter:1987:TTA

[Por87] James E. Porter. Truth in technical advertising: a case study. IEEE Transactions on Professional Communication, PC-30 (3):182–189, September 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Potter:1977:MMM

[Pot77] J. Leith Potter. Mathematical model for market forecasting by a technical society. IEEE Transactions on Professional Communication, PC-20 (2):121–125, September 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Potvin:1982:OPE

[Pot82] Janet H. Potvin. Other papers: Eight steps to better newsletters. IEEE Transactions on Professional Communication, PC-25(4):204–210, December 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Potvin:1983:POA

[Pot83] Janet H. Potvin. Planning and organizing an annual conference. IEEE Transactions on Professional Communication, PC-26(3):123–152, September 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Potvin:1984:UTR

[Pot84] Janet H. Potvin. Using team reporting projects to teach concepts of audience and written, oral, and interpersonal communication skills. IEEE Transactions on Professional Communication, PC-27(3):130–137, September 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Potsus:2000:CMS

[Pot00a] W. B. Potsus. Corporate memory: strategies for knowledge management [Book Review]. IEEE Transactions on Professional Communication, 43(4): 412–414, December 2000. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Potsus:2000:OGE

[Pot00b]  W. B. Potsus.  Outlining

goes electronic [Book Review]. IEEE Transactions on Professional Communication, 43(3): 336–338, September 2000. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Potsus:2001:RPE

[Pot01a] W. B. Potsus. Revising prose, 4th ed. [Book Review]. IEEE Transactions on Professional Communication, 44(3):216– 218, September 2001. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Potsus:2001:PHU

[Pot01b] W. B. Potsus. Pradeep Henry, User-centered information design for improved software usability [Book Review]. IEEE Transactions on Professional Communication, 44(2):155– 158, June 2001. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Potsus:2002:UWB

[Pot02] W. B. Potsus. Understanding writing blocks. IEEE Transactions on Professional Communication, 45(2):147–150, June 2002. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Power:1980:RYW

[Pow80] Ruth M. Power. Rate your writing skills. IEEE Transactions on Professional Communication, PC-23(2):84–86, June 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Power:1981:WNT

[Pow81] Ruth M. Power. Who needs a technical editor? IEEE Transactions on Professional Communication, PC-24(3):139–140, September 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Powell:1982:ETC

[Pow82a] Kenneth B. Powell. Effective technical communication. IEEE Transactions on Professional Communication, PC-25 (3):163, September 1982. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Powell:1982:WV

[Pow82b] Kenneth B. Powell. Writing with verve. IEEE Transactions on Professional Communication, PC-25(4):189–191, December 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Pierson:1997:BEK

[PP97] M. M. Pierson and B. L. Pierson. Beginnings and endings: keys to better engineering technical writing. IEEE Transactions on Professional Communication, 40(4):299– 304, December 1997. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Prasad:1987:EPB

[Pra87] Sudha Prasad. Effective presentations: a bibliography. IEEE Transactions on Professional Communication, PC-30 (1):39–40, March 1987. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Priest:1972:RCS

[Pri72] W. Curtiss Priest. Restructuring communication in science and technology: a PSYCHOM ’72 paper. IEEE Transactions on Professional Communication, PC-15(2):20–22, June 1972. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Price:1997:ISI

[Pri97] J. Price. Introduction: special issue on structuring complex information for electronic publication. IEEE Transactions on Professional Communication, 40(2):69–77, June 1997. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Price:1999:CAP

[Pri99] J. Price. Christopher Alexander’s pattern language. IEEE Transactions on Professional Communication, 42(2):117– 122, June 1999. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Procter:1994:YHH

[Pro94] J. Procter. You haven’t heard a word I said: getting managers to listen. IEEE Transactions on Professional Communication, 37(1):18–20, March 1994. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Prystash:1995:CAC

[Pry95] D. T. Prystash. Corporate attitudes and commitments to telecommuting. IEEE Transactions on Professional Communication, 38(2):95–99, June 1995. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Philbin:1996:PRS

[PS96] A. I. Philbin and M. M. Spirek.

The paradox of revision: a study of writing as a product in the revision of manuals. IEEE Transactions on Professional Communication, 39(1): 30–37, March 1996. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Peng:2012:FKS

[PS12] Yuan Peng and Juliana Sutanto. Facilitating knowledge sharing through a boundary spanner. IEEE Transactions on Professional Communication, 55(2):142–155, June 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Palmeri:2005:DNR

[PT05] J. Palmeri and P. M. Tuten. Dialogic negotiations: a reflective tale of collaboration across the academic-practitioner divide. IEEE Transactions on Professional Communication, 48(3): 313–323, September 2005. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Pullinger:1986:CCE

[Pul86] D. J. Pullinger. Chit-chat to electronic journals: Computer conferencing supports scientific communication. IEEE Transactions on Professional Communication, PC-29(1):23– 29, March 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Pendharkar:2004:DCM

[PY04] P. C. Pendharkar and K. Young. The development of a construct for measuring an individual’s perceptions of email as a medium for electronic communication in organizations. IEEE Transactions on Professional Communication, 47 (2):130–143, June 2004. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Qureshi:2005:AKT

[QK05] S. Qureshi and P. Keen. Activating knowledge through electronic collaboration: vanquishing the knowledge paradox.

IEEE Transactions on Professional Communication, 48(1): 40–54, March 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Queipo:1986:UEO

[Que86] Larry Queipo. User expectations of online information. IEEE Transactions on Professional Communication, PC-29(4):11–15, December 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Rosenbaum:1997:DWJ

[RA97] S. Rosenbaum and L. Anschuetz. Dennis Wixon and Judith Ramey, Eds., Field Methods Casebook For Software Design. New York: Wiley, 1996, 342 pp. [Book Review]. IEEE Transactions on Professional Communication, 40(3): 239–241, September 1997. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Rabin:1992:MCI

[Rab92] K. M. Rabin. Making company information accessible: how to develop an online information system. IEEE Transactions on Professional Communication, 35(1):2–6, March 1992. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Racker:1959:SWP

[Rac59] Joseph Racker. Selecting and writing to the proper level. IRE Transactions on Engineering Writing and Speech, EWS-2(1): 16–21, January 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Railton:1986:CST

[Rai86] Diana Railton. Communication skills training for engineering students in British universities. IEEE Transactions on Professional Communication, PC-29(2):7–14, June 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Raisanen:2020:UMR

[Räi20]  Tiina Räisänen.  The use

of multimodal resources by technical managers and their peers in meetings using English as the business lingua franca. IEEE Transactions on Professional Communication, 63(2): 172–187, June 2020. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Raju:2012:ICT

[Raj12] Ritu Raju. Intercultural communication training in IT outsourcing companies in India: a case study. IEEE Transactions on Professional Communication, 55(3):262–274, September 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ramey:1989:SBB

[Ram89] J. Ramey. A selected bibliography: a beginner’s guide to usability testing. IEEE Transactions on Professional Communication, 32(4):310– 316, December 1989. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ramsey:1998:ETT

[Ram98] R. D. Ramsey. The Emergence of A Tradition: Technical Writing In The English Renaissance, 1475–1640 [Book Review]. IEEE Transactions on Professional Communication, 41(3):223–225, September 1998. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ramsey:2000:MST

[Ram00] R. D. Ramsey. From Millwrights To Shipwrights To The Twenty-first Century: Explorations In A History of Technical Communication In The United States [Book Review]. IEEE Transactions on Professional Communication, 43(2): 230–232, June 2000. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ramirez:2003:RPI

[Ram03] L. M. Ramirez. Rune Pettersson, Information design: an introduction [Book Review]. IEEE Transactions on Professional Communication, 46(4):

333–334, December 2003. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ramirez:2005:VLP

[Ram05a] L. Ramirez. Value leadership: The 7 principles that drive corporate value in any economy. IEEE Transactions on Professional Communication, 48(1): 109–110, March 2005. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ramirez:2005:GIC

[Ram05b] L. M. Ramirez. Guide to interpersonal communication. IEEE Transactions on Professional Communication, 48(4): 439–440, December 2005. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ramirez:2007:PLS

[Ram07] Lino Ramirez. Putt’s Law and the Successful Technocrat: How to Win in the Information Age (Putt, A.; 2006). IEEE Transactions on Professional Communication, 50(1): 78–80, March 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Raskin:1986:QS

[Ras86] Robin Raskin. The quest for style. IEEE Transactions on Professional Communication, PC-29(3):10–18, September 1986. CODEN IEPCBU.

ISSN 0361-1434 (print), 1558-1500 (electronic).

Rath:1961:NTT

[Rat61] Gustave J. Rath. A new task for the technical writer programming teaching machines. IRE Transactions on Engineering Writing and Speech, EWS-4(1):3–4, January 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Ratliff:1980:PGO

[Rat80] Gerald Lee Ratliff. Performance guide for oral communication. IEEE Transactions on Professional Communication, PC-23(1):11–14, March 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Rathjens:1985:SCC

[Rat85] Dietrich Rathjens. The seven components of clarity in technical writing. IEEE Transactions on Professional Communication, PC-28(4):42–46, December 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Rathjens:1986:REO

[Rat86] Dietrich Rathjens. Reverse engineering: The outline as document restructuring tool. IEEE Transactions on Professional Communication, PC-29(3):19– 22, September 1986. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Raudsepp:1979:PYI

[Rau79] Eugene Raudsepp. Present your ideas effectively. IEEE Transactions on Professional Communication, PC-22(4): 204–210, December 1979. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Raudsepp:1981:TBM

[Rau81] Eugene Raudsepp. Toward better meetings: a psychologist’s view. IEEE Transactions on Professional Communication, PC-24(3):136–138, September 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Raudsepp:1983:SCT

[Rau83] Eugene Raudsepp. Stimulating creative thinking. IEEE Transactions on Professional Communication, PC-26(4):209–211, December 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Rau:2021:RDR

[Rau21] Gerald Rau. Rediscovery of developmental research articles in electrical engineering and description of their macrostructure. IEEE Transactions on Professional Communication, 64(4):427–443, December 2021. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ravin:1988:GES

[Rav88] Y. Ravin. Grammar errors and style weaknesses in a text-critiquing system. IEEE Transactions on Professional Communication, 31(3):108– 115, September 1988. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Raven:1989:SDC

[Rav89] M. E. Raven. A survey of diagrams in CAD documentation. IEEE Transactions on Professional Communication, 32(3): 200–204, September 1989. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ravilla:2006:EIE

[Rav06] Sumana Ravilla. The Elements of International English Style: a Guide to Writing English Correspondence, Reports, Technical Documents, and Internet Pages for a Global Audience, by Edmond H. Weiss. IEEE Transactions on Professional Communication, 49(3): 287–288, September 2006. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Robles:2019:UCM

[RB19] Vincent D. Robles and Matthew J.

Baker. Using case-method pedagogy to facilitate audience awareness. IEEE Transactions on Professional Communication, 62(2):192–207, June 2019.

CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Robson:1959:OL

[RC59] Ernest M. Robson and J. D. Chapline. The orchestra of the language. IRE Transactions on Engineering Writing and Speech, EWS-2(3):93, December 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Robinson:1989:NKI

[RC89] P. A. Robinson and S. S. Courter. A new kind of internship: technical writing for engineers. IEEE Transactions on Professional Communication, 32(3):150–153, September 1989. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Riley:1982:DEW

[RCB82] Michael W. Riley, David J. Cochran, and John L. Ballard. Designing and evaluating warning labels. IEEE Transactions on Professional Communication, PC-25(3):127–130, September 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Rainey:1990:TOR

[RCK^{+}90] K. T. Rainey, P. E. Conners, R. S. Kelly, J. L. Pevey, and H. J. Smith. Texts and other resources for training in technical and scientific communication. IEEE Transactions on Professional Communication, 33(4):205–219, December 1990. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Roberts:2002:PCG

[RCS02] T. L. Roberts, P. H. Cheney, and P. D. Sweeney. Project characteristics and group communication: an investigation. IEEE Transactions on Professional Communication, 45(2): 84–98, June 2002. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Robert:2005:PRC

[RD05] L. P. Robert and A. R. Dennis. Paradox of richness: a cognitive model of media choice. IEEE Transactions on Professional Communication, 48(1): 10–21, March 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Roberts:2016:IVC

[RD16] Nicholas Roberts and Michael Dinger. The impact of virtual customer community interactivity on organizational innovation: An absorptive capacity perspective. IEEE Transactions on Professional Communication, 59(2):110–125, June 2016. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Riley:2009:EDC

[RDJM09] Kathryn Riley, Michael Davis, Apryl Cox Jackson, and James Maciukenas. “Ethics in the Details”: Communicating engineering ethics via microinsertion. IEEE Transactions on Professional Communication, 52(1):95–108, March 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

RULL:2004:DPI

[RDRS04] R. K. Rull, S. J. D’Souza, D. Roy, and D. M. Sharp. Designing procedural illustrations. IEEE Transactions on Professional Communication, 47(1):27–33, March 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Read:1980:HPT

[Rea80] Nat B. Read. How to prepare for the TV interview. IEEE Transactions on Professional Communication, PC-23 (1):45–47, March 1980. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Rea:2022:CES

[Rea22] Ashley Rea. Coding equity: Social justice and computer programming literacy education. IEEE Transactions on Professional Communication, 65(1):87–103, March 2022. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Redish:1981:ULR

[Red81] Janice C. Redish. Understanding the limitations of readability formulas. IEEE Transactions on Professional Communication, PC-24(1):46– 48, March 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Redish:1989:RLD

[Red89] J. C. Redish. Reading to learn to do. IEEE Transactions on Professional Communication, 32(4):289–293, December 1989. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Redish:1998:CAE

[Red98] J. C. Redish. Cognitive Aspects of Electronic Text Processing, Vol. LVIII In The Series Advances In Discourse Processes [Book Review]. IEEE Transactions on Professional Communication, 41(1):77–79, March 1998. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Redish:2010:TCU

[Red10] J. Redish. Technical communication and usability: Intertwined strands and mutual influences. IEEE Transactions on Professional Communication, 53(3):191–201, September 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Reel:1964:PL

[Ree64]  Willard D. Reel.  Project

literacy. IEEE Transactions on Engineering Writing and Speech, EWS-7(3):3–8, December 1964. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Regan:1970:DAI

[Reg70] John E. Regan. The dynamic aspects of information flow within a society. IEEE Transactions on Engineering Writing and Speech, EWS-13(2): 65–73, September 1970. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Rehling:1994:ITM

[Reh94] Louise Rehling. “Is it theirs, mine, or ours?” Ownership, collaboration, and cultures. IEEE Transactions on Professional Communication, 37(1):42–49, March 1994. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Rehling:1998:NAR

[Reh98] L. Rehling. Nostalgic Angels: Rearticulating Hypertext Writing [Book Review]. IEEE Transactions on Professional Communication, 41(1): 80–81, March 1998. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Reich:1978:DS

[Rei78] Jack Reich. Designing slides.

IEEE Transactions on Professional Communication, PC-21 (3):108–110, September 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Reich:1979:HCS

[Rei79] Jack Reich. How to compose slides for lectures, Katharina Meyer-Hartwig, Walter Bleifeld, and Ulf Hegewald. IEEE Transactions on Professional Communication, PC-22 (3):174–175, September 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Reimold:1981:BWC

[Rei81] Cheryl Reimold. Business writing clear and simple. IEEE Transactions on Professional Communication, PC-24(4): 184–185, December 1981. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Reilly:1988:CQW

[Rei88a] A. D. Reilly. Clarification questions that work. IEEE Transactions on Professional Communication, 31(2):93–95, June 1988. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Reitman:1988:SYD

[Rei88b] P. Reitman. Streamlining your documentation using quick references. IEEE Transactions on Professional Communication, 31(2):75–83, June 1988. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Reilly:1999:ESS

[Rei99]  C. A. Reilly.  Essays In

The Study of Scientific Discourse: Methods, Practice, and Pedagogy [Book Review]. IEEE Transactions on Professional Communication, 42(4): 313–315, December 1999. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Reiff:2002:IAS

[Rei02] M. J. Reiff. Interacting with audiences: social influences on the production of scientific writing [Book Review]. IEEE Transactions on Professional Communication, 45(4): 311–313, December 2002. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Remus:1978:SPP

[Rem78] William Remus. Strategies for a publish or perish world. IEEE Transactions on Professional Communication, PC-21 (4):141–143, December 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Renze:1996:IFT

[Ren96] J. L. Renze. Important factors in the technical proposal process according to engineering faculty. IEEE Transactions on Professional Communication, 39(2):87–98, June 1996. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Reubens:1965:CGS

[Reu65] John B. Reubens. Challenge:

Good scientific writing. IEEE Transactions on Engineering Writing and Speech, EWS-8(2): 48–55, December 1965. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Ramey:1987:SBB

[RFH87] Judith Ramey, David Farkas, and Mark Haselkorn. Standards: Breaking down the barriers in electronic technical publishing. IEEE Transactions on Professional Communication, PC-30(4):258, December 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Rifkind:1992:CGI

[RH92] L. J. Rifkind and L. F. Harper.

Cross-gender immediacy behaviors and sexual harassment in the workplace: a communication paradox. IEEE Transactions on Professional Communication, 35(4):236– 241, December 1992. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Rice:1994:DCF

[RH94] R. P. Rice and J. T. Huguley. Describing collaborative forms: a profile of the team-writing process. IEEE Transactions on Professional Communication, 37(3):163–170, September 1994. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ruppel:2001:SKT

[RH01] C. P. Ruppel and S. J. Harrington. Sharing knowledge through intranets: a study of organizational culture and intranet implementation. IEEE Transactions on Professional Communication, 44 (1):37–52, March 2001. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Rhodes:1977:EPC

[Rho77] Sarah N. Rhodes. Editorial Processing Center — prelude to phase III. IEEE Transactions on Professional Communication, PC-20(2):102–105, September 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Rice:2009:DCK

[Ric09] J. A. Rice. Devising collective knowledges for the technical writing classroom: a course-based approach to using Web 2.0 writing technologies in collaborative work. IEEE Transactions on Professional Communication, 52(3): 303–315, September 2009. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ridgway:1987:RMM

[Rid87] Lenore S. Ridgway. Read my mind: What users want from online information. IEEE Transactions on Professional Communication, PC-30(2):87– 90, June 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Rife:2010:EPL

[Rif10] Martine Courant Rife. Ethos, pathos, logos, kairos: Using a rhetorical heuristic to mediate digital-survey recruitment strategies. IEEE Transactions on Professional Communication, 53(3):260–277, September 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Riley:1999:SHH

[Ril99]  K. Riley.  Seven habits

of highly effective writers. IEEE Transactions on Professional Communication, 42 (1):47–51, March 1999. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Riley:2000:BLG

[Ril00] K. Riley. Bureaucratic Language In Government and Business [Book Review]. IEEE Transactions on Professional Communication, 43(2):233– 235, June 2000. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Riley:2001:IBC

[Ril01] K. Riley. International Business Communication, 2nd Edition [Book Review]. IEEE Transactions on Professional Communication, 44(1):62– 63, March 2001. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Riley:2002:AKH

[Ril02a] K. Riley. Anges Kukulska-Hulme, Language and Communication: Essential concepts for user interface and documentation design. IEEE Transactions on Professional Communication, 45(1):54– 56, March 2002. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Riley:2002:DEW

[Ril02b] K. Riley. Designing effective web sites: a concise guide [Book Review]. IEEE Transactions on Professional Communication, 45(3):206– 208, September 2002. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Rines:1959:SLC

[Rin59] Robert H. Rines. Some legal consideration in presenting technical information. IRE Transactions on Engineering Writing and Speech, EWS-2(3): 84–88, December 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Rines:1963:CPP

[Rin63] Robert H. Rines. Current problems in the protection of creative ideas and writings. IEEE Transactions on Engineering Writing and Speech, EWS-6(2): 66–68, December 1963. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Risdon:1998:RDL

[Ris98] K. C. Risdon. Release 2.0:

a Design For Living In The Digital Age [Book Review]. IEEE Transactions on Professional Communication, 41(3): 221–222, September 1998. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Rivers:1985:BRM

[Riv85] William E. Rivers. Book reviews: Marya W. Holcombe and Judith K. Stein, Presentations for decision makers: Strategies for structuring and delivering your ideas. IEEE Transactions on Professional Communication, PC-28 (1):51, March 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ross:1991:CPP

[RK91] S. Ross and B. Karis. Communicating in public policy matters: addressing the problem of non-congruent sites of discourse. IEEE Transactions on Professional Communication, 34(4):247–254, December 1991. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Rojek:1994:DCS

[RK94] J. Rojek and A. Kanerva. A data-collection strategy for usability tests. IEEE Transactions on Professional Communication, 37(3):149–156, September 1994. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Robey:2000:SLC

[RKP00] D. Robey, Huoy Min Khoo, and C. Powers. Situated learning in cross-functional virtual teams. IEEE Transactions on Professional Communication, 43(1):51–66, March 2000. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Rondthaler:1978:SRO

[RL78] Edward Rondthaler and Edward J. Lias. Soundspel: a revised orthography of the English language. IEEE Transactions on Professional Communication, PC-21(1):25– 29, March 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Rosenberg:1984:DCD

[RL84] Jerry M. Rosenberg and Ed Lacy. Dictionary of computers, data processing & telecommunications. IEEE Transactions on Professional Communication, PC-27(4): 231–232, December 1984. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Reilly:1996:RCT

[RL96] C. A. Reilly and B. L’Eplattenier.

Redefining collaboration through the creation of World Wide Web sites. IEEE Transactions on Professional Communication, 39(4):215–223, December 1996. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Rauch:1997:EBM

[RLG97] T. Rauch, P. Leone, and D. Gillihan. Enabling the book metaphor for the World-Wide-Web: disseminating on-line information as dynamic Web documents. IEEE Transactions on Professional Communication, 40(2):111–128, June 1997. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Roberts:2006:EIS

[RLS06] T. L. Roberts, P. B. Lowry, and P. D. Sweeney. An evaluation of the impact of social presence through group size and the use of collaborative software on group member “voice” in faceto-face and computer-mediated task groups. IEEE Transactions on Professional Communication, 49(1):28–43, March 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Racker:1961:TWT

[RM61] Joseph Racker and H. B. M. Technical writing techniques for engineers. IRE Transactions on Engineering Writing and Speech, EWS-4(1):27, January 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Riley:2003:RDD

[RM03] K. Riley and J. Mackiewicz. Resolving the directness dilemma in document review sessions with nonnative speakers. IEEE Transactions on Professional Communication, 46(1):1–16, March 2003. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ritzenthaler:1996:WCC

[RO96] G. Ritzenthaler and D. H. Ostroff. The Web and corporate communication: potentials and pitfalls. IEEE Transactions on Professional Communication, 39(1):16–22, March 1996.

CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Robbins:1972:CAS

[Rob72] J. C. Robbins. Communication among scientists and engineers: a commentary: a PSYCHOM ’72 paper. IEEE Transactions on Professional Communication, PC-15(2):65–67, June 1972. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Robbins:1973:SFS

[Rob73] Jan C. Robbins. Social functions of scientific communication. IEEE Transactions on Professional Communication, PC-16(3):131–135, September 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Robinson:1978:SWS

[Rob78] Jeannine H. Robinson. Successful writing, a short course for professionals, George T. Vardaman and Patricia Black Vardaman. IEEE Transactions on Professional Communication, PC-21(3):129, September 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Robinson:1979:CCV

[Rob79a] Arthur L. Robinson. Communicating with computers by voice. IEEE Transactions on Professional Communication, PC-22(3):159–165, September 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Robinson:1979:WEM

[Rob79b] David Owen Robinson. Where the elite meet. IEEE Transactions on Professional Communication, PC-22(3):130–131, September 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Roberts:1995:RIT

[Rob95] S. S. Roberts. The rhetorical infrastructure of technology transfer as a source for professional growth. IEEE Transactions on Professional Communication, 38(1):41–44, March 1995. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Robinson:1996:PCK

[Rob96]  J. H. Robinson.  Profes-

sional communication in Korea: playing things by eye. IEEE Transactions on Professional Communication, 39(3): 129–134, September 1996. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Robles:2018:RDT

[Rob18] Vincent D. Robles. Resolving discourse at technical-support helpdesks. IEEE Transactions on Professional Communication, 61(3):275–294, September 2018. CODEN IEPCBU.

ISSN 0361-1434 (print), 1558-1500 (electronic).

Rockett:1959:PIF

[Roc59] Frank H. Rockett. Planning illustrations first, simplifies writing later. IRE Transactions on Engineering Writing and Speech, EWS-2(2):56–61, June 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Rohde:2019:DCE

[Roh19] Blake T. Rohde. David C.

Evans: Bottlenecks: Aligning UX Design With User Psychology [Book Review]. IEEE Transactions on Professional Communication, 62(2):208– 209, June 2019. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Rokey:2019:AMM

[Rok19] Laura Rokey. Aaron Marcus, Masaaki Kurosu, Xiaojuan Ma, and Ayako Hashizume: Cuteness Engineering: Designing Adorable Products and Services [Book Review]. IEEE Transactions on Professional Communication, 62(1):107–109, March 2019. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Romm:1981:WG

[Rom81] Ethel Grodzins Romm. Writing guide. IEEE Transactions on Professional Communication, PC-24(3):120–121, September 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Rosenthal:1964:HWT

[Ros64] Murray P. Rosenthal. How to write a technical book and get it published. IEEE Transactions on Engineering Writing and Speech, EWS-7(2):11–20, September 1964. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Rosenberg:1979:WDL

[Ros79] Marcy Rosenberg. War declared on lingo. IEEE Transactions on Professional Communication, PC-22(3):169–170, September 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Rosenburg:1983:EPS

[Ros83] Ronald C. Rosenburg. The engineering presentation some ideas on how to approach and present it. IEEE Transactions on Professional Communication, PC-26(4):191–193, December 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Rosenbaum:1989:UEV

[Ros89] S. Rosenbaum. Usability evaluations versus usability testing: when and why? IEEE Transactions on Professional Communication, 32(4):210– 216, December 1989. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ross:1994:EML

[Ros94] S. M. Ross. Electronic mail:

legal and ethical concerns in the United States and Canada. IEEE Transactions on Professional Communication, 37(4): 218–225, December 1994. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ross:2015:MPL

[Ros15]  Derek G. Ross.  Monkey-

wrenching plain language: Ecodefense, ethics, and the technical communication of ecotage. IEEE Transactions on Professional Communication, 58(2):154–175, June 2015. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Rothkopf:1978:PPP

[Rot78] Michael H. Rothkopf. Publication policy in a primitive scientific society. IEEE Transactions on Professional Communication, PC-21(4):139–140, December 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Roundtree:2017:CMT

[Rou17a] Aimee Kendall Roundtree.

Communicating Mobility and Technology: a Material Rhetoric for Persuasive Transportation [by Pflugfelder, E. H.; Book review]. IEEE Transactions on Professional Communication, 60(1):112–114, March 2017.

CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Roundtree:2017:CRB

[Rou17b] Aimee Kendall Roundtree. Carolyn R. Boiarsky, Risk Communication and Miscommunication: Case Studies in Science, Technology, Engineering, Government, and Community Organization. IEEE Transactions on Professional Communication, 60(3):332– 333, September 2017. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Roundtree:2019:DKB

[Rou19] Aimee Kendall Roundtree. David Kmiec and Bernadette Longo, Eds.: The IEEE Guide to Writing in the Engineering and Technical Fields [Book Review]. IEEE Transactions on Professional Communication, 62(1):110–111, March 2019. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Rowlett:1975:TPS

[Row75] Russell J. Rowlett. A total publication system for scientific information. IEEE Transactions on Professional Communication, PC-18(3):229–233, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Rowe:1978:LEF

[Row78] W. D. Rowe. Law of the estimated fact. IEEE Transactions on Professional Communication, PC-21(2):47–48, June 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ross:2019:PPD

[ROW19] Derek G. Ross, Brett Oppegaard, and Russell Willerton. Principles of place: Developing a place-based ethic for discussing, debating, and anticipating technical communication concerns. IEEE Transactions on Professional Communication, 62(1):4–26, March 2019. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Roy:2007:SCS

[Roy07] Debopriyo Roy. Significance of configuration and subassemblies in sequential procedural instructions and role of textgraphical aid: an explorative study. IEEE Transactions on Professional Communication, 50(2):147–162, June 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Riley:1998:PBV

[RP98] K. Riley and F. Parker. Parallels between visual and textual processing. IEEE Transactions on Professional Communication, 41(3):175–185, September 1998. CODEN IEPCBU.

ISSN 0361-1434 (print), 1558-1500 (electronic).

Redish:1993:MVA

[RR93] J. Redish and J. Ramey. Measuring the value added by professional technical communicators. IEEE Transactions on Professional Communication, 36(3):158–164, September 1993. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Rose:2017:CBU

[RRW^{+}17] Emma J. Rose, Robert Racadio, Kalen Wong, Shally Nguyen, Jee Kim, and Abbie Zahler. Community-based user experience: Evaluating the usability of health insurance information with immigrant patients. IEEE Transactions on Professional Communication, 60(2):214–231, June 2017. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ruark:1980:PDA

[Rua80] Henry C. Ruark. Planning and developing an audiovisual presentation. IEEE Transactions on Professional Communication, PC-23(4):179–181, December 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Rubinoff:1966:RPL

[Rub66] Morris Rubinoff. A rapid procedure for launching a microthesaurus. IEEE Transactions on Engineering Writing and Speech, EWS-9(1):8– 14, July 1966. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Rubens:1986:PCD

[Rub86a] Brenda Rubens. Personality in computer documentation: a preference study. IEEE Transactions on Professional Communication, PC-29(4):56– 60, December 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Rubens:1986:OIT

[Rub86b] Philip Rubens. Online information, traditional page design, and reader expectation. IEEE Transactions on Professional Communication, PC-29(4):75– 80, December 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Rubens:1987:CET

[Rub87] Philip Rubens. The cartoon and ethics: Their role in technical information. IEEE Transactions on Professional Communication, PC-30(3):196–201, September 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Rude:1992:RSI

[Rud92] C. D. Rude. The rhetoric of scientific inquiry. IEEE Transactions on Professional Communication, 35(2):88–90, June 1992. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ruiz:2020:WTM

[Rui20]  Allison Lee Ruiz.  Wear-

able Technology in Medicine and Health Care: Raymond Kai-Yu Tong [Book Review]. IEEE Transactions on Professional Communication, 63(3): 275–276, September 2020. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Russak:1977:PPS

[Rus77] Ben Russak. Publishing primary scientific journals during the next decade. IEEE Transactions on Professional Communication, PC-20(2):51– 54, September 1977. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Rutkowski:2002:CRV

[RVV^{+}02] A. F. Rutkowski, D. R. Vogel, M. Van Genuchten, T. M. A. Bemelmans, and M. Favier. ecollaboration: the reality of virtuality. IEEE Transactions on Professional Communication, 45(4):219–230, December 2002. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Rutkowski:2008:CVT

[RVvGS08] Anne-Françoise Rutkowski, Doug Vogel, Michiel van Genuchten, and Carol Saunders. Communication in virtual teams: Ten years of experience in education. IEEE Transactions on Professional Communication, 51(3):302– 312, September 2008. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Rosenbaum:1986:ADM

[RW86] Stephanie Rosenbaum and R. Dennis Walters. Audience diversity: a major challenge in computer documentation. IEEE Transactions on Professional Communication, PC-29(4):48–55, December 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ray:2024:MID

[RWAK24] Anirban Ray, Ian R. Weaver, G. Edzordzi Agbozo, and Yeqing Kong. Mapping interaction design in global health interventions: a comparative analysis of COVID-19 mHealth technologies. IEEE Transactions on Professional Communication, 67(2):173–191, June 2024. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Raymond:1990:CPP

[RY90] J. Raymond and C. Yee. The collaborative process and professional ethics. IEEE Transactions on Professional Communication, 33(2):77–81, June 1990. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ryder:1985:ICR

[Ryd85] John D. Ryder. An IEEE centennial retrospective. IEEE Transactions on Professional Communication, PC-28(2):2– 6, June 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Still:2010:ETC

[SA10] Brian Still and Michael J. Albers. Editorial: Technical communication and usability studies. IEEE Transactions on Professional Communication, 53 (3):189–190, September 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Shah:2023:CSL

[SA23] Samir Shah and Bay Arinze. Comparing student learning in face-to-face versus online sections of an information technology course. IEEE Transactions on Professional Communication, 66(1):48–58, March 2023. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sable:1961:USS

[Sab61] J. D. Sable. The use of semantic structures in information systems. IRE Transactions on Engineering Writing and Speech, EWS-4(3):83–90, December 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Sacks:1977:RJ

[Sac77] Jean W. Sacks. Rebirth of a journal. IEEE Transactions on Professional Communication, PC-20(2):84–85, September 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sachs:2001:PDE

[Sac01] H. L. Sachs. Paul Dombrowski, Ethics in technical communication [Book Review]. IEEE Transactions on Professional Communication, 44(2): 149–150, June 2001. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sall:1961:ROR

[Sal61] Chester W. Sall. Robert T.

Oliver and Rupert L. Cortright, Effective Speech. IRE Transactions on Engineering Writing and Speech, EWS-4(3): 102, December 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Sall:1962:EOL

[Sal62] Chester W. Sall. Editorial: an open letter to PGEWS members. IRE Transactions on Engineering Writing and Speech, EWS-5(2):37, December 1962. ISSN 0893-7869 (print), 2331-3692 (electronic).

Salton:1968:USD

[Sal68]  Gerard Salton.  The use

of standardized documentary data in automatic information dissemination. IEEE Transactions on Engineering Writing and Speech, EWS-11(2): 101–110, August 1968. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Salton:1980:NNH

[Sal80] Gerard Salton. The network nation: Human communication via computer, S. R. Hiltz and M. Turoff. IEEE Transactions on Professional Communication, PC-23(2):107–108, June 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Samuel:1958:LUO

[Sam58]  Rhys Samuel.  Let’s use

our heads. IRE Transactions on Engineering Writing and Speech, EWS-1(2):13–17, August 1958. ISSN 0893-7869 (print), 2331-3692 (electronic).

Samuelson:1970:FTI

[Sam70] Kjell Samuelson. Free-text indexing to meet computers halfway. IEEE Transactions on Engineering Writing and Speech, EWS-13(2):46–48, September 1970. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Samson:1988:PDR

[Sam88] D. Samson. A publications department’s role in marketing. IEEE Transactions on Professional Communication, 31(2): 88–92, June 1988. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Samson:2006:LDL

[Sam06] C. Samson. … Is different from, …: a corpusbased study of evaluative adjectives in economics discourse. IEEE Transactions on Professional Communication, 49(3): 236–245, September 2006. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sankar:1983:RDN

[San83] C. S. Sankar. Reduction of duplicate names in data entry forms. IEEE Transactions on Professional Communication, PC-26(4):204–208, December 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sanders:1986:PPP

[San86] Scott P. Sanders. Practice, proficiency, professionalism. IEEE Transactions on Professional Communication, PC-29(2):15– 18, June 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sanders:1987:BAT

[San87a] Scott P. Sanders. Balancing acts: a theme and one variation. IEEE Transactions on Professional Communication, PC-30(4):249, December 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sanders:1987:SOW

[San87b] Scott P. Sanders. Subjective objectivity: What I’m teaching now. IEEE Transactions on Professional Communication, PC-30(2):91–96, June 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sanders:1987:WCE

[San87c] Scott P. Sanders. What is communication education and training? IEEE Transactions on Professional Communication, PC-30(1):45–46, March 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sanford:1991:AGR

[San91] C. C. Sanford. Augmenting the gatekeeper’s role: a decision support system for a journal editor. IEEE Transactions on Professional Communication, 34(3):140–146, September 1991. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sapienza:2008:SMA

[Sap08] Filipp Sapienza. A shared meanings approach to intercultural usability: Results of a user study between international and American university students. IEEE Transactions on Professional Communication, 51(2):215–227, June 2008. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Spinuzzi:2020:GNG

[SAP20] Clay Spinuzzi, David Altounian, and Gregory P. Pogue. Go or no go: Learning to persuade in an early-stage student entrepreneurship program. IEEE Transactions on Professional Communication, 63(2):100–117, June 2020. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sargent:1972:RPA

[Sar72] Ann W. Sargent. Requests for proposal: Asking the right question: a PSYCHOM ’72 paper. IEEE Transactions on Professional Communication, PC-15(3):72–76, September 1972. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sarasohn:1973:TJW

[Sar73] Homer M. Sarasohn. The technical journal who needs it? IEEE Transactions on Professional Communication, PC-16 (3):129–130, September 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Saunders:1982:WEA

[Sau82] Arlan Saunders. Writing effective assembly procedures. IEEE Transactions on Professional Communication, PC-25 (1):20–21, March 1982. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sauer:1992:ERM

[Sau92]  B. A. Sauer.  The engi-

neer as rational man: the problem of imminent danger in a non-rational environment. IEEE Transactions on Professional Communication, 35(4): 242–249, December 1992. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sauer:1993:DMN

[Sau93] B. A. Sauer. Demand modeling, new mode problems, and the $64 (sic) question: technological utopianism in America’s race to develop high speed rail technology. IEEE Transactions on Professional Communication, 36(3):144–151, September 1993. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Savakinas:1988:RAW

[Sav88] R. C. Savakinas. Ready, aimwrite. IEEE Transactions on Professional Communication, 31(1):5–7, March 1988. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sawyer:1999:PJK

[Saw99] P. R. Sawyer. Profile of James Kalmbach. IEEE Transactions on Professional Communication, 42(1):61–63, March 1999. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Salton:1979:CSC

[SB79] G. Salton and D. Bergmark. A citation study of computer science literature. IEEE Transactions on Professional Communication, PC-22(3):146–158, September 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Schmidt:1984:CTR

[SB84] Steven Schmidt and Charles E.

Beck. Creating the technical report. IEEE Transactions on Professional Communication, PC-27(1):49, March 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sole:1991:DDK

[SB91] T. D. Sole and G. Bist. The design document: key to managing documentation development. IEEE Transactions on Professional Communication, 34(2):70–74, June 1991. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sole:1995:BTI

[SB95] T. D. Sole and G. Bist. Benchmarking in technical information. IEEE Transactions on Professional Communication, 38(2):77–82, June 1995. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Smart:2000:CCF

[SB00] K. L. Smart and C. Barnum. Communication in crossfunctional teams: an introduction to this special issue. IEEE Transactions on Professional Communication, 43(1): 19–21, March 2000. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Stibic:1981:PDP

[SC81] Vladimir Stibic and Patricia Chaffee. Personal documentation for professionals. IEEE Transactions on Professional Communication, PC-24 (2):106–107, June 1981. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sha:2012:RAR

[SC12] Xiqing Sha and Klarissa T. Chang. Research article the role of leadership and contextualization on citizenship behaviors in distributed teams: a relational capital perspective. IEEE Transactions on Professional Communication, 55(4): 310–324, December 2012. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Schindler:1963:WTE

[Sch63]  G. E. Schindler.  Web-

ster’s Third and the English language. IEEE Transactions on Engineering Writing and Speech, EWS-6(1):7–9, September 1963. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Schindler:1965:TOT

[Sch65] G. E. Schindler. The treasure of our tongue, by Lincoln Barnett. IEEE Transactions on Engineering Writing and Speech, EWS-8(1):36–37, June 1965. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Schultz:1966:STI

[Sch66] Louise Schultz. Scientific and technical information activities: National studies and machine systems. IEEE Transactions on Engineering Writing and Speech, EWS-9(2):36–39, December 1966. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Schindler:1967:WES

[Sch67] George E. Schindler. Why engineers and scientists write as they do — twelve characteristics of their prose. IEEE Transactions on Engineering Writing and Speech, EWS-10(2): 27–32, December 1967. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Schlesinger:1968:FBT

[Sch68] Emily K. Schlesinger. Five books on technical communication. IEEE Transactions on Engineering Writing and Speech, EWS-11(3):166–168, December 1968. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Schwarzlander:1970:ESS

[Sch70] Harry Schwarzlander. Encyclopedic storage of scientific and technical knowledge. IEEE Transactions on Engineering Writing and Speech, EWS-13(2):48–57, September 1970. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Schlesinger:1972:BR

[Sch72a] Emily K. Schlesinger. Book review. IEEE Transactions on Professional Communication, PC-15(4):123–124, December 1972. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Schlesinger:1972:RCS

[Sch72b] Emily K. Schlesinger. Relationships in communication by sound and sight. IEEE Transactions on Professional Communication, PC-15(4):111–114, December 1972. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Schlesinger:1972:WMT

[Sch72c] Emily K. Schlesinger. Words, messages, and technical writing. IEEE Transactions on Professional Communication, PC-15(3):87–88, September 1972. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Scherr:1973:IR

[Sch73a] George H. Scherr. Irreproducible research. IEEE Transactions on Professional Communication, PC-16(3):126–128, September 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Schlesinger:1973:BRa

[Sch73b] Emily K. Schlesinger. Book review. IEEE Transactions on Professional Communication, PC-16(1):23, March 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Schlesinger:1973:BRb

[Sch73c] Emily K. Schlesinger. Book review. IEEE Transactions on Professional Communication, PC-16(2):44–45, June 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Schlesinger:1973:BRc

[Sch73d] Emily K. Schlesinger. Book review. IEEE Transactions on Professional Communication, PC-16(3):182, September 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Schlesinger:1973:BRd

[Sch73e] Emily K. Schlesinger. Book reviews. IEEE Transactions on Professional Communication, PC-16(4):216–217, December 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Schindler:1975:WES

[Sch75a] George E. Schindler. Why engineers and scientists write as they do twelve characteristics of their prose. IEEE Transactions on Professional Communication, PC-18(1):5– 10, March 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Schlesinger:1975:TTW

[Sch75b] Emily Schlesinger. The teaching of technical writing, Ed. Donald C. Cunningham and Herman A. Estrin. IEEE Transactions on Professional Communication, PC-18(4): 325–326, December 1975. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Schlesinger:1975:CWM

[Sch75c] Emily Schlesinger. Correctness in writing: More about Johnny. IEEE Transactions on Professional Communication, PC-18(4):323–324, December 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Schwarzlander:1975:ESS

[Sch75d] Harry Schwarzlander. Encyclopedic storage of scientific and technical knowledge. IEEE Transactions on Professional Communication, PC-18(1):35– 44, March 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Schlesinger:1978:CGE

[Sch78a] Emily Schlesinger. A communicative grammar of English, Geoffrey Leech and Jan Svartvik. IEEE Transactions on Professional Communication, PC-21(2):86–87, June 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Schlesinger:1978:DTR

[Sch78b] Emily Schlesinger. Designing technical reports: Writing for audiences in organizations, J. C. Mathes and D. W. Stevenson. IEEE Transactions on Professional Communication, PC-21(1):38, March 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Schlesinger:1978:HWI

[Sch78c] Emily Schlesinger. How to write information mapping, Robert E. Horn. IEEE Transactions on Professional Communication, PC-21(1): 39, March 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Schlesinger:1978:SSW

[Sch78d] Emily Schlesinger. The science of scientific writing, Judson Monroe, Carole Meredith, and Kathleen Fisher. IEEE Transactions on Professional Communication, PC-21(2):85, June 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Schlesinger:1978:WCD

[Sch78e] Emily Schlesinger. Written communication for data processing, R. S. Smith. IEEE Transactions on Professional Communication, PC-21(3): 130, September 1978. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Schlesinger:1979:SIT

[Sch79] Emily Schlesinger. Scientific information transfer: The editor’s role, Miriam Balaban. IEEE Transactions on Professional Communication, PC-22 (4):218–219, December 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Schlesinger:1980:CMS

[Sch80a] Emily Schlesinger. Creativity: The magic synthesis, Silvano Arieti. IEEE Transactions on Professional Communication, PC-23(2):105–107, June 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Schlesinger:1980:WPH

[Sch80b] Emily Schlesinger. Writing with precision: How to write so that you cannot possibly be misunderstood, Jefferson D. Bates. IEEE Transactions on Professional Communication, PC-23 (2):110, June 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Schlesinger:1980:CC

[Sch80c]  Emily Schlesinger.  Cre-

ativity and communication. IEEE Transactions on Professional Communication, PC-23(2):101, June 1980. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Schoen:1980:LGB

[Sch80d] Robert Schoen. Let’s give better scientific and technical talks. IEEE Transactions on Professional Communication, PC-23(3):120–122, September 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Schiff:1981:SAF

[Sch81a] Peter M. Schiff. Speech: Another facet of technical communication. IEEE Transactions on Professional Communication, PC-24(2):89–90, June 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Schlesinger:1981:VC

[Sch81b] Emily Schlesinger. Verbal camouflage. IEEE Transactions on Professional Communication, PC-24(2):102, June 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Schoenfeld:1981:HDI

[Sch81c] Robert Schoenfeld. Human dimensions for inanimate objects? IEEE Transactions on Professional Communication, PC-24(4):182–183, December 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Schoenfeld:1981:WMN

[Sch81d] Robert Schoenfeld. When to make nouns go into action. IEEE Transactions on Professional Communication, PC-24 (3):123, September 1981. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Schoenfeld:1982:DCV

[Sch82] Robert Schoenfeld. Dollarsand-cents value of efficient presentation. IEEE Transactions on Professional Communication, PC-25(3):144–150, September 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Schlesinger:1983:EYB

[Sch83] Emily K. Schlesinger. Edit yourself, Bruce Ross-Larson.

IEEE Transactions on Professional Communication, PC-26 (2):87–88, June 1983. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Schell:1986:TOP

[Sch86] David A. Schell. Testing online and print user documentation. IEEE Transactions on Professional Communication, PC-29(4):87–92, December 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Schriver:1989:ETQ

[Sch89] K. A. Schriver. Evaluating text quality: the continuum from text-focused to reader-focused methods. IEEE Transactions on Professional Communication, 32(4):238–255, December 1989. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Schreiber:1993:AWJ

[Sch93] E. J. Schreiber. From academic writing to job-related writing: achieving a smooth transition. IEEE Transactions on Professional Communication, 36(4): 178–184, December 1993. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Schaefer:2010:RCL

[Sch10]  K. Schaefer.  Response-to-

complaint letter as a rhetorical genre. IEEE Transactions on Professional Communication, 53(2):158–163, June 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Schriver:2017:PLU

[Sch17] Karen A. Schriver. Plain language in the US gains momentum: 1940–2015. IEEE Transactions on Professional Communication, 60(4):343– 383, December 2017. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Schneider:2020:GNI

[Sch20] Jessica Ann Schneider. Grantsmanship for New Investigators: Thomas F. Hilton and Carl G. Leukefeld [Book Review]. IEEE Transactions on Professional Communication, 63(1): 98–99, March 2020. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Shi:2020:IMG

[SCW20] Xingsong Shi, Saul Carliner, and Wenjing Wan. Internetmediated genre studies: an integrative literature review (2005–2019). IEEE Transactions on Professional Communication, 63(4):279–295, December 2020. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sengupta:2006:QTU

[SD06] A. Sengupta and A. Dillon. Query by templates: using the shape of information to search next-generation databases. IEEE Transactions on Professional Communication, 49(2):128–144, June 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Shachak:2013:UMP

[SDB^{+}13] Aviv Shachak, Rustam Dow, Jan Barnsley, Karen Tu, Sharon Domb, Alejandro R. Jadad, and Louise Lemieux-Charles. User manuals for a primary care electronic medical record system: a mixedmethods study of user- and vendor-generated documents. IEEE Transactions on Professional Communication, 56(3): 194–209, September 2013. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Searle:1973:PPS

[Sea73] Shayle R. Searle. The publishor-perish syndrome. IEEE Transactions on Professional Communication, PC-16(3): 136–139, September 1973. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sedgwick:1969:CCS

[Sed69] H. D. Sedgwick. Computercontrolled storage, retrieval, update, and print out. IEEE Transactions on Engineering Writing and Speech, EWS-12 (2):57, May 1969. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Seideman:1967:IER

[Sei67] Irving M. Seideman. Improving engineering reports and talks. IEEE Transactions on Engineering Writing and Speech, EWS-10(1):3–9, July 1967. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Seideman:1972:BR

[Sei72a] Irving M. Seideman. Book review. IEEE Transactions on Professional Communication, PC-15(3):104–105, September 1972. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Seideman:1972:E

[Sei72b] Irving M. Seideman. Editorial. IEEE Transactions on Professional Communication, PC-15 (2):19, June 1972. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Seideman:1972:ENR

[Sei72c] Irving M. Seideman. Editorial: New recognition for the communications professional. IEEE Transactions on Professional Communication, PC-15 (4):107, December 1972. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sekey:1973:ACS

[Sek73] Andrew A. Sekey. Abstract, conclusions and summaries. IEEE Transactions on Professional Communication, PC-16 (2):25–26, June 1973. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Selber:1995:MPH

[Sel95] S. A. Selber. Metaphorical perspectives on hypertext. IEEE Transactions on Professional Communication, 38(2):59–67, June 1995. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Selber:1998:WNB

[Sel98] S. A. Selber. The Wired Neighborhood [Book Review]. IEEE Transactions on Professional Communication, 41(1): 82–83, March 1998. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Selber:1999:CTC

[Sel99]  S. A. Selber.  Computers

and technical communication: pedagogical and programmatic perspectives. IEEE Transactions on Professional Communication, 42(2):135–136, June 1999. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sencer:1962:RCP

[Sen62] Robert A. Sencer. Research in communication: a progress report. IRE Transactions on Engineering Writing and Speech, EWS-5(2):44–49, December 1962. ISSN 0893-7869 (print), 2331-3692 (electronic).

Siau:2010:ENC

[SEN10] Keng Siau, John Erickson, and Fiona Fui-Hoon Nah. Effects of national culture on types of knowledge sharing in virtual communities. IEEE Transactions on Professional Communication, 53(3):278– 292, September 2010. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Serna:1988:MSP

[Ser88] B. Serna. My student project leader experience. IEEE Transactions on Professional Communication, 31(4):195– 197, December 1988. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sheehan:1999:GRI

[SF99] R. Johnson Sheehan and A. Flood. Genre, rhetorical interpretation, and the open case: teaching the analytical report. IEEE Transactions on Professional Communication, 42(1):20–31, March 1999. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Spyridakis:2002:EIV

[SF02] J. H. Spyridakis and W. Fukuoka.

The effect of inductively versus deductively organized text on American and Japanese readers. IEEE Transactions on Professional Communication, 45(2):99–114, June 2002. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Strother:2009:LEI

[SFM09] Judith B. Strother, Zohra Fazal, and Melinda Millsap. Legal and ethical issues of the corporate blogosphere. IEEE Transactions on Professional Communication, 52(3):243– 253, September 2009. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Seiling:2024:BPP

[SGM^{+}24] Lukas Seiling, Rita Gsenger, Filmona Mulugeta, Marte Henningsen, Lena Mischau, and Marie Schirmbeck. Beware: Processing of personal data informed consent through risk communication. IEEE Transactions on Professional Communication, 67(1):4–25, March 2024. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sankar:1991:RUI

[SH91] C. S. Sankar and W. H.

Hawkins. The role of user interface professionals in large software projects. IEEE Transactions on Professional Communication, 34(2):94–100, June 1991. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Scholtz:1993:UTM

[SH93] J. Scholtz and M. Hansen. Usability testing a minimal manual for the Intel SatisFAXtion faxmodem. IEEE Transactions on Professional Communication, 36(1):7–11, March 1993. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Suchan:2001:CCV

[SH01] J. Suchan and G. Hayzak. The communication characteristics of virtual teams: a case study. IEEE Transactions on Professional Communication, 44(3): 174–186, September 2001. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sharp:1961:LLP

[Sha61] Harold S. Sharp. Library and laboratory: Partners in research. IRE Transactions on Engineering Writing and Speech, EWS-4(2):58–61, May 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Shatzkin:1968:SPE

[Sha68] Leonard Shatzkin. Some possible effects of computer composition technology on technical book publication. IEEE Transactions on Engineering Writing and Speech, EWS-11(2): 140–144, August 1968. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Sharma:1982:GER

[Sha82] B. C. Sharma. Graphic evaluation of readability scores. IEEE Transactions on Professional Communication, PC-25 (1):34–35, March 1982. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Shay:1989:GDU

[Sha89] L. A. Shay. The great debate over unclassified information: national security versus scientific freedom. IEEE Transactions on Professional Communication, 32(3):139– 148, September 1989. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Shand:1994:UMPa

[Sha94a] R. McIntosh Shand. User manuals as project management tools. I. Theoretical background. IEEE Transactions on Professional Communication, 37(2):75–80, June 1994. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Shand:1994:UMPb

[Sha94b] R. McIntosh Shand. User manuals as project management tools. II. Practical applications. IEEE Transactions on Professional Communication, 37(3):123–142, September 1994. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Shen:2008:LSO

[SHB08] Jia Shen, Starr Roxanne Hiltz, and Michael Bieber. Learning strategies in online collaborative examinations. IEEE Transactions on Professional Communication, 51(1):63–78, March 2008. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Shephard:1973:SED

[She73] David A. E. Shephard. Some effects of delay in publication of information in medical journals, and implications for the future. IEEE Transactions on Professional Communication, PC-16(3):143–147, September 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Shepley:2006:SSS

[She06]  N. Shepley.  Study skills:

a student survival guide. IEEE Transactions on Professional Communication, 49 (1):85–86, March 2006. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Shirrell:1975:NPO

[Shi75] Robert Shirrell. New production options for the journal publisher. IEEE Transactions on Professional Communication, PC-18(3):150–153, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Shinoda:1981:PJS

[Shi81] Yoshiaki Shinoda. Pitfalls for Japanese specialists in English technical writing. IEEE Transactions on Professional Communication, PC-24(3):151–152, September 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Shinoda:1982:TET

[Shi82]  Yoshiaki Shinoda.  Teach-

ing English technical writing to Japanese specialists. IEEE Transactions on Professional Communication, PC-25 (3):140–143, September 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Showstack:1980:DPN

[Sho80] Richard Showstack. Discourse punctuation: The next stage of printing. IEEE Transactions on Professional Communication, PC-23(3):143–147, September 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Shrimpton:1972:TCP

[Shr72] Rose S. Shrimpton. Technical communications and psychophysics: a PSYCHOM ’72 paper. IEEE Transactions on Professional Communication, PC-15(3):77–79, September 1972. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Shrivastava:2020:MCD

[Shr20] Archana Shrivastava. Are Millennials communication deficient? Solving a generational puzzle in an Indian context. IEEE Transactions on Professional Communication, 63(3): 259–271, September 2020. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Spyridakis:1997:MTS

[SHS97] J. H. Spyridakis, H. Holmback, and S. K. Shubert. Measuring the translatability of Simplified English in procedural documents. IEEE Transactions on Professional Communication, 40(1):4–12, March 1997. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Shultz:1993:IYW

[Shu93] K. Shultz. If you want to speak: imagination in speechmaking. IEEE Transactions on Professional Communication, 36(4): 201–205, December 1993. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Srinivasan:2006:CMT

[SI06] A. Srinivasan and G. Irwin. Communicating the message:translating tasks into queries in a database context. IEEE Transactions on Professional Communication, 49 (2):145–159, June 2006. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sides:1987:UAM

[Sid87] Charles H. Sides. Using attitude measurement to validate consultant-taught writing courses. IEEE Transactions on Professional Communication, PC-30(2):97–100, June 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Siegel:1978:LCL

[Sie78] Alan Siegel. To lift the curse off legalese simplify, simplify. IEEE Transactions on Professional Communication, PC-21 (1):31–32, March 1978. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Simpson:1996:ATS

[Sim96a] D. Simpson. American Technological Sublime [Book Review]. IEEE Transactions on Professional Communication, 39(4): 242, December 1996. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Simpson:1996:CCT

[Sim96b] D. Simpson. Computers and Culture: Two Views of the Twenty-First Century [Book Reviews]. IEEE Transactions on Professional Communication, 39(1):58, March 1996. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sims:2022:TOO

[Sim22] Michela Sims. Tools for overcoming oppression: Plain language and human-centered design for social justice. IEEE Transactions on Professional Communication, 65(1):11– 33, March 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Spruijt:1999:ITF

[SJ99] S. Spruijt and C. Jansen. The influence of task and format on reading results with an online text. IEEE Transactions on Professional Communication, 42(2):92–99, June 1999. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sangwan:2020:ACB

[SJD20] Raghvinder S. Sangwan, Kathryn W. Jablokow, and Joanna F. De-Franco. Asynchronous collaboration: Bridging the cognitive distance in global software development projects.

IEEE Transactions on Professional Communication, 63(4): 361–371, December 2020. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sun:2018:MUC

[SJYZ18] Ya Sun, Guangsa Jin, Yingli Yang, and Jingjing Zhao. Metaphor use in Chinese and American CSR reports. IEEE Transactions on Professional Communication, 61(3):295– 310, September 2018. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Skelton:1990:DCS

[Ske90] T. M. Skelton. Designing communication systems for decentralized organizations: a new role for technical communicators. IEEE Transactions on Professional Communication, 33(2):83–88, June 1990. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Skolnik:1979:HAP

[Sko79]  Herman Skolnik.  Histori-

cal aspects of patent systems. IEEE Transactions on Professional Communication, PC-22 (2):59–63, June 1979. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Skolnik:1981:CTI

[Sko81] Herman Skolnik. Communicating technical information. IEEE Transactions on Professional Communication, PC-24 (2):72, June 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Skopec:1983:PSA

[Sko83] Eric W. Skopec. A problemsolving approach to preparing professional presentations. IEEE Transactions on Professional Communication, PC-26 (1):30–35, March 1983. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sun:2023:INC

[SKZ23] Ya Sun, Deyi Kong, and Limei Zhai. To interact and to narrate: a categorical multidimensional analysis of Twitter use by US banks and energy corporations. IEEE Transactions on Professional Communication, 66(2):117–130, June 2023. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sparkes:1979:CDO

[SLP^{+}79] John J. Sparkes, David Liddell, Christopher A. Pinches, Graham A. Read, and Peter I. Zorkoczy. Communication developments in the open university. IEEE Transactions on Professional Communication, PC-22(4):182–186, December 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sampath:1966:TWI

[SM66] Pushpa Sampath and V. S.

Murthy. Technical writing in India. IEEE Transactions on Engineering Writing and Speech, EWS-9(2):30–32, December 1966. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Sherry:1998:DCD

[SM98] L. Sherry and K. M. Myers.

The dynamics of collaborative design. IEEE Transactions on Professional Communication, 41(2):123–139, June 1998. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Still:2010:BPT

[SM10] Brian Still and John Morris. The blank-page technique: Reinvigorating paper prototyping in usability testing. IEEE Transactions on Professional Communication, 53(2): 144–157, June 2010. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Smith:1960:OCT

[Smi60] Eric M. Smith. Organizing a complex technical proposal effort. IRE Transactions on Engineering Writing and Speech, EWS-3(1):28–30, April 1960.

ISSN 0893-7869 (print), 2331-3692 (electronic).

Smith:1979:VP

[Smi79]  Frank R. Smith.  Verbal

pollution. IEEE Transactions on Professional Communication, PC-22(3):165–169, September 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Smith:1984:MW

[Smi84] Frank Smith. Myths of writing. IEEE Transactions on Professional Communication, PC-27 (2):101–104, June 1984. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Smith:1985:MTT

[Smi85] Herb Smith. Methods for training the technical editor in interpersonal skills. IEEE Transactions on Professional Communication, PC-28(1):46– 50, March 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Smith:2003:SMW

[Smi03] A. M. K. Smith. Scientists must write: a guide to better writing for scientists, engineers, and students [Book Review]. IEEE Transactions on Professional Communication, 46(2):142–143, June 2003. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Smith:2010:DHC

[Smi10] Elizabeth O. Smith. Documentation: a History and Critique of Attribution, Commentary, Glosses, Marginalia, Notes, Bibliographies, Works-Cited Lists, and Citation Indexing and Analysis (Hauptman, R.) [Book Review]. IEEE Transactions on Professional Communication, 53(3):317– 319, September 2010. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Smith:2020:CAF

[Smi20] Jordan Smith. A content analysis of figure captions in academic journals from four disciplines. IEEE Transactions on Professional Communication, 63(4):341–360, December 2020. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sonka:2021:FTA

[SMP21] Kate Sonka, Casey McArdle, and Liza Potts. Finding a teaching A11y: Designing an accessibility-centered pedagogy. IEEE Transactions on Professional Communication, 64(3):264–274, September 2021. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Smart:1996:EDC

[SMS96] K. L. Smart, J. L. Madrigal, and K. K. Seawright. The effect of documentation on customer perception of product quality. IEEE Transactions on Professional Communication, 39(3): 157–162, September 1996. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Snodey:1964:IRC

[Sno64] S. Richard Snodey. Information retrieval — a comprehensive indexed bibliography of 1957–1961 world literature. IEEE Transactions on Engineering Writing and Speech, EWS-7(1):22–38, March 1964. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Spinuzzi:2014:MPE

[SNT^{+}14] Clay Spinuzzi, Scott Nelson, Keela S. Thomson, Francesca Lorenzini, Rosemary A. French, Gregory Pogue, Sidney D. Burback, and Joel Momberger. Making the pitch: Examining dialogue and revisions in entrepreneurs’ pitch decks. IEEE Transactions on Professional Communication, 57(3):158– 181, September 2014. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Spinuzzi:2015:RPR

[SNT^{+}15] Clay Spinuzzi, Scott Nelson, Keela S. Thomson, Francesca Lorenzini, Rosemary A. French, Gregory Pogue, Sidney D. Burback, and Joel Momberger. Remaking the pitch: Reuse strategies in entrepreneurs pitch decks. IEEE Transactions on Professional Communication, 58 (1):45–68, March 2015. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Shaw:1998:LSA

[SO98] P. Shaw and A. Okamura. The letter of submission: avoiding the promotional genre. IEEE Transactions on Professional Communication, 41(4): 274–276, December 1998. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Somerick:1989:UPI

[Som89] N. M. Somerick. Using the phase III public relations professional to improve corporate communication and to help prevent crises. IEEE Transactions on Professional Communication, 32(2):101–105, June 1989. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sondel:1962:TFT

[Son62] Bess Sondel. Toward a field theory of communication. IRE Transactions on Engineering Writing and Speech, EWS-5(2): 50–54, December 1962. ISSN 0893-7869 (print), 2331-3692 (electronic).

Sondel:1965:ORT

[Son65] Bess Sondel. Operations research: a two-stage semantic process. IEEE Transactions on Engineering Writing and Speech, EWS-8(1):34–35, June 1965. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Sophar:1975:VFU

[Sop75] Gerald J. Sophar. Vestigiality of fair use. IEEE Transactions on Professional Communication, PC-18(3):220–221, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Soroka:1975:LPP

[Sor75] Marguerite C. Soroka. Librarians and publishers: Partners. IEEE Transactions on Professional Communication, PC-18 (3):216–219, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Southworth:1972:WST

[Sou72] Mason P. Southworth. Why some technical professionals write and some don’t. IEEE Transactions on Professional Communication, PC-15(3):89– 94, September 1972. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Southworth:1975:WST

[Sou75] Mason P. Southworth. Why some technical professionals write and some don’t. IEEE Transactions on Professional Communication, PC-18(1):29– 30, March 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Souther:1985:IIN

[Sou85a] James W. Souther. Identifying the informational needs of readers; a management responsibility. IEEE Transactions on Professional Communication, PC-28(3):9–12, September 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Souther:1985:WR

[Sou85b] James W. Souther. What to report. IEEE Transactions on Professional Communication, PC-28(3):5–8, September 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Southard:1988:ELP

[Sou88]  S. Southard.  Experiential

learning prepares students to assume professional roles. IEEE Transactions on Professional Communication, 31(4): 157–159, December 1988. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Southard:1996:PAU

[Sou96] S. Southard. Practical Approaches to Usability Testing for Technical Documentation [Book Reviews]. IEEE Transactions on Professional Communication, 39(2):105, June 1996. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sampath:1986:ITP

[SP86] G. Sampath and Angelina Piccoli. An introduction to text processing. IEEE Transactions on Professional Communication, PC-29(3):36–37, September 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Smura:1987:TNB

[SP87] Edwin J. Smura and Archie D. Provan. Toward a new beginning: the development of a standard for font and character encoding to control electronic document interchange. IEEE Transactions on Professional Communication, PC-30 (4):259–264, December 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Spada:2014:HED

[Spa14] Kathleen Spada. Higher Education in the Digital Age [Book Review]. IEEE Transactions on Professional Communication, 57(2):150–153, June 2014. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Spence:1960:HGA

[Spe60] A. N. Spence. How the graphic arts can help engineering communication. IRE Transactions on Engineering Writing and Speech, EWS-3(1):16–19, April 1960. ISSN 0893-7869 (print), 2331-3692 (electronic).

Spitzer:1986:WSC

[Spi86] Michael Spitzer. Writing style in computer conferences. IEEE Transactions on Professional Communication, PC-29(1):19– 22, March 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Spinuzzi:1994:DKF

[Spi94] C. Spinuzzi. A different kind of forum: rethinking rhetorical strategies for electronic text media. IEEE Transactions on Professional Communication, 37(4):213–217, December 1994. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Spilka:1996:PWC

[Spi96] R. Spilka. Professional Writing in Context: Lessons from Teaching and Consulting in Worlds of Work [Book Review]. IEEE Transactions on Professional Communication, 39(4):240, December 1996. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Spinuzzi:2016:ISI

[Spi16]  Clay Spinuzzi.  Introduc-

tion to the special issue on entrepreneurship communication. IEEE Transactions on Professional Communication, 59(4):316–322, December 2016. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sparrow:1984:TBC

[SPR84] W. Keats Sparrow, Nell Ann Pickett, and Diana C. Reep. Technical and business communication in two-year programs. IEEE Transactions on Professional Communication, PC-27 (1):46, March 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sprey:1997:VCT

[Spr97] J. A. Sprey. Videoconferencing as a communication tool. IEEE Transactions on Professional Communication, 40(1): 41–47, March 1997. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Spyridakis:1998:KSD

[Spy98]  J. H. Spyridakis.  Karen

Schriver, Dynamics In Document Design. IEEE Transactions on Professional Communication, 41(2):152–154, June 1998. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sigband:1983:CMB

[SR83a] Norman B. Sigband and Michael J. Rossi. Communication for management and business. IEEE Transactions on Professional Communication, PC-26(2):91, June 1983. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Stibic:1983:TMT

[SR83b] V. Stibic and Michael J. Rossi.

Tools of the mind: Techniques and methods for intellectual work. IEEE Transactions on Professional Communication, PC-26(1):44, March 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Schulz:1997:DUG

[SRVR97] E. Schulz, J. Ramey, M. Van Alphen, and W. Rasnake. Discovering user-generated metaphors through usability testing. IEEE Transactions on Professional Communication, 40(4):255–264, December 1997. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Strate:1987:MLH

[SS87] Larry Strate and Skip Swerdlow. The maze of the law: How technical writers can research and understand legal matters. IEEE Transactions on Professional Communication, PC-30 (3):136–148, September 1987.

CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Scott:2010:ERC

[SS10] Christopher Scott and Saonee Sarker. Examining the role of the communication channel interface and recipient characteristics on knowledge internalization: a pragmatist view. IEEE Transactions on Professional Communication, 53(2): 116–131, June 2010. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Scheffler:1974:STA

[SSM74] F. L. Scheffler, H. H. Schumacher, and J. F. March. The significance of titles, abstracts, and other portions of technical documents for information retrieval. IEEE Transactions on Professional Communication, PC-17(1):1–8, March 1974. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sarker:2005:KTV

[SSNJ05] S. Sarker, S. Sarker, D. B. Nicholson, and K. D. Joshi. Knowledge transfer in virtual systems development teams: an exploratory study of four key enablers. IEEE Transactions on Professional Communication, 48(2):201–218, June 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sarker:2009:SRT

[SSS09] Saonee Sarker, Suprateek Sarker, and Christoph Schneider. Seeing remote team members as leaders: a study of US–Scandinavian teams. IEEE Transactions on Professional Communication, 52 (1):75–94, March 2009. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Siebenhandl:2013:UCD

[SSS^{+}13] Karin Siebenhandl, Günther Schreder, Michael Smuc, Eva Mayr, and Manuel Nagl. A user-centered design approach to self-service ticket vending machines. IEEE Transactions on Professional Communication, 56(2):138–159, June 2013. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Stenstrom:2008:OHG

[SSSC08] Eric Stenstrom, Philippe Stenstrom, Gad Saad, and Soumaya Cheikhrouhou. Online hunting and gathering: an evolutionary perspective on sex differences in Website preferences and navigation. IEEE Transactions on Professional Communication, 51(2):155–168, June 2008. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sabaj:2023:BAO

[SSV^{+}23] Omar Sabaj, Clay Spinuzzi, Germán Varas, Paula Cabezas, and Valentin Gerard. The basis of aaaalll of our program! The start-up Chile playbook as metagenre. IEEE Transactions on Professional Communication, 66(3):300–316, September 2023. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Soderlund:2017:TUA

[SSW17] Lars Söderlund, John Spartz, and Ryan Weber. Taken under advisement: Perspectives on advisory boards from across technical communication. IEEE Transactions on Professional Communication, 60(1):76–96, March 2017. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

StJohn:1985:TDT

[St.85] Mary St. John. The teradata document template: a planning tool. IEEE Transactions on Professional Communication, PC-28(1):29–33, March 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Scott:1999:CTU

[ST99a] C. R. Scott and C. E. Timmerman. Communication technology use and multiple workplace identifications among organizational teleworkers with varied degrees of virtuality. IEEE Transactions on Professional Communication, 42(4): 240–260, December 1999. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

StAmant:1999:WCR

[St.99b] K. St. Amant. When culture and rhetoric contrast: examining English as the international language of technical communication. IEEE Transactions on Professional Communication, 42(4):297–300, December 1999. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

StAmant:2000:CCR

[St.00a] K. St. Amant. Competitive Communication: a Rhetoric For Modern Business [Book Review]. IEEE Transactions on Professional Communication, 43(2):224–225, June 2000. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

StAmant:2000:ETU

[St.00b] K. St. Amant. Expanding translation use to improve the quality of technical communication. IEEE Transactions on Professional Communication, 43(3):323–326, September 2000. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

StAmant:2001:ERI

[St.01a] K. St. Amant. Exploring the rhetoric of international professional communication: an agenda for teachers and researchers [Book Review]. IEEE Transactions on Professional Communication, 44(3):221– 223, September 2001. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

StAmant:2001:TLT

[St.01b] K. St. Amant. Technology and Literacy In The Twentyfirst Century: The Importance of Paying Attention [Book Review]. IEEE Transactions on Professional Communication, 44(1):67–70, March 2001. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

StAmant:2001:WSR

[St.01c] K. St. Amant. Web.studies: rewiring media studies for the digital age [Book Review]. IEEE Transactions on Professional Communication, 44(4): 304–305, December 2001. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

StAmant:2002:LCC

[St.02a] K. St. Amant. Leading in a culture of change [Book Review]. IEEE Transactions on Professional Communication, 45(3): 198–200, September 2002. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

StAmant:2002:VRG

[St.02b] K. St. Amant. Visual revelations: graphical tales of fate and deception from Napoleon Bonaparte to Ross Perot [Book Review]. IEEE Transactions on Professional Communication, 45(4):307–308, December 2002. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

StAmant:2002:TSC

[St.02c] K. St. Amant. Translating into success: cutting-edge strategies for going multinational in a global age. IEEE Transactions on Professional Communication, 45(2):140–141, June 2002. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

StAmant:2003:BBW

[St.03] K. St. Amant. Beyond borders: Web globalization strategies [Book Review]. IEEE Transactions on Professional Communication, 46(3):241– 243, September 2003. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

StAmant:2004:CIG

[St.04] K. St. Amant. The cultural imperative: Global trends in the 21st Century. IEEE Transactions on Professional Communication, 47(1):78–79, March 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Siau:2005:TCI

[ST05a] Keng Siau and Xin Tan. Technical communication in information systems development: the use of cognitive mapping. IEEE Transactions on Professional Communication, 48(3): 269–284, September 2005. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

StAmant:2005:DG

[St.05b] K. St. Amant. In defense of globalization. IEEE Transactions on Professional Communication, 48(1):107–108, March 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Siau:2006:CMT

[ST06a] K. Siau and X. Tan. Cognitive mapping techniques for userdatabase interaction. IEEE Transactions on Professional Communication, 49(2):96–108, June 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

StAmant:2006:OEG

[St.06b] Kirk St. Amant. Online education: Global questions, local answers. IEEE Transactions on Professional Communication, 49(4):368–369, December 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

StAmant:2007:ISS

[St.07] Kirk St. Amant. Introduction to the special section on examining international outsourcing: Perspectives, practices, and projections. IEEE Transactions on Professional Communication, 50(2):81–84, June 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

StAmant:2009:RDJ

[St.09] Kirk St. Amant. Ronald Deibert, John Palfrey, Rafal Rohozinski, and Jonathan Zittrain, Editors, Access Denied: The Practice and Policy of Global Internet Filtering — Book Review. IEEE Transactions on Professional Communication, 52(4):413– 414, December 2009. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Staiger:1973:SAD

[Sta73] David L. Staiger. Separate article distribution as an alternate to journal publication. IEEE Transactions on Professional Communication, PC-16 (3):107–112, September 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Stamant:1999:GNC

[Sta99] K. Stamant. Guide To Nontraditional Careers In Science [Book Review]. IEEE Transactions on Professional Communication, 42(4):316– 317, December 1999. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Stafford:2008:ISI

[Sta08] Tom Stafford. Introduction to the special issue on instant messaging in the workplace. IEEE Transactions on Professional Communication, 51(4): 350–351, December 2008. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Stanchevici:2013:SGC

[Sta13] Dmitri Stanchevici. Stalinist Genetics: The Constitutional Rhetoric of T. D. Lysenko (Standchevici, D.) [Book Review]. IEEE Transactions on Professional Communication, 56(1):83–85, March 2013. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Stanton:2017:CER

[Sta17] Rhonda Stanton. Communicating with employees: Resisting the stereotypes of generational cohorts in the workplace. IEEE Transactions on Professional Communication, 60(3): 256–272, September 2017. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Starks:2020:TCA

[Sta20] Caitlin E. Starks. Technical Communication After the Social Justice Turn: Building Coalitions for Action: Rebecca Walton, Kristen R. Moore, and Natasha N. Jones [Book Review]. IEEE Transactions on Professional Communication, 63(2):205–206, June 2020. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Stevens:1962:BCC

[Ste62] Raymond Stevens. Bottoms up communications for creative innovation. IRE Transactions on Engineering Writing and Speech, EWS-5(2):39–43, December 1962. ISSN 0893-7869 (print), 2331-3692 (electronic).

Stettinius:1977:WVS

[Ste77] Wallace Stettinius. Web versus sheetfed printing. IEEE Transactions on Professional Communication, PC-20(2):68– 69, September 1977. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sternglass:1999:TKT

[Ste99] M. S. Sternglass. Time to know them: a longitudinal study of writing and learning at the college level. IEEE Transactions on Professional Communication, 42(2):139–140, June 1999. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

STEEHOUDER:2004:APK

[Ste04] M. Steehouder. Acquiring procedural knowledge of a technology interface: introduction to this special issue. IEEE Transactions on Professional Communication, 47(1):1–4, March 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Steward:2005:MPP

[Ste05] S. Steward. Management principles and practices for technical communicators. IEEE Transactions on Professional Communication, 48(2):226– 227, June 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Stephens:2008:OCW

[Ste08] Keri K. Stephens. Optimizing costs in workplace instant messaging use. IEEE Transactions on Professional Communication, 51(4):369–380, December 2008. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Stevens:2022:DGS

[Ste22] Hannah L. Stevens. I don’t have a gun stop shooting: Rhetorical analysis of law enforcement use of force policy documents. IEEE Transactions on Professional Communication, 65(1):104–117, March 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Still:1959:CTS

[Sti59] Edwin W. Still. Can they see what you say when you speak? IRE Transactions on Engineering Writing and Speech, EWS-2(1):24–29, January 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Singh:2009:COS

[STM09] A. Singh, A. Taneja, and G. Mangalaraj. Creating online surveys: some wisdom from the trenches tutorial. IEEE Transactions on Professional Communication, 52(2): 197–212, June 2009. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Stoakley:2005:GMR

[Sto05] C. D. Stoakley. Guide to media relations. IEEE Transactions on Professional Communication, 48(4):430–431, December 2005. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Stoakley:2008:WPD

[Sto08] Candace D. Stoakley. Web Portfolio Design and Applications (J. DiMarco; 2006) [Book Review]. IEEE Transactions on Professional Communication, 51(1):110–111, March 2008. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Stolley:2009:UMG

[Sto09] Karl Stolley. Using microformats: Gateway to the Semantic Web. IEEE Transactions on Professional Communication, 52(3):291–302, September 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Stone:2015:SRD

[Sto15]  Ronald L. Stone.  Slide

Rules: Design, Build, and Archive Presentations in the Engineering and Technical Fields (Nathans-Kelly, T. and Nicometo, C. G.) [Book Review]. IEEE Transactions on Professional Communication, 58(2):232–235, June 2015. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Shen:2022:UEV

[STP22] Zixing Shen, Songxin Tan, and Michael J. Pritchard. Understanding the effects of visual cueing on social media engagement with YouTube educational videos. IEEE Transactions on Professional Communication, 65(2):337–350, June 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Strem:1960:TMP

[Str60] Steve E. Strem. Technical motion pictures in science and engineering. IRE Transactions on Engineering Writing and Speech, EWS-3(2):36–42, July 1960. ISSN 0893-7869 (print), 2331-3692 (electronic).

Streeter:1964:STF

[Str64] James P. Streeter. Specifications: Their function, generation, and associated problems. IEEE Transactions on Engineering Writing and Speech, EWS-7(1):10–19, March 1964. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Strong:1984:EGB

[Str84] David Strong. Earl G. Bingham, Pocketbook for technical and professional writers. IEEE Transactions on Professional Communication, PC-27 (4):232, December 1984. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Stratton:1989:CWW

[Str89] C. R. Stratton. Collaborative writing in the workplace. IEEE Transactions on Professional Communication, 32(3): 178–182, September 1989. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Strosninder:1999:UTA

[Str99] M. Strosninder. User and task analysis for interface design. IEEE Transactions on Professional Communication, 42(3): 188–189, September 1999. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Strosnider:2001:VMC

[Str01] M. Strosnider. Video-mediated Communication [Book Review]. IEEE Transactions on Professional Communication, 44(1):60–61, March 2001. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Strother:2003:WBT

[Str03] J. B. Strother. Web-based training [Book Review]. IEEE Transactions on Professional Communication, 46(3):248– 250, September 2003. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Strother:2004:CCP

[Str04] J. B. Strother. Crisis communication put to the test: the case of two airlines on 9/11. IEEE Transactions on Professional Communication, 47(4): 290–300, December 2004. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Strantz:2021:UWS

[Str21] Adam Strantz. Using Web standards to design accessible data visualizations in professional communication. IEEE Transactions on Professional Communication, 64(3):288– 301, September 2021. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Stuckey:1969:AAP

[Stu69] R. D. Stuckey. Autoscript — an automated publications system. IEEE Transactions on Engineering Writing and Speech, EWS-12(2):29–33, May 1969. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Sturges:1992:OED

[Stu92] David L. Sturges. Overcoming the ethical dilemma: communication decisions in the ethic ecosystem. IEEE Transactions on Professional Communication, 35(1):44–50, March 1992. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Stypa:1986:PDP

[Sty86] John Stypa. Product documentation with a purpose. IEEE Transactions on Professional Communication, PC-29(3):29– 30, September 1986. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

StAmant:2009:EIE

[SU09] Kirk St. Amant and Jan M. Ulijn. Examining the information economy: Exploring the overlap between professional communication activities and information-management practices. IEEE Transactions on Professional Communication, 52(3):225–228, September 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Subramanyam:1976:CJC

[Sub76] K. Subramanyam. Core journals in computer science. IEEE Transactions on Professional Communication, PC-19(2):22– 25, June 1976. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Subramanyam:1979:LLL

[Sub79] K. Subramanyam. Lotka’s Law and the literature of computer science. IEEE Transactions on Professional Communication, PC-22(4):187–189, December 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Subbiah:1992:AND

[Sub92] M. Subbiah. Adding a new dimension to the teaching of audience analysis: cultural awareness. IEEE Transactions on Professional Communication, 35(1):14–18, March 1992. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sullivan:1989:BNC

[Sul89] P. Sullivan. Beyond a narrow conception of usability testing. IEEE Transactions on Professional Communication, 32(4): 256–264, December 1989. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Summers:1972:PCA

[Sum72] Clarence S. Summers. Problems in communications aggravated by the business recession: a PSYCHOM ’72 paper. IEEE Transactions on Professional Communication, PC-15 (2):60–64, June 1972. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Suojanen:2013:ICN

[Suo13]  Tytti Suojanen.  Intercul-

tural communication: a new approach to international relations and global challenges (Sadri, Houman A. and Flammia, Madelyn) [Book Review]. IEEE Transactions on Professional Communication, 56(1): 81–82, March 2013. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Surray:1996:PWC

[Sur96] A. M. Surray. Professional writers have a classroom role. IEEE Transactions on Professional Communication, 39(3): 163–165, September 1996. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Susskind:1961:PSP

[Sus61] Charles Susskind. Proposed system of periodic reports to government agencies. IRE Transactions on Engineering Writing and Speech, EWS-4(1): 19–20, January 1961. ISSN 0893-7869 (print), 2331-3692 (electronic).

Sussman:1985:CPT

[Sus85]  David Sussman.  Compos-

ing photographs for technical journals. IEEE Transactions on Professional Communication, PC-28(1):3–12, March 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sutherland:1977:ITS

[Sut77] William R. Sutherland. Impact of technology on scientific publishing. IEEE Transactions on Professional Communication, PC-20(2):56–58, September 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sivunen:2006:TLT

[SV06] A. Sivunen and M. Valo. Team leaders’ technology choice in virtual teams. IEEE Transactions on Professional Communication, 49(1):57–68, March 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sanz-Valdivieso:2024:TPM

[SV24] Lucía Sanz-Valdivieso. Technologypowered multilingual professional and technical writing: an integrative literature review of landmark and the latest writing assistance tools. IEEE Transactions on Professional Communication, 67(3): 301–315, September 2024. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Smith:2018:FCS

[SvI18] Jessica M. Smith and Tom van Ierland. Framing controversy on social media: #NoDAPL and the debate about the Dakota Access Pipeline on Twitter. IEEE Transactions on Professional Communication, 61(3):226–241, September 2018. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Swain:1980:GPW

[Swa80]  Philip W. Swain.  Giving

power to words. IEEE Transactions on Professional Communication, PC-23(3):135–137, September 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Swarts:2005:PMS

[Swa05] J. Swarts. PDAs in medical settings: the importance of organization in PDA text design. IEEE Transactions on Professional Communication, 48(2):

161–176, June 2005. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Swarts:2018:LDW

[Swa18] Jason Swarts. Locating and describing the work of technical communication in an online user network. IEEE Transactions on Professional Communication, 61(4):356– 371, December 2018. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Spyridakis:2005:IBR

[SWB^{+}05] J. H. Spyridakis, C. Wei, J. Barrick, E. C. Uddihy, and B. Maust. Internet-based research: providing a foundation for web-design guidelines. IEEE Transactions on Professional Communication, 48(3): 242–260, September 2005. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sun:2021:ESC

[SWR21] Ya Sun, Gongyuan Wang, and Hui Ren. To entertain or to serve: Chinese and US banks online identity based on a genre analysis of social media. IEEE Transactions on Professional Communication, 64(2): 121–136, June 2021. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Smith:2009:ACB

[SY09] Susann Smith and Anna Young. Adapting to change: Becoming a learning organization as a relief and development agency. IEEE Transactions on Professional Communication, 52(4):329–345, December 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sun:2022:COS

[SZLY22] Ya Sun, Limei Zhai, Wenbin Liu, and Kaiwen Yang. Corporations’ owned social media narrative. IEEE Transactions on Professional Communication, 65(2):280–293, June 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Szoka:1982:GCR

[Szo82] Kathryn Szoka. A guide to choosing the right chart type. IEEE Transactions on Professional Communication, PC-25 (2):98–101, June 1982. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Tallman:1977:INC

[Tal77] Johanna E. Tallman. Implications of the new copyright law for libraries and library users. IEEE Transactions on Professional Communication, PC-20 (3):178–184, November 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Tarnow:2001:RMV

[Tar01] E. Tarnow. A recipe for mission and vision statements. IEEE Transactions on Professional Communication, 44(2): 138–141, June 2001. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Taylor:1959:MT

[Tay59]  Paul Taylor.  Microphone

technique. IRE Transactions on Engineering Writing and Speech, EWS-2(1):30–32, January 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

Taylor:1973:SES

[Tay73] Robert L. Taylor. A study of the effects of security classification restrictions on technical communication. IEEE Transactions on Professional Communication, PC-16(4):194–199, December 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Taylor:1982:BSI

[Tay82] Richard A. Taylor. Brainstorming: In search of an idea. IEEE Transactions on Professional Communication, PC-25 (1):38–40, March 1982. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Taylor:2001:PRH

[Tay01]  T. N. Taylor.  Profile of

Richard Hill. IEEE Transactions on Professional Communication, 44(2):142–146, June 2001. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Taylor:2002:PTW

[Tay02] K. Taylor. The portable technical writer. IEEE Transactions on Professional Communication, 45(1):67–69, March 2002. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Thompson:1995:PRR

[TC95] L. H. Thompson and M. B.

Coney. Putting reader roles to the test: an ethnomethodological approach. IEEE Transactions on Professional Communication, 38(2):100–109, June 1995. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Teo:2006:PEF

[TCW06] Hock-Hai Teo, Hock Chuan Chan, and Kwok Kee Wei. Performance effects of formal modeling language differences: a combined abstraction level and construct complexity analysis. IEEE Transactions on Professional Communication, 49(2): 160–175, June 2006. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Tham:2018:UVR

[TDG^{+}18] Jason Tham, Ann Hill Duin, Laura Gee, Nathan Ernst, Bilal Abdelqader, and Megan McGrath. Understanding virtual reality: Presence, embodiment, and professional practice. IEEE Transactions on Professional Communication, 61(2):178–195, June 2018. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Thompson:1995:PMI

[TDS95] J. A. Thompson, K. B. DeTienne, and K. L. Smart. Privacy, e-mail, and information policy: where ethics meets reality. IEEE Transactions on Professional Communication, 38(3): 158–164, September 1995. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Tebeaux:1980:WMB

[Teb80]  Elizabeth Tebeaux.  What

makes bad technical writing bad? A historical analysis. IEEE Transactions on Professional Communication, PC-23 (2):71–76, June 1980. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Tell:1968:AAE

[Tel68] Bjorn V. Tell. ABACUS-AB atomic energy computerized user-oriented services: the mechanization of bibliographic list production. IEEE Transactions on Engineering Writing and Speech, EWS-11(2): 110–117, August 1968. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Ternent:1965:HHT

[Ter65] William A. Ternent. Here’s how to turn your engineers into authors. IEEE Transactions on Engineering Writing and Speech, EWS-8(1):27–29, June 1965. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Ternent:1972:WYN

[Ter72] William A. Ternent. Where are you, now that I need you: Or societal communication concepts: a PSYCHOM ’72 paper. IEEE Transactions on Professional Communication, PC-15 (3):100–104, September 1972. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Terrant:1973:PRJ

[Ter73] Seldon W. Terrant. Problems related to journal subject coverage, formats, and packaging. IEEE Transactions on Professional Communication, PC-16 (3):68, September 1973. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Terragno:1979:PTL

[Ter79] P. James Terragno. Patents as technical literature. IEEE Transactions on Professional Communication, PC-22(2): 101–104, June 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Teremetsky:1994:STT

[Ter94] V. E. Teremetsky. SOVAM TELEPORT: telecommunications in Russia and abroad. IEEE Transactions on Professional Communication, 37(2): 68–69, June 1994. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Thayer:2004:MCA

[Tha04] A. Thayer. Material culture analysis and technical communication: the artifact approach to evaluating documentation. IEEE Transactions on Professional Communication, 47(2): 144–147, June 2004. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Tham:2022:PFD

[Tha22] Jason Tham. Pasts and futures of design thinking: Implications for technical communication. IEEE Transactions on Professional Communication, 65(2):261–279, June 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Tham:2024:RVR

[Tha24] Jason Tham. Researching with virtual reality: Exploring the methodological affordances of VR for sociotechnical research and implications for technical and professional communication. IEEE Transactions on Professional Communication, 67(2):192–210, June 2024. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Theodos:1997:DCA

[The97] R. D. Theodos. Deborah C. Andrews, Ed., International Dimensions of Technical Communication. Arlington, Va: Soc. Tech. Commun., 1996, 135 pp., No Index. [Book Review]. IEEE Transactions on Professional Communication, 40(4):305–307, December 1997. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Thornton:1967:MCT

[Tho67]  Matthew Thornton.  Mak-

ing conventional troubleshooting charts uncharacteristically effective. IEEE Transactions on Engineering Writing and Speech, EWS-10(1):14–20, July 1967. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Thompson:1970:SUN

[Tho70] Charles W. N. Thompson.

Study of the user’s “Natural queue” of documents. IEEE Transactions on Engineering Writing and Speech, EWS-13(2):73–78, September 1970. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Thompson:1973:SAM

[Tho73] Charles W. N. Thompson. A systems approach to the management of a scientific journal. IEEE Transactions on Professional Communication, PC-16 (3):148–155, September 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Thomas:1975:EBS

[Tho75]  Garth J. Thomas.  Edit-

ing a basic science journal. IEEE Transactions on Professional Communication, PC-18 (3):186–189, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Thompson:2001:CTC

[Tho01] I. Thompson. Collaboration in technical communication: a qualitative content analysis of journal articles, 1990–1999. IEEE Transactions on Professional Communication, 44(3): 161–173, September 2001. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Tibbetts:1982:TRW

[Tib82]  Arn Tibbetts.  Ten rules

for writing readably. IEEE Transactions on Professional Communication, PC-25(1):10– 13, March 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Tilstra:1979:SIS

[Til79] C. F. Tilstra. Standards for information systems manuals. IEEE Transactions on Professional Communication, PC-22 (1):27–29, March 1979. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Tillery:2006:PNW

[Til06] Denise Tillery. The problem of nuclear waste: Ethos and scientific evidence in a high-stakes public controversy. IEEE Transactions on Professional Communication, 49(4): 325–334, December 2006. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Turetken:2011:EII

[TJQN11] Ozgur Turetken, Abhijit Jain, Brandi Quesenberry, and Ojelanki Ngwenyama. An empirical investigation of the impact of individual and work characteristics on telecommuting success. IEEE Transactions on Professional Communication, 54(1):56–67, March 2011. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Tong:2004:TAG

[TK04] A. Tong and E. Klecun. Toward accommodating gender differences in multimedia communication. IEEE Transactions on Professional Communication, 47(2):118–129, June 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Tebeaux:1992:IWT

[TL92] E. Tebeaux and M. Lay. Images of women in technical books from the English Renaissance. IEEE Transactions on Professional Communication, 35(4): 196–207, December 1992. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Timmerman:2008:POM

[TM08] C. Erik Timmerman and S. Naga Madhavapeddi. Perceptions of organizational media richness: Channel expansion effects for electronic and traditional media across richness dimensions. IEEE Transactions on Professional Communication, 51(1):18– 32, March 2008. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Tenbrink:2015:ECT

[TM15] Thora Tenbrink and Annika Maas. Efficiently connecting textual and visual information in operating instructions. IEEE Transactions on Professional Communication, 58(4): 346–366, December 2015. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Tompkins:1969:TSN

[Tom69] Howard E. Tompkins. Technical societies and the new information media. IEEE Transactions on Engineering Writing and Speech, EWS-12(2):58–61, May 1969. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Topf:1987:CDM

[Top87]  Mel A. Topf.  A course

in document management. IEEE Transactions on Professional Communication, PC-30 (4):254–257, December 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Tracey:1974:LE

[Tra74] Jim Tracey. Letters to the Editor. IEEE Transactions on Professional Communication, PC-17(2):41–42, June 1974. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Tracey:1983:TLS

[Tra83] James R. Tracey. The theory and lessons of stop discourse. IEEE Transactions on Professional Communication, PC-26 (2):68–78, June 1983. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Tracey:1992:SGS

[Tra92] J. R. Tracey. STOP, GO, and the state of the art in proposal writing. IEEE Transactions on Professional Communication, 35(3):143–155, September 1992. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Tran:2012:UCD

[Tra12]  Charity C. Tran.  User-

Centered Design for Personalization (van Velsen, L.; 2011) [Book Review]. IEEE Transactions on Professional Communication, 55(1):102–103, March 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Trice:2016:EMU

[Tri16] Michael Trice. Evaluating multilevel user skill expression in a public, unsupervised wiki: a case study. IEEE Transactions on Professional Communication, 59(3):261–273, September 2016. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Trowill:1966:GEM

[Tro66] John L. Trowill. The G-EWS Membership Survey. IEEE Transactions on Engineering Writing and Speech, EWS-9(2): 40–45, December 1966. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Trummel:1987:SSO

[Tru87] Paul Trummel. To see or not to see the other rhetoric.

IEEE Transactions on Professional Communication, PC-30 (1):30–31, March 1987. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Trummel:1988:RCI

[Tru88] P. Trummel. Rhetoric+typography:

creative interaction in modern communication. IEEE Transactions on Professional Communication, 31(3):124– 129, September 1988. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Trummel:1991:SCC

[Tru91] P. Trummel. Shape of concept: color percept … graphics, geometry, and gestalt. IEEE Transactions on Professional Communication, 34(3): 174–179, September 1991. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Tredennick:1981:SGS

[TS81] Nick Tredennick and Brion N.

Shimamoto. On systematic generation of scientific papers. IEEE Transactions on Professional Communication, PC-24 (3):124–127, September 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Teeni:2001:POC

[TSS^{+}01] Dov Te’eni, Abraham Sagie, David G. Schwartz, Nurit Zaidman, and Yair Amichai-Hamburger. The process of organizational communication: a model and field study. IEEE Transactions on Professional Communication, 44 (1):6–20, March 2001. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Tsui:1991:ECS

[Tsu91] Chia-Jung Tsui. English communication skills needs of professionals in Taiwan’s hightechnology industries. IEEE Transactions on Professional Communication, 34(2):79–82, June 1991. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Tukey:1962:KRC

[Tuk62] John W. Tukey. Keeping research in contact with the literature: Citation indexes and beyond. IRE Transactions on Engineering Writing and Speech, EWS-5(2):78–82, December 1962. ISSN 0893-7869 (print), 2331-3692 (electronic).

Turner:1974:C

[Tur74] Margaret S. Turner. Copyright. IEEE Transactions on Professional Communication, PC-17(3/4):43–47, September/ December 1974. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Turner:1979:ASP

[Tur79] Ruth H. Turner. 36 aids to successful proofreading. IEEE Transactions on Professional Communication, PC-22(1):19– 20, March 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Tusin:2005:AAO

[Tus05] J. Tusin. Approaches to audience: an overview of the major perspectives. IEEE Transactions on Professional Communication, 48(3):324– 326, September 2005. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Tusin:2006:GP

[Tus06] J. Tusin. Guide to presentations. IEEE Transactions on Professional Communication, 49(3):291–292, September 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Tran:2020:MHT

[TVRR20] Thi Tran, Rohit Valecha, Paul Rad, and H. Raghav Rao. Misinformation harms: a tale of two humanitarian crises. IEEE Transactions on Professional Communication, 63(4): 386–399, December 2020. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Tan:2000:DTE

[TWHN00] B. C. Y. Tan, Kwok-Kee Wei, W. W. Huang, and Guet-Ngoh Ng. A dialogue technique to enhance electronic communication in virtual teams. IEEE Transactions on Professional Communication, 43(2): 153–165, June 2000. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Truett:1991:IPP

[TYT91] L. F. Truett, T. G. Yow, and B. E. Tonn. Influencing public policy through information sharing. IEEE Transactions on Professional Communication, 34(4):228–232, December 1991. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ulijn:2001:TIG

[UC01] J. M. Ulijn and C. P. Campbell. Technical innovation and global business communication: an introduction. IEEE Transactions on Professional Communication, 44(2):77–82, June 2001. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ulijn:2001:NFF

[ULK01] J. M. Ulijn, A. Lincke, and Y. Karakaya. Non-face-to face international business negotiation: how is national culture reflected in this medium. IEEE Transactions on Professional Communication, 44(2):

126–137, June 2001. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ulrich:1984:WGT

[Ulr84] Gael D. Ulrich. Write a good technical report. IEEE Transactions on Professional Communication, PC-27(1):14– 19, March 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ummelen:2000:MRB

[UN00] N. Ummelen and R. Neutelings. Measuring reading behavior in policy documents: a comparison of two instruments. IEEE Transactions on Professional Communication, 43(3): 292–301, September 2000. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Underwood:1967:CRI

[Und67] William J. Underwood. A critique of reading improvement training. IEEE Transactions on Engineering Writing and Speech, EWS-10(2):52–56, December 1967. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Urbach:1977:AJA

[Urb77] Peter F. Urbach. Access to journal article copies through NTIS. IEEE Transactions on Professional Communication, PC-20(3):176–178, November 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ulijn:2002:ISI

[UVB02] J. M. Ulijn, D. R. Vogel, and T. M. A. Bemelmans. Introduction to the special issue. IEEE Transactions on Professional Communication, 45(4): 213–218, December 2002. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

VanVeen:1959:WBT

[Van59] Frederick T. Van Veen. Write better than you talk. IRE Transactions on Engineering Writing and Speech, EWS-2(1): 15, January 1959. ISSN 0893-7869 (print), 2331-3692 (electronic).

VanVeen:1962:GTL

[Van62] Frederick Van Veen. A guide to technical literature production [book review]. IRE Transactions on Engineering Writing and Speech, EWS-5(1):32, August 1962. ISSN 0893-7869 (print), 2331-3692 (electronic).

VanVeen:1963:IPT

[Van63] Frederick Van Veen. An index to 500 papers through 1962 on engineering writing and related subjects. IEEE Transactions on Engineering Writing and Speech, EWS-6(1):50–58, September 1963. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

VanHagan:1971:LST

[Van71] Charles E. Van Hagan. The long and short of technical writing courses. IEEE Transactions on Engineering Writing and Speech, EWS-14(3): 68–75, December 1971. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

VanCott:1975:DIJ

[Van75] Harold P. Van Cott. Derivative income from journal publishing. IEEE Transactions on Professional Communication, PC-18(3):103–107, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

VanDersal:1977:HWB

[Van77] William R. Van Dersal. How to write a book and get it published. IEEE Transactions on Professional Communication, PC-20(4):245–251, December 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Vanderbilt:1979:PII

[Van79] Byron M. Vanderbilt. The plight of the independent inventor. IEEE Transactions on Professional Communication, PC-22(2):63–67, June 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

VanDuyn:1983:TDP

[Van83] J. Van Duyn. The DP professional’s guide to writing effective technical communication. IEEE Transactions on Professional Communication, PC-26 (1):45, March 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Varley:1977:CPP

[Var77] H. Leslie Varley. Composition from publisher-prepared, machine-readable input. IEEE Transactions on Professional Communication, PC-20(2):65– 67, September 1977. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Vcherashnyi:1994:IIR

[Vch94] R. P. Vcherashnyi. Industrial information in Russia. IEEE Transactions on Professional Communication, 37(2):66–67, June 1994. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

vanderGeest:2013:ISS

[vdGRRv13] T. van der Geest, J. Ramey, S. Rosenbaum, and L. van Velsen. Introduction to the special section: Designing a better user experience for selfservice systems. IEEE Transactions on Professional Communication, 56(2):92–96, June 2013. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

vandenHaak:2006:CIA

[vdHdJS06] Maaike J. van den Haak, Menno D. T. de Jong, and Peter Jan Schellens. Constructive interaction: an analysis of verbal interaction in a usability setting. IEEE Transactions on Professional Communication, 49(4):311–324, December 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

vanderMeij:1996:DMH

[vdM96] H. van der Meij. Does the manual help? An examination of the problem-solving support offered by manuals. IEEE Transactions on Professional Communication, 39(3): 146–156, September 1996. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

vanderMeij:1997:BWE

[vdM97a] H. van der Meij. Brent Wilson, Ed., Constructivist Learning Environments: Case Studies In Instructional Design. Englewood Cliffs, Nj: Educational Technology Pub., 1996, 252 pp. With Index. [Book Review]. IEEE Transactions on Professional Communication, 40(4): 310–312, December 1997. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

vanderMeij:1997:IAU

[vdM97b] H. van der Meij. The ISTE approach to usability testing.

IEEE Transactions on Professional Communication, 40(3): 209–223, September 1997. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

vanderMeij:2007:GOG

[vdM07] Hans van der Meij. Goalorientation, goal-setting, and goal-driven behavior in minimalist user instructions. IEEE Transactions on Professional Communication, 50(4):295– 305, December 2007. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

vanderMeij:2004:FCP

[vdMG04] H. van der Meij and M. Gellevij. The four components of a procedure. IEEE Transactions on Professional Communication, 47(1):5–14, March 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

vanDeursen:2019:SNS

[vDVvL19] Alexander J. A. M. van Deursen, Carla Verlage, and Ester van Laar. Social network site skills for communication professionals: Conceptualization, operationalization, and an empirical investigation. IEEE Transactions on Professional Communication, 62 (1):43–54, March 2019. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Velte:1977:TTN

[Vel77] Charles Velte. The treatment of technical names. IEEE Transactions on Professional Communication, PC-20(1):23– 26, June 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Velotta:1987:SLW

[Vel87a] Christopher Velotta. Safety labels: What to put in them, how to write them, and where to place them. IEEE Transactions on Professional Communication, PC-30(3):121–126, September 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Velte:1987:DOE

[Vel87b] Charles E. Velte. Does online editing promote trespassing? IEEE Transactions on Professional Communication, PC-30 (3):179–181, September 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Velotta:1988:LBH

[Vel88] C. Velotta. Looking back:

how practical experience in the classroom taught me a professional approach on the job. IEEE Transactions on Professional Communication, 31(4): 160–162, December 1988. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Velte:1989:DUT

[Vel89] C. E. Velte. Document usability through objectives. IEEE Transactions on Professional Communication, 32(4):279– 282, December 1989. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Velte:1990:FIU

[Vel90] C. E. Velte. Face up to the interface (user interfaces). IEEE Transactions on Professional Communication, 33(3): 115–117, September 1990. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Venezky:1962:SSE

[Ven62] Richard L. Venezky. A smallscale experiment in the application of creativity to the teaching of technical communication. IRE Transactions on Engineering Writing and Speech, EWS-5(1):21–25, August 1962. ISSN 0893-7869 (print), 2331-3692 (electronic).

Vervalin:1980:CYF

[Ver80] Charles H. Vervalin. Checked your fog index lately? IEEE Transactions on Professional Communication, PC-23(2):87– 88, June 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Vealey:2016:RWC

[VG16] Kyle P. Vealey and Jeffrey M. Gerding. Rhetorical work in crowd-based entrepreneurship: Lessons learned from teaching crowdfunding as an emerging site of professional and technical communication. IEEE Transactions on Professional Communication, 59(4): 407–427, December 2016. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

vanHooijdonk:2008:IMP

[vHK08] Charlotte van Hooijdonk and Emiel Krahmer. Information modalities for procedural instructions: The influence of text, pictures, and film clips on learning and executing RSI exercises. IEEE Transactions on Professional Communication, 51(1):50–62, March 2008. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Vicente:2003:CWA

[Vic03] K. J. Vicente. Cognitive Work Analysis Toward Safe, Productive, and Healthy Computer-Based Work [Book Review]. IEEE Transactions on Professional Communication, 46(1): 63–65, March 2003. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Visscher:1977:COI

[Vis77] Maurice B. Visscher. Copyright and other impediments to scientific communication. IEEE Transactions on Professional Communication, PC-20 (3):149–151, November 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Vest:1996:EEP

[VLA96] D. Vest, M. Long, and T. Anderson. Electrical engineers’ perceptions of communication training and their recommendations for curricular change: results of a national survey. IEEE Transactions on Professional Communication, 39(1): 38–42, March 1996. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Vest:1995:RCT

[VLTP95] D. Vest, M. Long, L. Thomas, and M. E. Palmquist. Relating communication training to workplace requirements: the perspective of new engineers. IEEE Transactions on Professional Communication, 38(1): 11–17, March 1995. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

vanLuxemburg:2002:CEC

[vLUA02] A. P. D. van Luxemburg, J. M. Ulijn, and N. Amare. The contribution of electronic communication media to the design process: communicative and cultural implications. IEEE Transactions on Professional Communication, 45(4): 250–264, December 2002. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Verzella:2021:ETS

[VMM21] Massimo Verzella, Elisabet Arnó Macià, and Bruce Maylath. Engineers taking a stance on technical communication: Peer review of oral presentations via the Trans-Atlantic and Pacific Project. IEEE Transactions on Professional Communication, 64(1):66–83, March 2021. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

vanMeurs:2015:RUE

[vMPKG15] Frank van Meurs, Brigitte Planken, Hubert Korzilius, and Marinel Gerritsen. Reasons for using English or the local language in the genre of job advertisements: Insights from interviews with Dutch job ad designers. IEEE Transactions on Professional Communication, 58(1):86–105, March 2015. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Verdugo:2013:PUC

[VNC^{+}13] Renato Verdugo, Miguel Nussbaum, Magdalena Claro, Marcos Sepulveda, Boris Escobar, Ricardo Rendich, and Fernando Riveros. Preparing undergraduate computer science students to face intercultural and multidisciplinary scenarios. IEEE Transactions on Professional Communication, 56(1):67–80, March 2013. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Vogt:1983:YDA

[Vog83] Herbert E. Vogt. You don’t have to be an artist to produce an illustrated document. IEEE Transactions on Professional Communication, PC-26 (3):108–112, September 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Voland:1982:BGM

[Vol82] Gerard Voland. Bond graphing as a mode of technical communication. IEEE Transactions on Professional Communication, PC-25(1):35–37, March 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

VonEckardt:1978:WOM

[Von78]  Wolf Von Eckardt.  This

way out from our mad jumble of signs. IEEE Transactions on Professional Communication, PC-21(2):63–66, June 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

vanVelsen:2007:EUS

[vSdJ07] Lex S. van Velsen, Michaël F. Steehouder, and Menno D. T. de Jong. Evaluation of user support: Factors that affect user satisfaction with helpdesks and helplines. IEEE Transactions on Professional Communication, 50(3):219– 231, September 2007. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Varhelahti:2021:DCV

[VT21] Mervi Varhelahti and Tiia Turnquist. Diversity and communication in virtual project teams. IEEE Transactions on Professional Communication, 64(2):201–214, June 2021. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

VanHout:2016:LPE

[VV16] Tom Van Hout and Ellen Van Praet. Lookalike professional English. IEEE Transactions on Professional Communication, 59(4):398–406, December 2016. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Vogel:2001:ERR

[VVL^{+}01] D. R. Vogel, M. Van Genuchten, D. Lou, S. Verveen, M. Van Eekout, and A. Adams. Exploratory research on the role of national and professional cultures in a distributed learning project. IEEE Transactions on Professional Communication, 44(2):114–125, June 2001. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

vanWaes:2000:TAM

[vW00] L. van Waes. Thinking aloud as a method for testing the usability of Websites: the influence of task variation on the evaluation of hypertext. IEEE Transactions on Professional Communication, 43(3): 279–291, September 2000. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Waaijer:2009:HRO

[Waa09] Marlen K. Waaijer. Handbook of Research on Open Source Software: Technological, Economic, and Social Perspectives (St. Amant, K. and Still, B., Eds.; 2007) [Book Review]. IEEE Transactions on Professional Communication, 52(1): 109–112, March 2009. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Walterscheid:1964:WRC

[Wal64] Edward C. Walterscheid. Who is responsible for communication? IEEE Transactions on Engineering Writing and Speech, EWS-7(2):8–10, September 1964. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Waldhart:1974:USR

[Wal74] Thomas J. Waldhart. Utility of scientific research: The engineer’s use of the products of science. IEEE Transactions on Professional Communication, PC-17(2):33–35, June 1974. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Walter:1985:EVS

[Wal85] Gerard G. Walter. Estimating the vocabulary size of the disadvantaged reader. IEEE Transactions on Professional Communication, PC-28(4):21– 25, December 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Waldron:1986:IK

[Wal86] Vincent R. Waldron. Interviewing for knowledge. IEEE Transactions on Professional Communication, PC-29(2):31– 34, June 1986. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Walker:1987:ISO

[Wal87]  Janet H. Walker.  Issues

and strategies for online documentation. IEEE Transactions on Professional Communication, PC-30(4):235–248, December 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Wallace:1997:GIY

[Wal97] M. Wallace. Groupware: if you build it, they may not come. IEEE Transactions on Professional Communication, 40(1): 48–53, March 1997. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Walsh:1998:IOP

[Wal98] T. Walsh. Illustrations in oral presentations: photographs. IEEE Transactions on Professional Communication, 41(3): 209–212, September 1998. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Walker:1999:UGT

[Wal99] K. Walker. Using genre theory to teach students engineering lab report writing: a collaborative approach. IEEE Transactions on Professional Communication, 42(1):12–19, March 1999. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Walwema:2018:TIP

[Wal18] Josephine Walwema. Transliteracies in intercultural professional communication. IEEE Transactions on Professional Communication, 61(3):330– 345, September 2018. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Walwema:2021:DEI

[Wal21] Josephine Walwema. Diversity, equity, and inclusion in the transactions. IEEE Transactions on Professional Communication, 64(2):102–104, June 2021. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Walwema:2023:CSJ

[Wal23] Josephine Walwema. Centering social justice at ProComm Limerick 2022. IEEE Transactions on Professional Communication, 66(1):3–6, March 2023. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Wantman:1989:RCA

[Wan89] C. D. Wantman. Restructuring of corporate AV presentations: an analysis of current practice. IEEE Transactions on Professional Communication, 32(2):111–116, June 1989. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ward:1980:CYS

[War80] Jean Ward. Check out your sexism: a quiz for journalists. IEEE Transactions on Professional Communication, PC-23 (4):196–198, December 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Warlum:1988:IOM

[War88] M. F. Warlum. Improving oral marketing presentations in the technology-based company. IEEE Transactions on Professional Communication, 31(2): 84–87, June 1988. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Williams:2016:NPI

[WARAL16] Sean D. Williams, Gisela Ammetller, Inma Rodríguez-Ardura, and Xiaoli Li. A narrative perspective on international entrepreneurship: Comparing stories from the United States, Spain, and China. IEEE Transactions on Professional Communication, 59(4): 379–397, December 2016. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Williams:2020:NIW

[WARAL20] Sean D. Williams, Gisela Ammetller, Inma Rodríguez-Ardura, and Xiaoli Li. Narratives of international women entrepreneurs: an exploratory case study of identity negotiation in technology startups. IEEE Transactions on Professional Communication, 63 (1):39–51, March 2020. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Watts:1990:CEM

[Wat90] D. R. Watts. Creating an essential manual: an experiment in prototyping and task analysis. IEEE Transactions on Professional Communication, 33 (1):32–37, March 1990. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Watts:2020:UAS

[Wat20] Julie Watts. Using adapted studio critique to teach peer review in the document design classroom. IEEE Transactions on Professional Communication, 63(1):52–63, March 2020. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Walters:1992:DAS

[WB92] N. J. Walters and C. E.

Beck. A discourse analysis of software documentation: implications for the profession. IEEE Transactions on Professional Communication, 35(3): 156–167, September 1992. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Winckler:2013:IUE

[WBB13] Marco Winckler, Cedric Bach, and Regina Bernhaupt. Identifying user experience dimensions for mobile incident reporting in urban contexts. IEEE Transactions on Professional Communication, 56 (2):97–119, June 2013. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Wagner:1989:CEN

[WBW89] G. C. Wagner, W. Barfield, and T. R. Williams. A comparison of expert-novice text editing performance using dedicated and PC-based text editors.

IEEE Transactions on Professional Communication, 32(1): 49–55, March 1989. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

White:2015:CFS

[WC15] Ryman White and Ming Cheung. Communication of fantasy sports: a comparative study of user-generated content by professional and amateur writers. IEEE Transactions on Professional Communication, 58(2): 192–207, June 2015. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Walton:2009:VSD

[WD09] Rebecca Walton and Brian DeRenzi. Value-sensitive design and health care in Africa. IEEE Transactions on Professional Communication, 52(4): 346–358, December 2009. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Weaver:1985:TCC

[Wea85] Diane Gilman Weaver. Technical communications for computer users. IEEE Transactions on Professional Communication, PC-28(2):24–28, June 1985. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Weber:1965:APF

[Web65] Roger W. Weber. Associate processing of fragmentary information. IEEE Transactions on Engineering Writing and Speech, EWS-8(2):71–80, December 1965. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Weber:1975:LVM

[Web75] Hans H. Weber. The Librarian’s view of microforms. IEEE Transactions on Professional Communication, PC-18 (3):168–173, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Webb:1979:IPL

[Web79] Michael R. Webb. Improve project leadership with the aid of a communication infrastructure. IEEE Transactions on Professional Communication, PC-22(4):213–217, December 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Weedon:2019:RRE

[Wee19] Scott Weedon. The role of rhetoric in engineering judgment. IEEE Transactions on Professional Communication, 62(2):165–177, June 2019. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Wehner:1977:DJ

[Weh77] John M. Wehner. Death of a journal. IEEE Transactions on Professional Communication, PC-20(2):92–93, September 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Weil:1975:PIC

[Wei75] Ben H. Weil. Private information center aspects. IEEE Transactions on Professional Communication, PC-18(3): 222–223, September 1975. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Weil:1977:CAM

[Wei77] Ben H. Weil. Copying access mechanisms. IEEE Transactions on Professional Communication, PC-20(3):171–173, November 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Weiss:1989:SSS

[Wei89] L. A. Weiss. Selling science to society. IEEE Transactions on Professional Communication, 32(2):106–110, June 1989. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Weiss:1993:DDS

[Wei93] E. H. Weiss. Of document databases, SGML, and rhetorical neutrality. IEEE Transactions on Professional Communication, 36(2):58–61, June 1993. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Weiss:1995:PCO

[Wei95] E. H. Weiss. “Professional communication” and the “odor of mendacity”: the persistent suspicion that skilful writing is successful lying. IEEE Transactions on Professional Communication, 38(3):169– 175, September 1995. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Weiss:1998:TFW

[Wei98] E. H. Weiss. From Talmud folios to Web sites: HOT pages, COOL pages, and the information plenum. IEEE Transactions on Professional Communication, 41(2):97–106, June 1998. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Weiss:2001:PDE

[Wei01] E. H. Weiss. Preparing and delivering effective technical presentations: 2nd edition [Book Review]. IEEE Transactions on Professional Communication, 44(3):219–220, September 2001. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Weiss:2002:PWD

[Wei02a] E. H. Weiss. Principles of Web Design [Book Review]. IEEE Transactions on Professional Communication, 45(4): 319–320, December 2002. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Weiss:2002:PWC

[Wei02b]  E. H. Weiss.  Piet Wes-

tendorp, Carel Jansen, and Rob Punselie, Editors Interface Design & Document Design. IEEE Transactions on Professional Communication, 45(2): 138–139, June 2002. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Wells:1975:E

[Wel75] William F. Wells. Editorial. IEEE Transactions on Professional Communication, PC-18 (1):1, March 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Wells:1976:E

[Wel76] William F. Wells. Editorial. IEEE Transactions on Professional Communication, PC-19 (2):21, June 1976. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Wells:1982:HWB

[Wel82] Oliver C. Wells. How to write a book. IEEE Transactions on Professional Communication, PC-25(3):131–132, September 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Welhausen:2022:PDC

[Wel22] Candice A. Welhausen. Precarious data: Crack, opioids, and enacting a social justice ethic in data visualization practice. IEEE Transactions on Professional Communication, 65(1): 50–69, March 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Wentworth:1969:SCU

[Wen69] John W. Wentworth. Summary of the case for use of video tapes for technical training. IEEE Transactions on Engineering Writing and Speech, EWS-12(2):38–39, May 1969. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Wende:1985:DMS

[Wen85]  C. W. J. Wende.  Deca-

logue of the monthly summary. IEEE Transactions on Professional Communication, PC-28 (3):40, September 1985. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Werenfels:1971:IRP

[Wer71] Peter H. Werenfels. IEEE recommended practice: The international system of units. IEEE Transactions on Engineering Writing and Speech, EWS-14 (3):79, December 1971. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Westrum:1971:EIE

[Wes71] Ronald M. Westrum. The effect of improved electronic communications on business travel. IEEE Transactions on Engineering Writing and Speech, EWS-14(3):76–78, December 1971. CODEN IEE-WAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Wilson-Fetrow:2023:ESW

[WFSC^{+}23] Madalyn Wilson-Fetrow, Vanessa Svihla, Eva Chi, Catherine Hubka, and Yan Chen. Engineering students writing perceptions impact their conceptual learning. IEEE Transactions on Professional Communication, 66(2):186–201, June 2023. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Wright:1982:ITB

[WH82] William F. Wright and Donald T. Hawkins. Information technology a bibliography. IEEE Transactions on Professional Communication, PC-25 (1):43–50, March 1982. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Williams:1998:EVM

[WH98] T. R. Williams and D. A. Harkus. Editing visual media. IEEE Transactions on Professional Communication, 41(1): 33–46, March 1998. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Whalen:1983:DHB

[Wha83] Tim Whalen. Developing an in-house business and technical writing course. IEEE Transactions on Professional Communication, PC-26(4):160–161, December 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Whalen:1985:RWI

[Wha85] Tim Whalen. Renewal: Writing the incumbent proposal. IEEE Transactions on Professional Communication, PC-28 (2):13–16, June 1985. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Whalen:1986:PCM

[Wha86] Timothy Whalen. Proposal costs: Meeting the milestones on time, on budget. IEEE Transactions on Professional Communication, PC-29(3):25– 28, September 1986. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Whalen:2001:RFS

[Wha01] T. Whalen. Robert S. Frey, Successful proposal strategies for small businesses: winning government, private sector, and international contracts, 2nd ed. [Book Review]. IEEE Transactions on Professional Communication, 44(2): 147–148, June 2001. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Whalen:2002:TMC

[Wha02a] T. Whalen. Technical marketing communication [Book Review]. IEEE Transactions on Professional Communication, 45(4):317–318, December 2002. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Whalen:2002:WPR

[Wha02b]  T. Whalen.  Writing pro-

posals: rhetoric for managing change [Book Review]. IEEE Transactions on Professional Communication, 45(3): 201–202, September 2002. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Whalen:2004:AGG

[Wha04]  Tim Whalen.  Alan G.

Gross, Joseph E. Harmon, and Michael Reidym Communicating Science: The Scientific Article From the 17th Century to the Present. IEEE Transactions on Professional Communication, 47(3):224– 225, September 2004. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Wang:2012:RAP

[WHC^{+}12] Jingguo Wang, Tejaswini Herath, Rui Chen, Arun Vishwanath, and H. Raghav Rao.

Research article phishing susceptibility: an investigation into the processing of a targeted spear phishing email. IEEE Transactions on Professional Communication, 55(4): 345–362, December 2012. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Whittaker:1975:MAU

[Whi75] Della A. Whittaker. Modern American Usage by Wilson Follett. IEEE Transactions on Professional Communication, PC-18(4):325, December 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Whittaker:1976:BRa

[Whi76a] Della A. Whittaker. Book review. IEEE Transactions on Professional Communication, PC-19(1):20, March 1976. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Whittaker:1976:BRb

[Whi76b] Della A. Whittaker. Book review. IEEE Transactions on Professional Communication, PC-19(2):42, June 1976. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Whitby:1977:BSE

[Whi77a] Oliver W. Whitby. Background and status of the EPC-II experiment. IEEE Transactions on Professional Communication, PC-20(1):2–6, June 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Whittaker:1977:BR

[Whi77b] Della A. Whittaker. Book review. IEEE Transactions on Professional Communication, PC-20(2):137, September 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Whittaker:1977:HTW

[Whi77c] Della A. Whittaker. Handbook of technical writing, Charles T. Brusaw, Gerald J. Alred, and Walter E. Oliu. IEEE Transactions on Professional Communication, PC-20(4): 255, December 1977. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Whittaker:1977:HPP

[Whi77d] Della A. Whittaker. How to prepare professional design brochures, Gerre Jones. IEEE Transactions on Professional Communication, PC-20 (4):255–256, December 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Whittaker:1978:CBG

[Whi78a] Della A. Whittaker. Conceptual blockbusting: a guide to better ideas, James L. Adams.

IEEE Transactions on Professional Communication, PC-21 (3):128–129, September 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Whittaker:1978:PTT

[Whi78b] Della A. Whittaker. Photo technology, third edition Marshall Lacour and Irvin T. Lathrop. IEEE Transactions on Professional Communication, PC-21(3):128, September 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Whittaker:1978:SMW

[Whi78c] Della A. Whittaker. Scientists must write, Robert Barrass. IEEE Transactions on Professional Communication, PC-21 (4):170, December 1978. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Whittaker:1978:TIG

[Whi78d] Della A. Whittaker. Technical illustrating, George E. Morris. IEEE Transactions on Professional Communication, PC-21(1):39, March 1978. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Whittaker:1978:WPR

[Whi78e] Della A. Whittaker. Writing professional reports: a guide for students, Russell Rutter and Karl E. Gwiasda. IEEE Transactions on Professional Communication, PC-21(2):86, June 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

White:1979:MCE

[Whi79a] George R. White. Management criteria for effective innovation. IEEE Transactions on Professional Communication, PC-22 (2):77–82, June 1979. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

White:1979:PP

[Whi79b] Joseph F. White. Presenting papers. IEEE Transactions on Professional Communication, PC-22(4):179–182, December 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Whittaker:1979:BBP

[Whi79c] Della A. Whittaker. Basic business and professional speech communication, Ted Frank and David Ray. IEEE Transactions on Professional Communication, PC-22(3):173–174, September 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Whittaker:1979:HWW

[Whi79d] Della A. Whittaker. How to write for the world of work, Thomas E. Pearsall and Donald H. Cunningham. IEEE Transactions on Professional Communication, PC-22(1):39– 40, March 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Whittaker:1979:RTB

[Whi79e] Della A. Whittaker. Reading technical books, Anne Eisenberg. IEEE Transactions on Professional Communication, PC-22(3):175, September 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Whittaker:1979:TIT

[Whi79f] Della A. Whittaker. Technical illustration, third edition T. A. Thomas. IEEE Transactions on Professional Communication, PC-22(3):175–176, September 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Whittaker:1979:TRS

[Whi79g] Della A. Whittaker. Technical report standards: How to prepare and write effective technical reports, Lawrence R. Harvill and Thomas L. Kraft. IEEE Transactions on Professional Communication, PC-22 (1):40, March 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Whittaker:1979:TRW

[Whi79h] Della A. Whittaker. Technical report writing, James W. Souther and Myron L. White. IEEE Transactions on Professional Communication, PC-22 (1):40–41, March 1979. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Whittaker:1979:TTH

[Whi79i] Della A. Whittaker. Technology transfer: How to make it work, Hyman Olken. IEEE Transactions on Professional Communication, PC-22(3): 176, September 1979. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Whittaker:1980:CMP

[Whi80a] Della A. Whittaker. Communication for managers, Paul Preston. IEEE Transactions on Professional Communication, PC-23(2):108–109, June 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Whittaker:1980:PPE

[Whi80b] Della A. Whittaker. Practically painless English, Sally F. Wallace. IEEE Transactions on Professional Communication, PC-23(3):154, September 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Whittaker:1980:WYR

[Whi80c] Della A. Whittaker. Write for your reader. IEEE Transactions on Professional Communication, PC-23(4):170–173, December 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Whittaker:1982:GDD

[Whi82a] Della A. Whittaker. Guidelines for document designers. IEEE Transactions on Professional Communication, PC-25 (3):162, September 1982. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Whittaker:1982:L

[Whi82b] Della A. Whittaker. On language. IEEE Transactions on Professional Communication, PC-25(3):164, September 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Whitney:1987:CER

[Whi87] Margaret A. Whitney. Combining elegance and readability: Walker Gibson’s tough, sweet, and stuffy. IEEE Transactions on Professional Communication, PC-30(4):222–226, December 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Whitney:1989:ACC

[Whi89] M. A. Whitney. Analyzing corporate communications policy using ethnographic methods. IEEE Transactions on Professional Communication, 32(2): 76–80, June 1989. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Whitney:1991:TMD

[Whi91] M. A. Whitney. Time to make the donuts: the role of video in corporate technical training. IEEE Transactions on Professional Communication, 34(1): 38–41, March 1991. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Whiteside:2006:WSW

[Whi06] A. L. Whiteside. Writing selves/writing societies: Research from activity perspectives. IEEE Transactions on Professional Communication, 49(1):89–90, March 2006. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Whitfield:2014:NCE

[Whi14]  Melinda Whitfield.  Nego-

tiating cultural encounters. IEEE Transactions on Professional Communication, 57(3): 238–239, September 2014. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Wilson:1964:IRI

[Wil64] Lloyd B. Wilson. Information retrieval: Implementation activities for the IEEE. IEEE Transactions on Engineering Writing and Speech, EWS-7(3): 37–43, December 1964. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Williams:1975:LSD

[Wil75]  Gordon R. Williams.  Li-

brary subscription decisions. IEEE Transactions on Professional Communication, PC-18 (3):207–209, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Williams:1977:QRP

[Wil77a]  Joanna Williams.  Qual-

ity in the review process. IEEE Transactions on Professional Communication, PC-20 (2):131–132, September 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Williamson:1977:JGD

[Wil77b] Harry Williamson. Journal growth by division. IEEE Transactions on Professional Communication, PC-20(2):89– 92, September 1977. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Williams:1980:ICM

[Wil80] Donald E. Williams. Idea clarification: a matter of predesign. IEEE Transactions on Professional Communication, PC-23 (1):32–34, March 1980. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Williams:1982:IC

[Wil82] Donald E. Williams. Idea clarification. IEEE Transactions on Professional Communication, PC-25(4):172–174, December 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Wilcox:1984:HWR

[Wil84] Alan D. Wilcox. How to write a recommendation. IEEE Transactions on Professional Communication, PC-27(4):211–214, December 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Will:1989:AAC

[Wil89] J. W. Will. Age of the amateur communicator. IEEE Transactions on Professional Communication, 32(2):117–119, June 1989. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Williams:1997:IPC

[Wil97] C. Williams. Intel’s Pentium chip crisis: an ethical analysis. IEEE Transactions on Professional Communication, 40 (1):13–19, March 1997. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Williamson:1998:CPP

[Wil98] W. J. Williamson. The Computer and The Page: Publishing, Technology, & The Classroom [Book Review]. IEEE Transactions on Professional Communication, 41 (1):84–85, March 1998. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Willerton:2003:BMR

[Wil03] Russell Willerton. Barbara Mirel and Rachel Spilka, Reshaping Technical Communication: New Directions and Challenges for the 21st Century [Book Review]. IEEE Transactions on Professional Communication, 46(1):70–72, March 2003. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Williams:2010:EWS

[Wil10a] Julia M. Williams. Evaluating what students know: Using the RosE Portfolio System for institutional and program outcomes assessment tutorial. IEEE Transactions on Professional Communication, 53(1): 46–57, March 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Williams:2010:BCR

[Wil10b] Miriam F. Williams. From black codes to recodification: Removing the veil from regulatory writing book review. IEEE Transactions on Professional Communication, 53(4): 426–427, December 2010. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Willers:2024:BAD

[Wil24] Heidi Willers. Bridging the accessibility divide: Testing the efficacy of an accessible user experience model via a case study of Microsoft’s Inclusive Design Toolkit. IEEE Transactions on Professional Communication, 67(1):121–132, March 2024. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Winton:1973:JM

[Win73] R. C. Winton. Just a minute. IEEE Transactions on Professional Communication, PC-16 (4):206–208, December 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Winton:1983:PSA

[Win83] Robert C. Winton. Problems in speaking abroad. IEEE Transactions on Professional Communication, PC-26(4):194–196, December 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Winsor:1988:CFC

[Win88] D. A. Winsor. Communication failures contributing to the Challenger accident: an example for technical communicators. IEEE Transactions on Professional Communication, 31(3):101–107, September 1988. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Winn:1989:RGT

[Win89] W. Winn. The role of graphics in training documents: toward an explanatory theory of how they communicate. IEEE Transactions on Professional Communication, 32(4): 300–309, December 1989. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Winn:1990:ERI

[Win90]  W. Winn.  Encoding and

retrieval of information in maps and diagrams. IEEE Transactions on Professional Communication, 33(3):103– 107, September 1990. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Winn:1991:CDD

[Win91] W. Winn. Color in document design. IEEE Transactions on Professional Communication, 34(3):180–185, September 1991. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Winsor:1997:ELW

[Win97] D. A. Winsor. Electronic Literacies in the Workplace: Technologies of Writing [Book Reviews]. IEEE Transactions on Professional Communication, 40(1):54–55, March 1997. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Wing:1999:CDS

[Win99] L. Wing. Creating a doc spec.

IEEE Transactions on Professional Communication, 42(2): 128–131, June 1999. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Wiseman:1991:HAT

[Wis91] T. L. Wiseman. How to avoid the transitional ax in indirect bad news messages. IEEE Transactions on Professional Communication, 34(1): 20–23, March 1991. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Wisne:2015:SOI

[Wis15] Karen Wisne. Sharing Our Intellectual Traces (Bridgeford, T., Kitalong, K. S., and Williamson, B., Editors) [Book Review]. IEEE Transactions on Professional Communication, 58(2):230–231, June 2015. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Wittman:1977:BJN

[Wit77] Allan Wittman. Birth of a journal: The need. IEEE Transactions on Professional Communication, PC-20(2):79– 81, September 1977. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Wirtz:2013:IUE

[WJ13] Simone Wirtz and Eva-Maria Jakobs. Improving user experience for passenger information systems. Prototypes and reference objects. IEEE Transactions on Professional Communication, 56(2):120–137, June 2013. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Walker:1968:CMT

[WKSH68] Gordon L. Walker, Susumu Kuno, Barbara N. Smith, and Roland B. Holt. Chinese mathematical text analysis. IEEE Transactions on Engineering Writing and Speech, EWS-11 (2):118–128, August 1968. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Weger:2022:HED

[WLM^{+}22] Kristin Weger, Sheri Leder, Bryan Mesmer, Vineetha Menon, and Harald Schaub. How effectively do we communicate? An analysis of team reflexivity in transition and action phases of team collaboration. IEEE Transactions on Professional Communication, 65(3):392–410, September 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Waller:1982:DYN

[WLMR82] Robert Waller, Paul Lefrere, and Michael Macdonald-Ross. Do you need that second color? IEEE Transactions on Professional Communication, PC-25 (2):80–85, June 1982. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Walter:1987:LDD

[WM87] Charles Walter and Thomas F. Marsteller. Liability for the dissemination of defective information. IEEE Transactions on Professional Communication, PC-30(3):164–167, September 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Wareham:2007:CMU

[WMPBA07] Jonathan Wareham, Volker Mahnke, Sanjay Peters, and Niels Bjorn-Andersen. Communication metaphors-in-use: Technical communication and offshore systems development. IEEE Transactions on Professional Communication, 50 (2):93–108, June 2007. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Wong:2016:WCP

[WOD^{+}16] Louie H. M. Wong, Carol X. J. Ou, Robert M. Davison, Hui Zhu, and Cheng Zhang. Web 2.0 and communication processes at work: Evidence from China. IEEE Transactions on Professional Communication, 59(3):230–244, September 2016. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Woelfle:1975:REP

[Woe75] Robert M. Woelfle. The role of the engineer in preparing proposals. IEEE Transactions on Professional Communication, PC-18(1):31–34, March 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Woelfle:1979:REI

[Woe79] Robert M. Woelfle. Role of the engineer in improving the communication of technical information. IEEE Transactions on Professional Communication, PC-22(1):24–26, March 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Woelfle:1980:IWP

[Woe80] Robert M. Woelfle. The impact of word processing on engineering communications. IEEE Transactions on Professional Communication, PC-23 (4):159–163, December 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Wofford:2009:VDM

[Wof09] Tracey Norden Wofford. Visual Design for the Modern Web (McIntire, P.; 2008) [Book Review]. IEEE Transactions on Professional Communication, 52(3):316–317, September 2009. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Wolber:1979:BVP

[Wol79] William G. Wolber. The business value of patents. IEEE Transactions on Professional Communication, PC-22(2):73– 76, June 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Wolff:1982:WYG

[Wol82] Michael F. Wolff. When you have to get up and speak. IEEE Transactions on Professional Communication, PC-25 (4):202–203, December 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Wolf:1988:CVG

[Wol88] J. Wolf. Copyediting versus grading — a student perspective. IEEE Transactions on Professional Communication, 31(4):169–170, December 1988. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Wolfe:2006:MMR

[Wol06] J. Wolfe. Meeting minutes as a rhetorical genre: Discrepancies between professional writing textbooks and workplace practice tutorial. IEEE Transactions on Professional Communication, 49(4):354– 364, December 2006. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Woolsey:1979:EMW

[Woo79] Gene Woolsey. An essay on misused words: Sophisticated and elegant. IEEE Transactions on Professional Communication, PC-22(3):170–171, September 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Woolsey:1980:CSC

[Woo80]  Gene Woolsey.  Corporate

style, corporate substance, and the sting. IEEE Transactions on Professional Communication, PC-23(2):67–68, June 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Woolsey:1981:LWB

[Woo81] Gene Woolsey. On learning where to begin, how to listen, and when to stop. IEEE Transactions on Professional Communication, PC-24(4):197–198, December 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Woolever:1990:CLL

[Woo90] K. R. Woolever. Corporate language and the law: avoiding liability in corporate communications. IEEE Transactions on Professional Communication, 33(2):94–98, June 1990. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Worsley:1987:UDP

[Wor87] Michael D. Worsley. The UK Data Protection Act 1984 and international writers. IEEE Transactions on Professional Communication, PC-30(3): 208–211, September 1987. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Weiss:1982:WSS

[WP82] Edmond H. Weiss and Kenneth B. Powell. The writing system for scientists and engineers. IEEE Transactions on Professional Communication, PC-25(1):51, March 1982. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Weissbach:2018:CWC

[WP18] Robert S. Weissbach and Ruth C. Pflueger. Collaborating with writing centers on interdisciplinary peer tutor training to improve writing support for engineering students. IEEE Transactions on Professional Communication, 61 (2):206–220, June 2018. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Weimann:2013:ETP

[WPSB13] Peter Weimann, Michael Pollock, Elsje Scott, and Irwin Brown. Enhancing team performance through tool use: How critical technology-related issues influence the performance of virtual project teams. IEEE Transactions on Professional Communication, 56(4): 332–353, December 2013. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Williams:1983:STL

[WR83] Joseph M. Williams and William Rivers. Style: Ten lessons in clarity and grace. IEEE Transactions on Professional Communication, PC-26 (2):88–89, June 1983. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Wright:1979:PPU

[Wri79] Carl M. Wright. Publication, public use, and sale as bars to patenting. IEEE Transactions on Professional Communication, PC-22(2):89–92, June 1979. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Wright:1981:FST

[Wri81] Patricia Wright. Five skills technical writers need. IEEE Transactions on Professional Communication, PC-24(1):10– 16, March 1981. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Wuebben:2023:VRC

[WRTBRL23] Daniel Wuebben, Jose Luis Rubio-Tamayo, Manuel Gertrudix Barrio, and Juan Romero-Luis.

360^{∘} video for research communication and dissemination: a case study and guidelines. IEEE Transactions on Professional Communication, 66 (1):59–77, March 2023. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Wenger:1989:RRV

[WS89] M. J. Wenger and J. H. Spyridakis. The relevance of reliability and validity to usability testing. IEEE Transactions on Professional Communication, 32(4):265–271, December 1989. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Williams:1992:VDH

[WS92] T. R. Williams and J. H. Spyridakis. Visual discriminability of headings in text. IEEE Transactions on Professional Communication, 35(2):64–70, June 1992. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Williams:2004:ISI

[WS04] J. M. Williams and J. Strother. Introduction to the Special Issue on New Case Studies for Technical and Professional Communication Courses. IEEE Transactions on Professional Communication, 47(4):229–232, December 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Wu:2023:DCM

[WS23] Chenghui Wu and Ya Sun. Discursive construction of message credibility for Chinese stateowned enterprises on Twitter. IEEE Transactions on Professional Communication, 66(3): 236–258, September 2023. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Worley:2009:ITE

[WT09] Wanda L. Worley and Lee S.

Tesdell. Instructor time and effort in online and face-to-face teaching: Lessons learned. IEEE Transactions on Professional Communication, 52(2): 138–151, June 2009. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Wu:1988:PW

[Wu88]  V. W. Wu.  The prob-

lems that I have with writing. IEEE Transactions on Professional Communication, 31(4): 181–183, December 1988. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Warren:2006:CCC

[WX06] Thomas L. Warren and Kejun Xu. Cross-cultural communication: Perspectives in theory and practice. IEEE Transactions on Professional Communication, 49(4):370– 372, December 2006. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Xu:2024:MDW

[XS24] Wenjuan Xu and Xingsong Shi. Measuring the degree of website adaptation and its influencing factors: an empirical investigation of Chinese MNCs’ overseas websites. IEEE Transactions on Professional Communication, 67(1): 89–104, March 2024. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Xu:2007:TCT

[Xu07] Kejun Xu. Technical Communication (Tenth Edition) (Lannon, J. M.; 2006). IEEE Transactions on Professional Communication, 50(2):172– 174, June 2007. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Xu:2008:BME

[Xu08] Kejun Xu. Building and Managing Enterprise-Wide Portals (J. Polgar et al.; 2005) [Book Review]. IEEE Transactions on Professional Communication, 51(1):112–113, March 2008. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Xu:2010:ISI

[XZZ10] Yunjie Xu, Cheng Zhang, and Chenghong Zhang. Information seeking in an information systems project team. IEEE Transactions on Professional Communication, 53(4): 370–381, December 2010. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Yu:2017:GSC

[YB17] Danni Yu and Marina Bondi.

The generic structure of CSR reports in Italian, Chinese, and English: a corpus-based analysis. IEEE Transactions on Professional Communication, 60 (3):273–291, September 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Yeats:2009:OTC

[Yea09]  Dave Yeats.  Outsourcing

Technical Communication: Issues, Policies, and Practices (Thatcher, B. and Evia, C., Eds.; 2008) [Book Review]. IEEE Transactions on Professional Communication, 52(2): 213–214, June 2009. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Yeats:2010:NTK

[Yea10] Dave Yeats. Network: Theorizing Knowledge Work in Telecommunications (Spinuzzi, C.) [Book Review]. IEEE Transactions on Professional Communication, 53(3):320– 321, September 2010. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Yee:1988:TEP

[Yee88] C. Yee. Technical and ethical professional preparation for technical communication students. IEEE Transactions on Professional Communication, 31(4):191–194, December 1988. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Yee:1990:BMD

[Yee90]  C. Yee.  The brain/mind

and document design. IEEE Transactions on Professional Communication, 33(4):168– 171, December 1990. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Yerkey:1974:CER

[Yer74] A. Neil Yerkey. Common elements and requirements of information storage and retrieval systems. IEEE Transactions on Professional Communication, PC-17(1):23–28, March 1974. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Yin:2015:SMM

[YFW15] Juelin Yin, Jieyun Feng, and Yuyan Wang. Social media and multinational corporations corporate social responsibility in China: The case of ConocoPhillips Oil Spill Incident.

IEEE Transactions on Professional Communication, 58(2): 135–153, June 2015. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Yli-Jokipii:2001:LGE

[YJ01] Hilkka Yli-Jokipii. The local and the global: an exploration into the Finnish and English websites of a Finnish company. IEEE Transactions on Professional Communication, 44(2): 104–113, June 2001. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Yuan:2024:IDV

[YJY24] Wen Yuan, Yue Jin, and Yingli Yang. The influence of disciplinary variation and speaker characteristics on the use of hedges and boosters in Zhihu live talks. IEEE Transactions on Professional Communication, 67(1):105–120, March 2024. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Yin:2013:SHI

[YK13] Chun-Po Yin and Feng-Yang Kuo. A study of how information system professionals comprehend indirect and direct speech acts in project communication. IEEE Transactions on Professional Communication, 56(3):226–241, September 2013. CODEN IEPCBU.

ISSN 0361-1434 (print), 1558-1500 (electronic).

Yorks:1978:MPR

[Yor78] Lyle Yorks. Managing professional relationships: Communication skills. IEEE Transactions on Professional Communication, PC-21(2):56–60, June 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Young:1980:BAW

[You80]  Richard I. Young.  Build

an article writing program. IEEE Transactions on Professional Communication, PC-23 (3):130–132, September 1980. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Youngblood:2010:OCC

[You10] Susan Youngblood. Ongoing Crisis Communication: Planning, Managing, and Responding, 2nd Edition (Coombs, W. T.) and Handbook of Risk and Crisis Communication (Heath, R. L. and O’Hair, H. D., Eds.) [Book reviews]. IEEE Transactions on Professional Communication, 53(2):174–178, June 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Youngblood:2011:OBP

[You11] Norman E. Youngblood. Opposing broad patent: Scientific American’s response to Morse v. O’Reilly (1848).

IEEE Transactions on Professional Communication, 54 (1):31–42, March 2011. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Yoshizawa:2012:AIS

[YTY12] Sayuri Yoshizawa, Takao Terano, and Atsushi Yoshikawa. Assessing the impact of student peer review in writing instruction by using the normalized compression distance. IEEE Transactions on Professional Communication, 55(1): 85–96, March 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Yu:2010:DTC

[Yu10] Han Yu. Developing technical communication education for Chinese industry professionals: Preliminary findings and suggestions. IEEE Transactions on Professional Communication, 53(2):102–115, June 2010. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Yu:2011:IIC

[Yu11] Han Yu. Integrating intercultural communication into an engineering communication service class tutorial. IEEE Transactions on Professional Communication, 54 (1):83–96, March 2011. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Yu:2012:SES

[Yu12] Han Yu. A study of engineering students’ intercultural competence and its implications for teaching. IEEE Transactions on Professional Communication, 55(2):185–201, June 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Yu:2020:CEC

[Yu20] Danni Yu. Corporate environmentalism: a critical metaphor analysis of Chinese, American, and Italian corporate social responsibility reports. IEEE Transactions on Professional Communication, 63(3): 244–258, September 2020. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Yu:2023:CCG

[Yu23] Danni Yu. A cross-cultural genre analysis of leadership statements in Italian and American university sustainability reports. IEEE Transactions on Professional Communication, 66(1):26–47, March 2023. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Yu:2024:ILI

[Yu24] Minjing Yu. Introducing language and intercultural communication, 3rd edition: Jane Jackson: [book review]. IEEE Transactions on Professional Communication, 67(1):154– 155, March 2024. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Yuan:2021:CIC

[Yua21] Sheng Yuan. Comparing international communication of corporate social responsibility by Chinese and Korean firms on social media. IEEE Transactions on Professional Communication, 64(2):154–169, June 2021. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Zapiti:1980:MGP

[Zap80] Kallie Zapiti. Model guidelines for the preparation of cameraready typescripts by authors/ typists, M. O’Connor, Ed. IEEE Transactions on Professional Communication, PC-23 (4):204, December 1980. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Zaneski:1984:SMP

[ZB84] Richard Zaneski and R. John Brockmann. Software manual production simplified. IEEE Transactions on Professional Communication, PC-27(2): 109, June 1984. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Zollet:2015:CFI

[ZB15] Roman Zollet and Andrea Back. Critical factors influencing diffusion of interactivity innovations on corporate Websites. IEEE Transactions on Professional Communication, 58(1):2–19, March 2015. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Zemliansky:2017:WAC

[ZB17] Pavel Zemliansky and Landon Berry. A writing-across-the-curriculum faculty development program: an experience report. IEEE Transactions on Professional Communication, 60(3):306–316, September 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Zhang:2024:MAL

[ZC24] Huiyu Zhang and Feiyu Chen. A multidimensional analysis of leaders’ messages in Chinese and American corporate social responsibility reports. IEEE Transactions on Professional Communication, 67(1): 63–88, March 2024. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Zemliansky:2012:AEC

[Zem12] Pavel Zemliansky. Achieving experiential cross-cultural training through a virtual teams project. IEEE Transactions on Professional Communication, 55(3):275–286, September 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Zemliansky:2021:FDT

[Zem21] Pavel Zemliansky. Faculty development training in online instruction at a Norwegian university: an experience report. IEEE Transactions on Professional Communication, 64(1): 84–95, March 2021. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Zettler:2019:UDM

[Zet19]  Jessi Zettler.  UX Design

for Mobile: Pablo Perea and Pau Giner [Book Review]. IEEE Transactions on Professional Communication, 62(1): 112–113, March 2019. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Zhang:2021:CIM

[ZFWX21] Huiyu Zhang, Jieyun Feng, Yuanhong Wei, and Jiayu Xia. The construction of interpersonal meanings in the iPhone 1 product launch presentation: Integrating verbal and visual semiotics. IEEE Transactions on Professional Communication, 64(1):3–22, March 2021. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Zhang:2021:ESC

[Zha21] Yan Zhang. Effective Scientific Communication: The Other Half of Science: Cristina Hanganu-Bresch and Kelleen Flaherty: [Book Review].

IEEE Transactions on Professional Communication, 64 (1):96–97, March 2021. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Zhou:2005:EID

[Zho05] L. Zhou. An empirical investigation of deception behavior in instant messaging. IEEE Transactions on Professional Communication, 48(2): 147–160, June 2005. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Zhou:2008:PTP

[Zho08] Quan Zhou. Portfolios for Technical and Professional Communicators (Smith, H. J. and Haimes-Korn, K.; 2007) [Book Review]. IEEE Transactions on Professional Communication, 51(2):228–229, June 2008. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Ziegler:1969:TUP

[Zie69] J. C. Ziegler. Text/360 from a user’s point of view. IEEE Transactions on Engineering Writing and Speech, EWS-12 (2):33–35, May 1969. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Zimmerman:1983:URT

[Zim83] Muriel Zimmerman. The use of repetition in technical communication. IEEE Transactions on Professional Communication, PC-26(1):9–10, March 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Zimmerman:1988:CVG

[Zim88] D. E. Zimmerman. Copyediting versus grading — an alternative approach for critiquing students’ work. IEEE Transactions on Professional Communication, 31(4):163– 168, December 1988. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Zimmerman:1993:MFN

[Zim93] B. Zimmerman. Metaphor, frame, and nonverbal communication: an ethnographic study of a technical writing classroom. IEEE Transactions on Professional Communication, 36(3):137–143, September 1993. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Zimmerman:1999:FKM

[Zim99] Muriel Zimmerman. Fundable knowledge: the marketing of defense technology, by A. D. Van Nostrand. IEEE Transactions on Professional Communication, 42(2):143–144, June 1999. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Zimmerman:2000:WWO

[Zim00] M. Zimmerman. Weaving the web: the original design and ultimate destiny of the World Wide Web by its inventor [Book Review]. IEEE Transactions on Professional Communication, 43(2):217–218, June 2000. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Zimmerman:2001:WMU

[Zim01]  M. Zimmerman.  Writing

in a milieu of utility: the move to technical communication in American engineering programs, 1850–1950, 2nd ed. [Book Review]. IEEE Transactions on Professional Communication, 44(4):298– 299, December 2001. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Zimmerman:2013:DRT

[Zim13]  Muriel Zimmerman.  De-

sign Research Through Practice: From the Lab, Field, and Showroom, Ilpo Koskinen and John Zimmerman and Thomas Binder and Johan Redstrom and Stephan Wensveen. IEEE Transactions on Professional Communication, 56(3):262–263, September 2013. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Zhou:2017:HCB

[ZJG17] Shixiang Zhou, Heejin Jeong, and Paul A. Green. How consistent are the best-known readability equations in estimating the readability of design standards? IEEE Transactions on Professional Communication, 60(1):97– 111, March 2017. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Zhang:2017:IEH

[ZLCZ17] Chenghong Zhang, Tian Lu, Shoucong Chen, and Cheng Zhang. Integrating ego, homophily, and structural factors to measure user influence in online community. IEEE Transactions on Professional Communication, 60(3): 292–305, September 2017. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Zhou:2023:WCT

[ZM23] Quan Zhou and Zarah C.

Moeggenberg. What can technical and professional communication do for UX education: a case study of a userexperience graduate certificate. IEEE Transactions on Professional Communication, 66(4): 366–381, December 2023. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Zani:2022:MFS

[ZNG22] Azma Alina Ali Zani, Azah Anir Norman, and Norjihan Abdul Ghani. Motivating factors to self-disclosure on social media: a systematic mapping. IEEE Transactions on Professional Communication, 65(3): 370–391, September 2022. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Zollet:2014:ICW

[Zol14] Roman Zollet. Interactivity of corporate Websites: an integrative review of the literature. IEEE Transactions on Professional Communication, 57(1): 2–16, March 2014. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Zoetewey:2004:TAM

[ZS04] M. E. Zoetewey and J. Staggers. Teaching the Air Midwest Case: a stakeholder approach to deliberative technical rhetoric. IEEE Transactions on Professional Communication, 47(4):233–243, December 2004. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Zurkowski:1975:PCT

[Zur75] Paul G. Zurkowski. Photocopying and copyright technology confronts law. IEEE Transactions on Professional Communication, PC-18(3):224–228, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Zurkowski:1977:AEE

[Zur77] Paul G. Zurkowski. The approaching economic environment. IEEE Transactions on Professional Communication, PC-20(2):54–55, September 1977. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Zahed:2001:CFI

[ZVS01] F. Zahed, W. V. Van Pelt, and J. Song. A conceptual framework for international Web design. IEEE Transactions on Professional Communication, 44(2):83–103, June 2001. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Zha:2021:UTB

[ZWE21] Shenghua Zha, Shenghua Wu, and Julie M. Estis. Using team-based learning to promote engineering students performance and self-efficacy in a technical writing class. IEEE Transactions on Professional Communication, 64(4): 456–467, December 2021. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Zhang:2020:DDL

[ZWX20] Huiyu Zhang, Yicheng Wu, and Zihan Xie. Diversity or division: Language choices on international organizations official Websites. IEEE Transactions on Professional Communication, 63(2):139–154, June 2020. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Zhou:2005:HAE

[ZZ05] L. Zhou and Dongsong Zhang. A heuristic approach to establishing punctuation convention in instant messaging. IEEE Transactions on Professional Communication, 48(4): 391–400, December 2005. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Zhang:2022:ETB

[ZZL22] Yueyue Zhang, Cheng Zhang, and Miaomiao Liu. Effects of top-down, bottom-up, and horizontal communication on organizational commitment: Evidence from Chinese Internet firms. IEEE Transactions on Professional Communication, 65(3):411–426, September 2022. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).